A Complete Bibliography of Publications in Computers and Mathematics with Applications (2000–2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

(2 + 1) [KHN09, LWZ08, Ma09a, TG03, YZ02b, Yan03a, Zha07b]. (2, 2^{n} − 1, 2^{n} + 2^{n−1}1, 2^{n+1} + 2^{n} − 1) [ARS02]. (3 + 1) [BB08b, Asl11, HO00, TGH02]. (A, η) [CyL08, Jin07, Jin08, LCV06, yL08b, yL09c]. (A_{σ}) [Sav08a]. (α, A) [Lug01]. (α, β) [DSA09, Jun10, WI08]. (α ∈ N) [WS07]. (d, B) [Lin07]. ([digamma], α, ρ, θ) [Zal02]. (ϵ, ϵ ∨ q) [Jun09]. (F, α, π, d) [GA06]. (F ρ, θ) [Zal03]. (H, η) [FHT05]. (h, q) [OS08b]. (k, n − k) [LJL06, ZYW08]. (N + 1) [WYL08a]. (n + 2) [WWYL09]. (n, k, m) [AFI06]. (n, n + 1) [Den08]. (\overline{∈}, \overline{∈} ∨ \overline{q}) [MZD09]. (p, r) [Ant08]. 0 [SKU01]. 1 [Col02, PP08, Ram05, SKU01, YM04b].

$104.95 [Ano00-56]. $105 [Ano00x]. $105.00 [Ano04m]. $110 [Ano00-50]. $110.00 [Ano01d]. $115 [Ano04r]. $119.00 [Ano00q, Ano00r, Ano03-48]. $12.95 [Ano00-87]. $120 [Ano00-28, Ano03m]. $125.00 [Ano00-71]. $130 [Ano03-57]. $132 [Ano00d]. $14 [Ano00p]. $14.95 [Ano00-90, Ano03-47]. $140.00 [Ano01h]. $149.95 [Ano03-74]. $16.95 [Ano03p]. $165 [Ano00w]. $17.50 [Ano00-76, Ano00-77]. $18.75 [Ano00-66]. $18.95 [Ano03-27]. $189.00 [Ano00-48, Ano00-49]. $19.95 [Ano00-75]. $19.99 [Ano00z]. $193.00 [Ano00g, Ano00h]. 1: −3 [HRS08]. 2 [AC02, AM09, BR03, CM00, CM01b, DPU07, HJWL03, He00a, HLC00, HLCW07, ISK00, KKM00, Kru08, Ren03, YM04a, ZCM01a, Zha03b]. $20.00 [Ano00j, Ano00k].

$213.50 [Ano00-85]. $22.50 [Ano01v]. $22.95 [Ano03q]. $24.95 [Ano03-49, Ano00v, Ano00-57, Ano00-82, Ano00-86, Ano00-91, Ano03-50, Ano04-30]. $25 [Ano03-38, Ano03-90]. $25.00 [Ano00l, Ano00-42, Ano00m, Ano03l]. $250.00 [Ano03-43]. $26 [Ano03-76]. $26.00 [Ano00-35]. $27.50 [Ano00-80, Ano00-81]. $27.95 [Ano04l, Ano03-89]. $28 [Ano00p]. $29.95 [Ano00-74, Ano01u, Ano03-68, Ano00-36, Ano00-52, Ano00-53, Ano00-65, Ano00-72, Ano00-73, Ano00-92, Ano00-93, Ano01-27, Ano01-28, Ano03u, Ano03-85, Ano03-64, Ano03-72]. 2m^{th} [CH00c, GH03]. 2n [CZ08]. 3 [AMV08, Bae07, Bri00, BV01b, Deu01, HZZ02, HWHG08, ITR08, Kor00a, Kor07, NSB08, Nik06, OAVB08, WYL08b, WB09, YW03b, ZRS05]. $30 [Ano04e, Ano04g]. $30.00 [Ano00-54]. $315 [Ano00y]. $32.00 [Ano04x]. $34.95 [Ano00-39, Ano00-40, Ano00-37, Ano00-44, Ano00-45, Ano00-60, Ano00-61, Ano00-63, Ano00-62, Ano00-67, Ano00-68, Ano01s, Ano03-83]. $35 [Ano00-83, Ano00-88, Ano00-89, Ano03f, Ano03t, Ano03n, Ano03-79, Ano03-33, Ano03-63, Ano03-62]. $35.00 [Ano00-64, Ano00-30, Ano01m, Ano01t, Ano04-27]. $37.95 [Ano03x]. $38 [Ano03-78]. $39.95 [Ano00-34, Ano03w, Ano00-32, Ano00-33, Ano01y, Ano03-35, Ano04-29]. $40 [Ano03-53, Ano03-66]. $40.00 [Ano01q, Ano01r, Ano03-71]. $42.50 [Ano03-58]. $44 [Ano00-47]. $44.95 [Ano01f, Ano01k, Ano03e, Ano03-69, Ano00-27, Ano03-30]. $45 [Ano03-41, Ano03-87, Ano04h, Ano04-31]. $45.00 [Ano01g, Ano03v, Ano00e, Ano00f, Ano00-29, Ano03z, Ano03-81]. $49.50 [Ano03-49]. $49.95 [Ano00-46, Ano03-34, Ano03-54, Ano04j, Ano04o, Ano04p, Ano00-55, Ano00-70, Ano03i, Ano03y, Ano03-84, Ano03-56]. $50 [Ano00-38, Ano03-60, Ano03-70]. $50.00 [Ano01x]. $52.50 [Ano01p]. $54 [Ano00-58]. $54.95 [Ano01i, Ano03-75, Ano03h, Ano03r, Ano03-40, Ano03-45]. $55 [Ano03-86, Ano00n, Ano00-84, Ano03g, Ano03o, Ano03-28]. $55.00 [Ano01o]. $59.95 [Ano00-78, Ano00-79, Ano03-67, Ano00o]. $60 [Ano00-88, Ano00-89, Ano04d]. $60.00 [Ano00l, Ano00m, Ano03-80, Ano00-59, Ano00-69, Ano01n]. $62.00 [Ano01z, Ano01w]. $64.95 [Ano03-37]. $65 [Ano03s, Ano03-55, Ano03-61]. $65.00 [Ano03-68, Ano01e]. $67.50 [Ano04-32]. $69.95 [Ano03-74, Ano04z, Ano03-31, Ano03-46, Ano04f, Ano04y]. $70 [Ano03-36, Ano04n]. $70.00 [Ano03j, Ano04q]. $72p [Ano00-31]. $74.95 [Ano00-74]. $75 [Ano00-51, Ano03k, Ano03-29]. $79.95 [Ano03-82, Ano04i, Ano04v, Ano03-44, Ano03-51, Ano03-52, Ano03-88, Ano04w]. 8 [KHA07]. $80.00 [Ano00-41]. $80.95 [Ano00i]. $83.95 [Ano04u]. $84.95 [Ano03-42]. $85 [Ano04e, Ano04g, Ano00-43, Ano03-73]. $85.00 [Ano00-80, Ano00-81, Ano00s, Ano00t, Ano01l]. $89 [Ano00u]. $89.95 [Ano03-39, Ano04s]. $9.95 [Ano01j]. $90.00 [Ano03-77]. $95 [Ano04-28]. $95.00 [Ano03-59]. $97.00 [Ano04k]. $98 [Ano03-65]. $98.00 [Ano04t]. ^{∗} [YY07]. ^{[X]} [AHRT09, AM04, GS07b]. ^{B} [GLZ00, Gup02b]. ^{X} [LKH03]. _{2}F_{1} [IR08]. _{H} [AM04]. _{N} [AK06]. A [Ver08b, WC04, WX00b]. A − θ [KK06a]. A_{1}XB_{1} = C_{1} [NOY01]. A_{2}XB_{2} = C_{2} [NOY01]. A_{T,S}^{(2)} [SC07]. AB^{2} [JKY06]. a ≥ 3 [CC08b]. α [HX03, mLpW08, Xu02b, Zha04a]. AX + Y B = C [Liu06a]. AX − EXF = BY [WHD09]. AXB = C [CI06, Liu08e]. AXB = D [YD08]. AY B = E [DH08b]. B [AA05, Ant08, SJL09]. B(r, s, t) [FBA07]. B_{2} [YT00]. B_{n}^{(χ)} [LS06a]. B_{z} [KPSW06]. β [Hav08]. △ [DP08]. C [FBA07, KL02, LYG05, XL01]. C^{(1,1/3)} [JN07]. C^{1} [CL08b, He00a, LW09a, LHC05]. C^{2} [LF09]. C^{k} [Gou02]. C_{0} [FBA07]. χ_{n+1} = (χ_{n}, χ_{n−k}) [SX06]. χ_{n} + 1 = α + βχ_{n}A + Bχ_{n} + Cx_{n−1} [KLMR03]. CY D = F [DH08b]. D [HCLF01, LL05, HL07, JLP09]. Δ_{∞} [JN07]. E [SJL09, FQZ02]. e^{z}E_{1}(z) [Has02]. ϵ [Li07]. f [KC05a, qWjH07]. G [LWY09]. γ [Ant05]. GF(2^{m}) [LLY01, JKY06]. H [Din00b, HL08a, HLWT04, HN07, LCC08, Ver08a, Ver12, Wu00, WTY03, XH07, ZGY05a, ZGY07a, BV09]. H − η [LHH08]. H_{2} [Mah09]. H_{2}H_{∞} [YL04c]. H_{∞} [KLP09, PCE^{+}05, SN03]. H_{W}^{ω} (R^{n}) [PPA08]. hp [DMVR00, PLL03, Ses03, WFS03, ZRS05, BCS03]. I [ADN02, CS08, Ikh01, KT03b, LW06, LW08a, LW08b, Liu09a, SFL02, WF05, hYL09, ZL07b]. J [Ano01h, WS07]. K [CLJ09, KWHC05, Waz05c, DW02b, GSS02, Inc09]. K/(1 + χ^{n}) [NG04]. κ [HTZY03]. L [DP02, Din02, DPJ03, EGM^{+}02, JC05, Wen08]. L^{1} [Xu02a]. L^{2} [Yan05a]. L^{∞} [Bou03, JK01]. L^{ϕ}(μ) [Din04b]. L_{1} [WFLC05]. L_{∞}[a, b] [BCDF02]. l_{N×N}^{1} [PS03b]. LD [PS07b]. LU [MAK02, Pan00]. LUT [CWLC09]. M [CCRMRF00, FTC03, Sot01, WS07, ZHL05, AZCK05, DG05, LKC04, Ma01, Ma04, SSX09, Sun09a, XC06, Xie08, YW08b, ZL08c, ZYW08]. M(x) [Cha08]. M/M^{B}/1 [HR07]. M^{X}/G^{Y} /1/N [SG08]. C^{2} [HRS08]. R_{+}^{1} [DZ08]. N [LY08b, LWY09, PS08a, Qua01, TL00, AM08c, BS01b, Coh00, DLF09, DCLF09, Dra08, FW00, GK04a, GM09c, HYZ02, Jum01, KDY09, KL02, XC06, XC07, Yao04]. n + 2 [YWYL07]. n^{th} [Coh00, Guo01, Guo02, Liu06b]. n × n [YSP^{+}08]. O(n) [LU09a]. P [Net05, QS03, XX09a, AS03a, AM08c, BS03b, BH00, CG08, CD05a, CWZZ07, CT08, EA03b, Jia01c, JX04, JLZ09, KR07b, KLW09, LX03, Liu04c, MDG05, MA09b, PGT08, PDBH03, Sun09a, TS00b, WG06, WG07b, WG08a, WA08, WGZ06, XY04, XG07, YW08a, YWYL07, ZG09b]. p^{k} [Ben00, Cha01b]. P_{⋆}(κ) [CKL07]. Φ [HH03, Hua00, WS04a, AOS07, Bai01, Cab01, HGH02, KT09, LK03]. Φ_{1} [SL06c, SJL08]. π [BDT04, Cho08]. Q [SH02, Gau04, KR07b, MRS08, Rom01]. Q(\sqrt{3}) [Deu01]. QR [SV09]. R [Yav05a, XZ09a]. R^{m} [Ubh04]. R^{n} [LdSF00]. R_{0} [WHS08]. R_{3} [She04a]. R_{L} [ZL07b]. S [Ano04y, AK06, CCRMOG04]. σ [MM08d]. T [CC04a, LZZ04, ND05b, DLF09]. T_{31/2}L [Li04a]. τ [ND01]. θ [Gan06, GL03a, SYL05]. × [LS02]. U [DFFL09]. U(n + 1)q [Wan09c]. V [Zal02]. V_{p} [KHS00]. X [JAM02]. x′(t) + ax(t) + a_{1}x([t − 1]) = 0 [LGY09]. X(t_{0}) = I [Ada04]. X − A^{∗}X^{−n}A = Q [ES01]. x^{1/3} [Bel09]. x^{2} − kxy + y^{2} + x = 0 [MZ04]. X^{Δ} = AX [Ada04]. x_{n+1} = (1 − α)x_{n} + βe^{−γx_{n−k}} [MY05]. x_{n+1} = (a + x_{n}x_{n−k})/(x_{n} + x_{n−k}) [ASÇY08]. x_{n+1} = A + y_{n−m}/x_{n} [ZYEZ07]. x_{n+1} = α + x_{n−1}/x_{n} [Ste08].

\frac{x_{n+1}=A ∏_{i=l}^{k} x_{n−2i−1}}{B+C ∏_{i=l}^{k−1} x_{n−2i}} [HKA08].

x_{n+1} = max{1/x_{n}, A_{n}/x_{n−1}} [FM08b]. \frac{x_{n+1}=x_{n}+px_{n−k}}{x_{n}+q} [DDM08]. y′ = λy [WX00a]. y_{n+1} = A + x_{n−m}/y_{n} [ZYEZ07]. Z_{m} [Coo05]. z ← z^{α} + c [WS07]. |A|_{k} [Sav07a]. ||A^{−1}||_{∞} [Hua08b].

*congruence [ZYCI09].

-1 [XC07]. -accretive [AZCK05, LKC04, CyL08, Jin07, Jin08, LCV06, yL08b]. -adaptive [ZRS05]. -adic [KR07b]. -algebras [JLP09]. -ary [DLF09, DCLF09, KDY09]. -attracting [XX09a]. -averaging [Din04b]. -band [FTC03]. -based [CWLC09, KPSW06]. -binomial [Wan09c]. -calculus [Gau04]. -center [HLC00]. -coercive [MM08d]. -concavity [KL02, SJL08]. -conorms [DLF09]. -conservative [MM08d]. -constant [KHA07]. -convergence [DP08, HCLF01]. -Convex [CCRMOG04, BV09, CCRMRF00, DP02, Din02, DPJ03, JC05, SJL09, Wen08, Zal02, Zal03]. -convexity [SL06c, WS04a]. -cost [WI08]. -D [AC02, CM00, Col02, HJWL03, HZZ02, Kor00a, KKM00, Nik06, Ram05, WYL08b, YM04a, YM04b, YW03b, ZCM01a]. -decomposition [MAK02]. -DGFEM [WFS03]. -diagonal [HX03]. -difference [Rom01]. -dimensional [Asl11, BB08b, HO00, HL07, KHN09, LWZ08, Ma09a, TGH02, TG03, YZ02b, Yan03a, Zha07b, Qua01, WYL08a]. -disfocality [LYG05]. -divergence [KC05a]. -dominating [HLCW07]. -double [Sav08a]. -error [Bou03]. -Euler [KR07b]. -extension [OS08b]. -fields [PS07b]. -fuzzy [DSA09, Jun10, EGM^{+}02, Jun09, MZD09]. -graphs [Den08]. -groups [BH00]. -harmonic [WC04]. -hemicontractive [Bai01, HH03, Hua00]. -ideals [JLZ09]. -integrals [KR07b, MRS08]. -invariant [XX09a]. -invexity [Ant05, Ant08]. -irreducibility [XY04]. -Laplacian [AOS07, Cab01, CG08, CWZZ07, Jia01c, JX04, KLW09, LX03, Liu04c, MDG05, PGT08, Sun09a, WG07b, WG08a, WA08, XG07, YW08a, YWYL07, ZG09b]. -Laplacian-Like [WG06]. -Manifolds [Lug01]. -matchings [YW08b]. -matrices [HL08a, HLWT04, LCC08, ZHL05]. -matrix [Sot01]. -maximal [Ver12, Ver08b, Ver08a]. -means [DW02b, CLJ09, KWHC05]. -methods [Gan06, GL03a, SYL05]. -Moment [WGZ06]. -monotone [CC04a, FHT05, HN07, LZZ04, LHH08, XH07, ZGY05a, ZGY07a]. -mortar [BCS03]. -norm [CT08]. -normal [Xu02b]. -norms [DLF09]. -optimal [ADN02]. -Order [Liu06b, CH00c, GH03, Guo01, Guo02, WWYL09]. -Partitioning [HTZY03]. -person [KL02]. -plane [Hav08]. -point [DG05, Ma01, Ma04, SSX09, Sun09a, ZL08c, ZYW08]. -policy [LWY09, PS08a]. -predators [XC06]. -preinvexity [Ant05]. -preys [XC06]. -projection [qWjH07]. -queue [LWY09]. -regularity [JN07]. -regularized [XL01]. -resolvent [yL09c]. -sequences [YT00]. -smooth [AA05]. -space [Din00b]. -spaces [Wu00, WTY03]. -species [FW00, LY08b, TL00]. -stability [Zha04a]. -stable [Net05, WX00b]. -strictly [Inc09]. -strongly [HGH02, LK03, HH03]. -th [Dra08]. -topological [Li04a]. -Torus [BS01b]. -type [WHS08]. -valent [AM08c, MA09b]. -value [ND01]. -variational [Li07]. -version [AS03a, BS03b, CD05a, EA03b, PDBH03]. -vex [SJL09].

/1 [AM04, GLZ00, Gup02b]. /G [AM04]. /G/1 [AHRT09, LKH03]. /Geo/m/N [GS07b].

058 [Ano00-52, Ano00-53]. 059 [Ano00-72, Ano00-73]. 068 [Ano00-92, Ano00-93].

1 [AHRT09, AM04, GLZ00, Gup02b, Gup02a, KA00b, KA02, LKH03, LL06b, LWY09, LMSH04, Wan04a, WZ07b, WX09a]. 114 [Ano00w]. 137.95 [Ano03-86]. 14-th [Ano00-85]. 15 [Ano03-57]. 17.00 [Ano00-42]. 18.50 [Ano04l]. 19.95 [Ano01u]. 191.04 [Ano00-85]. 1947-2005 [GS05]. 196 [Ano00y]. 1999 [Ano00-85].

21-st [Ano00v]. 255.00 [Ano00-71]. 269.00 [Ano00-48, Ano00-49]. 27.00 [Ano01g]. 27.10 [Ano03w]. 28.00 [Ano01q, Ano03-69, Ano04o]. 2nd [Ano01w].

30 [Ano00-47]. 34.00 [Ano01i]. 35.00 [Ano01w]. 35.50 [Ano03-54]. 38.50 [Ano01o, Ano04j]. 39.50 [Ano01z].

4 [Ano00-74]. 42.00 [Ano03-80, Ano04p]. 44 [Liu09a]. 45 [Yan08]. 45.50 [Ano03-37]. 46 [Liu08b]. 49 [ID10]. 49.00 [Ano04i]. 4d [HHL^{+}06]. 4d-VAR [HHL^{+}06].

5 [Ano03-38]. 50 [ID08]. 52 [Ata09]. 52.00 [Ano03-82]. 52.50 [Ano04z]. 53 [KÖ08b, Zho08]. 54 [Ari09c, EM09b, ES08b, KÖ08a, NLL08, SW08]. 55 [Ver12]. 56 [Ano00u, Asl11, fDxZ11]. 58 [AR10, JY11]. 59.50 [Ano04v].

60.00 [Ano03-48]. 62.00 [Ano04k]. 63.50 [Ano04-28, Ano04u]. 68.25 [Ano00-56]. 69.00 [Ano00q, Ano00r].

7 [Ano00-86, Ano03-36]. 70 [Ano00-52, Ano00-53, Ano00-72, Ano00-73, Ano00-92, Ano00-93]. 70-058 [Ano00-52, Ano00-53]. 70-059 [Ano00-72, Ano00-73]. 70-068 [Ano00-92, Ano00-93]. 73 [Ano00x]. 75.00 [Ano01d]. 76 [Ano00-50].

83 [Ano00-28].

91.25 [Ano00d]. 92-05-29 [Guy00]. 92-05-31 [Guy00]. 925 [EM09b]. 98.00 [Ano01h].

= [LL05].

a-priori} [WWS^{+}06]. AAAI [Ano01r]. abatement [CS08]. Abbass [Ano03k]. Abbott [Ano00-54]. Abdelhamid [Ano03-81]. Abel [HM03a, HHL08, MCH01, MH02, MO01, PSS09]. Abel-type [MCH01, MO01]. Abelson [Ano00j, Ano00k]. Absence [PP08]. Absolute [XLL04, Zha04b, Che03, Sav07b, Sav08b, SL09]. absolutely [BCDF02, CCD07, ZC06]. absorptions [Che04e]. Abstract [Xie01, ZLS06, BPA03, BSK04, HM05b, HPG06, LX00, LQ04, MAH08, Par04a, WW03, WFC04a]. abundant [SSH09, Yan03a]. Academic [Ano00d, Ano00u, Ano00x, Ano00y, Ano00w, Ano00-28, Ano00-50, Ano00-56, Ano00-71, Ano01d, Ano01h, Ano04k]. Academic/Plenum [Ano00-71]. accelerate [HBP06]. Accelerated [Pao03, Wan00e, PR06]. accelerating [FAF09, RTS07]. acceleration [BNP05, PW00]. accelerators [VA04]. accessible [HYZ06]. accounts [SY05]. Accretive [CPC00, AZCK05, CKN04, CyL08, HH03, HCLJ00, HGH02, HF04, IP07, Jin07, Jin08, LCV06, yL08b, LK03, LKC04, MN01, ZCK00]. accretiveness [Dog01]. accumulation [MT09]. Accuracy [CSX01, Sch00, Yan03d, AY06, HSS01, JYDH02, KKS^{+}04a, Ryo03, SALK02, YX09b]. Accurate [GK04a, Gra08, LTWM08, De 02, Jog04, LS04a, WK05]. ACL2 [Ano01h]. Acoustic [IS06, SLL03, WDLP04, WDLP08, Ano03f]. acoustics [CFR03, LDK^{+}06]. Acquisition [HCTC03, PCA09]. action [CS00a, SK06a, Ano00-46]. activations [NC09, WL08c]. active [JZQZ07, bJQxC09, LYF09, Mar09, MRMRM02, SWS05, SHWF09]. active-set [JZQZ07]. Activities [Ano00-35]. activity [Kaw06, MS09]. actuator [TLS^{+}08]. actuators [WGF09]. acyclic [MPZ08]. Aczél [DBGSP08]. Aczél’s [Wu08c]. Adam [Ano03-60]. Adams [LT09a, Sha02]. adaptation [ZL09c]. adaptations [Pan05b]. Adaptive [CJL04, DH07, Lei02, MKK08, PL01a, PKP09, SL08, WLL06, ZYG^{+}09, ABH08, BI02, CY00, ES08b, HL02b, HLW07, Han08, HLP03, LML02, LW02, LNL08, LPvST07, MBI08, NLL07, NLL08, NST01, PLP09, PLL03, SSL07, SC00a, TAT07, TLP^{+}09, WWS^{+}06, YHS08, ZRS05, ZZ03a]. adaptivity [OP01, PDP05]. addition [IK05, Ung07]. additional [RS06]. Additive [FGH08, PIM^{+}08, IW06]. ADI [Deh02, LCL06]. adic [KR07b]. Adjoint [AN05, Ale06, GC06, PPP^{+}06, BN05, CLC01, DW08a, HHR07, LYG05, Shi03, VJ04, WXM09]. Adjoint-Based [Ale06]. adjustable [JCC^{+}09]. adjusting [HLP05]. administration [Ano01s]. admission [WLT07]. Adomian [GS06, JYDH02, LHC08, Les02, LW09c, ÖY08, SGM09]. adsorbents [VM09]. Advanced [Ano00-88, BUP09b, BBP09, DN03, FHZ08, LYL07, PS03b, SYL05, Yan01, Ano00d, Ano01d, Ano03g]. advanced-delay [BBP09]. Advances [YZWL09, Ano00e, Ano00f, Ano01e]. Advancing [BW06a]. Advection [STK08, BI02, BMM^{+}06, Che04f, Col02, CGPB08, Les05, PKSP08, PDRR08, Roo08]. Advection-diffusion [STK08, Les05, PDRR08]. advection-dispersion [Roo08]. advection-dominated [Che04f]. advertising [Bel01]. aerodynamic [DS02]. aesthetic [LLS09b]. Affine [PO04, FF01, GGCD05, XZ08a]. affinity [Ano03-40]. Africa [Ano00-32, Ano00-33]. after [Iac07, JYDH02, Ano01x]. after-treatment [JYDH02]. Agatha [Ano01x]. age [LGZ01, Li09b, Ano03-28]. age-structured [LGZ01, Li09b]. agent [Ano01n, KR06]. agents [Ano01t, Ano03w]. aggregation [Che04g, PIM^{+}08]. aggregations [MM08c]. aging [Kan07]. agreement [CYH05, LLS^{+}09a, YSH03]. AI [Ano00l]. aid [YCCY02]. aided [Ano03-69, BP09, Cip08, Ano01h]. Air [MS08, Gre09, WK08, Zla01]. aircraft [LNL08]. Akkermans [Ano00-38]. Al [Ano00m]. Alan [Ano00-71, Ano03-38, Ano03-55]. album [Ano03-88]. Aldo [Ano03-57]. Alexander [Ano03-82, Ano03-86]. Alexei [Ano03-70]. Alexey [Ano04s]. Alfred [Ano00-84]. Alfredo [Ano03-86]. Algebra [Che06, ABBD01, Ano00-58, Ano01w, BENT06, CKY07, GT00, LRLM^{+}01, NV08, Roj08, WY02, Ano00-31, Ano03e]. algebra-based [WY02]. Algebraic [MG05, BD01, Bai00, BPSW00, BCT00, DW08a, EW00b, GM00, HM04, Lin08c, MP02a, Mär05c, MRÁR08, Sha09, WZ08a, YXZ09, ZW09b]. algebras [Dha04, Jun08, Jun09, JLZ09, JLP09, MZD09, YWC08, ZX08]. Algorithm [AS07, BFL06, CWY02, LC06a, LY06b, Zha06, ZLS06, AC04, AM09, ABH08, AK08c, BHEKE07, Bog09, BCT00, BK07, CG05, CJ08, CM01a, CLJ09, CGY08a, Cha06, CWLC09, CML^{+}09, CZ09b, CCL02, CFY08, CKL07, CSLC07, CCP09, DH08b, DF06, DY05, DYZ08, DZ08, DW02b, EBVS06, Gal02a, GL05, Hau06, HLP05, HCG08, HLC00, HL02c, HLM04, Hua05, HY06, HN07, HLCW07, HL08c, IK05, JZ02, JCCT09, JT05, bJQxC09, Jin08, Kar01, KK06d, KB04, KA07, KKS04b, KGWL08, KD08, KH09, LPRH05, Lee01a, LTT02, LGGSP06, LWLJ09, LCL^{+}09, LW09c, LLJ09a, LCY05, Lin04a, Lin07, LU09a, LdSF00, LLL02, MRT05, MCRÁ05, MPZ08, Mur05a, NVAR03, OTB^{+}08, OS09, Pan00, PP01a, PMQR09, PSS07, PYLX08, PG06, PG08, RH09, RKS08, RZH03, Sha09, Sny00, SP09, SMS^{+}04, SB00, TZ08, TP03, Ubh04, ÜL00, Ver08b, Ver08a, Ver12]. algorithm [WWWL08, WZ08a, WHS08, WYL^{+}09, WNS09, Waz01, WFTL09, XH02, XXC^{+}09, XXLD04, XC04b, YY07, YZHY09, YC07a, YZWY08, ZGY05b, Zha01d, Zha02a, ZW03a, ZYY08, ZYY09, ZXTL09, ZLN06, ZZ06]. algorithm-based [LWLJ09]. Algorithmic [KCD05, Ano00-31, Ano03-45, Jin09, Mou02]. Algorithms [Ano00s, Ano00t, Ano00-70, Lan08, LX06, ABBDP04, Ano03-42, Bar01d, BY03, BD04c, BGS04, CLM03, CLY08, CSdS08, CHL09, CSZH09, CWZ06, DLL08, DAKO06, FHZ08, GPMPR03, HHJS03, Her01, HR05, IK04, JS08, Jin07, KPD^{+}09, Kur08, LCHL06, LZZ04, LRL09, LHW04, MMM04, MUHdH03, Noo00b, PSS09, PLD02, PGH^{+}08, PZ07, PRA02, RTS07, SKU01, SHWF05, TH02, Ver01a, Ver01b, WD08, XZDD09, XD02, XW05, Yan03f, ZGY07b, ZAY08, ZZZ02, Ano01g, Ano03p, Ano03-68, Ano03i, Ano03h]. Alienor [ZCM01b]. aligned [CO00, HW05c]. alignment [CSZH09]. alignments [RMC08]. Allee [JD07]. allelopathy [ZM02]. Allen [Ano00n, Ano03-78]. Alligator [Ano00-62]. allocation [EMP^{+}06, JWC08, LLS^{+}09c, QW09, WI08]. allowable [CHW07]. allowing [LGGSP06]. Almost [AB08a, AOW05, FW04a, LKC04, Sav07b, Son09, ZS08a]. along [Yav05d]. altering [Bal00]. Alternate [WW08a, SMO03]. alternately [LYL07]. Alternating [BNP05, Che04c, GK04b, HLW07, LY09a, WYH00, ZW09a, ZH09]. alternating-direction [LY09a]. alternation [FM05]. alternative [SH02]. alternatives [CL00, She05]. Alvin [Ano03-53]. always [MG08a]. Alzer [XX02]. Amanda [Ano03-86]. ambiguity [AS05]. amended [Pan05a]. America [Ano00-32, Ano00-33, Ano00-35, Ano00-66, Ano03-35, Ano03-58, Ano01-28]. American [CKKK06, Hon02]. among [DY09, LWLJ09, MM08c, OKCO08, ZY07]. amphiphilic [ITM08]. ample [YLZ08b]. amplitude [Blo08, Fan09, Hai03, ÖY07b, PB01, Zha09a, Zha09g]. amplitude-frequency [Zha09a]. Amsterdam [Ano00g, Ano00h, Ano00-85]. analog [TLS^{+}08, Ano04d]. analogies [NST01]. analogue [JSS01, RGTT03, ÜÇT08]. analyses [Che04a, He00a, RPO03]. Analysis [ACV03, AS03a, ACLL06, Bel06, CJ08, CCK03, CCKH07, DA06, EAE09, FP06, GC06, Gil00, GC02, Iac07, IMLF09, Jon07, KS00b, Lee04, LC06b, Ras06, SPW08, SW06, ZSPZ08, Abu01, AB05, Ano00-94, Ano03-73, Ano04i, Ano04m, Ano04p, Asi08, BSS98, Bal00, BJ02, BNH08, BP03, BX02, BMM^{+}06, CJR09, CM09, CCS08, CPP08, Che00a, Che04a, CC08a, CW08, CCF06, CFH05, DPU07, EA03b, EW00b, FT01, FS04, FM08c, GT00, Gla07a, GS07b, GMMB06, GMMB09, GYPL08, GHC09, Has02, Hav08, HC02, HHC08, IKY06, IDOK06, Jah04, JS04, KCD05, KOV08, Koi07, KZD03, KWHC05, LTYB09, LMA08, LGBS00, Li00a, LZZ04, tLW08, Li08a, LYJ09, Lin08a, LCC08, LB04, LLL02, MQ05, MH09, MS03, Min08, Mos08, Mou02, NS00, NTTS09, NKCS08, NC09, Noo02a, PSM02, PC08, PL05b]. analysis [PB02, RMC08, SANH08, SDK01, Ser00, SS09b, SSL07, SBSY06, SK06b, TS00a, TGIM02, TL06, TV07, TYW05, WK04, WZ06, WKZ08, WYL08b, Wei00, Win03, WX03c, Wu03, Xia07, YTCF08, Zha03a, Zha04b, ZYG^{+}09, ZZZ09, ZC08, ZZD08, Ano03-46, Ano03-52, Ano03-62, Ano04e]. analysis-based [Lin08a]. Analytic [CSPJ05, JPV01, RRM08, SW02a, SW02b, AIS08, AD09, BR05b, CCJ08, CK06, CDJ04, JI04, Jan03, LTWM08, MA09b, NCK^{+}08, PS03a, PMS07, QC06, SL07a, Waz07b, XA09]. Analytic-numerical [CSPJ05, RRM08]. Analytical [Das09, STH08, WG08b, Ari09b, BNH08, DS02, De 04, GGE09, HL09a, Lu09b, MH09, Wu09]. analyzing [XZ09b]. anarchy [HLY08]. Anatoli [Ano04w, Ano04v]. Anco [Ano04y, Ano04z]. Anders [Ano00-28]. Anderson [Ano03-54]. André [Ano04j]. Andrea [Ano04-29]. Andreas [Ano04p]. Andy [Ano00-63]. Aneja [MCC09]. aneurysms [BW09]. angle [LGGSP06, WB01, Köh00]. Angular [CGL01, AA08]. animal [Ano03-78]. animals [Ano03-36]. Anisotropic [PPP^{+}06, Bla07, CLMM06, EP04, Wan02c, ZS07]. anisotropy [Sar03]. Anjewierden [Ano00-38]. Anjo [Ano00-38]. Anne [Ano03-29]. annealing [ZYY08]. annuli [CYX09]. anonymous [Jua03]. answering [Ano04q]. ant [KWHC05, KS07, MWZ09]. Anthony [Ano00-58, Ano00-89, Ano03-32]. anthropometrical [SYHC06]. Anti [Gon09, Ou08, BCGA08, ES07]. anti-involutions [ES07]. Anti-periodic [Gon09, Ou08, BCGA08]. antigen [LMSH04]. Antiperiodic [LSN05, Jan04b]. antisymmetric [tLjW08]. Anupam [Ano00-50]. anurans [Ano03f]. any [Dra03b, TT09, lYsT09]. aorta [MP07b]. Apostol [LS06b, WJW08, ZY08a]. appearance [LLT09]. Appell [BCR04]. Appl [AR10, Asl11, Ata09, fDxZ11, ES08b, ID10, ID08, JY11, KÖ08a, KÖ08b, Liu08b, Liu09a, NLL08, SW08, Ver12, Yan08, Zho08]. applicants [KT03a]. Application [Asl11, BB09a, BBA09, BB08b, ES08b, GK02, GNM07, GAT07, GL09c, He09a, KWHC05, KH09, LWLJ09, NLL07, NLL08, ÖY07a, PdlF09, PS04, RD07, SGM09, Sam07, TL06, YLZG01, Yil09, YB07a, Zha06, Zha09a, ZWC^{+}09, dM06, Ano00-37, Ano03-82, Ano03-75, BGMB02, BGA08, BI01, BGS04, Can03, DTXW05, DE06a, ESNP06, FjH07, GSNDB^{+}08, GYPL09, HHS07, Hua01, ISAAM09, IDOK06, JSKE05, Kay04, KDC08, KC03b, LCWL06a, LCWL06b, LSZW07, MRB02, MAR05b, MVS^{+}07, Öze08, PPRG02, PP09, PKT07, PV01, PDBH03, RZH03, SC07, SL00b, SZ09a, Ung01, WK08, WD07a, WWHL09, XZ08b, XMW09, YL09b, YK00, Zha07b, ZL05, Zla01]. Applications [Ana06, Ari09c, BRC00, Che06, DLF09, EM09b, KH00, LCGX06, SM00, Ana04d, Ano03-48, Ano04m, Ano04w, Ano04v, Arg00a, Arg00b, AVA01, Bab09, BL01, BL08, BS01a, BCDF02, BBCD02, BBM08, Cap00, CD09, CTYW05, CLT09, De 04, Dha06, Dha07, DBEGB08, Din02, DSS00, Dra03a, Dra08, EMP^{+}06, EM03, FM06, Gla05b, He03, HW07, HHY09, Her01, HJL05, HM01, IL04, JSS01, JC05, Kir08, Lan04a, Lan04b, LCJL03, LCYL00, LC03, LWZ04, LTSW02, LLNS05, LW09e, LWG04, LCWL06c, Liu07b, LL08b, Liu08d, Mas07, NOY01, Nie09, OEV09, PG04, RV05, RSR00, Roj08, RFCCRM03, Sid08, THD08, Uje03, Uje04a, Uje07, VV08, Ver08a, Ver09, Ver12, Wan00b, WLS06, WLW09, Wan09a, Wen08, Wu00, WD07b, qWjH07, XG07, ZGY05a, ZYG08, ZC09b, Ano00y, Ano00-50, Ano00-71, Ano00-70, Ano01k, Ano03-59]. Applied [Ano03-46, IS06, Ano00-59, BT03, BDGR01, CCZK09, El 03, GW03a, JLP09, Mar00, MIPT08, MRdCV00, RW03, WC04, XWD04, ZX09, Ano00i, Ano03j]. Applying [BCD08, BHS^{+}08, CH05a, FL05, HCG08, Lin08b]. approach [AH09, AO01c, Ano01x, Ano03v, BGMB02, Bel01, BHS^{+}08, BS08, BPW00, BHCØ07, BV01b, CFM02, COS09, CH05a, CL01, CKKK06, CCP09, CM08, DCW08, Ehr02, FP02, FM08c, GM00, GZZ09, GLTS08, HSG^{+}06, HHS06, HCWC07, IN07, Isk04, JKY06, KLHC08, Kha01, LP08a, yL08a, LWY02, LWY04, LLNS05, Liu09b, Lu09b, LG09b, LHWW08, Mah09, MM08a, Mak09, MH09, MRP08, yNyLjZaZ08, OY06, Pal01, PP01a, PKM08, Par04a, Par04b, PP01b, PM05a, QW09, Ren03, Ryo07, SS09a, SW09, SM09, SWL^{+}04, Shi05, Sho09b, SB03, SMO03, TS00a, TLP^{+}09, WS00, WFLC05, WLS06, WSO08, Wan09d, WY02, XZ09b, Yan06b, YFT04, YGH09, Zha07a, ZMLG09, ZLN06, dSM02, Ano03-34, Ano04-27]. approaches [GPT03, Ind02, KK06a, Lin08b, MS08, YAVB08, Ano03-41]. appropriateness [LRLM^{+}01]. approximability [Ano00o]. approximable [BA08]. approximants [NMD09, ZP02]. approximatable [HYZ06]. Approximate [BNH08, CLY08, CHZ09, ETAJ09, FM06, Gu07, HHL08, KFMH06, Ma09a, Waz00, WW08c, Wu09, BAK09, BB01, CGY08a, CL01, DW02a, KC02, Kar03, LZ09, LTWM08, SA03, SV09, Tan03b]. Approximating [All02, DDF02, Dra08, udK07, HS00, IP07, CSPJ05, HH02b, Sya01a]. Approximation [AG09, Bao03, CKQ08, Dog05, DE06b, FLR05, Gal08a, GJQ07, yL09c, Li00b, LCWL06a, Mus06, Sar06a, YYZ07, YI06, ZC06, AC00, AG01b, Ana04a, Ana04b, Ana04c, Ana04e, AD08, Ano00o, Ano03h, Bal04, BB04, Bou00, BH01, Buh06, COEP00, CCP02, Can03, CD05a, Cav00, CGY08b, CGY09a, Che01b, CC09b, CZ00, fDxZ11, Dog01, DFSS08, ES05, GA07, Gan00, Gou02, GQL05, Gup03, HH04a, Her01, HM00, HZ07, JI04, Jan04c, KP01, KFMH06, KK08a, Kil04, KP06, KL03, LG08, LC06b, LCGX06, Lia05, LBW05, LT09b, Mur08a, Nie01b, NRL00, ODO03, PW00, PR07, QLD08a, Roo08, Sar06b, SKW02, STH08, Tel08, TCH07, Ubh04, Ver01a, WWS^{+}06, XX08, YC07a, YD08, Zen00, ZCK00, Ano00-59]. approximation-solvability [Ver01a]. Approximations [AA08, DDCB04, LT05, TDY08, ZQ00, Alb02, AG01c, CBJ01, FDWC02, KL06, Kot04, Kov05, Li05, OLS06, SM02]. approximationsto [Che04d]. approximators [HB04]. aquifers [CCK08]. Arbib [Ano03-86]. arbitrary [ED09, Kov00, LCHL06, Lin08b, RV01]. arc [KC05b]. Arcady [Ano03-36]. Architects [Ano00j, Ano00k]. Architecture [GPT03, Ano01i]. Architecture-based [GPT03]. architectures [Ano01y]. Arcs [HH00]. Area [WL09a, ASM09, CT02b, DZQ06, WL08b]. areas [Che01a]. argument [AB08b, ÇT05, LS03, SL07a, Yan01, ZT07]. argumental [NON^{+}08]. Arguments [WY06, Can08, CKZ01, DN03, Don09, Jan07, Jan09, JA03, Kop01, Liu08a, LYL07, SYL05, WZG00, Wan04b, Won08, ZD03]. Argyris [Ano00g, Ano00h]. Ariel [EM09b]. arising [AA08, ADDL04, BLS05, HCL07, Iva07, Lin08c, Mur05a, dZmClQ09]. Arithmetic [TS00b, Ano00s, Ano00t, Gra08, XZ06]. arithmetical [IW06]. ARMAX [WD08]. Arnold [Ano00-86, Ano03-80, Ano03-72]. Arnoldi [HBP06, Jia00a, JZ02]. Arokia [Ano00-48, Ano00-49]. arrangement [CS00b, KM03b]. array [HW05c, LLY01, MSN^{+}00, SMS^{+}04]. arrays [Åhl02, Ber00, Mär05c, Son00]. Arrhenius [dR03]. arrival [SG08, WCL02]. arrivals [GSS04, KK06c]. Arrow [PKM^{+}06]. Arrow-Head [PKM^{+}06]. art [Ano03-76, Ano03u, Ano04-31, LLS09b, WYL^{+}09]. arteries [Mur05a]. artery [ICM09, MG01, SS09e]. article [ZSKZ03]. Artificial [WZ08c, GOMA07, HH02a, HW05a, KA09a, Ano03k]. ary [DLF09, DCLF09, KDY09]. aspect [Mou02]. aspects [GOMA07, LF05, OSS07, SV08, SKG01, Ano00-48, Ano00-49]. asphalts [PM08]. assemblies [AK08a]. assessment [IY06, KPD^{+}09, Tsa03]. asset [LPvST07, MKK08]. assets [EBVS06]. assignment [Bog09, Kar01]. Assimilation [GCN06, Dim08, TZ08]. assisted [Cse06]. associated [Aou08, ASS08, AY09, AD09, BENT06, CBGVPP09, DGVP06, FZ07b, GSS02, GYS05, GM09a, HS07, KR07b, PS09, SM06, Sri03, WJS09, YD08]. Association [Ano00-32, Ano00-33, Ano00-35, Ano00-66, Ano03-35, Ano03-58, KS07, DY09, LCY05, ZY07]. associative [GSNDB^{+}08]. assuming [HLM04, Hua05, HLCW07, LHYM09, WZ07a]. assumption [CWY08]. assumptions [Dog01, YLZ08a]. astro [AF09]. astro-inertial [AF09]. astrophysical [Ano03-82]. Asymmetric [LHL07, LM03, AH08, HLY08, WYH00]. asymmetrical [Ano03-77]. asymptotes [Fro08]. Asymptotic [ATY08, APvZ01, Asi08, BD03, CP01a, CKR03, CK04, CP01b, Dib03, Dix05, HR07, HLY05, LC06c, Med01, Pen09, SS09b, SR03, SWL05, Ste07, Tan02a, TMG03, Tia08, Wan06c, Wan06b, WKZ08, Wan09e, Xu01, XLT09a, ZYH01, ASÇY08, AP03, CP02, CS02c, CD05c, DK04, Dua09, EIR03, GZ08, Gu08, Guo03, JT07a, Kra01, LYW04, LP08c, LZ05, LLT08, LB01, LG04a, LST02, MC02, Nab04, Pan07b, PS01, Pit03, Sun06a, Sun06b, TL00, TZ04b, Tur07, VR02, WZZ09, XLL03, YTC08, YH09, Zha01a, ZY05]. asymptotical [HLX07]. Asymptotically [yL06, CWY08, CZG04, Din04a, DS05b, udK07, Gu07, KH08, OS09, Tia05, TCH07, Wan07a]. Asymptotics [BCGA05, BC02, BC03, CC02, WWX04, YLZ08a, ZT08]. asynchronous [Bai00]. ATM [HCY00, NQHN03]. atomic [HLY08]. atomicity [HLM04, LHYM09]. attachment [DMP03]. attachment-line [DMP03]. attack [CGFS09]. attracting [Xu03a, XX09a, ZA03b, Zhu08b]. attraction [Gre01, HLZ09]. attractive [HZZ02]. Attractivity [JK09, MY05, CZ03b, CH05b, CWL07, EM03, GW03b, Huo05, LW05, LYW02, MHS01, Mur07b, SA02, SC00b, Wen02, XC06, ZS08a]. attractor [Pol09]. Attractors [DWL07, ZYZ08]. attribute [SYJJ09, XMW09]. attributes [Lin08b]. Atul [Ano03-53]. Auchmuty [Gal01]. auction [Che08b, Lon03]. augmented [LTTH09]. Ausiello [Ano00o]. authenticated [CYH05, YSH03]. authentication [LLS^{+}09a]. authenticator [CKY07]. Author [Ano00a, Ano00b, Ano01a, Ano01b, Ano02a, Ano02b, Ano03a, Ano06a, Ano06l]. auto [BS01b, ODO03]. auto-sleep [ODO03]. auto/endomorphisms [BS01b]. automata [Ano03n, CZ00, CSC^{+}09, OY06, WWHL09]. Automatic [ZCM01a, Ano00e, Ano00f, NQHN03]. autonomous [AF09, DH08c, tLW08, NTHP09, PP09, WZZ09, Yav05c]. autoregression [APvZ01]. autoregressive [EZ08, NWL08]. auxiliary [BK00, LV03, MYG09, SP08a, Zha07b]. availability [LHS08, TY03, XW05]. available [Ata09, Kum06, TH02]. average [CH05a, Gu08, GPL08, She00, She02]. averaged [Alb02]. averages [Man01, TÇ04]. averaging [CC05, Din04b, HCWC07, OS09]. Avi [Ano00i]. Avinash [Ano04u]. Avishai [Ano00-85]. avoidance [YF04, Yav05c]. Avresky [Ano00w]. aware [BK07]. AXB [LL05]. axiomatic [SW09]. axion [HY00]. axis [Yav05d, YG08]. axis-switching [YG08]. Axisymmetric [Ari07a, ASHA08, JZC06]. Ayala [CWL07, FW00]. Ayers [Ano03-55].

B [Ano00-82, Ano03-35, Ano03-33, Ano03-40, Ano03-60, FMFV08, HHK06, VNFF08, CC09a, CWB01, KA09a]. B-spline [CC09a, CWB01, KA09a]. Bachmann [Ano03e]. backhauls [WW09c]. backlogging [LCT08]. backtracking [YL08c]. backup [KS03a]. backward [HZ05, Jun01]. backwards [MRdCV00]. bacteria [ZSPZ08]. bacteria-immunity [ZSPZ08]. Bahcall [Ano03-70]. balance [BGMB02, BGAM09, CCZK09, Ehr02, EG02, ÖY07b, VSKF08, Zha09b]. balanced [Gos00, Whi00]. balancing [PLL03]. balking [ASESESA09, AP09]. ball [Hua04, Ung01]. Balla [GS05]. Baltes [Ano00-48, Ano00-49]. BAM [tLW08, Li08a, MG08c, WL08c, ZLLL08]. Banach [HPT05, SW08, ATT03a, AZCK05, AAI05, Arg00b, AP01, Bai01, BSD00, BA01, BPM05, BGG04, BP03, BHN01, CGY09b, CKY04, CPC00, CWY02, CL08a, CHZ09, CKQ08, CyL08, Dha04, DY05, DYZ08, FW08, udK07, GJM03, Guo01, Guo02, GJS01, fHCH09, HCLJ00, Hua00, Hua01, Jeo04, Jin07, Jin08, KK01, LT05, LCV06, yL08b, LKS08a, LS08, LK03, LWG04, LHH08, OS09, PPRG02, Pan07c, PXW09, PW06, SC04, TCH07, WLW07, qWjH07, XH07, XWL06, YL07, ZYG08, ZC09b, ZC09a, dSOGRMRL04, dSRGRMRL08]. band [FTC03, NV08]. banded [RH09]. bandlimited [CWW08, HS00]. Barbara [Ano00-90]. Bardos [Phu02]. Bargaining [ND01]. Barone [Ano03-44]. barriers [You04]. bars [Has03]. Bartels [Ano03-46]. Basant [Ano03-72]. base [KS00a]. Based [Ale06, BFL06, CW06, DA06, AB08c, ARS02, ARS03, Ano05f, AFI06, BM05b, CJL04, CM01a, COS09, CD05a, Cap07, CWLC09, CL01, CGGC02, CC05, CC08a, CML^{+}09, CSZH09, Che04g, CCHK08, CCP09, DD08, EMP^{+}06, ETH05, GMG09, GPT03, GYPL08, GYPL09, GLTS08, HLSL03, HCWC07, IK04, IK05, Jah04, JCCT09, Jun01, KK06a, KT09, KC03b, KX05, KPSW06, LTYB09, LRLM^{+}01, LF03, LS04a, LW04, LCL00, LYL00b, LMHW05, LCL08, LJ08a, LWLJ09, LCL^{+}09, LW09c, LSL09, LLJ09a, LWZ09, Lin08a, LCWL06b, LCWL06c, LSZW07, LW08a, LW08b, LLT09, Luo06, MWZ09, MJ02, MPZ08, MRdCV00, NLLL09, OLAS01, OF08, PLP09, PT07, SKW02, Sch01, SW09, SHWF05, SDZ06, SP09, SZ09b, TKC02, TLS^{+}08, Ver08b, Ver08a, Ver12, WWWL08, WDLP08, WL09a, WY09a, WNS09, WY02, WFTL09, WFT^{+}09, WSZ02, Wu04, WLT07, WS08b, XLL08, XZDD09, XDD09, YW03a, YFT04, YTEC02, YL04b]. based [YLZ08b, ZMLG09, ZZ09b, ZL09c, ZLD08, ZMM07, ZZ02b, ZZ06, ZL07c, Ano03w]. Basel [Ano03w]. bases [ID10, JW05, LBW05, MP07a, MEG08, PPRG02, Wan00c]. basic [Ano01-27, BBGG08, BGLT08]. Basis [DW08b, GX04, Sar06a, All02, AM08b, BSR02, BGG04, Bax02, Bax06, BI02, Boy01, CLMM06, CCF06, DS07, DR02, DF02, DH07, Dua08, Fly06, FDWC02, FWL04, FLW06, GK02, HMV04, Hon02, HHRA02, HLC05, KH00, LF03, LF05, LGGSP06, Lin00, MDPD02, PD04, PD06, PB02, SN02, SV06, Sar06b, SDZ06, TL06, VTR04, WHL02, WW08c, YTEC02, Ano03-65]. basis-deficiency-allowing [LGGSP06]. Baskakov [GQL05, ZG02, ZZ09a]. batch [GS07b, KK06c, SG08]. bathtub [SY05]. Bayes [Elk06, GZZ09, Jah04]. Bayesian [BGS04, JAM02, KLHC08, KC03b]. bazaar [Ano00-75]. BBK [HL08c]. BBM [Che04b]. BBM-equation [Che04b]. BCCB [CJ05]. BCH [Sol00]. BCI [JLZ09]. BCI-algebras [Jun08, Jun09, JLZ09]. BCK [Jun08, Jun09]. BCK/BCI [Jun08, Jun09]. BCK/BCI-algebras [Jun08, Jun09]. BD [DH03]. BDF [ZL02]. be [LL07, WW04b]. beam [CO07, GX04, LHC08, Liu00]. beat [Lin08a]. Beating [DFY06]. becoming [Ano01-28]. bed [VM09]. Bedau [Ano03k]. beds [Gla07b]. Beesack [CD09]. before [Ung03]. beginner [Ano00-74]. beginning [Ano01p]. Behavior [Wan06b, ATY08, CS02c, Dix05, Dua09, FDWC02, GL09b, HLS05, HLY05, JKN09, JT07a, KR06, Kra01, LLT08, LZ02, LG04a, LST02, Med01, MC02, Pen09, PP04, Qiu08, SZJ09, Tan02a, TL00, TMG03, Wan06c, Wan09e, XLL08, Xu01, ZW04, ZY05, ZYH01, ZY08b]. behavioral [LBZ^{+}09]. behaviors [Guo03, HC09]. behaviour [Dib03, EZ08, IWL04, IW06, MG01, Zha01a]. Behind [ANV00, Ano00-47]. Behrooz [Ano00s, Ano00t]. Beijing [Ano00u]. Beissinger [Ano03-62]. Bekoff [Ano03-78]. Bell [Ano00-71, BNR05, NR04, Ikh08, WW09a]. Beller [Ano00-64]. Bellman [BH01, DBGSP08]. bells [EPS05]. below [ZQC08]. Beltrami [BMH08, LXZ08, Xu04]. BEM [FP02, YW03b]. Ben [Ano03-50]. Benassy [Ano04-27]. Benjamin [YM09]. Benker [Ano00-58]. Benoit [Ano03-35]. Berger [Ano00q, Ano00r]. Bergman [RNM00]. Berlin [Ano00q, Ano00r, Ano00o, Ano00-58, Ano01i, Ano01g, Ano03-42, Ano03-82, Ano03e, Ano03-69, Ano04i, Ano04o, Ano04p, Ano04v, Ano04z]. Bernhard [Ano01g]. Bernoulli [AM04, KA02, LHC08, LS06b, MV08a, WJW08, ZY08a]. Bernstein [AG09, MP07a, PP01b, XX08]. Bernstein-like [MP07a]. Bernstein-type [PP01b]. Berrington [Ano00-71]. Bertsekas [Ver08b]. Bessel [LLNS05, Mal04, Rap01, Sim01, WLS06]. best [Gan00, KT03a]. Bestselling [Ano00-79, Ano00-78]. beta [Nad05, PG06, ZC06]. Bethe [He09a]. better [Lin04c, MG08a]. between [AO01b, AK07, AO09, CD03, CC08a, CKR03, CK04, CT02b, Coh00, Fro08, GGD00, HdLS03, HN07, Isk08, Jun01, KGB07, Lee01b, LS06b, MO07, ÖY08, PB02, TY03, WL08b, WY01, WB01, YL09b, YX09b, dSRGRMRL08]. Beverton [SZJ09]. beyond [ANV00, SA08]. Bézier [GJQ07, Gup04, KA00a, Las08, ZG02, ZL08a]. BFGS [Mur05a, XWW08, YL08c]. BFGS-type [XWW08]. BGK [BJK^{+}09, Mie01]. BGK/level [BJK^{+}09]. Bi [MCC09, YZHY09, BG01, KH09]. Bi-criteria [MCC09]. Bi-level [YZHY09, KH09]. bi-temperature [BG01]. bibliographical [Blo00]. biconjugate [VB01]. Bicriteria [WT00]. bicycle [Yav05c, Yav06, Yav07b]. Bid [Bel06, Che08b]. bidding [LTYB09]. bifinite [CZ06]. bifunctions [FH07]. Bifurcation [HD08, Ma07, DZ05, DL08, Far02, GK04a, HZ05, Ji02, Li05, Lu09c, NT08, TZ04a, ZHZ00]. Bifurcations [HYZ02, LC02, MM06]. big [CH00b]. biharmonic [AV00, CWLC07, Har01, LL02a]. Bilevel [LY05a, LHWW08, WWWL08]. bilinear [Bib06, GKM08, LBW09]. Bill [Ano00-39, Ano00-40]. billing [IDOK06]. bimodular [MVS^{+}07]. binary [ARS02, Ano05f, BB09b, Lee09, OY06, SS09d, SM06, YL04b]. binding [MD06, XY04]. Bingham [FF07]. Bingham-like [FF07]. Binnard [Ano00x]. binomial [HHS07, Sun09b, Wan09c]. bioacoustic [BHCØ07]. biobjective [Wu04]. biochemical [AMYS^{+}09, HYZ02, LMA08, TZ04a]. biofluids [BHS^{+}08]. biographical [Blo00]. bioimaging [FDY^{+}08]. bioinformatics [Ano03-74]. biological [BBH07, SD07]. biology [YAVB08]. biomaterials [MVS^{+}07]. biomedical [BFFL08, OTB^{+}08]. biomimetic [Ano03-55]. bioremediation [CCK08]. Biorthogonal [He04, Xie08]. bipancyclic [SLTH09]. bipancyclicity [SWKT08]. bipanpositionable [SLTH09]. Birkhäuser [Ano03w]. birth [LP00]. birth-death [LP00]. Births [BFL06]. Bispectral [TGIM02]. bisplit [MV08b]. bistable [SW07]. Bisymmetric [Wan05a, YLL08]. bit [JKY06]. bit-serial [JKY06]. Bivariate [DZZT06, LW09a, LHC05, AG01c, AG01b, CL08b, HL00, LF09, WFLC05, WYL08b]. Björn [Ano00-28]. BL [MZD09, ZX08]. BL-algebras [MZD09, ZX08]. black [CL02a, Nam02, AE08, CNPP08]. bladder [GMSRP08]. Blank [Ano01s]. Blank-Edelman [Ano01s]. Blasius [FLfFL08]. Blending [LH00b, CCS00, LL02a]. blending-type [CCS00]. blind [TLP^{+}09]. Bloch [XC04a]. Block [JL09b, MK03, BL01, CFM02, KP05a, LBW09, LS02, PV01, STH08, Whi00, WH09]. block-coding [CFM02]. blocking [HH00]. blockquasi [CPS05]. blockquasi-tridiagonal [CPS05]. blocks [JL09b, WSG08]. blood [AB09, BHS^{+}08, BW09, bFFlT09, LW05, MM00, MP08a, MG01, MP02b, Mur05a, SASHS09, SS09e, SM08, WB09]. Blow [Che04e, DL05, DXZ05, Wan02b, ZW06, DDX04, LX03, SW07]. Blow-Up [ZW06, Che04e, DL05, DXZ05, Wan02b, DDX04, LX03, SW07]. blowing [ES05, MM07b]. Bluman [Ano04y, Ano04z]. BMO [XD09]. board [Ano01-29, Ano02e, Ano02f, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano00-95, Ano00-96, Ano03-91, Ano03-92, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano05d, Ano05e, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v]. Bob [Ano00-38]. bodies [IN07, Sta07]. body [Ano03-65, HW05a, RB07, SYHC06]. Bogdanov [TZ04a]. boiling [JTRu09]. Boltzmann [AZB08, AS09b, AVBW08, AB09, Asi08, Bab09, BHS^{+}08, CJ08, CJR09, CM09, CLL09, CLT^{+}08, DS08, Dub08, HKL09, ITR08, IMLF09, JTRu09, JY08, JY09, KIXF09, KOCO08, KMK09, Lee09, ME00, MHTR08, Mie01, OKCO08, PW00, STK08, SP08c, SGR^{+}09, SM08, TBDL09, TK09, VRJ09, VM09, WB09, Yon09, YG08, dG09]. Bona [YM09]. Bonatti [Ano01n]. bond [Gre01]. bondage [HX06, HX07]. bonded [AS03a, EA03b]. bonus [Ano00-78, Ano00-79]. Book [Ano00c, Ano00e, Ano00j, Ano00f, Ano00k, Ano00d, Ano00i, Ano00-74, Ano00l, Ano00-75, Ano00q, Ano00s, Ano00m, Ano00r, Ano00t, Ano00u, Ano00o, Ano00p, Ano00n, Ano00z, Ano00x, Ano00y, Ano00w, Ano00v, Ano00g, Ano00-29, Ano00h, Ano00-28, Ano00-27, Ano00-76, Ano00-30, Ano00-77, Ano00-31, Ano00-32, Ano00-33, Ano00-35, Ano00-34, Ano00-36, Ano00-37, Ano00-38, Ano00-39, Ano00-40, Ano00-42, Ano00-43, Ano00-41, Ano00-44, Ano00-48, Ano00-45, Ano00-49, Ano00-47, Ano00-51, Ano00-46, Ano00-50, Ano00-52, Ano00-54, Ano00-53, Ano00-57, Ano00-55, Ano00-56, Ano00-60, Ano00-78, Ano00-61, Ano00-79, Ano00-58, Ano00-63, Ano00-62, Ano00-59, Ano00-64, Ano00-67, Ano00-66, Ano00-65, Ano00-68, Ano00-80, Ano00-71, Ano00-69, Ano00-81, Ano00-70, Ano00-85, Ano00-72, Ano00-73, Ano00-83, Ano00-84, Ano00-87, Ano00-82]. Book [Ano00-86, Ano00-88, Ano00-89, Ano00-92, Ano00-91, Ano00-93, Ano00-90, Ano01c, Ano01e, Ano01d, Ano01f, Ano01i, Ano01g, Ano01h, Ano01j, Ano01k, Ano01y, Ano01m, Ano01n, Ano01z, Ano01o, Ano01p, Ano01l, Ano01q, Ano01r, Ano01s, Ano01u, Ano01t, Ano01w, Ano01x, Ano01v, Ano01-27, Ano01-28, Ano02c, Ano02d, Ano03i, Ano03-76, Ano03g, Ano03f, Ano03-36, Ano03-77, Ano03j, Ano03k, Ano03h, Ano03l, Ano03t, Ano03m, Ano03n, Ano03p, Ano03u, Ano03s, Ano03r, Ano03-78, Ano03o, Ano03w, Ano03q, Ano03v, Ano03y, Ano03-27, Ano03-79, Ano03x, Ano03z, Ano03-29, Ano03-30, Ano03-28, Ano03-80, Ano03-31, Ano03-81, Ano03-39, Ano03-35, Ano03-32, Ano03-33, Ano03-38, Ano03-34, Ano03-37, Ano03-41, Ano03-42, Ano03-40, Ano03-85, Ano03-44, Ano03-82, Ano03-43, Ano03-83, Ano03-45, Ano03-46]. Book [Ano03-47, Ano03-50, Ano03-49, Ano03-48, Ano03-51, Ano03-52, Ano03-53, Ano03-84, Ano03-86, Ano03-56, Ano03-55, Ano03-54, Ano03-57, Ano03-58, Ano03-60, Ano03-61, Ano03-63, Ano03-62, Ano03-59, Ano03-64, Ano03-87, Ano03-65, Ano03-73, Ano03-72, Ano03-66, Ano03-70, Ano03-67, Ano03-68, Ano03e, Ano03-69, Ano03-71, Ano03-74, Ano03-75, Ano03-89, Ano03-88, Ano03-90, Ano03b, Ano03c, Ano03d, Ano04e, Ano04d, Ano04f, Ano04g, Ano04h, Ano04j, Ano04i, Ano04k, Ano04n, Ano04l, Ano04m, Ano04-27, Ano04o, Ano04p, Ano04r, Ano04q, Ano04-28, Ano04s, Ano04u, Ano04t, Ano04c, Ano04w, Ano04v, Ano04x, Ano04y, Ano04z, Ano04-30, Ano04-31, Ano04-29, Ano04-32, Ano04a, Ano04b, Ano05a, Ano05b, Ano05c, Ano06b, Ano08a, Nam02, Rod00, Rod01a, Rod01b, Rod02a, Rod02b, Rod04, Rod05, Rod06]. Books [Ano00-29, Ano00-59, Ano00-78, Ano00-79]. bookshelf [Ano00-78, Ano00-79]. Boolean [LSL09, CCS00, CSC^{+}09, JWSS04, Luo06, Wan00a]. Boolean-type [LSL09, Luo06]. bore [EW00b]. Borgese [Ano03-57]. Borjas [Ano01o]. Bornemann [Ano03-67]. Bose [ZZ09c]. Boston [Ano00d, Ano00x, Ano00w, Ano00-28, Ano00-50, Ano00-56, Ano01d, Ano01h]. both [FG00, Tan03b, ZSZ06, ZY07]. Boulder [Ano03g, Ano03n, Ano03s, Ano03-33, Ano03-70]. bound [CZ02, FB05, HL03a, SHWF09, ZQC08]. boundaries [BGR09, DLSB08, FDWC02, Pos08]. Boundary [CJ08, CLL09, GC06, GCN06, Liu06b, Mas07, NRL08, Ram05, ABBD01, ABBDP04, AS03b, ATT03b, AOS07, AN09, AM09, AS01, AD03, AL08b, AA01, And03, AK08b, AS08, AC03, ACH01, Ayk04, BG04a, BLD07, Bai08, BJBL03, BHNO04b, BM05b, BW05, BW06c, BFO02, Cab01, CH01, COE03, CC09a, CM09, CG08, CM04, CS02b, CK02a, CMS04, CT02a, CWZZ07, CO07, CL08a, CKKK06, CH00c, CW05, CZ08, CZ09c, Cui09, CM08, DPU07, DRR09, DR01a, DMP03, DL05, DHM05, DG05, Du08, DXZ05, Dua09, EGZ04, EGA03, ES08a, EBVS06, FFK02, FZG07, FR04, GNR05, GH03, Gra03, GHJ06, Ha01, HL02a, HH02a, HH02b, HW05a, He02, HG03, HH04b, HW09, HT02, HY03, HLMR08, HLC05, IS04, ITT07, IMLF09, Jan04a, Jan04b, Jan04c, Jan05, Jan07, Jia00b, Jia01b, Jia01c, Jia02]. boundary [JG04, JX04, JY09, KP01, KK09, KA09b, KK04, Kel06, KS00b, KL08b, KKZ02, KLW09, KKM00, Kru08, LP08a, LW09a, LR03, LLWY00, Li04b, LCGX06, LTWM08, LS08, LLJ09b, Lia04, Lia05, LG09a, LDG05, LJL06, Lin07, LT04, Liu02, Liu04a, LQ04, Liu04b, Liu04c, LHC05, LY05b, Liu07a, LLW07, LDK08, Lu07, LSN05, LJ08b, Ma00, Ma01, Ma04, MDG05, MZ09, MS04a, Mar07, McD07, MEG08, MS04b, NVAR03, NRL00, NMD07, NMD08, OIL08, OIL09, Pal01, PLG07, PGT08, Pao03, PKP09, PM05a, Pru09, RR08, Roo08, RSK02, Ryo04, SA05, SS09b, SR04, SC03, SW07, Slo09, Son08, Sta02, SPW08, SZL02, Sun09a, TLLM08, TD07, Tho01, TGS07, Top04, Tur07, VBG07, VNG04, VJ04, Wan00e, Wan03a, Wan03b, WG07a, WG07b, WLW07, WJA08, WWW08, WSL08, Wan08a, WG08a, WFY09, Wan09d, Waz00, Waz01, WA08, Won00]. boundary [WX03b, WCW08, Won08, WK05, WZ08c, Wu09, XXD06, Xu07a, XL09, XG07, Yan05a, Yan06a, Yan07, YW07, Yao04, Yao05, Yao07, Yas08, ZM04a, ZLY06, ZL08c, ZYW08, ZL08d, ZJ08, Zha09c, ZG09a, Zha09d, Zha08d, ZG09b, ZC09a, ZCC04]. boundary-element [BJBL03]. boundary-layer [XL09]. boundary-linked [EBVS06]. boundary-only [CT02a]. boundary-value [EGZ04, KA09b, Won08, ZG09a]. Bounded [DJ08, DN03, DZZT06, KK05, MRT04, PS01, Tan02b, BCGA05, CCD07, Che08a, Dra08, JY08, LZ04a, Liu00, THD08, XC07, Zen00, ZG02, ZC06, ZLH07]. Boundedness [LC07b, Zha01b, LLT08, WD05, ZSKZ03, ZLCM08]. Bounding [CD09, AN05, HHR07, ÜL02]. Bounds [BD02, BGNA07, CD05b, CD06, Coo05, DS04, PS07b, RSR00, AJ09a, AJ09b, CG00, CD03, CBJ01, CRK02, DSS00, Gal05, Hu08, JAM02, JMR00, Kum05, KC05a, KNP06, Li01a, PRA01, PJJP04, TK02, Uje03, Uje07, WWS^{+}06, Xu02a, ZL07c]. Bourne [Ano03-74]. Boussinesq [AETEZ07b, CC04b, LWZ08, LC06b, SSD07, Waz05b, Waz05d]. Box [Ano04c, Yan03f]. box-constrained [Yan03f]. Bradenbaugh [Ano00-37]. brain [Ano03x, Ano03-86, Ano04-28, Ano03-77, Ano03l]. Brain-wise [Ano03l]. branch [CZ02, FLfFL08]. branches [HY08, MP05, MP08b]. breakdown [WX09a]. breakdowns [LWY09, NQHN03, Wan04a]. breaking [Tao09, YZ02b]. breast [BHCØ07, IDM08]. Breazeal [Ano03y]. Breé [Ano03-52]. Brenner [Ano03-84]. Brezis [HZ08a]. Brian [Ano00z, Ano00-29, Ano00-69, Ano03-33]. bridge [GL08b]. Bridget [Ano03-36]. brief [Ree00]. Bringing [Ano03-71]. Broadband [SGR^{+}09]. broadcast [JWC08]. broadcasting [TJ01, ZW08]. Broer [RD07, KHN09, Ma09a]. Brookings [Ano04c]. Brooks [Ano00l, Ano00m]. Brosch [LYX09]. Brownian [AL08a, Gre04, HH09]. Bruce [Ano04k]. Bruijn [HX06, HX07, May00]. bubbles [BBB08, Kor00a]. buckling [CA09]. budget [Lin06b]. buffer [GM09b, GSS04, SG08]. Building [Pio00, LG09b, Ano01f]. buildings [KOV08]. Bulirsch [Ano03-46]. bulk [GM09b]. Bunce [Ano00-62]. bundle [SP08a]. bundle-type [SP08a]. Buoyancy [SHKL08, SBPD08]. burger [Zha01a]. Burgers [SGM09, BA09a, Hay06, KC02, Kay04, KX05, Ras06, SSH09, WZ08c, cW09b]. Burghardt.The [Ano03-78]. Burmeister [SZ09b]. Burr [JAM02, MJ02]. bus [SMO03]. buses [Ano00-90]. Butte [Ano03-53]. BVP [SSX09, SAS06, WWYL09, YWYL07]. BVPs [AORS03, DM08, GM09c, LS04b, WW03, Yil08].

C [Ano00-28, Ano00-35, Ano01z, Ano01r, Ano01x, Ano03p, Ano03s, Ano03-42, Ano03-46, Ano03-84, Ano03-59, Ano03-67, Ano04u, Ano04w, Ano04v, Ano04y, Ano04z, LL05, ASESESA09, FP06]. CA [Ano00-75, Ano00v, Ano00-27, Ano00-34, Ano00-36, Ano00-37, Ano00-39, Ano00-40, Ano00-44, Ano00-45, Ano00-52, Ano00-53, Ano00-57, Ano00-60, Ano00-78, Ano00-61, Ano00-79, Ano00-63, Ano00-62, Ano00-67, Ano00-65, Ano00-68, Ano00-72, Ano00-73, Ano00-87, Ano00-82, Ano00-86, Ano00-92, Ano00-91, Ano00-93, Ano01f, Ano01j, Ano01k, Ano01r, Ano01s, Ano01-27, Ano04q, Ano04-32, CCK03, CCKH07]. cache [BK07, CCHK08]. cache-aware [BK07]. CAD [Ano00-48, Ano00-49]. Cahn [CCK00, UK08, Ye02]. calculate [BI01]. Calculating [Bel05a]. Calculation [SBD^{+}00, AC04, AN06, Lam05, Mic09, She02, YW03b]. calculations [WK03]. Calculus [LG09a, BG07b, CGL01, Fer01, Gau04, JSS01, LTSW02, LTS03, LLNS05, SV01, SM00, TWS00, WLS06, Ano03-64]. calendar [Sch00]. Calendrical [Ano03m]. calibration [KYJ03]. call [CKKK06, WLT07]. callosum [Ano04-28]. Calude [Ano03-45]. Cambrian [Ano00l, Ano00m]. Cambridge [Ano00e, Ano00j, Ano00f, Ano00k, Ano00-74, Ano00l, Ano00m, Ano00n, Ano00-76, Ano00-30, Ano00-77, Ano00-38, Ano00-43, Ano00-41, Ano00-54, Ano00-59, Ano00-69, Ano00-83, Ano00-84, Ano00-88, Ano00-89, Ano01e, Ano01y, Ano01n, Ano01u, Ano01t, Ano01x, Ano01-28, Ano03-36, Ano03-77, Ano03j, Ano03k, Ano03l, Ano03t, Ano03m, Ano03-78, Ano03o, Ano03v, Ano03y, Ano03-79, Ano03x, Ano03z, Ano03-28, Ano03-81, Ano03-32, Ano03-38, Ano03-41, Ano03-85, Ano03-83, Ano03-47, Ano03-50, Ano03-53, Ano03-86, Ano03-55, Ano03-54, Ano03-60, Ano03-61, Ano03-63, Ano03-87, Ano03-66, Ano03-71, Ano03-89, Ano03-90, Ano04e, Ano04d, Ano04g, Ano04h, Ano04n, Ano04l, Ano04-27, Ano04-28, Ano04-30, Ano04-31, Ano04-29]. campaign [Bel01]. campaigns [Bel05b]. Campus [Ano04c]. can [LL07, WW04b]. cancer [IDM08, XMW09]. candidates [Bel05b]. canes [LCYL00]. Canonical [Rap01, Cho03, IK04]. capacitance [YW03b]. capacitated [Lin01, XD02]. capacitated-flow [Lin01]. capacity [Jai05, WK08]. capillary [OKCO08]. Caputo [Mur06, Mur08b]. car [FY04]. Carathéodory [YO06]. carcinoma [GMSRP08]. cardinal [All02, Sun00]. care [Ano03-63]. cargo [WK08]. CARI [BHCØ07]. Carl [Ano03f, Ano04n]. Carlo [Ano04i, Lei02, Wan02e]. carpet [Jia09]. cascade [LWLJ09]. case [Ano01h, Ano03-80, Ano04-32, BM09, BK07, CO07, DSS08, Kan04, KKM00, Lin04a, MAR05b, WB09]. cash [Cap07]. Castor [Mic09]. cat [GPAA^{+}09, DP08]. catalytic [GW01a, HRP01, XZ09b]. catastrophes [KA00b]. categories [EGM^{+}02, HLCL00]. category [CZ06, YL05, ZAAESG09]. cathedral [Ano00-75]. catheterized [SS09e]. Cau [Ano00-55]. Cauchy [CP04, El 03, Les05, LX00, Mar05a, SWX06, Yil09]. Cauchy-likesystems [CP04]. causal [PP06]. Caused [Wan06a]. cavities [ZRS05]. cavity [DS08, Gre03, MK08, SHKL08, Zha03b]. Cayley [CLS08, Pet04]. CD [Ano00-34, Ano00-39, Ano00-40, Ano00-46, Ano00-78, Ano00-79, Ano00-83, Ano01f, Ano01k, Ano03v, Ano03-34, Ano03-58, Ano03-67, Ano03e, Ano03-75, Ano03-88, Ano00-78, Ano00-79]. CD-ROM [Ano00-34, Ano00-39, Ano00-40, Ano00-46, Ano00-83, Ano01f, Ano01k, Ano03v, Ano03-34, Ano03-58, Ano03-67, Ano03e, Ano03-75, Ano00-78, Ano00-79]. Cebysev [CD09]. Ceder [Ano00-85]. Ceiling [Bel06, MO08]. cell [AACLM08, BI01, CL00, MM08c, SGR^{+}09]. cells [LW05, LMSH04]. Cellular [GY06, CSC^{+}09, YMI06, Zha03a, Zha04b, Ano03n]. censored [MJ02]. center [CR06, DL08, Gin02, HLC00, HLM04]. center-finding [HLM04]. centers [DLM08, GK04a, GL06, HRS08, ZHZ00]. central [MP08b]. centro [DH08b]. centro-symmetric [DH08b]. centroid [CT02b, WL08b]. centroidal [DW02b, DW05]. centrosymmetric [Wan05a, ZZH03]. century [Ano00v]. Ceramics [LYF09]. cerebral [BW09]. Certain [Aou08, CCS02, LTSW02, SBS06, WSZ09, YI06, AG01a, AÖS04, AIS08, ASS08, AY09, ÇT05, CK06, CK02c, EAA09, EGA03, Fro08, GYS05, Gre04, Hil01, HdLS03, HL05, HL03b, KK06a, LTS03, MA09b, NCK^{+}08, Sri03, TA04, Trz01, XA09, ZD01]. Cesaro [Kua08, Alb02]. CFD [Mas07]. CG [Yan03d]. CG-type [Yan03d]. Chain [Ano04i, MR06, OSS07, BLD07, DK05, LR03, LP00, Nak06, WK03]. chains [BC01, LXH05]. Chaitin [Ano03u]. Chalmers [Ano00-84]. chamber [Dem00]. Chan [LHR09]. change [CYH05, Kar08, Liu04b]. changes [BH03]. Changhee [KR07b]. changing [AS03b, DCS09, DG05, LS09, LLJ09b, SSX09, Yan06a, YW08a]. channel [AH08, CCZK09, Gla05b, HASA07, JS04, Kan04, Mak09]. Chaos [MS03, Ano00g, Ano00-29, Ano00h, GLTS08, MM06, TCX08, Ano00p]. chaos-evolutionary-programming [GLTS08]. chaotic [DW09, GLTS08, LW09f, TLP^{+}09, WY09a, XXC^{+}09, ZXTL09]. Chapman [Ben01]. character [ABH08, SX06]. Characteristic [Che04f, ZCK00, CZW08, Fer01, GY08]. Characteristic-nonconforming [Che04f]. Characteristics [DHY06, Ikh08]. Characterization [HLWT04, QC06, ZGY05a, Pra07, VDMK08]. Characterizations [Li04a, XM08].

Characterizing [BCA09, Mär05c, WGF09]. Charafi [Ano03-34]. Charles [Ano00-27, Ano03-54]. chasing [She04b]. Chebyshev [AK06, Bar01d, Boy01, Boy07, CBJ01, CGBJR03, CN04, DGVP06, Her01, KT08, RTW05, Ser00, YW03a]. checking [KNP06]. checkpointing [NFI03, NUN06]. Chee [Ano00-31]. chemical [Gre01, Jon07, NB08]. Chemically [BR06, MRMRM02]. Chemically-Reacting [BR06]. chemistry [GNM07, dR03]. Chen [LC06a, LLXC03]. Cheng [Ano04e]. Chengtian [He09a]. Cheung [Ano03-64]. Chicago [Ano00p, Ano00-42, Ano00-64, Ano00-80, Ano00-81, Ano01z, Ano01o, Ano01q, Ano01v, Ano03f, Ano03-29, Ano03-80, Ano03-57, Ano03-62, Ano04x]. Chichester [Ano00-55]. Chii [Ano04d]. Chinese [ARS03, WW02a]. Chircop [Ano03-57]. choice [Ano04g, Hoh01, YZHY09]. Cholesky [Més05]. choosability [CZW08]. Choosing [KT03a, HY09, HB04]. Choquet [Pos08]. Chor [EMD08]. chord [Gil00, HS02]. Chree [SK04]. Chris [Ano03-54]. Christodoulos [Ano00y]. Christoffel [Ano03-66]. Christoph [Ano03e]. Christopher [Ano03-90]. Christos [Ano04-30]. Chromatic [ISAAM09]. Churchland.The [Ano03l]. CICS [Ano01k]. circle [BCA09, CS04b, GM09a, SS06]. circuit [LL08a, MNIS06]. circuitry [DW09]. Circuits [Ano04d, YSP^{+}08]. Circulant [CZ00, JLW04, BFM04, FL05, JL09b, PSM02]. circular [Che01a, FMFV08, Gil00, Gre09, Mic08, SAS06, Son00, Zhu05, ZS08b, Zhu09b]. circumstance [ODO00]. Circumventing [KH00]. citations [Ano03-60]. city [Ano04l]. civilized [Ano01m]. Claes [Ano00-28]. claim [YLD08]. Clarendon [Ano01m, Ano01l]. clarifier [BKR05]. Clark [Ano00-80, Ano00-81]. Class [LET06, Mig06, PXXW06, Sat06, ZHZT06, Agr08, Bae07, Bai00, BCRS05, BN00, CYG06, CGY08b, Che08a, CC08b, DL05, DH03, DLM08, EGA03, FW08, FM07, FZG07, FLW06, GL06, Gon09, GM09c, Guo02, HH02a, Han04, Han06, HR05, Hua01, HZ07, HZ09b, JA03, JX04, KSZ02, Kot01, KC05a, KHS00, LDZW05, LBZ^{+}09, LWZ09, Lin08c, LC02, LC06c, MP07a, ND01, NRL08, OR01, OZY05, OSSY04, PS03a, PXGL09, PP06, RC07, SZ05, Sun05b, Sun05c, Tao06, TCX08, Ver01a, WYH00, Wan03b, WFC04a, WH08a, WYL08a, WYL^{+}09, WS04b, WL09b, Xia07, XTX01, XZ08b, XX09a, YAVB08, Zen09, ZT04, ZLH07, ZYZ08, ZG09a, ZZ05, ZXL07, ZYCI09, ZH02, ZHL04, ZH04]. Classes [PMS07, AR07, AY09, CCS02, DFFL09, Hil01, PWX08, Pie09, WSZ09, WHL^{+}09, ZCS04, Zha09f]. Classic [TSY09]. classical [BJ00, TB09, YL05]. Classification [Li01b, An03, BFFL08, Bil05, FM08c, GZZ09, XMW09, ZZ09b]. classifier [HJWL03, Lin08a, WYL^{+}09]. classifiers [SS09a]. CLAWPACK [FL03]. cleaning [HL09a, Ano03-31]. Clements [Ano00-30]. climates [KOV08]. clinical [FDY^{+}08]. clock [CGFS09]. clock-controlled [CGFS09]. cloning [Tan00]. close [NT08]. closed [LGH09]. closest [LG00]. closure [BJK^{+}08]. cloth [Ano00l, Ano00m, Ano00p, Ano00-64, Ano00-80, Ano00-81, Ano03-49, Ano03-68, Ano03-74]. clothing [LHLW07]. cloud [De 04]. Clough [LF09]. CLSI [Ano04-32]. Clues [May00]. cluster [GYPL08, HH09]. Clustering [CW06, IS06, FDY^{+}08, GYPL09, HLSL03, KPD^{+}09, KWHC05, PL01b, PCA09, SP09, YL09b]. clusters [Ano01f]. CO [Ano03g, Ano03n, Ano03s, Ano03-33, Ano03-70, AI06, KLZZ09, CWY02, DM09, FDY^{+}08, ZH09]. co-clustering [FDY^{+}08]. co-coercive [ZH09]. co-complementarity [CWY02]. Co-evolutionary [KLZZ09, DM09]. Co-Quasi-Variational-Like [AI06]. coagulation [BHS^{+}08]. coastal [WHL02]. coaxial [Lee01b]. cocoercive [CyL08, LCV06, yL09c]. code [AFI06, BBGG08, PG06, PG08, Pos00, Yan03b]. coded [ZXTL09]. codes [AMM00, AB08c, Ano00-54, BSM01, Mic08, MPP08, Ree00, Sha02]. codeword [AMM00]. Codimension [Lu09c]. coding [CFM02, LHW04, MK03]. Coefficient [Cha06, Pan05b, ACCMJ07, Fer01, GT00, GS01, GK04b, JP01, KZ02, KHN09, MV08a, PDRR08, STH08]. Coefficient-free [Pan05b]. Coefficients [Ikh01, Ikh02, Ikh04, AÖS04, AFI06, BB06, BBP09, Bib06, CN08a, EM09c, HZ08a, JKR08, KÖ08a, Kar09, Lai08, LZ02, LLD02, Öca07, OIL08, Sal08, SW04, SS06, TYP00, WZ00, Wu08a, ZhYL05, ZS08a, ZZ03b, ZLLL08, ZY08b, ZC01]. coercive [MM08d, ZH09]. Coexistence [LP04, WW02b]. cogeneration [She05]. cognition [Ano03-78]. cognitive [Ano03-78, Ano04-28, MPS04, TDHL09, Ano03o]. Cohen [Gon09, LYJ09, NC09, RC07, ZLH07]. coincidence [DP02, SP08b]. cold [KOV08, TG03]. cold-electron [TG03]. Colin [Ano03-78]. Collado [Ano03-41]. Collado-Vides [Ano03-41]. collapse [PP08]. Collected [Ano03n]. collection [IMPF08]. collections [Ano04q]. Colleen [Ano03o]. College [Bel05a]. Collision [KKS04b, Ano00-71, OAVB08, PW00, YF04, Yav05c, dG09]. collocation [BDGR01, CN04, DRR09, DS07, DH07, EBVS06, FFK02, GPMPR03, HLC05, HL06, KA09a, KT08, LP08a, LT04, LLTL05, MPS04, MEG08, PB02, RR03a, RR08, SV06, UHuI09, Ye02, YLB06]. collocation-interpolation [RR03a]. colloidal [OKCO08]. colony [KS07, MWZ09]. Color [Ano00-42, PLP09, LH00a, LLT09, Ano00-42]. coloring [HWC04]. column [LGGSP06]. columns [AC09, CA09]. Combination [JY11, YLY^{+}09, Ber05, Jun01]. Combinations [HL06, Ano04-32, BB08a]. Combinatorial [Ano00o, Ano00-41, Ano03h, HSG^{+}06, LW09b, Ano01g, Ano03p]. combinatorics [CGL01, DLL08, Ano03q]. combined [BJK^{+}09, DS02, Li01d, NS07, TZS03, mWmZ09]. Combining [VTR04, WYL^{+}09, bJQxC09]. combustion [CLT^{+}08, GW01a]. come [Ano00-90]. comfort [ES08b, LHLW07, NLL07, NLL08]. comma [MPP08]. comma-free [MPP08]. Comment [Asl11, EM09b, ES08b, Liu08b, NLL08, Ari09c, HPT05, NLL08]. commentary [Ano03-80]. Comments [Isk02, Ter00, hYL09, HM05b]. commerce [Jua03]. Common [Ano03f, ĆRRU08, yL06, Rad09, CWY08, CKQ08, udK07, KFMH06, NOY01, OS09, PKT07, TCH07, Wan07a, YYZ07]. commonKADS [Ano00-38]. Communication [LLS^{+}09a, YCCY02, Ano03f, DM09, KM03b, KYNI06, LI00c, RVKDF09, SMOK05, TLP^{+}09, XCH07, Ano03-85]. Communication-efficient [LLS^{+}09a]. communications [LW09f, YMI06]. communities [KPD^{+}09]. Community [RMC08, WSO08]. commutative [Ano03e]. Commutativity [TWS00]. commuting [BB08a, PV01]. Compact [LCL06, AG09, CGGC02, Gal08a, Gar01, KZ02, KO09a, LW03, PG06, PG08, SN02, WJA08, Waz04, Waz05b, Zha03b]. compactly [WW08c, YTEC02]. compactly-supported [YTEC02]. compactness [Bri00, DMSV00, Dog01]. compacton [WH07]. compacton-like [WH07]. compactons [Waz05d]. companion [KP05a]. comparative [Ano04u, BUP09a, Dim08, GC08]. Comparing [BCDF02, HB04, Ano00n]. Comparison [AP08, BJK^{+}08, CJR09, Isk02, KGB07, LR03, LX06, MHTR08, MO07, ÖY08, VA04, AB09, Bao03, CH01, CHY01, DGG^{+}04, GGE09, Gla05a, Gla06, Gla07a, Isk08, Jan05, JEP08, LRL09, LY04, PKSP08, PB02, SK04, SH06, SDZ06, Sma01, Sür08, TRG01a, WF07]. comparisons [CJ04, Hof03, HZ09a]. compatible [ĆRRU08]. competence [HCTC03]. competing [BGS04]. Competition [SSLL09, Ano04-27, FW00, GW03b, Mar09, MM08b, Wan01, ZLZ09]. Competitive [Bel05b, CF04, GY09, Hoh01, LY08b, LF02, LL03c, TL00, WW04b, YWZ03]. complementarity [CWY02, CKL07, HLP05, IL04, JT05, Lóp09, Tao06, XZ08a, YL09a]. complemented [CCK03, CCKH07]. complete [XH02, ZT08]. Completely [AI06, CGY08b, PL05b, Xu02a, ZGY05b]. completion [MS09, TT09, YC07a]. completion-time [MS09]. Complex [Mis05, RV05, AG09, Ano00u, BCT00, CLL09, DCW08, Gal08a, Has04, HR08, HJL05, MOP06, MA09b, ME03, Nic04, NB08, PS01, RRR07, Rap01, SS06, TaL09, UHuI09, WSO08, Waz05c, WSZ03, XCL^{+}08, XZ09a, ZL09a]. complex-free [WSZ03]. complexes [Dar08]. complexities [SDCL08]. Complexity [Ano01i, CKL07, XW05, Ano00-70, Ano03n, CG00, CWLC09, Lan08, Ano00o, Ano00p]. component [GK04a, LGBS00, MP05, WZ07a]. components [TBR01, ZZ09c]. composed [DH08a]. composite [EA03b, mLpW08, LY08d, LB04, MUHdH03, YW09]. composition [Ano00x, LYL00a]. Compositional [WZE^{+}06]. compositions [XWD04]. compound [HCTC03]. compressible [Ban00, Fei07, Gla04, Gla08, NP07, Pad06, ST07, Tan01]. compression [ABH08, AB08c, CLX08, KP03, Kor07, OY06, WL05, WLL06, ZLN06]. Comput [AR10, Asl11, Ata09, fDxZ11, ES08b, ID10, ID08, JY11, KÖ08a, KÖ08b, Liu08b, Liu09a, NLL08, SW08, Ver12, Yan08, Zho08].

Computable [JMR00]. Computation [Ano03-54, AMR05, JWSS04, Ano03-32, Ano04-30, Ari09a, Fan02a, FW04b, GK04a, GM00, HY00, LLCS02, Nab04, Rap01, SA08, TG03, Tun05, XLZ02, YZ02b, YZZC09, YB08, Ano00-54]. Computational [ABBD01, HH09, OSS07, Ano00-48, Ano00-49, Ano03j, Ano03-41, BN05, BS00b, BHCØ07, CFR03, CPP07, CPP08, CGGC02, Ind02, Isk08, Jon07, KM03a, LF05, PC08, RZH03, TB08, VB01, YAVB08, ZC08, Ano00q, Ano00r, Ano03s, Ano03r, Ano04f]. Computationalism [Ano03t]. computations [AH00, BLS05, CK02a, CY05, FHZ08, IR08, PIM^{+}08, Sid08, Ano03-39]. computed [LL07, WWH05]. Computer [BP09, GT00, Gre09, HK08, SHWF05, Ano00j, Ano00k, Ano00-58, Ano00-69, Ano01w, Ano03q, Ano03-69, Ano04t, Cip08, Jum01, Kor00b, LRLM^{+}01, MPS04, MG08c, NS00, ODO03, PS08b, Ano00s, Ano00t, Ano01i, Ano01h]. computer-aided [Cip08, Ano01h]. Computer-based [SHWF05]. Computers [Ari09c, EM09b, LLB05, Ano01y, Ano01x, Ano03-32]. Computing [ACCMJ07, Boy07, GSS04, PD04, SPV01, SV09, VBG07, AMYS^{+}09, Ano00-55, Ano03-50, Ano03-73, Ano03-67, Ano04o, Ano04p, Ano04s, CXZ09, DW02b, ES01, LLY01, LYC07, PST08, Ter00, WZ08a, YFT04, ZM94, ZWC^{+}09, Ano00w, Ano00-41, Ano03-56]. concave [DW08b, WY02, YHT04]. concavity [KL02, SJJ09, SJL08]. concentrating [MT00]. concentration [CGPB08, PCRR09]. concept [LSZW07, Zha02a]. concepts [Isk02, LCL00, Ano00i, Ano03-87]. conceptual [PCA09, WFT^{+}09]. concerning [YXZ09, Zha08b]. concurrent [CS00b, Dra06]. condensates [ZZ09c]. condensing [AO01a]. condition [BKN09, CCJ08, DR01a, Deh02, DS08, DXZ05, Dua09, JY09, KKM00, Kru05, Lia04, Lia05, LLW08, Mar07, OIL09, Phu02, Pru09, SANH08, SD08, Slo09, Sot01, Sun06a, Sun06b, VM09, Wan07b, Win03, WLZ07, YH09, YTCF08]. Conditional [Çin03]. conditioned [Ata09, CWZ03, Kum06, WSZ02]. conditioning [KH00, SN01]. Conditions [Gin02, AN09, ARM08, ADP08, AL08b, BM04a, BV01a, BN00, BHNO04b, BM05b, BFO02, BRMS01, COEP00, Cab01, COE03, CM09, CL07b, CLL09, CMS04, CSX01, CLT09, CR06, CWZ06, DRR09, DL05, DL08, ES08a, GS08, GM09c, HH02a, HH02b, HW05a, HH04b, HR02, HZ03, HL00, Hor08, Jan04a, Jan04b, Jan04c, JG04, LW09a, mLpW08, LX04, LHC05, Mas07, McD07, MS04b, Nie01a, NRL00, PWX08, PXGL09, Rad09, RR08, SS09b, SWX06, Son08, TL08, UA06, VNG04, VJ04, WG07a, Wu08a, YO06, Zal02, Zal03, ZSKZ03, ZG09a, ZCK00, ZCK05]. conducting [Fei07, HM05a]. conduction [MO08, Mur07a, NST01, OIL09, WZ09]. cone [Gil00, MS00, Rad09, RR09a, SZ09b, XZ08a]. cones [CPC00]. conference [Ano00-43, Ano01e]. configurations [NT08, WH08b]. confined [HZZ02, Qia08]. confinement [RGTT03]. confluent [HP02]. conformal [RNM00]. conformality [HL00]. conforming [GKM08]. confusion [KT09]. confusion-based [KT09]. congestion [HLY08, HSG^{+}06, KYNI06]. conic [QJZ09]. conjecture [Cip08, DW05, Gyi01, LH08]. conjectures [KLMR03]. Conjugacy [DP07, Far02]. Conjugate [ABBDP04, MLW09, Bax02, LJL06, Won00, ZYW08]. Conley [Zha02c]. connected [KLZ08, Kru03b, Kru08, LW03]. Connection [AO09]. connectionist [HXG03]. connections [AK07, YH03]. connectivity [HK07]. conorms [DLF09]. consciousness [Ano00-84]. consecutive [Coh00]. conservation [Ano03-80, Ano03-37, CS02b, Gos00, Kan02, PLD02, SD08, Soh05]. Conservative [CCK00, Gla00, Gla03, Gla04, Gla05b, Gla08, Gla09, Bel09, FP02, Gla05a, Gla06, Gla07a, MM08d, YS04]. consideration [KYNI06]. considerations [WH08a]. Considering [OK06]. consistency [PL03]. consistent [Bar01a, Lam05, Min08]. Constant [AOW05, AOTW07, LY06b, WY06, WS05, AB08b, AVO08, CA09, FQZ02, GK04b, KHA07, LS03, Sal08, Syn08, YT00]. Constant-sign [AOW05, AOTW07, WS05]. constantly [FAF09]. constants [CS00a, DH03, XC04a]. constitutive [MVS^{+}07]. Constrained [Din00a, Zha06, Din00b, DPJ03, GL08b, HSM00, JT05, JZT07, bJQxC09, KLZZ09, LT01, Sny00, SHWF09, SZ09b, WW02a, WY02, Yan03f, ZX09]. Constraint [ZXL07, COS09, HCG08, Lin04a, Lin06b, LHS08, RKS08, SHWF05, Yav00]. constraint-based [COS09]. constraints [FHH08, FY01, HL09b, KKY06, PKM08, Tao06, TLS^{+}08, Wan02a, Wan09b, XW05]. constricted [MG01]. construct [Wan09a]. constructed [HHS06, ID10, JW05]. Constructing [ACJ07, LW03, RR03b, YZ02b, Che04g, MAK03, cWcX09, cW09b, YJCZ08]. Construction [CC08b, JG07, Kov00, LCHL06, Xie08, ZA03b, CY00, CRK02, Chu07, PS04, PCA09, RSR01, RSE01, SC02]. Contact [Mig06, FS04, KOCO08, Sch01]. contact-line [KOCO08]. contaminant [DE06a]. content [HLY09]. Contents [Ano06l, Ano00-97, Ano00-98, Ano01-30, Ano01-31, Ano02g, Ano02h, Ano03-93, Ano03-94, Ano04-39, Ano05h, Ano06j]. context [LWZ09]. contextual [TDHL09]. Contiguous [IR08]. contingent [YLD08]. continua [BJBL03, OS08a]. Continuity [LW08a, Csa01, GA07, YWC08]. Continuous [CD06, DP02, DE06b, LL06a, PF01, ZYZ08, AC09, Bae07, BCDF02, BRMS01, BKR05, CCD07, CKZ01, HM03b, Hua04, JS03, KKS04b, KL08a, Kum04, Kum05, KU01, LYL07, Moo02, OLAS01, Pao01a, Saw03, SYL05, SF07, ÜL02, Wan09a, XM08, XTZ08, ZC06, Zhu08b]. continuous-discrete [OLAS01]. Continuous-Time [LL06a, BRMS01]. continuously [ZLLL08]. continuum [Che04a, Öze08]. contours [LT05]. contract [RS06]. contraction [DWL08, Han06, LW02]. contractions [MCQK09]. contractive [AA09c, GJM03, Inc09, Rad09, ZCK05]. contractors [JJ01]. Contribution [Gyi01]. contributions [Ano03e]. Control [BFL06, DZZT06, FY01, FY04, GC06, GCN06, LLML03, LET06, OK06, PXXW06, Sat06, XWL06, Yav00, YF04, Yav05a, Yav07a, YGH09, AHN06, ASM09, Ano05g, AK03, BG04b, Che08a, CLS07, DR04, Dra06, ES08a, ETH05, FY03a, FY03b, GX04, HRP01, Iva07, JCCT09, JCC^{+}09, Jum01, Kar05, KY00, KAKK00, KLP09, KB08, KX05, KL01, LC06b, LML02, LYW02, LY06a, MBI08, MM08a, PCE^{+}05, PKP09, PXGL09, Phu02, PR07, SDK09, Saw03, SN03, TLS07, TD07, Tur04, WLS01, WS08a, WWS^{+}06, WLT07, XCL^{+}08, XT08, YX07, YK00, Yav05b, Yav05c, Yav07b, YZZ^{+}08, ZLCM08, ZL09a, ZTH09]. Controllability [BSD00, BA01, BPM05, SBS06, GHJ06, HPT05, Phu02, SC04]. controllable [Pan05c]. controlled [AVBW08, CGFS09, ODO00, LC06a]. controller [VG08, YL04c]. controllers [Kar01, KT02, LLML03]. Controlling [WY09b]. controls [GW03b, JS03, NTHP09, PXW09, Wen00]. conundrum [Ano03-63]. Convection [ASL06, AM08a, BD04c, CAZAA00, GY08, Ili04, KZ02, KOV08, Kur08, LP08a, LY08a, MO08, MS04b, MS00, Ram05, SHKL08, SDDC08, VR02, VH06, Wan02c, ZDLH08]. convection-diffussion [VH06]. convection-dominated [GY08, Ili04, Wan02c]. Convergence [Ale06, AEÇ06, AB05, Ban00, BA08, CS02b, DE06b, HHS06, HH03, HCLJ00, Jin08, JY08, KZ02, KH08, Les02, Li00a, LZZ04, LZJ04, LRL09, LKC04, LCC08, LT09b, LY06b, MXZL09, Mie01, Ras06, Sal08, SH06, SM02, TS06, Tia05, Xu04, YH03, Zha06, ZLD08, ZHL04, Ana00a, Ana05, Arg00a, Arg00b, BM09, BA09c, BB01, CZG04, DP08, HR02, HCLF01, Hua04, HN07, HS04b, Inc09, IL04, JZT07, JDY08, Kur08, KI04, Li01c, Li02, LXZ08, MCQK09, NV08, OI00, PZ07, PV00, PPI03, PIR04, PR06, PW06, QJZ09, RL08, SDVS09, Sav08a, Ses03, ST07, Son08, SY09, ÜL00, Wan02d, WWW08, WW09b, WK03, WFS03, WF01, XCD07, XX09b, YS04, ZP02, ZG02, ZL08a, ZZ03a, ZT08, ZCK00, ZL05]. convergent [AD03, HZ09a, JWC08, SZ09b, TXC02, WXS01, Yan03f]. Conversations [Ano03u]. Converse [QLD08a]. Conversion [ID10, JW05]. converter [HRP01, HNR08]. converters [ARS02]. Convex [CCRMOG04, LK04, ZYL06, BV09, CCRMRF00, CHZ09, DP02, Din02, DPJ03, EAA09, udK07, GSS02, GMMB09, JC05, KDC08, Lan04b, LU09a, LdSF00, SM00, SL06a, SL06b, SJL09, Ubh04, WL08a, Wen08, WD07b, YHT04, YL04a, Zal02, Zal03, Ano00u]. Convexity [BW06b, SLJ04, SJJ09, Sya01b, SL06a, SL06c, WS04a, Xie01, YL07]. Convolution [HT04, RFCCRM03, Ana04a, Ana04b, DF06, Xu02a]. convolutional [AFI06]. convolutional-code-based [AFI06]. Convolutions [NG04, ND05b, ND05a, Pie09]. cookbook [Ano00-37]. cooling [LYF09, RDS00, SKU01]. Coombes [Ano00z]. Cooper [Ano00-42]. cooperation [Wan06c]. cooperation-diffusion [Wan06c]. cooperative [WW02b, ZWW04]. coordinate [AM08b, CO00]. coordinates [GGCD05, GAA06, Kan02]. coordination [HY08]. Core [AS05, AS07, AOS07, Ano00-34, Ano00-52, Ano00-53, Ano00-92, Ano00-93, CB02, bFFlT09, FF07, JTRu09, SPW08, WGF09]. Corless [Ano03-30]. Cornelias [Ano03-68]. corner [Kel06]. corpus [Ano04-28]. corrected [AC02, AC05, BD02, BX02]. correcting [AB08c, Pen03, Ree00]. Correction [Ikh08, AN05, FJ05, Lee04, WDLP04, WDLP08]. correctness [Ano01i]. corrector [PSS07]. correlated [FKM05]. correlation [BS01b, ZYG^{+}09]. correlations [RMC08]. corresponding [VNFF08]. Corrigendum [Ano00-94, Ver12]. corrosion [HHC08]. cosine [KGB07, Waz05c]. cosmological [HM03a]. cospectral [WHL^{+}09]. Cost [BFL06, HLY08, IDOK06, LW09d, Lin04a, Lin08b, PKM08, Par04b, SY05, TSY09, WI08, Ano03v]. costs [FKM05, HCTC03]. Couette [WL04]. Coulomb [AC05]. countable [YYZ07]. countably [AO01a]. counter [KM01]. counter-example [KM01]. Counterexamples [ZSKZ03]. counterpart [GMG09]. Counting [Lai08, XH02]. Coupled [PL01c, RL08, ZW06, AN09, BUP09b, CCJ08, CDJR03, CJO05, CK02b, CDJ04, CGPB08, HKL03, He02, JC00, JSF00, JPV01, LDK^{+}06, LHT07, MM08c, OR01, PJJP04, SAB07, SK09, WWW08, WLWC08, ZLD08]. Coupling [He02, MZ07, Li01d, MS04a]. couplings [LD08]. Cournot [CS05b]. covariance [BD04a]. coverage [ASM09, JCCT09, JCC^{+}09, WNS09]. coverage-control [ASM09]. covering [QW09, TDY08]. Covolume [KK02]. crack [DD00]. cracked [Kru05]. cracks [Kor07, Kru05]. Cramér [NWL08]. Cramér-von [NWL08]. Crank [ZW09a]. Crawford [Ano00-36]. Crawley [Ano03-73]. crawling [CML^{+}09]. Creating [Ano01-28]. credit [Cap07, ETH05, HL08b, LZS07]. credit-based [ETH05]. Crescenzi [Ano00o]. Cressman [Ano04h]. Cristian [Ano03-45]. Criteria [HY09, TZ01, TS06, AL04, BKK08, ÇT05, CZ03a, CS04b, CCF09, CH03, DJ08, DK05, DSS08, DAKO06, DR01b, EM09c, GS09, Hil01, Isk08, JLT04, JT07b, Kar09, LRLM^{+}01, LA00, Liu08b, LT09b, Luo02, LD02, MCC09, Man00, PG00, QSL08, SL00a, SHWF05, SS06, TX04, TÇ04, TBG05, WZG00, WWZ03, Wan04b, WL07b, XT04b, Xu08b, YYZ03, ZT04, ZWY08]. criterion [BGS04, Che08b, GZ08, HL08a, HLZ09, Kha09, Par04b, QZ05, WB01, WI08]. critical [CYG06, CA09, DSS08, Kan09, LP08c, LP08b, ZHZ00, Zha02c, ZZ00b]. critique [Ben01]. Croft [Ano04k]. Cross [Wan06a, CML^{+}09, CA09, Has03, MG03, RV01, WL07a]. cross-cut [WL07a]. Cross-Diffusion [Wan06a]. cross-language [CML^{+}09]. cross-section [Has03, MG03, RV01]. cross-sections [CA09]. crossing [MAR05b]. Crossings [DL06]. crossover [Cha06]. Crothers [Ano00-71]. Crouzeix [SB06]. CRT [ARS03, ARS02]. cruciform [DD00]. crude [SM02]. Crutchfield [Ano03-80]. Cryptanalysis [TJ01, WY09a]. Cryptographic [CKY07]. Cryptosystem [TJ01, EMD08, WY09a, ZW08]. cryptosystems [Sma01]. crystal [Tur04]. Csiszar [Ana07b]. Csiszár’s [KC05a]. CSLI [Ano04q, Ano04x]. CT [MRÁR08]. cube [CDV08, YY07]. cubes [Gor00, LTTH09]. cubic [AS01, Blo08, CL08b, DCW08, DSHS05, HSWL08, HRS08, LW09a, LG08, LC02, LHC05, LPSS02, PHAF09, TDS08, WFLC05, ZHZ00]. cubic-matrix [TDS08]. culture [Ano01m]. cumulative [QNN03]. Cure [CPP07]. currency [CM03]. Current [SK06c, CMS04, KK06a]. currents [Lee09]. curvature [Bao03, CB02, XX09b]. curve [ADN02, BV01b, CWB01, Hua08c, KU01, LW04, LHLD08, Sma01, Yav05a, Yav07a]. curved [APS07, RQH^{+}08]. curves [GPvZ06, KA00a, LH00b, LW03, MCZ00, Sya02]. curvilinear [GAA06]. Custance [Ano04-31]. customer [AP09, Hoh01, YZHY09, Ano03w]. Customer-based [Ano03w]. customers [KF05a]. cut [sLhYL09, SYJJ09, WL07a, hYLL09]. cut-points [SYJJ09]. cuts [Lin07]. CVS [Ano01j]. cyborg [Ano04l]. cycle [CHW07, GY09, LL03c, Mic09]. Cycles [ZHZT06, ANZ09, Blo08, DL08, GL06, GY09, HL04, LL08a, LC02, LPSS02, LL02b, LL03c, WH08b, XHZ05, ZA03b, ZL07a]. cyclic [OS09, WD07a]. Cyclone [HHL^{+}06]. cylinder [FMFV08, Gre09, Kan07]. cylinders [Lee01b]. cylindrical [HHK06]. Cynthia [Ano03y].

D [Ano00-87, Ano03v, Ari09c, EM09b, AC02, AM09, AMV08, Bae07, BR03, Bri00, BV01b, CM00, CM01b, Col02, DPU07, HJWL03, He00a, HZZ02, HWHG08, ITR08, ISK00, Kor00a, Kor07, KKM00, Kru08, NSB08, Nik06, OAVB08, PP08, Ram05, WYL08b, WB09, YM04a, YM04b, YW03b, ZRS05, ZCM01a, Zha03b]. D.C [Ano03-35]. DAEs [HR05, Lam05, MAR05b]. daily [YZWY08]. Dale [Ano03-28, Ano03-62]. Dalheimer [Ano00-67, Ano00-68]. D’Aluisio [Ano01u]. dam [AHN06]. damage [JSK^{+}04, JSKE05, QNN03, SO03, SK06a]. Damij [Ano03-75]. damped [CC04b, FS04, LY09b, WWZ03, ZP01]. damping [IJB00, LA00, ZWY08]. dams [ZYY08]. Danny [Ano00-43]. Dantzig [LLL02]. Daphnia [Gou01]. Darcy [MS00]. Dario [Ano03-36]. Dasu [Ano03-31]. Data [Can03, GCN06, PLL03, SYHC06, ABH08, AB08c, All02, An03, Ano00-50, Ano03-31, Ano03-43, Ano03-73, BFFL08, BR03, CSPJ05, Dim08, DY09, DE06a, FDY^{+}08, FM08c, HW08, IKY06, JWC08, Kru08, LGH09, MRP08, MJ02, OTB^{+}08, ROY01, SMW04, SWK07, TS00a, TD07, TZ08, WK04, WL06, WB01, Wu03, Ano04e]. data-driven [DE06a]. Database [KS07, Ano00d, NK01, NT02, ZCCB08, Ano00v]. databases [CS00b, CKY07]. David [Ano00-30, Ano00-36, Ano00-57, Ano00-60, Ano00-61, Ano00-84, Ano01f, Ano01s, Ano03-39]. Davidson [Ano03-77, FJ05]. Davies [Ano00-29]. Davis [Ano03-55, Ano03-70]. Dawn [Ano03s]. DBI [Ano00-62]. DC [Ano00-32, Ano00-33, Ano00-35, Ano00-66, Ano03-58]. DCNNs [ZS08a]. dead [AOS07, SPW08]. dead-core [SPW08]. death [Ano00v, LP00]. debates [Ano00-84]. Debian [Ano00-39]. debris [De 04]. debugging [Ano01d, TY03]. decay [AU03, BS01b, KLJ09]. Decentralized [TLS^{+}08]. Decimations [BZ00]. decipherable [BSM01]. Decision [MM05a, PRA02, Che08b, Ehr02, Eke02, ESNP06, KKY06, LYL00a, LCL00, LLL08b, LY05a, MRB02, QLD08a, SJL08, TS00a, WS04a, WL07b, XZ08b]. decision-making [Che08b, WL07b]. Decisions [Ano03x]. decoding [Pan03, Sol00]. Decomposition [CZLN06, HS07, AV00, AK03, BHEKE07, BD04c, Chu05, Gra03, GS06, GL00, Han08, JYDH02, JDH03, JDY08, Jin09, JR01, KP05a, Kay04, KPP00, LHC08, LCHL06, Les02, Les05, LY08a, LW09c, LT01, LLL02, Mar09, MAK02, OLAS01, Pan00, SGM09, SSSH01, Sch01, SC02, Waz00, WS04b, YS04, ZW03a]. decompositions [Sau07, SB06]. decreasing [CDM01, GN09, KA07, Wan09b, ZC09b]. Dedication [CEG04, Dey00a, GS05, SY03a, dKLR09]. defect [Lee04]. deficiency [LGGSP06]. Defined [CD06, AIS08, BDA02, CK06, FH07, Gal02b, KDC08, Kum04, NR01, POS03, Pie09, Sav08a, Sya02, XA09]. definite [ES01, LHL07, MOP06, Sun00, Xie00, YH04, YTEC02]. definition [Ely04, XCS^{+}08, YL04a]. Definitions [CHL09]. definitive [Ano00-27, Ano00-34, Ano00-65]. Deflation [PV01]. deflections [LM03]. deformable [FM08a, FF07, ZS07]. deformation [Bat07]. deforming [HRM08]. Degenerate [ZW06, DDX04, DXZ05, EA03a, GL06, KLJ09, LST02, SC00a]. degradable [NS00]. degree [DLM08, DL08, GK04a, HYZ02, KA00a, LJ08a, PP08, TL07]. degree- [GK04a]. degrees [LW08b, Mos08, PCRR09]. Delay [BK04, BBI05, Li08a, OK06, PCE^{+}05, Par04b, TZ01, TaL09, WY06, AR09, AR10, AE07, BJ02, BT03, BBP09, CJ05, CZ03b, CN08b, CP01b, DJ08, DSS08, DH08c, DFY02, DTZF03, FW04a, GF06, Gou01, Han04, HPG06, HC01, HCLF01, Hua08a, Huo04, Huo05, JLT04, KÖ08b, Kar09, Kha09, KAKK00, Kot01, Kot04, KZD03, LA03, LDZW05, Li08b, LX00, Lin05, LB01, LZ02, LLD02, LG04a, LG04b, LXK09, LS01a, LS03, MBI08, MHS01, Mur05b, Pen02, PHW03, Pen09, PS03b, PP04, Pit01, QSL08, RC09, SA02, Son09, Tan02b, TZ04b, TX04, Tia08, WZ06, WZ07c, WG08b, XC06, XT01, XTX01, XC00, Yan04c, YX07, YH09, Yen08, Zha01b, Zha01c, Zha02b, ZL02, ZA03a, Zha04b, ZFYS05, ZS05, ZLY06, ZTH09, ZM02, ZZ00b, ZYH01, ZZ01c, ZZ03b, Zho08, ZHL04, Zhu08b]. Delay-dependent [Li08a, PCE^{+}05, Par04b, TaL09, YH09]. delay-differential [Par04b]. Delayed [HLY06, CÁHCJL^{+}09, ÇT05, CH05b, DW09, HZ08b, Kar05, Kha05, MXZL09, NTTS09, PLG^{+}09, Raz00, TaL09, WZZ09, XT08, Zha03a, ZSPZ08, Zha08c, ZW03b]. Delays [GY06, LZL06, CH04, CS04a, Che08a, Dix05, GSZ02, HHW08, Lev06, tLW08, Li08a, LYJ09, LH06, LC06c, LC07b, Mah09, PCE^{+}05, Pao01a, RC07, RRM08, SC00b, SMO03, TLS07, WZ00, Wan09e, XCL^{+}08, XT04b, Xu01, XZZ01, XLL03, Xu03a, XCD07, XLT09b, YH09, YL04c, ZSG04, ZZ01b, ZZ02a, Zha04b, ZT04, ZLH07, ZSZ06, ZZ05, ZLLL08]. Delbrück [Ano04d]. demand [Che01a, DCR07, KH02c, SZ06, WCL02]. demands [HLY08, KH02b, YC07b]. demographic [Qiu08]. demon [Hua05, HLCW07]. dendritic [ZDLH08]. denoising [CVW00]. denominator [Kov05]. dense [LCY05, TRD08, ZMM07]. densifying [MCZ00]. density [GL03b, Jia09]. denumerable [KL08a]. Departure [ET09, TT03]. depend [Jan07]. Dependable [Ano00w]. Dependence [Lo 02, JS03, YLZ08a]. dependencies [NK01]. Dependent [BFL06, AS09a, AHN06, BK04, GT00, GS07a, HJW05, HPG06, Jai05, JPV02, KLP09, KY07a, KY07b, Li08a, LDHW05, NS07, PCE^{+}05, Par04b, Pen09, Pet04, PD06, PJJP04, RB06, TaL09, WWW^{+}09, WZ08b, XC00, YH09, Zha09e]. depending [LZL04, SL07a]. Derek [Ano00-76, Ano00-77]. Derivation [JS04, BB09b, RZ03, Yav06]. derivative [Ban07, BD02, DGPT08, FAF09, Jum06, Mär05c, SL07a, WX03c]. derivatives [Ana04d, BCDF02, BX02, CSC^{+}09, DM07, Dra08, ED09, Hua04, LTY05, MM05b, Mur06, Mur08b, SMMM02, WF01, WXS01, XC07, YLB06, ZL08c]. derived [CCK03, CCKH07]. Deriving [FGX05]. Derrick [Ano00-71]. Dershowitz [Ano03m]. Descartes [Ano00-62]. Descent [CKY04, JT05, SH01, ZCK00]. DeSchreye [Ano00-43]. description [BBH07, LLS09b]. descriptor [YL04c]. Design [CCHK08, GHC09, IMPF08, LLY01, Ano03-69, LMA08, MIPT08, NK01, PS08b, SS09a, Sun05b, Sun05c, Ver08b, Ver08a, Ver12, WFTL09, WFT^{+}09, YL04c, ZWC^{+}09, Ano00x]. Designated [YXZL09]. designed [Gos00]. designing [ZW03a, Ano01k, Ano03y]. designs [Whi00, Ano00s, Ano00t]. Desktop [Ano00-78, Ano00-79, Ano00-36, Ano00-86, Ano00-91]. Detachment [Ano00-42]. Detectability [BHCØ07]. Detecting [KHS00, Ros03, HHR07, WSO08]. Detection [AS07, BIKL06, GHC09, HHC08, IDM08, JSK^{+}04, JSKE05, KKS04b, ZMLG09]. detector [ARS03]. deteriorating [HY06, HL08b, MM08a, MM08b]. deterioration [AH03, LHS08, Wan09b]. determinants [HL05]. Determination [CA09, ÖY07b, TB09, Tou05, HSM00, HS02, Ikh08, Lam05, MM08b, NT08, RV01]. determined [SS09c, ZLY09]. Determining [MPTV06, HY06, YLZG01]. Deterministic [BBGG08, BBM08, LC07a, AC00, LCL08, MS08, Ano00y]. detonator [BT03]. detours [HSG^{+}06]. Deuflhard [Ano03-67, Ano04p]. Development [Ano01k, EA03b, PM08, TZ08, Ano03-75, CPP07, Dra06, HW07, Ree00]. developments [AW01]. deviated [Jan07]. deviating [Can08, Don09, Jan09, JA03, Kop01, Liu08a, SL07a, Wan04b, Won08, ZT07]. deviation [XXLD04]. deviations [AA08, BX07]. DGFEM [WFS03]. diabetes [SYHC06]. diagnosis [SYHC06, YY07]. diagonal [CT09, HX03]. diagonalization [CFY08]. diagram [DMVR00]. diagrams [KKS04b]. dialogue [Ano00-64]. Dialogues [Ano03-89]. Diameter [YMTE05, MP05, She00]. dichotomy [HM01]. dielectric [SBD^{+}00]. diesel [HK08]. Dieter [Ano03-71]. Difference [ATT03a, AEÇ06, CD06, TZ01, Wan06b, AG01a, AZ01, AOW05, AD03, AO05, AY06, AO09, AK01, AP03, ACH01, Ayk04, BP03, BGNA07, COEP00, Cab01, COE03, CYG05, CYG06, CHS02, CP01a, CDM01, CDM03, CD05b, CZ03b, CS04b, Che03, CKZ01, CKR03, CK04, CCK00, CK02b, CH03, Coh00, CP01b, CD05c, DJ08, DS04, DN03, fDxZ11, DK04, DP07, DR01b, DH03, DDM08, EM09a, EM03, EM09c, ERRY01, FH03, FT03, Gla00, Gla03, Gla04, Gla05a, Gla05b, Gla06, Gla07a, Gla08, Gla09, GJM03, GNR05, GH03, GW01b, HJW05, HW09, Hil01, HLX07, HL03b, HM01, HSS01, IS04, IJB00, JV01, JKN09, JK09, JA03, JLT04, JT07a, JT07b, JZ03, KK09, KC02, KZ02, Kar09, Kha09, KKZ02, Kop01, Kra01, KLP01, LFZ00, Li01b, LA03, LZL04, LS04b, LZ04a]. difference [LZ05, LYW02, LLT08, Lin05, Lin06a, LZ02, LLD02, LG04a, LC04, LG04b, LY08d, LL08b, LXK09, LYG05, LPvST07, LS01a, Luo02, MHS01, Med01, Mic03, MPV02, Mur08a, Nic04, Pao01a, PS04, Pen02, PHW03, PS01, PS03b, PP04, Pit01, Pit03, Roe09, Rom01, Sah01, SMMM02, SW04, SL00a, Shi03, SDZ06, Sie03, Ste08, SS06, SX06, SBD^{+}00, TYP00, Tan02a, Tan02b, TL08, TA01, TRG01a, TRG01b, TMG03, TK07, TZ04b, TX04, Trz01, WZ00, WJA08, WG08a, Wei05, WA01, WX03b, cWcX09, XTX01, XT04b, Xu01, Xu03a, YW07, YTC08, YZZC09, ZZ00a, Zha01b, ZP01, ZCZ01, ZD01, Zha01c, ZZ01b, ZZ02a, ZSKZ03, Zha03b, ZW03a, ZA03a, ZL03, ZD03, ZT04, ZM04a, ZY05, ZYEZ07, Zha09e, ZZ00b, ZYH01, ZZ01c, Zho01, ZZ03b, ZC01, ZHL04, Zhu08b]. Differences [CD03, Fer03, Jun01, LY08a]. different [Hav08, fHCH09, HB04, Kon09, Sma01, VB01, ZL09a]. Differentiability [Sya01b, HR02]. differentiable [Kir08, LV03, YHT04]. Differential [Che00a, HM04, Liu06b, NCK^{+}08, NON^{+}08, Rom01, SW02b, WY06, WS04b, ZLJW06, ZCW06, AETEZ07a, AB08a, AVLPN04, AP08, AML02, AL04, AN09, ATY08, AB08b, AJ09a, AJ09b, AU09, AIS08, AE07, Ano00-66, Ano03-42, Ano03-59, Ano03-67, Ano04f, Ano04y, Ano04z, AO08a, Ash05, ACH01, AT02, BMHV09, BUP09a, BUP09b, BF03, BX07, Bai08, BA01, BM01, BM02, BJ02, BK04, BHEKE07, BN00, BHN01, BHNO04a, BGO04, BHNO04b, BGO05, BBRS08, BBI05, BBP09, BG08, BA09b, BA09c, CCP02, CJ05, ÇT05, CCJ08, Can08, CH04, CH00a, CP02, CK08, CBJ01, CQJ03, CL08a, CK02c, CJV07, CHR09, DI08, DKS07, DD04, DJL04, DSHS05, DK05, DSS08, DHM05, DZ08, Dix05, Don09, Dor01, Du08, DFY02, DTZF03, Dub08, EA03a, ES08a, FF02, FFK02, FW04a, Fen08]. differential [Fro08, Gal08c, Gan06, GL09b, Gar01, GL03a, GLC09, GF06, GL06, GJ07, GMG09, GS08, GH03, GS09, GMMB06, GL00, Guo01, Guo03, GZ06, HMV04, HM03a, HH04b, HPG06, Hil01, HYLW00, HRM08, HYZ02, HdLS03, HCLF01, HC02, HL04, Hua08a, HLY05, HLP03, JKN09, JJ01, JI04, Jan03, Jan04a, Jan04b, Jan04c, Jan07, Jan09, Jay04, JKP02a, JLW04, JL09a, JSF00, KGB07, KH00, KÖ08a, KSZ02, KLW09, KD08, Kot01, Kot04, KZD03, Lev06, LH00b, LA00, LZJ06, LTWM08, LS08, LFZ08, LY08c, LW09c, LG09a, LX00, LXH05, LTSW02, LTS03, LLNS05, LDG05, LB01, Liu04a, LLTL05, LY06a, Liu07a, LX07, Liu08b, Liu08c, LS01b, LD02, LS03, LSN05, LJ08b, LYL07, MAH08, MCH01, MH02, Man00, Man01, MP02a, Mär05c, McG03, MZ07, MQ06, MO07, ME08, Mur05b, Mus06, NRL00, NRL06, Öca07, OM09, OLS06]. differential [OZY05, OSSY04, ÖZ07, PG04, PPRG02, PGT08, Par04b, PG00, Pen09, PXW09, PB02, QZ05, QSL08, RSD09, RC09, Roe09, SNT08, SNG09, Sah01, SERG09, SALK02, ST00, SR03, SR04, SW02a, SL07a, SWL05, SYL05, SWX06, SOW04, SZ05, Swe07, SK09, TD07, TD09, Tel08, Tia08, TÇ04, TA04, TBG05, VR02, Van07, WJX04, Wan00d, WWZ03, WZ06, WLS06, WG07a, WF07, WLW07, WSL08, WFY09, WWS^{+}06, WKL06, WWC08, WWX04, WH05, WD05, WGZ06, WG08b, WW08d, WL09b, cWcX09, XT01, XLZ02, XWD04, Xu08b, XX09a, Yan01, Yan04c, YYZ03, Yan03e, Yan03c, Yan04d, Yen08, Yil08, YZZC09, YZWY08, YXZ09, Zen09, Zha02b, ZS04, ZFYS05, ZLY06, ZJLW06, ZL07a, ZTW09, ZWY08, ZM02, ZY08b, Ano00z]. differential-algebraic [MP02a]. Differential-difference [Rom01, Sah01, cWcX09, YZZC09]. differential-functional [BK04]. differentiation [BR03, CS05b]. diffpack [Ano04f]. diffraction [Ano03-44]. diffuse [YLB06]. Diffusion [SPE06, Wan06a, AHN06, AMS08, AM08a, AAD03, AVBW08, BKM08, BD04c, BGM^{+}02, CCP02, CN01, CAZAA00, CAZ01, CVW00, Che04f, Che00b, CJK09, CSPJ05, Das09, DRR09, DWL08, DH08c, fDxZ11, Dog01, Dog05, GAT07, GY08, JC00, JPV01, JP01, JPV02, KZ02, Kur08, LP08a, Les05, Li02, LY08a, LST02, LC06c, MC02, Mic03, MS04b, Mur07a, Mur08a, Pao01a, PL01c, PDRR08, Ram05, RR08, RRM08, SG07, SV06, SR03, SDDC08, STK08, SK06b, Syn08, VR02, Wan02b, Wan02c, WYW05, Wan06c, WKZ08, Wan09e, Waz07a, WFS03, WW02b, XCD07, YV05, Yil09, YWZ03, Zha09e, ZSZ06]. diffusion-controlled [AVBW08]. diffusion-reaction [Ram05]. diffusion-type [CCP02]. diffusionless [WY09b]. Diffusive [Cav00, CW08, Gou01, Huo05]. Diffusivity [CKK06, CM01b]. diffussion [VH06]. Digital [Sar06a, Li01c]. digitized [SM08]. digits [BDT04]. digraphs [HX06, HX07]. dilatant [ZMM07]. dilation [CC08b, Xie08]. dimension [BI01, CPZ00, JG04, KA07, Zen09]. Dimensional [CS05a, DS06a, Liu00, WZE^{+}06, AS09a, AK08a, Ano00-94, Ari07b, Ari09c, Asl11, BSS98, BMH08, BM09, BA09a, BB08b, CK03, CCC08, Che00b, CHM06, Deh02, DS07, Den07, DW05, EM09b, FL03, FP02, GY09, HJW05, HO00, HL03b, HL07, Jia01c, JX04, JT07a, JT07b, JDH03, JP01, KC02, KO09a, Kay04, KHN09, Kor00b, LWZ08, LY09a, Li01b, LTWM08, sLhYL09, LC04, Liu04c, LZS07, LY08d, LL02b, LL03c, MQ05, Ma09a, Mar05a, MM08d, PG04, Qua01, RRR07, Raz00, RR08, Roo08, SN02, SD08, SKG01, SYHC06, TGH02, TG03, TCX08, VH06, Wan03b, WG07b, WYL08a, Waz05d, WK03, WZ08c, XYC^{+}09, YZ02b, Yan03a, Zha07b]. dimensionless [MO08]. dimensions [AC04, AC05, CD05a, CAZ01, CY00, DMSV00, ER07, GS07a, KU01, LYW04, Ren03, SS07b]. Dimiter [Ano00w]. Diophantine [MZ04]. Dirac [QZ05]. Direct [Slo07, Ste06, BMHV09, dSM02]. direction [HLW07, LY09a, MLW09, WYH00, Xie08, YX09a, ZH09]. Directional [Yav05b, HM00, LTY05, WNS09, ZMLG09]. directions [CG01, JZT07, Wan02d, Ano03t]. Dirichlet [AOS07, JN07, Kru05, Kru08, LWY02, SNG06, Sta02, Wei00, WYZY01, Yan06a]. disambiguation [Ano04-32]. disasters [LL06b]. disc [Ung03]. disc-transformation [Ung03]. discipline [PdlF07]. Disconjugacy [CH03, LYG05]. discontinued [dZmClQ09]. discontinuities [KM07, MP02a, Sho09b]. discontinuity [BGAM09, UA06, VC03, Zha09a]. Discontinuous [ASL06, SW06, Abu01, AA05, COEP00, CH00a, Che00b, Dha06, GMŚ07, Kru08, Li02, RW03, RPO03, WL08c, Yao07, Zha09e]. Discounted [KKY06]. discounting [DKO03, MM08a]. Discovering [LJSH09]. discovery [YZWL09, Ano03-43]. Discrete [AO01c, AGO03a, AP04, And03, AM04, CDJR03, CCS08, Dib01, Hil01, IY06, IDOK06, JC00, LKH03, VJ04, XX09b, AHRT09, AM08a, AGO03b, ATT03b, AlS08, All05, ACCMJ07, AA01, AAD03, Ano04x, AC03, BD01, BCJ08, BD03, CÁHCJL^{+}09, CG08, CJO05, CF04, Che01b, Che04a, CZ05a, CWL07, Cse06, DCW08, DMSV00, DCS09, Dib03, EIR03, GPvZ06, GW01a, GM09b, GHW04, GSS04, GK03, HHR07, HT02, HZ03, Hor08, HC01, HZ08b, HSS01, JSF00, Kar05, KC05b, LXZ08, Ma07, MXZL09, MBI08, Mah09, MRT04, McG03, MC02, Mie01, Mur07b, OLAS01, PW00, RGTT01, RGTT03, SBPD08, Son09, SF07, SZJ09, Sun05b, Sun06a, TLLM08, TCX08, Ubh04, ÜÇT08, Wan00e, Wan01, Wan03b, WZ07b, WLWC08, WZZ09, cWcX09, Xu04, YJCZ08, YH03, ZSG04, ZCM01a, ZWCW00, ZL09b, Zho03, ZW03b, ZH04, Ano04g]. discrete-continuous [SF07]. discrete-delay [HC01]. Discrete-time [AM04, IDOK06, LKH03, AHRT09, CÁHCJL^{+}09, GM09b, GHW04, GSS04, Kar05, MBI08, Mah09, Sun05b, Sun06a, WZ07b, WZZ09, YH03, ZW03b, ZH04]. discrete-velocity [Mie01]. Discretization [Lia04, YH04, AS09a, Far01, HCL07, HS04b, JMR00]. discretizations [PD06]. discretized [AMV08, AU03, Far02, IY06, You04]. discretizing [Gar01]. discrimination [Lin08a]. discussion [WWH05]. discussions [Ano00-84]. disease [AVAGP08, GPAA^{+}09]. disfocality [LYG05]. disjunctive [Amm07]. disk [Ari09b, BC02, BC03, GM00, Ikh08, KS03a, KY00, Tur07, Yav05a, Yav05b, Yav05d, Yav07a]. disks [AK08a, AG09, Gal08a, LM03]. dispatch [LCL^{+}09]. dispersal [CCW00, Cui02]. dispersion [CGPB08, MKSB08, MS00, Roo08, SC00b]. dispersion-advection [CGPB08]. dispersive [VNG04, Yan03a, ZTW09, ZW09a]. displacement [De 02, LY08d, MB08]. displacements [AK08a]. dissection [RKS08]. dissipation [KLJ09, Man05]. Dissipative [All05, Dor01, GS08, HM03a, Phu02, SWX06, You04]. dissipative-type [GS08]. Dissipativity [Gan06]. dissolution [AVBW08]. Distance [Fro08, MUHdH03, Bax06, She02, She04b, YT00, ZZ09b]. distillation [HM03b]. distinct [PT00, YT00]. distress [CCF06]. distributed [Ano00-54, Can08, CLS07, DR04, Don09, HL02c, Hua05, HLCW07, LPRH05, LDZW05, tLW08, LHYM09, LC07b, Raz00, RB06, Sha09, Wan04b, WFC04a, WL06, YL04c, ZLLL08]. Distribution [Ata09, DA06, LC06a, Wu08b, Gil00, IWL04, IUCR06, Kaw06, Kum06, LWLJ09, Nab04, ND05b, ND05a, SS09d, Wan02e, YC07a]. Distributions [CD06, SNG06, ZHZT06, Ata09, GSS04, Kon09, Kum06, LWZ04, Nad05]. district [SKU01]. disturbance [BRC00, Kar05, Sun06a, Sun06b]. disturbance-rejection [Kar05]. disturbances [MM07a]. divergence [KC05a]. divergent [HZ09a]. diverse [Ano03f]. dividend [BCJ08]. Divisible [KM03b, MR06, MYR09, SMO03, SMOK05]. division [TS00b, Tan03b]. divisions [CG01]. divisor [KFMH06]. divisors [Coo05]. Dix [Ano01n]. DM [Ano00q, Ano00r, Ano00-48, Ano00-49, Ano00-47, Ano00-46, Ano01i, Ano01g, Ano01w]. DNA [Dar08, RM08, SK06a, Tan00, XXC^{+}09, ZWC^{+}09]. do [Ano00-90]. document [GZZ09]. Dollarization [Ano03z]. domain [AVO08, ADH^{+}03, AV00, BD04c, CSX01, Gra03, GL00, HJW05, HLZ09, ITT07, JZ08, JR01, KPP00, Kru03b, Kru05, Kru08, Lev06, LY08a, LK04, Lin08a, Liu00, MRMRM02, Sch01, WY07, WS04b, WS08c, Xu07a, YS04, ZW03a, Ano04q]. domains [CZ06, Den07, Din04b, HH02a, HH02b, JY08, Ren03, XM08]. dominance [CL01, HX03]. dominant [Yu04]. dominated [BCGP03, BMM^{+}06, Che04f, GY08, Ili04, PDBH03, Wan02c]. dominating [HHJS03, HLCW07]. domination [HX07]. Dominono [Gar00]. Donald [Ano04x]. done [Bat07]. Dordrecht [Ano00u, Ano00y]. Dortmund [Ano03-88]. dot [HW05c]. Double [BG07b, CWZZ07, GH03, CB02, LP08a, LF09, Lu09c, Moo02, NFI03, ÖY09, Sav08a, SH06, Waz05a, ZAAESG09, Zhu05, ZMM07]. double-sequence [Moo02]. double-step [Zhu05]. doubling [MM06]. Douglas [Ano00n, Ano04d]. Dover [Ano03p, Ano03q]. DPR1 [PKM^{+}06, PMR^{+}08]. Dr. [Ben01]. drain [DPU07]. drainage [DK09]. Drawing [Kor05]. DRBEM [YTEC02]. Driscoll [Ano03-66]. Drive [Ano04c]. Driven [DHY06, DE06a, GPvZ06, MRP08, WSG08, Zha03b]. DRM [FP02]. drop [Kor00b]. dry [Gla07b]. drying [CM06]. DSeSC [Pan05a]. Dual [Mal04, Ano03-68, CKL06, DD08]. dual-primal [CKL06]. dual-rate [DD08]. Duality [ZhYL05, CLT09, GHHD08, GA06, KT03b, Zal02, Zal03]. duals [TB09]. duct [Gla08, Gla09]. due [Ari07a, El 03, FAF09, Nik06, RB06, RB07, VSKF08]. Duelling [Ano03-27]. Duffing [Fan09, ÖY07b, PHAF09, WG06]. Duffing-harmonic [ÖY07b]. duopoly [LLL08b]. Durand [BGP04]. durations [Lin08b]. dusty [GGD00]. dV [Waz05c]. D’Yakonov [SSSH01]. Dynamic [DH08c, GL09b, HC09, KN04, KWT02, LTYB09, LCL^{+}09, Mig06, SDCL08, SZJ09, SZ06, YC07b, AR09, AR10, Ano03-79, Ano03-51, Ano04-27, AO09, CKCG03, CL07b, CWZZ07, CS02c, DM07, DE06a, Gal06a, HNR08, HT05, JEP08, KÖ08b, LW09b, MQ05, MH09, Nak06, SA03, Wan08a, WSZ02, WZ07c, Yav06, ZS05, ZLZ09, Zho08, SH02]. Dynamic- [SH02]. Dynamical [LYJ09, Qiu08, CZ05a, DWL07, Dog05, Don04, Hau06, HHRA02, LBW09, Noo02b, SNA09, TaL09, XCL^{+}08, ZL09a, Zho03]. Dynamics [DDM08, GPAA^{+}09, WW04b, Zha09f, ZH04, AACLM08, AlS08, Ano03-54, Ano03-65, Ano04h, BBCE09, BRC00, CSC^{+}09, JSGP08, KLMR03, LP04, Lo 02, MRT04, Qia08, Raz00, RZH03, SA02, SASHS09, TL07, XZZ01, YK00, Ano03r]. dynamo [RF09]. dysarthria [KT09]. dysfunctions [ZCCB08]. Dziok [PMS07].

e-marketplace [KC03b]. each [Ano00-88, Ano00-89]. early [Ano00l, Ano00m, GSS04, IDM08]. earning [DAKO06]. Earth [HWHG08]. easily [Wei00]. easily-generated [Wei00]. Eastaway [Ano00-90]. eave [KOV08]. ECG [Lin08a]. Eckhaus [DS05a]. Eckstein [Ver08b]. ecological [GAT07, XC06]. Econometrics [Ano03-28]. Economic [Ano03-28, ZLS06, Ano00-30, JZ03, MGG05, YE01, Ano03-29]. economic-lot-size [MGG05]. economics [Ano01o, Ano03j, Ano03-80]. economy [Nak06, Ano03-29, Ano03-60]. eddy [DMP03, KK06a, Lun03, dR03]. Edelman [Ano01s]. Edge [MV08b, Ano00p, BIKL06, CCP09, NSB08, Pla07, TYS03, ZMLG09]. edge-based [CCP09]. edible [FM08a]. Edited [Ano00e, Ano00f, Ano00q, Ano00r, Ano00w, Ano00-76, Ano00-77, Ano00-43, Ano00-50, Ano00-54, Ano00-55, Ano00-56, Ano00-66, Ano00-80, Ano00-71, Ano00-81, Ano00-85, Ano00-84, Ano01e, Ano01h, Ano01y, Ano01z, Ano01o, Ano01q, Ano01r, Ano01x, Ano03-36, Ano03-77, Ano03k, Ano03t, Ano03-78, Ano03o, Ano03z, Ano03-29, Ano03-80, Ano03-81, Ano03-35, Ano03-32, Ano03-33, Ano03-38, Ano03-41, Ano03-43, Ano03-55, Ano03-57, Ano03-61, Ano03-62, Ano03-70, Ano03-71, Ano03-74, Ano04f, Ano04k, Ano04-28, Ano04x, Ano04-30, Ano04-29, Ano00j, Ano00k]. Edition [Ano00-46, Ano03h, Ano03p, Ano03r, Ano03-39, Ano03-85, Ano03-45, Ano03-46, Ano03-84, Ano03-56, Ano03-59, Ano04j, Ano04z, Ano00i, Ano00p, Ano00-51, Ano00-59, Ano00-67, Ano00-68, Ano00-69, Ano00-86, Ano00-89, Ano01w, Ano04e, Ano04f, Ano04i, Ano04n, Ano04m, Ano04p, Ano04r]. Editorial [Ano01-29, Ano02e, Ano02f, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano00-95, Ano00-96, Ano03-91, Ano03-92, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano05d, Ano05e, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09v]. Eduardo [Ano03z]. education [Ano03-35]. educational [Ano00-32, Ano00-33]. Edward [Ano03m]. Effect [ICM09, KKS^{+}04a, Wan07b, CCW00, Cui02, HHK06, HAKA07, HNSA09, HSS01, KY07b, Mak09, MCC09, Sar03, SB03, SMO03, TRD08, TG03, Tur07, WWW^{+}09, WH05, YWC08]. Effective [GL08b, PST08, Tan04, EA03a, FGX05, HW05c, SNG09, SHS05, STH08]. effector [Yav00]. Effects [Bat07, Lee09, APS07, GMS08, IMLF09, JD07, LZJ06, MB08, MM07b, MS00, SBPD08, SDCL08, XLL04, ZJLW06]. Efficiency [Zal02, NSQ02, Pos08]. Efficient [BMH08, MP02a, PSS09, RRR07, RDS00, Sch01, ZY07, BCJ08, Bog09, CP04, CZ09b, DF06, HNR08, HWX07, HRP04, HY00, JK01, JCCT09, KS03a, LLS^{+}09a, LCY05, MAK03, MQ06, NMD07, OM09, RZH03, SS09a, SNA09, Tan04, dSOGRMRL04]. EHW [DM09]. Eigen [PMR^{+}08]. Eigen-solving [PMR^{+}08]. Eigenfrequency [RV01]. eigenfunctions [ACCMJ07, VJ04]. eigenmodes [PD04]. eigenpair [Xie00]. eigenproblems [Jia01a, JZ02]. Eigenstructure [VJ04]. Eigenvalue [PD06, SB03, ZL08c, BS03b, HW05b, Hof03, IN07, JSK^{+}04, Kar01, LHC08, LUZ05, PSES07, SB00, YLL08]. Eigenvalues [WS08c, ZQC08, DMSV00, GS04, HHR07, Hu08, LR03, LH05a, LYC07, PV01, RSR00, Wu08b, XZ09a, YFT04]. eigenvectors [Kor05, YFT04]. eight [Pla07]. eight-tetrahedra [Pla07]. eikonal [QS03]. Einstein [Åhl02, Ung00, Ung05, Ung07, ZZ09c]. EinSum [Åhl03]. Eiter [Ano01n]. either [KT03a]. El-Mistikawy [Ari09c]. Elastic [AC09, LLS^{+}09c, AK08a, CO07, CA09, DH08a, DD00, GGD00, GAA06, HLY08, HGS07, IN07, KL01, MG03, RB06, Ste07, YLB06]. elasticity [Ara01, BCS03, CK03, JKR08, Liu00, RB06, RB07]. elasto [NLR00]. elasto-plastic [NLR00]. elastoplastic [HD08]. elastostatics [Mar05a]. election [Bel05b]. elective [Ano00-72, Ano00-73]. Electoral [Bel05a]. electric [MH09]. electrical [KPSW06]. electrically [HM05a, Mos08]. electrokinetically [WSG08]. electromagnetic [ISK00, Tan07, ZRS05]. Electron [Gre01, TG03]. Electronic [Ano03-88, Ano00-76, Ano00-77, Jua03]. Electroosmotic [DHY06]. electrorheological [ER07]. electrostatic [KL05]. Element [ASL06, SK06c, AS03a, ABBD01, ABBDP04, AM09, Ano03-84, Ban00, BT03, BJBL03, Bed02, BCS03, Bla04, Bla07, BG04b, BH01, BKK08, BLS05, CJL04, Che00a, Che04f, Che04a, Che00b, Coo05, DW08a, DMVR00, EP04, FK01, FR04, GY08, Gir03, He02, HL06, HGS07, IN07, KK02, KPP00, KLS06, KP06, Li00a, LL02a, Li02, LW09d, Lia05, NS07, NSB08, OP01, OLS06, OIL08, OIL09, PLL03, PDBH03, SN01, SMH08, Ses03]. element-finite [NS07]. elementary [He09b]. elements [Ano03-34, AMV08, HM00, KK06b, MS04a, Pet04, SN02, SB06, TV07, Wu00]. eleven [ZL07a]. Eliasmith [Ano03-54]. elimination [Lee09]. Elisabeth [Ano03-57].

Ellen [Ano00-78, Ano00-79, Ano03-76]. elliptic [ACM05, AK08c, AMV08, Bou03, CL07a, CK02a, CGGC02, EMM09, FK01, FFK02, HLC05, HZ07, Kan09, KH00, KF05b, KR07a, KP06, LF03, LW04, Li01d, LP04, LHLD08, MPS04, Pao03, PW05, PD04, Ren03, RW03, SERG09, Sir00, Sma01, TLLM08, Wac00, WLS01, WKL06, WS08c, Xu08b, YZZC09, ZL04, ZC09b, ZH02]. Ellsworth [Ano00-42]. Elsevier [Ano00-85]. email [Ano00-60, Ano00-61]. Embedding [LTTH09, KM01, ZZ09b]. embeddings [DBX04]. emerge [AA08]. emphasize [NKCS08]. Empirical [Ano03-78, Jah04, Ano03v, Cho08, Hau06]. employing [WF01]. EMQ [LCT08]. emulation [Yan03b]. enclosure [MO08]. encoding [Pan03, Sol00, XXC^{+}09, ZWC^{+}09]. encoding/decoding [Pan03, Sol00]. encountered [PDBH03]. encyclopedia [Ano01v]. end [FG00, Yav00]. end-effector [Yav00]. endomorphisms [BS01b]. Energetic [VSKF08]. energetics [TYS03]. Energy [HW05c, JCCT09, KLJ09, ÖY07b, TXC02, Tan07, WH08a, Zha09b, Ano03-44]. Engine [SGR^{+}09]. engineering [Ano00-28, Ano00-38, Ano01x, Ano03-85, Ano03-54, KL09, LMA08, LKL08, PDBH03, Ano04w, Ano00-55, Ano04v]. engineers [Ano00g, Ano00h]. engines [Nam01a]. enhanced [HRM08, YZWY08]. Enigma [COS09, Cse06]. Enlarged [Ano03p]. enlightened [Ano00-80, Ano00-81]. enough [Ano03-58]. Enriched [YLB06]. enrollments [LC07a]. ensembles [AMYS^{+}09]. ENSO [HX09]. enterprise [Ano00-36, Ano00-92, Ano00-93, Ano03-81]. Entire [ZCL08, CZLC08, Fan02b, Leu00, Zha08b]. entrance [XLL08]. entropic [Han08]. Entropy [AU03, CW06, MRdCV00, Bal00, DSS00, Dra03a, Pra07, WYW05, YW08c]. Entropy- [MRdCV00]. enumeration [HHS07]. envelopment [TS00a]. environment [LLL08b, PL01b, WM04, WI08, XC00, YYNK03]. environmental [Fer03]. environments [Bra08, Col04]. enzymatic [BBM08]. EOQ [MM08b, PKM08]. EP [LWZ04]. epidemic [HZ05, LGZ01, NTTS09, PLG^{+}09, Qiu08, TL07]. epidemiological [AMC08, DLB03]. epiderivatives [CLT09]. epsilon [RZ03]. equal [YB07b]. equal-width [YB07b]. equalization [TLP^{+}09]. Equation [Liu06a, Ras06, SW02b, Ste06, WY06, dM06, AETEZ07b, ACCMJ07, ACM05, AM08b, AO08a, AK06, AS08, AV00, Bab09, BUP09a, BP09, BM01, BM02, BM04b, BPP04, BB06, BS00b, BRC00, BGM^{+}02, Cav00, CK08, CAZ01, CQT02, CZ03b, Che04b, CCK00, CC04b, Col02, CGPB08, CSPJ05, CI06, DCW08, DK09, Das09, DR01a, DJL04, Deh02, DS07, DSS08, Dix05, Dog05, Dor01, DP07, DDM08, Dua09, ES01, ESR06, FLfFL08, FJ05, Fen08, FQZ02, FFXL05, Gal08c, GAT07, GT00, GF06, GMG09, GC08, GS09, GHJ06, GK04b, HH02a, HJW05, HM03a, He09a, Hem08, HPG06, HO00, Hor01, HL06, HLX07, HHL08, Ibr09, Iva07, IJB00, JKN09, JG07, JKP02a, JG04, JL09a, JP01, Kan04, KLJ09, KZ02, KL05, Kar03, Kay04, Kha05, KT08, Kha09, KD08, KÖ09b, Kru03b]. equation [KLP01, KX05, LWZ08, Lee09, Lev06, LX03, LZ05, LWY06, LCGX06, Li08b, LY08c, LL05, LLT08, LWY02, LWY04, LLNS05, LG04a, LX07, Liu08a, Liu08e, LXK09, LGY09, LYG05, Lu09b, MYG09, MH02, Man00, MV08a, MZ04, MHS01, MC02, Mie01, MPV02, MRdCV00, ME03, MO01, NSB08, OLS06, ÖY07a, PGT08, PKP09, Pen09, PKSP08, Phu02, Pit01, Pit03, Pol09, QFZ03, RC09, Roo08, RZ03, SSH09, SD09, SGM09, SALK02, SD08, SK04, SM09, SDDC08, SDB09, SW02a, SL07a, SV08, Sim01, Slo09, Ste08, SOW04, SX06, Tao09, TBDL09, TGH02, TG03, UHuI09, UK08, Wac08, WT00, WC04, WYW05, WLS06, WG07a, WY07, WKZ08, mWmZ09, Waz04, Waz05a, Waz05b, Waz05d, WX00a, WZ08c, WHD09, cW09b, Xu02a, Xu07a, XYC^{+}09, YZ02b, Yan04c, Ye02, YM09, YTCF08, YD08, YB07a, YB07b, Zha01a, ZZ01a, ZW03a, ZW04, Zha07b, ZYEZ07]. equation [Zha07a, ZTW09, ZXlD09, ZW09a, ZQ00, ZHKW05, ZT07, ZZD08]. Equations [AO01c, GC06, GCN06, LCL06, Liu06b, PXXW06, RL08, SBS06, Sat06, TZ01, Wan06b, WG06, XWL06, ZLJW06, ZCW06, AETEZ07a, AB08a, AH00, AP08, AM08a, AR09, AR10, AG01a, AZ01, AML02, ATT03a, AGO03a, AGO03b, ATT03b, AL04, AOW05, AOTW07, AAI05, AN09, ATY08, AB08b, AJ09a, AJ09b, AU09, AE07, AL08a, AL08b, AAD03, AE08, Ano00z, Ano00-66, Ano03-59, Ano04f, ABB05, AU03, Ash05, AO05, AY06, AO09, Asl11, AK01, AP03, ACH01, AT02, Ayk04, BMHV09, BKM08, Bai00, BF03, BX07, Bai08, BPA03, BSK04, BM04a, BAK09, BS01a, BCRS05, BJ02, BD03, BNH07, BP03, BHEKE07, BB09a, BB09b, BHN01, BHNO04a, BBRS08, BGNA07, BBI05, BBP09, BPM09, BG07a, BG08, BBA09, BA09a, BA09b, BA09c, BG11, Bil05, BG04b, BPSW00]. equations [BOP07, BH01, BB08b, COEP00, CCP02, COE03, CJ05, CYG05, CYG06, CL07a, ÇT05, Can08, CH04, CH00a, CP01a, CP02, CDM01, CDM03, CPC00, CKN04, CAZAA00, CHY01, CQJ03, Che04e, CP04, CS04b, Che04d, CWZZ07, CWLC07, CO07, CL08a, CZ09b, CyL08, CKZ01, CLS07, Chu07, CJK09, CK02b, CH03, CNPP08, Con03, CK02c, CJV07, CCZK09, CN08b, CD05c, CHR09, DI08, DKS07, DGR04, DJ08, DH08b, DK05, Den07, Deu01, DN03, Dib01, Dib03, DL05, DHM05, DZ08, fDxZ11, Dog01, Don09, DK04, DR01b, DH03, DM07, Du08, DFY02, Dub08, EA03a, EM09a, EM03, EM09c, El 03, ERRY01, Fan02a, FGX05, FFK02, FW04a, FP02, FH03, Fro08, FT03, GK02, Gan06, GTJ07, GGA09, GL09b, Gar01, GL03a, GS01, GLC09, GSN08, Gir03, Gla00, Gla03, Gla05a, Gla06, Gla09, GJ07]. equations [GMS09, GS08, GJM03, GH03, GW01b, GL00, Guo01, Guo02, Guo03, GZ06, Gyi00, GJS01, Had03, Hav08, HH04b, HL09a, HM04, HT05, HZZ02, HYLW00, HRM08, HdLS03, HHRA02, HCLF01, HC02, Hua04, Hua08a, HM01, HLY05, HLP03, IP07, ITT07, JJ01, JI04, Jan03, Jan04a, Jan04b, Jan04c, Jan07, JK09, Jay04, JKP02b, Jeo04, JEP08, JA03, JLT04, JT07b, JZ08, JDH03, JDY08, JLW04, JZ03, KGB07, KH00, Kan02, KÖ08a, KÖ08b, Kar09, KLS06, KP06, KSZ02, KKZ02, KLW09, Kop01, Kot01, Kot04, KZD03, Kra01, Kru03a, KMK09, Kur08, LT05, LYW04, yL09c, LKS^{+}06, LC06b, Leu00, LH00b, LA00, LFZ00, Li01d, LL02a, LA03, LZL04, LZ04a, LMHW05, LZJ06, LTWM08, LY08a, LS08, mLpW08, LFZ08, tLjW08, LW09c, LSL09, Lia04, LX00, LX04, LXH05, LTSW02, LTS03, Lin05, LDG05, Lin06a, Lin08c, LB01, LZ02]. equations [LLD02, LK03, LWG04, LG04b, LDHW05, LLTL05, Liu07a, Liu08b, LN08, LL08b, LN07, LST02, LS01a, LS01b, Luo02, LD02, LS03, LSN05, Luo06, LJ08b, LYL07, MAH08, MCH01, Mal04, Man01, Mar07, ME00, Mär05c, McG03, Med01, MZ07, MLW09, MVS^{+}07, MQ06, MO07, MN01, MCZ00, MRMRM02, MS04b, Mur08a, Mur05b, Mur07b, Mus06, NOY01, Nic04, yNyLjZaZ08, NRL00, NRL06, NRL08, NW09, NP07, Öca07, OM09, OI00, OR01, OZY05, OSSY04, ÖZ07, PPRG02, PSS09, Pao01a, Pao01b, Par04a, PS04, PKT07, PG00, Pen02, PHW03, PXW09, PS01, PS03b, PP04, PHAF09, PR07, PB02, QS03, QSL08, RD07, RR09b, RF09, RSD09, RGTT01, Ren03, Roe09, Rom01, SDVS09, SNT08, SNG09, SG07, Sah01, SERG09, SMMM02, ST00, SW04, SHY08, SR03, SR04, Sha04, SL00a, Shi03, Sid08, Sie03, SWL05]. equations [SYL05, SWX06, SS07b, SSD07, SZ05, SS06, Swe07, SAB07, SNA09, SK09, TKC02, TYP00, Tan02a, Tan02b, Tan07, TL08, TD07, TD09, Tel08, TA01, TRG01a, TRG01b, TMG03, TZ04b, TX04, Tia08, TÇ04, TA04, TBG05, Tou05, Trz01, VR02, Van07, WJX04, Wan00e, WZG00, Wan00d, Wan02c, WW03, WWZ03, Wan04b, Wan05a, Wan05b, WZ06, WLW07, WLWC08, WSL08, Wan08a, WYL08a, WG08a, WX09b, WFY09, Wan09a, Waz05c, Waz07b, Waz07a, Wei05, WY09b, WZ09, WKL06, WA01, WX03b, WS05, WWC08, WF01, WSZ02, WS04b, WWX04, WGZ06, WZ07c, WG08b, cWcX09, XH02, XT01, XTX01, XLZ02, XT04b, XDD09, XWZ04, Xu01, Xu03a, Xu07b, Xu08b, XX09a, Yan01, Yan03a, YYZ03, Yan03c, Yan04d, YW07, YTC08, lYsT09, Yen08, Yil08, YZZC09, You04, YLD09, YL08c, YXZ09, YB08, Yus09, Zen09, ZYG08, ZZ00a, ZP01, ZCZ01, ZD01]. equations [ZZ01b, Zha02b, ZZ02a, ZL02, ZSKZ03, ZA03a, ZL03, ZT04, ZS04, ZL04, ZM04a, ZFYS05, ZY05, ZS05, ZLY06, ZJLW06, Zha07b, ZTW09, ZWY08, ZC09b, ZM02, ZYCI09, ZCK00, ZZ00b, ZZ01c, Zho01, ZH02, ZZ03b, ZLD08, ZY08b, Zho08, ZC01, ZHL04, Zhu08b, ZCC04, dG09, HM05b, Ano03-42, Ano03-67, Ano04y, Ano04z]. equator [LC02]. equilateral [Köh00]. Equilibria [Wu01, AMR05, Din00b, DPJ03, JPS09, KY02, KWT02, NP07, SA03]. Equilibrium [ZLS06, Ano04-27, CGY09b, CKY04, Din00c, DP02, FHH08, HL09b, KK08a, KL02, LF02, Mou02, NT08, NSQ02, PY09, SWL05, Wan02a, Wu00, ZL09a]. equitability [BW06a]. equivalence [BKK08, CWLC07, CKR03, CK04, WHL^{+}09]. Equivalency [HN07]. Equivalent [AGS09, Dub08, ZCK05, CZ05a, SMOK05, YL05]. Eran [Ano04-28]. Erdik [NLL08]. ergodicity [Gu08]. Eric [Ano00-75, Ano00-34]. Erlang [WCL02].

Errata [Ano05f, Ano05g, Eng04]. Erratum [Ata09, BM02, ID10, KÖ08a, KÖ08b, Liu09a, SW02b, SW08]. Error [Ale06, Bed02, DGG^{+}04, Has02, HHC08, OLS06, PPP^{+}06, Uje04a, WYL08b, WX03c, Xia07, AB08c, AN05, AN06, Bla04, Bou03, BMM^{+}06, CK03, CD05a, CBJ01, CGBJR03, El 03, FFXL05, Gal01, Had03, HL03a, Ikh08, JMR00, KKS^{+}04a, Koi07, Liu08d, OP01, PMQR09, PJJP04, Ree00, RW03, RPO03, Uje07, VH06, WWS^{+}06, Xu02a]. error-free [PMQR09]. Errors [GSZ02, yL06, TS06, Bar01a, CZG04, DW02a, HCLJ00, Hua00, HGH02, HF04, KK01, yL08b, LKC04, MT09, PW06, dG09]. errors-in-variables [Bar01a]. Ervin [Ano04c]. Esko [Ano00-47]. essay [Ano00-42]. Essential [Ano03-30]. essentials [Ano00-52, Ano00-53, Ano04o]. Establishing [GYPL08]. estimate [Bla04, Bou03, CD05a, DKO03, Gal01, ZL08a, ZZ01c]. Estimates [BS01b, Gar01, BK04, Bed02, CK03, CS00a, FFXL05, GA07, Had03, ID05, ID08, KLJ09, OLS06, RW03, SM02, VH06, XWD04]. Estimation [Ale06, APV02, BPvZ03, CKK06, Hua08b, KDKS01, MW00, SK06c, AN06, BGA08, CÁHCJL^{+}09, Cha06, CKKK06, Col02, CM06, DAKO06, El 03, FM06, GL03b, HSG^{+}06, LP08b, LG09b, OP01, She00, SL08, Wu08b]. Estimations [HL05]. estimator [Bar01a]. estimators [BB01, Elk06]. Ethernet [Ano00-27]. Euclidean [Bax06, Cip08, Gor00]. Euler [AC09, BOP07, DMP01, Far01, Gla03, Gla06, Gla09, GS09, KR07b, Kot01, LHC08, LS06b, MV08a, Mat03, OS08b, PV00, WJW08, Xu06, Yu04]. Euler-Grüss [Mat03]. Euler-like [PV00]. Eulerian [Gir06, VTR04, WZE^{+}06]. EUR [Ano03w, Ano03-37, Ano03-82, Ano03-48, Ano03e, Ano03-69, Ano04i, Ano04k, Ano04p, Ano04v, Ano04z]. EUR49.95 [Ano04j]. euro [Ano00-85]. Europe [Ano00-80, Ano00-81]. European [LPRH05, LPvST07]. evader [She04b]. Evaluate [Wu04, Bar01d, CH05a, Lin04a]. Evaluating [HW08, Wac00]. Evaluation [Hua06, Tan00, XYC^{+}09, AHRT09, BY03, DF06, GYPL08, GYPL09, Has02, HT04, Isk08, ID10, JG04, JW05, KPD^{+}09, Lin06b, Mur06, She05, TBR01, WL06, YMI06]. evaluations [WX03c]. evaporation [Kor00b]. Even [DM07, SC03, WWW08]. Even-order [DM07, WWW08]. Event [ST00]. eventual [FM08b]. eventually [WZ00]. everyday [Ano00-90]. evidence [BCGP03]. Evolution [Ano01u, DD03, Mig06, PXXW06, Sat06, AH00, AL08a, Asl11, BPM05, BB09a, Bil05, BB08b, DGR04, FS07, GTJ07, GJS01, HM01, IDM08, JKP02b, SSLL09, WW03, YZWY08]. Evolutionary [Bib06, LBZ^{+}09, CKCG03, CSZH09, DM09, GLTS08, KLZZ09, KX05, LRL09, LHW04, RTS07, XXC^{+}09, Ano04h]. Ewald [Ano01l]. Ewing [Ano00-88, Ano00-89]. Exact [AETEZ07a, AH00, CDJ04, DCW08, Fly06, GSN08, HH02b, Hay06, JP01, KF05a, PSM02, SG07, SW07, Slo09, SK09, TEB06, Waz05a, WZ09, XLZ02, ZXlD09, Bae07, BCA09, BRC00, CS01, FMFV08, GM00, GHJ06, KHN09, LHT07, McG03, Phu02, Sah01, SHY08, SK04, SDB09, VNFF08, WX00a, cWcX09, Yan03a, lYsT09, YB08, ZTW09]. exam [Ano00-52, Ano00-53, Ano00-72, Ano00-73, Ano00-92, Ano00-93]. example [KM01, XCS^{+}08]. Exchange [CJ08, Ano00-44, Ano00-45]. excitation [BB04]. excitations [ABS09]. exclusion [CF04]. exhaust [CS08]. exhibiting [NVAR03]. Existence [AN09, AAD03, AC03, Ayk04, BF03, BG04a, BPA03, BSK04, BM01, BM02, BM04a, BV05, BN00, BW05, COEP00, CCP02, CYG05, CL07b, CH05b, CZ08, CZ09c, Den07, Din00b, Din00c, DY05, DYZ08, EMM09, ER07, Fen08, GL09a, GY06, Guo01, GLZ00, HYZ06, HG03, HW09, HT02, HY03, HPG06, HYLW00, HLY06, HZ09b, Huo04, JS03, Jan04a, Jan07, Jan09, JA03, KK08a, KL08b, LP08c, LP08b, LW05, LJL06, LWG04, Liu04a, LY05b, LDK08, Liu08a, MDG05, MZ09, NRL00, PZ07, SWX06, Sun05a, SL07b, WG07a, WG07b, WW08b, WA08, Wen00, WCW08, WWC08, WWYL09, WLW08b, WL08c, Yan04d, Yan05b, Yan07, YW08a, Yao04, YWYL07, Zen09, ZFYS05, ZLS06, ZLH07, ZS08a, ZYW08, Zha02c, ZZ03b, ZHKW05, BHN01, CH01, CK02c, CHR09, DSS08, FW04a, Gin02, GS08, GN09, HZZ02, Ibr09, KL02, KY02, LDZW05]. existence [LZJ06, tLW08, mLpW08, MRMRM02, OEV09, Son00, Sot01, WJX04, Wu00, WLZ07, Zha01a, ZJLW06, HM05b]. exogenous [Sun06a, Sun06b]. Exp [Asl11, BB09a, BA09a, BB08b, GGA09, KHN09, KÖ09b, SDB09, mWmZ09, WZ09, WH07, XYC^{+}09, ZTW09]. Exp-function [Asl11, BB09a, BA09a, BB08b, GGA09, KHN09, KÖ09b, SDB09, mWmZ09, WZ09, WH07, XYC^{+}09, ZTW09]. expanding [HCTC03]. expansion [CM09, HHC08, LDK^{+}06, SERG09, VJ04]. expansions [AC06, CGBJR03, CC02, Sri03]. expectation [ABCD05, BDA02]. expectations [LH05b]. expected [KNP06, XZ08b, XTZ08]. Experience [YX09b]. Experience-oriented [YX09b]. experiences [Ara01]. experiencing [LM03]. experiment [Ano00-29]. Experimental [OAVB08, PM08]. Experimentation [Ano00-28]. Experiments [Pet01, CL02a, Jon07, Kur08]. experimentsregarding [DR04]. expert [LRLM^{+}01]. experts [Ano01x]. Explicit [AJ09a, DSS08, Gra03, JPV02, LY08a, LS06a, OS09, Trz01, Tsi02, Tsi03, Wan07a, WWS^{+}06, WBSP08, lYsT09, AM08a, AJ09b, GGA09, GW03a, JMR00, LDHW05, LHT07, Lun03, SSSH01, SV06, SH01, Van07, mWmZ09, WX00a, WX00b, YS04]. explicit-implicit [YS04]. Explicit/implicit [LY08a]. exploit [MRÁR08]. exploration [Ano00g, Ano00h]. explorations [Ano00-32, Ano00-33]. Exploratory [Ano03-31]. Exploring [Ano00-66, Ano03-32, CL00, Ano00-29]. Exponential [AK01, GHW04, HM01, TZ04b, WFS03, WH05, ZSZ06, ZC01, Has02, HHW08, KL03, tLW08, LC07b, WWW^{+}09, XLT09b, ZLH07, ZZ05, ZLLL08]. Exponentially [KP03, HL08b, KK09, KS03b, Wan02c, Yu04, ZL07c]. exponentially-fitted [KS03b]. exponentials [BCR04]. Exponentiated [Nad05]. exponents [Kan09, LWZ08, LC06a]. Expression [WW00, BM00, FM08c, LP08c]. expressions [ZQC08]. Extended [Ari09a, CAZAA00, Che04a, Sau07, WD08, Ano01w, CAZ01, Che00a, DD03, DCW08, DD08, HCG08, HX06, HX07, YK00, Ano03-45]. Extenics [Hua06]. Extension [AFI06, BC03, BA09a, Cha08, HLCL00, CX04, CL09, JDY08, LCL00, MM05b, NR04, OS08b, Wu03]. extensions [Ano00-70, CCS00]. extensive [Ano04h]. exterior [Den07, Kru05, Kru08, LWY02, LWY04]. external [ABS09]. extinction [Bra08, CF04, LY08b]. extracting [WDLP04]. Extraction [LH00a, Gre04, LLS09b, QLD08a, WDLP08]. extragradient [BXFZ09, PY09, WXW01]. Extrapolating [Tsa03]. Extrapolation [WK03, SS09b, Sid08, Ikh04]. Extremal [Cab01, LY09c, MSS08]. Extremality [CH00a]. extreme [Elk06, HHR07, KKS04b, STV08]. extreme-value [Elk06]. extremes [EZ08]. EYR [Ano04o].

F [Ano00-30, Ano00-71]. facilities [PS08b]. Facility [WI08, Che01a]. Fackler [Ano03j].

Factor [AS07, LYL00a, LCYL00, LCL00, WK04, CC08b, LYL00b, MD06, Sav07b, Xie08]. Factor-2 [AS07]. factoring [Tan03b]. factorisation [KM03a]. Factorization [RTW05, JV01, Més05, Pan07a, WH09]. factorizations [Gal05, SV09]. Factors [Sav07a, LYL00a, Sav08b, SL09]. Facts [CJV07, KLMR03]. Faedo [BAK09]. Fahle [Ano03-61]. failed [Lin06b]. Failure [KI04, SY05]. failures [WZ07b]. fairness [CZ05b]. Faith [Ano01u]. FALSCAL [CSLC07]. Fame [Ano03-35]. families [AHS00, ASS08, BL08, CjBN09, ĆRRU08, GYS05, HdLS03, LTS03, YZ02b, Yan03a, YB08]. Family [TA06, YI06, AÖS04, AB05, BFM04, CWY08, CKQ08, Gu07, NSQ02, OS09, PIR04, PR06, Sha04, TCH07, Wan07a, XHZ05, YTC08, YYZ07, Ano04-29]. Fan [DBEGB08, LH05a]. far [YZ02a]. Farid [Ano03v]. Farlye [Ano00-36]. Farquhar [Ano00-57]. Fast [JG04, RSR01, Tun05, WS00, AC04, MRT05, PKSP08, YY07, YW03b, RR03b]. fastened [AS03a]. Fatma [Ano01n]. Fault [DM09, FL05, GHC09, SMS^{+}04]. fault-tolerant [FL05, SMS^{+}04]. Faust [Ano00g, Ano00h]. Favard [Gal08a]. FBP [HL08c]. FCM [LCL08]. FCM-based [LCL08]. FD [CC04b]. feasibility [Wan02a]. feasible [JT05, JZT07, bJQxC09]. FEAST [TBR01]. feature [BFFL08, Can03, OTB^{+}08]. features [Ano00-88, Lin08a, TGH02]. Federico [Ano03z]. Fedkiw [Ano03-51]. Feedback [PXXW06, AM04, DGPT08, GW03b, Kar01, KT02, Kar05, Kha05, KLP09, KX05, LC06b, LCL00, LYW02, LY06a, LWY09, NTHP09, PKP09, Raz00, SDK09, SN03, SA03, Wen00, ZW03b]. feedforward [LL03a]. Feenstra [Ano01z]. Fejér [AG01b, BV09]. Fekete [SV09]. FEM [BS03b, GKM08]. Fensel [Ano03-71]. Fermat [Cha01b, Ben00]. Fernando [Ano03-33]. Ferziger [Ano03r].

FETI [CKL06]. Feynman

[Ano03-33, FM03, Ano03-32, Ano03-33]. FFT [CC04b]. FFT-Perturbation [CC04b]. FGrp [YL05]. fiber [DDCB04, RQH^{+}08]. Fibonacci [KN08, Wan00a]. Field [ASL06, TS00b, Aro08, CO00, CKY07, CGPB08, Har01, HAKA07, LWZ09, RF09, TK09, YZ02a]. fields [AL08a, Ano04i, CHM06, Fer01, GPvZ06, Gal08c, HSS01, JKY06, JZC06, PS07b, Sma01, Tan07, YLZ08b]. fifth [Wu09, YW08b, Zha09c]. fifth-order [Wu09, Zha09c]. figures [Ano04i, Köh00]. filamentary [Bae07]. filled [Dem00, ZX09]. filling [LW03]. film [ES05, LWY06, MM07b, SANH08, SL00b]. Filon [BS00a]. Filter [SK06c, Cha06, GHC09, HK08, yNyLjZaZ08, SMW04, WWH05]. Filtered [Con03]. Filtering [BIKL06, Sar06a, AF09, AC06, AL08a, AHNV01, OLAS01, Ter00, ZM94]. filters [CY05, Dim08, Lun03, MZD09, ZX08]. filtration [HK08]. final [Wan02d]. finance [Ano03j]. Financial [CCF06, AAK08, Ano00n, BGMB02, BBGG08, Syn08]. financing [HL08b]. finders [Pan05b]. Finding [GGA09, PS07a, BGP04, BPW00, CGY08a, CjBN09, HLM04, HLCW07, KI04, LCY05, LdSF00, Pan05a, PIR04, PR06, WXS01, WX03c, WSZ03, YC06]. finds [HLC00]. fine [FBA07]. Finite [ASL06, CCC08, CD06, GNR05, Jai05, KC02, Lia05, LY08d, Pao01a, Raz00, SMH08, SK06c, AHN06, AS03a, AMT03, Ano03-84, AMV08, AO09, ACH01, Ban00, BDA02, BJBL03, Bed02, BCS03, BM05b, Bla04, Bla07, BG04b, BH01, BKK08, BMM^{+}06, BLS05, CJL04, CHS02, CWY08, Che04f, Che00b, CHM06, CCP09, DW08a, DMVR00, fDxZ11, DDF02, EP04, FK01, FR04, GY08, GM09b, Gu07, GSS04, HJW05, He02, HL06, HGS07, HSS01, IS04, IN07, JKN09, JKY06, KK09, KPP00, KLS06, KP06, Kum04, Li00a, LL02a, Li02, LPvST07, MS04a, Mic03, Mur08a, NS07, NSB08, Nic04, OP01, OLS06, OY06, OS09, PLL03, Pet04, PDBH03, Roe09, Rub07, Ses03, SDZ06, SG08, Sma01, Son09, Ste07, SB06, SBD^{+}00, TCH07, TCX08, Wan02c, Wan07a, WJA08, Xu03b, Zha03b, ZW03a, Zha09e, Ano03-34]. finite-buffer [GSS04]. Finite-difference [GNR05, AO09, HJW05, IS04, Nic04, Roe09, SDZ06, SBD^{+}00, ZW03a]. Finite-dimensional [Raz00, CHM06, TCX08]. Finite-Element [ASL06, SK06c, Lia05, AS03a, BJBL03, BCS03, Bla04, BG04b, BLS05, CJL04, Che04f, FR04, HL06, KLS06, KP06, OLS06, PLL03, PDBH03, Ses03]. finite-strain [Ste07]. finite-volume [BM05b, BMM^{+}06]. fire [MKF09]. firm [LZS07]. First [Ano00i, PWX08, ZLJW06, AVO08, Ano01y, Ano03-70, BX07, BHN01, BHNO04b, COEP00, Can08, CH00a, Dix05, Don09, GS09, GJS01, JLT04, JT07b, KÖ08b, Liu07a, Lon03, LSN05, LJ08b, NF03, NRL06, PHW03, SL07b, STH08, TL08, Uje03, VNFF08, Wan00d, WW08b, Wan08a, WZ07c, Yan01, ZFYS05, ZLY06, Zho08]. first- [NF03]. First-Order [ZLJW06, PWX08, BHN01, BHNO04b, COEP00, Can08, CH00a, Dix05, GS09, GJS01, JLT04, KÖ08b, LSN05, LJ08b, NRL06, PHW03, SL07b, STH08, TL08, Wan00d, Wan08a, Yan01, ZFYS05, ZLY06]. Fischer [SZ09b]. Fisher [Cav00, Mic03, cW09b]. fit [Sür08]. fitted [KP03, KK09, KS03b, Sim01, Wan02c]. Fitting [KU01]. five [Ano00j, Ano00k, DFFL09, FM08b]. Fixed [AO01b, AO01a, AA09c, COE03, CWY08, CCSL04, Dha06, yL06, RR09a, ZC09b, AO01c, AP01, Bai01, Bou00, CC04a, CKQ08, ĆRRU08, Dha03, Dha04, Dha07, ED09, FT05, udK07, Gu07, GN09, HH04a, Hua00, IK03, Kam09, Lan04b, LTTH09, LLX05, Mar00, MSN^{+}00, Moo02, OI00, OS09, PKT07, Rad09, Roe09, Sya01a, TCH07, Wan07a, WA01, Won08, Xie01, XWZ04, XG07, YYZ07, Yav05b, ZP02, Zha08b]. Fixed-Point [yL06, AO01b, AO01a, CCSL04, Dha06, AO01c, Bou00, Dha03, Dha04, FT05, IK03, Lan04b, Xie01, XWZ04]. fixed-points [CC04a, Roe09]. fixed-sign [WA01, Won08]. fixed-size [MSN^{+}00]. flags [BH00]. flame [BT03]. Flanagan [Ano00-36]. Flashcards [Ano00-52, Ano00-53, Ano00-72, Ano00-73, Ano00-92, Ano00-93]. flat [DF02, FWL04, LF05, XL09]. fleets [HY08]. flexible [HTKM09, XW05]. floating [Gra08, PMQR09]. floating-point [PMQR09]. Floreano [Ano03-36]. Florentine [Son00]. Floudas [Ano00y]. Flow [BR06, MM00, PM08, WZE^{+}06, ASHA08, AZB08, AH08, AS09b, Ari07a, Ari07b, Ari09a, Ari09b, Ari09c, AB09, Bat07, BW09, BRC00, Bla07, BC08, BFO02, Cap07, CMS04, CCZK09, DS08, EM09b, ETH05, FMFV08, FAF09, FTC03, GGE09, GGD00, Gre04, Gre09, HW05a, HASA07, HNSA09, ICM09, INP08, JZC06, KK06b, KHA07, Lee01b, Lee04, Lia05, Lin01, Lin06b, LY08d, Mak09, MAR05b, MP08a, MKSB08, MG03, Mur05a, NMD09, PGH^{+}08, Pru09, SMH08, SANH08, Sar03, SASHS09, ST07, SHI09, SS09e, SGR^{+}09, SR07, SZL02, SM08, STH08, Tan01, TYS03, TK09, Tur07, VSKF08, WL04, WSG08, WB09, XL09, YZHY09, YTEC02, ZZJ^{+}04, Zho07, ZMM07]. flowfield [AN06]. Flows [Hay06, HLMR08, AA05, APS07, Bab09, Ban00, BC00, BJK^{+}09, BCGP03, CLL09, Dog08, Fer03, Gla04, Gla05b, Gla07b, Gla08, Gla09, GMŚ07, HHK06, JS04, JY08, LLCS02, LLC08, MM07a, MG03, NS07, Pad06, TRD08]. flows-theory [Gla05b]. flowshop [XW05]. flowshops [XXLD04]. fluctuations [AK08a, Iac07]. Fluid [BR06, Hay06, SGR^{+}09, AS09a, ASHA08, AZB08, AH08, Ano03r, Ano03-88, AO09, Bat07, BBCE09, Con03, FMFV08, FAF09, GGD00, GS07a, GMŚ07, HM05a, HHK06, HASA07, HAKA07, HNSA09, HHRA02, ICM09, INP08, KK06b, KHA07, Lee04, Lia05, MM00, MAR05b, MP08a, MG01, MG03, MRMRM02, OM09, RZH03, SMH08, SANH08, ST07, SHI09, Sta07, VSKF08, ZZJ^{+}04]. fluid-dynamic [AO09]. Fluids [HLMR08, APS07, ER07, Fei07, FS07, FF07, Lee09, Qia08, SZL02, VNFF08, XL09]. flux [Gos00]. flux-vector [Gos00]. foam [DK09]. focal [AA01, And03, BW05, BW06c, CW05, Ely04, Hof03, LDK08]. Focus [NKCS08]. focused [CML^{+}09]. Fokker [AU03]. fold [Far02]. folded [HK07]. folding [LY09c, RM08]. Folkmar [Ano03-67]. following [Qua01]. forbindelser [Bu00]. force [Bel09, GX04, Zha09g]. Forced [ÖZ07, ÇT05, GZ06, KÖ08a, LZ04a, LG04a, LX07, Öca07, Xu08b, ZY08b]. forces [CJ08, CM09, RB07, XXD06]. forcing [LL08b]. forecasting [BBGG08, HW08, LC07a, LCL08, NLLL09, TZS03, Ano00-30]. foreign [CM03]. forests [SS09c]. fork [ET09]. fork-join [ET09]. form [AN05, Ano00-42, Ano04h, GS09, Köh00, WY01, WX00b, XX02, YJCZ08, YMI06, Tun05]. Formal [XM08, BV01b, TXC02]. formalism [PM01]. formation [CL00, De 04, HH09]. forming [CPP08]. Forms [LW06, DBX04, HL08c, Sie03, WF05, XWD04]. Formula [RKS08, BD02, Ber05, CAZ01, HR08, Liu08d, Lon08, Uje04a]. formulae [Ana07b, BD03, CP01a, KNP06, MCS08]. formulas [BCA09, CBGVPP09, DGVP06, HS07, HHS06, HHY09, LS06a, OSS07, YW03a]. formulation [BBL05, BGLT08, CW09a, EG02, Fan09, Hau06, LT04, MG08a, Man05, Pla06, QS03, Zha09a, ZXlD09, Zha09g, ZXX05]. formulations [HL00]. Fortin [Bla04]. FORTRAN [Ano03s]. Forward [GPvZ06, AK01, Jun01, LZS07, YC07b]. forwarding [XCH07]. foundation [PC08]. Foundations [Ano00-54, Ano01l]. Four [ARS02, GY09, HZ08b, AVO08, Ano05g, DLM08, FY03a, FY04, HW00, Jan09, MP05, MM08d, PLG07, ZW09a, ZG09b]. four-constant [AVO08]. four-links [Ano05g, FY03a]. Four-moduli [ARS02]. four-order [ZW09a]. four-point [Jan09, PLG07, ZG09b]. Fourier [Ano03-52, RR03b, Abu01, AVO08, ADH^{+}03, BHHA08, Boy07, BJ00, CJO05, KHS00, Mal04, MJD08, ME03, Ras06, Ye02]. Fourt [TA01]. Fourt-order [TA01]. Fourth [Swe07, Bai08, BD02, BGNA07, BG07a, CYG05, CZ09c, Cui09, FT03, KZ02, LZ05, LST02, McG03, NMD07, OLS06, SANH08, SALK02, SR03, SR04, Slo09, WJA08, YW08b, Yil08, Zha03b, ZL08c]. fourth-order [Bai08, BGNA07, BG07a, CYG05, CZ09c, Cui09, FT03, KZ02, LZ05, LST02, McG03, NMD07, OLS06, SALK02, SR03, SR04, WJA08, Yil08, Zha03b, ZL08c]. Fox [PS09]. Fractal [CFM02, SWK07, AT05, BI01, Mac06, WL05, WLL06, WYL08b, ZLN06]. fractals [Ano03-40, Ano03-35]. fraction [Bel05a]. Fractional [WY06, AN09, AC06, AL08a, Amm07, Ana04d, Ana07a, Ant08, BKM08, CE02, Das09, fDxZ11, FAF09, FR04, HRM08, HdLS03, Isk04, JSS01, Jum06, Jun01, LT09a, LW09c, LTSW02, LTS03, LLNS05, MQ06, MO07, Mur06, Mur07a, Mur08a, Mur08b, OM09, Roo08, SGM09, SM00, SZ05, TWS00, WLS06, Zal02, Zal03]. fractional-calculus [LLNS05, WLS06]. fractional-diffusion [Mur07a]. fracture [PDBH03, YLB06]. fracture-dominated [PDBH03]. fractured [MKSB08]. fractures [Kor07]. fragile [RMC08, ZTH09]. frame [ADH^{+}03, dSM02]. Frames [SKHK06]. framework [DE06a, JTRu09, KPSW06, LGBS00, LLL08a, Noo02a, SL00b]. Frank [Ano04w, Ano04v]. Franklin [Ano00n].

Franz [Ano03f]. Frazer [Ano00-87]. fréchet [LV03, AO01a]. fréchet-differentiable [LV03]. Fredholm [BMHV09, CK08, DI08, Had03, LN07, Pao01b, YLD09]. Free [GC06, GCN06, BI02, CT02a, CKKK06, DPU07, FG00, bJQxC09, LHC08, MPP08, Pad06, PB01, Pan05b, PMQR09, Ryo04, UHuI09, WX03c, WSZ03, WK05, Zal03]. freedom [Mos08, PP08]. Frequency [Lin08a, AVA01, CW09a, CSX01, Fan09, JZ08, ÖY07b, Zha09a, ZXlD09, Zha09g]. frequency-amplitude [Fan09, ÖY07b, Zha09g]. Frequency-domain [Lin08a]. Frequent [XT04b, LGH09, ZY07]. friction [BOP07, SBPD08, TRD08]. Frictional [Mig06, ZMM07]. frictionless [FS04]. friendly [Ano00-87]. Frobenius [CBJ01, WW08a]. frog [Sny00]. front [Gou01, Ano03-37]. Frontiers [Ano03-38]. frying [FM07, FM08a]. Fuchsian [LTS03]. fugacity [CGPB08]. Full [SC07, Ano03-71, Gal05, JH09]. Full-rank [SC07]. Fully [Ste06, LUZ05, SSSH01]. Function [Sar06a, AS05, AS07, AM08b, Aou08, AY09, AD09, Asl11, BSR02, Ban07, BB09a, BA09a, BB08b, Can03, CLMM06, Cha08, CCF06, DCW08, Dra03b, DH07, Elk06, FDWC02, GGA09, GPL08, HRP01, HS00, KHN09, KÖ09b, KHS00, LTYB09, LF03, LH05b, Mer05, MF01, OS08b, PD06, PS09, PB02, RGTT01, SERG09, SV06, Sar06b, Sav08a, SDB09, SDZ06, TD07, VTR04, WYL08b, mWmZ09, WJS09, Wei00, WZ09, WH07, Xie08, XYC^{+}09, XTZ08, YC07a, YZZC09, YTEC02, ZTW09, ZX09, ZLN06]. function-based [SDZ06]. function-value [MF01]. Functional [HLY06, HL07, LHP07, MCZ00, SW02b, AB08a, AML02, Ayk04, BF03, BX07, BSD00, BA01, BPA03, BM01, BM02, BK04, BHN01, BHNO04a, BGO04, Cab01, CCP02, CFM02, CH00a, CP01a, CP02, CD09, CWZZ07, CD05c, CHR09, DHM05, Du08, ERRY01, GYS05, GL09b, GZ06, HPT05, HM05b, HZ03, HYLW00, HZ08b, HC02, Huo05, Ibr09, KKZ02, LZL04, LZJ06, LFZ08, LX00, LL08b, LD02, LJ08b, Mar09, Mus06, Nie01a, NRL00, NRL06, NRL08, SW02a, SL07a, SDCL08, TBG05, WJX04, WZG00, Wan00d, WSL08, WWC08, WH05, WGZ06, XX09a, Zen09, ZZ01a, ZS04]. functional-differential [HC02]. functionals [Cap00, PS04]. Functions [BFL06, DE06b, AO01b, AC02, AÖS04, AIS08, Ana04d, Ana08, AHS00, Aou08, ASS08, AM08c, AY09, AD09, BMHV09, Bal00, BKN09, BBCD02, Bax02, Bax06, BI02, BCGA08, BD07, Bib06, BV09, BPW00, Boy01, BR03, Buh06, CYX09, CCS02, CGGC02, CZ06, CWW08, CZLC08, CC09b, CK06, CCHK08, Chu07, CC08b, DGVP06, DW08b, DGR04, DS07, DR02, Don04, DS06b, Dra08, DF02, Dua08, EAA09, FF02, Fan02b, Fly06, FWL04, FLW06, GK02, Gal02b, GSS02, Gan00, GA06, HL02a, HYZ06, HS07, HHS07, Hof03, Hon02, HF00, IWL04, IW06, Jah04, JSS01, JWSS04, Jum01, Jum06, Käm00, KH00, KDC08, KLP09, Kov00, KT03b, yL08a, LF05, Li00b, Li01c, LLML03, LL03a, LTY05, LWY02, LWY04, Liu08c, LU09a, Mis05, MA09b, MDPD02, Nes01, Nob06, NCK^{+}08, Oro02, OF08, PdlF09, PS03a]. functions [PMS07, Pie09, PG08, PD04, Rap01, SN02, SM00, Sri03, Sun00, SJL09, TZS03, TL06, Ubh04, UA06, VBG07, Wac00, Wan00a, WYL08b, Wan08b, WSZ09, WJS09, WHL02, WW08c, XC04a, Xu02b, Xu03b, XC07, Xu08a, XA09, YO06, YW07, YW08a, YH11, Zal02, Zal03, Zen00, ZG02, ZC06, ZCL08, Zha08b, ZL08b, Zha08d, ZS08b, Zhu09b, dSOGRMRL04]. functors [EGM^{+}02]. fundamental [CWLC07, SKG01, YTCF08, Ano00-31]. Fundamentals [Ano03-39]. Further [TX04, Wan02a, WA01, BD04a, Jum06, Kir08, LYX09, MPU00]. fusion [WWHL09]. future [Ano00-76, Ano00-77]. Fuzzy [AEÇ06, Ana04a, Ano05f, CC05, CC08a, DCLF09, Eke02, ESNP06, EMP^{+}06, Fot09, GL05, HLSL03, KDY09, LZL06, LLS09b, LCWL06b, She05, Shi05, SL06a, SL06b, TYW05, WS04a, WLW09, XMW09, YL04b, AVLPN04, AH09, Ana04c, Ana04e, AD08, Ano00-47, AA09b, BM01, BM02, BM04a, BW06b, CCRMRF00, CCRMOG04, CCSL04, CH05a, CL01, CL02b, CT08, CZ09a, CCL02, Che04g, Cho03, CSLC07, CT02b, CL09, DLF09, DLB03, DZQ06, DSA09, DCR07, EGM^{+}02, ES08b, GS08, GYPL08, GPL08, GYPL09, HJWL03, HLCL00, HCWC07, HF00, HLY09, HL09b, Hua08c, HLZ09, Isk02, Isk04, Isk08, Jun09, Jun10, JY11, KH02b, KH02c, KLP09, KC05b, yL08a, LYL00a, LCYL00, LCL00, LML02, LLML03, LL03b, LMHW05, LC07a, LCL08, mLpW08, LJ08a, LSL09, LLJ09a, sLhYL09, LLL08b, LH05b, LCWL06a, LCWL06c, LWL08, LNL08, LT09b, Liu09a, Luo06, LHLW07, MZD09, MM05a, MGG05]. fuzzy [MM08a, NLL07, NLL08, PL01a, PL01b, PKM08, Pan05c, Pan07b, PdlF07, PdlF09, RFCCRM03, SW09, SSL07, SN03, SFL02, SWL05, SWX06, Sya01b, SLJ04, SL06c, SJL08, TB09, Tan03a, TAT07, Tia01, WT00, WL07b, WK08, WML08, WL08b, WHS08, WL09a, WI08, Wu03, Wu04, Wu08a, WWHL09, XZ08b, XTZ08, YLY^{+}09, YL09b, YYNK03, YW08c, YLD08, YZWL09, YX09b, YL04a, YL05, hYL09, hYLL09, ZAAESG09, ZDS08, ZX08, ZhYL05, ZYL06, ZL06, ZL07b, ZXTL09, ZL05]. fuzzy-random [Ana04e]. Fuzzy-rough [XMW09]. fuzzy-stochastic [PKM08]. fuzzy-wavelet [Ana04c]. fuzzy-weighted-average [CH05a]. fuzzy/fuzzy [PKM08]. fuzzy/fuzzy-stochastic [PKM08]. FVS [LW09a].

G [AHRT09, Ano00o, Ano03-32, Ano03-33, Ano03-42, Ano03-85, AM04, Gup02a, JC05, KA02, LKH03, LL06b, LWY09, Wan04a, WZ07b, WX09a, GLZ00]. Gabriel [Ano03i]. Gagne [Ano00i]. Gale [Ano00n]. Galerkin [ABBD01, ABBDP04, AVO08, BAK09, BLS05, EGZ04, GW01a, Gir03, LY09a, MEG08, MS04b, RW03, RPO03, SW06]. Galois [TS00b]. Galvin [Ano00i]. Gambosi [Ano00o]. Game [CE02, EAE09, FM08c, TOH03]. games [Ano00-57, AMR05, CS02c, Din00a, Din00b, DPJ03, Gal08c, HLY08, KL02, MS03, ND01, NSQ02, PRA01, PRA02, SA03, Wu00, Wu01, Ano03-49, Ano04h]. Gamma [Ata09, Kum06, Sri03, Gal02a, Mer05, PG08]. gaps [HHR07]. Garber [Ano03-38]. GARCH [GT08]. Garfinkel [Ano00j, Ano00k, Ano00v]. Garrett [Ano00z]. Gas [Tan01, BRC00, CS08, Dem00, Gla04, Gla09, Kor00a, KW03, NMD09, PP08, YLZG01]. Gas-kinetic [Tan01]. gas-phase [KW03]. gases [Gla07a, Tan01, dSM02]. Gasmi [Ano03v]. Gauker [Ano03-90]. Gauss [AG03, CLM03, Kov00, LZJ04, LCC08, MC07, Sha09]. Gaussian [CH00b, CHM06, EPS05, HL07, XX09b, Ano03-40]. Gautschi [Ano03-46]. Gaver [VA04]. GBP [Ano00d, Ano00q, Ano00r, Ano00u, Ano00x, Ano00y, Ano00w, Ano00-28, Ano00-47, Ano00-46, Ano00-50, Ano00-56, Ano00-71, Ano01d, Ano01i, Ano01g, Ano01h, Ano01w, Ano03-37, Ano03-82, Ano03-48, Ano03-69, Ano04j, Ano04i, Ano04k, Ano04o, Ano04p, Ano04z]. GDQ [Che04a, TV07]. GDSS [SHS05]. Geert [Ano03-89]. Gegenbauer [FF01]. gene [AMYS^{+}09, FM08c, Ano03-41]. General [DL08, Roe09, WW09a, ZAY08, AHRT09, AZ01, Ano04-27, AHK00, Boy07, CHY01, Din00a, GGA09, Hoh01, HCLF01, HGS07, JZQZ07, JL09b, KA02, LV03, LP04, LLNS05, MP07a, MH02, MRT05, NOY01, Noo00a, Noo02a, PS04, PG06, Sun05b, Sun05c, TRG01b, TaL09, Wan05b, WLS06, WS07, Wan08b, XH07, XD02, XDD09, XX02, YH11, YXZ09, Zha04a, ZXL07, Zhu08a]. generalisation [Ano03-48]. generalised [KFMH06, SB03]. Generalization [LY04, WD07a, Kir08, Lon03, MJD08, MPU00, PCRR09, ÜL02]. Generalizations [Jun09, SJL09, THD08, Wan09c, Xie01]. Generalized [AAI05, AI06, AEÇ06, AA09a, Ant08, BH00, CyL08, Din02, FjH07, GSS02, fHCH09, HBCK00, HX03, Jeo04, yL06, LCJL03, LKS08a, LKS08b, Nie01a, Pie09, RR08, SB00, Tia01, Ver08b, WX09b, Yan03a, YD08, ZDS08, AOLL04, Aou08, AY09, AD09, APS07, Ash05, AAR03, BZN07, BGAM09, BDGR01, BCR04, CLY08, CGY08b, CGY09b, CWY02, CWW08, CHZ09, DLSB08, DRR09, DH08b, DGPT08, DP02, DYZ08, Don04, EGM^{+}02, FW00, FM06, FF01, FMFV08, FAF09, FS07, GL09a, GPL08, GA06, HSG^{+}06, Hua01, IUCR06, JG07, JZ02, JZT07, KK08a, KT08, KY02, Kov00, KT03b, KI04, LLWY00, LZZ04, LDK08, LHT07, Mah09, MWH08, Mis05, MK08, NC09, PP01a, PSS07, PWX08, Pat05, PPA08, PL05b, PZ07, PY09, Pla06, RNM00, RB06, RB07, SD09, SERG09, Sav07b, SP08a, SC07, SWKT08, Sot02, SDCL08, TDY08, Wan00a]. generalized [WZG04, WJS09, Waz04, Waz05c, WA08, WlW08a, WZ09, qWjH07, WLW08b, XCS^{+}08, YV05, lYsT09, Yu04, YL07, Zal02, Zal03, Zen00, ZGY05b, ZGY07a, ZGY07b, ZCM01a, ZX08, Zha01a, Zha07b, Zha07a, ZY08a, ZCK00, ZXX05]. Generalizing [WL07b, ZZD08, DS06b]. generally [HSM00, Ter00, ZM94]. generate [Lin07, ODO03]. generated [AT05, CLS08, NSQ02, TC09, WYL08b, Wei00]. generating [CCS02, HHS07, MCH01, MH02, MCZ00, PXW09]. generation [CJL04, GG06, LBZ^{+}09, NKCS08, YZWY08]. Generator [LLB05, LW04, TK02, Tan03b, Tan04]. generators [CGFS09]. Generic [DTZF03, NK01, WB09]. genes [Mic08, MD06, ZCCB08]. Genetic [SA03, SP09, WWWL08, Cha06, CCL02, Hau06, JCCT09, Lee01a, LLJ09a, MMM04, SKU01, WNS09, WFTL09, XD02, YZHY09, ZXTL09, ZLN06]. genomic [FDY^{+}08, TB08]. genomics [RMC08, Ano03-53]. Genuinely [Ili04]. Geo [GS07b, AHRT09, AM04, LKH03, WZ07b]. Geo/G/1 [WZ07b]. Geometric [CLX08, DA06, OSSY04, Ana06, Ano00-41, CC05, Chu07, Gal08a, LG00, PKM08, Phu02, WZG04, WFLC05, WZ07a, Ano03-42]. geometric-mean [CC05]. geometrical [Zha08c]. geometrically [SJJ09]. geometries [HHK06, Nic04, RRR07]. geometry [CLL09, Cip08, MUHdH03, PG04, Ung00, Ung01, Ung05, Ung07, Ano00-32, Ano00-33, Ano01m]. George [Ano01o, Ano04y, Ano04z]. geosphere [VTR04]. geostatistics [VTR04]. Gerdes [Ano00-32, Ano00-33]. Gerhard [Ano03e, Ano04i]. Gerhardt [Ano03f]. Germany [Ano03-88, Ano04f]. Gersho [DW05]. Gert [Ano03e]. Gert-Martin [Ano03e]. GF [TS00b]. GI [FP06, GS07b]. GI/M/s/c [FP06]. Giacomo [Ano04d]. Gill [Ano03q]. Gillian [Ano03-36]. Gilpin [CWL07, FW00]. Gini [BP09, CD05b, CD06, SJJ09]. Ginzburg [DCW08]. given [SA05]. glass [BB06]. Glenn [Ano00-91]. Glimcher [Ano03x]. Global [Ana05, BGA08, BPSW00, BFL06, CZ03b, DK04, FW00, Gal02a, GS08, GW03b, HLX07, HZ05, ID05, ID08, Kr05, LL02a, LX03, LYW02, LLT08, LG04a, LXK09, LY06b, MAH08, MY05, MHS01, Mur05b, Mur07b, PG06, Pit03, Pol09, TL07, Wan01, WZG04, WC04, WWL06, WL08a, Wen02, XC06, XWD04, XZZ01, XLL03, XCD07, XLT09b, YTC08, Zha01a, Zha03a, ZY05, ZZ05, ZLLL08, ZCM01b, ZZ02b, Ano00y, BGG04, BGS04, COEP00, CS05a, CH05b, CWL07, CSC^{+}09, CZ02, Csa01, CWZ06, DL05, EG02, EM03, FS07, GK02, GGG03, HW05a, HZZ02, HWHG08, Huo05, Hur00, JK09, LYW04, Lei02, LW05, LZ05, tLW08, LWG04, PG08, SA02, SC00b, SY09, VG08, Wan02e, Wen00, YH09, ZS08a, ZX09, ZZ06]. Globally [Yan03f, SZ09b]. Gloria [Ano04-31]. Glowinski [WFC04b]. Glynn [Ano00-74]. GMRES [CJ04, HBP06, MAK03]. GMRES/LS [HBP06]. GMRES/LS-Arnoldi [HBP06]. GNU [Ano00-39]. GNU/Linux [Ano00-39]. Goal [OP01, PDP05, PS08b]. Goal-oriented [OP01, PDP05]. goals [LI00c]. Golay [Sol00]. Gol’dberg [YXZ09]. golden [KN08, SA08]. Goldstein [HS04a, SS07a]. Goldstein-Levitin [HS04a]. Golinski [Ano00-80, Ano00-81]. Golomb [Blo00]. Golombs [Bu00]. Gompertz [LP08c, LP08b]. Gong [Ano00u]. Good [KA00a, Tou05]. goodness [Sür08]. goodness-of-fit [Sür08]. goods [Bel01, YC07b]. Gordon [Ano03-78, SHY08, WX09b, Lu09b, mWmZ09, Waz05a]. Gorges [ZYY08]. GPB [Ano04v]. GPS [BGA08, KKS^{+}04a]. Graceful [MP05, MP08b]. grade [AS09a, ASHA08, GS07a, HASA07, HNSA09, MP08a, SMH08, SANH08, SHI09, VSKF08]. grade-two [AS09a, GS07a]. Gradient [XDD09, ABBDP04, FK01, GMS08, JZT07, KL05, KP06, SS07a, WD08, ZLD08]. gradient-based [ZLD08]. gradient-finite [FK01]. gradients [Bax02, KF05b, MPW07]. grading [Che04g]. Grain [ASL06]. grains [PP08]. Gram [GLR05]. granular [BS08, LLC08, Pei02, TRD08, Wan00b, ZMM07]. granularity [CGM08]. Graph [LMA08, RMC08, BR05a, Das04, Guo07, KC03a, KYAA08, PM05b, WH08a]. graphical [DKO03, LW09b]. graphics [Ano00-57, Ano03-35]. graphs [CPZ00, CZW08, CLS08, Den08, FL05, ISAAM09, KLZ08, Kor05, LL07, LL08a, MWH08, MV08b, NT02, WH08a, WHL^{+}09, ZLY09, Ano03i, Ano03-83]. Grau [Ano04-31]. gravitation [Ano03-33]. gravity [Har01, RQH^{+}08]. Gray [Ano00-74]. greatest [KFMH06]. Green [HL02a, Hof03, LWY02, LWY04, WC04, XD09, YW07, Zha08d]. Greeorv [Ano03-87]. Greg [Ano00i]. Gregor [Ano01j]. Gregory [Ano03u]. Greuel [Ano03e]. grey [Lin08a]. Grid [BI02, MYR09, CJL04, CWZ03, SC00a]. Grid-free [BI02]. grids [BMH08, CHS02, CKL06, CY00, HS04b, LY08d, NST01, SMW04, STK08, TSZ02]. grids/projection [CG05]. Groetsch [Ano00-35]. Gropp [Ano00-88, Ano00-89]. Grossberg [Gon09, LYJ09, NC09, RC07, ZLH07]. groundwater [AS09b, DPU07]. Group [Che04g, Ano03-58, Ano05f, BENT06, CL00, CCKH07, Don04, ES08b, FF01, NLL07, NLL08, PL01a, TSY09, Ung03, WL07b, YL04b]. grouping [TSY09]. groups [AK03, AA09b, BH00, Öze08, YL05]. growing [BBH07, BB09b]. Growth [ASL06, Bra08, bFFlT09, Ano03-28, BBI05, Boy01, CCW00, Cui02, GMG09, Tur04, YXZ09, ZC01, ZDLH08]. Grp [YL05]. GRPM [HZ09a]. GRPM-style [HZ09a]. Grüss [Ana07b, Che01c, DMP01, Liu07b, LN09, MPU00, Mat03, Nie09, ÖY09, Uje03, Xu06, Yan06b]. guaranteed [Par04b, PPI03, Ryo03]. guess [KLS06]. guide [Ano00-74, Ano00-82, Ano03-85, FG00, Ano00-27, Ano00-34, Ano00-65, Ano03-72]. Gunter [Ano00g, Ano00h, Ano03w]. Gurland [Mer05]. Guus [Ano00-38]. GXH [LL05]. gyration [DZQ06].

H [Ano00-55, Ano03f, Ano03r, Ano03-54, Ano04-30]. Haase [Ano00g, Ano00h]. Habib [Ano04v]. Hadamard [BV09, EAA09]. haemodynamic [MG03]. Hagen [Pru09]. Hahn [Ano03-65]. Haib [Ano04w]. Hairer [Ano03-42]. Hal [Ano00j, Ano00k]. Half [Buh06, DLSB08, DM08, DR01b, HM01, JEP08, KK08b, LG09a, MZ09, ÖZ07, TRG01a, TGS07, Yan04b, Zha09d]. half-line [DM08, HM01, MZ09, TGS07, Zha09d]. half-linear [DR01b, JEP08, KK08b, TRG01a]. Half-plane [Buh06, Yan04b]. halftoning [CCP09]. halibut [Ano03-80]. Hall [Ano00-34, Ano03-59, Ano04r, HHK06]. Hallam [Ano03-36]. Hamdy [Ano04r]. Hamel [GGE09, MM07a]. Hameroff [Ano00-84]. Hamilton [Ano01-27, BH01, LL08a]. Hamiltonian [All05, AG03, CLM03, CZ03a, DCS09, GK03, HK07, KLZ08, LTTH09, LD08, MWH08, MS02, PM01, WXM09, YJCZ08, ZHZT06, ZQ00, ZQC08]. Hamiltonian-connected [KLZ08]. Hamiltonian-connectivity [HK07]. Hamiltonian-laceability [HK07]. Hammerstein [WD08]. Hamming [AB08c, Sol00]. Hammond [Ano00-63]. handbook [EA03b, Ano03-43]. Handling [PGH^{+}08, OM09, Waz07a]. Hankel [BS00a, Roj08]. Hans [Ano00-38, Ano00-58, Ano01w, Ano03e, Ano04f]. haplotyping [BGLT08]. hard [Ano03h, KS03a, LW09b]. hardback [Ano00-74, Ano04e, Ano04g]. hardware [Ano00s, Ano00t]. Hardy [Ano01w, Li00b, Har01, HZ08a, Kan09]. Harmless [SC00b]. Harmonic [Cat05, BGAM09, CW09a, DBX04, Fan09, Kru05, Leu00, Liu09c, ÖY07b, VG08, WC04]. harmony [BM00]. Harold [Ano03-49]. Harry [Ano00-42]. Hartman [Che01b]. harvest [ZWCW00]. harvest/stock [ZWCW00]. harvesting [BBI05, HZ08b, WW04a, WLW08b]. Harvey [Ano00-69]. hash [CCHK08]. Hashagen [Ano01y]. Hatfield [Ano03-33]. having [AO01b, KP01, Mic03, PCRR09, WHL^{+}09, WSZ03]. Hayes [Ano03-36]. Head [PKM^{+}06]. Health [KS07, Ano03-38, Ano03-63]. heap [CNK00, CRK02]. heap-mergesort [CNK00]. Heat [BC00, GC08, Bat07, Che04e, Dua09, Fei07, FQZ02, GL09c, GW06, HH02a, Hor01, MO08, MRdCV00, Mur07a, NST01, OIL08, OIL09, Phu02, QFZ03, RB06, Slo09, WZ09, XL09, ZW03a]. heat-conduction [NST01]. heated [JZC06, SHKL08]. heatfunction [MO08]. heating [SKU01]. Hebrew [Sch00]. Heck [Ano04j]. hedging [BM06]. Heidelberg [Ano00-47, Ano04f]. Heilbron [Ano01m]. Helge [Ano03-37, Ano03-74]. helical [FMFV08]. helicopter [VG08]. Helmholtz [DR01a, Kru03b, LWY02, LWY04, Sau07]. Helmut [Ano00q, Ano00r]. hematopoiesis [Wen02, WLZ07]. hematopoietic [AACLM08]. hemicontractive [Bai01, HH03, Hua00]. hemirings [DSA09, Jun10]. Hemivariational [Mig06, HZ09b]. Hendler [Ano03-71]. Hendry [Ano00-30]. Henrik [Ano03-37]. Henry [Ano00-48, Ano00-49, Ano03-71]. hereditary [JJ01]. Hermite [AG01b, BV09, GW06, KL06, Mak09, MC07, Sun00]. Hermitian [Jia00a, LP08a, LHL07]. heteroclinic [GY09, LL03c]. Heterogeneous [SDK01, Ano04s, Col04, EAE09, XCL^{+}08, Ano01n]. heterozygous [Lan08]. Heuristic [XD02, Ind02, WW09c, WYSS09]. heuristics [Ano03h]. Hey [Ano03-32]. Hibbert [Ano00-71]. hidden [Ano00-90, LWZ09, May00, Ros03]. hidden-class-label [LWZ09]. Hierarchical [GAA06, STK08]. hierarchies [OS08a]. hierarchy [LD08]. High [AC02, AC05, Ana04b, Ana04c, CM04, IS04, Li01c, LCL06, LYF09, NJ08, AC04, BOP07, GGA09, GPPRM01, HNR08, Liu09b, Net07, Ou08, RC09, SCF^{+}09, SR07, TGIM02, VNG04, WF01, XLL03, XLT09a, XLT09b, Ano03-44].

High-energy [Ano03-44]. High-Order [LCL06, AC02, AC05, Ana04b, Ana04c, CM04, IS04, NJ08, AC04, GPPRM01, Net07, Ou08, RC09, SCF^{+}09, VNG04, WF01, XLL03, XLT09a, XLT09b]. high-speed [TGIM02]. Higher [AK08b, BJBL03, CLT09, WWW08, AGO03a, AORS03, ADP08, BNR05, CH04, EM09a, GW01b, Jia00b, Kra01, LFZ00, LA03, Li04b, Lin05, LDG05, Liu04a, Mas07, Pal01, PGT08, Pit03, RC07, SC03, SB06, WJA08, Waz00, WA01, Won08, Yan07, Zen09, ZL08d, ZG09b, Zho01, ZZ03b, ZY08b]. higher-dimension [Zen09]. Higher-order [AK08b, BJBL03, AGO03a, AORS03, ADP08, CH04, GW01b, Jia00b, LFZ00, LA03, Li04b, Lin05, LDG05, Pal01, Pit03, RC07, SC03, SB06, WJA08, WA01, ZG09b, Zho01, ZZ03b]. highly [CXZ09, SNT08, WK05]. highway [SWK07]. Hilbert [Ano01l, BV05, CGY09a, CZ09b, DW02a, DDF02, FHT05, HS09, Inc09, IL04]. Hill [BBI05, GK04b]. Hilliard [CCK00, UK08, Ye02]. Himmelberg [Xie01]. Hirota [LHT07, Wan09a]. histogram [LG09b]. History [Ano01y, Ano01m, Ano00l, Ano00m, Mus06, Ree00]. history-valued [Mus06]. Hit [SK06a]. HIV [LMSH04]. HIV-1 [LMSH04]. HJB [KX05]. HLI [Ste06]. Hm [Ano01f]. HMM [PT07, WLT07]. HMM-based [PT07]. Hoboken [Ano03-31, Ano03-74, Ano04m, Ano04s, Ano04u, Ano04t]. Hocking [Ano04m]. HODEs [CKCG03]. Hofestädt. [Ano03-41]. Hogendijk [Ano03-81]. Hohenberg [Pol09]. Hohmann [Ano04p]. Holden [Ano03-37]. Hölder [Hua04, HL07, LHP07, Mae00]. Hölder-type [Hua04]. hole [Bs03a, EW00b, Nam02]. Holland [Ano00g, Ano00h]. Holling [SDCL08]. Holt [SZJ09]. homeomorphism [SSX09]. homestead [Ano04-29]. homoclinic [DCS09, DZ05, Lu09c]. homogeneous [BM09, El 03, Gin02, HRS08, KIXF09, KW03, LW09a, RSD09, RR08, ZC01]. Homology [MPZ08]. homomorphism [SSX09]. homotopic [Pan07a]. homotopic/factorization [Pan07a]. Homotopy [Bel09, BG08, NMD08, Ari07b, Ari09a, Ari09b, Ari09c, BNH08, BBA09, BA09c, BG11, CJK09, EM09b, GTJ07, GSN08, He09b, HRM08, KGB07, Lu09b, MZ07, MO07, MEA09, ÖY08, SNG09, SANH08, Sho09a, SHI09, SK09, Xu07a, Yil08, Yil09, Yus09, dZmClQ09, ZZD08]. honeycomb [SWKT08, YMTE05]. Hoog [Ano00-38]. Hopfield [GHW04, MXZL09, Ou08, WZZ09, XZZ01, XLL03, XLT09a, XLT09b, YH09]. Hoppensteadt [Ano04w, Ano04v]. horizontal [SMH08, Yav05d]. Horswill [Ano01k]. Horwood [Ano00-55]. Hosoya [Den08, YW08b]. Höst [Ano00-28, JD07, MM06, Qiu08]. host-parasite [MM06]. host-parasitoid [JD07]. hot [RDS00]. hot-rolled [RDS00]. Hough [LWZ04]. HP [PDP05]. HP-adaptivity [PDP05]. Hromkovic [Ano03h]. Hsiao [Ano04e]. Hu [Ano03p]. hub [QW09]. Huber [Ano03f]. Hubert [Ano03-65]. Huffman [AMM00, MK03]. Hugdahl [Ano03-77]. Hughes [Ano01x]. Human [Ano03-50, PS08b, BI01, GPAA^{+}09]. Human-computer [PS08b]. Hunt [Ano00z]. Hurewicz [WI08]. Hursley [Ano01k]. Hussein [Ano03k]. Huygens [Zhu09a]. Hybrid [CGY09a, Dha07, Gir06, HS09, SLL03, Ano03-82, CPP07, CCKH07, GL05, GLTS08, Han08, HBP06, HKL09, tLW08, MPW07, PKM08, PC08, SS09a, WL05, WLL06, XXC^{+}09, YZHY09, ZYY08, ZL09c, ZZ06]. hydraulic [Dor01]. hydro [YZWY08]. hydrocarbon [CS08]. hydrodynamics [JTRu09, MG08a]. hydromagnetic [GGD00]. hydropower [LWLJ09]. Hyper [MWH08, KYAA08, Man05]. hyper-dissipation [Man05]. Hyper-Hamiltonian [MWH08]. hyper-Wiener [KYAA08]. Hyperbolic [Ung00, Ung01, ADDL04, Ano03-37, AO05, AY06, BB09b, BG08, CDJR03, Che04d, CCZK09, Din04b, FL03, Gar01, Gos00, KK05, LY09a, PLD02, PJJP04, Soh05, Ung05, Ung07, WZG00, Zha09e, ZS08b]. hyperbolic-parabolic [AO05, AY06]. hypercube [KK06d, KM01, SLTH09, ZXX05]. hypercubes [HK07]. hypergeneralized [BB08a]. hypergeometric [Aou08, AY09, AD09, CCS02, GSS02, HS07, IR08, WJS09]. hypergroups [DFFL09]. hyperideals [ZDS08]. Hyperinterpolation [CDV08]. hyperlogistic [LG04b]. hypermodules [Fot09]. hyperrings [ZDS08]. Hypersingular [Kru03a, KR01]. Hyperstructures [SM06]. hypothesis [Cha08, WZ08b].

i.i.d [PL05a]. Iacoboni [Ano04-28]. IBM [Ano01k]. iCAM [LLT09]. iconic [Ano03-83]. ideal [BKR05, Gla07a, WL09a, WWH05]. ideals [CZ09a, DSA09, JLZ09, JLP09, Jun10]. Identification [BFO02, CM00, HD09, YM04a, BFKL08, Col04, DD08, GT08, HMCR00, MV08a, PP06, WFC04a, WD08, WTZY05, YM04b]. Identifying [TD07, JKR08, MD06]. identities [WW09a, ZY08a]. identity [AP04, Ano01-28, CD05b, Pad06, YJCZ08]. ideology [STV08]. idiots [Ano03-27]. IHCP [CM00, CM01b, Mur08b, YM04a, YM04b, ZCM01a]. II [ARS03, Ano01l, Ano01x, Ano01-28, CD06, ARS02, ARS03, CCK00, Dha04, Har01, Ikh02, Köh00, LH00b, Lia04, PS01, SYHC06, Won00]. III [Ano00-84, Ikh04, RB06, RB07, SDCL08]. IIR [Cha06]. Ill [SN01, WSZ02, CWZ03, HB04, KH00].

Ill-conditioned [WSZ02, CWZ03]. Ill-conditioning [SN01, KH00]. ill-posed [HB04]. Illiad [Ano00-87]. illusion [Ano04-31]. illustrated [Ano03-72]. Illustrations [Ano00-90, Ano04p]. ILU [MAK03]. Image [CVW00, AMS08, ADH^{+}03, BC00, BIKL06, BFKL08, CFM02, HCG08, Iac07, Li01c, LLS09b, LWZ09, MWZ09, OY06, PLP09, ZLN06, Ano04i]. images [LH00a, OY06]. imaginary [LYC07]. imaging [IDM08, KPSW06]. immersion [FM08a, Ano04-31]. immigration [Ano01o]. immiscible [BJK^{+}09, Qia08]. Immune [Zha06, LWLJ09, WWL06]. immunity [ZSPZ08]. impact [Ano01z, AP09, DZ05]. impairments [YCCY02]. impatience [AP09]. impedance [Kru03a, Kru03b, KPSW06]. imperfect [Ano04-27, BM06, LCT08, TY03]. Implementation [GPPRM01, VB01, BD01, DW09, HWHG08]. implementations [MHTR08]. implication [ZL07b]. implications [LSL09, LW08a, Luo06]. Implicit [fDxZ11, For01, Gir06, MM05b, MVS^{+}07, Mur08a, Nko00, Ste06, Ano03-51, AG03, CGY09b, CCZK09, Fen08, GPPRM01, Hua01, JPV01, KO09a, yL08a, LY08a, NF03, Net05, Net07, PL05b, Rub07, Tia01, Ver03, YS04, ZZZ02]. implicitization [WW08c]. implicitly [Gal02b, Sya02]. importance [BK07]. imprecise [GHHD08]. Improper [CZW08]. Improved [FFK02, YSH03, ZYY08, Waz05b, YHS08, ZLN06]. Improvement [Che01c, Kir08, MPU00, Mat03, TD09, BG11, LH05a, WY09a, WSZ02, Yus09]. Improvements [CCZK09, Jin09, Wu08c]. Improving [AJ09b, Chu05, JYDH02, MRP08, WK03]. impulse [LZJ06, ZJLW06]. impulses [DTZF03, GF06, Huo04, tLW08, Li08b, LFZ08, LYJ09, LX07, LS01b, Luo03, PG00, Pen02, WW08b, WD05, WGZ06, Zen09].

Impulsive [BHNO04a, De 02, LET06, PXXW06, Sat06, YZZ^{+}08, ZLJW06, ZCW06, ATY08, ADN02, BHN01, BGO04, BHNO04b, CQJ03, CWZZ07, CK02c, CHR09, DHM05, DFY02, GS09, Guo02, HH04b, HC09, HPG06, Jan09, KZD03, LDZW05, Li08a, LS08, LB01, LSN05, LJ08b, MAH08, MG08c, NRL08, ÖZ07, PXW09, SDCL08, WF07, Wan08a, WFY09, Wei05, WH05, WTZY05, XLL04, XT08, XX09a, XLT09a, XLT09b, Yan01, Yan04c, YX07, ZS04, ZLY06, ZL09b, Zhu08b]. Impulsively [ZL09a]. incidence [HWC04, NTTS09]. incident [MP08b]. included [Ano00-34, Ano00-39, Ano00-40, Ano00-46, Ano01f, Ano01k, Ano03v, Ano03-34, Ano03-58, Ano03-67, Ano03e, Ano03-75]. Includes [Ano00-42, Ano00-78, Ano00-79]. Including [GC06, GCN06, MM07b]. Inclusion [CK06, LC03, GYS05, PV00, YL09b]. Inclusions [AI06, AVLPN04, AAI05, BV05, BN00, BGO04, BHNO04b, BGO05, CGY08a, CGY08b, CKN04, CL07b, DY05, FHT05, Hua01, HF04, HN07, Jeo04, Jin07, Jin08, KB04, LCV06, yL08b, LKS08a, LHH08, Noo02a, PL05b, PZ07, STH08, Tia01, Ver09, XH07, ZGY05a, ZGY07b, ZAY08]. Incomplete [LN08, ZM04b, DKO03, PRA01, PRA02, Sri03, TS00a, Whi00]. incompressibility [Lee09]. Incompressible [HLMR08, Bla07, DH08a, GMŚ07, KK06b, MRMRM02, Sta07, SZL02]. Inconsistent [Tan03a]. Incorporating [zGtM08]. increasing [Dha07, SSX09, Sav07b, SL09]. increasingly [DF02, LF05]. Increment [KL03]. incremental [GZZ09]. increments [CH00b, HL07]. indefinite [HL08c, KK08b, LZ09]. independent [AN06, AK08c, HHJS03, KK05, LGBS00, Lü08, dG09]. Inderjeet [Ano00e, Ano00f]. Index [Ano06l, Lam05, Ano00a, Ano00b, Ano01a, Ano01b, Ano02a, Ano02b, Ano03a, Ano06a, Den08, DDCB04, HR05, KT09, KYAA08, XCH07, YW08b, Zha02c]. index-2 [HR05]. indexing [SP09, Ano00d]. Indian [Ano03-29]. indicator [KT09, Pla06]. indices [RVKDF09, TBR01]. indices-realistic [TBR01]. indirect [AVBW08, SK06a]. Indiveri [Ano04d]. individual [GMG09, TDHL09]. individual-based [GMG09]. indoor [KYJ03]. induced [Dog05]. industry [She05]. inelastic [MB08, PP08]. Inequalities [Kum04, LX06, WlW08a, WD07b, XD09, Xu06, ABCD05, Alb02, AP04, Ana04d, Ana07a, Ana07b, Ana08, BK00, BZN07, BDA02, BD04a, BCDS09, BXFZ09, BV09, Bou03, CLY08, CWY08, CCD07, Che01c, DMP01, DBGSP08, DYZ08, DDF02, FW08, FH07, GL09a, GYS05, Gau04, GK03, HF00, HBCK00, JC05, Kha01, KDC08, Kir08, Kua08, Kum05, Lan04b, yL08a, LKS08b, Li09a, LW02, LY05a, Liu07b, Liu08d, Lü08, MRS08, MPU00, Mat03, Noo00a, Noo00b, NCK^{+}08, ÖY09, POS03, RN02, SP08a, THD08, Uje04a, Uje04b, ÜÇT08, Ver01a, Ver01b, Ver03, WYH00, WC04, XH02, XWZ04, XC07, YHS08, YHT04, Yan06b, ZGY05b, ZGY07a, ZL05, ZS08b, Zhu09b, Zhu09a]. Inequality [Mig06, AL08b, BBCD02, CD09, DBEGB08, Dra03a, DS06b, FY01, GL08a, HL02b, HS04a, Han06, HLW07, Han08, fHCH09, HZ07, HZ09b, JT05, JZT07, LN09, Nie09, PSS07, Par04b, PCRR09, SZ09b, Uje03, WZ00, WXW01, WD07a, Wu08c, XX02, ZH09, Zhu08a, dSRGRMRL08]. inequity [LJ08a]. inertia [RF09]. inertial [AF09, HTKM09, You04]. inexact [Arg00a, YL08c]. inf [mLpW08]. inf- [mLpW08]. infant [JSGP08]. infected [WS08a]. infection [WWL06]. inference [APvZ01, BGLT08, ES08b, LNL08, MJ02, NLL07, NLL08, WML08, ZL06]. Infiltration [CKK06, Jon07]. Infine [Nob06]. Infinite [dM06, AO01c, AMT03, Ano03-76, Boy01, CKQ08, CDJ04, CP01b, DH08c, ES08a, Fer01, Fly06, GGD00, GW06, JPV01, JPV02, Kil04, LX00, LHP07, NP07, PG04, PWX08, PP04, RB06, SN01, SN02, Sav07a, Sav08b, SL09, WK03, YYZ07, Zha02b, dORM08]. infinite-dimensional [PG04]. Infinitely [MZ04, HF00, LDZW05]. infinity [Guo03, HL04, Ma07, ZL07a]. inflation [MM08a]. inflow [HM03b]. Influence [MB08, Mar09, Pru09, SANH08, dSM02]. Information [BFKL08, Gal08b, Gal09a, Ano00j, Ano00k, Ano01e, Ano03-28, Ano03-75, Ano04k, Ata09, CC05, HCWC07, HLY09, KC05a, Kum06, LGBS00, LYL00a, LCYL00, LCL00, LCWL06c, LJSH09, MRP08, PRA01, PRA02, QLD08b, TS00a, Ter00, TDHL09, XMW09, ZM94, ZY02, ZM04b, Ano03-45]. information-theoretic [LGBS00]. informative [HX09]. infrequent [DY09, ZY07]. inherent [Mak09]. Initial [KK09, ACD01, AP08, AOLL04, BN00, CH01, CCJ08, CS02b, CL07b, CBJ01, Che04b, DS02, FS07, KLS06, KS03b, Lam05, LX04, LWG04, MMM04, MM05b, NF03, Net05, Net07, OEV09, PF01, SZL02, WWS^{+}06]. initial-value [AP08, Che04b, KS03b]. initialization [CLJ09, CZ02, RTS07, VRJ09]. initializations [VDMK08]. initialize [MCRÁ05]. inner [HSS01, LMHW05, MT00]. innovation [HD09, YWZ03]. innovations [Ano03-60]. innovative [DY09, LHC08]. Input [PP06, AHN06, GC02, GM09b, HD09, KLP09, LBZ^{+}09, Mac06, MW00, Wu03]. input-and-output [MW00]. input-output [PP06]. input/output [LBZ^{+}09]. inputs [CS01, GLTS08]. insects [Ano03f]. inspection [SI03, WM09]. inspired [LRL09, ZXTL09]. Instabilities [MM06]. Instability [GK04b, HT05, SL08]. instances [AGS09]. insulated [DLSB08]. Insurance [KS07, LP06]. integer [BD01, Coh00, LHW04, LU09a, SHWF05, SF01]. integers [Coh00]. Integrability [GL06, EPS05, Sah01]. integrable [Käm00, LD08, SSD07, YJCZ08]. Integral [Gau04, Man01, Oro02, TÇ04, YI06, AOTW07, AMT03, AC06, Ana07b, ABB05, AO08a, AVA01, BS01a, BM04b, BPP04, BCRS05, BM05a, Bar01b, BCDF02, BCDS09, BA09b, BG07b, CM01a, CD03, CD09, CCRMRF00, CK08, CT08, CZ09b, CWZ06, Deh02, ESR06, El 03, GK02, GYS05, GSN08, GMS09, Had03, HaLC08, Has02, He02, Hor01, HHL08, Ibr09, JSS01, JJ01, Kan04, Kar03, Kru03a, Kum05, LT05, LT04, LN07, Mar07, MO01, OF08, PSS09, PKT07, Ram05, SDVS09, SNG09, SD08, SNA09, WG07a, WLW07, WS05, XC07, Xu07b, YLD09, ZL02, ZG09a]. integral-differential [BA09b, WLW07]. Integral-Type [YI06]. integrals [All02, AA09a, CCD07, CXZ09, DF06, FM03, KR07b, KR01, MRS08, ÖY09]. integrands [BCDS09, CCD07]. Integrated [Sar06b, AF09, HY06, KHS00]. Integration [Ste06, AACLM08, Ano03-42, Ano04y, Ano04z, CAZAA00, CT02a, HKL03, Jay04, Li01c, MQ05, MP02a, Rub07, VAR03]. integration-free [CT02a]. Integrative [DFY06, Ano03-53]. integrator [HLS05]. Integrators [Gir06, BCDS09, CCD07]. integro [ATY08, AO08a, BMHV09, BUP09a, CK08, DI08, Gan06, Guo01, HH04b, Hua08a, JKN09, JL09a, KGB07, LS08, MQ06, Pao01b, PXW09, SNT08, SNG09, Swe07, WG08b, Yen08, Yil08, ZJLW06]. integro-differential [ATY08, AO08a, BMHV09, BUP09a, CK08, DI08, Gan06, Guo01, HH04b, Hua08a, JKN09, JL09a, KGB07, LS08, MQ06, PXW09, SNT08, SNG09, Swe07, WG08b, Yen08, Yil08, ZJLW06]. integro-parabolic [Pao01b]. integrodiferential [SC04].

Integrodifferential [SBS06, XWL06, BSD00, BPA03, BSK04, BPM05, BM04a, BV05, BAK09, Guo02, HPT05, HM05b, LX04, LWG04, Wen00, ZCC04]. intelligence [Ano00l, Ano00m]. intelligent [GL05, PC08]. intelligentsia [Ano03-89]. Intense [Ano00-42]. interaction [JD07, LMSH04, PS08b, SO03, SZ06, Wan06c]. Interactions [RB07, FM03, MK08, OKCO08, RB06]. interactive [MM05a, XZDD09, Yan03b]. interception [YZHY09]. interconnected [TLS07]. interest [Ano04-40, Ano04-41, CK03, CM03, GPvZ06]. interface [BLS05, DD03, VC03, Ano00-88, Ano00-89]. interfaces [BOP07]. interference [LW09f]. interior [Ano03-68, CKL07, HLP05, bJQxC09, Kru03b, Més05, Wan02d]. interior-point [Ano03-68, bJQxC09, Wan02d]. internal [PP08]. international [Ano00-43, Ano00-85, Ano01z]. Internet [Ano00-60, Ano00-61, Ano04-29, Tsa03, Ano00-34]. interpolants [All02, CLMM06, CM04, FW04b, FWL04, GQL05, TSZ02]. interpolated [SALK02]. Interpolating [DZZT06]. Interpolation [DF02, LL03a, LL03b, OS08b, SAS06, Sya02, Bal04, Bal02, CD05a, CWB01, CC08b, DR02, DH07, FP02, FZ07a, He09a, HHRA02, HSWL08, KK06d, LG08, LF05, Lin04c, LLTL05, OSS07, RR03a, Sha04, She04a, Sun00, TK09, WYL08b, WB01, WlW08a]. interpolations [LF09]. interpretability [YX09b]. interpretation [SV01]. interrogation [Ano03-69, FGX05]. Interscience [Ano03-39, Ano04m, Ano04s, Ano04u, Ano04t]. intersection [Bao03, Lan04a, WY01]. intersection-form [WY01]. interstitial [MG01]. intertemporal [SZ06]. Interval [AL04, GYPL09, LA00, Liu08b, QLD08b, SOW04, TBG05, WZ08a, YYZ03, AO01c, BPP04, BDA02, BCGA05, Boy01, CC08a, Csa01, DLF09, GLTS08, HW08, HLZ09, JY11, Kum04, LL07, MC07, NUN06, XZ06, YLY^{+}09, YLD08]. Interval-valued [GYPL09, CC08a, JY11, YLY^{+}09]. Intervals [CD06, BN00, GK04b, Kil04, LC03, NFI03]. Introducing [Ano00-69, Zha02a]. Introduction [Ano03h, AA09b, FHZ08, Ano00g, Ano00h, Ano01w, Ano03-30, Ano03-82, Ano03-73, Ano03e, He09b, Ano00-28, Ano04i, Ano04p, Ano04r, Ano03-46, Ano04j]. Intuitionistic [CZ09a]. Intuitive [Nam00, Nam01b, WWH05]. invariance [BP09]. Invariant [AP03, Xu03a, Zhu08b, Ana04b, Gar01, Ter00, XX09a, ZM94]. Invariants [Nes01]. inventory [DCR07, Hoh01, HY06, KH02b, KH02c, MM05a, MM08a, Saw03, WCL02, XZ08b]. Inverse [BGP04, CPP08, DS06a, MV08a, SPE06, Yan02, Yan04a, YLL08, CCP09, GL09c, GS06, JKR08, JSK^{+}04, KR07a, LZ09, LLW08, Mur07a, NC09, ÖY07a, PST08, PSES07, SC07, Slo07, Sot02, TT09, WW00, Xie00, Yan04b, YK00, YTCF08, ZZH03, Ano00-35]. Inverses [LS02, Wac00]. Inversion [HP02, De 02, DWL08, Jog04, LS04a, MRT05, MG05, RH09, TKC02, VA04, WYZY01]. invert [JZ02]. investigating [HGS07]. Investigation [CSC^{+}09, BR05b, FGH08, MT09, Tur07, Wan02a, WDLP08, YV05]. investigations [AF02, CM09]. Investor [LWL08]. invexity [Ant05, Ant08]. inviscid [De 04]. involutions [ES07]. Involutive [CG01]. involving [Ana04d, CG08, Dha03, Ehr02, GA06, HF04, JSS01, Kan09, KT03b, Kum04, LCV06, LHH08, MN01, PMS07, Sav07b, SL09, Ver03, WlW08a, XZZ01]. IOM [Jia00a]. ion [HMCR00]. ionic [Sam07]. IP [Ano03w, Ano00-72, Ano00-73]. IPA [MPW07]. iris [MWZ09]. irrational [Zha09g]. irreducibility [XY04]. Irregular [CM09]. irregularity [Wan02e]. irreversibility [Mak09]. Isaac [Ano03-53]. Isaacs [Gal08c]. Ishikawa [SW08, IL04, KK01, LKC04, Pan07c, Tia05]. Islam [Ano03-81]. isochronicity [WL09b]. isochronous [DLM08]. isolate [BCT00]. isolated [WHL^{+}09]. Isolating [ZW09b]. isolation [XZ06]. Isomorphism [DFFL09]. isothermal [DLSB08, VM09]. isotropic [KIXF09]. Israel [Ano00-85]. issue [FHZ08, TB08]. issues [ABBD01, Ano04-40, Ano04-41, KM03a, Ano01o]. item [HY06, MM05a, PKM08]. items [DY09, HL08b, MM08a, MM08b, ZY07]. itemsets [LGH09]. iterated [CLC01, CM02, Lin04b, WYL08b, WYSS09, ZLN06]. iterated-subspace [WYSS09]. iterates [Agr08]. Iteration [LN07, TS06, AETEZ07a, AETEZ07b, AU09, AC09, BNH07, BG07a, CA09, Das09, Dra03b, ES01, Fas02, GAT07, GTJ07, GL09c, GLC09, GJ07, HWX07, HW07, Hem08, HGH02, HN07, HS09, Ikh08, IL04, JG07, KA09b, Li09b, LBW05, Lu07, MDG05, MH07, MZ07, MO07, Moo02, NMD09, OM09, Pan07a, RD07, SDVS09, SD09, SNT08, SG07, SGM09, Sal08, SD07, Slo07, Slo09, SZ09a, Swe07, SAB07, TD09, Wan07a, mWmZ09, Waz07c, Waz07b, Waz07a, WF01, Wu09, Xu07b, YW09, Yao04, Yao05, Yao07, YLD09, YB07a, ZH07, Zha09c, Zha09d, Zhu05]. Iterations [Iva07, Bai00, BGP04, CT09, CZG04, KK01, Pao03, PW06, Wan02d]. Iterative [Ale06, BD04c, BGM^{+}02, CGY08a, CGY08b, Chu05, Hua00, HF04, Jin07, KB04, KA07, KR07a, yL06, LG00, LK03, LHH08, MPS04, MN01, SW02b, WLWC08, ZGY05b, Arg00a, Bai01, CCS08, Chu07, DH08b, EGA03, FGH08, Has04, HH04b, HCLJ00, Inc09, Jin08, KZ02, KK08a, KI04, yL08b, LTT02, Lin08c, LKC04, LdSF00, MCQK09, Min08, NSB08, Noo00a, NW09, PP09, Pan07c, PZ07, Pet05, RR09b, SSSH01, SW02a, SP08b, SW08, Tia05, Ver01a, Wan00e, WW03, Wan03b, WZ08a, WSZ02, XDD09, XWZ04, YTEC02, ZGY07a, ZGY07b, ZAY08, ZHL05, ZZZ02, ZLD08]. IV [CGPB08]. IVPs [CM02, Ikh01, Ikh02, Ikh04, Tsi02]. IVSA [Zha02a]. Iwańska [Ano01r]. IWO [ZWC^{+}09].

J [Ano00d, Ano00z, Ano00-51, Ano00-54, Ano00-66, Ano00-84, Ano00-87, Ano01i, Ano01m, Ano01o, Ano01-27, Ano03-77, Ano03j, Ano03u, Ano03r, Ano03-27, Ano03-32, Ano03-46, Ano03-53, Ano03-73, Ano04l]. Jacobi [BH01, AAR03, CL07a, CCS02, DJL04, FJ05, Guo03, HZ03, JSKE05, KA00a, WSZ03, YZZC09]. Jacobi-free [WSZ03]. Jacobian [LN08]. Jacobians [ID08, ID05]. Jacques [Ano03v]. Jaffe [Ano03-60]. James [Ano00-59, Ano03-80, Ano03-85, Ano03-71, Ano04t]. Jan [Ano00-80, Ano00-81, Ano03-81, Ano03-43]. Java [Ano04u, Ano00-36, Ano01p, Ano03-47]. JavaScript [Ano00-37]. Jean [Ano03-36, Ano03v, Ano04-27]. Jean-Arcady [Ano03-36]. Jean-Jacques [Ano03v]. Jean-Pascal [Ano04-27]. Jeffery [GGE09, MM07a]. Jens [Ano01g]. Jensen [PCRR09]. Jeremy [Ano00-90]. Jerry [Ano00-74, Ano00-37]. Jerusalem [Ano00-85]. jet [JZC06]. jets [De 04, YG08]. Jim [Ano00-36]. Jimbo [HO00, TGH02]. Jiming [Ano03-68]. Jin [Ano00-50]. Jing [Ano00-50]. job [Nak06]. jobs [KK03]. Joel [Ano03-55]. John [Ano00i, Ano00z, Ano00g, Ano00h, Ano00-90, Ano01k, Ano03-36, Ano03-79, Ano03-31, Ano03-58, Ano03-73, Ano03-70, Ano03-74, Ano04k]. Johnson [Ano03-31, Ano04t, HM05a]. join [ET09]. joining [KK03, Zha02c]. Joint [FDY^{+}08, Mae00, MGG05]. joints [EA03b, FY01]. Jonathan [Ano00-82]. Joo [Ano03-83]. Jordan [Zhu08a]. Jordan-type [Zhu08a]. Jörg [Ano04d]. Jorgenson [Ano03-28]. Joseph [Ano03-55, Ano03-63, Ano04-29]. Joshi [Ano00-50]. Jr [Ano03-70]. Julia [Dra03b]. Julio [Ano03-41]. July [Ano00-85]. Jump [Syn08, WLWC08, ZLD08]. Jump-diffusion [Syn08]. Juraj [Ano03h]. Jürgen [Ano01n]. just [Ano03-58].

Kac [Ano03-64]. Kak [Ano04u]. Kallaher [Ano00-66]. Kalle [Ano00-67, Ano00-68]. Kalman [Ter00, Dim08, SK06c, ZM94]. Kalman-filtering [Ter00, ZM94]. Kamenev [DK05]. Kamenev-type [DK05]. Kann [Ano00o]. Kant [Ano01l]. Kantorovich [AG09, ZZ09a]. Kantorovitch [Gup04]. Kaplan [Ano01q, Ano03-76]. Kármán [PKP09]. Kaszniak [Ano00-84]. Katalin [GS05]. Kaufman [Ano00-67, Ano00-68]. Kaufmann [Ano01h]. Kaup [KHN09, Ma09a, RD07]. Kautz [HX06, HX07, ZXX05]. Kavanaugh [Ano04-29]. Kawahara [YM09, YB08]. KdV [Kay04, LHT07, SGM09, Waz04, ZQ00]. Keener [Ano00-59]. Keith [Ano00-71]. Kelly [Ano00-42]. Kelvin [SV08]. Keng [Ano00-31]. Kennet [Ano03v]. Kenneth [Ano00-71, Ano03-77, Ano04g]. Kernel [El 03, RNM00, SL08, Xu02a]. kernels [BBCE09, MOP06]. Kerner [BGP04]. Kevin [Ano00z]. key [CYH05, LLS^{+}09a, YSH03]. Kho [Ano03-53]. Kiat [Ano01w]. Kilian [Ano03-88]. Kiltinen [Ano03-58]. kind [CZ09b, DWL07, Kar03, LY09a, LMHW05, Liu08a, LN07, RN02, Slo09, YW03a, ZZ01a, ZT07]. kinds [KK03, LCYL00, MZD09, NQHN03]. kinematic [Yav06]. Kinetic [BS08, AO09, BPM09, BC08, Dog05, Lo 02, Tan01]. King [Ano01p]. KINSOLVER [AMYS^{+}09]. Kirchhoff [ACM05]. KKM [Din02, FjH07, JC05, Wen08]. Klaus [Ano01e]. Klaus-Robert [Ano01e]. Klein [SHY08, WX09b, mWmZ09]. Klösgen [Ano03-43]. Kluwer [Ano00d, Ano00u, Ano00x, Ano00y, Ano00w, Ano00-28, Ano00-50, Ano00-56, Ano00-71, Ano01d, Ano01h, Ano04k]. knapsack [Shi05]. knowledge [Ano01r, Ano03-43, Ano03-60, Ano04-32, WK08, YZWL09, Ano00-38]. known [SAS06]. Knudsen [Ano00-82]. Knuth [Ano04x]. Kohane [Ano03-53]. Kolmogorov [LPSS02]. Komleva [CC02]. Komleva-type [CC02]. König [GJQ07, Zen00]. Koonin [Ano03s]. Korte [Ano01g]. Korteweg [LRL06, HKL03, ME03, Ras06, UHuI09, Zha07b]. Kramer [Ano04d]. Kraus [Ano01n]. Kris [Ano00-36]. Kronecker [BUP09b]. Krueger [Ano03-29]. Krylov [CL02a, LBW09, Tou05]. Kuhn [Pla07]. Kuhn.Princeton [Ano03-49]. Kukles [Gin02]. Kuramoto [KT08, LRL06, SGM09]. Kurt [Ano00-41]. Kutta [CM02, PF01, XT04a, AG03, BJ02, CFR03, CLM03, GSZ02, GW03a, GPPRM01, GPMPR03, LGY09, LYL07, XT02, Xia07]. Kuznetsov [MYG09, BBA09]. Ky [DBEGB08, LH05a].

L [Ano00j, Ano00k, Ano00z, Ano00-57, Ano00-71, Ano01m, Ano03j, Ano03y, Ano03-35, Ano03-85, Ano03-84, Ano03-55, Ano03-87, Ano04t, Ano04-29, JC05, Ano03-36, Din02, FjH07]. L-KKM [Din02, FjH07]. label [LWZ09]. labelled [Lon08]. laboratory [Ano00j, Ano00k, Ano03-82]. laceability [HK07]. Lacunary [AEÇ06]. Lafferty [Ano04k]. Laffont [Ano03v]. lag [CH00b]. Lagrange [AN05, Bal04, Gir03, KK06d, LF09]. Lagrangian [BI02, Gir06, VTR04, WZE^{+}06, YAVB08]. Laguerre [BCR04, Don04, Dra03b, PPI03].

Laguerre-like [PPI03]. Laguerre-type [BCR04]. Lai [LC06a]. Lallemand [dKLR09]. Lamé [JKR08]. Laminar [XL09, Gre04, YG08]. Lanchester [SS07b]. Lanczos [SB00]. Landau [DCW08, LWY06, Liu09c]. landing [LH00b, LNL08]. landmark [GSNDB^{+}08, LGH09]. Langtangen [Ano04f]. language [Ano01r, Ano03-86, BV01b, CML^{+}09, Ano00-83, Ano01r, Ano04k]. Languages [VV08]. Laplace [AETEZ07a, BMH08, CWLC07, De 02, LS04a, LCGX06, LXZ08, MG05, VA04, Xu04, YTCF08]. Laplacian [AOS07, Cab01, CG08, CWZZ07, Das04, GS04, Guo07, Jia01c, JX04, KLW09, Kru08, LLWY00, LX03, Liu04c, MDG05, PGT08, SS09c, Sun09a, WG06, WG07b, WG08a, WA08, WS08c, XG07, YW08a, YWYL07, ZLY09, ZG09b]. laptop [Ano03-50]. Lar [Ano00-67, Ano00-68]. Large [DMP03, HCL07, JKN09, PB01, PW05, ZW04, dR03, ACD01, AMYS^{+}09, AR07, Ano04q, CKL07, FHZ08, Jia00a, Jia01a, JZ02, KDKS01, Lun03, LM03, MAK02, MAK03, OY06, Sid08, TLS07, XWW08]. Large-eddy [DMP03, dR03]. large-scale [FHZ08, TLS07, XWW08]. Large-time [ZW04]. large-update [CKL07]. largest [Den08, LYC07]. laser [CPP08]. Lastovetsky [Ano04s]. late [GSS04]. latent [SP09]. Lattice [AS09b, AVBW08, CM09, ITR08, KIXF09, OKCO08, SM08, VM09, WB09, dG09, AZB08, AB09, Asi08, BJK^{+}09, BHS^{+}08, CJ08, CJR09, CLL09, CLT^{+}08, DWL07, DS08, Dub08, HKL09, IMLF09, JTRu09, JY08, JY09, KOCO08, KMK09, Lee09, LY09c, MHTR08, MKK08, STK08, SP08c, SGR^{+}09, TBDL09, TK09, VRJ09, WYZY01, Yon09, YG08]. Lattice-Boltzmann [SM08, MHTR08]. lattices [Fly06, GGG03, LSZW07, Pei02, SW09, Wan00b]. Laurence [Ano00-54]. Lauricella [HS07]. law [BCGP03, FZ07b, Gal08b, GMŚ07, KK06c, Lin04b, PL05a, Ung07, YLZ08a]. laws [CS02b, CCZK09, Gos00, OIL09, PLD02, Soh05, Ano03-37]. layer [DMP03, ICM09, KP01, LT04, PM05a, Roo08, Tur07, Xu07a, XL09, dR03]. layered [BDO04, BD04b, ICM09]. layers [Kel06, LTWM08, NVAR03, TBDL09]. layout [PS08b]. LBGK [SDDC08]. LBM [SS09b]. LCD [KS00b]. leading [HM04]. leap [Sny00]. leap-frog [Sny00]. Learning [An03, LY06b, WZ08b, ZYY09, CCF06, GSNDB^{+}08, HCTC03, KY07b, MCC09, PYLX08, SS09a, Wan08b, WWW^{+}09, YH11, Zha02a, ZL07c, Ano03-61, Ano00-39, Ano00-40, Ano03-48]. Least [BB04, BG04b, FR04, LL05, Xie00, ZZH03, Bed02, BM05b, CJL04, Che04d, CLS07, DW02a, DD08, KN04, KL06, KU01, Liu08e, LU09a, MS04a, Mur05a, SDZ06, Tan07, XZDD09, Yan03d]. Least-square [BB04, ZZH03, BM05b]. Least-squares [BG04b, LL05, Xie00, CJL04, Che04d, CLS07]. least-squares/penalty [CLS07]. Lebeau [Phu02]. Lebesgue [OF08]. lectures [Ano03-33, Ano03-49]. LEDA [Ano00-41]. Lee [Tan03a]. Leebaert [Ano00-76, Ano00-77]. Leen [Ano01e]. left [CL01]. leg [XT01]. Legendre [GW06, HZ03]. LEGO [Ano00-82]. Leibniz [JSS01, TWS00]. Lemma [Pal01, KM01, LCJL03, Pen03]. Lencevicius [Ano01d]. length [Gil00, GSS04]. lengths [AMM00, Gil00, KC05b]. Leonardo [Ano03-50]. lesions [BHCØ07]. Leukocytes [SASHS09]. Level [DL06, OK06, Ras06, BJK^{+}09, CGPB08, IDM08, KO09a, KM03b, KH09, LLW08, MKF09, NUN06, SHWF05, Sir00, VDMK08, YZHY09, Ano03-51]. level-2 [LLW08]. levels [BFM04]. Levitin [HS04a]. Levy [Ano03z, Çin03]. Lewin [Ano00p]. Lexical [PCA09]. lexicon [PCA09]. LFOPC [Sny00]. Li [Tan03a]. Liapunov [Dib03, Nes01]. Liaw [TJ01, ZW08]. library [YSP^{+}08]. Lidstone [Pal01, WJA08, WX03b]. lie [AK03, BENT06, CZ09a]. Lieberman [Ano03-71]. Liénard [CR06]. Life [Ano00p, Ano03k, CHW07, SBSY06, Ano00-90]. Lifshitz [LWY06]. Lifshitz-type [LWY06]. light [AA08, Gal08b, Iac07, WGF09]. light-weight [WGF09]. Like [AI06, OSS07, WG06, Arg00a, CGY08a, CGY09a, DY05, DYZ08, DR02, FF07, GL08a, GL09a, GMŚ07, KB04, Lin00, MP07a, MM08c, PV00, PPI03, Ryo00b, SP08a, Ung03, WH07, YZ02b, Yon09, ZGY05b, ZGY07b, ZAY08, dSRGRMRL08]. likelihood [BB01, CS01, Hur00]. likesystems [CP04]. Limit [CHM06, QZ05, SS06, ZHZT06, ANZ09, Blo08, CS04b, Din04a, DKO03, DAKO06, DF02, DL08, FWL04, GL06, GY09, HL04, HL07, LHP07, LC02, LPSS02, LL02b, LL03c, Mic09, Phu02, WK03, WH08b, XHZ05, XLL08, ZL07a]. limit-circle [SS06]. Limit-point [QZ05, SS06]. limited [KK06c, Lin06b, Saw03, XWW08]. limited-flow [Lin06b]. limits [Ano03u, Ano03-32, Pit01]. Lindsey [Ano00-51]. Lindstedt [Mar00]. Line [Ano00-42, Bal04, CCSL04, DMP03, DM08, He00b, HM01, KOCO08, LG09a, MZ09, MO01, yNyLjZaZ08, SS07a, SH01, TGS07, WW04b, ZY02, Zha09d, Ano00-42]. Linear [DLL08, Liu06a, Wu03, ZP01, ABBD01, AO01b, Agr08, ATY08, AJ09a, AJ09b, AD08, Ano04m, ASS08, Ash05, AO08b, AHK00, BB08a, Bal02, Bar01a, BD03, BBI05, BG11, BZ00, CG00, CC09a, CJ05, CCS08, Cap00, CDM01, CDM03, Cha01a, CZ03a, CP04, CS04b, CCK03, CCKH07, CKL07, CK04, CNPP08, CC02, De 02, DD04, DP07, DR01b, DM07, EA03a, EM09c, EGR03, FB05, FM06, GSZ02, Gan06, GS01, GOMA07, GW01b, Gup03, GK03, Han04, HLW07, HLP05, HBP06, HCL07, HRM08, HF00, HCLF01, HL08c, Hua08a, IS04, Isk04, Isk08, IJB00, JKR08, Jia00a, JEP08, JT07b, JL09b, JH09, Jun01, KGB07, KT02, KAKK00, KDC08, KK08b, Kop01, KU01, KH09, yL08a, LG08, LLY01, LC06b, Lev06, LZ04a, LHL07, LY08c, LLJ09a, LW02, LSZ08, LX07, LHS08, Luo02, LLL02, MS02, MSN^{+}00]. linear [MAK02, MAK03, MO07, ME08, ÖZ07, PCE^{+}05, Par04a, Pet04, PP04, Pit01, Pit03, QZ05, RSD09, SDK09, Sal08, Sha09, SHWF05, SSLL09, SOW04, SS06, STH08, SL06c, TRG01a, TL06, Tia08, Tsi02, WWWL08, Wan09b, Waz07c, Waz07b, Waz07a, WSZ02, Wu08a, XDD09, XA09, Yan04d, YL09a, Yus09, ZZ00a, Zha04a, ZL04, ZFYS05, ZHL05, ZLCM08, Zha08d, ZZ09b, Zho01]. linear-quadratic [WWWL08]. Linear-time [DLL08]. linearity [Nko00]. linearizable [JK09]. linearization [TLP^{+}09]. Linearized [DFY02, He03, PHW03, GL03a, WK05]. linearly [JWC08, WY02]. lineof [KKS^{+}04a]. lines [Kot04, SD08]. linex [Jah04]. link [FKM05, FY01, Yav00, YF04]. linked [AETEZ07a, EBVS06]. Linking [PKM^{+}06]. links [Ano05g, FY03a]. Linux [Ano01f, Ano00-39, Ano00-40, Ano00-67, Ano00-68]. Liouville [AU09, CH00c, DH03, Jum06, KM07, KK08b, LR03, LS09, Ma07, UA06, VJ04]. Lipatov [Ano03-82]. Lipschitz [Csa01, NC09, XD09]. Lipschitzian [PW06, KK01, BCDS09, ZP02]. Lipsman [Ano00z]. liquid [AVO08, Kor00a, Kor00b, LLCS02, Nik06, Sam07]. Liquids [BBB08, Gre03]. List [Ano00-97, Ano00-98, Ano01-30, Ano01-31, Ano02g, Ano02h, Ano03-93, Ano03-94, Ano04-39, Ano05h, Ano06j]. Liu [Ano04d]. Lloyd [Ano03-66]. LMI [GHC09]. Load [MR06, CA09, HTZY03, KM03b, LWLJ09, NLLL09, PLL03, SMOK05]. loads [CCF09, MYR09, SMO03]. lobsters [MP05, MP08b]. Local [Arg00a, DGPT08, FS07, Kan02, LB04, NWL08, RS09, SL07a, CK03, ER07, KLJ09, Kr05, SV06, SBPD08, Sun06a, Sun06b, VG08, WW09b]. Localization [IS06, BCD08]. localized [PL01c, WKZ08, XX08]. locally [AA09c, CLMM06, DPJ03, KK06b, Yan04b, YY07, ZG02, ZZ09b]. locate [HL08b]. location [AGS09, Che01a, LS04a, ST00, TOH03, ÜL02, WI08]. location-allocation [WI08]. lock [EW00a, Eng04]. Log [RW03, Syn08]. log-return [Syn08]. logarithm [FZ07b, Lin04b, PL05a, YLZ08a]. logarithmic [AC02, Dra03a]. logarithmo [PR01]. logarithmo-technia [PR01]. Logic [BFFL08, Ano00-43, Ano03-83, LL03b, VV08, WHS08, Zha08a, Ano00q, Ano00r, Ano00-47, Ano00-43]. logistic [CS04a, HY08, Li08b, LYW02, Mur07b]. logit [CCF06]. London [Ano01w, Ano03u, Ano03-48, Ano03-72]. long [BS00b, NLLL09, TG03, Yan03a, YB07a, ZTW09, ZTH09]. longest [Pla07]. longest-edge [Pla07]. loop [EW00a, Eng04]. loops [KP05b, Lu09c]. Lorenz [LLXC03, WY09b]. loss [AFI06, GLR05, Jah04, Pan03]. loss-resilient [Pan03]. Lossen [Ano03e]. lossless [AB08c]. Lot [KH02b, HTKM09, HY06, MGG05, WM09, XD02, YE01]. lot-sizing [HTKM09, HY06, XD02]. Lotka [ADN02, BNH07, CF04, GY09, Huo04, LDZW05, LY08b, LL02b, LL03c, NTHP09, TL00, WW04b, Wan09e, ZSG04, ZLZ09, ZZD08]. Louis [Ano04c]. Lovink [Ano03-89]. low [CWLC09, GW03a, PCRR09]. low-complexity [CWLC09]. low-order [GW03a]. Lower [Li01a, PRA01, TK02, AOLL04, BHNO04b, COEP00, Jia01b, Jia01c, Jia02, LH08, SA05, SAS06, Wu01, ZJ08, ZQC08]. LS-Arnoldi [HBP06]. LSFD [SDZ06]. LSLR [LY04]. LU [WH09]. Lubich [Ano03-42]. Lucja [Ano01r]. Lurie [XLL04, YZZ^{+}08]. Lusk [Ano00-88, Ano00-89]. Lyapunov [LC06a, Ash05, GK03, KLP09, KZD03, LT05, PS04, ÜÇT08, Wac08, WLWC08, ZLD08]. Lyapunov-type [ÜÇT08]. lytic [WWL06].

M [ASESESA09, Ano00o, Ano01i, Ano01r, Ano01-28, Ano03m, Ano03p, Ano03r, Ano03-78, Ano03o, Ano03-30, Ano03-38, Ano03-43, Ano03-48, Ano03-69, FP06, Jai05, KA00b, GS07b, ASESESA09, GLZ00, Gup02b, Gup02a, Jai05, KA00b, KA02, LL06b, LWY09, Wan04a, WX09a]. M/G [GLZ00]. M/G/1 [Gup02a, KA02, LL06b, LWY09, Wan04a, WX09a]. M/M [Gup02b]. M/M/1 [KA00b]. M/M/c [ASESESA09]. M/M/r [Jai05]. MA [Ano00e, Ano00j, Ano00f, Ano00k, Ano00l, Ano00m, Ano00n, Ano00-29, Ano00-76, Ano00-30, Ano00-77, Ano00-38, Ano00-43, Ano00-54, Ano00-59, Ano00-69, Ano00-83, Ano00-84, Ano00-88, Ano00-89, Ano01e, Ano01y, Ano01n, Ano01u, Ano01t, Ano01x, Ano01-28, Ano03-36, Ano03-77, Ano03j, Ano03k, Ano03l, Ano03t, Ano03-78, Ano03o, Ano03v, Ano03y, Ano03-79, Ano03x, Ano03z, Ano03-28, Ano03-81, Ano03-32, Ano03-38, Ano03-41, Ano03-85, Ano03-83, Ano03-47, Ano03-50, Ano03-53, Ano03-86, Ano03-55, Ano03-54, Ano03-60, Ano03-61, Ano03-63, Ano03-87, Ano03-71, Ano03-89, Ano03-90, Ano04d, Ano04h, Ano04n, Ano04l, Ano04-27, Ano04-28, Ano04-30, Ano04-31, Ano04-29]. machine [CCP09, He00b, KC03a, KY07a, KY07b, LHS08, MCC09, MH09, Pio00, SL00b, TH02, Wan08b, WWW^{+}09, Wan09b, YH11]. machines [XXLD04, ZL07c]. Mackey [BB06]. Mackey-glass [BB06]. MacQueen’s [DW02b]. macroeconomics [Ano04-27]. macroelements [He00a]. macroscopic [AO09, BBH07, BB09b, Dog08]. macrotransport [VM09]. Maekawa [Ano03-69]. magic [dG09]. magna [Gou01]. magnetic [Aro08, HAKA07, SWS05]. magnetized [TG03].

Magnetohydrodynamic [INP08]. magnetohydrodynamics [BG01]. magnetosphere [HWHG08]. Magnus [Ano00-28]. Magnusson [Ano00-36]. Mahony [YM09]. maintenance [AH03, HY08, MS09, NQHN03, QNN03]. Majewski [Ano04o]. maker [Ehr02]. makespan [LHS08]. making [Ano00-64, Che08b, Ehr02, Eke02, ESNP06, HT04, LYL00a, LCL00, MRB02, SJL08, WS04a, WL07b, XZ08b]. making-involving [Ehr02]. management [Ana06, Ano00-38, Ano00-50, Ano01x, Lin08b, LNL08, LLT09, Yan03b]. managing [LRLM^{+}01, Ano00-44, Ano00-45]. Mandelbrot [Ano03-35, Ano03-40]. MANETs [LJSH09]. Maneuvering [Ano05g, FY03a]. Manfred [Ano03-61]. Mani [Ano00e, Ano00f]. Manifolds [Lug01, AP03, DBX04, Far01, Gar01, LW03, You04]. manipulator [FY01, KL01, Yav00, YF04]. manipulators [HS02]. Mann [Ano03-57, SW08, KK01, MCQK09, Pan07c]. Manolios [Ano01h]. Manolopoulos [Ano00d]. Manuel [Ano03w, Ano03-60]. manufacturing [CPP07, CL00, Ano03-69]. many [AO01b, HF00, HY08, KFMH06, MZ04]. MAOR [HL03a]. map [Leu00, PL01b, WY09a]. Maple [Ano03-34, PG06, PG08, Ano04j]. Mapping [SMS^{+}04, AZCK05, DWL08, Dra03a, fHCH09, Jin08, PS07b, RNM00, WS07, YLZ08b, ZYL06, Ano03-66]. Mappings [yL06, TS06, AO01b, Ano00u, Bai01, BW06b, BCDF02, BA08, CWY08, CGY09a, CCRMRF00, CZG04, CKQ08, CyL08, Dha07, GJM03, Gu07, HCLJ00, HGH02, HF04, HN07, HS09, IK03, Inc09, ID05, ID08, Jin07, KH08, KK01, Kir08, LCV06, Li07, LV03, Liu09c, Pan07c, PY09, PW06, RR09a, SW08, Son08, Sya01b, SL06b, SL06c, Tia01, Tia05, TCH07, THD08, Ver03, Wan07a, YHT04, YYZ07, ZGY07a, ZA03b, ZYL06, Zho06]. maps [AO01a, AA09c, ĆRRU08, udK07, HLY08, Lan04a, Lan04b, OS09, ZP02]. Mara [Ano00-64]. Marc [Ano03-78]. Marchetti [Ano00o]. Marchetti-Spaccamela [Ano00o]. Marco [Ano04-28]. Margaret [Ano01-28]. Maria [Ano00g, Ano00h]. marine [Ano03-80]. Mario [Ano03j]. Marius [Ano04q]. Mark [Ano00e, Ano00f, Ano00-63, Ano03k, Ano03v, Ano04-32]. Marked [DL06]. marker [HX09]. market [KWT02, LTYB09, LLL08b]. marketplace [KC03b, Ano00-76, Ano00-77]. markets [AAK08, Ano04-27, BGMB02, BM06]. Markov [Ano04i, EZ08, KKY06, KL08a, KL03, LWZ09, MR06, Nak06, SCF^{+}09]. Markov-switching [EZ08, SCF^{+}09]. Markovian [KCD05, TEB06, WLWC08, ZLD08]. Martin [Ano00-28, Ano03e]. mass [FP02, HCG08, HW05c, NP07, VC03]. mass-constraint [HCG08]. match [CS00b]. Matching [Hua06, Lee01a, TBDL09, YAVB08]. matchings [YW08b]. material [DH08a, De 02, ZMM07]. materials [MVS^{+}07]. Math [AR10, Ano03u, Asl11, Ata09, fDxZ11, ES08b, ID10, ID08, JY11, KÖ08a, KÖ08b, Liu08b, Liu09a, NLL08, SW08, Ver12, Yan08, Zho08]. MathCAD(R) [Ano00-58]. Mathematica [Ano00-74, BDGR01, Mar00]. Mathematical [Ano00-32, Ano00-33, Ano00-35, Ano00-66, Ano03-35, Ano03-84, Ano03-58, AVAGP08, Bel06, CCK08, Fei07, ISK00, KPSW06, WS08a, AACLM08, Ano00-58, Ano04i, Ant05, AMC08, CCS08, FM07, HL09b, LMA08, MG01, MP02b, MG03, MD06, RRM08, Ros03, STV08, SDK01, SV08, Tao06, Wan02a, ZhYL05, ZXL07, Ano03-52]. Mathematica(R) [Ano00-46]. mathematician [Ano03u]. mathematicians [Ano01-28].

Mathematics [Ano03-46, Ari09c, EM09b, Ano00-59, Ano00-90, Ano03-35, Cse06, RM07a, RM07b, Ano01l, Ano03-76, Ano04x, Ano00-47]. Mathieu [PS09, dM06]. Mathieu-type [PS09]. Matlab [ANV00]. MATLAB [Ano00z]. Matrices [Che06, AK07, BL01, Bar01b, Bax02, Bax06, DH08b, FLR05, GS01, Gyi00, Has04, HL08a, HP02, HLWT04, Jog04, KP05a, KDKS01, LZ09, Lin00, LYC07, LH08, LCC08, LS02, MRT05, MSS08, MM08d, PMR^{+}08, Pan05c, PL03, PV01, PSES07, RV05, RR03b, Roj08, SN01, SH06, SV09, Sot02, Ter00, Wu08b, WH09, Xie00, YLL08, ZM94, ZHL05, ZZH03]. Matrix [DGR04, Liu06a, MSN^{+}00, Pan03, PKM^{+}06, dM06, Ada04, AK07, Ano00-94, Ano03-39, BSS98, CLC01, CFY08, CH03, CI06, DJL04, DSHS05, DH08b, FT01, FH03, HR08, HRP04, Jia01a, JSKE05, LH05a, tLjW08, LL05, Liu08e, MZ07, Min08, MRÁR08, NOY01, PIM^{+}08, Pan07b, Par04b, PV01, RH09, RSR01, RSE01, Ser00, Sot01, SMS^{+}04, TB09, TT09, TDS08, Wan05a, Wan05b, Wan07b, WXM09, WHD09, XDD09, XZ09a, YFT04, YD08, ZYCI09]. matroids [LL08a]. Matt [Ano00-67, Ano00-65, Ano00-68, Ano01h]. matter [BS08]. matters [Tun05]. Matthew [Ano00-69]. Matthias [Ano00-67, Ano00-68, Ano03t]. Maud [Ano03-52]. Mavle(R) [Ano03-30]. max [CRK02, EG02, SM09]. max-min [SM09]. Maximal [AMM00, HHJS03, mLpW08, QW09, Ver08b, Ver08a, Ver12, Wu00, Zho06]. Maximality [CZ06]. maximizing [HTZY03]. Maximum [Hur00, Jia09, Bal00, BB01, Hor08, Kov05, MPV02]. maximum-minimum [Hor08]. maximum-norm [Kov05]. Maxwell [DMSV00, FAF09, NSB08, Nic04]. Maybury [Ano00e, Ano00f]. MCA [PYLX08]. McCarty [Ano00-39, Ano00-40]. McConnell [Ano03-57]. McCullouch [Ano03-62]. MCDM [CL09]. McOwen [Ano03-59]. MCSE [Ano00-52, Ano00-53, Ano00-72, Ano00-73, Ano00-92, Ano00-93]. MDOF [MB08]. MDS [ZZ06]. MDS-based [ZZ06]. Me [Ano04l]. Mean [CD06, CJV07, HHW08, IWL04, IW06, SA05, APV02, Bao03, BPvZ03, CD05b, CCRMRF00, CC05, HCWC07, KN08, Kua08, PLP09, RF09]. meaning [Ano03-90]. means [Ana07b, BP09, BCDF02, CD03, DW02b, GPL08, KWHC05, SJJ09, CLJ09]. measurable [BA08]. measure [BLD07, BC01, Bri00, DK05, KC05a, LR03, LXH05, OF08, PLG^{+}09, PL03, Wan02e]. measurement [CZ05b, SCF^{+}09, TY03]. measurements [SN03, ZS04, Ano00-51]. Measures [PPP^{+}06, BCA09, CC08a, YL09b]. Measuring [SKHK06, Gal08b]. MEBDF [ACD01]. Mechanical [DFY06]. Mechanics [RM07a, RM07b, Che04a, OM09, Öze08]. mechanism [LLT09, MT00]. mechanisms [Ano03-65]. Media [Ano03-88, WZE^{+}06, GMS08, LY08d, Mar09, NB08]. medical [LRLM^{+}01]. medium [Bae07, BS08, CM06, CDJ04, DS05a, HHK06, IUCR06, JPV02, MK08, NMD09, YW03b]. medullary [MT00]. meets [RGTT03]. Mehlhorn [Ano00-41]. Mellin [BJ00]. Melnikov [DZ05]. melt [ZDLH08]. melting [DD03]. Members [Ano01k]. membership [HF00, yL08a, XTZ08]. membrane [SAS06]. membranes [Win07]. memories [CCHK08]. Memory [SS07a, JZT07, Sha09, XWW08]. Menlo [Ano01r]. Menten [Ber05, Huo05]. Menzel [Ano01u]. Meredith [Ano03s]. Mergers [Ano01q]. mergesort [CNK00]. Meromorphic [Xu03b, Ban07, BD07, CYX09, EAA09, FF02, Liu08c, WSZ09, Xu08a, YXZ09, ZCL08, ZL08b]. meromorphically [Aou08, AY09, WJS09]. Mersenne [Tan03b]. Mertens [Cha08].

Mesh [CN01, PPP^{+}06, AK08c, BT03, CLX08, FL03, HLP03, IJB00, KP01, She02, UHuI09]. mesh-free [UHuI09]. meshes [FL05, Hor08, HWC04, LXZ08, Nko00]. Meshfree [SV06, BBCE09, YLB06]. meshless [BJBL03, CT02a, Dua08, Mar05a, WHL02]. Mesoscale [HHL^{+}06, ITM08]. Mesoscopic [KL09, LKL08, MKSB08]. message [Ano00-88, Ano00-89]. message-passing [Ano00-88, Ano00-89]. metabolic [LMA08]. metabolism [Ano03-41]. Meteorological [Ste06]. Method [AC06, BUP09b, DS06a, Kot04, LET06, LW06, Nie06, Ras06, SPE06, Ste06, SK06c, YC06, AETEZ07a, AETEZ07b, Abu01, ADDL04, AS03a, ABBD01, ABBDP04, AW01, AOLL04, AM09, AVO08, AZB08, AN06, ACCMJ07, AU09, AD03, AE07, Ano00-94, AMC08, Ari07b, Ari09a, Ari09b, Ari09c, AO08a, AK06, Asl11, AC09, AV00, BMHV09, BSS98, BNP05, BSR02, Bal00, BCJ08, BBL05, BS06, BNH08, BNH07, BN05, BJK^{+}09, BB09a, Bel09, BGAM09, BCS03, BHNO04b, BENT06, BS00b, BM09, BPM09, BG07a, BG08, BBA09, BA09a, BA09c, BG11, BI01, Bla04, Bla07, Bou00, BB08b, BLS05, CC09a, CJL04, CYG06, CL07a, CJR09, CLJ09, CJO05, CL00, Cha01a, CK08, Che04c, CX04, CWLC07, CHZ09, CXZ09, Che00b, Che04g, CLS07, CC04b, Chu05, CN08a, CJK09, CA09, Das09, DS07, DZQ06, DWL08]. method [DD08, DKO03, DZ05, DS08, Dua08, EP04, EGZ04, EGA03, EM09b, FK01, Far01, FFK02, FJ05, GW01a, GAT07, GTJ07, GGA09, GY08, GS01, GL09c, GLC09, GSN08, GJ07, GS06, GL08b, zGtM08, GL00, GHJ06, GK00, Guo03, GW06, Ha01, HS04a, HJW05, HLW07, Han08, Har01, Has04, HS02, HLSL03, HBP06, HWX07, HW07, He09b, fHCH09, Hem08, HR02, Hon02, HRM08, HL03a, Hua04, HGS07, HZ07, HHC08, HL08b, HS09, HLP03, ITR08, Inc09, JG07, JWC08, JZQZ07, Jia00a, JCZ09, JZT07, bJQxC09, Jia01c, Jia02, JYDH02, JDH03, JDY08, JH09, JY09, KP01, KA09a, KGB07, KK02, Kan04, KA09b, Kay04, KYJ03, Kha05, KHN09, KLS06, Koi07, KS03b, KÖ09b, Kot01, KZD03, KW03, KMK09, KI04, LHC08, LY09a, LW09b, Lee04, Lee09, Les02, Les05, LZJ04, LUZ05, LDK^{+}06, LY08a, LT09a, Li09b, LW02]. method [LT04, LT01, LLTL05, LCWL06c, LZS07, LSZW07, LCC08, LN08, LW08b, LLS^{+}09c, Liu09a, LY09b, LGY09, LPvST07, Lu07, Lu09b, MYG09, MS04a, MH02, MCC09, Mar00, Mar05a, MH07, ME00, MCQK09, MHTR08, MM05b, MAK03, MQ06, MO07, ME08, MEA09, MS04b, ME03, NS07, Nic04, NRL06, Noo00a, NMD07, NMD08, NMD09, OP01, OM09, ODO03, OIL08, OIL09, ÖY07a, ÖY07b, ÖY08, Pan05a, Par04a, PS04, PV00, PPI03, PST08, PT07, PDRR08, QJZ09, Qua01, RS09, RD07, RTS07, Ram05, Rap01, RNM00, RR03a, RR08, RW03, RPO03, Ros03, Ryo00b, SDVS09, SD09, SNT08, SNG09, SG07, SGM09, SSSH01, SERG09, SANH08, Sal08, SV06, SD07, SD08, SP08a, SC00a, SS07a, SDB09, Sho09a, SHI09, Slo07, Slo09, SZ09a, SKG01, SH01, SH02, SFL02, SF07, SP08c, SGR^{+}09, SY09]. method [SHWF09, SZ09b, SB00, SBD^{+}00, Swe07, SAB07, SNA09, SK09, Sya01a, TKC02, Tan03a, Tao06, Tao09, TD09, TK09, TV07, UHuI09, ÜL02, VR02, VA04, Van07, VDMK08, WYH00, WXW01, Wan02d, Wan02c, Wan03a, WDLP04, WL05, WF05, WLL06, WZ06, Wan07a, WJA08, WML08, WWWL08, WDLP08, WL08b, mWmZ09, WW09c, WW09b, Wan09a, WFC04b, Waz00, Waz05a, Waz07c, Waz07b, Waz07a, WY01, WZ09, WX00a, WX00b, WXS01, WSZ03, WS04b, WK05, WH07, WZ08c, WS08b, Wu09, WYSS09, XWW08, Xu07a, Xu07b, XZ08a, XYC^{+}09, YHS08, YH04, YV05, YL09a, YW09, Ye02, YM09, Yil08, Yil09, Yon09, YLB06, YTCF08, YLD09, YW03b, YG08, YL08c, hYL09, YB07a, Yus09, ZGY07a, ZZ03a, ZL06, Zha07b, ZH07, ZJ08, Zha09e, ZTW09, ZH09, Zha09c, ZX09, Zha09b, ZL07b, Zhu05, dZmClQ09, ZCM01b, ZZD08]. methodology [Ano00-38, Ano03-75]. Methods [Ano03-59, LCL06, Sar06a, SW06, AVLPN04, AP08, AM08a, Alb02, AB05, Ano00y, Ano03r, Ano03-51, Ano04f, Ano04g, Ano04i, Ano04w, Ano04v, Ano04y, Ano04z, AG03, AS09b, Ara01, Arg00a, Arg00b, AO09, BUP09a, BJ02, BK04, BT03, BJBL03, Bed02, BBCE09, BDGR01, BX02, BBRS08, BM05b, BXFZ09, BG04b, BC01, BC08, BPSW00, Boy01, BMM^{+}06, CCP02, CLM03, CCS08, CL02a, CPP07, CPP08, CM04, CGY09a, CKY04, CKL06, CQT02, Che04f, Che04a, CWZ03, Che08b, Chu05, CM02, CN04, Csa01, DRR09, DH07, FT01, FGH08, FL03, For01, FT05, FM05, FFXL05, GSZ02, Gan06, GNM07, GGE09, GL03a, GW03a, Gir03, GPPRM01, GPMPR03, GNR05, HL02b, Han06, He02, HL09a, HKL09, HR05, HLC05, HL06, HCLF01, HC02, Hua08a, HZ09a, HSS01, Ikh04, Ikh08, Jan03, Jan05]. methods [JK01, Ji02, Jia01a, JZ08, JMR00, JY08, JR01, KZ02, KT08, KS00b, KL09, LF03, Lei02, LH00b, Li00a, Li01d, LL02a, Li02, LCGX06, LHL07, LJ08a, LLJ09a, LY04, Lin08c, LW08a, LN07, LHH08, LKL08, LYL07, Ma09a, MKF09, Mar09, MZ07, Més05, Min08, MF01, MEG08, MKK08, NF03, Net05, Net07, NW09, Pao01b, PF01, PLP09, PLL03, PY09, PIR04, PR06, PB02, PDBH03, RR09b, RDS00, Sar06b, Ses03, SR03, Sha04, SDZ06, Sid08, Sim01, Sir00, SYL05, TB08, Ter00, Tho01, Tia08, TP03, Tou05, Tsi02, Tsi03, Ver01b, VM09, VB01, VAR03, Wan00e, Wan02e, Wan03b, Waz05c, WBSP08, WK03, WF01, XT01, XT02, XT04a, Xia07, XWZ04, Yan03d, YYZ07, ZM94, ZL02, ZZJ^{+}04, Zha04a, ZHL05, Ano01d, Ano03-84, Ano04m]. metric [Din04b, Rad09, RR09a, WHS08]. metrics [CHL09]. MEW [XYC^{+}09]. Meyer [Ano03-36, GJQ07, Zen00]. Meyer-König [GJQ07, Zen00]. MHD [INP08, CMS04, HASA07, HWHG08, KHA07, MM07a]. MIC [AMV08]. Michael [Ano00x, Ano00-30, Ano00-52, Ano00-53, Ano00-66, Ano00-72, Ano00-73, Ano00-92, Ano00-93, Ano01t, Ano03-35, Ano03-86, Ano03-73]. Michaelis [Ber05, Huo05]. micro [Bab09, VM09]. microaggregation [DFSS08]. microanalyis [ADH^{+}03]. Microarrays [Ano03-53]. microchannels [WSG08]. microcirculation [SASHS09]. Microflows [DHY06]. microgravity [DS08]. Microlocal [AHNV01]. micromagnetics [LWY06]. micropolar [AH08, ICM09, INP08]. microprocessor [IYN03]. microscopic [BBH07]. Microsoft [Ano00-44, Ano00-45]. microstructure [BM00, OS08a]. Microtransducer [Ano00-48, Ano00-49]. microwave [IDM08]. mid [TXC02]. mid-point [TXC02]. Middle [NLLL09]. Middle-long [NLLL09]. migration [VTR04]. mild [BN00, HR02]. Milekic [Ano00-83]. Milne [Wan09c]. Milstein [WZ06]. Mimetic [BC01, CHS02, HCL07, HSS01, HS04b]. MIMO [LBW09]. Min [EG02, CRK02, SM09]. Min-max [EG02, CRK02]. MINDSTORMSTM [Ano00-82]. Mineola [Ano03p, Ano03q]. minihypers [HH00]. minima [Qua01]. Minimal [CCS00, YSP^{+}08, Bel05a, HHJS03, HLCW07, KN04, Li04a, Lin07, QNN03, SY05, YW08b, ZQC08]. Minimality [WH08a]. minimax [Ant08, JC05, Lan04b, Mis05]. minimization [BC00, CWZ06, LT01, SH01, SHWF09, WY02, WI08, WYSS09]. minimize [MS09]. minimizers [HYZ06]. Minimizing [LW09f, LHS08, She04b, Kar01, TC09].

Minimum [CFR03, XXLD04, GL08b, HTZY03, Hor08, LL05, LT09b, RNM00]. minimum-norm [LL05]. minimum-risk [LT09b]. Mining [LGH09, Ano03-31, Ano03-43, DY09, KS07, LCY05, SYHC06, WK04, ZY07]. Minkowski [POS03]. minmax [Zal03]. Minty [GL08a, LCJL03]. Mirakjan [Gal08a]. Miranda [Ano03j]. MIRK [CM04]. Miroslaw [Ano04o]. miscible [LY08d]. Mises [NWL08]. Mistikawy [Ari09c]. Mitchell [Ano04l]. Miwa [HO00, TGH02]. Mixed [CMS04, DHY06, GW06, KPP00, LET06, Mou02, ACJ07, BCS03, BBP09, BMM^{+}06, CDJR03, CJO05, CGY08a, CDJ04, DY05, DYZ08, DM07, ET09, ETAJ09, FM06, HM00, HBCK00, JPS09, JC00, JSF00, JP01, KK08a, KS00b, KP06, KLW09, Kru05, LS08, LHW04, LWG04, Man05, MS04b, Noo00a, PZ07, PXW09, PY09, SMMM02, SF07, WLW07, Xu08b, YLD09, YL04c, ZGY05b, ZGY07a, ZGY07b, ZAY08]. mixed-coding [LHW04]. mixed-integer [LHW04]. mixing [DS02, WSG08, ZL07c, dR03]. mixture [Asi08, LWZ09, Mas07, Sam07]. mixtures [BB09b, Sam07]. Mizoguchi [Kam09]. MKZ [ZL08a]. MKZ-Bézier [ZL08a]. MM [YY07]. MO [Ano04c]. mobile [Ano03w, KR06, KYNI06, YMI06, Ano00-50]. Möbius [Ung03]. mod [Ben00, Cha01b]. mode [KM03b, LB04, SMOK05]. Model [ACLL06, CZLN06, DFY06, HHL^{+}06, KNP06, Wan06a, AS09a, ADDL04, AACLM08, ADN02, Ano00-94, AMC08, AVBW08, AVA01, Bab09, BSS98, BBGG08, BG01, CL02a, CLJ09, Cap07, CS01, CH05b, CWL07, CW08, CSZH09, CCP09, CGPB08, Cui02, DD03, DW09, DCR07, FW00, FM07, FM06, GW01a, GT08, GS07a, GMG09, GPAA^{+}09, Gou01, GSNDB^{+}08, GLZ00, Gup02a, HR07, Hau06, HX09, HY09, Hoh01, HLCW07, IKY06, ITM08, JAM02, KH02b, KH02c, KOCO08, KWT02, KPSW06, LP06, LGZ01, LW05, LC07a, LCL08, LYW02, LCT08, LY09b, LLT09, LHLW07, MGG05, MM00, MP08a, Mic09, Mie01, MM08b, MG01, MP02b, MG03, MD06, MF01, MRP08, Mos08, MJ02, NTTS09, Nik06, PKM08, PdlF07, PdlF09, PLP09, PS08b, PLG^{+}09, Pio00, QNN03, Qiu08, RS06, ROY01, SA02, Sam07, STV08, SD07, SDDC08, SCF^{+}09]. model [SDCL08, SZJ09, TL07, TAT07, TYW05, Ung00, Ung01, VNG04, Wan01, WCL02, WM04, Wan06c, WWL06, WI08, Wen02, WHL02, WW02b, WLZ07, WLW08b, XC06, XLL08, XCD07, XZ08b, YY07, YZHY09, Yav06, YWZ03, ZWCW00, ZSPZ08, Zha08c, ZM02, ZMM07, ZZ02b]. model-based [PLP09]. Modeling [ASL06, CZLN06, CGM08, CS08, Dar08, GMSRP08, JSGP08, KR06, MKF09, TDHL09, ZZJ^{+}04, ZDLH08, AK08a, Ano04k, AS09b, AVAGP08, Asi08, BJBL03, CKCG03, CCK08, ESNP06, Kor00a, LYF09, MS08, MKSB08, MVS^{+}07, RM08, TB08, VTR04, YX09b, ZZZ09, Ano03o, Ano03-86]. Modelling [BC08, CGPB08, FM08a, FF07, LMSH04, Tur04, WSG08, Yav07b, CPP08, CCC08, Gre03, ISK00, IUCR06, Mar09, MUHdH03, RZ03, SBPD08, ZBS03]. Models [HLY06, LYL00b, MR06, Ste06, TA06, ACV03, AMS08, AAK08, Ano03v, Ano04m, AB09, Bar01a, BBM08, BJK^{+}08, CCJ08, CM01a, CCS08, Col04, DPU07, DRR09, DLB03, DSHS05, Dog08, Eke02, ETH05, EZ08, FM06, GAA06, HM03a, HXG03, HZ05, Ili04, IY06, JB05, JS04, LMA08, Li09b, LI00c, Lo 02, Mas07, MT00, MKK08, Mur07b, NWL08, PdlF09, PDRR08, PM08, RRM08, SK06a, SDK01, Saw03, SN03, Syn08, Ter00, TDS08, ÜL02, WS08a, ZM94, Zal02, Zal03, Zla01, dG09, Ano00-51]. modern [Ano00-66, Ano04p, TBR01]. Modification [VAR03, CN08a, Wan07b]. Modifications [BS06, DGVP06]. Modified [AETEZ07b, BXFZ09, Jum06, XWZ04, ZHL05, BFM04, CM01a, EGR03, GSN08, HS04a, KS00a, Koi07, LY08a, LWY02, LWY04, Lu09b, LHR09, MCQK09, MP08a, ME08, MEA09, ME03, Noo02b, PSS07, Pol09, SHI09, SY09, SZJ09, UHuI09, Wan02d, Waz00, Waz05c, YB08, ZZZ09, Zhu05]. modular [Deu01, LLY01, LTT02, NFI03]. modules [DM09]. Moduli [Gan00, AK08a, ARS02, HNR08]. modulus [GA07, Tan03b]. modus [WHS08]. Moira [Ano03-57]. molecular [Gre09, Kor07, MUHdH03, Qia08]. molecules [ITM08]. mollification [AM08a, Col04, ZCM01a]. Moment [WGZ06, Ana06, HJWL03, Kum05, Lin04b, Nab04, ZT08, ZL05]. moment-bounds [Kum05]. Moments [Ata09, BD04a, CQJ03, Kum04, Kum06, WH05]. Momentum [CJ08, LLC08, CGL01]. Moncur [Ano00-52, Ano00-53, Ano00-72, Ano00-73, Ano00-92, Ano00-93]. Monetary [Ano04n]. monic [JCZ09]. monitoring [Ano03w, NQHN03]. Monod [Ber05]. Monolayer [Qia08]. monoliths [XZ09b]. Monotone [HH04b, Jan03, LX06, NRL06, SMMM02, WW03, ABB05, CC04a, Dha07, FHT05, HN07, Kra01, LZZ04, LZ04a, LW02, LDK08, LHH08, MDG05, Noo00b, OI00, OF08, Pao03, PY09, Wan00e, Wan03b, XH07, ZGY05a, ZGY07a, Zho06]. Monotonic [BM04b, BPP04, BCRS05, Dra06, ZZZ02, DS06b, Xu02a]. monotonicity [BZN07, GL09a, Ver08b, Ver08a, Ver12]. Monte [Ano04i, Lei02, Wan02e]. Montgomery [AP04]. Moore [Ano01h, LLW08, PST08, WW00]. MOPSO [YZLL09, ZL09c]. Moringo [Ano03-33]. morphological [BH03]. Morris [GSNDB^{+}08]. Morse [Yan03b]. mortar [BCS03]. MOSEL [ZBS03]. most [LGGSP06]. most-obtuse-angle [LGGSP06]. Motility [DFY06]. Motion [BFKL08, Bao03, Dem00, FY04, KY00, Mar00, Sta07, Yav05a, Yav05b, Yav05c, Yav05d, Yav06, Yav07a, Yav07b, Ano03-88]. motions [ER07, Gal09b, Gre04, KOCO08]. motivic [BM00]. Mott [Ano00-91]. motz [LCGX06]. Move [HHL^{+}06]. movement [MG01]. Moving [FL03, AM09, BT03, BIKL06, HLP03, ITR08, KL06, MP05, Nko00, SD03, YF04, Yav05d, Yav07b]. moving-least-squares [KL06]. Moving-mesh [FL03]. MPI [Ano00-88, Ano00-89]. MPI-2 [Ano00-88]. MREIT [KPSW06]. MTOR [Cha01a]. Mueller [Ano01i]. Müller. [Ano01e]. Mullin [HHS07]. Multi [JCC^{+}09, MT00, PKM08, WSL08, dORM08, BJK^{+}08, CQT02, Dha07, Du08, GZZ09, GK00, HD09, HC09, HGS07, JS08, JCCT09, LRL09, LW09d, LZS07, LLS^{+}09c, LPvST07, MM05a, MM08a, MG08b, MKK08, PS08b, RR09b, WG07b, WYL^{+}09, XZ08b, YZLL09, ZL08d, Zha09d]. multi-asset [LPvST07, MKK08]. multi-class [WYL^{+}09]. multi-dimensional [LZS07]. multi-goal [PS08b]. multi-innovation [HD09]. multi-input [HD09]. Multi-item [PKM08, MM05a]. multi-items [MM08a]. Multi-nephron [MT00]. Multi-objective [JCC^{+}09, dORM08, JCCT09, LRL09, LW09d, MM08a, XZ08b, YZLL09]. multi-p [GK00]. Multi-point [WSL08, Du08, WG07b, ZL08d, Zha09d]. multi-product [MG08b]. multi-species [HC09]. multi-step [RR09b]. multi-structures [HGS07]. multi-symplectic [CQT02]. multi-target [LLS^{+}09c]. multi-trees [JS08]. multi-valued [Dha07]. multi-vasa [MT00]. multi-view [GZZ09]. multiagent [KC03b]. multicast [IKY06]. multicellular [BBH07, BB09b]. multichannel [CS00b]. Multiciliary [DFY06]. multicolor [LB04]. multicommodity [Lin01, Lin04a]. Multicomponent [SW06]. multicomputer [ISAAM09]. Multicriteria [SBSY06, Ehr02, SKW02, TS00a]. multidelays [JLW04]. Multidimensional [BNR05, Åhl02, CSLC07, FL05, PLD02]. multidomain [FP02]. multiexponent [LH08]. multifactorial [LYL00a]. multilane [BC08]. multilayer [WHL02]. multilayered [YZ02a, Yan04a]. Multilevel [CGGC02, GKM08, NSB08, Zha01d, GL05, HTKM09, OTB^{+}08, SWL^{+}04, Shi05]. multimedia [WLT07, Ano00-57]. multimodal [ZYY09]. Multiobjective [Zha06, BW06a, CSZH09, Din00a, Din00b, DPJ03, EG02, ESNP06, EMP^{+}06, GMMB06, GA06, Isk04, Isk08, KT03b, MM05a, Nob06, SWL^{+}04, SL06a, WFTL09, YYNK03, Zal02]. multiplayer [PRA01, PRA02]. Multiple [AA01, CG08, CH00c, Cui09, DS02, DM08, Gor00, HZ08a, HLY06, Jia00b, JX04, Liu07a, PLG07, RM08, Won08, Yas08, ZCW06, AC06, Asi08, BF03, CYX09, Cha06, CX04, CML^{+}09, CCL02, Che08b, CjBN09, EMM09, HT02, Kan09, LI00c, Ma07, NJ08, QW09, SHWF05, SK06b, SJL08, TK02, Tan03b, Tan04, WZ00, WL07b, Win03, WXS01, WX03c, YC06, YW08a, YSH03, YW03b, ZJLW06]. multiple-criteria [WL07b]. multiple-criterion [Che08b]. multiple-key [YSH03]. multiple-relaxation-time [Asi08]. multiplication [Cho03, HRP04, JKY06, LLY01, LTT02, MSN^{+}00, SMS^{+}04, Wei00]. Multiplicative [XL01, GL03b, IWL04, IW06, YL09a]. Multiplicity [Ma00, SZ05, ZLJW06, HW09, HW00, KL08b, SW07, SL07b, Zen09]. multiplier [CK06, HRMS09, LT01, Tan03b]. multipliers [TK02, Tan04]. multiply [Kru03b, Kru08, Xia07]. multiply-stiff [Xia07]. multipoint [Jan04a, Jan05, LDG05, MDG05]. multipolyhedron [Lee01a]. multiprocessor [Pet01, YY07]. Multiquadric [LT04, FFK02, FW04b, Har01, KH00, Lin04c, Sar06b, WKL06]. multiquadric-biharmonic [Har01]. multiquadrics [Fas02]. multirate [HD09]. multiresolution [He00a]. multiscale [Ano03-82]. multiscaling [CC08b]. Multiserver [GM09b, GSS04, KCD05, PS08a, ZBS03]. Multiset [BSM01]. Multisoliton [OR01]. multisolution [CL00]. multispecies [BNH07, Ram05]. multisplitting [Ano00-94, BSS98, Cha01a, XZ08a, YL09a]. multistep [For01, GSZ02, Hua08a, JMR00, MF01, TP03, XT02, ZC08]. multiunit [Lon03]. multivalent [AÖS04, AHS00, Aou08, ASS08, AY09, PS03a, PMS07, WSZ09, WJS09, XA09]. multivalently [AIS08]. Multivalued [AI06, CKN04, Dha04, AA09c, Bai01, BA08, BGO05, Dha06, HH03, Hua00, HGH02, Lan04a, Pan07c, SW08, ZCK00]. multivariable [HS07, KT02]. Multivariate [Ana07a, HHRA02, WL07a, All02, Ana00a, Ana04b, APV02, CSX01, DFSS08, LF05, SMW04]. multiwavelet [LCHL06]. Multiwavelets [PPA08, AHNV01, HW00]. MuPAD [Ano04o]. Muriel [Ano03-85]. Murphy [Ano03-87]. Murray [Ano01-28]. MUSCL [Soh05]. music [BM00]. mutation [ZL09c]. mutual [CF04, XMW09]. mutualism [Wan09e]. mutually [SHKL08]. MWF [BM09, BPM09].

N [Ano01j, Ano01p, Ano01q, Ano01s, Ano03-66, Ano03-70, GS07b]. Nachum [Ano03m]. Näher [Ano00-41]. Nahin [Ano03-27]. Nair [MCC09]. Naive [GZZ09]. nanotechnology [dZmClQ09]. narrow [SS09e]. Nash [JPS09, KL02, LF02, NSQ02, SA03]. Nash-type [LF02]. Nathan [Ano00-48, Ano00-49, Ano00-78, Ano00-79]. nation [Ano00v]. National [KS07]. natural [Ano00g, Ano00h, KOV08, MO08, Ano01r]. nature [May00, SWK07]. Navier [AL08b, FP02, GC06, GCN06, HLMR08, ITT07, KLS06, Lia04, Man05, NP07, PR07, Tou05]. navigation [AF09, KYJ03, KKS^{+}04a, ZYY08]. NDDEs [Zha04a]. Near [ZHZT06, FDWC02, JS04, Käm00, SL07a, VBG07]. Near-Hamiltonian [ZHZT06]. nearest [LdSF00]. Necessary [mLpW08, PXGL09, TL08, BRMS01, Sun06a, Sun06b, WLZ07]. Nédélec [GGCD05]. needed [Kru03a]. needs [Ano03-50]. negative [AÖS04, BBP09, EM09c, KÖ08a, Kar09, Kha05, LWZ08, LP08b, LZ02, LLD02, Öca07, SW04, TYP00, WL09a]. negatively [FZ07b]. Neggers [BKN09]. neighborhood [CLJ09]. Neighborhoods [AÖS04, AIS08, MA09b]. neighbour [LL07]. neighbour-scattering [LL07]. nephron [MT00]. Netherlands [Ano04k]. nets [IUCR06]. network [ASM09, Ana00a, Ana04c, Ana04e, Ano00w, CH05a, CZ05b, CCF06, ETH05, ET09, GOMA07, GHHD08, GMS09, HY00, Iac07, IYN03, JCC^{+}09, KYNI06, KC03b, KL01, LMA08, Lin01, Lin06b, Lin08b, LI00c, LH06, LYC07, LLS^{+}09c, LHLW07, LHWW08, MPW07, PL01b, PSM02, SSL07, SWL^{+}04, SMOK05, WYL^{+}09, WLD09, YH03, ZXX05, ZLLL08]. networked [Ano04l, ZTH09]. Networking [Ano00-52, Ano00-53]. Networks [GY06, AMYS^{+}09, AM08b, Ano03-89, BBM08, Can03, Che08a, CZ00, FKM05, Gon09, GHW04, HH04a, HaLC08, HCY00, HHW08, HLZ09, ISAAM09, IMPF08, JCCT09, KM03b, KB08, LL03a, tLW08, Li08a, LLL08a, LYJ09, LW09f, LC06c, MXZL09, MG08c, MPTV06, NC09, Ou08, PS07a, RC07, RVKDF09, SA03, SM08, SMO03, TGIM02, TaL09, WSO08, WNS09, WZZ09, WLT07, WL08c, XCL^{+}08, XZZ01, XLL03, XCH07, XCS^{+}08, XLT09a, XLT09b, YH09, YFT04, YMI06, Zha03a, Zha04b, ZLH07, ZYZ08, ZL09a, ZSZ06, ZXTL09, ZZ05, ZW03b, ZH04, Ano03-48, Ano04s]. networks-applications [BBM08]. Neuburg [Ano00-65]. Neumann [COE03, KKM00, LWY04, RR08, Sim01, WA08].

Neural [Ano03-86, GY06, KL01, YFT04, Ana00a, Ana04c, Ana04e, Ano00-54, Ano01e, Ano03-48, Can03, Che08a, DW09, ES08b, GOMA07, Gon09, GHW04, HH04a, HHW08, HLZ09, Iac07, KB08, LL03a, tLW08, Li08a, LYJ09, LYC07, LNL08, LLS^{+}09c, LW09f, LC06c, LHLW07, LHWW08, MXZL09, MPTV06, NLL07, NLL08, NC09, Ou08, PL01b, RC07, SWL^{+}04, SA03, WZZ09, WL08c, WLD09, XZZ01, XLL03, XLT09a, XLT09b, YH09, Zha03a, Zha04b, ZLH07, ZYZ08, ZSZ06, ZXTL09, ZZ05, ZW03b, ZLLL08, ZH04, Ano00-54, Ano03-54]. Neuro [CCL02]. Neuro-fuzzy [CCL02]. neurobiological [Ano03-54]. neuroeconomics [Ano03x]. neuron [LH06, NC09, WL08c, Ano03-79]. neurons [LYL00b, YH03]. neurophilosophy [Ano03l]. neuroscience [Ano04-28]. Neurotechnology [Ano03-55]. Neutral [Wan06b, AB08a, AML02, BSD00, BA01, BPA03, BSK04, BM01, BM02, BK04, Can08, CH04, CHR09, Don09, GL09b, GS09, HPT05, HM05b, Huo04, KÖ08a, Kar09, KSZ02, Li08b, Lin05, Lin06a, LY06a, Liu08b, LXK09, LS01b, Luo02, Mus06, Öca07, OZY05, Par04b, Pen02, PHW03, QSL08, SW04, TYP00, Tan02a, TL08, TMG03, TK07, Wan00d, Wan04b, WG08b, YYZ03, Zha01b, ZCZ01, Zha01c, ZL03, ZFYS05, ZS07, ZZ03b, ZY08b]. neutrinos [Ano03-70]. neutron [AK06, BGM^{+}02]. Nevanlinna [RGTT03]. never [Ung03]. Newhouse [Ano03-63]. Newmark [Rub07]. Newton [Arg00a, Arg00b, Chu05, CN08a, Fas02, For01, FT05, FM05, Gal08b, Hua04, KLS06, KI04, LZJ04, LN08, NR01, Pan07a, Ryo00b, SHWF09, Tao06, WW09b, WS08b, Zhu05]. Newton-like [Ryo00b]. Newton-like-iterative [Arg00a]. Newton-type [Tao06]. Newton/structured [Pan07a]. Newtonian [APS07, AB09, BW09, ES05, FS07, MM00, MG01, SZL02, WB09, XL09]. NGP [Zha04a]. NHPP [IY06]. Nicholas [Ano03-69]. Nicolson [ZW09a]. Nigel [Ano00-38]. nilpotent [BENT06]. Nils [Ano03-37]. nine [HL04]. ninth [DLM08]. Nirenberg [ADP08, HZ08a]. Nirenberg-type [ADP08]. Nizhnik [SDB09]. NJ [Ano03-31, Ano03-49, Ano03-59, Ano03-68, Ano03-74, Ano04m, Ano04r, Ano04s, Ano04u, Ano04t]. NL [KC03a]. NLG [Ano00d, Ano00u, Ano00x, Ano00y, Ano00w, Ano00-28, Ano00-50, Ano00-56, Ano00-71, Ano00-85, Ano01d, Ano01h]. no [AS05, SH01, XXLD04]. no-wait [XXLD04]. node [Li05]. nodes [Lin06b, ZL09a]. noise [AL08a, CS05a, CS02a, SNA09]. noises [Jum01]. noisy [AC06, BR03, LLXC03, SMW04]. Non [AMC08, Aro08, BBCE09, BW09, DH03, ZTH09, AK07, Ano00-30, AB09, CWW08, DH08c, FM08a, GKM08, Gla07a, HLY08, HL08a, Jia00a, KK01, LKH03, tLW08, LTS03, MG01, ME08, MS00, NTHP09, PP09, PW06, RR09a, RSD09, ST07, SZL02, Tan03b, TCH07, VM09, VJ04, WB09, WZZ09, XL09]. non-atomic [HLY08]. non-autonomous [DH08c, tLW08, NTHP09, PP09, WZZ09]. non-bandlimited [CWW08]. non-conforming [GKM08]. non-Darcy [MS00]. non-deformable [FM08a]. Non-fragile [ZTH09]. non-Fuchsian [LTS03]. non-Hermitian [Jia00a]. non-homogeneous [RSD09]. non-ideal [Gla07a]. non-isothermal [VM09]. non-linear [ME08]. non-Lipschitzian [PW06, KK01]. non-Mersenne [Tan03b]. Non-Newtonian [BW09, AB09, MG01, SZL02, WB09, XL09]. non-parameter [HL08a]. non-partitioned [AK07]. Non-planar [Aro08]. non-preemptive [LKH03]. non-self [RR09a, TCH07]. non-self-adjoint [VJ04]. Non-smooth [BBCE09]. Non-standard [AMC08]. non-stationary [Ano00-30, ST07]. nonautonomous [AP03, CP01b, LY08b, LG04b, LC07b, Mur07b, Nes01, Pit01, Sie03, SC00b, TL00, ZWW04]. nonblocking [KM03b, SMOK05]. Noncentral [SNG06]. nonclearing [Ano04-27]. noncoercive [Bou00, Bou03]. noncompact [BN00, Din00c, IL04, Waz04, Waz05b, Wu00]. nonconforming [Che04f, KK02, Ses03]. Nonconvex [BM06, ARM08, BGG04, HSM00, HS02, PSS07]. Nonconvolution [ABB05]. noncooperative [AMR05, PRA01, PRA02]. Nondecreasing [ESR06]. nondifferentiable [Jum06, ZXL07]. Nonempty [TEB06]. nonequilibrium [BRC00]. nonergodicity [KL08a]. Nonexistence [WZ00]. Nonexpansive [yL06, TS06, CWY08, CGY09a, CZG04, CKQ08, udK07, HS09, KH08, PY09, Son08, Tia05, TCH07, Wan07a, YYZ07]. Nonhomogeneous [ZLY06, BD03, LP06, Ma04, WC04, ZBS03]. nonhyperbolicity [BV01a]. Nonhysteretic [CKK06]. nonincreasing [EM03]. nonintegrable [TGH02]. noninvertible [AK01]. Nonlinear [AO01c, AI06, BGAM09, BR06, CB02, HaLC08, HW05b, Kar08, LCV06, LS08, NST01, Nie06, PXXW06, RL08, Sah01, Sat06, SPE06, Sun05b, Sun05c, Ver03, Waz05b, Win07, ZCS04, AETEZ07a, AH00, AA08, AL04, AN09, AC06, AL08a, Amm07, AR07, AE08, AK08c, ABB05, Asl11, AP01, AT02, BMHV09, Bai00, BPA03, BSK04, BSR02, BM01, BM02, BV05, BM05a, BP03, BHEKE07, BB09a, Bel09, BDGR01, BHNO04b, BB04, BGR09, BMH08, BA09b, BS03b, BR05b, BPSW00, BPW00, BFO02, BB08b, COEP00, Cab01, CÁHCJL^{+}09, CYG06, ÇT05, CP01a, CP02, Cat05, CVW00, CBJ01, CHY01, CQT02, CZ03b, CX04, Che04d, CyL08, CCK00, Chu07, CjBN09, CJK09, CNPP08, CM08, DH08a, DI08, DRR09, DJ08, DN03, DYZ08, DTXW05, Dor01, DZ05, DFY02, DXZ05, Dua09]. nonlinear [EA03a, EGZ04, ETAJ09, EAE09, ERRY01, EIR03, Fan02a, Fas02, FY03b, Fro08, FS07, GNM07, GTJ07, GGA09, GGE09, GT00, GY08, GSN08, GMS09, Gra03, Ha01, Had03, Han04, HMCR00, He03, HW09, HL09a, HM05b, HZZ02, HBCK00, HF04, HL03b, Ili04, IJB00, JKN09, Jan04a, Jan04b, Jan04c, JKP02b, JZQZ07, JT05, JZT07, JDH03, JDY08, Jin07, KF05b, KB04, KK05, KPP00, KL08b, KLW09, KD08, KB08, LPRH05, LY09a, yL08b, yL09c, LKS08a, LA00, LFZ00, Li01b, LML02, LA03, LZL04, LS04b, LLX05, LLJ09b, Li09b, LX00, LHW04, Lin06a, LSZ08, LLD02, Liu02, LK03, LWG04, LC04, LQ04, Liu08b, LN08, LXK09, Liu09b, Lu07, LS01b, LHWW08, Ma01, MCH01, Man01, MH07, MHS01, MZ07, Mic03, MLW09, MF01, MEA09, MN01, Mur05a, Mur05b, MO01, NTTS09, NJ08, yNyLjZaZ08, NRL00, NW09, ÖY07a, Pao01b, Pao03, PKT07]. nonlinear [PL01c, PG00, Pen02, PHW03, PL05b, PZ07, PXW09, PP06, PS01, PS03b, QSL08, RR09b, RF09, Ren03, RR03a, SNG09, SG07, SERG09, SKU01, SA02, SSX09, SAS06, Sha04, She00, Sho09a, Sho09b, Son09, SK06b, SZ05, SL07b, Sun06a, Sun06b, SAB07, SK09, Tan07, TL08, TA01, TL06, Tia01, TÇ04, TA04, Trz01, ÜÇT08, Ver01a, Ver01b, Ver09, WZG00, Wan00d, Wan03a, WWZ03, WFC04a, WLW07, WKZ08, Wan08a, WYL08a, Wan09d, WWS^{+}06, Waz01, Waz07c, Waz07b, Waz07a, WA08, WZ09, WWYL09, WF01, WWX04, WW08d, XLZ02, Xu01, YYZ03, Yan03e, Yan03c, lYsT09, Yao05, Yas08, YZZC09, YLD09, YMMC03, YL08c, ZGY05b, ZGY07a, ZP01, ZCZ01, ZZ01b, ZL02, ZSKZ03, ZY05, ZS05, ZYEZ07, Zha09a, ZTW09, ZXlD09, ZTH09, Zha09b, ZWY08, ZC09b, Zha09g, ZC09a, ZW03b, ZY08b, ZC01, ZH04]. nonlinearities [AS03b, DG05, EM03, HRS08, KP06, Kon09, LLJ09b, PHAF09, Xu08b, Yan06a]. Nonlinearity [HX09, LS09, Liu09b, SSX09]. Nonlocal [AS08, ZW06, ADDL04, BM04a, BN00, CL07b, Che04e, DDX04, GNR05, GM09c, HMCR00, JKP02a, LX03, LX04, Liu04a, Öze08, PGT08, XCD07, ZHKW05]. nonlytic [WWL06]. nonmatching [CKL06]. Nonmonotone [Mig06, CSdS08, zGtM08, SY09, ZZ03a]. nonmonotonic [HZ08b]. Nonnegative [AS03b, MSS08, CZW08, GLZ00, RSR01, RR03b, Xie00, ZHL05, ZHKW05]. Nonnegativity [HZ03]. nonnormal [CT08]. nonobtuse [Kr05]. Nonorthogonal [AC00]. Nonoscillation [DR01b, KSZ02, BBP09, Kar09, TYP00, ZZ00a]. Nonoscillatory [AGO03b, Yan03e, BGO05, HLY05, Li01b, TMG03, Wei05, ZFYS05, ZZ03b]. nonparabolic [HW05c]. nonpolynomial [Bil05]. nonpreemptive [GC02]. Nonreflecting [DR01a, JG04, VNG04]. Nonresonance [BGO04]. Nonself [TS06, IK03, KH08, PW06, Son08, Wan07a]. Nonself-Mappings [TS06, PW06, Son08]. nonsimultaneous [TH02]. nonsingular [TT09]. Nonsmooth [BGG04, BRMS01, GA06, KT03b, Nob06, PWX08, ZX09]. nonstandard [Mic03, Roe09]. nonstationary [BPvZ03]. nonstiff [CM02]. nonsuccessive [TK02]. nonsymmetric [AK08c, Lin08c, MAK02, MAK03]. Nontrivial [TLLM08, ZL08d, EMM09, NT08, ZHKW05]. nonuniform [LHC05]. nonuniformly [ZW04]. Nonzero [FW08]. Noor [HS09]. Nörlund [LS06a]. Norm [SKW02, CT08, Deu01, HL07, KN04, Kar01, Kov05, Kua08, LL05, LHP07]. Norm-based [SKW02]. Normal [KM03a, Sie03, FS04, KGWL08, Xu02b]. normalized [ZT08]. normally [Gar01]. normals [WB01]. norms [DLF09, TC09, XD09]. Norske [Bu00]. North [Ano00g, Ano00h]. North-Holland [Ano00g, Ano00h]. Norton [Ano01p]. nose [Pos00]. Note [Jun10, LLB05, LW06, Xu06, Ada04, AR10, CC04a, Dua08, FQZ02, GT08, HSWL08, Jun01, JY11, Kan07, LSL09, LCT08, Pen03, QSL08, RSE01, SR07, Yan08, YH11, ZY02]. Notes [BG11, fDxZ11, Lin00, Liu09a, Pei02, hYL09, Blo00]. notice [ID08]. Novel [GOMA07, YH09, AB08c, Ano04-30, BHS^{+}08, DW09, JKY06, KÖ09b, LDK^{+}06, RTS07, WYL^{+}09, WW09c, XCS^{+}08, YZLL09, ZWC^{+}09]. Novikov [SDB09]. NP [LW09b]. NP-hard [LW09b]. NPD [Dra06]. NT [Ano00-72, Ano00-73, Ano00-92, Ano00-93]. Nth [GM09a]. NTP [ID10, JW05]. nuclear [IUCR06]. nucleolus [ND01]. Null [Phu02, Ano06k]. Number [Ind02, LLB05, ZHZT06, AS05, AT05, ANZ09, EG02, ES08a, GG06, GM00, HT04, IYN03, KT03a, LW04, LL07, LLJ09a, LLW08, Lon08, MPTV06, RW03, Tan04, TY03, Tsi03, Wan07b, XHZ05, YTCF08, Ano03-56]. Numbers [AEÇ06, AH09, AS07, CL01, CL02b, CC08a, CT08, Che04g, Cho03, CT02b, DZQ06, HWC04, HX06, HX07, Isk02, KR07b, LH05b, LWL08, MIPK04, OS08b, TB09, TAT07, WL08b, WL09a, YLZG01, ZZ01c, dG09]. numeric [SYJJ09]. Numerical [AVLPN04, AACLM08, AAK08, ACLL06, AM08b, Ara01, BMHV09, BKM08, BH03, BCGP03, BR05b, BR03, CL02a, CO00, CK02a, CKL06, CY00, CSdS08, CKK06, CC04b, CJK09, CNPP08, DPU07, DI08, DSHS05, DR04, DW02b, ES05, FS04, HRP01, HL02a, HKL03, HMCR00, Ji02, JZC06, KK06b, KHA07, KT08, KOV08, KR01, KW03, Kur08, LW09c, LSZ08, LET06, NS07, NLR00, Pao01b, PJJP04, RNM00, Roo08, Ryo00b, Ryo00a, RN02, RSK02, SD07, SPE06, SDZ06, SWS05, SK06b, TDS08, UHuI09, VC03, WL04, WDLP08, YV05, YB07b, Zha03b, ACD01, AP08, ADDL04, AM09, AS01, ACJ07, AB05, AE07, AE08, Ano03-42, Ano03-46, Ano04f, AMC08, Bab09, BUP09a, BBL05, CM09, CDJR03, CJO05, Cha08, CJ04, CDJ04, CSPJ05, CJV07, CM08, DKS07, De 02, DS02, DS07, Dim08, Dog08, FGH08, FHZ08]. numerical [GK02, GGE09, GHJ06, HW05a, HC01, HC02, Jan03, Jay04, JC00, JSF00, JPV01, KP03, KS00b, KS03b, LF03, LS04a, LH00b, Li01c, Li05, LWY02, LWY04, MM08a, MM05b, MAK02, MT09, Mur06, Mur07a, NVAR03, Nie09, OAVB08, PPRG02, Pao03, PM05a, PDRR08, PB02, RL08, RRM08, Ryo04, Ryo07, SDK01, SR03, Sim01, SPW08, SD03, Tia08, VR02, VA04, Win03, WG08b, WLD09, XCS^{+}08, ZCM01a, ZZJ^{+}04, Zha09c, dG09, LH00b, Ano04p]. Numerov [AW01, Tsi03, Van07, Wan03a]. Numerov-type [Van07]. Nutrient [JB05]. Nutrient-plankton [JB05]. Nutshell [Ano00-78, Ano00-79, Ano00-36, Ano00-86, Ano00-91]. NY [Ano03p, Ano03q, Ano04g, Ano04j]. Nyström [BBRS08, CM02, PF01, XT04a].

O [Ano03-29, Ano03-58]. O.D.E. [BB04]. OBDD [JWSS04]. obesity [JSGP08]. object [Ano01w, Ano03-75, Ano04u, BIKL06, HJWL03, PGH^{+}08, ROY01, Ano00-55]. object-oriented [Ano01w, Ano03-75, Ano04u, ROY01, Ano00-55]. objective [JCCT09, JCC^{+}09, LRL09, LW09d, MM08a, SJL08, XZ08b, YZLL09, dORM08]. objectives [BGMB02]. objects [Ano04u, LG00, Ros03]. Obrechkoff [NF03, Net07]. Observability [AHK00, SBS06]. observable [Nak06, TG03]. observation [AL08a]. Observations [FDWC02, AC06, CÁHCJL^{+}09, CS05a, FWL04, LLXC03, OLAS01]. observer [GLTS08, Sun05b, Sun05c, TLS^{+}08]. observer-based [GLTS08, TLS^{+}08]. observers [ZCS04]. obstacle [CG05, LZZ04, Ryo00b, Ryo03, Ryo07, XC04b, Yan04a, YF04, Zha01d, ZZZ02]. obstacles [KKM00]. obtain [BENT06, MCC09]. obtained [MP05, YW09]. obtaining [Waz01, YC07a]. obtuse [LGGSP06]. occasion [dKLR09]. occlusion [PGH^{+}08]. occurring [ZZ06]. Ocean [CZLN06, Ano03-57]. odd [LLT08]. odd-order [LLT08]. ODE [ANV00, LG08, MT09, Pet01, SSD07, Waz05a, Zla01]. ODEs [Ano00-66, BBL05, CM08, Ikh01, Ikh02, Ikh04, Jia00b, Sal08]. off [LW09d, YX09b]. offline [WM09]. offs [Dra06]. Ohlsson [Ano00-28]. oil [CL02a, FM08a, YLZG01]. Olaf [Ano03e]. Oldroyd [AVO08, FMFV08, HHK06, KHA07, VNFF08]. Oldroyd-B [FMFV08, HHK06, VNFF08]. oligomer [WK03]. oligopolies [CS05b, SZ06]. oligopoly [PP09]. Oliver [Ano04-31]. on-line [He00b, ZY02]. One [Ano04c, CS05a, DS06a, Dor01, GS06, Liu04c, Ano00-94, BSS98, BS00a, BM09, Bs03a, CCC08, CAZAA00, CZLC08, Che00b, DWL07, FL03, HJW05, HN07, Ikh01, Ikh02, Ikh04, Jia01c, JG04, JX04, JDH03, JP01, MQ05, Min08, NWL08, OAVB08, Roo08, SD08, She04b, Slo07, SZ09a, SKG01, VH06, WG07b, WK03, WX00b, WS08b, XT01, Xu03b, Xu08a, YLD08, ZCL08, ZL08b, ZX09]. One-Dimensional [CS05a, DS06a, Liu04c, Ano00-94, BSS98, BM09, CCC08, Che00b, FL03, HJW05, Jia01c, JX04, JDH03, JP01, MQ05, Roo08, SD08, SKG01, VH06]. one-leg [XT01]. one-parameter [ZX09]. one-period [YLD08]. One-phase [GS06, Slo07, SZ09a]. one-step [CAZAA00, HN07, Ikh01, Ikh02, Ikh04, Min08, WX00b]. ones [BD04a, SS09d]. online [LGH09]. only [CT02a, WSZ03]. Onsager [Yon09]. Onsager-like [Yon09]. onthe [CJL04]. onto [SMS^{+}04]. Open [GC06, GCN06, CCZK09, Gla05b, GM00, IMLF09, Ano04q]. Open-domain [Ano04q]. operated [Mos08]. operating [Ano00i, Ano00-56]. Operation [SKU01, YWC08, Ano05f, Cho03, Dor01, EW00a, Eng04, LJ08a, YL04b]. operations [AFL^{+}09, AVA01, KYAA08, LW04, LNL08, ZYL06, Ano04r]. Operator [CN08b, Hip06, WG06, AIS08, BSR02, BS01a, CLC01, FGH08, FBA07, Gup02b, HHS06, HHY09, IP07, JV01, yL09c, LZZ04, LXZ08, MDG05, Oro02, OI00, PMS07, PW00, SALK02, UA06, WC04, WA08, qWjH07, XH07, XD09, XA09, YZ02a, ZYG08, ZL07b, ZCK00]. operator-integral [BS01a]. Operators [YI06, AZCK05, Agr08, Ana00a, AG01c, Ana04a, Ana04b, Ana04c, Ana04e, Ana05, AD08, ASS08, ABB05, AO08b, AP01, BM05a, Bou03, CCP02, CC04a, CHS02, Cap00, CC05, CC09b, CC02, DD04, Dha03, Dha04, Dha06, DBX04, FHT05, Gal08a, GYS05, GJQ07, GN09, Gup03, Gup04, HMV04, HHR07, Hil01, HH03, HCWC07, Hua00, KF05b, Kua08, yL09c, LLX05, LTSW02, LTS03, LK03, LKC04, LHH08, MN01, PXW09, PP01b, PD04, SZ09b, XWD04, Xu04, Zen00, ZG02, ZGY05a, ZC06, ZL08a, ZZ09a, ZC09b, ZQC08]. Opial [Ana04d]. opinion [Che04g]. opportunities [DCR07]. optical [FTC03, JS08, PGH^{+}08, Raz00]. Optimal [AH03, AHN06, CD05a, Che01a, CHW07, HM00, IYN03, Käm00, KM03b, KB08, LET06, MNIS06, NFI03, NUN06, ODO00, OK06, PXXW06, PR07, SI03, SO03, Saw03, TLS07, WW04a, WM09, XWL06, AC06, Ana06, ADN02, BBGG08, BR05a, Bog09, CY00, CLS07, DKO03, DFSS08, Dra06, DW05, FFXL05, Gal08c, HRP01, He00b, HL08b, JS03, Kar05, KK03, LCL^{+}09, LCT08, LT09b, MM08a, MCC09, NQHN03, ODO03, PXGL09, PXW09, TLP^{+}09, VG08, WLS01, WS00, WZ07a, WTZY05, YZLL09, YZWY08, YD08, ZM04b]. Optimal-order [HM00]. Optimality [ARM08, CWZ06, ES08a, KT03b, Wu08a, Amm07, CLT09, CCF09, Gal06a, HYZ06, PWX08, PXGL09, Zal03]. optimally [HY06]. Optimisation [PPP^{+}06, FT05]. Optimization [BFL06, Gal02b, HM03b, LDHW05, LY09c, MIPT08, Zha06, ARM08, ASM09, Ano00o, Ano00y, Ano01g, Ano03h, BGMB02, BGG04, BM06, BG04b, BGS04, BRMS01, CS05a, CS02a, CPP07, Che04c, CHZ09, CCL02, CCF09, CZ02, Csa01, DTXW05, EG02, FHH08, GK02, Gal02a, GGG03, zGtM08, GA06, GL00, HTKM09, HRMS09, IMPF08, JZQZ07, JCC^{+}09, JT05, JZT07, bJQxC09, KLZZ09, KT03b, KH09, Lei02, LJ08a, LWLJ09, LCL^{+}09, LLJ09a, LI00c, LG09b, MWZ09, MLW09, NLLL09, Nob06, PWX08, PG06, PG08, SYJJ09, Sny00, SH02, SF01, SF07, SL06a, Wan02e, WZG04, WL08a, WFTL09, Wu04, WS08b, WYSS09, XWW08, Yan03f, YZLL09, YL07, ZX09, ZYY09, ZZZ09, ZCM01b, ZZ02b, ZZ06, dSOGRMRL04, dSRGRMRL08]. optimize [PdlF09]. optimized [LWLJ09]. Optimizing [PdlF07, Ano00-57]. optimum [LLL08b, YLZG01]. Option [TAT07, BCJ08, CKKK06, CNPP08, DLL08]. optional [Wan04a, WZ07b, WX09a]. options [CM03, Hon02, LPRH05, LPvST07]. oracles [YXZL09]. orbits [AlS08, DCS09, Din04a, DS05b, Kot01, VAR03, WM04, Zho03, ZW03b]. Order [CZLN06, LCL06, LBW09, Liu06b, SW02b, XT04a, ZLJW06, AG01a, AGO03a, ATT03b, AORS03, AL04, AC02, AC04, AC05, AS01, AJ09a, AJ09b, ADP08, Ana04b, Ana04c, AA01, And03, AK08b, AW09, AY06, AC03, AT02, Ayk04, BG04a, BX07, BLD07, Bai08, BA01, BPA03, BPM05, BJBL03, BN00, BHN01, BHNO04b, BBRS08, BGNA07, BNR05, BG07a, COEP00, CH01, COE03, CC09a, CYG05, ÇT05, CHS02, Can08, CH04, CH00a, CM04, CL02b, CQJ03, CS04b, CL08a, CLT09, Chu07, CjBN09, CH00c, CM02, Coo05, CZ08, CZ09c, Cui09, DJ08, DK05, DCS09, DHM05, Dix05, Dog08, Don09, DK04, DP07, DR01b, DM07, EM09a, EM09c, ED09, FB05, Fen08, FR04, FT03, FFXL05, GM09a, GW03a, GF06, GPPRM01, GH03, GM09c, GW01b, GS09, Guo01, Guo02, GZ06, GJS01, HJWL03, HAKA07]. order [HW09, HT02, HY03, HM05b, HM00, HLX07, HLY05, IS04, Jan09, JEP08, Jia00b, JLT04, JT07b, Jum01, KA09b, KZ02, KÖ08b, KK04, KKZ02, KLW09, Kra01, LA00, LFZ00, LA03, LS04b, Li04b, LZ04a, LZ05, LTY05, LS08, LLT08, Lin05, LLNS05, LDG05, Liu04a, Liu04b, Liu07a, LX07, Liu08b, LDK08, LWL08, LXK09, LYG05, LST02, Lu07, Luo02, LD02, Luo03, LSN05, LJ08b, Ma00, MQ05, Man00, Man01, Mas07, McG03, MA09b, NF03, Net07, NJ08, NWL08, NRL06, NMD07, NMD08, OLS06, Ou08, Pal01, PLG07, PWX08, PGT08, PG00, Pen02, PHW03, PXW09, Pit03, QSL08, RSD09, Rap01, RC07, RC09, SALK02, Sha02, SR03, SR04, Shi03, SC03, SW02a, SCF^{+}09, SB06, SOW04, Sun05a, SS06, SL07b, SZ09b, STH08, Swe07, Tan02b, TL08, TA01, TRG01a, TRG01b, TÇ04, TA04, TBG05, TK09, Top04, Tsi02]. order [TDS08, VR02, VNG04, Van07, Wan00e, Wan00d, WWZ03, Wan04b, WLS06, WG07a, WLW07, WW08b, WJA08, WWW08, WSL08, Wan08a, Waz00, WA01, WWC08, Won08, WWYL09, WX00b, WF01, WWX04, WD05, WZ07c, WW08d, Wu09, XLL03, XZ08a, Xu08b, XLT09a, XLT09b, Yan01, Yan04c, YYZ03, Yan03c, Yan07, YW07, lYsT09, Yao05, Yao07, Yen08, Yil08, YWYL07, ZZ00a, ZP01, ZCZ01, Zha03b, ZFYS05, ZS05, ZLY06, ZL08c, ZL08d, ZJ08, ZW09a, Zha09c, Zha09d, ZWY08, Zha08d, ZC09b, ZG09b, Zho01, ZZ03b, ZY08b, Zho08]. ordered [AP01, Dha06, QLD08b, ZYG08, ZC09b]. ordering [SL06c]. orders [BS00a]. Ordinal [TH02]. Ordinary [Jan04b, Liu06b, AP08, Ano03-42, Ano03-67, Ash05, ACH01, CCP02, CL08a, DKS07, Gar01, GL03a, HdLS03, LDG05, Liu04a, LSN05, MCH01, PGT08, ST00, SR03, SR04, Tel08, VR02, ZTW09]. O’Reilly [Ano00-75, Ano00v, Ano00-27, Ano00-34, Ano00-36, Ano00-37, Ano00-39, Ano00-40, Ano00-44, Ano00-45, Ano00-52, Ano00-53, Ano00-57, Ano00-60, Ano00-78, Ano00-61, Ano00-79, Ano00-63, Ano00-62, Ano00-67, Ano00-65, Ano00-68, Ano00-72, Ano00-73, Ano00-87, Ano00-82, Ano00-86, Ano00-92, Ano00-91, Ano00-93, Ano01f, Ano01j, Ano01k, Ano01s, Ano01-27]. organizing [PL01b]. orientation [ID05, ID08]. oriented [Ano01w, Ano03-75, Ano04u, OP01, PDP05, ROY01, YX09b, Ano00-55]. Orienting [CLS08]. original [CT02b, WL08b]. Orlicz [Sav08a]. Ornstein [APvZ01, AF02, BPvZ03, BB01]. Orthogonal [CZLN06, HW00, BC02, BC03, BCGA05, Boy07, CO00, He00a, MCS08, SHKL08, VBG07]. orthogonality [APvZ01, GLR05, Jay04, VJ04]. orthogonalization [GLR05, JH09]. orthotropic [CB02, DD00]. Osborn [Ano00z]. Oscillating [DL06, Van07, ZY08b]. Oscillation [AG01a, AML02, AT02, ÇT05, Can08, CZ03a, Don09, FT03, GF06, GW01b, GS09, HL03b, JT07a, JT07b, KÖ08a, Kop01, Lin05, Lin06a, LLD02, LX07, LS01b, Luo02, LD02, Man00, MS02, Öca07, OZY05, PG00, Pen02, SA02, SW04, Shi03, TYP00, TZ01, TRG01a, TRG01b, TK07, TA04, WZG00, Wan00d, WWZ03, Wan04b, WY06, WWX04, WD05, XTX01, Xu08b, Yan01, Yan04c, ZZ00a, ZZ01a, ZCZ01, ZT04, ZS05, ZWY08, AL04, CZ03b, CKZ01, DJ08, DK05, DH03, DFY02, EM09c, GZ06, JEP08, JLT04, Kar09, Kha09, LA00, Liu08b, LS01a, LS03, Luo03, Man01, ÖZ07, QSL08, SL00a, SOW04, Tan02b, TL08, TX04, TÇ04, TBG05, WXM09, YYZ03, ZP01, ZA03a, ZC01]. Oscillations [LC04, ZZ02a, ZZ00b, Zho01, DTZF03, HM05a, LGY09, MH07, PHW03, SD03, TA01, Zha09b]. oscillator [Bel09, BGAM09, CW09a, ETAJ09, Fan09, ÖY07b, Zha09a]. oscillators [DZ05, He03, Liu09b, ME08, MEA09, Sho09a, Sho09b, Zha09g]. Oscillatory [BGO05, LZ02, ZY08b, CXZ09, Fly06, FLW06, GZ06, RSE01, XT04b, ZL04]. Oseen [He02, PR07]. Osher [Ano03-51]. Ostrovsky [KÖ09b]. Ostrowski [AP04, Ana07a, Ana07b, BBCD02, BCDS09, CCD07, Che01c, DS06b, KDC08, Kum05, Liu07b, LN09, Lü08, MPU00, Nie09, THD08, Uje03, Uje04b, Yan06b]. Ostrowski-Grüss [Che01c, Liu07b, LN09, MPU00, Nie09, Uje03, Yan06b]. other [Ano03-27, Ano03-58, PKSP08]. outerthickness [PM05b]. Output [YMMC03, JWSS04, LBZ^{+}09, LLL08b, MW00, PKP09, PP06, SN03, Wu03]. Oval [Ano03-58]. over-relaxed [Ver08a, Ver12]. overbooking [WK08]. overhead [KY00]. overrelaxation [BNP05]. Overspecification [DS06a]. overspecified [TD07]. Oxford [Ano00s, Ano00t, Ano00-31, Ano00-51, Ano01m, Ano01l, Ano03-76, Ano03-43]. Özcan [Ano01n].

P [Ano00o, Ano00-30, Ano00-59, Ano01x, Ano01-27, Ano03-81, Ano03-33, Ano03-63, Ari09c, EM09b, Net07, GK00]. P-stable [Net07]. Pachpatte [Li09a]. Pacific [Ano03-80, CZLN06]. package [ACD01]. packed [VM09]. packet [AFI06]. packets [LCHL06]. packing [RM08]. packings [OAVB08]. Pade [NMD09, AETEZ07a, Mak09]. PageRank [Sid08]. Pages [Ano04-27, Ano04v, Ano00e, Ano00j, Ano00f, Ano00k, Ano00d, Ano00i, Ano00-74, Ano00l, Ano00-75, Ano00q, Ano00s, Ano00m, Ano00r, Ano00t, Ano00u, Ano00o, Ano00p, Ano00n, Ano00z, Ano00x, Ano00y, Ano00w, Ano00v, Ano00g, Ano00-29, Ano00h, Ano00-28, Ano00-27, Ano00-76, Ano00-30, Ano00-77, Ano00-31, Ano00-32, Ano00-33, Ano00-35, Ano00-34, Ano00-36, Ano00-37, Ano00-38, Ano00-39, Ano00-40, Ano00-43, Ano00-41, Ano00-44, Ano00-48, Ano00-45, Ano00-49, Ano00-47, Ano00-51, Ano00-46, Ano00-50, Ano00-54, Ano00-57, Ano00-55, Ano00-56, Ano00-60, Ano00-61, Ano00-58, Ano00-63, Ano00-62, Ano00-59, Ano00-64, Ano00-67, Ano00-66, Ano00-68, Ano00-80, Ano00-71, Ano00-69, Ano00-81, Ano00-70, Ano00-85, Ano00-83, Ano00-84, Ano00-87, Ano00-82, Ano00-86, Ano00-88, Ano00-89, Ano00-91, Ano00-90, Ano01e, Ano01d]. pages [Ano01f, Ano01i, Ano01g, Ano01h, Ano01j, Ano01k, Ano01y, Ano01m, Ano01n, Ano01z, Ano01o, Ano01p, Ano01l, Ano01q, Ano01r, Ano01s, Ano01u, Ano01t, Ano01w, Ano01x, Ano01v, Ano01-27, Ano01-28, Ano03i, Ano03-76, Ano03g, Ano03f, Ano03-36, Ano03-77, Ano03j, Ano03k, Ano03h, Ano03l, Ano03t, Ano03m, Ano03n, Ano03p, Ano03u, Ano03s, Ano03r, Ano03-78, Ano03o, Ano03w, Ano03q, Ano03v, Ano03y, Ano03-27, Ano03-79, Ano03x, Ano03z, Ano03-29, Ano03-30, Ano03-28, Ano03-80, Ano03-31, Ano03-81, Ano03-39, Ano03-35, Ano03-33, Ano03-38, Ano03-34, Ano03-37, Ano03-41, Ano03-42, Ano03-40, Ano03-85, Ano03-44, Ano03-82, Ano03-43, Ano03-83, Ano03-45, Ano03-46, Ano03-47, Ano03-50, Ano03-49, Ano03-48, Ano03-51, Ano03-52, Ano03-53, Ano03-84, Ano03-86, Ano03-56, Ano03-55, Ano03-54, Ano03-57]. pages [Ano03-58, Ano03-60, Ano03-61, Ano03-63, Ano03-62, Ano03-59, Ano03-64, Ano03-87, Ano03-65, Ano03-73, Ano03-72, Ano03-66, Ano03-70, Ano03-67, Ano03-68, Ano03e, Ano03-69, Ano03-71, Ano03-74, Ano03-75, Ano03-89, Ano03-90, Ano04e, Ano04d, Ano04f, Ano04g, Ano04h, Ano04j, Ano04i, Ano04k, Ano04n, Ano04l, Ano04m, Ano04o, Ano04p, Ano04r, Ano04q, Ano04-28, Ano04s, Ano04u, Ano04t, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-30, Ano04-31, Ano04-29, Ano04-32, Ano00-78, Ano00-79, Ano03-32]. Painlevé [GT00, RGTT01, RGTT03]. pair [CRK02, DH08b, HHY09, HY00, HCY00, IK03, LdSF00]. pairs [AHK00, CFY08]. pairwise [LY04, PT00]. Panagiotis [Ano01h]. panel [Ano04e]. Panovsky [VAR03]. pantograph [SD09]. Papadimitriou [Ano04-30]. paper [Ano00l, Ano00p, Ano00-80, Ano00-81, Ano03-49, Ano03-68, Ano03-74, JY11, Liu08b, Pen03]. paperback [Ano00-74, Ano04e, Ano04g]. papers [Ano04x, HPT05, Ano03n]. Parabolic [DS06a, LCL06, SBS06, ZW06, ACJ07, AO05, AY06, BG07a, CJO05, CDJ04, Deh02, DR04, DL05, DDX04, DXZ05, EA03a, FFXL05, GJ07, HH02b, Hor08, JZ08, KO09a, KPP00, Mar07, OZY05, PB01, Pao01b, SM00, TKC02, TD07, Tia08, WFC04a, WY07, WS04b, YS04, ZW04, ZHKW05]. paradigm [Ano03-68]. Paradis [Ano03-85]. Parallel [AMV08, CY05, CM02, HWHG08, LKS^{+}06, SP08c, Wan03b, ZZ06, Ano00-89, Ano04-28, Bai00, Bar01d, BY03, BGS04, Cha01a, GS01, GGD00, GK00, HS02, He00b, HBP06, KK06d, LUZ05, LCL^{+}09, LLCS02, LLL02, MPS04, PLL03, Pet05, PIR04, RKS08, RZH03, Sha09, SL00b, Sir00, TH02, XXLD04, Zhu05, Ano04s]. Parallel-iterated [CM02]. parallelization [HKL09]. parallelogramic [YMTE05]. Parameter [CKKK06, Col02, Col04, CM06, WFC04a, Zal03, CXZ09, DLB03, DWL08, FM08b, FW04b, HL08a, HHC08, KGWL08, LP08c, LP08b, LKS^{+}06, LZL04, LFZ08, Liu08d, Shi03, WTZY05, ZX09]. Parameter-free [Zal03]. parameterization [CTYW05]. Parameterizations [KT02]. parameterized [WW03]. parameters [BS06, CS00a, CSdS08, CM00, EMD08, FZ07a, GAT07, HB04, JS03, Lo 02, MV08b, PKM08, Syn08, WKL06, WW08d]. Parametric [FH07, ABS09, ID10, JPS09, JW05, NWL08, PL05b, Sun05b, Sun05c, WY09b, WW08c]. paramonotonicity [CLY08]. parasite [MM06]. parasitic [Lee09, YW03b]. parasitoid [JD07]. paraxial [QS03]. Pareto [YZLL09, Din00b, DPJ03, SM02]. Pareto-optimal [YZLL09]. Parhami [Ano00s, Ano00t]. parity [Boy07]. Park [Ano01r]. Parker [Ano03-70]. parsimony [BGLT08]. parsing [KC03a]. Part [CS08, AVA01, BDO04, HRM08, LYC07, SSLL09, Ikh08]. Partial [AG01c, AG01b, LY08b, Ste06, AETEZ07a, AML02, Ano04f, Ari07a, BG08, BA09c, BS04, Bs03a, CKZ01, DS04, Dub08, FFK02, FT03, GL03a, GJ07, GS07b, GL00, HPG06, HLP03, JSF00, KH00, KHA07, LZ02, LLD02, LC04, LLTL05, MAH08, MO07, OM09, PS03b, PB02, SERG09, SMH08, SK09, TD07, TZ04b, TX04, WKL06, WX03b, WG08b, XTX01, XLZ02, XT04b, ZD01, ZY02, ZA03a, ZD03, ZT04, Ano03-59]. partial- [GS07b]. partial-difference [DS04]. partially [Nak06, Ver03]. Particle [Gre03, LLCS02, MIPT08, AK08a, AF09, Ano03-44, BX02, HTKM09, Kor00a, KLZZ09, KW03, KH09, LWLJ09, LCL^{+}09, MG08a, MLW09, NLLL09, SP08c, YZLL09, ZZJ^{+}04, ZZZ09]. particle-fluid [ZZJ^{+}04]. particles [HK08, ITR08, Mar09, OKCO08]. particular [BD03]. Particulate [SS09e, HK08]. partition [Pla07, WLL06, WL07a]. partitioned [AK07, PV01, TT09]. Partitioning [Zla01, LLCS02, MP05, OY06, PST08, SB00, WSO08, ZZ02b, HTZY03]. partitions [BKK08, Kr05]. Parts [PL01b]. party [LLS^{+}09a]. Pasca [Ano04q]. Pascal [Ano04-27]. Passification [LZL06]. passing [Ano00-88, Ano00-89]. Passivity [LZL06]. password [CYH05]. past [Ari07b, Ari09a, Ari09b, Ari09c, EM09b, SMH08, WWW^{+}09]. past-sequence-dependent [WWW^{+}09]. patch [XC00]. patches [Gou02]. patents [KWT02, Ano03-60]. Path [QW09, HCY00, HL02c, Hua05, KC05b, MP08b, PS07a]. paths [BR05a, FKM05, LTTH09]. Patricia [Ano03l]. Patrikalakis [Ano03-69]. pattern [WYL^{+}09, WS08b, WYSS09]. patterned [WSG08]. Patterns [Wan06a]. Patwardhan [Ano00-78, Ano00-79]. Paul [Ano00-44, Ano00-45, Ano01i, Ano03j, Ano03-27, Ano03x]. Paulus [Ano00-32, Ano00-33]. payment [BCJ08, Jua03]. payoff [YLD08]. PBVPs [SL07b, WW08b]. PCA [LY06b]. PCG [Ara01, Kur08, NV08]. PDE [Cat05, FFK02, Mic03]. PDEs [AR07, BNH08, BR05b, Fas02, LF03, Waz07c]. peaky [Abu01]. Pearson [ND05a]. pedestrian [Dog08]. peer [WBSP08]. Peirce [Ano03-83]. Penalty [HZ07, Nie06, CLS07]. pencil [EGR03]. pendulum [Dor01]. penetration [De 04]. Peng [Ano03-68]. Penrose [LLW08, PST08, WW00]. peptide [Bru08]. Perceptual [Ano03-61]. perfect [MM08b, SC02]. perfectly [TBDL09]. Performability [NS00, Nab04, Pla06]. Performance [AHRT09, AG03, GS07b, LLL02, VG08, YMI06, ZBS03, BM00, GYPL08, GYPL09, HW08, KZ02, LRL09, Pet05, WL06, WGF09, WFT^{+}09, ZL07c]. Pergamon [Ano00-85]. Perić [Ano03r]. Period [MM06, DCR07, FM08b, KH02c, Pan07b, YLD08]. Period-doubling [MM06]. period-five [FM08b]. Periodic [AZ01, AlS08, BX07, CYG06, CL07a, CH04, Che04b, DHM05, DE06b, GY06, HLY06, Kot01, LH06, Liu06b, LJ08b, MH07, PG04, RC07, WG06, ZWCW00, ZL03, ZWW04, ZLJW06, ZLZ09, Zha09b, ZM02, Zho03, ZT07, AB08a, Abu01, AOW05, AB08b, AC03, BCGA08, CYG05, CZ03b, CH05b, CO07, CK02c, FW00, FW04a, Gon09, HH04b, HC09, HZ08b, Huo04, Ji02, JA03, JY08, KK05, KS03b, Les02, LZL04, Li04b, LW05, LDZW05, LZJ06, tLW08, Li08b, LFZ08, LY06a, Liu07a, Liu08a, LL08b, MNIS06, Net05, Net07, NRL00, Ou08, PF01, SA02, Son09, Sta02, SX06, Ter00, Top04, WJX04, WM04, Wan08a, WFY09, Waz05d, WY09b, Wen00, Wen02, WLZ07, WH07, WLW08b, WL08c, WW08d, WLD09, cWcX09, cW09b, XC06, XXD06, Yan05b, ZM94, ZSG04, Zen09, ZJLW06, ZLH07, ZSPZ08, ZS08a, ZW03b, ZHKW05, ZLLL08, ZC01, ZCC04]. periodical [AD03]. Periodicity [BS04, CS04a, AR09, AR10, AZ01, Bs03a, FM08b, HZ05, YH03, ZSZ06, ZLLL08, ZHL04]. periods [ZHZ00]. peripheral [ICM09]. perishable [KS00a]. Peristaltic [AH08, HAKA07, MP02b]. Perl [Ano00-79, Ano00-78, Ano00-62, Ano01s]. Permanence [CWL07, Huo05, NTHP09, TL00, CF04, Cui02, LY08b]. permanental [Gyi00]. permeability [STH08]. permeable [MM00]. Permutation [JS08, Ano03-58]. permutations [Köh00]. Perseus [Ano00-29, Ano00-59]. Persistence [WM04, XC00, Wei05]. person [KL02]. Personal [LLB05, Blo00]. perspective [Ano03-45]. perspectives [Ano03-78, Ano03-81]. Perturbation [Gal05, Ano04w, Ano04v, Ari07b, Ari09a, Ari09b, Ari09c, BN05, Bel09, BG08, BBA09, BA09c, BG11, CL07a, CJK09, EM09b, GSZ02, GTJ07, GSN08, Guo07, HLSL03, He03, He09b, HRM08, IS04, KP03, KA09a, KGB07, LDK^{+}06, LSZ08, Lu09b, MZ07, MO07, MEA09, NVAR03, NMD08, ÖY08, SNG09, Sho09a, SHI09, SK09, WW00, WY09b, Xia07, Xu07a, Yil08, Yil09, Yus09, dZmClQ09, CC04b]. perturbations [AZCK05, Che01b, DGPT08, Han04, JKP02b, WF07, Wei05, XL01]. Perturbed [Bai01, JJ01, AL04, AE07, CN01, KP01, KK09, Li00a, Li02, LTWM08, Pit03, SR03, SR04, Sir00, SK06b, SNA09, VR02, WYL08b, Yan04b, YMMC03, ZCS04, Zho06]. pest [WS08a]. pesticide [CGPB08]. Peter [Ano00i, Ano01u, Ano03-47, Ano03-70, Ano03-67, Ano04p]. Petersen [MWH08]. petri [IUCR06]. petrochemical [She05]. Petrov [BLS05]. Petrowsky [ES08a]. Petter [Ano04f]. Pfister [Ano03e]. PH [LWZ04]. pharmaceutical [BBM08]. Phase [ASL06, LL06a, WZE^{+}06, AORS03, AP03, BJK^{+}09, BGR09, BJK^{+}08, EW00a, Eng04, GS06, KW03, LZ09, LGGSP06, MAR05b, Nik06, PP01a, QC06, Slo07, SZ09a, SF01, WCL02, WW09c, ZZJ^{+}04]. phase-1 [PP01a]. Phase-Field [ASL06]. phase-I [LGGSP06]. phase-lock [EW00a, Eng04]. phases [DP07]. phasing [Lan08]. phenomena [AA08, CQJ03, MAR05b, MKSB08, dSM02]. phenomenon [FZ07a]. Philip [Ano03g, Ano03-74]. Phillips [Ano03g]. Physica [Ano00-47]. Physica-Verlag [Ano00-47]. Physical [Ano00-48, Ano00-49, Ish00, LWZ08, PM01, YCCY02]. physics [Ano03s, Ano03g]. physiological [Zho07]. Picard [Ana05, AA09a, DWL08]. pictorial [CS00b]. Piecewise [GL03a, WY06, AB08b, GLC09, GK04b, HF00, ITT07, KU01, LS03, LYL07, Nie01a, SYL05, WZ08a, ZW09b]. Piecewise-linearized [GL03a]. Piero [Ano01n]. Pierre [Ano03-52, dKLR09]. Pinning [XCL^{+}08]. pipe [HHC08]. pipeline [MG08b]. pipes [APS07]. Pironneaufor [WFC04b]. pixel [BFKL08, Jum01]. pixellation [MRÁR08]. placement [ISAAM09]. planar [Aro08, HSM00, HS02, Lee00, Liu09c, Yan05b]. Planck [AU03]. plane [Buh06, DD00, Hav08, MOP06, OS08a, PS01, Yan04b, Yav05b, Yav05d, Yav07a, Yav07b]. plankton [JB05, ZM02]. planning [Bel01, SKU01]. plant [AGS09]. plants [IUCR06, SKU01]. plasmas [TG03, Ano03-82]. plastic [NLR00]. plate [FAF09, GW06, HM05a, PKP09, SMH08, Ste07, SZL02, VSKF08]. plates [GGD00, PB01, SHKL08]. platform [Ano00-41, WFTL09]. platforms [HSM00]. players [EAE09]. pleasures [Ano03-76]. Plenum [Ano00-71]. plus [Blo00, NV08, Roj08, Ano03-73]. PM [WS00]. Pochhammer [SK04]. Pocket [Ano01j]. POD [KX05]. POD-based [KX05]. Poggio [Ano03-61]. Poincaré [Ana08, DBX04, Mar00, Ung01, WC04]. Point [DZZT06, FY03b, yL06, LX06, Yav05c, AO01b, AO01a, AO01c, ATT03b, AORS03, AN09, And03, AK08b, Ano03-68, Bai08, BW05, BW06c, Bou00, CYG06, CM04, CWY08, CCSL04, CS04b, CL08a, Che03, CKL07, CT02b, ĆRRU08, CM08, DS04, Dha03, Dha04, Dha06, Dha07, DG05, Du08, EGA03, Ely04, FR04, FT05, FM03, Gra08, Gu07, Ha01, HL02a, HG03, HLP05, IK03, INP08, Jan09, bJQxC09, KP01, KK09, Kam09, Kan02, KH02b, Kar08, Lan04b, LLJ09b, Liu02, Liu04b, Liu04c, LLTL05, LLW07, LDK08, Ma00, Ma01, Ma04, Mar00, Mar07, Més05, MEG08, NT08, PMQR09, PLG07, POS03, PKT07, QZ05, RR09a, SSX09, SPW08, SS06, Sun09a, TXC02, TGS07, Ver08b, Ver08a, Ver09, Ver12, Wan02d, Wan03a, WG07b, WSL08, Wan09d, Won08, Xie01, XWZ04, XG07, Yao04, Yao05, Yas08, ZM04a, ZL08c, ZYW08, ZL08d, ZJ08]. point [Zha09d, Zha08d, ZG09b]. Point-to-point [FY03b, Yav05c]. point-wise [Kan02]. points [AP01, AA09c, Bai01, Bao03, CC04a, CKQ08, FG00, udK07, GN09, Hua00, KM07, KDY09, LLX05, Lin07, LHLD08, LdSF00, Moo02, OI00, OS09, PT00, Rad09, Roe09, SYJJ09, SV09, Sya01a, TCH07, Wan07a, WL08b, WL09a, YYZ07, ZP02, Zha08b, Zha02c, ZC09b]. Pointwise [CK03, GQL05, GX04, LW08b]. Poiseuille [Pru09]. Poisson [AM08b, BSR02, HL06, Jay04, JDH03, KK06c, KL03, LKS^{+}06, Mac06, PS08a, SALK02]. Pokman [Ano03-64]. polar [AM08b, KP05a]. poles [LS04a, VBG07]. policies [CHW07, Dra06, HY06, ODO00, QNN03, SI03, SO03, Saw03]. policy [Ano03v, Ano03-29, Ano03-80, Ano03-38, KS00a, KS03a, KK03, LCT08, LWY09, MNIS06, PS08a, WS00, WW04a, WZ07a, WM09, YE01, Ano04n]. Polk [Ano03o]. polling [KK06c]. polluted [WM04]. pollution [CS08, Zla01]. poly [YX09a]. poly-scale [YX09a]. Polyak [HS04a]. Polycrystalline [ASL06]. polygon [GS04, Has03]. Polygonal [Köh00]. polygons [DW08b, Köh00]. polygroups [DLF09, KDY09]. polyhedral [Lóp09]. polyhedron [WY01]. polymer [MUHdH03]. polymers [WK03]. Polynomial [Bar01b, She04a, ANZ09, BY03, BGLT08, BGP04, Boy01, Boy07, BCT00, CBJ01, DJL04, DFSS08, Fly06, GL06, GM00, HYZ02, HSWL08, HL04, Ikh08, JCZ09, Lai08, LL07, LC02, MCRÁ05, Nie01b, Pan01, Pan05a, Pan05b, PV00, Tay01, Tel08, XH02, XHZ05, ZL07a]. polynomial-growth [Boy01]. polynomial-time [DFSS08]. polynomialdifferential [TZ04a]. polynomials [AG01b, AG09, AAR03, BNR05, BC02, BC03, BCGA05, BCA09, BCR04, CCS02, CBGVPP09, FF02, FF01, GM09a, Gra08, KM03a, KFMH06, KA00a, KR07b, KI04, LS06a, Liu08c, LS06b, MPTV06, NR01, NR04, NMD09, PCRR09, PV01, Rap01, RTW05, SPV01, Sya01a, Trz01, WJW08, WW09a, WlW08a, Win03, WSZ03, XY04, XX08, YC06, YW03a, ZY08a]. ponens [WHS08]. pool [GSNDB^{+}08]. Popoviciu [DBGSP08, Wu08c]. popular [Bel05a]. Population [BFL06, Lo 02, CWL07, CCW00, Cui02, GMG09, Li09b, LP04, MMM04, RTS07, SD07, WW04a, WS08a, ZWCW00, Ano03-62]. populations [AVAGP08, Bra08, GPAA^{+}09, Gou01]. Porous [WZE^{+}06, CM06, GMS08, HHK06, HASA07, LY08d, MG03, NMD09, SMH08, SR07, VM09]. Portable [Ano00-89].

Portela [Ano03-34]. portfolio [Ana06, Hua08c]. posed [HB04]. posedness [FH07, FHH08, Gup02a, WX09a]. posets [BKN09]. position [LTTH09]. positioning [BGA08, Hur00]. positions [KYJ03]. Positive [AOS07, ACM05, AW09, BW06c, CDM01, DG05, ERRY01, FZG07, HLY06, KKZ02, LA03, LZL04, LS04b, Li04b, LLX05, Li08b, LFZ08, LS09, LLJ09b, LXH05, Liu02, LQ04, Liu04b, Liu04c, LY06a, LLW07, Ma01, Ma04, MO01, Pal01, SSX09, SC03, Son09, Sta02, TGS07, Wan08a, WFY09, WW08d, XXD06, Yan03c, Yan06a, ZYG08, ZLJW06, ZG09a, ZG09b, ZH02, ZCW06, Agr08, AD08, AA01, AO08b, AP01, Ayk04, BF03, BG04a, BLD07, Bai08, BBP09, BW05, Cap00, CH05b, CWZZ07, CZ08, Den07, DSS08, DM08, EM09c, ES01, FLR05, Fer01, Gup03, HYLW00, HZ08b, Huo04, Jan09, Jia00b, JA03, JX04, JZ03, KÖ08a, Kar09, KLW09, Lai08, LWZ08, LLWY00, Leu00, LFZ00, LW05, LDZW05, LZJ06, LHL07, LLT08, LBW05, LZ02, LLD02, LG04b, LY05b, LDK08, LL08b, Ma00, MDG05, MZ09, MZ04, MOP06, Öca07, PLG07, PL03, RSE01, SW04]. positive [Sot02, SPW08, Sun00, Sun05a, SZ05, SX06, SL07b, Sun09a, TYP00, WJX04, WZ00, WG07b, WG08a, WL09a, Waz01, Wen00, Won00, WCW08, WWC08, WWYL09, WLZ07, WLW08b, Yan04d, Yan07, YW07, YW08a, Yao04, Yao05, Yao07, Yas08, YTEC02, YWYL07, Zen09, ZJLW06, ZYW08, Zha09d, ZC09a]. positive-definite [YTEC02]. positivity [GW03a]. positone [ZYW08]. possess [Dra08]. possibilities [CJV07]. possibility [Isk02, Isk04, Isk08]. Possibly [BFL06]. post [Ano01-28]. post-World [Ano01-28]. Postaenomic [Ano03-41]. Posteriori [Ale06, Had03, PPP^{+}06, RW03, AN06, Bla04, VH06]. postman [WW02a]. Postmodern [PR01]. Postprocessing [Sar06a]. postprocessor [AN06]. postservice [Kaw06]. postshock [SD03]. potential [AC05, All02, BRC00, CMS04, DCS09, HZ08a, KL05, Ano03-71]. potentials [GZ06, HZZ02, KK06a]. Powell [CL08b]. power [BGP04, Coh00, GMŚ07, HCWC07, ISAAM09, IUCR06, LTYB09, LCL^{+}09, LK04, NWL08, NLLL09, Pan05a, SL09, Sür08, Trz01, Ung05, YE01]. power-law-like [GMŚ07]. power-mean [HCWC07]. power-of-two [YE01]. powerful [Waz07a]. powers [Pan07b]. Powersums [Ben00, Cha01b]. Practical [ZS04, BL01, CL00, Jua03, Ano00-58]. practice [Ano03-72, Kor05, PDBH03]. precedence [Wan09b]. Precise [FZ07b, KYJ03, PL05a, YLZ08a, ZT08, VAR03]. precisely [Ano03-47]. Preconditioned [Bax02, Gra03, LCC08, Tou05, Yan03d]. preconditioner [LHR09, MAK03]. Preconditioners [CKL06, JR01, TC09, BFM04, CJ05, CT09, JLW04, JL09b, NV08]. Preconditioning [Hip06, KF05b, AMV08, CZ00, GKM08, KL05, LZ09, PIM^{+}08]. Predator [HLY06, Wan06a, ADN02, CW08, Col04, Cui02, DH08c, HC09, HZ08b, Huo05, NTHP09, SC00b, SDCL08, WLW08b, XC00, XCD07]. predators [XC06]. Predazzi [Ano03-44]. predicting [LHLW07]. Prediction [Gla07b, BCD08, CKCG03, CML^{+}09, CCF06, GPT03, HXG03, JAM02, KT09, MRP08, WLT07]. predictions [Har01, SLL03]. predictor [PSS07]. predictors [Sya02]. preemptive [LKH03, LWY09]. Preface [AO01d, Ana00b, Ana04f, Ano03-95, Bar01c, CP03, Dey00b, DLO06, Fas06, Gal02c, Gal06b, GKKL06, JAV08, KH02a, LD06, Net06, Osa03, PC09, RXY06, SAVA03, SY03b, SBGS01, SBGS05, Tay00]. preference [CL02b, ESNP06, LH05b, LWL08]. Preinvex [dSOGRMRL04]. Preinvexity [SL06c, Ant05]. Prentice [Ano03-59, Ano04r]. preproximity [ZAAESG09]. prescribed [MSS08, RSR01, RSE01, Sot01]. presence [AL08a, COEP00, RRR07]. presentation [Ano04u, LTS03]. presentations [Tay01]. presenting [LMSH04]. Preservation [LGY09, PP01b, ZZ09a, MG08c, Nko00]. Preserving [Jay04, AG01c, AG01b, Ano03-42, CLMM06, ID05, ID08]. Presidency [Bel05a]. Presidential [Bel05b]. Press [Ano00e, Ano00j, Ano00f, Ano00k, Ano00-74, Ano00l, Ano00s, Ano00m, Ano00t, Ano00u, Ano00p, Ano00n, Ano00-76, Ano00-30, Ano00-77, Ano00-31, Ano00-38, Ano00-42, Ano00-43, Ano00-41, Ano00-51, Ano00-54, Ano00-64, Ano00-80, Ano00-69, Ano00-81, Ano00-83, Ano00-84, Ano00-88, Ano00-89, Ano01e, Ano01y, Ano01m, Ano01n, Ano01z, Ano01o, Ano01l, Ano01q, Ano01r, Ano01u, Ano01t, Ano01x, Ano01v, Ano01-28, Ano03-76, Ano03g, Ano03f, Ano03-36, Ano03-77, Ano03j, Ano03k, Ano03l, Ano03t, Ano03m, Ano03n, Ano03s, Ano03-78, Ano03o, Ano03v, Ano03y, Ano03-27, Ano03-79, Ano03x, Ano03z, Ano03-29, Ano03-28, Ano03-80, Ano03-81, Ano03-32, Ano03-33, Ano03-38, Ano03-41, Ano03-85, Ano03-43, Ano03-83, Ano03-47, Ano03-50, Ano03-49, Ano03-53, Ano03-86, Ano03-55, Ano03-54, Ano03-57, Ano03-60]. Press [Ano03-61, Ano03-63, Ano03-62, Ano03-87, Ano03-66, Ano03-70, Ano03-68, Ano03-71, Ano03-89, Ano03-90, Ano04e, Ano04d, Ano04g, Ano04h, Ano04n, Ano04l, Ano04-27, Ano04-28, Ano04-30, Ano04-31, Ano04-29]. Press/The [Ano01r]. Pressure [DHY06, JY09, Bla07, HD08, STV08]. Pressure-Driven [DHY06]. pressure-sensitive [HD08]. pressure-stabilized [Bla07]. pressures [SR07]. prestarlike [AM08c]. previous [Ano04-40, Ano04-41]. Prey [HLY06, Wan06a, ADN02, CW08, Col04, Cui02, DH08c, HC09, HZ08b, Huo05, NTHP09, SC00b, SDCL08, WLW08b, XC00, XCD07]. preys [XC06]. Price [MM08b, HLY08, Lon03]. priceless [Ano03-63]. Pricing [YLD08, BM06, BCJ08, CKKK06, CNPP08, DLL08, Hon02, YC07b, Ano03-63]. Primal [KP06, Ano03-68, CKL06, LGGSP06]. primal-dual [Ano03-68]. primary [KCD05]. prime [Tan03b]. primes [BDT04]. Primitive [Ste06, LH08]. Princeton [Ano03-27, Ano03-49, Ano03-68]. principal [LJ08a]. principle [BK00, BHHA08, Dib01, Hor08, Li07, LV03, WF07, Wu03]. principles [AA05, Ano03-52, HLCL00, Ano04d, Ano00-59]. Priori [RPO03, CBJ01, PJJP04, WWS^{+}06]. Priority [WNS09, GC02, LKH03, PdlF07]. Priority-based [WNS09]. priority-discipline [PdlF07]. prisms [AA08]. privacy [Ano00v]. pro [Ano04o]. probabilistic [CHL09, Ind02, MT09, PP01b, Ren03]. Probabilities [CF04]. Probability [Ata09, Kum06, Ano03-27, BDA02, Hoh01, JWSS04, LWZ04, YC07a, ZZ09b, Ano04t]. probability-based [ZZ09b]. Problem [CKK06, DS06a, GC06, GCN06, SPE06, Ada04, AVO08, AGS09, ARM08, AD03, AA01, AK08b, Ano00-94, Ari09b, BK00, BSS98, BLD07, BC00, BGLT08, BD04c, BIKL06, Cab01, CG08, Che01b, Che04b, CX04, CO07, CHZ09, Che08b, DD00, DS02, De 04, DHM05, DG05, ES05, FLfFL08, FS04, FR04, FM08c, FS07, GY08, Gre03, GS06, HRP01, HMCR00, Has04, He02, HG03, HZ08a, Hof03, HL02c, HLM04, Hua05, JKR08, JN07, JKP02a, Jia01b, Jia01c, Jia02, JZ03, KA09a, KK02, KK06a, KH00, KK04, KS00b, KLW09, Kru03a, Kru03b, Kru05, Kru08, KC05b, KH09, LHC08, LUZ05, LCGX06, LS08, LLJ09b, LG09a, LX00, LWY02, LV03, LWY04, Liu02, LQ04, LY05b, LCWL06a, LDK08, LSN05, LJ08b, LHWW08, Ma01, MDG05, MZ09, MM08a, MRB02, MCC09, Man05, Mar07, Mar05a, MV08a, Mur05a, Mur07a, Nak06]. problem [OTB^{+}08, OEV09, PLG07, PW05, PSES07, QW09, Ryo00a, Ser00, SP08a, SW07, SL00b, Slo07, SZ09a, SWX06, Sot02, Sta07, SZL02, Sun09a, TCH07, TGS07, VNFF08, Wan00e, WWWL08, WW09c, WFC04b, WYZY01, WI08, WTY03, WTZY05, Xie00, XCH07, XZ08a, Yan02, Yan04b, Yan05a, Yan07, YZHY09, YL09a, YE01, Yao04, Yao07, Yil09, YD08, ZLY06, ZL08c, ZYW08, ZJ08, Zha09d, ZG09b, ZZZ02, dSOGRMRL04]. Problems [Guy00, LET06, Liu06b, Mig06, SW06, ZLS06, ZLJW06, ACD01, AP08, AS03b, ATT03b, AOLL04, AOS07, AM09, AS01, ACJ07, Amm07, And03, AW09, Ano00o, Ano00-31, Ano00-35, Ano00-58, Ano03f, Ano03h, AMV08, AS08, AC03, ACH01, Ayk04, BG04a, Bai08, BSR02, BN05, BCS03, BDGR01, BN00, Bla07, BG04b, BW05, BW06c, BS03b, Bou00, BRMS01, BMM^{+}06, BLS05, CH01, CG05, CC09a, CDJR03, CS02a, CM04, CN01, CGY09b, CKY04, CS02b, CK02a, CBJ01, CWY02, CGGC02, Che04f, Che04a, CWZZ07, CL08a, Che00b, CE02, CKL07, CLS07, CH00c, CW05, CDJ04, CZ08, CZ09c, Cui09, CM08, DR04, Din00c, DW02a, DE06a, DH07, Du08, EGZ04, EGA03, EMM09, ES08b, EBVS06, FHH08, FK01, FL03, FZG07, Gal02a, Gal08c, Gal08b, GSZ02, GNM07, GL09c, GKM08, GOMA07, GHHD08, GNR05]. problems [GH03, Gra03, GL00, GHJ06, Ha01, HL02a, HH02b, HL02b, HS04a, Han06, HLW07, Han08, HTKM09, HTZY03, HLP05, HW09, HT02, HY03, HW05b, HB04, Hor08, HLC05, HZ07, HZ09b, IS04, Ili04, IN07, Jan05, Jan07, Jan09, Ji02, Jia00b, JX04, JC00, JPV01, JP01, JPV02, KM07, KP01, KP03, KK09, KN04, Kan09, KA09b, KO09a, Kar08, KF05b, KK08a, Kel06, KR07a, KPP00, KL08b, KKZ02, KK08b, KS03b, KLZZ09, KY07a, KY07b, LP08a, LY09a, Lan04b, LR03, LW09b, LLWY00, LC06b, Les05, LH00b, Li00a, Li02, LZZ04, Li04b, LCGX06, LS09, LHW04, LDG05, LJL06, LSZ08, LW09e, LT04, LT01, LWG04, Liu04a, Liu04b, Liu04c, LCWL06b, LCWL06c, Liu07a, LLW07, Lóp09, Lu07, Ma00, Ma04, Ma07, MS04a, MM05b, MW00, MEG08, Mou02, NLL07, NLL08, NLR00, NVAR03, NF03, Net05, Net07, NRL08, Nob06]. problems [NMD07, NMD08, OEV09, Pal01, PG04, PGT08, Pao03, PF01, PXGL09, PY09, Pet04, PD06, PDBH03, Ram05, RMC08, RR08, RW03, Ryo00b, RSK02, Ryo03, Ryo04, Ryo07, SV06, SA05, SR04, SC00a, SC03, SWL^{+}04, Shi05, Sir00, SF01, SF07, Sta02, SPW08, SBSY06, SK06b, SZ09b, SB00, SL06a, TS00a, TLLM08, Tho01, TL06, TT09, Top04, VJ04, VH06, VC03, WLS01, WXW01, WW02a, Wan03a, Wan03b, WG07b, WLW07, WJA08, WWW08, WSL08, Wan08a, WG08a, WFY09, Wan09d, Waz00, Waz01, WA08, WFS03, Won00, WX03b, WCW08, Won08, WX00b, Wu04, WK05, Wu08a, Wu09, Xia07, XXD06, XD02, XZ08b, XC04b, XG07, YAVB08, Yan03d, Yan06a, YW07, Yao05, Yas08, YTCF08, YL07, YS04, YLL08, Zal02, Zal03, Zha01d, ZZJ^{+}04, ZM04a, ZhYL05, ZL08d, Zha09e, ZH09, Zha09c, ZG09a, Zha08d, ZC09a, ZXL07]. problems [ZZH03, dZmClQ09, ZCC04, ZZ06, dSRGRMRL08]. procedure [JSK^{+}04, PW00, SF01, Ver09, YS04]. procedures [HGH02, SP08b]. Proceedings [Ano00-43, Ano01e, Ano00-85]. process [AF02, AO08b, Bai01, BB01, CPP07, CPP08, Che04g, CH00b, ET09, GMSRP08, GLR05, HM03b, Hua00, Jon07, LP00, Mac06, MM05a, Moo02, TT03, WCL02, WZ07a, WLW09, YW09, Yu04, Zha08c]. Processes [yL06, LL06a, Ano00-71, APV02, BPvZ03, Bib06, CCRMOG04, Çin03, CSPJ05, ES01, FM07, GL03b, HCLJ00, HF04, HL07, KKY06, KL08a, KL03, yL08b, LHP07, Pan07c, SHS05, SW08, Syn08, Tur04, XLL08]. processing [AMS08, Ano01e, Ano01r, Ano03-52, CFM02, Can03, Iac07, IYN03, KK03, KY07a, LYL00a, LCYL00, LCL00, LG09b, MWZ09]. Processor [HRP04, LL06b, TBR01]. Processor-efficient [HRP04]. processor-sharing [LL06b]. produced [TK02]. producing [Tan04, YLZG01]. Product [Che06, WFTL09, BUP09b, CHW07, CS05b, Coh00, Dha03, Dra06, EAA09, MG08b, YMI06]. production [HY06, MM05a, SL00b, WS00, WCL02]. production-inventory [HY06, MM05a, WCL02]. production/PM [WS00]. productivity [Ano01q]. products [AK07, DCR07, HSS01, Sri03, YWZ03]. professional [Ano01-28]. Professor [Ano04c]. Profile [ZW06, DDCB04, DDX04]. profit [PdlF09]. programmers [Ano03-30]. Programming [Bel06, Nie06, Amm07, Ano00-43, Ano00-63, Ano00-89, Ano01k, Ano01p, Ano03-34, Ano04u, Ant05, Ant08, BW06a, DTXW05, Gal06a, GL05, GOMA07, GMMB06, GMMB09, GLTS08, Isk02, Isk04, Isk08, KN04, KH09, LMA08, LW09b, LHW04, LT09b, LHWW08, LLL02, Mis05, PKM08, SKU01, SKW02, SHWF05, SWL^{+}04, SF01, SF07, WZG04, WFLC05, WWWL08, WL08a, Wu04, Wu08a, YH04, YZHY09, Zal02, Zal03, ZhYL05, ZXL07, dORM08, Ano00-43, Ano01w, Ano01-27, Ano04f, Ano00-60, Ano00-61, Ano00-62, Ano04u]. programs [HL09b, Isk04, Isk08, Tao06, Wan02a]. progression [GMSRP08]. progressive [LBW05]. progressively [MJ02]. Project [Lin08b, ZYY08, YC07a]. projected [Noo02b, SHWF09, WW04b]. Projection [Noo00b, Ver01b, Alb02, CG05, CD05a, HL02b, HS04a, Han06, fHCH09, HJL05, Jia01a, JZT07, LW02, qWjH07, YHS08, Zha01d]. Projection-splitting [Noo00b]. projections [CCS00, PO04]. projective [Pan00]. projectors [BB08a]. proliferation [BI01]. Prolog [Cse06]. Prolog-assisted [Cse06]. promising [HWX07, SNT08]. proof [Cip08, Pen03]. propagating [BT03]. propagation [DS05a, GMS08, Ikh08, MKF09]. propagative [Kru03b]. propelled [Sta07]. Proper [CZLN06]. properly [HM04]. Properties [AVA01, BR05a, BV09, FH03, Jin09, Kar03, Nak06, PS03a, SNG06, TA06, qWjH07, XT02, XA09, BM05a, CK06, CP01b, DJL04, DK04, EM09a, Gal08a, Hor01, IWL04, LCHL06, LZJ04, LZ04a, NT02, OSSY04, PP01b, QJZ09, RTW05, XT01, XT04a, Yan04c, Yen08, ZSKZ03, Zha08a, ZZ09a, Ano00o]. Property [DZZT06, DH03, Gu08, GX04, IL04, ND01, NON^{+}08, RGTT03, SLTH09, Tan03a]. propositional [Zha08a]. propylene [CS08]. prosodic [NKCS08]. Protasi [Ano00o]. protected [CYH05]. protein [BCD08, Dar08, LY09c, MRP08]. protocol [CYH05, WGF09, YSH03]. protocols [Ano00-34, LLS^{+}09a]. prototyping [Ano00x]. provider [HY08]. Proximal [Alb02, LX06, CLY08, CHZ09, Han08, Ver08b, Ver08a, Ver09, Ver12]. proximal-type [CHZ09]. proximity [Pan01]. proxy [Ano03v, YXZL09]. Pseudo [LLB05, WL09b, CM02, Inc09, MCQK09]. pseudo-contractions [MCQK09]. pseudo-contractive [Inc09]. Pseudo-isochronicity [WL09b]. Pseudo-Random [LLB05]. pseudoconservation [KK06c]. pseudocontractions [Moo02, OI00]. pseudocontractive [Gu07, HCLJ00, HGH02, OS09, ZP02]. pseudogenes [RMC08]. pseudolite [KYJ03, KKS^{+}04a]. pseudomonotone [Ver03]. pseudomonotonicity [CLY08]. Pseudorandom [GGG03]. Pseudospectral [Ras06]. PSO [IMPF08]. Publications [Ano03p, Ano04q, Ano04x, Ano04-32]. publicity [PdlF09]. Publishers [Ano00-71, Ano04k]. Publishing [Ano00-55]. Pulsatile [BR06, MP08a, MG03]. pulsating [JZC06]. Pulse [CQJ03, PLG^{+}09]. Pultrusion [CPP07]. pupil [Iac07]. purchase [YC07b]. purchaser [MGG05]. Purdy [Ano01j]. Pure [XC06]. Pure-delay [XC06]. Puri [Ano03-72]. puzzle [Cse06]. puzzlers [Ano03-27]. puzzles [Ano03-58]. Python [Ano00-63].

Q.I. [CM01a]. QBD [LL06a]. QP [bJQxC09]. QP-free [bJQxC09]. Quadratic [CWB01, Jan04c, AT05, BUP09b, BM04b, BCRS05, Che00b, ESR06, GOMA07, Hau06, He00a, HZ03, Jah04, JCZ09, Sha04, SH01, WWWL08, YJCZ08]. quadratic-form [YJCZ08]. Quadratically [WXS01]. Quadrature [CBGVPP09, DGVP06, ED09, AC02, AC05, BCA09, Che00a, KD08, Kov00, Liu08d, MC07, MCS08, Nie09, Uje04a, Uje07, WX03a, WS04b, YW03a]. quadratures [KR01]. quadrilateral [CLX08, LXZ08]. qualification [ZXL07]. Qualitative [CW08, EM09a, FT01, Min08, Mos08]. quality [Dra06, KT09, LCT08, MS08, MRP08, WFTL09, WFT^{+}09]. quantitative [LCY05, MT09]. quantities [CK03]. quantity [KS00a, LLL08b]. quantization [BD01, Käm00]. quantum [HHRA02, HW05c, LRL09, XXC^{+}09, ZXTL09, Ano00-64, Ano03-64]. quantum-inspired [LRL09, ZXTL09]. quarantine [PLG^{+}09]. Quasi [AI06, yL06, Lei02, LHYM09, MMM04, Bou03, CDM01, CDM03, CGY08a, CGY08b, CKN04, De 02, Din00c, Din04b, DY05, DL08, For01, FT05, FM05, GL09a, GC02, GQL05, HYZ06, Hon02, HBCK00, LKS08a, LZ04a, LLJ09a, Lin04c, LU09a, Luo02, Noo02a, PL05b, PZ07, QS03, SL09, SHWF09, Tia01, Tia05, TCH07, Ubh04, Wan02e, WK03, WS08b, Yan04d, YW03b, ZGY07a, ZGY07b, ZAY08, ZL04]. quasi- [QS03]. quasi-convex [LU09a, Ubh04]. quasi-equilibrium [Din00c]. quasi-hyperbolic [Din04b]. quasi-interpolants [GQL05]. quasi-linear [CDM01, CDM03, De 02, LZ04a, LLJ09a, Luo02, Yan04d, ZL04]. Quasi-Monte [Lei02, Wan02e]. quasi-multiple [YW03b]. quasi-multiquadric [Lin04c]. quasi-Newton [For01, FT05, FM05, SHWF09, WS08b]. Quasi-Nonexpansive [yL06, Tia05]. quasi-nonexpansive-type [TCH07]. quasi-one-dimensional [WK03]. quasi-power-increasing [SL09]. quasi-radial [Hon02]. Quasi-random [MMM04, GC02]. Quasi-self-stabilization [LHYM09]. quasi-symmetric [DL08]. quasi-variational [Bou03, CGY08b, CKN04, HBCK00, LKS08a, Noo02a, PL05b, PZ07, Tia01, ZGY07a]. quasi-variational-like [CGY08a, DY05, GL09a, ZGY07b, ZAY08]. Quasiequilibrium [WTY03]. quasigroups [KP05b]. Quasilinear [Kan09, ACM05, CE02, DL05, DXZ05, Dua09, HG03, HW05b, LD02, Luo03, TRG01b, ZHKW05]. Quasilinearity [Bil05]. quasilinearization [YV05]. quasireversibility [TKC02]. Quaternion [ES07, BHHA08, CPS05, tLjW08, SPV01, Wan05a, Wan05b, Wu08b]. quaternionic [CFY08, DD04, JCZ09, WS07]. quenching [SC00a]. query [LG09b]. quest [Cho08]. question [Ano04q]. queue [AHRT09, ASESESA09, AM04, GC02, GS07b, GM09b, GSS04, Jai05, KCD05, KK03, KA00b, KA02, LKH03, LL06b, LWY09, Mac06, PS08a, SG08, Wan04a, WZ07b, WX09a]. queue-dependent [Jai05]. Queueing [ETH05, GLZ00, Gup02a, HR07, Jai05, Kaw06, ODO00, PdlF07, TT03, WLW09]. Queues [TEB06, FP06, GSS04]. queuing [PdlF09, PSM02]. Quick [Ano00-78, Ano00-79, Ano00-36, Ano00-86, Ano00-91]. Quintic [ZHZT06, ANZ09, BN05, LF09, PHAF09, WH08b, ZL07a]. quorum [MM08c, TT03, ZSPZ08].

R [Ano00z, Ano00w, Ano00-56, Ano00-84, Ano03-46, Ano03-62, Ano04m, Jai05]. R74 [RW03]. Radial [HLC05, LLTL05, Sar06a, ZL04, AM08b, BSR02, Bax02, Bax06, BI02, Buh06, CLMM06, CGGC02, CCF06, DS07, DR02, DF02, DH07, Dua08, Fly06, FDWC02, FWL04, FLW06, GK02, Hon02, HHRA02, Kan04, KH00, LF03, LF05, MDPD02, PD04, PD06, PB02, SV06, Sar06b, SDZ06, TL06, VTR04, WHL02, WW08c, YTEC02]. radial-basis-like [DR02]. radially [ASHA08, BR05b]. radiation [Bat07, CS02a, HNSA09, Mak09]. radiative [Dog01]. radionuclide [VTR04]. radius [DZQ06, Guo07, JCC^{+}09, Mae00]. Raimondas [Ano01d]. Rajeev [Ano00-88]. Rajkumar [Ano00-56]. Ralf [Ano03-41]. ramp [WWH05]. Random [CCJ08, DL06, LLB05, AL08a, Ana04e, Ano04i, BDA02, Blo00, CyL08, CHM06, CJV07, DS05a, FZ07b, GPvZ06, GG06, GL03b, GC02, HM03b, KS00a, KT03a, KDKS01, Kum04, Kum05, KNP06, yL09c, LW04, Lei02, LWZ09, Lin08b, MMM04, ODO00, PL05a, RRR07, SA08, SY05, Tan04, WLW09, WI08, WH05, WD05, WGZ06, YLZG01, YLD08, YXZL09, ZL05, Ano04w, Ano04v]. Randomization [Jum01, Ano03h]. randomly [Bra08, CÁHCJL^{+}09, NSQ02]. randomness [Ano03-45]. range [CLC01, HNR08]. rank [Gal05, SC07, Wan07b, WS08b, XZ09a]. rank- [XZ09a]. rank-one [WS08b]. Ranking [CT08, CT02b, DZQ06, LH05b, AH09, CL01, Tan03a, WL08b, WL09a]. Ranks [Liu06a, Liu08e]. Rannacher [AMV08]. rapid [Ano00x, OAVB08]. Rasch [FM06]. Rate [Ana00a, Gup03, LY06b, ST07, ZG02, ZW06, BBI05, DD08, DAKO06, GPvZ06, HM03b, KT09, NTTS09, SY05, Wan02b]. Rates [BB01, Zen00, AHN06, AC00, CM03, FZ07b, HZ09a, Li01c, PL05a]. ratio [Mer05, SA08, XC00]. ratio-dependent [XC00]. Rational [Boy01, DZZT06, Kil04, Ano01t, Bal02, CK02b, DGVP06, FH03, HLX07, Ikh01, Ikh02, Ikh04, Iva07, KLP01, Li00b, LZ04b, LZ05, LLT08, TSZ02, VBG07, ZSG01]. ratios [MB08]. Rauch [Phu02]. Raúl [Ano01y]. Rayleigh [Liu08a, ZT07]. Raymond [Ano00-75, Ano03-70]. rays [AA08]. RBF [CT02a, FZ07a, GMS09]. RBFFD [SDZ06]. RCGM [LY04]. RCNNs [LC07b]. re [Ano00-55]. re-engineering [Ano00-55]. Reacting [BR06, dR03]. reaction [BKM08, CN01, DWL08, GAT07, Li02, LC06c, MC02, Pao01a, Ram05, RR08, SR03, SK06b, TZ04a, Wan02b, Wan09e, XCD07, YV05, Yil09, ZSZ06]. reaction-diffusion [BKM08, CN01, DWL08, GAT07, Li02, LC06c, MC02, RR08, SR03, SK06b, Wan09e, XCD07, Yil09, ZSZ06]. reactions [HYZ02, KW03, NB08, PL01c, XZ09b]. Reactive [SW06, LCL^{+}09]. reactor [JTRu09]. read [HLM04, LHYM09]. read/write [HLM04, LHYM09]. readers [Ano04-40, Ano04-41]. reading [Ano00-29]. Real [DD04, XZ06, ZXTL09, BD01, Bal04, CCSL04, CFY08, Dra08, Gla04, Gla09, GM00, Lai08, tLjW08, LYC07, MPTV06, SBSY06, Sun05b, Sun05c, Sya01a, Tan01, Wan05a, Wan05b, WSZ03, Wu08b, XH02, ZW09b]. Real-coded [ZXTL09]. real-life [SBSY06]. real-valued [BD01]. REALbasic [Ano00-65]. realistic [TBR01]. Realization [AF09, Don04, LW09b]. Reanalysis [Mic09]. rearrangement [Nik06]. reason [Ano03u]. reasoning [Ano01h, LCWL06a, LCWL06b, Liu09a, SFL02, WML08, hYL09, ZL06, ZL07b, Ano01t]. receiver [Hur00]. reciprocal [MPV02, OEV09, PL03]. recirculating [JS04]. recoding [IK04]. recognition [HJWL03, KT09, Lee00, PT07, WYL^{+}09]. reconstruct [WYL^{+}09]. Reconstruction [LLXC03, NT02, OIL08, OIL09, CLMM06, CM01b, JSKE05, Lee01a, LCHL06, MRÁR08, RS09]. record [Jah04]. Records [DA06]. recovery [AFI06, NUN06]. Recruitment [SHS05]. recta [MT00]. rectangle [KK02]. Rectangular [KP05b, CFY08, Hor08, Las08, ÜL02, YG08]. recurrence [FB05, GMSRP08, NR01, dG09]. recurrence-progression [GMSRP08]. recurrences [FM05]. recurrent [HHW08, LYC07, ZYZ08]. recurring [BZ00]. recursions [Jia00a]. Recursive [CÁHCJL^{+}09, BL01, Bar01b, HKA08, LZ04b, Par04a, SB00, TK02, Tan03b, Tan04, ZSG01, Zha09f]. recycling [JB05]. Red [Ano00-40, LW05]. redefined [ZW08]. Redheffer [ZS08b]. Redheffer-type [ZS08b]. Reduced [CZLN06, Tsi03, YX09b]. Reduced-Order [CZLN06]. reduced-set [YX09b]. reducible [Rom01]. Reduction [LCWL06c, LSZW07, AVA01, CTYW05, HY00, KA00a, KGWL08, LBW09, Liu00, PMR^{+}08, XMW09, YGH09]. reductions [SSD07]. reducts [TDY08]. redundancy [NFI03]. Reference [Ano00-78, Ano00-79, dSM02, Ano00-86, Ano00-36, Ano00-91, Ano01j]. refinability [YX09a]. refinable [CSX01, Xie08]. refined [FJ05, Jia01a, JZ02, WB01]. Refinement [CW06, CN01, Kr05, Zhu08a]. Refinements [DBGSP08, LF09]. reflected [DZ08]. Reflection [DLSB08, Gal08b]. Reflections [Gal06a, Blo00]. reflexive [CI06, FW08, YD08]. reforms [Ano03-29]. refractive [DDCB04]. Regarding [Cha01b, Tel08, FWL04, Gyi00]. regenerator [SWS05]. region [CSdS08, zGtM08, JZQZ07, JSGP08, LH00a, QJZ09, ZZ03a]. Regional [AS09b]. regions [LCY05]. registration [LG00]. Regnell [Ano00-28]. Regression [Ano04m, BBGG08, JK01, LU09a, SSL07, WT00, Wu03]. Regular [JKP02a, Jin09, LY08c]. Regularity [JKP02b, XT01, BKK08, FS07, JN07, Mae00, Ano03-68]. Regularization [FTC03, FFXL05, HB04, Kha01, KR07a, MRdCV00, QFZ03]. Regularized [BK00, BS00b, TG03, XL01, YB07a]. regularized-long-wave [TG03]. regularizing [BFM04]. Regularly [MM08d]. regulation [Ano03-41, YMMC03, Ano03v]. regulatory [AMYS^{+}09]. Reingold [Ano03m]. rejection [GS07b, Kar05]. rejuvenation [IDOK06]. related [Ana06, BC00, BV09, CS00a, CCS02, Che04a, DS06b, Gal08c, Gal08b, KDY09, LCHL06, LW09e, TB09, XWZ04, ZCCB08]. relation [GYPL08, GYPL09, LMHW05, LSL09, LW09e, Luo06, NR01, ÖY07b, TYW05, Yon09]. relation-based [GYPL08, GYPL09]. relational [CKY07, mLpW08, Lin08a, NT02]. relations [FB05, HLY09, IR08, SM06, dSRGRMRL08]. Relationship [TY03, CZ05a]. relationships [GYS05, HdLS03, LS06b]. Relative [DW02a, DSS00]. Relativistic [HM03a, CW09a, SM09, Ung00]. relativity [Gal09a, Gal09b, Ung05]. relaxation [Asi08, BNP05, BK04, BKR05, CS02b, Jan05, JZ08, LB04, SB03, Wan03b]. Relaxed [HL08c, Sat06, XWL06, CyL08, LCV06, yL09c, Ver03, Ver08a, Ver09, Ver12]. release [AHN06, Ano00-86]. Relevance [AK08a, CML^{+}09]. relevant [JZ03]. Reliability [IKY06, IUCR06, KYNI06, ADDL04, Elk06, GPT03, HXG03, HY00, HCY00, IY06, Lin01, Lin04a, LI00c, RVKDF09]. Reliable [HWX07, Waz01, Waz07b]. relinking [QW09]. Remainder [ARS03, BD02]. remark [Pra07]. Remarks [Dha03, PM05b, Sri03, WLS06, Zho08, ZCC04, Sun09b]. removable [ODO00]. renal [MT00]. reneging [ASESESA09]. renewal [AH03, GM09b, WLW09]. renouncement [PdlF09]. rent [CS02c]. rent-seeking [CS02c]. reorder [KH02b]. reordering [DCR07]. repair [DKO03, QNN03, SY05, WZ07a, Zha08c]. repair-time [DKO03]. repairable [KLHC08, Zha08c]. repairs [AS03a, DAKO06, LWY09]. repairs-time [DAKO06]. repeated [Ano00-51, NSQ02]. Replacement [QNN03, DAKO06, SO03, WZ07a]. replenishment [CHW07, LCT08]. replication [WL06]. replications [Che04c]. Reply [Ari09c, Ben01]. report [Ano04c, Ano04a, Ano04b]. reports [Ano00c, Ano01c, Ano02c, Ano02d, Ano03b, Ano03c, Ano03d, Ano05a, Ano05b, Ano05c, Ano06b, Ano08a, Rod00, Rod01a, Rod01b, Rod02a, Rod02b, Rod04, Rod05, Rod06]. Representation [BKN09, HMV04, Ana07b, Ano01r, Ano03-54, BV01b, Cha08, Cho03, JPS09, MWZ09, McG03, NOY01, SC07]. representations [Ano00-54, Gal08b, LL03b]. represented [JWSS04]. representing [Ben00, Cha01b]. reproduction [Fly06]. Repulsive [ZCW06, XXD06]. required [LTTH09]. rescaled [Lin04b]. rescaling [LST02]. Research [KS07, WL06, WFT^{+}09, Ano03-38, Ano04r, BBM08, CSdS08, LW09d, LLT09]. Reservoir [VDMK08]. reservoirs [BH03, YLZG01]. reset [IYN03]. residual [DD08, DH07, RNM00, VB01, XZDD09]. Residuals [Jia01a, DW02a]. residuated [SW09]. residue [ARS02]. residues [Ben00, Cha01b]. resilient [Pan03]. Resolution [LSL09, Luo06]. Resolvability [CPZ00]. Resolvent [Gup02b, AAI05, HR08, Jeo04, yL09c]. Resonance [YW09, ZH07, CO07, Du08, LDG05, SL07a]. Resonant [HM05a, HRS08]. resource [ISAAM09, PKM08]. resources [Ano03-80, EMP^{+}06]. respect [DLF09]. respiration [CH05b]. respiratory [SA02]. response [Che04c, CCL02, FS04, HZ08b, Huo05, SDCL08]. Responses [HLY06, LM03, WWL06]. Restarted [JH09]. restoration [BC00]. restorations [TY03]. restraint [AC09]. restricted [Kov05, XCH07]. restriction [Tan04]. result [BHN01, Bs03a, XY04, YXZ09]. resultants [KFMH06, Sya01a]. resulting [El 03]. results [AN09, AC03, Bal02, CH01, CCP02, CH00a, CL07b, CHM06, DS06b, Fen08, GGE09, HPG06, Jum06, KL08b, LGZ01, LZ04a, LY08c, Lóp09, LYX09, LS01a, LS03, OAVB08, OEV09, WJW08, WA01, Zho06]. resume [LWY09]. retarded [BP03, CD05c, JKP02a, LYL07, ZD03]. retirement [dKLR09]. retract [Dib01]. RETRACTED [ID05, Pen09]. Retraction [ID08]. retrial [AHRT09, AM04, GC02, KF05a, KA02, LWY09, WZ07b]. retrials [AP09]. retrieval [CS00b, CC05, HCG08, HCWC07, LLS09b, TDHL09, Ano04k]. return [Syn08]. Reverse [ZL07b, Che08b, GL08b, HNR08, WL08a]. reversible [YSP^{+}08, ZHZ00]. Review [Ano00e, Ano00j, Ano00f, Ano00k, Ano00d, Ano00i, Ano00-74, Ano00l, Ano00-75, Ano00q, Ano00s, Ano00m, Ano00r, Ano00t, Ano00u, Ano00o, Ano00p, Ano00n, Ano00z, Ano00x, Ano00y, Ano00w, Ano00v, Ano00g, Ano00-29, Ano00h, Ano00-28, Ano00-27, Ano00-76, Ano00-30, Ano00-77, Ano00-31, Ano00-32, Ano00-33, Ano00-35, Ano00-34, Ano00-36, Ano00-37, Ano00-38, Ano00-39, Ano00-40, Ano00-42, Ano00-43, Ano00-41, Ano00-44, Ano00-48, Ano00-45, Ano00-49, Ano00-47, Ano00-51, Ano00-46, Ano00-50, Ano00-52, Ano00-54, Ano00-53, Ano00-57, Ano00-55, Ano00-56, Ano00-60, Ano00-78, Ano00-61, Ano00-79, Ano00-58, Ano00-63, Ano00-62, Ano00-59, Ano00-64, Ano00-67, Ano00-66, Ano00-65, Ano00-68, Ano00-80, Ano00-71, Ano00-69, Ano00-81, Ano00-70, Ano00-85, Ano00-72, Ano00-73, Ano00-83, Ano00-84, Ano00-87, Ano00-82, Ano00-86]. Review [Ano00-88, Ano00-89, Ano00-92, Ano00-91, Ano00-93, Ano00-90, Ano01e, Ano01d, Ano01f, Ano01i, Ano01g, Ano01h, Ano01j, Ano01k, Ano01y, Ano01m, Ano01n, Ano01z, Ano01o, Ano01p, Ano01l, Ano01q, Ano01r, Ano01s, Ano01u, Ano01t, Ano01w, Ano01x, Ano01v, Ano01-27, Ano01-28, Ano03i, Ano03-76, Ano03g, Ano03f, Ano03-36, Ano03-77, Ano03j, Ano03k, Ano03h, Ano03l, Ano03t, Ano03m, Ano03n, Ano03p, Ano03u, Ano03s, Ano03r, Ano03-78, Ano03o, Ano03w, Ano03q, Ano03v, Ano03y, Ano03-27, Ano03-79, Ano03x, Ano03z, Ano03-29, Ano03-30, Ano03-28, Ano03-80, Ano03-31, Ano03-81, Ano03-39, Ano03-35, Ano03-32, Ano03-33, Ano03-38, Ano03-34, Ano03-37, Ano03-41, Ano03-42, Ano03-40, Ano03-85, Ano03-44, Ano03-82, Ano03-43, Ano03-83, Ano03-45, Ano03-46, Ano03-47, Ano03-50, Ano03-49, Ano03-48]. Review [Ano03-51, Ano03-52, Ano03-53, Ano03-84, Ano03-86, Ano03-56, Ano03-55, Ano03-54, Ano03-57, Ano03-58, Ano03-60, Ano03-61, Ano03-63, Ano03-62, Ano03-59, Ano03-64, Ano03-87, Ano03-65, Ano03-73, Ano03-72, Ano03-66, Ano03-70, Ano03-67, Ano03-68, Ano03e, Ano03-69, Ano03-71, Ano03-74, Ano03-75, Ano03-89, Ano03-88, Ano03-90, Ano04e, Ano04d, Ano04f, Ano04g, Ano04h, Ano04j, Ano04i, Ano04k, Ano04n, Ano04l, Ano04m, Ano04-27, Ano04o, Ano04p, Ano04r, Ano04q, Ano04-28, Ano04s, Ano04u, Ano04t, Ano04w, Ano04v, Ano04x, Ano04y, Ano04z, Ano04-30, Ano04-31, Ano04-29, Ano04-32, Mic08]. Revised [Ano00-59, Ano03-45, Ano01w, WL08b, WW09b]. revisited [PB01]. revolution [Ano00-64]. Revolutions [Ano00-66]. reward [KNP06]. rework [LCT08]. rezoning [HLP03]. RFDEs [Far01]. Rham [DMVR00]. Rheinboldt [Ano03-67]. rheology [GMŚ07]. Rhoades [IK03]. Rhoades-type [IK03]. rhythm [BM00]. Riccati [BUP09b, GLC09, Iva07, Lin08c, WXM09]. rice [CGPB08]. Richard [Ano03-77, Ano03-33]. Richardson [VAR03]. Ricker [SZJ09]. riderless [Yav06, Yav07b]. Ridgway [Ano03-84]. Riemann [CD03, CD09, Cha08, Jum06]. Riesz [BGMB02]. Right [KK08b, And03, BW05, BW06c, CL01, DFFL09, Gil00, LDK08]. right-focal [And03]. Right-indefinite [KK08b]. rigid [PM08, Sta07, Ano03-65]. ring [HaLC08]. rings [Coo05]. Rio [Ano03-87]. Risebro [Ano03-37]. Risk [Hua08c, AVA01, Cap07, CC08a, LP06, LW09d, LT09b, SCF^{+}09, Tsa03, XLL08]. risk-time-cost [LW09d]. River [Ano03-59, Ano04r]. rivers [BH03]. Rivest [EMD08]. RK [HR05]. RNS [AS05, AS07, ARS03, TS00b]. road [RGTT03]. Rob [Ano00-90]. Robbins [Ano00-86]. Robert [Ano00-38, Ano01e, Ano01n, Ano01z, Ano03-76, Ano03-30, Ano03-59]. Robichaux [Ano00-44, Ano00-45]. Robin [Ben01, DL05]. Robinson [Ano00-63]. Robo [Ano01u]. robot [Ano05g, FY03a]. robots [Ano00-82, Ano03y, Ano03-55]. Robust [Che08a, Han04, MBI08, XT08, YX07, ZLCM08, ZL09b, BGA08, Bla04, Dra06, GHC09, HJWL03, HYZ06, PCE^{+}05, PG08, Qua01, WW09b, YL04c]. rockfill [SBPD08]. rod [FG00, Yav05d]. Rodin [Ano04c]. Rodney [Ano00l, Ano00m, Ano04d]. rods [JSK^{+}04, JSKE05]. Roe [BG01]. Roger [Ano00p, Ano03-70]. Rogers [CBGVPP09]. Rojas [Ano01y]. Role [LV03, Pad06, WKL06]. rolled [RDS00]. rollin [Yav05a]. rolling [Ano05g, FY03a, KY00, Yav05b, Yav05d, Yav07a, Yav07b].

ROM [Ano00-34, Ano00-39, Ano00-40, Ano00-46, Ano00-83, Ano01f, Ano01k, Ano03v, Ano03-34, Ano03-58, Ano03-67, Ano03e, Ano03-75, Ano03-88, Ano00-78, Ano00-79]. Ronald [Ano00z, Ano03-51]. roofs [KOV08]. Roos [Ano03-68]. root [BGP04, KI04, Pan05a, Pan05b, PR06, Ikh08]. root-finders [Pan05b]. root-finding [BGP04, Pan05a, PR06]. roots [CjBN09, JCZ09, MPTV06, NJ08, Nie01b, Win03, WXS01, YC06, Zha08a, ZW09b]. Ross [Ano01n, Ano04h]. Rota [HHS07]. rotated [GKM08]. Rotating [Hay06, AL08b, Ari09b, HM05a, KL01, LM03, MH09, RZ03, Tur07]. rotating-disk [Tur07]. rotation [SB03]. rotational [Kan02]. rotationally [Leu00]. rotations [AK08a, Rub07]. rotor [Dem00, YW09, Yav05b]. rotor-spun [YW09]. rotors [KY00]. Rough [CW06, Guo03, HY09, HS04b, LCWL06b, LCWL06c, LSZW07, PC08, QLD08a, SW09, TDY08, XZ08b, XMW09, YGH09]. rounding [MT09]. routing [WW09c]. RPICM [LLTL05]. RS [Lin04b]. RSV [AMC08]. RTDs [LKS^{+}06]. Rudd [Ano00-58]. Rudolph [Ano03-55]. rule [JSS01, KS07, LZ05, LGGSP06, MC07, QLD08a, TXC02, Uje07, ZY07]. Rules [Bel06, AC02, AC05, An03, BCD08, DY09, ED09, HRMS09, HLY09, Kov00, LCY05, NR01, Nie09, TWS00, WX03a]. Runeson [Ano00-28]. Runge [AG03, BJ02, CFR03, CLM03, CM02, FZ07a, GSZ02, GW03a, GPPRM01, GPMPR03, LGY09, LYL07, PF01, XT02, XT04a, Xia07]. Running [Ano00-67, Ano00-68]. runs [SS09d]. Russell [Ano03k].

S [Ano00-75, Ano00-71, Ano01n, Ano03q, Ano03-39, Ano03-82, Ano03-45, Ano03-53, FP06, KP01, JC05, Ano03-73]. S-G-L-KKM [JC05]. S-Plus [Ano03-73].

Sabin [CL08b]. Sabra [Ano03-81]. Saddle [Ano03-59, Ano04r, Li05]. Safer [EMD08]. sale [LZS07]. Salehi [Ano04w, Ano04v]. same [Ikh08]. sample [GG06, WZ08b]. sampled [SN03]. samples [EW00b, HS00, Ter00, ZM94]. sampling [CWW08, CZ02, GL03b, KDKS01]. sandwich [CB02, PB01]. sapiens [Ano01u]. SAR [LWZ09]. Sara [Ano01e]. Sarit [Ano01n]. satellite [LW09f, VAR03]. satisfaction [RKS08]. satisfiability [Bru08]. satisficing [MM05a]. satisfying [Nie01a]. Satsuma [LHT07]. saturating [GLTS08, TLS^{+}08, Wan06c]. saturation [KLP09]. Scalar [HRMS09, Bil05, DSS08, OS08a, WX00a, dR03]. scale [ACV03, AS09b, CW05, FHZ08, FY04, HY03, LCHL06, SI03, Sun05a, TLS07, XWW08, YX09a]. scaled [JAM02]. scales [ADP08, AK08b, AW09, BW05, BW06c, BG07b, CL07b, CWZZ07, EPS05, FZG07, Hof03, HT05, JEP08, KÖ08b, Li09a, LN09, OEV09, ÖY09, PXGL09, SSX09, SL07b, Sun09a, Top04, WW08b, WWW08, Wan08a, WS05, WZ07c, Yas08, ZS05, ZLZ09, Zho08]. Scaling [Bax06, CSLC07, Koi07]. scanning [SYHC06]. Scattered [Sun00, All02, WB01]. scatterer [YZ02a]. Scattering [Dor00, KKM00, AC04, ISK00, Kru03a, LL07, Yan02, Yan04a, Yan04b, Yan05a, ZRS05]. Schaffer [Ano00-80, Ano00-81]. Schauder [PPRG02, WY07]. schedule [ODO03]. schedules [KA02]. Scheduling [MR06, MS09, DAKO06, He00b, HY08, KM03b, KY07a, KY07b, LCWL06a, LCWL06b, LCWL06c, MCC09, MG08b, MYR09, SMO03, SMOK05, TH02, Wan08b, WWW^{+}09, Wan09b, XW05, YE01, YH11, YZWY08, ZY02, ZYY08]. schema [ROY01]. scheme [AB08c, Ano00-69, BG01, BKR05, CCC08, CS00b, CLT^{+}08, Che03, CCK00, DS04, Dub08, EA03a, Gos00, HJL05, JKN09, Jua03, KK09, KZ02, KL03, Lee00, LHW04, LKC04, LXZ08, LT09b, MQ05, MWZ09, Mic03, MRÁR08, OM09, PKSP08, RL08, Rub07, SALK02, SDDC08, Soh05, STK08, Tia05, Waz07a, WLT07, YXZL09, ZW09a]. schemes [AO05, AY06, Asi08, AO09, BS00a, BGM^{+}02, CFR03, CS02b, CAZAA00, CGGC02, CCZK09, GW03a, Gla00, Gla03, Gla04, Gla05a, Gla05b, Gla06, Gla07a, Gla08, Gla09, Gra03, Ikh01, Ikh02, Ikh04, IL04, IJB00, KO09a, Li01b, LDHW05, LY08d, LHLD08, LB04, NUN06, Nko00, PKSP08, Roe09, SMMM02, SC02, Tan01, VRJ09, VTR04, XX09b, Zha03b, ZC08]. Scheutz [Ano03t]. Schiestel [RZ03]. Schmidt [GLR05]. Schneiderman [Ano03-50]. Scholes [AE08, CNPP08]. Schönemann [Ano03e]. Schreiber [Ano00-38]. schrödinger [HJW05, LDHW05, AS08, CK08, CQT02, DS07, Den07, GT00, HZZ02, IJB00, JG04, Kan04, KD08, ÖY07a, RL08, Sim01, SAB07, Tan07, WX09b, ZXlD09]. Schumann [BM00]. Schur [SJJ09]. Schur-convexity [SJJ09]. Schur-geometrically [SJJ09]. Schurer [AG09]. Schwarz [JZ08, LZZ04, Sir00, Ano03-66]. Schwichtenberg [Ano00q, Ano00r]. Science [Ano00u, Ano00-85, Ano04w, Ano00j, Ano00k, Ano00-84, Ano03u, Ano03x, Ano03-81, Ano03-85, Ano04v, KL09, LKL08, SA08, Ano00-69, Ano03q, Ano04t, Ano01v]. sciences [Ano00-80, Ano00-81, BDGR01, TB08]. scientific [Ano03-30, Ano04p, Ano03-67]. scientists [Ano00g, Ano00h]. scope [KPD^{+}09]. Scott [Ano03-84]. screening [Can03]. screens [KKM00]. SDOF [MB08]. Sealed [Bel06, Che08b]. sealed-bid [Che08b]. Search [CX04, MIPK04, Ano00-70, Lei02, MCRÁ05, Nak06, Nam01a, yNyLjZaZ08, OTB^{+}08, SS07a, SHWF09, Wan02d, WS08b, WYSS09]. searches [SH01]. Searching [BDT04, Qua01, YZLL09]. seasonal [YC07b]. Sebastapol [Ano00-75, Ano00-27, Ano00-34, Ano00-57, Ano00-62, Ano00-87, Ano01j, Ano01k, Ano01-27]. Sebastopol [Ano00v, Ano00-36, Ano00-37, Ano00-39, Ano00-40, Ano00-44, Ano00-45, Ano00-52, Ano00-53, Ano00-60, Ano00-78, Ano00-61, Ano00-79, Ano00-63, Ano00-67, Ano00-65, Ano00-68, Ano00-72, Ano00-73, Ano00-82, Ano00-86, Ano00-92, Ano00-91, Ano00-93, Ano01f, Ano01s]. secant [HR02]. secession [Mar09]. Second [Ano00p, Ano00-51, Ano00-46, Ano00-69, Ano00-89, Ano03h, Ano03p, Ano03-39, Ano03-85, Ano03-45, Ano03-84, Ano03-56, Ano03-59, Ano04e, Ano04f, Ano04i, Ano04n, Ano04m, Ano04p, Ano04z, GZ06, LTY05, Luo03, PXW09, SW02b, TK09, Top04, ATT03b, AL04, ASHA08, AS01, AJ09a, AJ09b, AW09, AY06, AC03, AT02, Ayk04, BG04a, BLD07, BA01, BPA03, BPM05, BN00, BBRS08, COE03, CC09a, ÇT05, CHS02, CQJ03, CS04b, CL08a, CZ09b, CM02, DJ08, DK05, DCS09, DHM05, Dog08, DP07, DR01b, EM09c, Gal08b, GF06, HASA07, HNSA09, HW09, HT02, HM05b, HLX07, HLY05, Jan09, JEP08, Kar03, KK04, KT03a, KLW09, LA00, LZ04a, LS08, Lin04b, Liu04b, LX07, Liu08b, LXK09, LYG05, LN07, Lu07, Luo02, LD02, Ma00, MQ05, Man00, Man01, MP08a, NF03, PLG07, PG00, Pen02]. second [QSL08, RSD09, RN02, Shi03, SW02a, SOW04, Sun05a, SS06, SZ09b, Tan02b, TRG01a, TRG01b, TÇ04, TA04, TBG05, Tsi02, TDS08, Van07, VSKF08, Wan00e, WWZ03, Wan04a, Wan04b, WZ07b, WLW07, WWW08, WSL08, WX09a, WWC08, WWX04, WD05, WW08d, XZ08a, Xu08b, Yan04c, YYZ03, YW03a, Yan03c, Yao05, Yen08, ZZ00a, ZP01, ZCZ01, ZS05, ZJ08, Zha09d, ZWY08, Zha08d, ZC09b]. second-grade [ASHA08, MP08a, VSKF08]. Second-Order [SW02b, GZ06, LTY05, Luo03, PXW09, Top04, ATT03b, AL04, AS01, AW09, AY06, AC03, AT02, Ayk04, BG04a, BLD07, BA01, BPA03, BPM05, BN00, BBRS08, COE03, CC09a, ÇT05, CHS02, CQJ03, CS04b, CL08a, CM02, DJ08, DK05, DHM05, DR01b, GF06, HT02, HM05b, HLY05, JEP08, KK04, KLW09, LA00, LZ04a, Liu04b, Liu08b, LYG05, Lu07, Luo02, LD02, Ma00, MQ05, Man00, Man01, NF03, PLG07, PG00, Pen02, Shi03, SW02a, SOW04, Sun05a, SS06, SZ09b, Tan02b, TRG01a, TRG01b, TÇ04, TBG05, Tsi02, TDS08, Van07, Wan00e, WWZ03, Wan04b, WLW07, WWW08, WSL08, WWX04, WD05, WW08d, XZ08a, Yan04c, YYZ03, Yan03c, Yao05, ZZ00a, ZP01, ZCZ01, ZS05, Zha09d, ZWY08, Zha08d]. secondary [KCD05]. secret [SC02]. section [Has03, KN08, MG03, RV01]. sections [CA09, Lan04b, RDS00]. Security [ZW08, CH05a]. sedimentation [BKR05, RQH^{+}08]. Seeing [Ung03]. seeking [CS02c]. Segalman [HM05a]. segment [ZW09a]. segmentation [CVW00, LWZ09, MWZ09, PLP09]. Seidel [LCC08, Sha09]. Seifert [Ano03o]. SEIQR [PLG^{+}09]. SEIR [LGZ01]. Sejnowski [Ano00-54]. Selected [Ano04x]. Selecting [BL08]. Selection [SYJJ09, BFFL08, BFKL08, DP02, Hua08c, LZS07, OTB^{+}08, SHS05, WT00, WTZY05]. selections [ZM04b]. selectivity [LG09b]. selectors [BA08]. Self [HHJS03, Ano03-40, Ano04l, BN05, Dua09, HL02b, HLW07, Han08, HHR07, HLP05, HLC00, HL02c, HLM04, Hua05, HLCW07, LW02, LHYM09, LYG05, MNIS06, PL01b, RR09a, Shi03, Sta07, TCH07, VJ04, WXM09, YHS08, YH03, ZT08, ZL09c, Ano03-68]. self-adaptation [ZL09c]. self-adaptive [HL02b, HLW07, Han08, LW02, YHS08]. self-adjoint [BN05, HHR07, LYG05, Shi03, WXM09]. self-adjusting [HLP05]. self-affinity [Ano03-40]. self-connections [YH03]. self-normalized [ZT08]. self-organizing [PL01b]. self-propelled [Sta07].

Self-regularity [Ano03-68]. self-similar [Dua09]. Self-stabilizing [HHJS03, HLC00, HL02c, HLM04, Hua05, HLCW07]. self-testing [MNIS06]. semantic [Ano03-71, SP09]. semantics [LLS09b]. Semi [Gir06, BI02, CDJ04, CCZK09, GGE09, GZZ09, JPV01, JPV02, ÖY07a, PWX08, WW09b, YH04, ZYW08]. semi-analytical [GGE09]. semi-definite [YH04]. Semi-Implicit [Gir06, CCZK09]. semi-infinite [CDJ04, JPV01, JPV02, PWX08]. semi-inverse [ÖY07a]. semi-Lagrangian [BI02]. semi-local [WW09b]. semi-positone [ZYW08]. semi-supervised [GZZ09]. semicontinuity [BW06b, SLJ04]. semicontinuous [Wu01]. semicycle [LZ05]. Semicycles [ZZ01b, ZZ01c]. semidiscrete [Xu02a]. semidiscretization [AU03]. semigroups [IW06, XL01]. Semilinear [LX04, BA01, BM04a, BAK09, CK02a, CO07, DR04, EMM09, HPT05, JKP02a, LYW04, PW05, SC04, SC00a, Sir00, Wan02b, WW03]. Semilocal [HR02]. semipositone [AGO03a, AW09, MZ09, XXD06]. semirings [FJZ08]. Semitensor [Che06]. Sen’ [NLL08]. sense [Ano04-32, MGG05]. senses [PCA09]. sensing [ASM09, JCC^{+}09, MM08c, ZSPZ08]. sensitive [HX09, HD08]. Sensitivity [GC06, GMMB06, GMMB09, Mou02, Noo02a, PL05b, BLS05, CSdS08, VG08]. sensor [ASM09, JCCT09, JCC^{+}09, LLS^{+}09c, WNS09]. Separable [Wan00a, AG03, Mur05b]. separate [HLM04]. separated [Waz05a]. septic [WL09b]. sequence [AA09c, Gu07, Gup03, HKA08, LP00, Moo02, Sav07b, Sav08a, SA08, VA04, WWW^{+}09, ZSG01]. Sequences [AEÇ06, Bar01b, BCT00, BIKL06, BFKL08, BZ00, LZ04b, LBZ^{+}09, MCC09, May00, Nie01a, SL09, Sun09b, TB09, Tay01, Wan00a, WD07b, YT00, ZWC^{+}09, Zha09f, ZL07c]. sequencing [Bru08, KM03b]. Sequential [CT09, BCD08, CFH05]. serial [JKY06]. Series [SK04, BBGG08, Bar01d, Boy07, BJ00, CKCG03, CGBJR03, DK09, Has02, IR08, Jum06, KHS00, LC07a, LCL08, LHP07, Mal04, MJD08, MM05b, PS09, Sav07a, Sav08b, SL09, Sri03, TYW05, WZ07a, ZYG^{+}09, Ano00-30]. server [Ano00-92, Ano00-93, GC02, LWY09, ODO00, SG08, Wan04a, WX09a, Ano00-44, Ano00-45]. servers [Jai05]. service [Ano03w, GM09b, HY08, KK06c, KLHC08, RS06, SG08, Wan04a, WZ07b, WX09a]. services [Bel01, KCD05, Ano00-56]. sessions [Guy00]. Sestoft [Ano03-47]. Set [BFL06, CW06, EGM^{+}02, GJM03, Yan08, AO01b, ARS02, AFL^{+}09, ARM08, Ano00-88, Ano00-89, Ano03-51, BJK^{+}09, BSM01, CGY08b, EMP^{+}06, FW08, GK04a, Gup02b, HNR08, HLCL00, HY09, HRMS09, HCLJ00, Hua01, HF04, HLCW07, IDM08, Jeo04, JZQZ07, bJQxC09, JLP09, JY11, Li07, LCWL06b, LCWL06c, LSZW07, MBR03, MKF09, MCC09, MN01, Nam00, Nam01b, PdlF07, QLD08a, SV01, SM02, SHWF09, VDMK08, WW08a, Xu03b, YLY^{+}09, YGH09, YX09b, YLZ08b, ZGY05b, ZGY07a]. Set-Based [BFL06, CW06]. Set-contractive [GJM03]. set-theoretic [SV01]. set-valued [ARM08, CGY08b, FW08, HRMS09, HCLJ00, Hua01, HF04, Jeo04, Li07, MN01, YLZ08b, ZGY05b, ZGY07a]. Sets [PT00, CTYW05, Din04a, Dra03b, Eke02, HH00, HHJS03, HCTC03, ISAAM09, KGWL08, Lan04b, sLhYL09, MRB02, POS03, PC08, PG06, PG08, RFCCRM03, SW09, SLJ04, TB09, TDY08, WS07, Xu03a, XX09a, YL09b, YZWL09, hYLL09, Zhu08b]. setting [YLD08]. settings [GPvZ06]. settlements [SBPD08]. setup [WWW^{+}09]. Seventh [Ano04r, DL08]. Several [Liu08d, ZY08a, Ano00u, CH04, CS04a, JSS01, JA03, KKM00, LW08a, Sta07, XT04b, ZZ01b, ZZ02a]. SFR [Ano03-37, Ano03-82, Ano03-48, Ano03-69, Ano04j, Ano04i, Ano04o, Ano04p, Ano04v, Ano04z, Ano00q, Ano00r, Ano00-48, Ano00-49, Ano00-47, Ano00-46, Ano01i, Ano01g, Ano01w, Ano03w]. Shadbolt [Ano00-38]. shadowing [Gu08]. shallow [BFO02, CB02, Gir03, Gla00, Gla05a, Gla07b, Hav08]. Shane [Ano03-69]. Shannon [CW06, Dra03a]. Shape [CS02a, CLMM06, AG01c, AG01b, FW04b, FZ07a, Lee00, WKL06]. shaped [BKN09, WY07]. shapes [Bri00]. Shapiro [Ano01r]. share [CZLC08]. sharing [Ban07, BD07, CZ05b, Fan02b, LL06b, LHLD08, LYX09, SC02, Xu03b, Xu08a, ZCL08, ZL08b]. Sharkey [Ano03v]. Sharp [Hu08, KLJ09, KDC08, Uje04b, ZZ09c, Lü08]. shear [BCGP03, NS07, SASHS09, VSKF08]. shear-thinning [SASHS09]. sheared [Sar03]. shearing [ZMM07]. shed [KOV08]. sheet [ASHA08, AL08a, Ari07a, Ari07b, Ari09a, Ari09c, Bat07, EM09b, FLfFL08, XL09]. sheets [APvZ01, BPvZ03]. Sheffer [Nie01a]. shell [Ano01-27, CB02]. shells [GAA06]. Sheng [Ano00u]. Shepard [AG01c]. Shi [Ano01w]. shift [Ana04b, JZ02, PLP09]. shift-and-invert [JZ02]. shift-invariant [Ana04b]. Shiftability [SKHK06, Wan00c]. shifted [JH09]. Shih [Ano04d]. Shih-Chii [Ano04d]. Shin.The [Ano03-83]. Shing [Ano03p]. Shock [BD04b, Aro08, BDO04, LYF09, SO03, SD03]. shooting [EGA03, Ha01]. Shore [Ano00-90]. short [XZ09b]. shortages [CHW07]. Shortest [FKM05, HL02c, Hua05, KC05b, PS07a]. shrinking [FLfFL08]. SI [DLB03]. sided [ZZZ02]. sidewall [TRD08]. sideways [FQZ02, FFXL05, QFZ03]. Sierpinski [Jia09]. Siever [Ano00-78, Ano00-79]. sight [KKS^{+}04a]. sigmoidal [CC09b]. Sign [AS07, TA06, AS03b, AOW05, AOTW07, ARS03, Bal00, COE03, DCS09, DG05, Kar08, LS09, LLJ09b, Liu04b, SSX09, WA01, WS05, Won08, Yan06a, YW08a]. sign-altering [Bal00]. sign-changing [Yan06a, YW08a]. signal [Ano03-52, CFM02, JWSS04]. signals [AC00, CÁHCJL^{+}09, CWW08, QC06, WDLP04]. signature [YXZL09]. signatures [NFI03]. signless [SS09c]. Signorini [Ryo00a]. Silberschatz [Ano00i]. Silvia [Ano01i]. Similar [WYW05, Dua09]. Similarities [Fer03]. similarity [BCGP03, CC08a, WML08, YL09b]. Simon [Ano00-80, Ano00-81]. Simple [CYH05, AGS09, CGFS09, CLT^{+}08, Gra08, HLW07, HL08b, Kel06, LLNS05, LHS08, Nik06, WLS06]. simplex [Pan00, PP01a, WWWL08]. simplical [CZ02]. simplicial [ZZ02b]. simplified [LLL08a, Rub07, Ryo00a]. simplifies [ARS02]. Simply [Ano00-69]. Simpson [Lü08, Uje04b, Uje07]. Simson [Ano00j, Ano00k, Ano00v]. simulate [BHS^{+}08, JTRu09]. Simulated [HHL^{+}06, Tan03b, ZYY08]. Simulating [AZB08]. Simulation [Fer03, ITR08, MPW07, Ano03-82, Ano03-86, BH03, BW09, Che04c, CLT^{+}08, DKS07, DMP03, DE06a, GOMA07, HK08, ITM08, KK06b, Kor00b, Kor07, Lun03, ME00, Mar09, Mos08, NB08, OKCO08, Qia08, RR03a, SWS05, SPW08, SP08c, SGR^{+}09, SM08, Sür08, VM09, WL04, WB09, WLD09, XZDD09, XCS^{+}08, YB07b, Zha03b, ZZJ^{+}04, dR03, Ano04g]. simulations [AB09, Ban00, BC08, CLL09, DGG^{+}04, HW05a, HWHG08, IMLF09, JZC06, KIXF09, KD08, LY09c, MG08c, NS07, dSM02]. simulator [AMYS^{+}09]. Simultaneous [CFY08, CM01b, LW09e, SDK09, Bal04, Ikh08, KI04, Nie01b, PV00]. Simultaneously [Pan05c, WSZ03]. Sinc [CWW08, EGZ04, RR03a, MEG08, MS04b]. sinc-Galerkin [MS04b]. sine [KGB07, Lu09b, Waz05c]. sine-cosine [KGB07, Waz05c]. sine-Gordon [Lu09b]. Single [Kan04, KY07a, KY07b, RQH^{+}08, WZE^{+}06, Wan08b, WWW^{+}09, Wan09b, YH11, CS08, Che01a, DCR07, GC02, GSNDB^{+}08, Gup02a, KH02c, KM03b, LHS08, MCC09, MM06, WW04a, ZWCW00, ZJLW06, Zhu05]. single-level [KM03b]. Single-machine [KY07a, KY07b, Wan08b, YH11]. single-period [DCR07, KH02c]. Single-Phase [WZE^{+}06]. single-server [GC02]. single-species [MM06]. single-step [Zhu05]. Singular [ZCW06, AS03b, AOLL04, AOS07, Ana05, AA09a, Ano03e, AK03, BN05, BBRS08, CCP02, CM09, CM01a, CK02a, CZ09b, CZ08, CZ09c, DM08, EGA03, GK02, GSZ02, HZ08a, Hil01, Hua08b, IS04, Jia00b, Jia01b, Jia01c, Jia02, JX04, KP03, KA09a, KK04, KDKS01, KR01, KLW09, KB08, LT05, LFZ00, Li01a, LC03, LUZ05, LXH05, LJL06, LSZ08, LQ04, Liu04c, LY05b, LLW07, MSS08, NVAR03, PSS09, Phu02, QZ05, RV05, SC03, Sta02, SPW08, SS06, WYW05, Xia07, XXD06, Yan03c, Yan06a, Yao04, ZYW08, ZJ08, ZQC08]. singular-perturbation [KA09a]. singularities [AORS03, Gan00, Kel06, KHS00, Li01d, LCGX06]. Singularity [MAR05b, AC02, RGTT03]. singularly [AE07, CN01, KP01, KK09, Li00a, Li02, LTWM08, SR03, SR04, Sir00, SK06b, SNA09, VR02, YMMC03]. sinh [Waz05a]. sinh-Gordon [Waz05a]. sink [Bat07]. sintering [Nik06]. sinusoidal [AZB08]. SIR [NTTS09]. Sirospun [ZH07]. SIS [HZ05, WM04]. site [MB08]. sites [MD06, RMC08]. Sivashinsky [LRL06, KT08]. Six [ZS08b, DS04, LPSS02]. six-point [DS04]. sixth [NMD08, WX00b]. sixth-order [NMD08, WX00b]. size [AVAGP08, DFFL09, KH02b, MGG05, MSN^{+}00, WM09, YHS08]. size-reorder [KH02b]. sizing [HTKM09, HY06, XD02]. skew [LHL07, tLjW08]. skew-antisymmetric [tLjW08]. skew-Hermitian [LHL07]. skills [HCTC03]. Skjellum [Ano00-89]. Skorokhod [Ano04w, Ano04v]. Slavko [Ano00-83]. sleep [ODO03]. slender [DH08a]. slide [FG00]. sliding [PM05a, ZYG^{+}09]. slip [Ari07a, HLMR08, KHA07, Lia04, Lia05, SMH08, SANH08]. slopes [DPU07, PT00]. slowly [SD03]. Small [Blo08, GG06, Ban07, BRC00, FY04, Gil00, MM07a, MP02b, Wan07b]. small-scale [FY04]. smallest [Hua08b, Li01a]. smart [Bog09]. Smirnov [Ano03-70]. Smith [Ano03l]. Smooth [ADH^{+}03, VRJ09, AA05, Ana05, BBCE09, Hua00, ITT07, LK03, ZW04]. smoothed [MG08a]. smoothing [BMH08, SMW04]. smoothness [Ana05, Gan00, HL00, Lin04c]. Smythies [Ano03-79]. Sobolev [Ana08, BC02, BC03, BCGA05, DBX04, EP04, Kan09, KF05b, KL05, Pat05, PPA08]. sociable [Ano03y]. social [Lo 02]. society [Ano00j, Ano00k, STV08]. Soft [FJZ08, Jun08, JLZ09, JLP09, MBR03, Yan08, AFL^{+}09, AA09b, CTYW05, FGX05, JY11, KGWL08, MRB02, YLY^{+}09, ZS07]. software [Ano00-28, Ano03w, GPT03, HXG03, IY06, TY03, TBR01]. Solar [Ano03-70]. Solaris [Ano00-86]. solenoidal [KK06b]. solid [Ano03g, Kor07, LLCS02, Mar00, MO08, RB06, VC03]. solid-water [VC03]. solids [MKSB08, PM08, Ste07, SR07]. Solitary [DH08a, WH07, JG07, MP07b, WX09b, Zha07a]. soliton [DS05a, Ma09a, OR01, Tao09, Wan09a, cWcX09, cW09b, YZ02b]. soliton-like [YZ02b]. solitonic [HO00, TGH02]. solitons [MP07b, Waz05d]. Solla [Ano01e]. Solomon [Blo00]. solutal [ZDLH08]. solute [ZS07]. Solution [BDGR01, WZE^{+}06, Yil08, YTEC02, ACD01, Ada04, AOLL04, AS01, ASESESA09, AM08b, AE08, Ari09b, BMHV09, BP09, BSR02, BM01, BM02, BM04a, BBL05, BD03, BHEKE07, CL07a, CCJ08, CDJR03, CJO05, Cat05, CP04, CH05b, CWZ03, CDJ04, CNPP08, CJV07, Cse06, DI08, DK09, Das09, DRR09, De 04, DSHS05, DXZ05, ETAJ09, FLfFL08, Fas02, FR04, FP02, FS07, GHHD08, GNR05, GLZ00, HR07, HMCR00, HBP06, HL09a, HC01, HHL08, HL08b, Jan07, JG07, JDH03, JC00, JSF00, JPV01, JP01, JPV02, KP03, KF05a, Kar03, KS03b, KÖ09b, KA00b, LP08a, LF03, LRL06, LL05, LWY02, LWY04, LY05b, MH02, MPS04, MN01, MT09, Mur07a, NOY01, NSB08, NF03, NST01, OR01, Par04a, PS08a, PW05, PJJP04, PB02, QS03, RF09, SMH08, SD07, SD08, SK04, Sid08]. solution [Sim01, Slo09, SZ09a, Trz01, UHuI09, Wan05b, WLW07, Wen00, Wen02, WSZ02, WL08c, XZ06, Xie00, Xu02a, Yan03f, YZLL09, Yao05, Yao07, ZRS05, ZCM01a, Zha01a, Zha07a, ZS08a, ZYW08, ZXlD09, Zha09c, ZLD08]. Solutions [AO08a, BW06c, GY06, Gyi00, Hay06, HLY06, Liu06a, MCH01, ME08, PS03b, SW02b, TEB06, UK08, Wan06b, WG06, WHD09, ZLS06, ZLJW06, Zha08d, ZCW06, AETEZ07a, AB08a, AH00, AZ01, AML02, AGO03b, AS03b, AOW05, AOTW07, AOS07, AB08b, AJ09a, AJ09b, ACJ07, ACM05, AA01, AAD03, AW09, ACH01, Ayk04, Bae07, BKM08, BF03, BG04a, BX07, BLD07, Bai08, BPA03, BSK04, BV05, BAK09, Ban00, BM04b, BPP04, BCRS05, BNH08, BN00, BHNO04b, BGO05, BRC00, BW05, BR05b, Boy01, COEP00, Cab01, COE03, CYG05, CYG06, CG08, CH04, CM04, CDM01, CDM03, CWY08, CGY08a, CGY08b, CX04, CWZZ07, CWLC07, CC04b, CH00c, CW05, CK02c, CP01b, CHR09, CZ08, Cui09, CI06, DCW08, DJL04, DS04, Den07, DCS09, DN03, Dib03, DSS08, Din00c, DW02a, DL05, DY05, DYZ08, Dix05, DM08]. solutions [Dog08, DG05, Don09, ES01, ESR06, EMM09, ERRY01, EA03b, ER07, FW00, Fan02a, FW08, FL05, FW04a, FZG07, FMFV08, FH03, Fro08, GK02, GL09a, GGA09, GSN08, GS08, Gon09, GC08, Gou01, GH03, Guo01, HG03, HZ08a, HW09, HT02, HY03, HM05b, HYLW00, HO00, HZ08b, HZ09b, Huo04, HLY05, Ibr09, IP07, JKN09, JI04, Jan04a, Jan04c, Jan07, Jan09, JKP02a, JKP02b, Jia00b, Jia01b, Jia01c, Jia02, JA03, JX04, JZ03, KN04, KC02, KK08a, KHN09, KT08, KK05, KKZ02, KLW09, LYW04, Lai08, LG08, LLWY00, Leu00, LFZ00, Li01b, LX03, LA03, LZL04, LS04b, Li04b, LZ04a, LW05, LDZW05, LZJ06, LTWM08, tLW08, mLpW08, Li08b, LFZ08, LY08c, tLjW08, LS09, LLJ09b, LXH05, LLT08, LLNS05, LJL06, Lin06a, LSZ08, Liu02, LK03, LWG04, Liu04a, LG04a, LQ04, LG04b, Liu04b, Liu04c, LY06a, LH06, Liu07a, LLW07]. solutions [LDK08, Liu08a, LL08b, Liu08e, LT09b, LHT07, LS01b, MAH08, Ma00, Ma01, Ma04, MDG05, Ma07, Ma09a, MZ09, MRT04, MH07, MZ04, MC02, MCRÁ05, MRMRM02, Mus06, MO01, NLR00, Nic04, NRL00, OI00, Ou08, OSSY04, Pal01, PPRG02, PLG07, Pao01a, Pao03, PF01, PZ07, Pen09, PS01, PP04, Pit01, Pra07, RGTT01, RC07, RRM08, Ryo00b, Ryo00a, RN02, RSK02, Ryo07, SSH09, SG07, SSX09, SA05, SAS06, SHY08, SK04, SC03, SW07, SDB09, SW02a, SL07a, SKG01, Son09, Sta02, SPW08, SK06b, Sun05a, SZ05, SX06, SL07b, Sun09a, SK09, Tan02a, Tan07, TLLM08, TMG03, TGS07, TDS08, Van07, VNFF08, WJX04, WZ00, WC04, Wan05a, WYW05, Wan06c, WLS06, WG07a, WG07b, WW08b, WWW08, WLWC08, Wan08a, WYL08a, WG08a, Wan09e, WX09b, mWmZ09, WFY09, Waz00, Waz01, Waz05a, Waz05d, Wei05, WA08, WZ09, Won00, WA01]. solutions [WX03b, WS05, WCW08, WWC08, Won08, WWYL09, WLZ07, WH07, WLW08b, WW08d, Wu09, WLD09, cWcX09, cW09b, XH02, XXD06, XLZ02, XDD09, YZ02b, Yan03a, Yan03e, Yan03c, Yan04d, Yan05b, Yan06a, Yan07, YW07, YW08a, lYsT09, Yao04, Yas08, YZZC09, YTCF08, YWYL07, YD08, YXZ09, YB08, Zen09, ZYG08, ZWCW00, ZCZ01, ZZ01b, ZL03, ZW04, ZWW04, ZL04, ZFYS05, ZJLW06, ZLH07, ZL08d, ZJ08, ZTW09, ZLZ09, ZG09a, Zha09b, Zha09d, ZG09b, ZC09a, ZM02, ZYCI09, ZH02, ZZ03b, ZZH03, ZHKW05, ZT07, ZHL04, dSOGRMRL04, Ano03f]. solutionsof [Ryo04]. Solvability [BS01a, Du08, LDG05, Liu06b, PGT08, Sta07, yL09c, LXK09, Sah01, Ver01a]. solve [Ano03-58, Chu07, EA03a, KLZZ09, LHW04, PSS09, Pet04, ZZD08]. solved [EBVS06, GS06, Slo07]. solvent [ITM08]. solver [DS02, LHC08, NJ08, Pet01, SALK02, Wac08]. solvers [BMH08, CjBN09, NSB08, Pet05, Sch01]. Solving [AR07, Ano00-58, BGLT08, BS03b, Bru08, Che08b, GMS09, HTKM09, HF00, HL09b, Ryo03, SPE06, Tao09, WYH00, Wan02c, WTZY05, Ara01, BUP09a, BNH07, BG08, BA09b, BG11, BXFZ09, BPSW00, CC09a, CJ05, CN01, CZ09b, DH08b, EGZ04, GAT07, GOMA07, GJ07, Gra03, GJS01, Han06, Han08, Has04, Hem08, HHRA02, Iva07, IJB00, JDY08, JLW04, KP01, KGB07, KK06a, KZ02, KX05, KH09, yL08a, LW09b, LZZ04, Lin08c, LHH08, Lu07, LHWW08, MYG09, Mar05a, MAK02, MAK03, MLW09, MQ06, MEA09, NMD08, NW09, PMR^{+}08, PSS07, PKSP08, PDRR08, RR09b, SDVS09, SD09, SNT08, SNG09, SERG09, Sal08, Sha04, SP08a, SF07, SZ09b, SNA09, TD09, WWWL08, WHS08, Waz07c, Waz07b, WKL06, WF01, Xu07b, XC04b, YHS08, YL09a, Yil09, YLD09, Yus09, ZGY07b, ZAY08, Zha01d]. Some [AW01, AK07, Ano04-40, Ano04-41, AHS00, ASS08, AAR03, BL01, Bal02, BM05a, BDA02, BD04a, Blo00, Cap00, CS00a, Che06, CN08a, DMP01, DSS00, FMFV08, FWL04, GYS05, GN09, HM05b, HL00, Hor01, HdLS03, IN07, Jia00a, Kar09, Kel06, KHN09, LWZ08, LZ04a, LY08c, Li09a, Liu07b, LYX09, LS06b, MZD09, MP08b, NW09, PKSP08, PS09, RR09b, SL00a, SDB09, SV08, SP08b, SKG01, Sun09b, TOH03, WJW08, WYL08a, WJS09, Whi00, Wu08c, XC07, YHT04, YL07, ZX08, Zha08a, Zho06, ZL05, Zhu09a, dSRGRMRL08, AETEZ07a, AH00, AP08, ABCD05, AFL^{+}09, Ana00a, AK08c, BG04a, BLD07, BRC00, BR05b, Bou00, CGFS09, CL02a, CGGC02, Che01c, DBEGB08, EM09a, FW04a, LF03, LLNS05, LST02, Ma07, MH07, MRS08, Mat03, Nie09, RSK02, SERG09, SS09c, Sri03, SK09, Tan00, TK02, Tho01, Wan00c]. some [WLS06, Wen08, XLZ02, Yan03c, Yan06b, ZL04, Zha09b, Zha08d, ZYCI09, dSOGRMRL04]. Song [Ano03-56]. Sonin [CD05b]. Sons [Ano00i, Ano00z, Ano00-90, Ano03-31, Ano03-73, Ano03-74]. soot [HK08]. Soulie [Bla04]. sound [JZC06, WDLP04]. Source [IS06, YM04b, Ano01l, Bat07, BOP07, CCJ08, CM01b, DDX04, GL09c, Gos00, LX03, PLD02, SA08, SDDC08, WDLP08, YM04a, Zhu09b]. source/sink [Bat07]. sourcebook [Ano04-29]. Sources [ZW06, Che04e, MYR09, RB06, Ano04-32]. Spaccamela [Ano00o]. Space [LT01, LPvST07, MDPD02, XWL06, BV05, BGMB02, BGG04, BCA09, CS01, CKY04, CPC00, CAZ01, CL08b, CM01b, DLSB08, Din00a, Din00b, DW02a, DW05, Ehr02, Gor00, JZ08, JC05, LYW04, Li00b, LYL00a, LCYL00, LCL00, LYL00b, LS08, LK04, LW03, LQ04, LHC05, Mar07, MCZ00, Pat05, PPA08, PS03b, Roo08, UA06, WL07a, WHS08, ZC09a, Ano00-70]. Space-decomposition [LT01]. space-densifying [MCZ00]. space-filling [LW03]. space-frequency [JZ08]. Space-time [LPvST07, MDPD02]. spaces [AO01b, AO01a, ATT03a, AZCK05, AAI05, Arg00b, AP01, Bai01, BSD00, BA01, BPM05, BGMB02, BP03, BHN01, CGY09b, CGY09a, CWY02, CL08a, CHZ09, CKQ08, CyL08, DMVR00, Dha06, DP08, Din00c, DP02, Din02, DPJ03, DY05, DYZ08, EP04, FW08, FHT05, FjH07, udK07, GJM03, GGCD05, Guo01, Guo02, GJS01, fHCH09, HPT05, HCLJ00, Hua00, Hua01, HS09, Inc09, IL04, Jeo04, Jin07, Jin08, KDC08, KK01, Köh00, LT05, LCV06, yL08b, LW09a, LKS08a, LCYL00, Li04a, LK03, LWG04, LHH08, OS09, PG04, PPRG02, Pan07c, PXW09, PW06, Rad09, RR09a, SC04, Sav08a, SW08, TCH07, WLW07, Wen08, Wu00, WTY03, qWjH07, WZ08b, XH07, YL07, ZAAESG09, ZYG08, ZC09b, ZCK05, dSOGRMRL04, dSRGRMRL08]. Spain [JSGP08]. Spainhour [Ano00-78, Ano00-79]. sparse [ACD01, HRP04, HCL07, LZ09, MAK02, MAK03]. spatial [APvZ01, CS00b, CY00, NT02]. spatially [Col04, FZ07a, NB08]. spatiotemporal [TCX08]. SPEA2 [WFT^{+}09]. Special [FHZ08, Gal09b, LCGX06, RGTT01, TB08, Ano04-40, Ano04-41, Boy01, DGR04, Gal09a, Ung05, WL08a, WYL08a]. specially [JSKE05, XZ09a]. Species [Wan06a, FW00, HC09, LY08b, MM06, TL00, XC00, ZWCW00, Ano01u]. specimen [FM08a]. Spector [Ano01f]. spectra [Hil01, ZLY09]. Spectral [Ano00-94, GL03b, You04, AK06, BSS98, Boy01, Gir03, Guo03, GW06, Guo07, Hai03, Has04, KT08, LTY05, Mae00, RF09, Ser00, TK02]. spectral- [AK06]. Spectrum [KAKK00, Das04, EGR03, FBA07, HJWL03, HHR07, IN07, RSR01, RSE01, SS09c, Sot01, Sot02]. Speech [KT09, PT07]. Speed [CCP09, Gal08b, Gal09b, TGIM02]. spelling [PT07]. Sperm [DFY06]. Sperner [Pal01]. SPH [SL08]. sphere [OAVB08]. spheres [HD08]. spherical [DGG^{+}04, KKS04b, MK08, OS08a, SH01, TV07, WL04]. spinning [YW09, ZH07, Ano03-71]. spiral [Lee01b]. Spline [GJS01, BN05, CC09a, CL08b, CWB01, He04, HM00, KP01, KP03, KA09a, LW09a, Li01c, LHC05, WL07a]. Splines [McD07, AS01, CM01a, DSHS05, HL00, TDS08, WFLC05]. split [ME03]. split-step [ME03]. splitting [AK01, BNP05, BSR02, CCS08, CN08b, DZ08, FT01, FGH08, GW03a, Gos00, LHL07, Noo00b]. splitting-radial [BSR02]. splitting-step [DZ08]. splittings [CFH05, Hav08, Min08, SH06]. spot [YC07b]. Spreading [MM07b, Qia08]. spring [RM07a, RM07b]. Springer [Ano00q, Ano00r, Ano00o, Ano00-48, Ano00-49, Ano00-46, Ano00-58, Ano00-70, Ano01i, Ano01g, Ano01w, Ano03i, Ano03h, Ano03u, Ano03r, Ano03-30, Ano03-34, Ano03-37, Ano03-42, Ano03-40, Ano03-44, Ano03-82, Ano03-45, Ano03-46, Ano03-48, Ano03-51, Ano03-52, Ano03-84, Ano03-56, Ano03-64, Ano03-65, Ano03-67, Ano03e, Ano03-69, Ano03-75, Ano03-88, Ano04f, Ano04j, Ano04i, Ano04o, Ano04p, Ano04w, Ano04v, Ano04y, Ano04z]. Springer-Verlag [Ano00q, Ano00r, Ano00o, Ano00-48, Ano00-49, Ano00-46, Ano00-58, Ano00-70, Ano03-37, Ano03-82, Ano03-48, Ano03e, Ano03-69]. SPSS [Ano03-72]. spun [YW09]. Spurgeon [Ano00-27]. Spurious [HLS05]. SQP [JT05, SH02, SY09]. SQSD [SH01]. Square [EPS05, BB04, BM05b, CJV07, HHW08, SDZ06, Zha03b, ZZH03]. square-based [SDZ06]. squares [Bed02, BG04b, CJL04, Che04d, DW02a, DD08, FR04, KN04, KL06, KU01, LL05, Liu08e, LU09a, MS04a, Mur05a, XZDD09, Xie00, Yan03d]. squares/penalty [CLS07]. squaring [CWLC09, Koi07, LLY01]. Srivastava [PMS07]. SSC [KY02]. St [Ano04c, Ano00v, XC07]. Stability [AR09, AR10, AO05, Bar01d, BY03, BJ02, BP03, BX02, BDO04, BMM^{+}06, Che03, CN04, Dem00, GY06, HT05, HC02, Hua08a, JL09a, KÖ08b, yL08b, LRL06, Lev06, Li05, LC02, Lóp09, LYL07, NTTS09, NC09, Noo02b, PP09, Ses03, SYL05, TCX08, WLS01, WZ07c, WH09, Yen08, YS04, Zha01c, ZL02, ZSKZ03, Zho08, ZC08, ASÇY08, Ash05, AY06, AC09, BV01a, BD04b, CB02, EIR03, GZ08, GHW04, HR07, Han04, HLX07, HHW08, HZ05, JK09, Jin08, KZD03, Lee01b, LGZ01, LZ05, tLW08, Li08a, LB01, LKC04, LC07b, MM07a, MG08c, Mur05b, NTHP09, PS01, Pit03, PD06, Pru09, RL08, SP08b, SWL05, Sun06a, Sun06b, TZ04b, Tia08, Wan01, WZ06, WWL06, WF07, WZZ09, Wen00, WH05, WD05, WGZ06, WG08b, WL08c, XC00, XLL03, XLL04, XLT09a, XLT09b, YX07, YTC08, YH09, YG08, ZD01]. stability [Zha02b, Zha03a, ZA03a, ZD03, Zha04b, Zha04a, ZS04, ZLH07, ZL09b, ZSZ06, ZZ05]. Stabilization [AM08a, KY00, NP07, Yav05d, GX04, LHYM09, Par04b, PKP09, Tur07, XT08]. Stabilized [GPMPR03, Ban00, Bla07, DPU07, RPO03]. stabilizing [HHJS03, HLC00, HL02c, HLM04, Hua05, HLCW07]. Stable [FW04b, Mur07a, ZW03b, ACJ07, CZ06, Din04a, DS05b, JSF00, LML02, Mur06, Net05, Net07, PYLX08, PDRR08, Sar03, WX00b]. Stackelberg [TOH03]. stage [CCW00, SMW04, WW04a, WFTL09, XXLD04, XW05, XCD07, ZWW04]. stage-structure [CCW00]. stage-structured [ZWW04]. stages [Tsi03]. stagnation [INP08]. Stan [Ano00-46]. standard [AMC08, MG08a]. Standish [Ano03k]. Stanford [Ano04q, Ano04-32]. Stanley [BKN09]. Stanly [Ano03-51]. starlike [Ano00u, EAA09, KM01, Oro02, SM00]. start [KY07a, SMO03]. start-time [KY07a]. start-up [SMO03]. Starting [CLM03, HR05, GPMPR03, TP03, WZ07b]. State [ACLL06, WWHL09, AHN06, Ano03g, Ban00, CS01, CCP09, EZ08, HPG06, Kar01, KT02, Kar05, KLP09, KMK09, LP08a, MBI08, Pen09, PM01, SDK09, SL07a, TLS^{+}08, ZZ00b, Ano00-70]. state-delay [MBI08]. state-dependent [HPG06, Pen09]. state-feedback [KLP09]. State-space [Ano00-70]. stated [HM04]. states [HW05c, LYL00a, WW02b]. static [DD00, HLP03, Ji02]. station [KLHC08]. Stationary [Wan06a, Ano00-30, BPvZ03, CMS04, GL03b, KAKK00, ST07]. stations [LWLJ09]. Statistical [AEÇ06, AD08, DAKO06, DA06, DE06b, MJ02, AAK08, AF02, PLP09, Sav08a, Win03, ZZ02b, Ano03-73]. statistics [Ano04t, Jah04, Lin04b]. Steady [ACLL06, ITT07, Ban00, Kon09, LP08a, ST07, WB09]. steady-state [Ban00]. Steeb [Ano01w]. steel [RDS00]. steepest [SH01, ZCK00]. steerable [FY04]. steering [HH09]. Stefan [Ano00-41, Ano03-88, GS06, Slo07, SZ09a]. Steffensen [AB05]. stem [AACLM08]. stenosed [ICM09]. stenosis [WB09]. Step [TS06, AB05, BJ02, CAZAA00, CZG04, CM02, DZ08, FT05, FM05, HN07, Ikh01, Ikh02, Ikh04, Min08, ME03, PW06, RR09b, Tou05, Tsi02, WBSP08, WX00a, WX00b, YHS08, ZGY07b, Zhu05]. step-size [YHS08]. Stephen [Ano00-78, Ano00-79, Ano03n, Ano04y, Ano04z]. stepsize [JMR00, Xia07]. stereoregular [WK03]. Steven [Ano00-83, Ano01q, Ano03s, Ano03-62]. Stevenson [Ano04-32]. Stewart [HSM00]. Stieltjes [BS01a, BM05a, BCDS09, CD03, CCD07, CD09]. stiff [ACD01, AP08, DKS07, HCLF01, WX00b, Xia07]. stiffness [RDS00, SN01]. stimulus [Iac07]. Stirling [MIPK04]. Stochastic [AL08a, NB08, AAK08, AC06, BV05, BAK09, BBM08, CS01, CZ00, CM03, DZ08, ETAJ09, EBVS06, FKM05, Gyi00, HHW08, IUCR06, Isk04, Iva07, JJ01, JI04, KL03, KW03, KB08, LLL08a, MS08, PKM08, PS07a, RS06, Sot02, VDMK08, WZ06, WD08, XX09a, YYNK03]. stock [KS00a, ZWCW00]. stocks [KWT02]. Stoer [Ano03-46]. Stokes [AVO08, AL08b, BCS03, BG04b, CLS07, FP02, GC06, GCN06, ITT07, JY08, KK02, KLS06, Lia04, Man05, NP07, PR07, Tou05, VNFF08, WFC04b, YTEC02]. stopping [BGS04, YYNK03]. storage [CFR03, CS00b, WL06]. strain [GMS08, Ste07]. strapdown [AF09]. strategies [Bel05b, CN01, CML^{+}09, CZ02, zGtM08, LLCS02, Tan00]. strategy [Han08, HHR07, IDM08, JZQZ07, LZ09, SY05, YZLL09]. stratification [Sar03]. stratified [SSD07, VNG04]. stream [DS02, Kan02, ZYG^{+}09]. stream-wise [Kan02]. streamline [Zha09e]. streams [LGH09]. strength [KFMH06]. strengthened [HZ03]. stress [CMS04, Kor07, VSKF08]. stretching [ASHA08, Ari07a, Ari07b, Ari09a, Ari09c, Bat07, EM09b, XL09]. strict [JT05, MCQK09]. Strictly [MOP06, Dha07, Inc09, OS09]. string [FG00]. strings [SS09d]. strip [KK05]. Strong [BC02, Gir03, Inc09, MCQK09, PW06, ZP02, CZG04, ER07, FS07, JZT07, MRMRM02, OI00, Son08]. Strongly [AEÇ06, PXXW06, Sat06, BUP09b, CJO05, DYZ08, He03, HH03, HK07, HGH02, IP07, JZT07, JC00, JSF00, LK03, MH07, MEA09, PL05b, TG03, Wan03a, ZGY05b, Zha09b, ZL07c]. Strother [Ano01h]. Structural [BV01a, LY09b, Lo 02, ODO03, TL06, TV07, Ano03-74]. Structure [Ano03-42, TYS03, AT05, Arg00b, BM00, CCW00, DY09, Jay04, MAK02, Pan03, Qiu08, Ren03, TDHL09, WW04a, WSO08, XCD07]. Structure-preserving [Ano03-42]. structured [JSKE05, KK06c, LGZ01, Li09b, NST01, Pan07a, ZWW04]. Structures [PKM^{+}06, BDO04, BD04b, GYPL08, HGS07, LWZ08, LDK^{+}06, LD08, PLL03, SDK01, She00, She02, Waz04, Waz05b, YJCZ08].

Stuart [Ano00-84, Ano01r]. Stuck [Ano00z]. Studies [Ano03l, DW02b, Gre09, XZDD09, Ano01h]. Study [BA09c, DD00, GGE09, LJ08a, LLJ09a, Lin01, OAVB08, YG08, Ano03-80, BUP09a, BT03, BK07, CKL06, CPS05, Dim08, Gal01, HRP01, HL02a, HK08, HPT05, HXG03, KPD^{+}09, KHA07, KW03, LF03, LKS^{+}06, MAR05b, MG01, MP02b, MD06, MP07b, MM08c, MK08, SC04, SB03, SD03, Sür08]. studying [Ikh01, Ikh02, Ikh04, LWY06]. Sturm [AU09, BCT00, CH00c, DH03, KM07, KK08b, LR03, LS09, Ma07, UA06, VJ04]. Sturzenegger [Ano03z]. style [HZ09a]. sub [CZ09a, JZT07]. sub-feasible [JZT07]. sub-superalgebras [CZ09a]. subalgebras [Jun09]. subclass [AM08c]. Subclasses [AD09, AIS08, Aou08, ASS08, CK06, WJS09]. subcritical [DMP03]. subdifferential [AL08b]. Subdivision [CZ02]. subgrid [ACV03]. Subharmonic [DCS09]. Subject [DS06a, AH03, ASM09, HLMR08, Lin06b, LWY09, SD08, VSKF08]. sublinear [LS09, Man00]. submerged [Dor00, HW05a]. Suboptimal [OLAS01, YL04c]. subordination [NON^{+}08, Pie09]. subpatches [Las08]. subpolygroups [DCLF09]. Subrahmanian [Ano01n]. subsampling [DH07]. subset [YL04a, Zal02, Zal03]. subsets [IL04, LdSF00, MOP06]. subsoil [Ros03]. subsonic [BGR09]. subspace [CL02a, MPZ08, Tou05, WYSS09]. subspaces [LBW09, LHC05]. substitution [IK04, IK05]. substitution-based [IK04]. subsystems [Mar09]. Successive [Yao05, Yao07, Zha09d, BNP05]. suction [MM07b, Tur07]. Sufficiency [GA06]. Sufficient [Hor08, YO06, mLpW08, Sot01, TL08, WLZ07, Zal03]. sufficiently [AO01b]. Sugeno [OF08]. suggested [LP00]. Suitable [Ste06]. suite [AF09, ANV00]. sum [CCS00, Dha03, NR01, WY01]. sum-form [WY01]. summability [Sav07a, Sav07b, Sav08b, SL09]. summarization [Ano00e, Ano00f]. Summation [AO08b, YI06, Åhl02, Agr08, HHS06, HHY09, PMQR09]. summation-type [Agr08]. sums [POS03, RSR00, WW08a, ZT08]. Sun [Ano03-83]. Sun-Joo [Ano03-83]. super [CL08b, JZT07, LX07, NF03, Net05, Net07, ÖZ07]. super-half-linear [ÖZ07]. super-implicit [NF03, Net05, Net07]. super-linear [LX07]. super-memory [JZT07]. superalgebras [CZ09a]. Supercomputing [Ano00-71]. Superconvergence [Li01d, LL02a]. Superlinear [NV08, ZCW06, AK08c, CZ08, Jia00b, Jia01b, Jia01c, Jia02, Kur08, Lin05, SC03, XXD06]. superlinearly [Yan03f]. superposition [DS02]. supersaturated [bFFlT09]. supersplines [LF09]. Supervised [ZZ09b, GZZ09]. supply [KS00a, LTYB09, MIPT08, Saw03]. Support [SSL07, BCGA05, Cha08, SN02]. supported [WW08c, YTEC02]. Supporting [Åhl03]. suppression [ABS09]. Surface [DZZT06, GC06, GCN06, BGR09, CLMM06, Gou02, INP08, ID10, MM07b, Qua01, SYHC06]. surfaces [AZB08, BI01, JW05, Las08, LL02a, WW08c, Xu04, Ano03-51]. survey [ABCD05]. survivable [BR05a]. survival [LW05, PC08]. Susanne [Ano03-84, Ano03-88]. suspension [SS09e]. suspensions [HH09, SP08c]. sustaining [LW08b]. SVM [WYL^{+}09]. Swarm [MIPT08, DTXW05, HTKM09, KLZZ09, KH09, LWLJ09, LCL^{+}09, MLW09, NLLL09, XXC^{+}09, YZLL09, ZZZ09]. Swift [Pol09]. Switched [Mah09, XT08, ZLCM08]. Switching [LET06, EZ08, KL03, SCF^{+}09, YG08, ZL09b]. Sylvester [WW08a]. symbiotic [Zha02c].

Symbol [Ish00]. Symbolic [IK04, TG03, YZZC09, YB08, AH00, Ano03-34, Fan02a, HHS06, IK05, Tun05, XLZ02, YZ02b]. SymbolicC [Ano01w]. Symbolization [HHS07]. Symmetric [LWZ04, tLjW08, WX03a, Blo08, BR05b, CWZ03, DH08b, DL08, Has04, HL08c, KC05a, Leu00, LW02, MG08a, Net05, PB02, RSR01, RR03b, TG03, WX03b, WS08b, Yao04, YFT04, YLL08, YL08c, ZCK05]. symmetrical [PSES07, XZ08a]. symmetries [Åhl03, Boy07, MRÁR08]. symmetrization [Pan07a]. symmetry [Ish00, OS08a, Öze08, Ano04y, Ano04z]. Symplectic [CQT02, XT02, Ely04, HLS05, LDHW05, XT04a]. symposium [Ano00-85]. Synchronization [MM08c, ET09, HaLC08, LC06c, TaL09, XCS^{+}08]. synthesis [MS03, ZL06]. System [Ano00-86, CR06, Hua06, OK06, ZW06, ADDL04, AOW05, AOTW07, AN09, AS01, ABS09, Ano00i, Ano01s, Ano03-86, ANZ09, BCGA08, BPM09, BG11, Bou03, CC09a, CO00, CCC08, Che04e, CS04a, CK02b, CM06, DWL07, DH08c, Dog05, DLM08, DL08, DTZF03, DDX04, DXZ05, ES08a, ES08b, FHT05, GL09b, GMS09, GW03b, GY09, HKL03, HC09, fHCH09, HYZ02, HF00, HZ08b, HL02c, HL04, HY06, Huo04, Hur00, IYN03, Ish00, Jai05, Jin07, Jin08, KS00a, KC02, KF05a, KK06c, Kaw06, KB04, KLHC08, KKS^{+}04a, KHN09, KYNI06, KC03b, KB08, KS07, LRLM^{+}01, yL08b, LRL06, LML02, LDK^{+}06, LY08b, LG09a, Lin04a, LP04, LHYM09, LY06a, LCWL06c, LNL08, LLT09, LPSS02, LHH08, LL03c, LHT07, Lu07, Ma09a, MM05a, ME00, McG03, MG08b, MW00, NLL07, NLL08, NQHN03, yNyLjZaZ08, NW09, ODO00, PP09, PC08, PZ07]. system [Pet01, PM01, RS09, SNT08, SO03, SN03, SC00b, SSD07, TT03, TZ04a, Ver01b, Wan02b, Wan05b, WZ07a, WK08, Wan09e, WA01, WHL02, WS05, Won08, WSZ02, WH08b, XC06, XC00, YCCY02, Yan03b, YW08a, Yus09, ZSG04, ZGY07b, ZM04b, ZWW04, ZYEZ07, ZL07a, Zha08c, ZLZ09, ZQ00, Zha02c, ZLN06, Ano00-58]. system-application [KC03b]. systematic [Wan09a]. Systems [LZL06, ZHZT06, AH03, ACD01, AML02, AT05, AAK08, AF09, All05, ADP08, AR07, Ano00-55, Ano00-56, Ano01e, Ano01x, Ano03-75, AK08c, AG03, Ara01, AO08a, AP09, Ash05, AK03, BUP09b, Bai00, BSD00, BA01, BPM05, BK04, BNH08, BBH07, BA09a, BA09c, BFM04, Blo08, BPSW00, BPW00, CJ05, CLM03, CCS08, CJO05, CF04, Cha01a, CH05a, CZ03a, CZ05a, CHL09, CWZ03, CE02, CKR03, CK04, CPS05, CP01b, DKS07, DWL08, DCS09, DGPT08, Dib01, Dib03, DD08, EIR03, Ely04, FY03b, Gin02, GL06, Gos00, Gu08, GLTS08, GK03, Han04, HD09, Hau06, HBP06, HW05b, HPT05, HCL07, HC01, HC02, HL08c, HL03b, Huo05, IP07, IUCR06, Ish00, Jia00a, JT07a, JT07b, JL09b, Jin09, JH09, JSF00, Jum01, KT02, Kar05, KYJ03, KAKK00, KK05, KLP09, Kov00, Lai08]. systems [Li01b, LL03b, LS04b, LDZW05, LHL07, tLjW08, LBW09, LC02, Liu04a, LC04, LLW07, LY09b, LL02b, LLXC03, LHR09, MBI08, Mah09, MP02a, MPS04, MPW07, MUHdH03, MS02, MCS08, MM06, MAK02, MAK03, MLW09, MB08, MQ06, MIPT08, MCRÁ05, NS00, Nes01, NTHP09, Noo02b, NV08, ODO03, PL01a, PG04, PCE^{+}05, PLD02, Par04b, PL01c, PP06, PJJP04, QZ05, QLD08b, Raz00, SDK09, SC04, Sal08, Sha09, Sid08, Son09, Sun05a, Sun05b, Sun05c, Sun06a, Sun06b, SBD^{+}00, SNA09, TLS07, TD09, TL00, TK07, TCX08, TLS^{+}08, ÜÇT08, WW04b, WFC04a, Wan05a, WL06, WF07, WYL08b, WD08, WXM09, WWS^{+}06, Waz07c, Wen00, WH05, WD05, WW08d, WL09b, XH02, XHZ05, XLZ02, XLL04, XT08, Yan03e, Yan05b, YV05, YY07, YX07, YJCZ08, YMMC03, Yu04, YZZ^{+}08, YL04c, ZCS04, ZHZ00, Zha01b, Zha01c, ZD03, ZHL05, ZL06]. systems [ZLCM08, Zha08a, ZTH09, ZL09b, ZYH01, Zho03, Zla01, ZBS03, Ano00n, Ano01n, Ano03-54, Ano01x]. Systolic [BD01, LLY01, MSN^{+}00, SMS^{+}04]. Szász [Gal08a]. Szego [CBGVPP09, GM09a].

T [Ano00e, Ano00f, Ano00-71, Ano03p, Ano03-53, LMSH04]. T-cells [LMSH04]. T. [LHR09]. tables [QLD08a]. tabu [OTB^{+}08]. Tabular [Ano03-75]. tabulations [Ano03m]. Taha [Ano04r]. Tail [EZ08, ABS09]. Taiwan [KS07, Tsa03]. Takahashi [Kam09]. Takashi [Ano03-69]. Takens [TZ04a]. Talib [Ano03-75]. Tamraparni [Ano03-31]. Tan [Ano01w]. tangles [Dar08]. tanh [DCW08, Waz05a, Waz05c]. tanh-function [DCW08]. Tanigs [Ano03-68]. tapered [MP07b]. target [LLS^{+}09c, SK06a, WNS09, WWHL09]. targets [Bog09]. tariff [SSLL09]. Tarkan [NLL08]. task [LLS^{+}09c]. tau [Par04a, Rap01, DS06a, Kha05]. Taylor [Ano00-71, Ana07b, BBL05, HR08, Jum06, MM05b]. TCOM [DY09]. TCP [Ano00-72, Ano00-73]. TCP/IP [Ano00-72, Ano00-73]. Team [Ano01k]. technia [PR01]. Technique [WZE^{+}06, AETEZ07a, BHS^{+}08, BA09b, CT02a, Deh02, Dib03, He03, HH04b, JYDH02, KK09, MM05a, MCH01, MZ07, PdlF09, SR04, SY09, TLP^{+}09, WDLP08, Zha02b, Ano01m]. Techniques [Bel06, BRC00, CJR09, CO00, GJS01, IS04, Kr05, LRLM^{+}01, Li01d, MRdCV00, Pan07a, Sch01, WW03, WXM09, ZW03a]. technological [Fer03]. technologies [TBR01, Ano03-50]. technology [Ano00-55, Ano01v, Ano03-82, CL00, ES08b, KWT02, NLL07, NLL08, PL01a, WL06, Ano00-66]. telecommunication [AP09, IDOK06]. telecommunications [Ano03v]. telegraph [ACCMJ07]. Temperature [DS06a, LYF09, BG01, Kon09]. temperatures [Les02]. Temporal [AU03, MM07a, ROY01]. tensile [SL08]. tension [MM07b]. tensioned [GX04]. tensor [Åhl03]. tensors [CHS02]. Terlaky [Ano03-68]. term [AN05, BOP07, CCJ08, CM01b, IJB00, LFZ00, Lin06a, LL08b, NR01, SDDC08, WLW08b, YM04b, ZP01, ZCZ01]. Terminal [HCY00, HY00]. Terminal-pair [HCY00, HY00]. termination [Csa01]. terms [BUP09b, BD02, EGM^{+}02, Gos00, HS00, HM04, HZ03, Lin07, LC06c, PLD02, lYsT09, YM04a, ZS04, ZSZ06, ZZ01c, ZHKW05]. Terrence [Ano00-54]. tessellations [DW02b, DW05]. test [NWL08, Pan01, WX00a]. testing [LBZ^{+}09, MNIS06, NT02, TSY09]. Tests [KL08a, ME00, PM08, Sür08, TK02]. tetrahedra [Pla07]. tetrahedral [BKK08, Kr05]. text [Ano00e, Ano00f, Ano04q, SP09]. Texts [Ano03-46]. texture [MWZ09]. TFT [KS00b]. TFT/LCD [KS00b]. th [Ano00-85, Dra08, GM09c, CZ08]. th-order [GM09c, CZ08]. Thad [Ano03o]. Thai [PT07]. Thakur [Ano00-88]. Their [Bel06, Ano00o, BL08, BS01a, Cap00, CSdS08, CFH05, DJL04, FGH08, HHY09, IR08, ISAAM09, JC05, Köh00, LS04a, LWZ04, LD08, sLhYL09, LTSW02, NT02, SS09c, THD08, VV08, Wac00, WD07b, XX09b, ZLY09]. them [Ano03-58]. Theodore [Ano00-74, Ano03-31]. Theodoridis [Ano00d]. Theorem [ARS03, AK03, BM09, CHY01, CWW08, DP02, GL08b, Kam09, Lan04a, LH05a, LCWL06a, Liu09c, LYX09, PKT07, RR09a, Sav07b, SA05, SL06b, WW09b, Wen08, YXZ09]. theorems [AO01a, AZCK05, AT02, CC04a, Cap00, CCSL04, CZG04, ĆRRU08, CZ09c, Dha03, Dha04, Dha06, DP08, Din02, FjH07, FT03, GN09, HH03, HL07, IK03, Inc09, IN07, JEP08, JC05, KH08, Lan04b, LF02, LHP07, LWG04, OI00, Pen02, PW06, SH06, SP08b, Son08, SOW04, TRG01a, TRG01b, TA04, Wan00d, Wan09c, Wu00, Xie01, XG07, Zha08b, ZL08b]. theoretic [LGBS00, MRP08, SV01]. Theoretical [Ano03-65, CJ04, LF05, Ano03-78, FM08c, MM00, MM08c, PC08]. theories [hYLL09, Zha08a]. Theory [Ano00y, Ano00-29, Ano01g, CW06, Hua06, Nam01a, SBS06, Wan00b, Yan08, dM06, AO01b, AFL^{+}09, Ana06, Ano00-85, Ano03-49, Ano03-84, Ano03-56, Ano03-58, Ano04n, BC08, BFKL08, CP02, CGL01, CD05c, Dha07, Fei07, Gla05b, HD09, HHS07, Ind02, IL04, JLP09, Kor05, Kru03a, LYW04, LYL00a, LCYL00, LCL00, LZJ06, LW09d, Lin08c, LSZW07, MBR03, Mas07, MP07b, Nam00, Nam01b, PdlF07, Pei02, QLD08a, RMC08, Ser00, Ste07, WJX04, WF07, WY07, YLZG01, ZhYL05, ZJLW06, Zho07, Ano00-83]. theory-probabilistic [Ind02]. there [Dra03b]. Thermal [DHY06, LYF09, MS00, ES08b, HNSA09, LHLW07, Mak09, NLL07, NLL08]. Thermo [RB06, RB07]. Thermo-elastic [RB06]. thermo-elasticity [RB06, RB07]. thermocapillary [DS08]. thermoelastic [DLSB08, MK08]. thermoelasticity [SB03]. thick [MO08]. thickener [BKR05]. thickness [ICM09, PM05b]. thin [BT03, LWY06, LM03, MM07b, SANH08, Ste07, SZL02]. thin-plate [Ste07, SZL02]. thinning [SASHS09]. Third [Ano03r, Ano03-46, Ano04j, HJWL03, AG01a, AA01, And03, Ano00-67, Ano00-68, Ano00-84, Ano00-86, CH01, CjBN09, DK04, Fen08, HAKA07, HY03, KKZ02, LS04b, LDK08, SMH08, SHI09, VR02, WG07a, YW07, Yao07]. Third-order [HJWL03, AG01a, AA01, And03, CH01, CjBN09, DK04, Fen08, HY03, KKZ02, LS04b, VR02, YW07, Yao07]. Thirty [Ano00j, Ano00k, Ano03-70]. Thirty-five [Ano00j, Ano00k]. Thomas [Ano01n, Ano01x]. Three [ATT03b, AP01, CL08a, FT05, Kor00b, Lee01a, LL03c, Ras06, SN02, TS06, WZE^{+}06, Wan06a, WG08a, WX03b, hYLL09, ZGY07b, ZM04a, ZYY08, AC05, AN09, And03, AK08b, Ari07b, Ari09c, BG04a, Bai08, BMH08, CD05a, Che03, CZG04, Chu07, DW05, EM09b, ER07, FY01, GY09, HG03, HM00, KO09a, LYW04, LY09a, LLS^{+}09a, LLJ09b, sLhYL09, Liu02, Liu04b, Liu04c, LLW07, LL02b, LYX09, Ma00, Mar07, Mar05a, MP08b, NR01, PW06, Roe09, SS07b, SYHC06, TGS07, TT09, Won08, Yao05, Yas08, Yav00, YF04, YWZ03, ZJ08, Zha08d]. Three-Dimensional [WZE^{+}06, Kor00b, SN02, Ari07b, Ari09c, BMH08, DW05, EM09b, GY09, KO09a, LY09a, sLhYL09, LL02b, LL03c, Mar05a, SYHC06]. three-directional [HM00]. Three-Level [Ras06]. three-link [FY01, Yav00, YF04]. three-party [LLS^{+}09a]. Three-point [ATT03b, CL08a, ZM04a, AN09, And03, AK08b, Bai08, Che03, LLJ09b, Liu02, Liu04b, Liu04c, LLW07, Ma00, Mar07, TGS07, Won08, Yao05, Yas08, ZJ08, Zha08d]. Three-Species [Wan06a]. Three-Step [TS06, FT05, ZGY07b, CZG04, PW06]. three-term [NR01]. Three-view [Lee01a]. threes [Ano00-90]. Threshold [HZZ02, LGZ01]. thresholds [ZZ09c]. thrombus [bFFlT09]. throttling [WLT07]. throttling-based [WLT07]. Tidal [SK06c]. tight [ADH^{+}03]. Tikhonov [MRdCV00]. Tikhonov-based [MRdCV00]. tilings [AT05, Gor00]. Tim [Ano00-62, Ano00-91]. Time [BFL06, Gir06, GY06, Kar05, LL06a, LZL06, LST02, Mur08b, OK06, Ste06, WW02a, AS09a, AHRT09, ACJ07, ADP08, AK08b, AW09, Ano00-30, Asi08, AM04, BBGG08, Bib06, BW05, BW06c, BG07b, BRMS01, CÁHCJL^{+}09, CKCG03, CL07b, CAZAA00, CWZZ07, Che08a, CW05, DLL08, fDxZ11, DKO03, DAKO06, DFSS08, Dra06, EPS05, FZG07, GPvZ06, Gal08b, Gal09a, GT00, GS07a, GM09b, GHW04, GSS04, Han04, HJW05, HY03, Hof03, HT05, HZ08b, HHW08, IDOK06, JKN09, JEP08, JPV02, KÖ08b, KK06c, Kaw06, KL08a, KY07a, KY07b, KNP06, LKH03, LC07a, LL07, Li08a, LCL08, LYJ09, LW09d, Li09a, LDHW05, LN09, LPvST07, LC06c, MBI08, Mah09, McG03, MS09, Mur08a, MDPD02, NQHN03, NTTS09, OEV09, ÖY09, PCE^{+}05, Pao01a, PXGL09, PXW09, Pet04, PD06, PJJP04, RB06, SSX09, Saw03, SB03, Sun05a, SL07b, Sun09a]. time [Sun05b, Sun06a, TLS07, Ter00, Top04, Tou05, TYW05, WZ07b, WW08b, WWW08, Wan08a, Wan09e, WWW^{+}09, WZZ09, WS05, WZ07c, XC00, XLL03, YX07, YC07a, Yas08, You04, YH03, ZM94, ZW04, ZS05, ZS08a, ZYG^{+}09, Zha09e, ZTH09, ZLZ09, ZW03b, Zho08, ZH04]. time- [You04]. Time-constrained [WW02a]. Time-Delay [OK06, ZTH09]. time-delayed [Kar05]. time-delays [LYJ09, TLS07]. Time-Dependent [BFL06, AS09a, GS07a, HJW05, KY07b, LDHW05, PD06, RB06, WWW^{+}09, Zha09e]. time-dependent-coefficient [GT00]. time-integration [CAZAA00]. Time-Integrators [Gir06]. time-invariant [Ter00, ZM94]. time-limited [KK06c]. Time-optimal [Kar05]. time-quality [Dra06]. time-stable [ACJ07]. Time-Vary [GY06]. time-varying [Bib06, Che08a, Han04, HHW08, Li08a, LC06c, Mah09, PCE^{+}05, PXW09, Ter00, YX07, ZM94, ZS08a]. timely [LJSH09]. times [AHRT09, BHNO04a, BGO04, KA02, TH02, WWW^{+}09]. Timing [BM00]. tissue [FGX05, ZS07]. tissues [BBH07]. Titchmarsh [Cha08]. Tobias [Ano04d]. Tobin [Ano03-66]. Tocher [LF09]. Todd [Ano01e]. Toeplitz [Bax02, BFM04, CWZ03, GS01, LHR09, MRT05, NV08, Roj08]. Toeplitz-plus-Hankel [Roj08]. tolerance [DM09]. tolerant [FL05, SMS^{+}04]. Tomaso [Ano03-61]. Tomographic [Bal00]. tomography [WWH05]. tone [NKCS08]. tonometer [Hai03]. tool [SNG09, Waz07b]. tools [GHC09, SBSY06]. Topological [LI00c, Tho01, AO01b, AT05, Din00a, Din00c, FjH07, Li04a, XM08]. Topology [CCF09, Nik06]. TOPSIS [WL07b]. TOR [Cha01a]. tori [SWKT08]. Torsion [Has03, NLR00]. Torsional [SHI09]. Torus [BS01b, Dor00, YMTE05]. Total [Sar06a, GS07b, HX07, MS09]. total-batch [GS07b]. totality [CZ06]. Totally [LBW05, Gal08c, JPS09]. Toxoplasmosis [AVAGP08, GPAA^{+}09]. TPMACA [CCK03]. TPR1 [PKM^{+}06]. trace [PS07b]. track [Ano03-58]. tracked [DGG^{+}04]. tracker [GLTS08, TLS^{+}08]. tracking [Ano03-37, Kar05, LNL08, LLS^{+}09c, PGH^{+}08, TLS07, WWHL09]. Trade [YX09b, Ano01z, Dra06, HL08b, LW09d]. Trade-off [YX09b, LW09d]. trade-offs [Dra06]. traffic [Ano00-85, BC08, SWK07, TGIM02]. Trail [Wac08]. trailing [TYS03]. Train [Ano04g]. training [ZXTL09]. traits [Lan08]. Trajectories [XWL06, FY01, Gal08c, Zha02c]. Trajectory [LNL08]. Trajtenberg [Ano03-60]. transcription [MD06]. transfer [Bat07, Bib06, Dog01, GW06, Jum01, OIL08, VC03, XL09]. transferring [WY01]. transform [AMT03, AO08a, BHHA08, BJ00, De 02, DDF02, FTC03, LWZ04, LCWL06a, MG05, ME08, Pet04, VA04, ZMLG09, RR03b]. Transformation [Ano00-59, RSD09, CO00, CK06, Ely04, Li01c, LMHW05, Ung03, WX03c]. transformations [GM09a, Kan02, TB09, Wan09a, Yan03a]. transforms [AETEZ07a, BD01, BS00a, LS04a]. Transient [ASESESA09, KA00b, PS08a, TEB06, AS09b, BSR02, Che04a, PSM02]. transiently [LW09f]. transition [DMP03]. transitive [FLR05]. Translated [Ano03-46, Ano03-67, Ano04-31, Ano00-58]. translating [GX04]. Translation [SKHK06]. translational [Yav05d]. transmissible [CCF09]. transmission [AMC08, DLB03, IKY06, LW09d, MS04a, Yan02]. transmittal [Gal08b, Gal09a]. transonic [BRC00]. Transport [DRR09, SW06, AK06, Col02, DE06a, HMCR00, HAKA07, Lin08c, LLC08, MP02b, MO08, dSM02]. Transport-diffusion [DRR09]. Transportation [ZS07, Ano00-85, HY08, Ano00-85]. transposition [CLS08]. TRAP [PKSP08]. trapezoid [BD02, BCDS09, DDF02]. trapezoidal [AH09, AC02, AC05, BS00a, CAZ01, CT08]. Träumerei [BM00]. Traveling [Fan02a, KÖ09b, SHY08]. travelling [YB08]. treatment [AAK08, BM05b, JYDH02, Kha05, MG08c]. treatments [VC03]. tree [HLC00, KM03b]. Trees [YW08b, Ano03i, CLS08, Hu08, JS08, KM01, Lon08, MIPK04, MP05]. Trefethen [Ano03-66]. Trefftz [CWLC07]. trenches [DPU07]. triangle [WX03a]. Triangular [Las08, BMHV09, BKK08, Che04g, Cho03, Gal05, LY08d, TSZ02]. triangulation [CL08b, LW09a, LHC05]. triangulations [Pla07]. trichloroethylene [CCK08]. tridiagonal [CPS05, GS01, PSES07, WH09, YLL08]. Trigonometric [MCS08, Bal02, BCGA08, MJD08]. trigonometrically [KS03b]. trigonometrically-fitted [KS03b]. trigonometry [Ung00, Ung01]. triopoly [EAE09]. Triple [BLD07, Bai08, CW05, KLW09, LW06, LW08b, Liu09a, SFL02, Sun09a, WF05, Won00, ZC09a, BW05, Kar08, LW08a, Mal04, WG07b, hYL09, ZL07b]. triple-point [Kar08]. trivial [ZL09a]. Tropical [CZLN06, HHL^{+}06]. Troy [Ano00-91]. Truesdell [Sam07]. truncation [AVO08, AN05, CGBJR03, El 03]. truss [WGF09]. truss-core [WGF09]. trust [CSdS08, zGtM08, JZQZ07, QJZ09, ZZ03a]. trust-region [JZQZ07, QJZ09]. Tsotras [Ano00d]. TU [ND01, LD08]. TU-games [ND01]. tube [DH08a]. Tucson [Ano00-84]. Tumor [ACLL06, LH00a]. tumour [MM08c]. tumour-like [MM08c]. tuning [BK07]. Turán [Kov00]. turbulence [KIXF09, RZ03, Sar03]. turbulent [TYS03]. Turek [Ano03-88, AMV08]. Turing [Pio00, Ano04-30]. turning [KP01]. Turow [Ano04-29]. Turunen [Ano00-47]. TVD [Soh05]. twelfth [KA09b]. twelfth-order [KA09b]. twice [LV03]. Twin [ACH01, ABS09]. twin-tail [ABS09]. twisted [Lu09c, OS08b, YY07]. Two [AORS03, Bel06, BW06c, ES01, FFXL05, HL02b, ICM09, KO09a, LCYL00, LZ04b, LX06, LD08, LL02b, PY09, Sir00, YX09a, AS09a, AK08a, AC04, AB05, Ana04d, BP09, BB08a, BCDF02, BJ02, BJK^{+}09, BA09a, BFM04, BW05, CG05, CK03, CT09, CM04, CAZ01, CY00, CGBJR03, CWZ03, CKR03, CK04, CM02, CM08, DS02, Deh02, DS07, DMSV00, Den07, Dha03, EGA03, EZ08, FR04, FP02, FM05, GGD00, GS07a, Ha01, HL02a, HG03, HN07, HL03b, Isk08, ID10, JW05, JT07a, JT07b, KM07, KK09, KC02, KF05a, Kay04, KY00, KK05, KK03, KU01, Lan04b, LZ09, Lev06, Li01b, LTWM08, LC04, LH06, LDK08, Liu08a, LY08d, LdSF00, Lü08, Mar07, MAR05b, MYR09, MEG08, NQHN03, NUN06, PWX08, RR09a, RRR07, RR08, RZ03, SMW04, SO03, SSLL09, SF01, SPW08, TT09, Tsi02]. two [WCL02, Wan03a, Wan03b, WZ07a, WW09c, Wan09d, Waz05d, WFTL09, WBSP08, WX00a, WZ08c, WW08d, XXLD04, XW05, Xie08, XC00, XYC^{+}09, YE01, Yao04, YH03, Zen00, ZCS04, ZZ01a, ZZJ^{+}04, ZS04, ZYY08, Zha09f, ZZ09c, ZZZ02]. two-by-two [TT09]. two-component [WZ07a]. two-components [ZZ09c]. two-dimensional [AS09a, AK08a, BA09a, CK03, Deh02, DS07, Den07, FP02, HL03b, JT07a, JT07b, KC02, Kay04, Li01b, LTWM08, LC04, LY08d, RRR07, RR08, Wan03b, Waz05d, WZ08c, XYC^{+}09]. Two-direction [YX09a, Xie08]. two-equation [RZ03]. two-grid [CWZ03]. two-grids [CG05]. two-grids/projection [CG05]. Two-layered [ICM09]. Two-level [KO09a, Sir00, NUN06]. two-levels [BFM04]. two-neuron [LH06]. two-part [SSLL09]. two-patch [XC00]. two-phase [BJK^{+}09, LZ09, MAR05b, SF01, WCL02, WW09c, ZZJ^{+}04]. Two-point [AORS03, BW06c, BW05, CM04, CM08, EGA03, FR04, Ha01, HL02a, HG03, LDK08, MEG08, SPW08, Wan03a, Wan09d, Yao04]. two-sided [ZZZ02]. two-space-variables [Mar07]. two-species [XC00]. two-stage [SMW04, WFTL09, XXLD04, XW05]. two-state [EZ08]. two-step [AB05, BJ02, CM02, FM05, HN07, Tsi02, WBSP08, WX00a]. two-stream [DS02]. two-type [KF05a]. two-unit [SO03]. two-variable [BP09, CGBJR03]. two-variables [ZZ01a]. Type [GA06, ND05a, Xu06, YI06, AH03, AML02, Agr08, AP04, ACM05, ADP08, Ana04a, Ana04b, Ana04c, Ana04d, Ana07a, Ana07b, Ana08, BUP09b, BM04b, BCRS05, Bar01b, BCDS09, BBI05, Ber05, BCR04, CCP02, CCD07, CCS00, CZ03a, CS04a, CWL07, CHZ09, CZ09b, Che01c, CC02, DFFL09, DJ08, DMP01, DK05, Dib03, Din02, DDF02, ES08a, ESR06, FjH07, GA07, GS08, GJQ07, Has02, He04, HCLJ00, HF04, HL07, Ibr09, IK03, JAM02, Jan04b, Kam09, Kan04, KF05a, KDC08, KKS04b, KL06, KLW09, KT03b, Kum05, LKS08b, LDZW05, LWY06, LS08, LSL09, LLJ09b, Li09a, LF02, Lin05, LWG04, LHC05, Liu07b, LN09, Lü08, Luo06, MCH01, MPU00, Mat03, MO01, NWL08, Nie09, Nko00, OI00, ÖY09, PSS09, Pao01b, PXW09, PP01b, PS09, RC07, SR03, SP08a, SYHC06, SDCL08, Sun09b]. type [Tan02b, Tao06, TOH03, TK07, Tia05, TCH07, THD08, Tsi03, Uje03, Uje04b, ÜÇT08, VR02, Van07, WLW07, WHS08, XC06, XWW08, Xie01, XLL03, XLT09a, XLT09b, Yan03d, Yan06b, YL09b, Zen00, ZG02, Zhu08a, ZS08b, Zhu09a, ZCC04, Mis05, Hua04]. type-2 [LHC05, YL09b]. Type-I [Mis05]. types [HB04, ID10, JW05, Kon09, MRS08, MP08b, WYL08a, ZX08].

U.K [Ano00-74, Ano00-41]. U.S. [Bel05a, Bel05b]. U.S.A. [Ano04c]. Uhlenbeck [APvZ01, AF02, BPvZ03, BB01]. Ulf [Ano01y]. Ulrich [Ano00q, Ano00r, Ano03-70]. ultimate [ZLCM08]. Ultra [LYF09]. Ultra-High [LYF09]. ultrasonography [BHCØ07]. ultrasound [FGX05]. umbral [CGL01, Fer01, SV01, Tay01]. umbrella [LU09a]. Unbounded [CDM03, Leu00, LFZ00, BPP04, Boy01, Den07, FB05, HH02a, HH02b, HJW05, MRMRM02, MK08, Ren03, XTX01, Xu07a, Zha04b, ZLH07, ZSZ06]. Uncanny [Ano03-89]. uncapacitated [HTKM09]. uncertain [CCJ08, MBI08, PCE^{+}05, PG08, WK08, XT08, YX07, YL04c, ZLCM08, ZL09b]. uncertainties [Che08a, Gal02a]. uncertainty [Ano03x, BHHA08, CS05b, CSPJ05, Gal08b, Gal09a, Sun05b, Sun05c]. unconditionally [ACJ07]. unconfined [De 04]. unconstrained [FT05, zGtM08, SH01, WS08b, WYSS09, XWW08]. undergraduates [Ano00-35]. undetermined [CN08a]. unification [KC03a]. Unified [LW06, WF05, BM05b, BGS04, LTS03, MG08c, Yan06b, ZXX05]. Uniform [AE07, DDX04, EIR03, Li02, LB01, Xu02a, Ana05, Has03]. uniformisation [FP06]. uniformity [GG06]. Uniformly [DS05b, cWcX09, cW09b, AD03, Din04a, udK07, GSS02, Hua00, LK03, MN01, Sar03, SM00, ZLCM08]. unifying [LGBS00]. unilateral [RSK02]. unimodal [LU09a]. Unimodular [Ber00]. unique [LBZ^{+}09, WLW07]. Uniqueness [BD07, CH01, CZLC08, FF02, Fan02b, LLWY00, Liu08c, NSQ02, OEV09, Xu08a, Yan04b, Yan05a, Zha08b, ZL08b, AAD03, AC03, BM01, BM02, BM04a, CYX09, CZ09c, Fen08, GS08, GN09, GLZ00, Hua04, LZJ04, LJL06, Liu04a, LY05b, Liu08a, MRMRM02, SWX06, Yan07]. unit [Ana00a, BC02, BCA09, GM09a, GM00, Gor00, KKS^{+}04a, SO03]. units [BKR05]. Univariate [Ana04e, BD04a, BGP04, Lin04c]. universal [MWZ09, YSP^{+}08]. universe [LML02]. University [Ano00-74, Ano00s, Ano00t, Ano00p, Ano00-31, Ano00-42, Ano00-41, Ano00-51, Ano00-64, Ano00-80, Ano00-81, Ano01z, Ano01o, Ano01q, Ano01v, Ano03-76, Ano03f, Ano03m, Ano03-27, Ano03-29, Ano03-80, Ano03-43, Ano03-49, Ano03-57, Ano03-62, Ano03-66, Ano03-68, Ano04e, Ano04g, Ano04c]. UNIX [Tun05, Ano00-86]. unknown [Jan07]. Unleashing [Ung05]. unofficial [Ano00-82]. Unreliability [Lin06b]. Unreliable [WCL02, KLHC08, ZBS03]. unrestricted [Tan03b]. unscented [AF09]. unspecified [TL07]. Unsplit [PLD02]. Unstable [Far01, DJ08, Lin05, Tan02b]. Unsteady [ASHA08, FAF09, De 04, IMLF09, NMD09, WFC04b]. unstructured [BMH08, CHS02, SMW04]. unsymmetric [JZ02, PB02]. unsymmetrical [CB02]. untracked [DGG^{+}04]. update [CKL07, WS08b]. updated [XZ09a]. updates [For01]. updating [LY09b]. Upper [Ano03-59, Ano04r, BHNO04b, CZLN06, Jia01b, Jia01c, Jia02, ZJ08, AOLL04, COEP00, CG00, CPZ00, DSS00, HYZ06, Lin07, SA05, SAS06, SLJ04]. Upwind [AS09a, Gla00, Gla03, Gla04, Gla05a, Gla05b, Gla06, Gla07a, Gla08, Gla09, LY08a]. Upwinding [BOP07]. Urysohn [AO01c, BPP04, ESR06, Ibr09]. use [AS01, Ano00-58, FM05, Lun03, Mär05c]. used [AAK08]. useful [Pio00]. user [KPD^{+}09, Ano00-87]. users [Tsa03, YCCY02]. Using [CW06, HHL^{+}06, HRM08, Pet04, WI08, AETEZ07a, AH00, AS07, AM09, AZB08, AN05, AF09, ACCMJ07, AIS08, AM08b, Ano01w, Ano03-34, Ano03-73, AS09b, AO08a, AC09, BMHV09, BN05, BI02, BGLT08, BS03b, BHCØ07, Boy01, BCT00, CZLN06, CLJ09, CLMM06, CM04, Che00a, CKY07, CWW08, CSC^{+}09, CCP09, Col04, CA09, Cse06, DSHS05, DS07, DZQ06, Dim08, DR02, Dua08, ED09, Fan02a, FL03, FP06, GA07, GHC09, HL02a, HD09, HTKM09, Hau06, HHS06, IUCR06, Isk02, Isk08, KA09a, KK09, KK06b, KY00, KHN09, KA00a, KKS04b, KDKS01, KÖ09b, KB08, LHC08, LH00a, Lee00, Lee01a, Li01c, LGGSP06, Lin08a, LBW09, LNL08, LF09, LW09f, Lu09b, MP08a, MP07b, MO08, MEG08, NMD09, NV08, OY06, OIL08, OIL09, Pan00, PdlF07, PLP09, PD04, RNM00, Ryo00b, SALK02, SD07, Sha04, SDB09]. using [SHI09, SGR^{+}09, SK06c, SBD^{+}00, Swe07, SK09, TS00b, TV07, Trz01, TDS08, VDMK08, Wan02d, mWmZ09, WNS09, WZ09, Wu03, WH07, WW08c, XZ06, XLZ02, XYC^{+}09, YZ02b, YW09, Yav05b, Yav05d, Yil08, YTEC02, YG08, Zha03b, ZH07, ZLN06, ZBS03, Ano00-88, Ano00-89]. utility [KKY06]. utilizing [SN02].

V [Ano00o, Ano00-86, Ano01n, Ano03-44, Ano04w, Ano04v]. vacation [ODO00, SG08]. vacations [Gup02a]. vaccination [PLG^{+}09]. Valencia [JSGP08]. valent [AM08c, MA09b]. validation [BJK^{+}08, PM08]. Valiente [Ano03i]. valley [Qua01]. Valuation [EBVS06, CM03, TAT07]. Value [Liu06b, ACD01, AP08, AS03b, ATT03b, AOLL04, AOS07, AS01, AD03, AA01, And03, AK08b, AS08, AC03, ACH01, Ayk04, BG04a, BLD07, Bai08, BN00, BD07, BW05, BW06c, CH01, CC09a, CG08, CM04, CS02b, CK02a, CBJ01, Che04b, CWZZ07, CO07, CT08, CL08a, CZLC08, CH00c, CW05, CZ08, CZ09c, Cui09, CM08, DS02, DHM05, DG05, Du08, EGZ04, EGA03, Elk06, EBVS06, Fan02b, FZG07, FR04, GNR05, GH03, Gra03, Ha01, HL02a, HG03, HW09, HT02, HY03, HLC05, Hua08b, IS04, Jan05, Jan07, Jia00b, Jia01b, Jia01c, Jia02, JX04, KK09, KA09b, KK04, Kel06, KS00b, KL08b, KKZ02, KLW09, KS03b, LR03, LLWY00, Li01a, Li04b, LS08, LLJ09b, LG09a, LDG05, LJL06, Liu02, LWG04, Liu04a, LQ04, Liu04b, Liu04c, LY05b, Liu07a, LLW07, LDK08, Lu07, LSN05, LJ08b, Ma00, Ma01]. value [Ma04, MDG05, MZ09, Mar07, MM05b, MF01, MEG08, ND01, NF03, Net05, Net07, NRL00, NRL08, NMD07, NMD08, OEV09, Pal01, PG04, PLG07, PGT08, Pao03, PF01, RSK02, SA05, SR04, SC03, Sta02, SPW08, SZL02, Sun09a, TLLM08, Tho01, TGS07, Top04, Tou05, Wan00e, Wan03a, Wan03b, WG07b, WLW07, WJA08, WWW08, WSL08, Wan08a, WG08a, WFY09, Wan09d, WWS^{+}06, Waz00, Waz01, WA08, Won00, WX03b, WCW08, Won08, Wu09, XXD06, XZ08b, Xu08a, XG07, XTZ08, Yan06a, Yan07, YW07, Yao04, Yao05, Yao07, Yas08, ZM04a, ZLY06, ZL08c, ZCL08, ZYW08, ZL08d, ZJ08, ZL08b, Zha09c, ZG09a, Zha09d, Zha08d, ZG09b, ZC09a, ZCC04]. value-sharing [BD07, Fan02b]. valued [AO01b, ARM08, Ana08, BD01, BGMB02, BBCD02, CGY08b, CC08a, DLF09, Dha07, FW08, GYPL09, HRMS09, HCLJ00, Hua01, HF04, Jeo04, JY11, Li07, MN01, Mus06, TSZ02, YLY^{+}09, YLZ08b, ZGY05b, ZGY07a]. values [CYX09, FW04b, KDKS01, Lam05, LC03, LYX09, MSS08, SYJJ09, TK02, Yan05a]. Vandermonde [HP02, Jog04, SV09]. vapor [Jon07]. vapour [BBB08]. VAR [HHL^{+}06]. Variable [IJB00, KP01, LML02, LL06a, Sha02, ZL06, ACCMJ07, BP09, BHNO04a, BGO04, BB06, CQJ03, CGBJR03, Che04c, CA09, DSS08, Dix05, FZ07a, Gal09b, JMR00, JP01, KZ02, KHN09, Kum04, Mak09, MHS01, MG03, PDRR08, RDS00, TZS03, WT00, WZ00, WF07, Waz05a, Xia07, XZZ01, XTZ08, Zha04b, ZSZ06, ZZ05, ZZ03b]. variables [AORS03, Bar01a, BDA02, ES08a, FZ07b, JSS01, KK05, Kum05, Lü08, Mar07, PL05a, ZZ01a, ZL05, Zhu08b, Ano00u]. variance [ABCD05, BDA02, Ano04m]. variant [Gup04, SSSH01, ZCM01b]. Variants [Waz04, Waz05d, PKSP08, Waz05b]. Variation [Sar06a, BS06, CCD07, Dra08, THD08, XC07, Zen00]. Variational [AI06, AU09, AL08b, BNH07, GTJ07, HW07, Hem08, KA09b, KZD03, LX06, Li09b, Lu07, NMD09, SD09, Sho09b, SAB07, XH07, Xu07b, Zha07a, Zho07, AETEZ07a, AETEZ07b, AAI05, Alb02, AC09, BK00, BZN07, BG07a, BXFZ09, Bou00, Bou03, CLY08, CWY08, CGY08a, CGY08b, CKN04, CA09, Das09, DY05, DYZ08, FW08, FHT05, FH07, GL08a, GL09a, GAT07, GL09c, GLC09, GJ07, HL02b, HS04a, Han06, HLW07, Han08, HWX07, fHCH09, HBCK00, Hua01, HF04, HN07, HZ07, JG07, Jeo04, Jin07, KB04, Kha01, LCV06, yL08a, yL08b, LKS08a, LKS08b, Li07, LW02, LW09e, LY05a, Liu09b, LHH08, Man05, MH07, MZ07, MO07, Noo00a, Noo00b, Noo02a, OM09, PSS07, PL05b, PZ07, RD07, RN02, SDVS09, SNT08, SG07, SGM09, Sal08, SD07, SP08a, Slo07, Slo09, SZ09a, SZ09b, Swe07, Tao09, TD09, Tia01, Ver01a, Ver01b]. variational [Ver03, Ver09, WYH00, WXW01, mWmZ09, Wan09d, Waz07c, Waz07b, Waz07a, Wu09, XWZ04, XZ09b, YHS08, YW09, YM09, YLD09, YB07a, ZGY05a, ZGY05b, ZGY07a, ZGY07b, ZAY08, ZH07, ZH09, Zha09c, dSRGRMRL08]. variational-like [DYZ08, GL08a, KB04, SP08a, ZGY05b, dSRGRMRL08]. variational-type [LKS08b]. variations [BG07b, LG09a, VB01]. Varieties [MPP08]. variety [WZ08a, ZW09b]. various [WCW08, XM08]. VARMA [Ter00, ZM94]. Vary [GY06]. varying [AVAGP08, Bib06, Bra08, Che04c, Che08a, Han04, HHW08, Li08a, LC06c, Mah09, PCE^{+}05, PXW09, Ter00, YX07, ZM94, ZS08a]. vasa [MT00]. vascular [CCC08, ZCCB08]. Vassilis [Ano00d]. Vector [Sid08, TSZ02, AO01b, Ana08, BBCD02, CO00, CGY09b, CSX01, CHZ09, GL08a, GL09a, Gos00, HRP04, HSS01, KKS^{+}04a, LCJL03, LKS08b, Liu04a, MSN^{+}00, Qiu08, Rap01, SSL07, Shi03, VJ04, WSO08, WX00b, YL07]. vector-host [Qiu08]. vector-valued [Ana08, BBCD02]. Vectorial [DR02, dSOGRMRL04, dSRGRMRL08]. vectors [TK02, YW08c, YX09b]. vehicle [Mos08, WW09c]. Velde [Ano00-38]. velocity [CMS04, HT04, Mie01, PW00, Ung00, Ung07, Yav00]. Velte [SB06]. vendor [MGG05]. Verification [PDP05, NLR00, Ryo00b, Ryo00a, RN02, RSK02, Ryo04, Ryo07]. verifier [YXZL09]. Verlag [Ano00q, Ano00r, Ano00o, Ano00-48, Ano00-49, Ano00-47, Ano00-46, Ano00-58, Ano00-70, Ano03-37, Ano03-82, Ano03-48, Ano03e, Ano03-69]. version [AS03a, Ano00-74, BS03b, CD05a, EA03b, LCJL03, PDBH03, SKG01, WXW01, Ano03s]. versioned [ROY01]. versus [DW02a, NF03]. Vertex [SWKT08, KKS04b, LTTH09, WHL^{+}09]. Vertex-bipancyclicity [SWKT08]. vertical [Fro08, FG00, INP08, ZL09c]. vertically [HW05c]. vertices [MP08b]. very [KDKS01]. vessel [MM00, MG03, SM08]. Vesselov [SDB09]. vessels [MP02b]. vex [SJL09]. Via [DE06b, BMHV09, BD01, Bax06, BA08, CYG06, CM09, CD05b, CD09, CVW00, CWZ03, DCW08, Dib03, DDF02, FFK02, HRM08, JS03, Jum01, KLP09, LC06a, LWY02, LTS03, LWY04, LY05a, MM05a, NST01, Pal01, PMR^{+}08, RNM00, RR03b, Ser00, Sol00, XMW09, YX09b, Zha02c, ZZZ09, ZW06]. viability [Ano03-62]. vibrating [FG00]. Vibration [ABS09, TV07, CB02, LHC08, PB01, VG08]. Victor [Ano03-64]. video [WL05, WLL06]. Vides [Ano03-41]. Vidyasagar [Ano03-48]. Vienna [Ano00-48, Ano00-49]. Vietoris [LK04]. view [GZZ09, Lee01a]. VII [ND05a]. VIII [Ano03k]. viral [WWL06]. virtual [Ano03-89, Ano03-88, HCY00, SL00b, Ano04-31]. vis [MG01]. vis-a-vis [MG01]. visco [GGD00]. visco-elastic [GGD00]. Viscoelastic [Mig06, Bat07, De 02, FS04, Kan07, Lee04, Lia05, NS07, Win07]. Viscosity [APS07, dG09, CGY09a, KA09a, Mak09, QS03, Zha01a]. viscosity-like [CGY09a]. viscous [Con03, Dem00, Fei07, MG03, MRMRM02, Pad06, ST07, Sta07, SZL02]. vision [Jum01]. visual [BBGG08, ZZZ09, Ano01-27]. Visualization [MO08, Nam01b, Nik06, ZZ06]. vivo [LMSH04]. VLSI [Ano04d]. voice [RS09]. voids [IN07]. Volatility [TA06, LPRH05]. voltage [IUCR06]. Volterra [AOTW07, ADN02, ABB05, BMHV09, BS01a, BM04b, BCRS05, BM05a, BNH07, CF04, CK04, CP01b, DI08, EIR03, GK02, GY09, Had03, Huo04, Kan04, LDZW05, LY08b, LL02b, LL03c, Med01, NTHP09, PS04, PP04, TL00, WW04b, Wan09e, Xu02a, Xu03a, YLD09, ZSG04, ZL02, ZJLW06, ZLZ09, ZCC04, ZZD08]. Volume [Ano03-38, Ano06l, Ano03-28, BM05b, BMM^{+}06, CCC08, GY08, NS07, RM07a, RM07b, Wan02c]. Volumes [Ano01l]. Voronoi [DW02b, DW05]. vote [Bel05a]. VRG [ZCCB08]. Vries [HKL03, LRL06, ME03, Ras06, UHuI09, Zha07b]. VSVO [BBL05]. vulnerability [MV08b]. Vygen [Ano01g].

W [Ano00-35, Ano00-84, Ano01p, Ano03v, Ano03x, Ano03-28, Ano03-46, Ano03-49, Ano04k, Ano04y, Ano04z]. Waerden [Gyi01]. wages [Ano01z]. Wagner [Ano03-33]. Wagon [Ano00-46]. Wahlster [Ano03-71]. wait [XXLD04]. Waiting [Kaw06, GSS04]. waiting-time [GSS04]. Walk [DL06]. walks [RRR07]. wall [JS04, MM00]. Walsh [Ano04n]. Walter [Ano00-38, Ano00-91]. Wanner [Ano03-42]. War [Ano01x, Ano01-28]. Waring [Cha01b, Ben00]. warning [KS03a]. Warranty [SY05, RS06]. Washington [Ano00-32, Ano00-33, Ano00-35, Ano00-66, Ano03-35, Ano03-58, Ano04c]. wastewater [IMPF08]. Water [OK06, BFO02, Dor00, Gir03, Gla00, Gla05a, Gla07b, Hav08, JTRu09, LWZ08, MIPT08, PM08, VC03]. Watkins [Ano03-39]. Wave [Gou01, Hav08, OS08a, BS00b, CJK09, Fan02a, GMS08, GHJ06, HT04, Hem08, JG07, KLJ09, KÖ09b, LYW04, Phu02, SD08, SHY08, TG03, VNG04, WYL08a, WX09b, Waz07b, Yan03a, YB07a, YB07b, Zha07a, ZTW09]. wave-making [HT04]. Waveform [Jan05, BNP05, BK04, JZ08]. waveforms [Abu01]. wavefronts [MRT04]. waveguide [SBD^{+}00]. waveguides [RV01]. Wavelet [Ano03-52, AC00, Ana04c, BJBL03, CWZ03, EBVS06, FTC03, HMV04, KGB07, LWZ09, Lin00, MS03, Wan00c, Xie08, ZMLG09]. wavelet-collocation [EBVS06]. wavelet-like [Lin00]. Wavelets [Kar03, Pat05, QFZ03, ADH^{+}03, Cat05, FQZ02, He04, LH00a, Lee00, Mae00]. waves [Aro08, BGR09, BDO04, BD04b, DH08a, DLSB08, Dor00, KKM00, MP07b, RR03a, SD03]. Weak [CZG04, OI00, ZHZ00, CF04, CLT09, Gir03, NVAR03, XXD06, dSOGRMRL04]. weaker [GZ08]. Weakly [BGR09, ĆRRU08, GK02]. weapons [Bog09]. web [Ano03-71]. wedge [WY07]. wedge-shaped [WY07]. Weierstrass [KI04, SERG09]. weight [BKN09, DGVP06, HH04a, HTKM09, WGF09, YW08a]. weight-shaped-posets [BKN09]. Weighted [Bal04, Ban07, Che04d, CFH05, SC02, YI06, AP04, CH05a, DW02a, GA07, GPL08, HJWL03, HCWC07, JH09, Kua08, LLW08, LU09a, LYX09, MS09, OY06, PST08, TZS03, THD08, WW00, ZL06]. weighting [KLP09, LH05b]. weighting-dependent [KLP09]. Weirstrass [SL06b]. Weisler [Ano00-83]. Weissig [Ano03-74]. Weiszfeld [ÜL00]. Weitzenfeld [Ano03-86]. Weixiong [Ano00-70]. Well [FHH08, Gup02a, FH07, Gos00, RM07a, RM07b, WX09a, YLZG01]. well-balanced [Gos00]. Well-posedness [FHH08, Gup02a, FH07, WX09a]. well-spring [RM07a, RM07b]. Welsh [Ano00-67, Ano00-68]. WENO [CCC08]. Werner [Ano03-67]. Wesslén [Ano00-28]. Westview [Ano03g, Ano03s, Ano03-32, Ano03-33, Ano03-70]. Weyl [Sau07]. wheel [ZLY09]. wheels [FY04]. Which [ZLY09, Dra08, HLC00, Jan07]. while [She04b]. White [Nam02]. Whitham [RD07]. Whittaker [AMT03]. whole [MO01, Nam02]. whose [BCDF02]. Widder [Ana07b]. Wide [Ano03-71]. width [YB07b]. Wielinga [Ano00-38]. Wiener [KYAA08, WD08]. Wigner [LKS^{+}06]. wildland [MKF09]. Wiley [Ano00i, Ano00z, Ano00-90, Ano03-31, Ano03-39, Ano03-73, Ano03-74, Ano04m, Ano04s, Ano04u, Ano04t]. Willi [Ano01w, Ano03-43]. Willi-Hans [Ano01w]. William [Ano00-36, Ano00-80, Ano00-81, Ano00-88, Ano00-89, Ano01l, Ano03v, Ano03-33, Ano04l]. Williamson [Ano03q]. win [Bel05a]. Win32 [Ano00-63]. window [Ano03-60, LGH09]. Windows [Ano00-72, Ano00-73, Ano00-57, ZYG^{+}09, Ano00-92, Ano00-91, Ano00-93]. Winkler [Ano04i]. Wintner [Che01b]. wired [Ano04-29]. wireless [JCCT09, JCC^{+}09, LLS^{+}09c, WLT07]. wise [Kan02, Ano03l]. Wishart [RV05]. within [GPL08]. without [Ano03-90, CWY08, GL09a, JT05, Lin04b, Mär05c, Son08, WF01, WXS01, YXZL09]. Witzgall [Ano03-46]. Wohlin [Ano00-28]. Wolfe [LLL02]. Wolfgang [Ano01i, Ano03-71]. Wolfram.Westview [Ano03n]. Wolstenholme [WD07a]. Women [Ano01-28]. Wood [Ano00-60, Ano00-61]. Wooldridge [Ano01t]. word [CSZH09, PCA09, Ano04-32]. wordlength [ABH08]. words [BS04, Bs03a, Ano03-90]. work [Bat07, Pad06]. workflow [LLL08a]. workspaces [HSM00, HS02]. World [Ano01x, Ano01-28, Ano03-71]. Worm [OSS07]. Worm-Like-Chain [OSS07]. Wright [Ano00-69, PS09]. write [HLM04, LHYM09]. Wyndham [Ano00-90]. WZ [CJ04].

XOR [CCHK08]. XOR-based [CCHK08]. xxiii [Ano03-32].

Yan [Ano03-56, CZ03a]. Yannis [Ano00d]. Yap [Ano00-31]. yarn [YW09, ZH07]. Yau [CY05]. yearbook [Ano03-57]. years [Ano00j, Ano00k, Ano03-70]. yeast [GMG09]. Yeyati [Ano03z]. yielding [Kr05]. Yorick [Ano01w]. York [Ano00i, Ano00s, Ano00t, Ano00z, Ano00-31, Ano00-51, Ano00-46, Ano00-71, Ano00-70, Ano00-90, Ano01p, Ano03i, Ano03-76, Ano03h, Ano03m, Ano03r, Ano03-30, Ano03-39, Ano03-34, Ano03-37, Ano03-40, Ano03-44, Ano03-43, Ano03-45, Ano03-46, Ano03-51, Ano03-52, Ano03-84, Ano03-56, Ano03-64, Ano03-65, Ano03-73, Ano03-67, Ano03-75, Ano04e, Ano04g, Ano04j, Ano04w].

Zaidel [Ano04-28]. Zakharov [BBA09, JG07, MYG09, SHY08, lYsT09, Zha07a]. Zedan [Ano00-55]. Zekal [NLL08]. Zeller [GJQ07, Zen00]. Zellner [Ano03-80]. zero [BS00a, Coo05, PIR04, TK02]. zero-finding [PIR04]. Zeros [AZCK05, BPW00, Boy07, BCT00, GM00, Gra08, Ikh08, Pan01, PV00, SPV01, WX03c, WSZ03, YW03a]. Zhang [Ano00-70]. Zhukovskij [Din04a, DS05b]. Zimmerman [Ano03-85]. ZK [MYG09, XYC^{+}09]. Zytkow [Ano03-43].

References

Anderson:2001:MPS

[AA01] D. Anderson and R. I. Avery. Multiple positive solutions to a third-order discrete focal boundary value problem. Computers and Mathematics with Applications, 42(3–5):333–340, August/September 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                                       http://www.sciencedirect.
                                                       com/science/article/pii/
                                                       S0898122101001584.

Akalin:2005:PSD

[AA05] E. Akalin and M. U. Akhmet. The principles of B-smooth discontinuous flows. Computers and Mathematics with Applications, 49(7–8):981–995, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500115X.

Abelman:2008:ANP

[AA08] Shirley Abelman and Herven Abelman. Approximations of nonlinear phenomena arising in angular deviations of light rays that emerge from prisms. Computers and Mathematics with Applications, 55 (3):408–422, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107003732.

Anastassiou:2009:GPS

[AA09a] George A. Anastassiou and Ali̇ Aral. Generalized Picard singular integrals. Computers and Mathematics with Applications, 57(5):821–830, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005877.

Aygunoglu:2009:IFS

[AA09b] Abdülkadir Aygünoğlu and Halis Aygün. Introduction to fuzzy soft groups. Computers and Mathematics with Applications, 58(6): 1279–1286, September 2009.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210900501X.

Azam:2009:FPS

[AA09c] Akbar Azam and Muhammad Arshad. Fixed points of a sequence of locally contractive multivalued maps. Computers and Mathematics with Applications, 57(1):96–100, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006299.

Adimy:2008:NIM

[AACLM08] Mostafa Adimy, Oscar Angulo, Fabien Crauste, and Juan C. López-Marcos. Numerical integration of a mathematical model of hematopoietic stem cell dynamics. Computers and Mathematics with Applications, 56(3):594–606, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000515.

Anderson:2003:EUS

[AAD03] D. R. Anderson, R. I. Avery, and J. M. Davis. Existence and uniqueness of solutions to discrete diffusion equations. Computers and Mathematics with Applications, 45(6–9):1075–1085, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000877.

Ahmad:2005:GVI

[AAI05] R. Ahmad, Q. H. Ansari, and S. S. Irfan. Generalized variational inclusions and generalized resolvent equations in Banach spaces. Computers and Mathematics with Applications, 49(11–12):1825– 1835, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001999.

Alawneh:2008:NTS

[AAK08] Ameen Alawneh and Kamel Al-Khaled. Numerical treatment of stochastic models used in statistical systems and financial markets. Computers and Mathematics with Applications, 56(10):2724–2732, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004124.

Atia:2003:SGJ

[AAR03] M. J. Atia, J. Alaya, and A. Ronveaux. Some generalized Jacobi polynomials. Computers and Mathematics with Applications, 45(4–5):843–850, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000452.

Amat:2005:CNA

[AB05] S. Amat and S. Busquier. Convergence and numerical analysis of a family of two-step Steffensen’s methods. Computers and Mathematics with Applications, 49(1):13–22, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000039.

Abbas:2008:APS

[AB08a] S. Abbas and D. Bahuguna. Almost periodic solutions of neutral functional differential equations. Computers and Mathematics with Applications, 55(11):2593– 2601, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007596.

Akhmet:2008:PSD

[AB08b] M. U. Akhmet and C. Büyükadali.

On periodic solutions of differential equations with piecewise constant argument. Computers and Mathematics with Applications, 56(8):2034–2042, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003155.

Al-Bahadili:2008:NLD

[AB08c] Hussein Al-Bahadili. A novel lossless data compression scheme based on the error correcting Hamming codes. Computers and Mathematics with Applications, 56(1):143– 150, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008334.

Ashrafizaadeh:2009:CNN

[AB09] Mahmud Ashrafizaadeh and Hani Bakhshaei. A comparison of non-Newtonian models for lattice Boltzmann blood flow simulations. Computers and Mathematics with Applications, 58(5):1045–1054, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001102.

Arias:2005:NNI

[ABB05] M. R. Arias, R. Benítez, and V. J. Bolós. Nonconvolution nonlinear integral Volterra equations with monotone operators. Computers and Mathematics with Applications, 50 (8–9):1405–1414, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003901.

Ademoyero:2001:CLA

[ABBD01] O. O. Ademoyero, M. C. Bartholomew-Biggs, and A. J. Davies. Computational linear algebra issues in the Galerkin boundary element method. Computers and Mathematics with Applications, 42(10– 11):1267–1283, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002383.

Ademoyero:2004:CGA

[ABBDP04] O. O. Ademoyero, M. C. Bartholomew-Biggs, A. J. Davies, and S. C. Parkhurst. Conjugate gradient algorithms and the Galerkin boundary element method.

Computers and Mathematics with Applications, 48(3– 4):399–410, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210484063X.

Agarwal:2005:SSI

[ABCD05] R. P. Agarwal, N. S.

Barnett, P. Cerone, and S. S. Dragomir. A survey on some inequalities for expectation and variance. Computers and Mathematics with Applications, 49(2–3):429–480, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000441.

Al-Bahadili:2008:ACW

[ABH08] Hussein Al-Bahadili and Shakir M. Hussain. An adaptive character wordlength algorithm for data compression. Computers and Mathematics with Applications, 55(6):1250–1256, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005925.

Amer:2009:VST

[ABS09] Y. A. Amer, H. S. Bauomy, and M. Sayed. Vibration suppression in a twintail system to parametric and external excitations. Computers and Mathematics with Applications, 58(10):1947–1964, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006166.

Abuelmaatti:2001:NMF

[Abu01] M. T. Abuelma’atti. A new method for Fourier analysis of discontinuous and peaky periodic waveforms. Computers and Mathematics with Applications, 41(10–11):1417–1423, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001067.

Anastassiou:2000:NWA

[AC00] G. A. Anastassiou and S. Cambanis. Nonorthogonal wavelet approximation with rates of deterministic signals. Computers and Mathematics with Applications, 40(1):21– 35, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001371.

Aguilar:2002:HOC

[AC02] J. C. Aguilar and Y. Chen.

High-order corrected trapezoidal quadrature rules for functions with a logarithmic singularity in 2-D. Computers and Mathematics with Applications, 44 (8–9):1031–1039, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002122.

Atici:2003:EUR

[AC03] F. M. Atici and A. Cabada.

Existence and uniqueness results for discrete secondorder periodic boundary value problems. Computers and Mathematics with Applications, 45(6–9):1417– 1427, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210300097X.

Aguilar:2004:HOF

[AC04] J. C. Aguilar and Yu Chen.

A high-order, fast algorithm for scattering calculation in two dimensions. Computers and Mathematics with Applications, 47(1):1–11, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900016.

Aguilar:2005:HOC

[AC05] J. C. Aguilar and Yu Chen. High-order corrected trapezoidal quadrature rules for the Coulomb potential in three dimensions. Computers and Mathematics with Applications, 49(4):625–631, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000556.

Amirdjanova:2006:NMO

[AC06] A. Amirdjanova and S. Chivoret. New method for optimal nonlinear filtering of noisy observations by multiple stochastic fractional integral expansions. Computers and Mathematics with Applications, 52(1–2):161– 178, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210600188X.

Atay:2009:ESE

[AC09] Mehmet Tarik Atay and Safa Bozkurt Coskun. Elastic stability of Euler columns with a continuous elastic restraint using variational iteration method. Computers and Mathematics with Applications, 58(11–12):2528–2534, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001990.

Aloy:2007:CVC

[ACCMJ07] R. Aloy, M. C. Casabán, L. A. Caudillo-Mata, and L. Jódar. Computing the variable coefficient telegraph equation using a discrete eigenfunctions method. Computers and Mathematics with Applications, 54(3):448–458, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002283.

Abdulla:2001:MPN

[ACD01] T. J. Abdulla, J. R. Cash, and M. T. Diamantakis. An MEBDF package for the numerical solution of large sparse systems of stiff initial value problems. Computers and Mathematics with Applications, 42(1–2):121– 129, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001377.

Avery:2001:TSB

[ACH01] R. I. Avery, Chuan Jen Chyan, and J. Henderson. Twin solutions of boundary value problems for ordinary differential equations and finite difference equations. Computers and Mathematics with Applications, 42(3–5):695–704, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001882.

Aloy:2007:CUT

[ACJ07] R. Aloy, M. C. Casabán, and L. Jódar. Constructing unconditionally timestable numerical solutions for mixed parabolic problems. Computers and Mathematics with Applications, 53(11):1773–1783, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001526.

Allegretto:2006:NAT

[ACLL06] W. Allegretto, G. Cao, G. Li, and Y. Lin. Numerical analysis of tumor model in steady state. Computers and Mathematics with Applications, 52(5):593–606, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002471.

Alves:2005:PSQ

[ACM05] C. O. Alves, F. J. S. A.

Corrêa, and T. F. Ma. Positive solutions for a quasilinear elliptic equation of Kirchhoff type. Computers and Mathematics with Applications, 49(1):85–93, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500009X.

Abba:2003:ASS

[ACV03] A. Abbà, A. C. Cercignani, and L. Valdettaro. Analysis of subgrid scale models. Computers and Mathematics with Applications, 46(4):521–535, August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900149.

Amiraliyev:2003:UCD

[AD03] G. M. Amiraliyev and H. Duru. A uniformly convergent difference method for the periodical boundary value problem. Computers and Mathematics with Applications, 46(5–6):695–703, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901350.

Anastassiou:2008:SFA

[AD08] George A. Anastassiou and Oktay Duman. Statistical fuzzy approximation by fuzzy positive linear operators. Computers and Mathematics with Applications, 55(3):573–580, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107004920.

Aouf:2009:SAF

[AD09] M. K. Aouf and H. E. Darwish. Subclasses of analytic functions associated with the generalized hypergeometric function. Computers and Mathematics with Applications, 57(3):356–366, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006858.

Adamec:2004:NMS

[Ada04]  L. Adamec.  A note

on the matrix solution of the problem X^{Δ} = AX, X(t_{0}) = I. Computers and Mathematics with Applications, 48(10– 11):1485–1489, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004031.

Ackleh:2004:NMN

[ADDL04] A. S. Ackleh, K. Deng, J. Derouen, and Wei Li. A numerical method for a nonlocal hyperbolic model arising from a reliability system. Computers and Mathematics with Applications, 47(1):135–147, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900120.

Ashino:2003:STF

[ADH^{+}03] R. Ashino, S. J. Desjardins, C. Heil, M. Nagase, and R. Vaillancourt. Smooth tight frame wavelets and image microanalyis in the Fourier domain. Computers and Mathematics with Applications, 45(10–11):1551–1579, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001366.

Angelova:2002:OCI

[ADN02] J. Angelova, A. Dishliev, and S. Nenov. Ioptimal curve for impulsive Lotka–Volterra predator– prey model. Computers and Mathematics with Applications, 43(10–11):1203–1218, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102000925.

Amster:2008:NTC

[ADP08] Pablo Amster, Pablo De Nápoli, and Juan Pablo Pinasco. On Nirenbergtype conditions for higherorder systems on time scales. Computers and Mathematics with Applications, 55(12):2762–2766, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007766.

Amiraliyev:2007:UNM

[AE07] G. M. Amiraliyev and F. Erdogan. Uniform numerical method for singularly perturbed delay differential equations. Computers and Mathematics with Applications, 53(8):1251– 1259, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000831.

Ankudinova:2008:NSN

[AE08] Julia Ankudinova and Matthias Ehrhardt. On the numerical solution of nonlinear Black– Scholes equations. Computers and Mathematics with Applications, 56(3):799–812, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000783.

Altin:2006:LSL

[AEÇ06] Y. Altin, M. Et, and R. Çolak. Lacunary statistical and lacunary strongly convergence of generalized difference sequences of fuzzy numbers. Computers and Mathematics with Applications, 52(6– 7):1011–1020, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002744.

Abassy:2007:ESS

[AETEZ07a] Tamer A. Abassy, Magdy A.

El-Tawil, and H. El-Zoheiry. Exact solutions of some nonlinear partial differential equations using the variational iteration method linked with Laplace transforms and the Padé technique. Computers and Mathematics with Applications, 54(7–8):940–954, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002568.

Abassy:2007:MVI

[AETEZ07b] Tamer A. Abassy, Magdy A. El-Tawil, and H. El-Zoheiry. Modified variational iteration method for Boussinesq equation. Computers and Mathematics with Applications, 54(7–8):955– 965, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700257X.

Arato:2002:NSI

[AF02] M. Arató and S. Fegyverneki. New statistical investigations of the Ornstein–Uhlenbeck process. Computers and Mathematics with Applications, 44(5–6):677–692, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001827.

Ali:2009:RAI

[AF09] Jamshaid Ali and Jiancheng Fang. Realization of an autonomous integrated suite of strapdown astroinertial navigation systems using unscented particle filtering. Computers and Mathematics with Applications, 57(2):169–183, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006056.

Arai:2006:ECC

[AFI06] Masayuki Arai, Satoshi Fukumoto, and Kazuhiko Iwasaki. Extension of coefficients for (n, k, m) convolutional-code-based packet loss recovery. Computers and Mathematics with Applications, 51(2):247–256, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004621.

Ali:2009:SNO

[AFL^{+}09] M. Irfan Ali, Feng Feng, Xiaoyan Liu, Won Keun Min, and M. Shabir. On some new operations in soft set theory. Computers and Mathematics with Applications, 57(9):1547– 1553, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006871.

Agarwal:2001:OCT

[AG01a] R. P. Agarwal and S. R. Grace. Oscillation of certain third-order difference equations. Computers and Mathematics with Applications, 42(3–5):379–384, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001626.

Anastassiou:2001:PSPb

[AG01b] G. A. Anastassiou and S. G. Gal. Partial shape preserving approximation by bivariate Hermite–Fejér polynomials. Computers and Mathematics with Applications, 42(1–2):57– 64, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001304.

Anastassiou:2001:PSPa

[AG01c] G. A. Anastassiou and S. G. Gal. Partial shape preserving approximations by bivariate Shepard operators. Computers and Mathematics with Applications, 42(1–2):47– 56, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001298.

Antohe:2003:PGI

[AG03] V. Antohe and I. Gladwell. Performance of Gauss implicit Runge– Kutta methods on separable Hamiltonian systems. Computers and Mathematics with Applications, 45 (1–3):481–501, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800329.

Anastassiou:2009:ACB

[AG09] George A. Anastassiou and Sorin G. Gal. Approximation by complex Bernstein– Schurer and Kantorovich– Schurer polynomials in compact disks. Computers and Mathematics with Applications, 58(4):734–743, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003058.

Agarwal:2003:DSH

[AGO03a] R. P. Agarwal, S. R. Grace, and D. O’Regan. Discrete semipositone higher-order equations. Computers and Mathematics with Applications, 45(6–9):1171–1179, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000798.

Agarwal:2003:NSD

[AGO03b] R. P. Agarwal, S. R. Grace, and D. O’Regan. Nonoscillatory solutions for discrete equations. Computers and Mathematics with Applications, 45(6–9):1297–1302, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001019.

Agratini:2008:ICS

[Agr08]  Octavian Agratini.  On

the iterates of a class of summation-type linear positive operators. Computers and Mathematics with Applications, 55(6):1178– 1180, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005858.

AlBdaiwi:2009:EIS

[AGS09] Bader F. AlBdaiwi, Boris Goldengorin, and Gerard Sierksma. Equivalent instances of the simple plant location problem. Computers and Mathematics with Applications, 57(5):812–820, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006421.

Abdel-Hamid:2000:ESS

[AH00]  B. Abdel-Hamid.  Ex-

act solutions of some nonlinear evolution equations using symbolic computations. Computers and Mathematics with Applications, 40(2–3):291–302, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001619.

Abdel-Hameed:2003:OMS

[AH03] M. Abdel-Hameed. Optimal maintenance of systems subject to deterioration of the renewal type. Computers and Mathematics with Applications, 46(7):987–992, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390112X.

Ali:2008:PFM

[AH08] Nasir Ali and Tasawar Hayat. Peristaltic flow of a micropolar fluid in an asymmetric channel. Computers and Mathematics with Applications, 55(4):589–608, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107004944.

Abbasbandy:2009:NAR

[AH09] S. Abbasbandy and T. Hajjari. A new approach for ranking of trapezoidal fuzzy numbers. Computers and Mathematics with Applications, 57(3):413–419, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006238.

Ayala:2000:OGL

[AHK00] V. Ayala, A. Hacibekiroğlu, and E. Kizil. Observability of general linear pairs. Computers and Mathematics with Applications, 39(1–2):35–43, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003119.

Aahlander:2002:ESM

[Åhl02]  K. Åhlander.  Einstein

summation for multidimensional arrays. Computers and Mathematics with Applications, 44(8– 9):1007–1017, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002109.

Aahlander:2003:STS

[Åhl03] K. Åhlander. Supporting tensor symmetries in Ein-Sum. Computers and Mathematics with Applications, 45(4–5):789–803, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000415.

Abdel-Hameed:2006:OCF

[AHN06] M. S. Abdel-Hameed and Y. A. Nakhi. Optimal control of a finite dam with diffusion input and state dependent release rates. Computers and Mathematics with Applications, 51(2):317–324, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004694.

Ashino:2001:MFM

[AHNV01] R. Ashino, C. Heil, M. Nagase, and R. Vaillancourt. Microlocal filtering with multiwavelets. Computers and Mathematics with Applications, 41(1–2):111–133, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101850112.

Aboul-Hassan:2009:PED

[AHRT09] Abdel-Karim Aboul-Hassan, Sherif I. Rabia, and Fowzia A. Taboly. Performance evaluation of a discrete-time Geo^{[X]} /G/1 retrial queue with general retrial times. Computers and Mathematics with Applications, 58 (3):548–557, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109002661.

Aouf:2000:SFM

[AHS00] M. K. Aouf, H. M. Hossen, and H. M. Srivastava. Some families of multivalent functions. Computers and Mathematics with Applications, 39(7–8):39– 48, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000638.

Ahmad:2006:CGM

[AI06] R. Ahmad and S. S. Irfan. Completely generalized multivalued nonlinear coquasi-variational-like inclusions. Computers and Mathematics with Applications, 52(6–7):1183–1191, September/October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002872.

Altintas:2008:NCS

[AIS08] Osman Altintas, Hüseyin Irmak, and H. M. Srivastava. Neighborhoods for certain subclasses of multivalently analytic functions defined by using a differential operator. Computers and Mathematics with Applications, 55(3):331–338, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003409.

Almenar:2009:EBS

[AJ09a] Pedro Almenar and Lucas Jódar. Explicit bounds for the solutions of second order linear differential equations. Computers and Mathematics with Applications, 57(5):789–798, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004860.

Almenar:2009:IEB

[AJ09b] Pedro Almenar and Lucas Jódar. Improving explicit bounds for the solutions of second order linear differential equations. Computers and Mathematics with Applications, 57(10):1708– 1721, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002302.

Aulbach:2001:EFS

[AK01] B. Aulbach and J. Kalkbrenner. Exponential forward splitting for noninvertible difference equations. Computers and Mathematics with Applications, 42(3–5):743–754, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001948.

Ayala:2003:DTS

[AK03] V. Ayala and W. Kliemann. A decomposition theorem for singular control systems on lie groups. Computers and Mathematics with Applications, 45 (4–5):635–646, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000232.

Asadzadeh:2006:CSM

[AK06] M. Asadzadeh and A. Kadem. Chebyshev spectral-S _{N} method for the neutron transport equation. Computers and Mathematics with Applications, 52(3–4):509–524, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002161.

AlZhour:2007:SNC

[AK07] Zeyad Al Zhour and Adem Kiliçman. Some new connections between matrix products for partitioned and non-partitioned matrices. Computers and Mathematics with Applications, 54(6): 763–784, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107002672.

Agnolin:2008:EMT

[AK08a] I. Agnolin and N. P. Kruyt.

On the elastic moduli of two-dimensional assemblies of disks: Relevance and modeling of fluctuations in particle displacements and rotations. Computers and Mathematics with Applications, 55(2):245–256, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003513.

Anderson:2008:HOT

[AK08b] Douglas R. Anderson and Ilkay Yaslan Karaca. Higherorder three-point boundary value problem on time scales. Computers and Mathematics with Applications, 56(9):2429–2443, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003829.

Antal:2008:MIS

[AK08c] I. Antal and J. Karátson. A mesh independent superlinear algorithm for some nonlinear nonsymmetric elliptic systems. Computers and Mathematics with Applications, 55(10):2185– 2196, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007444.

Agarwal:2004:IOC

[AL04] R. P. Agarwal and Wan-Tong Li. Interval oscillation criteria for second-order nonlinear perturbed differential equations. Computers and Mathematics with Applications, 47(4–5):751– 765, February/March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104900624.

Amirdjanova:2008:SEE

[AL08a] Anna Amirdjanova and Matthew Linn. Stochastic evolution equations for nonlinear filtering of random fields in the presence of fractional Brownian sheet observation noise. Computers and Mathematics with Applications, 55(8):1766– 1784, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006505.

An:2008:VIR

[AL08b] Rong An and Kaitai Li.

Variational inequality for the rotating Navier–Stokes equations with subdifferential boundary conditions. Computers and Mathematics with Applications, 55 (3):581–587, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004932.

Alber:2002:PPM

[Alb02]  Ya. I. Alber.  Proxi-

mal projection methods for variational inequalities and Cesáro averaged approximations. Computers and Mathematics with Applications, 43(8–9):1107–1124, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800165.

Alekseev:2006:ABP

[Ale06]  A. K. Alekseev.  An

adjoint-based a posteriori estimation of iterative convergence error. Computers and Mathematics with Applications, 52 (8–9):1205–1212, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003142.

Allasia:2002:API

[All02] G. Allasia. Approximating potential integrals by cardinal basis interpolants on multivariate scattered data. Computers and Mathematics with Applications, 43(3–5):275–287, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002863.

Allahverdiev:2005:DDH

[All05] B. P. Allahverdiev. Dissipative discrete Hamiltonian systems. Computers and Mathematics with Applications, 49(7–8):1139–1155, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001288.

AlSharawi:2008:POP

[AlS08]  Ziyad AlSharawi.  Peri-

odic orbits in periodic discrete dynamics. Computers and Mathematics with Applications, 56(8):1966–1974, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002745.

Atencia:2004:DTG

[AM04] I. Atencia and P. Moreno.

Discrete-time Geo^{[X]} /G_{H} /1 retrial queue with Bernoulli feedback. Computers and Mathematics with Applications, 47(8–9): 1273–1294, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104901228.

Acosta:2008:SEM

[AM08a] Carlos D. Acosta and Carlos E. Mejía. Stabilization of explicit methods for convection diffusion equations by discrete mollification. Computers and Mathematics with Applications, 55(3):368–380, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003689.

Aminataei:2008:NSP

[AM08b] A. Aminataei and M. M. Mazarei. Numerical solution of Poisson’s equation using radial basis function networks on the polar coordinate. Computers and Mathematics with Applications, 56(11):2887–2895, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004616.

Aouf:2008:SVP

[AM08c] M. K. Aouf and A. O. Mostafa. On a subclass of n–p-valent prestarlike functions. Computers and Mathematics with Applications, 55(4):851–861, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005160.

Ahmed:2009:NNA

[AM09] S. G. Ahmed and S. A. Meshrif. A new numerical algorithm for 2D moving boundary problems using a boundary element method. Computers and Mathematics with Applications, 58(7):1302–1308, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004556.

Arenas:2008:NSN

[AMC08] Abraham J. Arenas, José Antonio Moraño, and Juan Carlos Cortés. Non-standard numerical method for a mathematical model of RSV epidemiological transmission. Computers and Mathematics with Applications, 56(3):670–678, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000576.

Agarwal:2002:OSS

[AML02] R. P. Agarwal, Fan Wei Meng, and Wei Nian Li. Oscillation of solutions of systems of neutral type partial functional differential equations. Computers and Mathematics with Applications, 44(5–6):777–786, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001906.

Abu-Mostafa:2000:MCL

[AMM00] Y. S. Abu-Mostafa and R. J. McEliece. Maximal codeword lengths in Huffman codes. Computers and Mathematics with Applications, 39(11):129– 134, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210000119X.

Ammar:2007:ONF

[Amm07] E. E. Ammar. On the optimality of nonlinear fractional disjunctive programming problems. Computers and Mathematics with Applications, 53(10):1527– 1537, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001216.

Azhar:2005:CEN

[AMR05] S. Azhar, A. McLennan, and J. H. Reif. Computation of equilibria in noncooperative games. Computers and Mathematics with Applications, 50(5–6):823–854, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003202.

Aboulaich:2008:NDM

[AMS08] R. Aboulaich, D. Meskine, and A. Souissi. New diffusion models in image processing. Computers and Mathematics with Applications, 56(4):874–882, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000886.

Al-Musallam:2003:FIW

[AMT03] F. Al-Musallam and V. K.

Tuan. A finite and an infinite Whittaker integral transform. Computers and Mathematics with Applications, 46(12):1847–1859, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902410.

Arbenz:2008:PMP

[AMV08] P. Arbenz, S. Margenov, and Y. Vutov. Parallel MIC(0) preconditioning of 3D elliptic problems discretized by Rannacher– Turek finite elements. Computers and Mathematics with Applications, 55(10): 2197–2211, May 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107007432.

Aleman-Meza:2009:KSC

[AMYS^{+}09] Boanerges Aleman-Meza, Yihai Yu, Heinz-Bernd Schüttler, Jonathan Arnold, and Thiab R. Taha. KIN-SOLVER: a simulator for computing large ensembles of biochemical and gene regulatory networks. Computers and Mathematics with Applications, 57(3):420–435, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006287.

An:2003:LCR

[An03] A. An. Learning classification rules from data. Computers and Mathematics with Applications, 45 (4–5):737–748, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000348.

Alekseev:2005:ACB

[AN05] A. K. Alekseev and I. M. Navon. Adjoint correction and bounding of error using Lagrange form of truncation term. Computers and Mathematics with Applications, 50(8– 9):1311–1332, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003834.

Alekseev:2006:PEE

[AN06] A. K. Alekseev and I. M.

Navon. A posteriori error estimation by postprocessor independent of method of flowfield calculation. Computers and Mathematics with Applications, 51(3–4):397–404, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005080.

Ahmad:2009:ERC

[AN09] Bashir Ahmad and Juan J.

Nieto. Existence results for a coupled system of nonlinear fractional differential equations with three-point boundary conditions. Computers and Mathematics with Applications, 58(9):1838–1843, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005975.

Anastassiou:2000:RCS

[Ana00a] G. A. Anastassiou. Rate of convergence of some multivariate neural network operators to the unit. Computers and Mathematics with Applications, 40(1):1– 19, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210000136X.

Anastassiou:2000:P

[Ana00b] George A. Anastassiou. Preface. Computers and Mathematics with Applications, 40(1):xiii, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100001358.

Anastassiou:2004:FAF

[Ana04a] G. A. Anastassiou. Fuzzy approximation by fuzzy convolution type operators. Computers and Mathematics with Applications, 48(9):1369–1386, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003815.

Anastassiou:2004:HOA

[Ana04b] G. A. Anastassiou. Highorder approximation by multivariate shift-invariant convolution type operators. Computers and Mathematics with Applications, 48(9):1245–1261, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003736.

Anastassiou:2004:HOF

[Ana04c] G. A. Anastassiou. Highorder fuzzy approximation by fuzzy-wavelet type and neural network operators. Computers and Mathematics with Applications, 48(9):1387–1401, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003827.

Anastassiou:2004:OTI

[Ana04d] G. A. Anastassiou. Opial type inequalities involving fractional derivatives of two functions and applications. Computers and Mathematics with Applications, 48 (10–11):1701–1731, November/December 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004237.

Anastassiou:2004:UFR

[Ana04e] G. A. Anastassiou. Univariate fuzzy-random neural network approximation operators. Computers and Mathematics with Applications, 48(9):1263–1283, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003748.

Anastassiou:2004:P

[Ana04f] George A. Anastassiou. Preface. Computers and Mathematics with Applications, 48(9):xiii, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104003724.

Anastassiou:2005:GSU

[Ana05] G. A. Anastassiou. Global smoothness and uniform convergence of smooth Picard singular operators. Computers and Mathematics with Applications, 50 (10–12):1755–1766, November/December 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122105004220.

Anastassiou:2006:AGM

[Ana06] G. A. Anastassiou. Applications of geometric moment theory related to optimal portfolio management. Computers and Mathematics with Applications, 51(9–10):1405–1430, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000885.

Anastassiou:2007:MFO

[Ana07a] George A. Anastassiou.

Multivariate fractional Ostrowski type inequalities. Computers and Mathematics with Applications, 54 (3):434–447, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210700226X.

Anastassiou:2007:TWR

[Ana07b] George A. Anastassiou.

Taylor–Widder representation formulae and Ostrowski, Grüss, integral means and Csiszar type inequalities. Computers and Mathematics with Applications, 54(1):9–23, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001678.

Anastassiou:2008:PST

[Ana08] George A. Anastassiou. Poincaré and Sobolev type inequalities for vectorvalued functions. Computers and Mathematics with Applications, 56(4):1102– 1113, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001119.

Anderson:2003:DTO

[And03]  D. R. Anderson.  Dis-

crete third-order three-point right-focal boundary value problems. Computers and Mathematics with Applications, 45(6–9):861–871, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103801578.

Anonymous:2000:AIa

[Ano00a] Anonymous. Author index. Computers and Mathematics with Applications, 39 (12):xi–xii, June 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100001346.

Anonymous:2000:AIb

[Ano00b] Anonymous. Author index. Computers and Mathematics with Applications, 40(12):X–XI, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210180002X.

Anonymous:2000:BR

[Ano00c]  Anonymous.  Book re-

ports. Computers and Mathematics with Applications, 39(9–10):241–248, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000985.

Anonymous:2000:BRAe

[Ano00d] Anonymous. Book review:

Advanced database indexing: By Yannis Manolopoulos, Yannis Theodoridis and Vassilis J. Tsotras. Kluwer Academic, Boston. (2000). 286 pages. $132, NLG 310, GBP 91.25. Computers and Mathematics with Applications, 40(2–3):416, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901755.

Anonymous:2000:BRAa

[Ano00e] Anonymous. Book review: Advances in automatic text summarization: Edited by Inderjeet Mani and Mark T. Maybury. MIT Press, Cambridge, MA. (1999). 434 pages. $45.00. Computers and Mathematics with Applications, 39(3–4):261, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900178.

Anonymous:2000:BRAc

[Ano00f] Anonymous. Book review: Advances in automatic text summarization: Edited by Inderjeet Mani and Mark T. Maybury. MIT Press, Cambridge, MA. (1999). 434 pages. $45.00. Computers and Mathematics with Applications, 39(7–8): 264, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902694.

Anonymous:2000:BREa

[Ano00g] Anonymous. Book review: An exploration of chaos: an introduction for natural scientists and engineers: By John Argyris, Gunter Faust and Maria Haase. North-Holland, Amsterdam. (1994). 751 pages. $193.00.

Computers and Mathematics with Applications, 39 (3–4):262, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210090035X.

Anonymous:2000:BREc

[Ano00h] Anonymous. Book review:

An exploration of chaos: an introduction for natural scientists and engineers: By John Argyris, Gunter Faust and Maria Haase. North-Holland, Amsterdam. (1994). 751 pages. $193.00. Computers and Mathematics with Applications, 39(7– 8):265, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210090289X.

Anonymous:2000:BRAf

[Ano00i]  Anonymous.  Book re-

view: Applied operating system concepts: First edition. By Avi Silberschatz, Peter Galvin and Greg Gagne. John Wiley & Sons, New York. (2000). 840 pages. $80.95. Computers and Mathematics with Applications, 40(2–3):420, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100902190.

Anonymous:2000:BRAb

[Ano00j]  Anonymous.  Book re-

view: Architects of the information society: Thirtyfive years of the laboratory for computer science at MIT: By Simson L. Garfinkel, edited by Hal Abelson. MIT Press, Cambridge, MA. (1999). 72 pages. $20.00. Computers and Mathematics with Applications, 39(3–4):263, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100900415.

Anonymous:2000:BRAd

[Ano00k]  Anonymous.  Book re-

view: Architects of the information society: Thirtyfive years of the laboratory for computer science at MIT: By Simson L. Garfinkel, edited by Hal Abelson. MIT Press, Cambridge, MA. (1999). 72 pages. $20.00. Computers and Mathematics with Applications, 39(7–8):264– 265, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100902797.

Anonymous:2000:BRCa

[Ano00l]  Anonymous.  Book re-

view: Cambrian intelligence: the early history of the new AI: By Rodney A. Brooks. MIT Press, Cambridge, MA. (1999). 199 pages. $60.00 (cloth); $25.00 (paper). Computers and Mathematics with Applications, 39(3–4):261, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900257.

Anonymous:2000:BRCe

[Ano00m] Anonymous. Book review:

Cambrian intelligence: the early history of the new Al: By Rodney A. Brooks. MIT Press, Cambridge, MA. (1999). 199 pages. $60.00 (cloth); $25.00. Computers and Mathematics with Applications, 39(7–8): 265, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902852.

Anonymous:2000:BRCk

[Ano00n]  Anonymous.  Book re-

view: Comparing financial systems: By Franklin Allen and Douglas Gale. The MIT Press, Cambridge, MA. (2000). 507 pages. $55. Computers and Mathematics with Applications, 40(2–3):419, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210090213X.

Anonymous:2000:BRCi

[Ano00o] Anonymous. Book review: Complexity and approximation: Combinatorial optimization problems and their approximability properties: By G. Ausiello, P. Crescenzi, G. Gambosi, V. Kann, A. Marchetti-Spaccamela and M. Protasi. Springer-Verlag, Berlin. (1999). 524 pages. $59.95. Computers and Mathematics with Applications, 40(2–3):417, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901834.

Anonymous:2000:BRCj

[Ano00p] Anonymous. Book review: Complexity: Life at the edge of chaos: Second edition. By Roger Lewin. The University of Chicago Press, Chicago. (1999). 234 pages. $28, £20 (cloth); $14, £10 (paper). Computers and Mathematics with Applications, 40(2–3):417, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901895.

Anonymous:2000:BRCc

[Ano00q]  Anonymous.  Book re-

view: Computational logic: Edited by Ulrich Berger and Helmut Schwichtenberg. Springer-Verlag, Berlin. (1999). 444 pages. $119.00, DM 179.00, öS 1307.00, sFr 162.00, GBP 69.00. Computers and Mathematics with Applications, 39 (3–4):263, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100900452.

Anonymous:2000:BRCf

[Ano00r]  Anonymous.  Book re-

view: Computational logic: Edited by Ulrich Berger and Helmut Schwichtenberg. Springer-Verlag, Berlin. (1999). 444 pages. $119.00, DM 179.00, öS 1307.00, sFr 162.00, GBP 69.00. Computers and Mathematics with Applications, 39(7– 8):265, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902876.

Anonymous:2000:BRCd

[Ano00s] Anonymous. Book review:

Computer arithmetic: Algorithms and hardware designs: By Behrooz Parhami. Oxford University Press, New York. (2000). 490 pages. $85.00. Computers and Mathematics with Applications, 39(3–4):264, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900518.

Anonymous:2000:BRCg

[Ano00t]  Anonymous.  Book re-

view: Computer arithmetic: Algorithms and hardware designs: By Behrooz Parhami. Oxford University Press, New York. (2000). 490 pages. $85.00. Computers and Mathematics with Applications, 39(7–8): 266, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902955.

Anonymous:2000:BRCh

[Ano00u]  Anonymous.  Book re-

view: Convex and starlike mappings in several complex variables: By Sheng Gong. Science Press, Beijing; Kluwer Academic, Dordrecht. (1998). 199 pages. $89, NLG 165, GBP 56. Computers and Mathematics with Applications, 40(2–3):415, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210090169X.

Anonymous:2000:BRDe

[Ano00v]  Anonymous.  Book re-

view: Database nation: the death of privacy in the 21-st century: By Simson Garfinkel. O’Reilly, Sebastopol, CA. (2000). 312 pages. $24.95. Computers and Mathematics with Applications, 40(2–3):417, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901810.

Anonymous:2000:BRDd

[Ano00w] Anonymous. Book review:

Dependable network computing: Edited by Dimiter R. Avresky. Kluwer Academic, Boston. (2000). 462 pages. $165, NLG 385, GBP 114. Computers and Mathematics with Applications, 40(2–3):415, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901676.

Anonymous:2000:BRDb

[Ano00x]  Anonymous.  Book re-

view: Design by composition for rapid prototyping: By Michael Binnard. Kluwer Academic, Boston. (1999). 139 pages. $105, NLG 245, GBP 73. Computers and Mathematics with Applications, 40(2–3):414, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100901639.

Anonymous:2000:BRDc

[Ano00y] Anonymous. Book review: Deterministic global optimization: Theory, methods and applications: By Christodoulos A. Floudas. Kluwer Academic, Dordrecht. (2000). 739 pages. $315, NLG 595, GBP

196. Computers and Mathematics with Applications, 40(2–3):415, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100901652.

Anonymous:2000:BRDa

[Ano00z] Anonymous. Book review: Differential equations with MATLAB: By Kevin R. Coombes, Brian R. Hunt, Ronald L. Lipsman, John E. Osborn and Garrett J. Stuck. John Wiley & Sons, New York. (2000). 256 pages. $19.99. Computers and Mathematics with Applications, 39(3–4):262, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900336.

Anonymous:2000:BREe

[Ano00-27]  Anonymous.  Book re-

view: Ethernet: the definitive guide: By Charles E. Spurgeon. O’Reilly, Sebastapol, CA. (2000). 498 pages. $44.95. Computers and Mathematics with Applications, 40(2–3):419, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902153.

Anonymous:2000:BREd

[Ano00-28]  Anonymous.  Book re-

view: Experimentation in software engineering: an introduction: By Claes Wohlin, Per Runeson, Martin Höst, Magnus C. Ohlsson, Björn Regnell and Anders Wesslén. Kluwer Academic, Boston. (2000). 204 pages. $120, NLG 280, GBP 83. Computers and Mathematics with Applications, 40(2–3):418, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100902037.

Anonymous:2000:BREb

[Ano00-29]  Anonymous.  Book re-

view: Exploring chaos: Theory and experiment: By Brian Davies. Perseus Books, reading, MA. (1999). 237 pages. $45.00. Computers and Mathematics with Applications, 39(3–4):265, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100900634.

Anonymous:2000:BRFb

[Ano00-30]  Anonymous.  Book re-

view: Forecasting nonstationary economic time series: By Michael P. Clements and David F. Hendry. MIT Press, Cambridge, MA. (1999). 362 pages. $35.00. Computers and Mathematics with Applications, 39(3–4):266, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100900853.

Anonymous:2000:BRFd

[Ano00-31]  Anonymous.  Book re-

view: Fundamental problems of algorithmic algebra: By Chee Keng Yap. Oxford University Press, New York. (2000). 511 pages. $72p. Computers and Mathematics with Applications, 40(2–3):414, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901615.

Anonymous:2000:BRGa

[Ano00-32]  Anonymous.  Book re-

view: Geometry from Africa: Mathematical and educational explorations: By Paulus Gerdes. The Mathematical Association of America, Washington, DC. (1999). 210 pages. $39.95. Computers and Mathematics with Applications, 39 (3–4):264, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100900592.

Anonymous:2000:BRGb

[Ano00-33]  Anonymous.  Book re-

view: Geometry from Africa: Mathematical and educational explorations: By Paulus Gerdes. The Mathematical Association of America, Washington, DC. (1999). 210 pages. $39.95. Computers and Mathematics with Applications, 39(7– 8):264, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100902773.

Anonymous:2000:BRIb

[Ano00-34]  Anonymous.  Book re-

view: Internet core protocols: the definitive guide: By Eric A. Hall. O’Reilly, Sebastapol, CA. (2000). 449 pages. $39.95 (CD-ROM included). Computers and Mathematics with Applications, 40(2–3):419, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100902050.

Anonymous:2000:BRIa

[Ano00-35]  Anonymous.  Book re-

view: Inverse problems: Activities for undergraduates: By C. W. Groetsch. The Mathematical Association of America, Washington, DC. (1999). 222 pages. $26.00. Computers and Mathematics with Applications, 39(3–4):265, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100900658.

Anonymous:2000:BRJa

[Ano00-36]  Anonymous.  Book re-

view: Java enterprise in a nutshell: a desktop quick reference: By David Flanagan, Jim Farlye, William Crawford and Kris Magnusson. O’Reilly, Sebastopol, CA. (1999). 604 pages. $29.95. Computers and Mathematics with Applications, 39(3–4):263, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900373.

Anonymous:2000:BRJb

[Ano00-37]  Anonymous.  Book re-

view: JavaScript application cookbook: By Jerry Bradenbaugh. O’Reilly, Sebastopol, CA. (1999). 462 pages. $34.95. Computers and Mathematics with Applications, 39(3–4):266, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900798.

Anonymous:2000:BRK

[Ano00-38]  Anonymous.  Book re-

view: Knowledge engineering and management: the commonKADS methodology: By Guus Schreiber, Hans Akkermans, Anjo Anjewierden, Robert de Hoog, Nigel Shadbolt, Walter Van de Velde, and Bob Wielinga. The MIT Press, Cambridge, MA. (2000). 455 pages. $50. Computers and Mathematics with Applications, 40(2–3):418, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901998.

Anonymous:2000:BRLa

[Ano00-39] Anonymous. Book review: Learning Debian GNU/ Linux: By Bill McCarty. O’Reilly, Sebastopol, CA. (1999). 343 pages. $34.95 (CD-ROM included). Computers and Mathematics with Applications, 39(3–4): 265, February 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100900713.

Anonymous:2000:BRLb

[Ano00-40] Anonymous. Book review: Learning Red Hat Linux: By Bill McCarty. O’Reilly, Sebastopol, CA. (1999). 378 pages. $34.95 (CD-ROM included). Computers and Mathematics with Applications, 39(3–4):265, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900695.

Anonymous:2000:BRLe

[Ano00-41]  Anonymous.  Book re-

view: LEDA: a platform for combinatorial and geometric computing: By Kurt Mehlhorn and Stefan Näher. Cambridge University Press, Cambridge, U.K. (1999). 1018 pages. $80.00. Computers and Mathematics with Applications, 40(2–3):420, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902177.

Anonymous:2000:BRLc

[Ano00-42]  Anonymous.  Book re-

view: Line form color: By Ellsworth Kelly. University of Chicago Press, Chicago. (1999). Includes “An Intense Detachment: Ellsworth Kelly’s Line Form Color”, an essay by Harry Cooper. $25.00, £17.00. Computers and Mathematics with Applications, 39(3–4):266, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210090083X.

Anonymous:2000:BRLd

[Ano00-43] Anonymous. Book review:

Logic programming: Proceedings of the 1999 international conference on logic programming: Edited by Danny DeSchreye. The MIT Press, Cambridge, MA. (1999). 622 pages. $85. Computers and Mathematics with Applications, 40(2–3):414, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901597.

Anonymous:2000:BRMa

[Ano00-44]  Anonymous.  Book re-

view: Managing Microsoft exchange server: By Paul Robichaux. O’Reilly, Sebastopol, CA. (1999). 701 pages. $34.95. Computers and Mathematics with Applications, 39(3–4):263, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900439.

Anonymous:2000:BRMc

[Ano00-45]  Anonymous.  Book re-

view: Managing Microsoft exchange server: By Paul Robichaux. O’Reilly, Sebastopol, CA. (1999). 701 pages. $34.95. Computers and Mathematics with Applications, 39(7–8):265, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902839.

Anonymous:2000:BRMg

[Ano00-46] Anonymous. Book review:

Mathematica(R) in action: Second Edition. By Stan Wagon. Springer-Verlag, New York. (2000). 592 pages. $49.95, DM 98, öS 716, sFr 89.50, GBP 34 (CD-ROM included). Computers and Mathematics with Applications, 40(2–3):420, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902256.

Anonymous:2000:BRMe

[Ano00-47]  Anonymous.  Book re-

view: Mathematics behind fuzzy logic: By Esko Turunen. Physica-Verlag, Heidelberg. (1999). 191 pages. $44, DM 78, öS 570, sFr 71, GBP 30. Computers and Mathematics with Applications, 40(2–3):416, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901731.

Anonymous:2000:BRMb

[Ano00-48] Anonymous. Book review:

Microtransducer CAD: Physical and computational aspects: By Arokia Nathan and Henry Baltes. Springer-Verlag, Vienna. (1999). 427 pages. $189.00, öS 2086.00, DM 298.00, sFr 269.00. Computers and Mathematics with Applications, 39 (3–4):264, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210090049X.

Anonymous:2000:BRMd

[Ano00-49] Anonymous. Book review: Microtransducer CAD: Physical and computational aspects: By Arokia Nathan and Henry Baltes. Springer-Verlag, Vienna. (1999). 427 pages. $189.00, öS 2086.00, DM 298.00, sFr 269.00. Computers and Mathematics with Applications, 39(7– 8):266, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902931.

Anonymous:2000:BRMh

[Ano00-50] Anonymous. Book review: Mobile data management and applications: Edited by Jin Jing and Anupam Joshi. Kluwer Academic, Boston. (1999). 118 pages. $110, NLG 255, GBP 76. Computers and Mathematics with Applications, 40(2–3):420, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210090227X.

Anonymous:2000:BRMf

[Ano00-51]  Anonymous.  Book re-

view: Models for repeated measurements: Second edition. By J. K. Lindsey. Oxford University Press, New York. (1999). 515 pages. $75. Computers and Mathematics with Applications, 40(2–3):418, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901950.

Anonymous:2000:BRNa

[Ano00-52]  Anonymous.  Book re-

view: Networking essentials: Flashcards for MCSE core exam 70-058: By Michael Moncur. O’Reilly, Sebastopol, CA. (1999). $29.95. Computers and Mathematics with Applications, 39(3–4):261, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900233.

Anonymous:2000:BRNc

[Ano00-53]  Anonymous.  Book re-

view: Networking essentials: Flashcards for MCSE core exam 70-058: By Michael Moncur. O’Reilly, Sebastopol, CA. $29.95. Computers and Mathematics with Applications, 39(7– 8):264, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210090275X.

Anonymous:2000:BRNb

[Ano00-54] Anonymous. Book review: Neural codes and distributed representations: Foundations of neural computation: Edited by Laurence Abbott and Terrence J. Sejnowski. MIT Press, Cambridge, MA. (1999). 345 pages. $30.00. Computers and Mathematics with Applications, 39(3–4):262, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100900294.

Anonymous:2000:BROb

[Ano00-55] Anonymous. Book review: Object-oriented technology and computing systems reengineering: Edited by H. Zedan and A. Cau. Horwood Publishing, Chichester. (2000). 197 pages.

$49.95. Computers and Mathematics with Applications, 40(2–3):417, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901913.

Anonymous:2000:BROc

[Ano00-56] Anonymous. Book review:

Operating systems and services: Edited by R. Rajkumar. Kluwer Academic, Boston. (1999). 197 pages. $104.95, NLG 215, GBP 68.25. Computers and Mathematics with Applications, 40(2–3):418, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901937.

Anonymous:2000:BROa

[Ano00-57]  Anonymous.  Book re-

view: Optimizing windows for games, graphics, and multimedia: By David L. Farquhar. O’Reilly, Sebastapol, CA. (2000). 278 pages. $24.95. Computers and Mathematics with Applications, 40(2–3):415, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901718.

Anonymous:2000:BRPe

[Ano00-58]  Anonymous.  Book re-

view: Practical use of Math-CAD(R): Solving mathematical problems with a computer algebra system: By Hans Benker, translated by Anthony Rudd. Springer-Verlag, Berlin. (1999). 505 pages. $54. Computers and Mathematics with Applications, 40(2–3):417, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100901871.

Anonymous:2000:BRPh

[Ano00-59] Anonymous. Book review: Principles of applied mathematics: Transformation and approximation: Revised edition. By James P. Keener. Perseus Books, Cambridge, MA. (2000). 603 pages. $60.00. Computers and Mathematics with Applications, 40(2–3):420, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100902219.

Anonymous:2000:BRPa

[Ano00-60]  Anonymous.  Book re-

view: Programming Internet email: By David Wood. O’Reilly, Sebastopol, CA. (1999). 362 pages. $34.95.

Computers and Mathematics with Applications, 39 (3–4):264, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100900555.

Anonymous:2000:BRPc

[Ano00-61]  Anonymous.  Book re-

view: Programming Internet email: By David Wood. O’Reilly, Sebastopol, CA. (1999). 362 pages. $34.95. Computers and Mathematics with Applications, 39(7– 8):266, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902992.

Anonymous:2000:BRPg

[Ano00-62] Anonymous. Book review:

Programming the perl DBI: By Alligator Descartes and Tim Bunce. O’Reilly, Sebastapol, CA. (2000). 346 pages. $34.95. Computers and Mathematics with Applications, 40(2–3):420, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902232.

Anonymous:2000:BRPf

[Ano00-63]  Anonymous.  Book re-

view: Python programming on Win32: By Mark Hammond and Andy Robinson. O’Reilly, Sebastopol, CA. (2000). 652 pages. $34.95. Computers and Mathematics with Applications, 40(2–3):418, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901974.

Anonymous:2000:BRQ

[Ano00-64] Anonymous. Book review: Quantum dialogue: the making of a revolution: By Mara Beller. University of Chicago Press, Chicago. (1999). 365 pages. $35.00, £24.50 (cloth). Computers and Mathematics with Applications, 39(3–4):266, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900816.

Anonymous:2000:BRRc

[Ano00-65] Anonymous. Book review: REALbasic: the definitive guide: By Matt Neuburg. O’Reilly, Sebastopol, CA. (1999). $29.95. Computers and Mathematics with Applications, 39(3–4):266, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900774.

Anonymous:2000:BRRb

[Ano00-66]  Anonymous.  Book re-

view: Revolutions in differential equations: Exploring ODEs with modern technology: Edited by Michael J. Kallaher. The Mathematical Association of America, Washington, DC. (1999). 89 pages. $18.75. Computers and Mathematics with Applications, 39(3–4):264, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900579.

Anonymous:2000:BRRa

[Ano00-67]  Anonymous.  Book re-

view: Running Linux, third edition: By Matt Welsh, Matthias Kalle Dalheimer and Lar Kaufman. O’Reilly, Sebastopol, CA. (1999). 730 pages. $34.95. Computers and Mathematics with Applications, 39(3–4):263, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900476.

Anonymous:2000:BRRd

[Ano00-68]  Anonymous.  Book re-

view: Running Linux, third edition: By Matt Welsh, Matthias Kalle Dalheimer and Lar Kaufman. O’Reilly, Sebastopol, CA. (1999). 730 pages. $34.95. Computers and Mathematics with Applications, 39(7–8):265– 266, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902918.

Anonymous:2000:BRSc

[Ano00-69]  Anonymous.  Book re-

view: Simply scheme: Introducing computer science, Second edition: By Brian Harvey and Matthew Wright. MIT Press, Cambridge, MA. (1999). 579 pages. $60.00. Computers and Mathematics with Applications, 39(3–4):266, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900877.

Anonymous:2000:BRSe

[Ano00-70] Anonymous. Book review: State-space search: Algorithms, complexity, extensions, and applications: By Weixiong Zhang. Springer-Verlag, New York. (1999). 201 pages. $49.95. Computers and Mathematics with Applications, 40(2–3):417, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901858.

Anonymous:2000:BRSb

[Ano00-71]  Anonymous.  Book re-

view: Supercomputing, collision processes, and applications: Edited by Kenneth L. Bell, Keith A. Berrington, Derrick S. F. Crothers, Alan Hibbert and Kenneth T. Taylor. Kluwer Academic/Plenum Publishers, New York. (1999). 284 pages. $125.00, GBP 81.25, NLG 255.00. Computers and Mathematics with Applications, 39(3–4):262, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900312.

Anonymous:2000:BRTb

[Ano00-72]  Anonymous.  Book re-

view: TCP/IP Windows NT 4.0: Flashcards for MCSE elective exam 70-059: By Michael Moncur. O’Reilly, Sebastopol, CA. (1999). $29.95. Computers and Mathematics with Applications, 39(3–4):261, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210090021X.

Anonymous:2000:BRTc

[Ano00-73]  Anonymous.  Book re-

view: TCP/IP Windows NT 4.0: Flashcards for MCSE elective exam 70-059: By Michael Moncur. O’Reilly, Sebastopol, CA. $29.95. Computers and Mathematics with Applications, 39(7– 8):264, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100902736.

Anonymous:2000:BRB

[Ano00-74]  Anonymous.  Book re-

view: The beginner’s guide to Mathematica(R) version 4: By Jerry Glynn and Theodore Gray. Cambridge University Press, Cambridge, U.K. (2000). 434 pages. $74.95 (hardback); $29.95 (paperback). Computers and Mathematics with Applications, 40(2– 3):418, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100902013.

Anonymous:2000:BRCb

[Ano00-75]  Anonymous.  Book re-

view: The cathedral and the bazaar: By Eric S. Raymond. O’Reilly, Sebastapol, CA. (1999). 268 pages. $19.95. Computers and Mathematics with Applications, 39(3–4):263, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900397.

Anonymous:2000:BRFa

[Ano00-76] Anonymous. Book review:

The future of the electronic marketplace: Edited by Derek Leebaert. MIT Press, Cambridge, MA. (1998). 383 pages. $17.50. Computers and Mathematics with Applications, 39(3–4):260, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900154.

Anonymous:2000:BRFc

[Ano00-77]  Anonymous.  Book re-

view: The future of the electronic marketplace: Edited by Derek Leebaert. MIT Press, Cambridge, MA. (1998). 383 pages. $17.50. Computers and Mathematics with Applications, 39(7– 8):264, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902670.

Anonymous:2000:BRPb

[Ano00-78] Anonymous. Book review:

The perl CD bookshelf: 6 bestselling books on CD-ROM: O’Reilly, Sebastopol, CA. (1999). $59.95. Includes bonus book, Perl in a Nutshell: a Desktop Quick Reference. By Ellen Siever, Stephen Spainhour and Nathan Patwardhan. O’Reilly, Sebastopol, CA. (1999). 654 pages. Computers and Mathematics with Applications, 39(3–4):264, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900531.

Anonymous:2000:BRPd

[Ano00-79] Anonymous. Book review: The Perl CD bookshelf: 6 Bestselling books on CD-ROM: O’Reilly, Sebastopol, CA. (1999). $59.95. Includes bonus book, Perl in a Nutshell: a Desktop Quick Reference. By Ellen Siever, Stephen Spainhour and Nathan Patwardhan. O’Reilly, Sebastopol, CA. (1999). 654 pages. Computers and Mathematics with Applications, 39(7–8): 266, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902979.

Anonymous:2000:BRSa

[Ano00-80]  Anonymous.  Book re-

view: The sciences in enlightened Europe: Edited by William Clark, Jan Golinski and Simon Schaffer. The University of Chicago Press, Chicago. (1999). 566 pages. $85.00, £59.50 (cloth); $27.50, £19.50 (paper). Computers and Mathematics with Applications, 39(3–4):261, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900270.

Anonymous:2000:BRSd

[Ano00-81] Anonymous. Book review:

The sciences in enlightened Europe: Edited by William Clark, Jan Golinski and Simon Schaffer. The University of Chicago Press, Chicago. (1999). 566 pages. $85.00, £59.50 (cloth); $27.50, £19.50 (paper). Computers and Mathematics with Applications, 39(7– 8):265, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902815.

Anonymous:2000:BRUb

[Ano00-82]  Anonymous.  Book re-

view: The unofficial guide to LEGO(R) MINDSTORM-STM robots: By Jonathan B. Knudsen. O’Reilly, Sebastopol, CA. (1999). 247 pages. $24.95. Computers and Mathematics with Applications, 39(3–4):265, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900737.

Anonymous:2000:BRTd

[Ano00-83]  Anonymous.  Book re-

view: Theory of language: By Steven E. Weisler and Slavko Milekic. The MIT Press, Cambridge, MA. (2000). 344 pages. $35 (CD-ROM $35). Computers and Mathematics with Applications, 40(2–3):416, July/ August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901779.

Anonymous:2000:BRTe

[Ano00-84] Anonymous. Book review: Toward a science of consciousness III: the third Tucson discussions and debates: Edited by Stuart R. Hameroff, Alfred W. Kaszniak and David J. Chalmers. The MIT Press, Cambridge, MA. (1999). 504 pages. $55. Computers and Mathematics with Applications, 40(2–3):416, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100901792.

Anonymous:2000:BRTa

[Ano00-85] Anonymous. Book review:

Transportation and traffic theory, proceedings of the 14-th international symposium on transportation and traffic theory; Jerusalem, Israel, 20–23 July, 1999: Edited by Avishai Ceder. Pergamon, Elsevier Science, Amsterdam. (1999). 796 pages. $213.50, NLG 421.00, euro 191.04. Computers and Mathematics with Applications, 39(3–4):260, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900130.

Anonymous:2000:BRUc

[Ano00-86] Anonymous. Book review:

UNIX in a nutshell: a desktop quick reference for System V release 4 and Solaris 7, third edition: By Arnold Robbins. O’Reilly, Sebastopol, CA. (1999). 598 pages. $24.95. Computers and Mathematics with Applications, 39(3–4):265, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100900750.

Anonymous:2000:BRUa

[Ano00-87] Anonymous. Book review: User friendly: By J. D. “Illiad” Frazer. O’Reilly, Sebastapol, CA. (1999). 122 pages. $12.95. Computers and Mathematics with Applications, 39(3–4):265, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100900671.

Anonymous:2000:BRUd

[Ano00-88] Anonymous. Book review: Using MPI-2: Advanced features of the message-passing interface: By William Gropp, Ewing Lusk and Rajeev Thakur. The MIT Press, Cambridge, MA. (1999). 382 pages. $35 (each); $60 (set). Computers and Mathematics with Applications, 40(2–3):419, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100902098.

Anonymous:2000:BRUe

[Ano00-89]  Anonymous.  Book re-

view: Using MPI: Portable parallel programming with the message-passing interface: Second edition. By William Gropp, Ewing Lusk and Anthony Skjellum. The MIT Press, Cambridge, MA. (1999). 371 pages. $35 (each); $60 (set). Computers and Mathematics with Applications, 40(2–3):419, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902074.

Anonymous:2000:BRWd

[Ano00-90] Anonymous. Book review:

Why do buses come in threes?: the hidden mathematics of everyday life: By Rob Eastaway and Jeremy Wyndham, Illustrations by Barbara Shore. John Wiley & Sons, New York. (1998). 156 pages. $14.95. Computers and Mathematics with Applications, 40(2–3):419, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902116.

Anonymous:2000:BRWb

[Ano00-91] Anonymous. Book review:

Windows 98 in a nutshell: a desktop quick reference: By Tim O’Reilly, Troy Mott and Walter Glenn. O’Reilly, Sebastopol, CA. (1999). 618 pages. $24.95. Computers and Mathematics with Applications, 39(3–4):265, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100900610.

Anonymous:2000:BRWa

[Ano00-92] Anonymous. Book review: Windows NT server 4.0 in the enterprise; flashcards: MCSE core exam 70-068: By Michael Moncur. O’Reilly, Sebastopol, CA. (1999). $29.95. Computers and Mathematics with Applications, 39(3–4):261, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100900191.

Anonymous:2000:BRWc

[Ano00-93] Anonymous. Book review: Windows NT server 4.0 in the enterprise; flashcards: MCSE core exam 70-068: By Michael Moncur. O’Reilly, Sebastopol, CA. $29.95. Computers and Mathematics with Applications, 39(7–8):264, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100902712.

Anonymous:2000:CSA

[Ano00-94] Anonymous. Corrigendum to “Spectral analysis of the matrix multisplitting method for one-dimensional model problem”. Computers and Mathematics with Applications, 39(5–6):259, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000596.

See [BSS98].

Anonymous:2000:EBa

[Ano00-95]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 39(1–2):iii–v, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100900099.

Anonymous:2000:EBb

[Ano00-96]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 40(1):iii–v, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100900117.

Anonymous:2000:LCa

[Ano00-97] Anonymous. List of contents. Computers and Mathematics with Applications, 39(12):iii–x, June 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100900014.

Anonymous:2000:LCb

[Ano00-98] Anonymous. List of contents. Computers and Mathematics with Applications, 40(12):III–IX, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101800018.

Anonymous:2001:AIa

[Ano01a] Anonymous. Author index. Computers and Mathematics with Applications, 41 (12):xi–xii, June 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101800262.

Anonymous:2001:AIb

[Ano01b] Anonymous. Author index. Computers and Mathematics with Applications, 42(12):XI–XII, December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101902202.

Anonymous:2001:BR

[Ano01c]  Anonymous.  Book re-

ports. Computers and Mathematics with Applications, 41(7–8):1077–1084, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003412.

Anonymous:2001:BRAb

[Ano01d]  Anonymous.  Book re-

view: Advanced debugging methods: By Raimondas Lencevicius. Kluwer Academic, Boston. (2000). 183 pages. $110.00, NLG 280.00, GBP 75.00. Computers and Mathematics with Applications, 41(3–4):542, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900495.

Anonymous:2001:BRAa

[Ano01e] Anonymous. Book review:

Advances in neural information processing systems 12: Proceedings of the 1999 conference: Edited by Sara A. Solla, Todd K. Leen and Klaus-Robert Müller. The MIT Press, Cambridge, MA. (2000). 1080 pages. $65.00. Computers and Mathematics with Applications, 41(3–4):539–541, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101900318.

Anonymous:2001:BRB

[Ano01f] Anonymous. Book review: Building Linux clusters: By David Hm Spector. O’Reilly, Sebastopol, CA. (2000). 332 pages. $44.95 (CD-ROM included). Computers and Mathematics with Applications, 41(3–4):538, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101900276.

Anonymous:2001:BRCb

[Ano01g] Anonymous. Book review: Combinatorial optimization: Theory and algorithms: By Bernhard Korte and Jens Vygen. Springer, Berlin. (2000). 530 pages. $45.00, DM 79.00, öS 577.00, sFr 72.00, GBP 27.00. Computers and Mathematics with Applications, 41(3–4):538, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101900239.

Anonymous:2001:BRCc

[Ano01h] Anonymous. Book review: Computer-aided reasoning:

ACL2 case studies: Edited by Matt Kaufmann, Panagiotis Manolios and J Strother Moore. Kluwer Academic, Boston. (2000). 337 pages. $140.00, NLG 355.00, GBP 98.00. Computers and Mathematics with Applications, 41(3–4):541–542, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900392.

Anonymous:2001:BRCa

[Ano01i]  Anonymous.  Book re-

view: Computer architecture: Complexity and correctness: By Silvia M. Mueller and Wolfgang J. Paul. Springer, Berlin. (2000). 553 pages. $54.95, DM 98.00, öS 716.00, sFr 89.50, GBP 34.00. Computers and Mathematics with Applications, 41(3–4):536, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900094.

Anonymous:2001:BRCd

[Ano01j] Anonymous. Book review:

CVS: Pocket reference: By Gregor N. Purdy. O’Reilly, Sebastapol, CA. (2000). 75 pages. $9.95. Computers and Mathematics with Applications, 41(3–4):542, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101900410.

Anonymous:2001:BRD

[Ano01k]  Anonymous.  Book re-

view: Designing and programming CICS applications: By John Horswill and Members of the CICS Development Team at IBM Hursley. O’Reilly, Sebastapol, CA. (2000). 398 pages. $44.95 (CD-ROM included). Computers and Mathematics with Applications, 41(3–4):541, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101900355.

Anonymous:2001:BRK

[Ano01l] Anonymous. Book review: From Kant to Hilbert: a source book in the foundations of mathematics, Volumes I and II: By William Ewald. Clarendon Press, Oxford. (1999). 1340 pages (in 2 volumes). $85.00. Computers and Mathematics with Applications, 41(3–4):538, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900252.

Anonymous:2001:BRG

[Ano01m] Anonymous. Book review:

Geometry civilized: History, culture, and technique: By J. L. Heilbron. Clarendon Press, Oxford. (2000). 309 pages. $35.00. Computers and Mathematics with Applications, 41(3–4):537, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900197.

Anonymous:2001:BRH

[Ano01n]  Anonymous.  Book re-

view: Heterogeneous agent systems: By V. S. Subrahmanian, Piero Bonatti, Jürgen Dix, Thomas Eiter, Sarit Kraus, Fatma Özcan, and Robert Ross. The MIT Press, Cambridge, MA. (2000). 580 pages. $60.00. Computers and Mathematics with Applications, 41(3–4):542, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900513.

Anonymous:2001:BRIb

[Ano01o] Anonymous. Book review:

Issues in the economics of immigration: Edited by George J. Borjas. University of Chicago Press, Chicago. (2000). 399 pages. $55.00, £38.50. Computers and Mathematics with Applications, 41(3–4):536, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900057.

Anonymous:2001:BRJ

[Ano01p]  Anonymous.  Book re-

view: Java programming: From the beginning: By K. N. King. W. W. Norton, New York. (2000). 788 pages. $52.50. Computers and Mathematics with Applications, 41(3–4):541, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900331.

Anonymous:2001:BRM

[Ano01q] Anonymous. Book review: Mergers and productivity: Edited by Steven N. Kaplan. University of Chicago Press, Chicago. (2000). 337 pages. $40.00, £28.00. Computers and Mathematics with Applications, 41(3–4):536, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900070.

Anonymous:2001:BRN

[Ano01r] Anonymous. Book review:

Natural language processing and knowledge representation: Language for knowledge and knowledge for language: Edited by Lucja M. Iwańska and Stuart C. Shapiro. AAAI Press/The MIT Press, Menlo Park, CA. (2000). 459 pages. $40.00. Computers and Mathematics with Applications, 41(3–4):536–537, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900112.

Anonymous:2001:BRP

[Ano01s]  Anonymous.  Book re-

view: Perl for system administration: By David N. Blank-Edelman. O’Reilly, Sebastopol, CA. (2000). 430 pages. $34.95. Computers and Mathematics with Applications, 41(3–4):538, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900215.

Anonymous:2001:BRRb

[Ano01t]  Anonymous.  Book re-

view: Reasoning about rational agents: By Michael Wooldridge. The MIT Press, Cambridge, MA. (2000). 227 pages. $35.00. Computers and Mathematics with Applications, 41(3–4):542, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900471.

Anonymous:2001:BRRa

[Ano01u] Anonymous. Book review: Robo sapiens: Evolution of a new species: By Peter Menzel and Faith D’Aluisio. The MIT Press, Cambridge, MA. (2000). 239 pages. $29.95, £19.95. Computers and Mathematics with Applications, 41(3–4):537, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900173.

Anonymous:2001:BRSc

[Ano01v] Anonymous. Book review: Science & technology encyclopedia: the University of Chicago Press, Chicago. (1999). 572 pages. $22.50. Computers and Mathematics with Applications, 41 (3–4):542, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0898122101900458.

Anonymous:2001:BRSa

[Ano01w] Anonymous. Book review:

SymbolicC++: an introduction to computer algebra using object-oriented programming, 2nd extended and revised edition: By Tan Kiat Shi, Willi-Hans Steeb and Yorick Hardy. Springer, London. (2000). 671 pages. $62.00, DM 110.00, öS 803.00, sFr 100.50, GBP 35.00. Computers and Mathematics with Applications, 41(3–4):537, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101900136.

Anonymous:2001:BRSb

[Ano01x] Anonymous. Book review:

Systems, experts, and computers: the systems approach in management and engineering, World War II and after: Edited by Agatha C. Hughes and Thomas P. Hughes. The MIT Press, Cambridge, MA. (2000). 513 pages. $50.00. Computers and Mathematics with Applications, 41(3–4):538, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210190029X.

Anonymous:2001:BRF

[Ano01y]  Anonymous.  Book re-

view: The first computers — History and architectures: Edited by Raúl Rojas and Ulf Hashagen. The MIT Press, Cambridge, MA. (2000). 457 pages. $39.95. Computers and Mathematics with Applications, 41(3–4):537, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210190015X.

Anonymous:2001:BRIa

[Ano01z] Anonymous. Book review: The impact of international trade on wages: Edited by Robert C. Feenstra. University of Chicago Press, Chicago. (2000). 406 pages. $62.00, £39.50. Computers and Mathematics with Applications, 41(3–4):535–536, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101900033.

Anonymous:2001:BRV

[Ano01-27]  Anonymous.  Book re-

view: Visual basic shell programming: By J. P. Hamilton. O’Reilly, Sebastapol, CA. (2000). 373 pages. $29.95. Computers and Mathematics with Applications, 41(3–4):541, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900379.

Anonymous:2001:BRW

[Ano01-28] Anonymous. Book review:

Women becoming mathematicians: Creating a professional identity in post-World War II America: By Margaret A. M. Murray. The MIT Press, Cambridge, MA. (2000). 277 pages. $29.95. Computers and Mathematics with Applications, 41(3–4):542, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101900434.

Anonymous:2001:EB

[Ano01-29]  Anonymous.  Editorial

board. Computers and Mathematics with Applications, 41(1–2):FM3–FM5, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210185001X.

Anonymous:2001:LCa

[Ano01-30] Anonymous. List of contents. Computers and Mathematics with Applications, 41(12):iii–x, June 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122101800250.

Anonymous:2001:LCb

[Ano01-31] Anonymous. List of contents. Computers and Mathematics with Applications, 42(12):III–X, December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101902196.

Anonymous:2002:AIa

[Ano02a] Anonymous. Author index. Computers and Mathematics with Applications, 43 (12):X–XI, June 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122102800256.

Anonymous:2002:AIb

[Ano02b] Anonymous. Author index. Computers and Mathematics with Applications, 44(12):XI–XII, December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800281.

Anonymous:2002:BRa

[Ano02c]  Anonymous.  Book re-

ports. Computers and Mathematics with Applications, 43(8–9):1195–1202, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800232.

Anonymous:2002:BRb

[Ano02d] Anonymous. Book reports.

Computers and Mathematics with Applications, 44 (10–11):1479–1492, November/December 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102002729.

Anonymous:2002:EBa

[Ano02e]  Anonymous.  Editorial

board. Computers and Mathematics with Applications, 43(1–2):FM3–FM5, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800013.

Anonymous:2002:EBb

[Ano02f]  Anonymous.  Editorial

board. Computers and Mathematics with Applications, 44(1–2):FM3–FM5, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800268.

Anonymous:2002:LCa

[Ano02g] Anonymous. List of contents. Computers and Mathematics with Applications, 43(12):III–IX, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122102800244.

Anonymous:2002:LCb

[Ano02h] Anonymous. List of contents. Computers and Mathematics with Applications, 44(12):III–X, December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210280027X.

Anonymous:2003:AI

[Ano03a] Anonymous. Author index. Computers and Mathematics with Applications, 46(12):XII–XIV, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902458.

Anonymous:2003:BRk

[Ano03b] Anonymous. Book reports.

Computers and Mathematics with Applications, 45 (4–5):851–860, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000464.

Anonymous:2003:BRx

[Ano03c] Anonymous. Book reports.

Computers and Mathematics with Applications, 46 (8–9):1455–1460, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902331.

Anonymous:2003:BRy

[Ano03d]  Anonymous.  Book re-

ports. Computers and Mathematics with Applications, 46(12):1875–1880, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902434.

Anonymous:2003:BRSg

[Ano03e]  Anonymous.  Book re-

view: A singular introduction to commutative algebra: By Gert-Martin Greuel and Gerhard Pfister, with contributions by Olaf Bachmann, Christoph Lossen and Hans Schönemann. Springer-Verlag, Berlin. (2002). 588 pages. $44.95, EUR 39.95 (CD-ROM included). Computers and Mathematics with Applications, 46(5–6):978, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103901635.

Anonymous:2003:BRAd

[Ano03f] Anonymous. Book review: Acoustic communication in insects and anurans: Common problems and diverse solutions: By H. Carl Gerhardt and Franz Huber. The University of Chicago Press, Chicago, IL. (2002): 531 pages. $35. Computers and Mathematics with Applications, 46(2–3):516, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900678.

Anonymous:2003:BRAc

[Ano03g] Anonymous. Book review: Advanced solid state physics: By Philip Phillips. Westview Press, Boulder, CO. (2002). 386 pages. $55. Computers and Mathematics with Applications, 46(2–3):506–507, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900502.

Anonymous:2003:BRAi

[Ano03h] Anonymous. Book review:

Algorithms for hard problems: Introduction to combinatorial optimization, randomization, approximation, heuristics: Second Edition. By Juraj Hromkovic. Springer, New York. (2003). 544 pages. $54.95. Computers and Mathematics with Applications, 46(5–6):986, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901829.

Anonymous:2003:BRAa

[Ano03i]  Anonymous.  Book re-

view: Algorithms on trees and graphs: By Gabriel Valiente. Springer, New York. (2002). 490 pages. $49.95. Computers and Mathematics with Applications, 45(10–11):1778, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210380144X.

Anonymous:2003:BRAg

[Ano03j] Anonymous. Book review:

Applied computational economics and finance: By Mario J. Miranda and Paul L. Fackler. The MIT Press, Cambridge, MA. (2002). 510 pages. $70.00. Computers and Mathematics with Applications, 46(5–6):983, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901775.

Anonymous:2003:BRAh

[Ano03k]  Anonymous.  Book re-

view: Artificial life VIII: Edited by Russell K. Standish, Mark A. Bedau, and Hussein A. Abbass. The MIT Press, Cambridge, MA. (2003). 434 pages. $75. Computers and Mathematics with Applications, 46(5–6):984–985, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901805.

Anonymous:2003:BRB

[Ano03l]  Anonymous.  Book re-

view: Brain-wise: Studies in neurophilosophy: By Patricia Smith Churchland.The MIT Press, Cambridge, MA. (2002). 471 pages. $25.00. Computers and Mathematics with Applications, 46(5–6):980, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901714.

Anonymous:2003:BRCb

[Ano03m] Anonymous. Book review:

Calendrical tabulations: By Edward M. Reingold and Nachum Dershowitz. Cambridge University Press, New York. (2002). 606 pages $120. Computers and Mathematics with Applications, 45(10–11):1777, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801396.

Anonymous:2003:BRCc

[Ano03n]  Anonymous.  Book re-

view: Cellular automata and complexity: Collected papers: By Stephen Wolfram.Westview Press, Boulder, CO. (1994). 596 pages. $35. Computers and Mathematics with Applications, 45(10–11):1779, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801475.

Anonymous:2003:BRCi

[Ano03o]  Anonymous.  Book re-

view: Cognitive modeling: Edited by Thad A. Polk and Colleen M. Seifert. The MIT Press, Cambridge, MA. (2002). 1270 pages. $55. Computers and Mathematics with Applications, 46(2–3):511–512, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900605.

Anonymous:2003:BRCd

[Ano03p] Anonymous. Book review: Combinatorial algorithms: Enlarged Second Edition. By T. C. Hu and M. T. Shing. Dover Publications, Mineola, NY. (2002). 354 pages $16.95. Computers and Mathematics with Applications, 45(10–11):1779, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801463.

Anonymous:2003:BRCk

[Ano03q]  Anonymous.  Book re-

view: Combinatorics for computer science: By S. Gill Williamson. Dover, Mineola, NY. (2002). 479 pages. $22.95. Computers and Mathematics with Applications, 46(2–3):517, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900691.

Anonymous:2003:BRCg

[Ano03r] Anonymous. Book review:

Computational methods for fluid dynamics: Third Edition. By J. H. Ferziger. and M. Perić. Springer, New York. (2002). 423 pages. $54.95. Computers and Mathematics with Applications, 46(2–3):503–504, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900460.

Anonymous:2003:BRCf

[Ano03s] Anonymous. Book review:

Computational physics: FOR-TRAN version: By Steven E. Koonin and Dawn C. Meredith. Westview Press, Boulder, CO. (1990). 639 pages. $65. Computers and Mathematics with Applications, 46(2–3):502, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900435.

Anonymous:2003:BRCa

[Ano03t]  Anonymous.  Book re-

view: Computationalism new directions: Edited by Matthias Scheutz. The MIT Press, Cambridge, MA. (2002). 209 pages. $35. Computers and Mathematics with Applications, 45(10–11):1776, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801372.

Anonymous:2003:BRCe

[Ano03u]  Anonymous.  Book re-

view: Conversations with a mathematician: Math. art. science and the limits of reason: By Gregory J. Chaitin. Springer, London. (2002). 158 pages. $29.95. Computers and Mathematics with Applications, 45(10–11):1781, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801499.

Anonymous:2003:BRCl

[Ano03v]  Anonymous.  Book re-

view: Cost proxy models and telecommunications policy: a new empirical approach to regulation: By Farid Gasmi, D. Mark Kennet, Jean-Jacques Laffont and William W. Sharkey. The MIT Press, Cambridge, MA. (2002). 257 pages. $45.00 (CD-ROM included). Computers and Mathematics with Applications, 46(5–6):982, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901763.

Anonymous:2003:BRCj

[Ano03w] Anonymous. Book review:

Customer-based IP service monitoring with mobile software agents: By Manuel Gunter. Birkhäuser, Basel. 160 pages. $39.95, sFr. 42, EUR 27.10. Computers and Mathematics with Applications, 46(2–3):516–517, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390068X.

Anonymous:2003:BRDd

[Ano03x]  Anonymous.  Book re-

view: Decisions uncertainty. and the brain: the science of neuroeconomics: By Paul W. Glimcher. The MIT Press, Cambridge, MA. (2003). 375 pages. $37.95. Computers and Mathematics with Applications, 46(5–6):978, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901611.

Anonymous:2003:BRDa

[Ano03y]  Anonymous.  Book re-

view: Designing sociable robots: By Cynthia L. Breazeal. The MIT Press, Cambridge, MA. (2002). 263 pages. $49.95. Computers and Mathematics with Applications, 45(10–11):1774, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103801293.

Anonymous:2003:BRDe

[Ano03z]  Anonymous.  Book re-

view: Dollarization: Edited by Eduardo Levy Yeyati and Federico Sturzenegger. The MIT Press, Cambridge, MA. (2003). 341 pages. $45.00. Computers and Mathematics with Applications, 46(5–6):979, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901659.

Anonymous:2003:BRDb

[Ano03-27] Anonymous. Book review: Duelling idiots and other probability puzzlers: By Paul J. Nahin. Princeton University Press, Princeton. (2000). 269 pages. $18.95. Computers and Mathematics with Applications, 45(10–11):1781, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801517.

Anonymous:2003:BREc

[Ano03-28] Anonymous. Book review:

Econometrics volume 3: Economic growth in the information Age: By Dale W. Jorgenson. The MIT Press, Cambridge, MA. (2002). 462 pages. $55. Computers and Mathematics with Applications, 46(2–3):507, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900526.

Anonymous:2003:BREa

[Ano03-29]  Anonymous.  Book re-

view: Economic policy reforms and the Indian economy: Edited by Anne O. Krueger. The University of Chicago Press, Chicago, IL. (2002). 377 pages. $75. Computers and Mathematics with Applications, 45(10–11):1778–1779, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801451.

Anonymous:2003:BREb

[Ano03-30] Anonymous. Book review: Essential Mavle(R) 7: an introduction for scientific programmers: By Robert M. Corless. Springer, New York. (2002). 282 pages. $44.95. Computers and Mathematics with Applications, 46(2–3):507, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900514.

Anonymous:2003:BREe

[Ano03-31] Anonymous. Book review: Exploratory data mining and data cleaning: By Tamraparni Dasu and Theodore Johnson. John Wiley & Sons, Hoboken, NJ. 203 pages. $69.95. Computers and Mathematics with Applications, 46(5–6):980, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901702.

Anonymous:2003:BRFc

[Ano03-32]  Anonymous.  Book re-

view: Feynman and computation: Exploring the limits of computers: Edited by Anthony J. G. Hey. Westview Press, Cambridge MA. (2002). xxiii + 438 pages. Computers and Mathematics with Applications, 45(10–11):1781–1782, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801529.

Anonymous:2003:BRFd

[Ano03-33]  Anonymous.  Book re-

view: Feynman lectures on gravitation: By Richard P. Feynman, Fernando B. Moringo, and William G. Wagner. Edited by Brian Hatfield. Westview Press, Boulder, CO. (2003). 232 pages. $35. Computers and Mathematics with Applications, 45(10–11):1782, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801530.

Anonymous:2003:BRFf

[Ano03-34]  Anonymous.  Book re-

view: Finite elements using Maple: a symbolic programming approach: By A. Portela and A. Charafi. Springer, New York. (2002). 325 pages. $49.95. (CD-ROM included). Computers and Mathematics with Applications, 46(2–3):515–516, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900666.

Anonymous:2003:BRFb

[Ano03-35]  Anonymous.  Book re-

view: Fractals, graphics. & mathematics education: Edited by Michael L. Fame & Benoit B. Mandelbrot. The Mathematical Association of America, Washington, D.C. (2002). 206 pages. $39.95. Computers and Mathematics with Applications, 45(10–11):1776, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103801360.

Anonymous:2003:BRAe

[Ano03-36]  Anonymous.  Book re-

view: From animals to animals 7: Edited by Bridget Hallam, Dario Floreano, John Hallam,l Gillian Hayes, and Jean-Arcady Meyer. The MIT Press, Cambridge, MA. (2002). 420 pages. $70. Computers and Mathematics with Applications, 46(2–3):518–519, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210390071X.

Anonymous:2003:BRFg

[Ano03-37] Anonymous. Book review:

Front tracking for hyperbolic conservation laws: By Helge Holden and Nils Henrik Risebro. Springer-Verlag, New York. (2002). 363 pages. $64.95, EUR 64.95, SFR 108.00, GBP 45.50. Computers and Mathematics with Applications, 46 (5–6):978, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103901581.

Anonymous:2003:BRFe

[Ano03-38]  Anonymous.  Book re-

view: Frontiers in health policy research. Volume 5: Edited by Alan M. Garber. The MIT Press, Cambridge, MA. (2002). 166 pages. $25. Computers and Mathematics with Applications, 46(2–3):508, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900563.

Anonymous:2003:BRFa

[Ano03-39] Anonymous. Book review:

Fundamentals of matrix computations: Second Edition. By David S. Watkins. Wiley–Interscience, New York. (2002). 618 pages. $89.95. Computers and Mathematics with Applications, 45(10–11):1775–1776, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103801335.

Anonymous:2003:BRGc

[Ano03-40]  Anonymous.  Book re-

view: Gaussian self-affinity and fractals: By B. B. Mandelbrot. Springer, New York. (2002). 654 pages. $54.95. Computers and Mathematics with Applications, 46(2–3):505, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900472.

Anonymous:2003:BRGa

[Ano03-41]  Anonymous.  Book re-

view: Gene regulation and metabolism: Postaenomic computational approaches: Edited by Julio Collado-Vides and Ralf Hofestädt. The MIT Press, Cambridge, MA. (2002). 310 pages. $45. Computers and Mathematics with Applications, 45(10–11):1781, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103801505.

Anonymous:2003:BRGb

[Ano03-42]  Anonymous.  Book re-

view: Geometric numerical integration: Structurepreserving algorithms for ordinary differential equations: By E. Hairer, C. Lubich, and G. Wanner. Springer, Berlin. (2002). 515 pages. $84.95. Computers and Mathematics with Applications, 45(10–11):1782– 1784, May/June 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103801554.

Anonymous:2003:BRHc

[Ano03-43] Anonymous. Book review:

Handbook of data mining and knowledge discovery: Edited by Willi Klösgen and Jan M. Zytkow. Oxford University Press, New York. (2002). 1026 pages. $250.00. Computers and Mathematics with Applications, 46(5–6):981–982, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901751.

Anonymous:2003:BRHa

[Ano03-44] Anonymous. Book review:

High-energy particle diffraction: By V. Barone and E. Predazzi. Springer, New York. (2002). 407 pages. $79.95. Computers and Mathematics with Applications, 46(2–3):510–511, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900599.

Anonymous:2003:BRIb

[Ano03-45]  Anonymous.  Book re-

view: Information and randomness: an algorithmic perspective: Second Edition, Revised and Extended. By Cristian S. Calude. Springer, New York. (2002). 467 pages. $54.95. Computers and Mathematics with Applications, 45(10–11):1778, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801438.

Anonymous:2003:BRIc

[Ano03-46] Anonymous. Book review: Introduction to numerical analysis, Texts in Applied Mathematics 12: Third Edition. By J. Stoer and R. Bulirsch. Translated by R. Bartels, W. Gautschi, and C. Witzgall. Springer, New York. (2002). 744 pages. $69.95. Computers and Mathematics with Applications, 46(2–3):509–510, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900587.

Anonymous:2003:BRJ

[Ano03-47] Anonymous. Book review:

Java precisely: By Peter Sestoft. The MIT Press, Cambridge, MA. (2002). 118 pages. $14.95. Computers and Mathematics with Applications, 46(2–3):517–518, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900708.

Anonymous:2003:BRLc

[Ano03-48]  Anonymous.  Book re-

view: Learning and generalisation: With applications to neural networks: By M. Vidyasagar. Springer-Verlag, London. (2003). 488 pages. $119.00, EUR 99.95, SFR 166.00, GBP 60.00. Computers and Mathematics with Applications, 46(5–6):981, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390174X.

Anonymous:2003:BRLb

[Ano03-49]  Anonymous.  Book re-

view: Lectures on the theory of games: By Harold W. Kuhn.Princeton University Press, Princeton, NJ. (2003). 107 pages. $24.95, £17.95 (paper); $49.50, £35.00 (cloth). Computers and Mathematics with Applications, 46(5–6):979, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901647.

Anonymous:2003:BRLa

[Ano03-50]  Anonymous.  Book re-

view: Leonardo’s laptop: Human needs and the new computing technologies: By Ben Schneiderman. The MIT Press, Cambridge, MA. (2002) 269 pages. $24.95. Computers and Mathematics with Applications, 45(10–11):1774, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801281.

Anonymous:2003:BRLd

[Ano03-51]  Anonymous.  Book re-

view: Level set methods and dynamic implicit surfaces: By Stanly Osher and Ronald Fedkiw. Springer, New York. (2003). 273 pages. $79.95. Computers and Mathematics with Applications, 46(5–6):983–984, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901799.

Anonymous:2003:BRMa

[Ano03-52] Anonymous. Book review:

Mathematical principles of signal processing: Fourier and Wavelet analysis: By Pierre Breé Maud. Springer, New York. (2002). 269 pages. $79.95. Computers and Mathematics with Applications, 45(10–11):1777, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801414.

Anonymous:2003:BRMb

[Ano03-53]  Anonymous.  Book re-

view: Microarrays for an integrative genomics: By Isaac S. Kohane, Alvin T. Kho, and Atul J. Butte. The MIT Press, Cambridge, MA. (2002). 306 pages. $40. Computers and Mathematics with Applications, 46(2–3):505–506, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900496.

Anonymous:2003:BRNd

[Ano03-54]  Anonymous.  Book re-

view: Neural engineering: Computation, representation, and dynamics in neurobiological systems: By Chris Eliasmith and Charles H. Anderson. The MIT Press, Cambridge, MA. (2003). 356 pages. $49.95, £35.50. Computers and Mathematics with Applications, 46(5–6):985–986, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901817.

Anonymous:2003:BRNc

[Ano03-55]  Anonymous.  Book re-

view: Neurotechnology for biomimetic robots: Edited by Joseph Ayers, Joel L. Davis, and Alan Rudolph. The MIT Press, Cambridge, MA. (2002). 636 pages. $65. Computers and Mathematics with Applications, 46(2–3):505, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900484.

Anonymous:2003:BRNb

[Ano03-56] Anonymous. Book review: Number theory for computing: Second Edition. By Song Y. Yan. Springer, New York.(2002). 435 pages $49.95. Computers and Mathematics with Applications, 46(2–3):502–503, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900447.

Anonymous:2003:BROa

[Ano03-57]  Anonymous.  Book re-

view: Ocean yearbook 15: Edited by Elisabeth Mann Borgese, Aldo Chircop, and Moira McConnell. The University of Chicago Press, Chicago. (2002). 1018 pages. $130. Computers and Mathematics with Applications, 46(2–3):507–508, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900538.

Anonymous:2003:BROb

[Ano03-58]  Anonymous.  Book re-

view: Oval track and other permutation puzzles: And just enough group theory to solve them: By John O. Kiltinen. Mathematical Association of America, Washington, DC. (2003). 305 pages. $42.50 (CD-ROM included). Computers and Mathematics with Applications, 46(5–6):979, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901672.

Anonymous:2003:BRPe

[Ano03-59] Anonymous. Book review: Partial differential equations: Methods and applications: Second Edition. By Robert C. McOwen. Prentice Hall, Upper Saddle River, NJ. (2003). 452 pages. $95.00. Computers and Mathematics with Applications, 46(5–6):978, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901593.

Anonymous:2003:BRPa

[Ano03-60]  Anonymous.  Book re-

view: Patents, citations & innovations: a window on the knowledge economy: By Adam B. Jaffe and Manuel Trajtenberg. The MIT Press, Cambridge, MA. (2002). 478 pages. $50. Computers and Mathematics with Applications, 45(10–11):1774, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210380127X.

Anonymous:2003:BRPb

[Ano03-61]  Anonymous.  Book re-

view: Perceptual learning: Edited by Manfred Fahle and Tomaso Poggio. The MIT Press, Cambridge, MA. (2002). 455 pages. $65. Computers and Mathematics with Applications, 45(10–11):1775, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801323.

Anonymous:2003:BRPd

[Ano03-62]  Anonymous.  Book re-

view: Population viability analysis: Edited by Steven R. Beissinger and Dale R. McCullouch. The University of Chicago Press Chicago (2002) 577 pages $35. Computers and Mathematics with Applications, 45(10–11):1777, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801402.

Anonymous:2003:BRPc

[Ano03-63]  Anonymous.  Book re-

view: Pricing the priceless: a health care conundrum: By Joseph P. Newhouse. The MIT Press, Cambridge, MA. (2002). 258 pages. $35. Computers and Mathematics with Applications, 45(10–11):1776–1777, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103801384.

Anonymous:2003:BRQ

[Ano03-64]  Anonymous.  Book re-

view: Quantum calculus: By Victor Kac and Pokman Cheung. Springer, New York. (2002). 112 pages. $29.95. Computers and Mathematics with Applications, 45(10–11):1776, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103801347.

Anonymous:2003:BRRb

[Ano03-65]  Anonymous.  Book re-

view: Rigid body dynamics of mechanisms 1. Theoretical basis: By Hubert Hahn. Springer, New York. (2002). 336 pages. $98. Computers and Mathematics with Applications, 46(2–3):519–520, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900721.

Anonymous:2003:BRSc

[Ano03-66]  Anonymous.  Book re-

view: Schwarz–Christoffel mapping: By Tobin A. Driscoll and Lloyd N. Trefethen. Cambridge University Press, Cambridge. (2002). 132 pages. $40. Computers and Mathematics with Applications, 46(2– 3):513, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103900629.

Anonymous:2003:BRSe

[Ano03-67]  Anonymous.  Book re-

view: Scientific computing with ordinary differential equations: By Peter Deuflhard and Folkmar Bornemann. Translated by Werner C. Rheinboldt. Springer, New York. (2002). 485 pages. $59.95. (CD-ROM included). Computers and Mathematics with Applications, 46(2–3):515, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900642.

Anonymous:2003:BRSf

[Ano03-68]  Anonymous.  Book re-

view: Self-regularity: a new paradigm for primal-dual interior-point algorithms: By Jiming Peng, Cornelias Roos, and Tanigs Terlaky. Princeton University Press, Princeton, NJ. (2002). 185 pages. $65.00 (cloth); $29.95 (paper). Computers and Mathematics with Applications, 46(2–3):520, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900733.

Anonymous:2003:BRSh

[Ano03-69] Anonymous. Book review: Shane interrogation for computer aided design and manufacturing: By Nicholas M. Patrikalakis and Takashi Maekawa. Springer-Verlag, Berlin. (2002). 408 pages. $44.95, EUR 39.95, SFR 66.50, GBP 28.00. Computers and Mathematics with Applications, 46(5–6):979, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901684.

Anonymous:2003:BRSd

[Ano03-70]  Anonymous.  Book re-

view: Solar neutrinos: the first thirty years: Edited by John N. Bahcall, Raymond Davis, Jr., Peter Parker, Alexei Smirnov, and Roger Ulrich. Westview Press, Boulder, CO. (2002). 462 pages. $50. Computers and Mathematics with Applications, 46(2–3):513–514, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900630.

Anonymous:2003:BRSi

[Ano03-71] Anonymous. Book review:

Spinning the semantic web: Bringing the World Wide Web to its full potential: Edited by Dieter Fensel, James Hendler, Henry Lieberman and Wolfgang Wahlster. The MIT Press, Cambridge, MA. (2003). 479 pages. $40.00. Computers and Mathematics with Applications, 46(5–6):980, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901726.

Anonymous:2003:BRSb

[Ano03-72] Anonymous. Book review:

SPSS in practice: an illustrated guide: By Basant K. Puri. Arnold, London. (2002). 175 pages. $29.95. Computers and Mathematics with Applications, 45(10–11):1776, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801359.

Anonymous:2003:BRSa

[Ano03-73] Anonymous. Book review: Statistical computing: an introduction to data analysis using S-Plus: by Michael J. Crawley. John Wiley & Sons, New York. (2002). 761 pages. $85. Computers and Mathematics with Applications, 45(10–11):1775, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103801311.

Anonymous:2003:BRSj

[Ano03-74]  Anonymous.  Book re-

view: Structural bioinformatics: Edited by Philip E. Bourne and Helge Weissig. John Wiley & Sons, Hoboken, NJ. (2003). 649 pages. $149.95 (cloth); $69.95 (paper). Computers and Mathematics with Applications, 46(5–6):983, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901787.

Anonymous:2003:BRT

[Ano03-75]  Anonymous.  Book re-

view: Tabular application development for information systems: an object-oriented methodology: By Talib Damij. Springer, New York. (2001). 190 pages. $54.95. (CD-ROM included). Computers and Mathematics with Applications, 46(2–3):503, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900459.

Anonymous:2003:BRAb

[Ano03-76] Anonymous. Book review:

The art of the infinite: the pleasures of mathematics: By Robert Kaplan and Ellen Kaplan. Oxford University Press, New York. (2003). 324 pages. $26. Computers and Mathematics with Applications, 46(2–3):502, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900423.

Anonymous:2003:BRAf

[Ano03-77] Anonymous. Book review:

The asymmetrical brain: Edited by Kenneth Hugdahl and Richard J. Davidson. The MIT Press, Cambridge, MA. (2003). 796 pages. $90.00. Computers and Mathematics with Applications, 46(5–6):980–981, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901738.

Anonymous:2003:BRCh

[Ano03-78] Anonymous. Book review: The cognitive animal: Empirical and theoretical perspectives on animal cognition: Edited by Marc Bekoff, Colin Allen and Gordon M. Burghardt.The MIT Press, Cambridge, MA. (2002). 482 pages. $38. Computers and Mathematics with Applications, 46(2–3):508–509, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900575.

Anonymous:2003:BRDc

[Ano03-79]  Anonymous.  Book re-

view: The dynamic neuron: By John Smythies. The MIT Press, Cambridge, MA. (2002). 150 pages. $35. Computers and Mathematics with Applications, 46(2–3):515, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900654.

Anonymous:2003:BREd

[Ano03-80]  Anonymous.  Book re-

view: The economics of marine resources and conservation policy: the Pacific halibut case study with commentary: Edited by James A. Crutchfield and Arnold Zellner. The University of Chicago Press, Chicago. (2003). 226 pages. $60.00, £42.00. Computers and Mathematics with Applications, 46(5–6):978, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901623.

Anonymous:2003:BREf

[Ano03-81] Anonymous. Book review:

The enterprise of science in Islam: New perspectives: Edited by Jan P. Hogendijk and Abdelhamid I. Sabra. The MIT Press, Cambridge, MA. (2003). 386 pages. $45.00. Computers and Mathematics with Applications, 46(5–6):980, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901696.

Anonymous:2003:BRHb

[Ano03-82]  Anonymous.  Book re-

view: The hybrid multiscale simulation technology: An introduction with application to astrophysical and laboratory plasmas: By Alexander S. Lipatov. Springer-Verlag, Berlin. (2002). 403 pages. $79.95, EUR 73.95, SFR 123.00, GBP 52.00. Computers and Mathematics with Applications, 46(5–6):978, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210390160X.

Anonymous:2003:BRIa

[Ano03-83]  Anonymous.  Book re-

view: The iconic logic of Peirce’s graphs: By Sun-Joo Shin.The MIT Press, Cambridge, MA. (2002). 208 pages. $34.95. Computers and Mathematics with Applications, 45 (10–11):1774–1775, May/ June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210380130X.

Anonymous:2003:BRMc

[Ano03-84] Anonymous. Book review: The Mathematical theory of finite element methods: Second Edition. By Susanne C. Brenner and L. Ridgway Scott. Springer, New York. (2002). 361 pages. $49.95. Computers and Mathematics with Applications, 46(2–3):512–513, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900617.

Anonymous:2003:BRGd

[Ano03-85] Anonymous. Book review:

The MIT guide to science and engineering communication: Second Edition. By James G. Paradis and Muriel L. Zimmerman. The MIT Press, Cambridge, MA. (2002). 324 pages. $29.95. Computers and Mathematics with Applications, 46(2–3):508, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900551.

Anonymous:2003:BRNa

[Ano03-86]  Anonymous.  Book re-

view: The Neural simulation language: a system for brain modeling: By Alfredo Weitzenfeld, Michael A. Arbib, and Amanda Alexander. The MIT Press, Cambridge, MA. (2002). 439 pages. $55, 137.95. Computers and Mathematics with Applications, 45(10–11):1779–1781, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801487.

Anonymous:2003:BRRa

[Ano03-87]  Anonymous.  Book re-

view: The Rio book of concepts: By Greeorv L. Murphy The MIT Press, Cambridge, MA. (2002). 555 pages. $45. Computers and Mathematics with Applications, 45(10–11):1778, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103801426.

Anonymous:2003:BRV

[Ano03-88] Anonymous. Book review: The virtual album of fluid motion: By Susanne Kilian and Stefan Turek. Springer Electronic Media, Dortmund, Germany. (2002). CD Rom. $79.95. Computers and Mathematics with Applications, 45(10–11):1782, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103801542.

Anonymous:2003:BRU

[Ano03-89]  Anonymous.  Book re-

view: Uncanny networks: Dialogues with the virtual intelligentsia: By Geert Lovink. The MIT Press, Cambridge, MA. (2002). 374 pages. $27.95. Computers and Mathematics with Applications, 46(5–6):979, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901660.

Anonymous:2003:BRW

[Ano03-90] Anonymous. Book review:

Words without meaning: By Christopher Gauker. The MIT Press, Cambridge, MA. (2003). 299 pages. $25. Computers and Mathematics with Applications, 46(2–3):508, July/ August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390054X.

Anonymous:2003:EBa

[Ano03-91]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 45(1–3):iii–v, January/February 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800007.

Anonymous:2003:EBb

[Ano03-92]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 46(1):iii–v, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103900745.

Anonymous:2003:LCa

[Ano03-93] Anonymous. List of contents. Computers and Mathematics with Applications, 45(12):III–XIV, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900125.

Anonymous:2003:LCb

[Ano03-94] Anonymous. List of contents. Computers and Mathematics with Applications, 46(12):III–XI, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902446.

Anonymous:2003:P

[Ano03-95] Anonymous. Preface. Computers and Mathematics with Applications, 45(6–9): xv, March/May 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103801566.

Anonymous:2004:BRb

[Ano04a]  Anonymous.  Book re-

port. Computers and Mathematics with Applications, 47(10–11):1787–1808, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839919.

Anonymous:2004:BRf

[Ano04b]  Anonymous.  Book re-

port. Computers and Mathematics with Applications, 48(1–2):305–334, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840161.

Anonymous:2004:BRPe

[Ano04c]  Anonymous.  Book re-

port: Professor Ervin Y. Rodin, Campus Box 1040, Washington University in St. Louis, One Brookings Drive, St Louis, MO 63130, U.S.A. Computers and Mathematics with Applications, 48(3–4):649–664, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840847.

Anonymous:2004:BRAb

[Ano04d]  Anonymous.  Book re-

view: Analog VLSI: Circuits and principles: By Shih-Chii Liu, Jörg Kramer, Giacomo Indiveri, Tobias Delbrück, and Rodney Douglas. The MIT Press, Cambridge, MA. (2002). 434 pages. $60. Computers and Mathematics with Applications, 47(6–7):1150, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901058.

Anonymous:2004:BRAa

[Ano04e]  Anonymous.  Book re-

view: Analysis of panel data: Second edition. By Cheng Hsiao. Cambridge University Press, New York. (2003). 366 pages. $85 (hardback), $30 (paperback). Computers and Mathematics with Applications, 47(6–7):1147–1148, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900995.

Anonymous:2004:BRC

[Ano04f]  Anonymous.  Book re-

view: Computational partial differential equations numerical methods and diffpack programming: Second edition. Edited by Hans Petter Langtangen. Springer. Heidelberg. Germany. (2003). 855 pages. $69.95. Computers and Mathematics with Applications, 47(8–9):1484–1486, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901496.

Anonymous:2004:BRD

[Ano04g]  Anonymous.  Book re-

view: Discrete choice methods with simulation: By Kenneth E. Train. Cambridge University Press, New York, NY. (2003). 334 pages. $85 (hardback), $30 (paperback). Computers and Mathematics with Applications, 47(6–7):1148, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901009.

Anonymous:2004:BRE

[Ano04h]  Anonymous.  Book re-

view: Evolutionary dynamics and extensive form games: By Ross Cressman. The MIT Press, Cambridge, MA. (2003). 316 pages. $45. Computers and Mathematics with Applications, 47(6–7):1148–1149, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901022.

Anonymous:2004:BRIb

[Ano04i]  Anonymous.  Book re-

view: Image analysis, random fields and Markov Chain Monte Carlo methods: a mathematical introduction: Second edition with 75 figures. By Gerhard Winkler. Springer, Berlin. (2003). 388 pages. $79.95, EUR 69.95, SFR 116.50, GBP 49.00. Computers and Mathematics with Applications, 47(8–9):1481, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901423.

Anonymous:2004:BRIa

[Ano04j]  Anonymous.  Book re-

view: Introduction to Maple: Third Edition. By André Heck. Springer, New York, NY. (2003). 828 pages. $49.95, EUR49.95, SFR 86.00, GBP 38.50. Computers and Mathematics with Applications, 47 (6–7):1149–1150, March/ April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901034.

Anonymous:2004:BRL

[Ano04k]  Anonymous.  Book re-

view: Language modeling for information retrieval:

Edited by W. Bruce Croft and John Lafferty. Kluwer Academic Publishers, The Netherlands. (2003). 245 pages. $97.00, EUR 99.00, GBP 62.00. Computers and Mathematics with Applications, 47(8–9):1482, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901435.

Anonymous:2004:BRMb

[Ano04l]  Anonymous.  Book re-

view: Me++: the cyborg self and the networked city: By William J. Mitchell. The MIT Press, Cambridge, MA. (2003). 259 pages. $27.95, £18.50. Computers and Mathematics with Applications, 47(6–7):1148, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901010.

Anonymous:2004:BRMc

[Ano04m]  Anonymous.  Book re-

view: Methods and applications of linear models. Regression and the analysis of variance, Second edition: By R. R. Hocking. Wiley–Interscience, Hoboken, NJ. (2003). 741 pages. $105.00. Computers and Mathematics with Applications, 47(6–7):1150–1152, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901071.

Anonymous:2004:BRMa

[Ano04n] Anonymous. Book review: Monetary theory and policy: Second edition. By Carl E. Walsh. The MIT Press, Cambridge, MA. (2003) 612 pages. $70. Computers and Mathematics with Applications, 47(6–7):1145–1146, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900971.

Anonymous:2004:BRMe

[Ano04o] Anonymous. Book review: MuPAD pro computing essentials: By Miroslaw Majewski. Springer, Berlin. (2002). 456 pages. $49.95, EYR 39.95, SFR 68.50 GBP 28.00. Computers and Mathematics with Applications, 47(8–9):1480–1481, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901411.

Anonymous:2004:BRN

[Ano04p] Anonymous. Book review:

Numerical analysis in modern scientific computing: an introduction: Second edition with 65 illustrations. By Peter Deuflhard and Andreas Hohmann. Springer, Berlin. (2003). 360 pages. $49.95, EUR 59.95, SFR 99.50, GBP 42.00. Computers and Mathematics with Applications, 47(8–9):1483–1484, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901472.

Anonymous:2004:BROb

[Ano04q]  Anonymous.  Book re-

view: Open-domain question answering from large text collections: By Marius Pasca. CSLI Publications. Stanford, CA. (2003). 149 pages. $70.00. Computers and Mathematics with Applications, 47(8–9):1483, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901460.

Anonymous:2004:BROa

[Ano04r] Anonymous. Book review:

Operations research: an introduction: Seventh edition. By Hamdy A. Taha. Prentice Hall, Upper Saddle River, NJ. (2003). 830 pages. $115. Computers and Mathematics with Applications, 47(6–7):1146, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900983.

Anonymous:2004:BRPb

[Ano04s]  Anonymous.  Book re-

view: Parallel computing on heterogeneous networks: By Alexey Lastovetsky. Wiley–Interscience, Hoboken, NJ. (2003). 423 pages. $89.95. Computers and Mathematics with Applications, 47(6–7):1145, March/ April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210490096X.

Anonymous:2004:BRPd

[Ano04t]  Anonymous.  Book re-

view: Probability and statistics for computer science: By James L. Johnson. Wiley–Interscience, Hoboken, NJ. 760 pages. $98.00. Computers and Mathematics with Applications, 47(8–9):1482–1483, April/ May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901459.

Anonymous:2004:BRPc

[Ano04u]  Anonymous.  Book re-

view: Programming with objects: a comparative presentation of object-oriented programming with C++ and Java: By Avinash C. Kak. Wiley–Interscience, Hoboken, NJ. (2003). 1115 pages. $83.95, £63.50. Computers and Mathematics with Applications, 47(6–7):1150, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210490106X.

Anonymous:2004:BRRb

[Ano04v]  Anonymous.  Book re-

view: Random perturbation methods with applications in science and engineering: By Anatoli V. Skorokhod, Frank C. Hoppensteadt, and Habib Salehi. Springer, Berlin. (2002) 489 Pages $79.95, EUR 84.95, SFR 141.00, GPB 59.50. Computers and Mathematics with Applications, 47(8–9):1478–1479, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901393.

Anonymous:2004:BRRa

[Ano04w] Anonymous. Book review:

Random perturbation methods with applications in Science and Engineering: By Anatoli V. Skorokhod, Frank C. Hoppensteadt and Haib Salehi. Springer, New York. (2002). 488 pages. $79.95. Computers and Mathematics with Applications, 47(6–7):1153–1154, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901095.

Anonymous:2004:BRS

[Ano04x]  Anonymous.  Book re-

view: Selected papers on discrete mathematics: Edited by Donald E. Knuth. CSLI Publications. Chicago, IL. (2003). 812 pages. $32.00. Computers and Mathematics with Applications, 47(8–9):1486, April/ May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901502.

Anonymous:2004:BRSa

[Ano04y]  Anonymous.  Book re-

view: Symmetry and integration methods for differential equations: S By George W. Bluman and Stephen C. Anco. Springer. (2003). 419 pages. $69.95. Computers and Mathematics with Applications, 47 (6–7):1152–1153, March/ April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901083.

Anonymous:2004:BRSb

[Ano04z]  Anonymous.  Book re-

view: Symmetry and integration methods for differential equations: Second Edition. By George W. Bluman and Stephen C. Anco. Springer, Berlin. (2002). 421 pages, $69.95, EUR 74.95, SFR 124.50, GBP 52.50. Computers and Mathematics with Applications, 47(8–9):1478, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901381.

Anonymous:2004:BRMd

[Ano04-27]  Anonymous.  Book re-

view: The macroeconomics of imperfect competition and nonclearing markets:, A dynamic general equilibrium approach: By Jean-Pascal Benassy. The MIT Press, Cambridge, MA. (2002) 272 Pages. $35.00. Computers and Mathematics with Applications, 47(8–9):1479–1480, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210490140X.

Anonymous:2004:BRPa

[Ano04-28] Anonymous. Book review: The parallel brain: the cognitive neuroscience of the corpus callosum: Edited by Eran Zaidel and Marco Iacoboni. The MIT Press, Cambridge, MA. (2003). 551 pages. $95, £63.50. Computers and Mathematics with Applications, 47 (6–7):1144–1145, March/ April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900958.

Anonymous:2004:BRWa

[Ano04-29]  Anonymous.  Book re-

view: The wired homestead: an MIT press sourcebook on the Internet and the family: Edited by Joseph Turow and Andrea L. Kavanaugh. The MIT Press, Cambridge, MA (2003). 502 pages. $39.95. Computers and Mathematics with Applications, 47(8–9):1482, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901447.

Anonymous:2004:BRT

[Ano04-30]  Anonymous.  Book re-

view: Turing (a novel about computation): Edited by Christos H. Papadimitriou. The MIT Press. Cambridge, MA. (2003) 284 pages. $24.95. Computers and Mathematics with Applications, 47(8–9):1486, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901514.

Anonymous:2004:BRV

[Ano04-31] Anonymous. Book review:

Virtual art: From illusion to immersion: By Oliver Grau. Translated by Gloria Custance. The MIT Press, Cambridge, MA. (2003). 416 pages. $45. Computers and Mathematics with Applications, 47(6–7):1150, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901046.

Anonymous:2004:BRWb

[Ano04-32] Anonymous. Book review:

Word sense disambiguation: the case for combinations of knowledge sources: By Mark Stevenson. CLSI Publications, Stanford, CA. (2003). 175 pages. $67.50. Computers and Mathematics with Applications, 47(8–9):1484, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901484.

Anonymous:2004:EBa

[Ano04-33]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 47(1):iii–v, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900004.

Anonymous:2004:EBb

[Ano04-34]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 48(1–2):CO1, iii– v, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104839920.

Anonymous:2004:EBc

[Ano04-35]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 48(5–6):iii–v, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104003323.

Anonymous:2004:EBd

[Ano04-36]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 48(7–8):iii–v, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003621.

Anonymous:2004:EBe

[Ano04-37]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 48(9):iii–v, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104003888.

Anonymous:2004:EBf

[Ano04-38]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 48(10–11):iii–v, November/December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004328.

Anonymous:2004:LC

[Ano04-39] Anonymous. List of contents. Computers and Mathematics with Applications, 48(12):III–XV, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210400464X.

Anonymous:2004:SPSa

[Ano04-40] Anonymous. Some previous special issues of interest to readers. Computers and Mathematics with Applications, 47(12): vii–viii, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840173.

Anonymous:2004:SPSb

[Ano04-41] Anonymous. Some previous special issues of interest to readers. Computers and Mathematics with Applications, 48(1–2):vii– viii, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839932.

Anonymous:2005:BRa

[Ano05a]  Anonymous.  Book re-

ports. Computers and Mathematics with Applications, 49(5–6):953–962, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001045.

Anonymous:2005:BRb

[Ano05b]  Anonymous.  Book re-

ports. Computers and Mathematics with Applications, 49(9–10):1585–1622, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001719.

Anonymous:2005:BRc

[Ano05c] Anonymous. Book reports.

Computers and Mathematics with Applications, 50 (10–12):1801–1824, November/December 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105004281.

Anonymous:2005:EBa

[Ano05d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 49(1):iii–v, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000180.

Anonymous:2005:EBb

[Ano05e]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 50(1–2):iii–v, July 2005. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105002592.

Anonymous:2005:EFG

[Ano05f]  Anonymous.  Errata to

“Fuzzy group based on fuzzy binary operation”. Computers and Mathematics with Applications, 49(11– 12):1951–1952, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105002117. See

[YL04b].

Anonymous:2005:EMC

[Ano05g] Anonymous. Errata to “Maneuvering control of a rolling four-links robot”. Computers and Mathematics with Applications, 50(3–4):653, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002737.

See [FY03a].

Anonymous:2005:LC

[Ano05h] Anonymous. List of contents. Computers and Mathematics with Applications, 49(11–12):III–XIII, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002233.

Anonymous:2006:AI

[Ano06a] Anonymous. Author index. Computers and Mathematics with Applications, 51(12):XIII–XV, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106801951.

Anonymous:2006:BR

[Ano06b]  Anonymous.  Book re-

ports. Computers and Mathematics with Applications, 51(6–7):1149–1162, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000484.

Anonymous:2006:EBa

[Ano06c]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 51(1):iii–v, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004815.

Anonymous:2006:EBb

[Ano06d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 52(1–2):iii–v, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106002252.

Anonymous:2006:EBc

[Ano06e]  Anonymous.  Editorial

board. Computers and Mathematics with Applications, 52(5):iii–v, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210600294X.

Anonymous:2006:EBd

[Ano06f]  Anonymous.  Editorial

board. Computers and Mathematics with Applications, 52(6–7):iii–v, September/October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003063.

Anonymous:2006:EBe

[Ano06g]  Anonymous.  Editorial

board. Computers and Mathematics with Applications, 52(8–9):iii–v, October/November 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0898122106003592.

Anonymous:2006:EBf

[Ano06h]  Anonymous.  Editorial

board. Computers and Mathematics with Applications, 52(10–11):iii–v, November/December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106003737.

Anonymous:2006:EBg

[Ano06i]  Anonymous.  Editorial

board. Computers and Mathematics with Applications, 52(12):iii–v, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000077.

Anonymous:2006:LC

[Ano06j] Anonymous. List of contents. Computers and Mathematics with Applications, 51(12):III–XII, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106001775.

Anonymous:2006:N

[Ano06k] Anonymous. null. Computers and Mathematics with Applications, 52(8–9):xvii– xxi, October/November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106003269.

Anonymous:2006:VCA

[Ano06l]  Anonymous.  Volume

contents and author index. Computers and Mathematics with Applications, 52(12):I–XIII, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000028.

Anonymous:2007:EBa

[Ano07a]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(1):iii–v, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003112.

Anonymous:2007:EBb

[Ano07b]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(2):iii–v, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003203.

Anonymous:2007:EBc

[Ano07c]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(3–4):iii–v, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210700380X.

Anonymous:2007:EBd

[Ano07d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(5):iii–v, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107003902.

Anonymous:2007:EBe

[Ano07e]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(6):iii–v, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004002.

Anonymous:2007:EBf

[Ano07f]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(7):iii–v, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004087.

Anonymous:2007:EBg

[Ano07g]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(8):iii–v, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004178.

Anonymous:2007:EBh

[Ano07h]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(9):iii–v, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004270.

Anonymous:2007:EBi

[Ano07i]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(10):iii–v, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004403.

Anonymous:2007:EBj

[Ano07j]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(11):iii–v, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004464.

Anonymous:2007:EBk

[Ano07k]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 53(12):iii–v, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004579.

Anonymous:2007:EBl

[Ano07l]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 54(1):iii–v, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210700466X.

Anonymous:2007:EBm

[Ano07m]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 54(2):iii–v, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107004750.

Anonymous:2007:EBn

[Ano07n]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 54(3):iii–v, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005214.

Anonymous:2007:EBo

[Ano07o]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 54(4):iii–v, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005329.

Anonymous:2007:EBp

[Ano07p]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 54(5):iii–v, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107005445.

Anonymous:2007:EBq

[Ano07q]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 54(6):iii–v, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210700613X.

Anonymous:2007:EBr

[Ano07r]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 54(7–8):iii–v, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006748.

Anonymous:2007:EBs

[Ano07s]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 54(9–10):iii–v, November 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107007158.

Anonymous:2007:EBt

[Ano07t]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 54(11–12):iii–v, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007250.

Anonymous:2008:BR

[Ano08a] Anonymous. Book reports. Computers and Mathematics with Applications, 56(2): 583–589, July 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107007468.

Anonymous:2008:EBa

[Ano08b]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(1):iii–v, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007900.

Anonymous:2008:EBb

[Ano08c]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(2):iii–v, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107008000.

Anonymous:2008:EBc

[Ano08d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(3):iii–v, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107008401.

Anonymous:2008:EBd

[Ano08e]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(4):iii–v, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107008504.

Anonymous:2008:EBe

[Ano08f]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(5):iii–v, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108000370.

Anonymous:2008:EBf

[Ano08g]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(6):iii–v, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108000643.

Anonymous:2008:EBg

[Ano08h]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(7):iii–v, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108001302.

Anonymous:2008:EBh

[Ano08i]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(8):iii–v, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108001430.

Anonymous:2008:EBi

[Ano08j]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(9):iii–v, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108001995.

Anonymous:2008:EBj

[Ano08k]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(10):iii–v, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108002071.

Anonymous:2008:EBk

[Ano08l]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(11):iii–v, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108002216.

Anonymous:2008:EBl

[Ano08m]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 55(12):iii–v, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108002320.

Anonymous:2008:EBm

[Ano08n]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(1):iii–v, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108002411.

Anonymous:2008:EBn

[Ano08o]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(2):iii–v, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108002848.

Anonymous:2008:EBo

[Ano08p]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(3):iii–v, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002927.

Anonymous:2008:EBp

[Ano08q]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(4):iii–v, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003088.

Anonymous:2008:EBq

[Ano08r]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(5):iii–v, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108003209.

Anonymous:2008:EBr

[Ano08s]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(6):iii–v, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108003313.

Anonymous:2008:EBs

[Ano08t]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(7):iii–v, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004227.

Anonymous:2008:EBt

[Ano08u]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(8):iii–v, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004690.

Anonymous:2008:EBu

[Ano08v]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(9):iii–v, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108005014.

Anonymous:2008:EBv

[Ano08w]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(10):iii–v, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108005117.

Anonymous:2008:EBw

[Ano08x]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(11):iii–v, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005993.

Anonymous:2008:EBx

[Ano08y]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 56(12):iii–v, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006548.

Anonymous:2009:EBa

[Ano09a]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(1):iii–v, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006925.

Anonymous:2009:EBb

[Ano09b]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(2):iii–v, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108007013.

Anonymous:2009:EBc

[Ano09c]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(3):iii–v, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108007128.

Anonymous:2009:EBd

[Ano09d]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(4):iii–v, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109000066.

Anonymous:2009:EBe

[Ano09e]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(5):iii–v, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109000145.

Anonymous:2009:EBf

[Ano09f]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(6):iii–v, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109000789.

Anonymous:2009:EBg

[Ano09g]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(7):iii–v, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109001175.

Anonymous:2009:EBh

[Ano09h]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(8):iii–v, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109001254.

Anonymous:2009:EBi

[Ano09i]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(9):iii–v, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109002193.

Anonymous:2009:EBj

[Ano09j]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(10):iii–v, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109002715.

Anonymous:2009:EBk

[Ano09k]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 57(11–12):iii–v, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002995.

Anonymous:2009:EBl

[Ano09l]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(1):iii–v, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109003137.

Anonymous:2009:EBm

[Ano09m]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(2):iii–v, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109003393.

Anonymous:2009:EBn

[Ano09n]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(3):iii–v, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003666.

Anonymous:2009:EBo

[Ano09o]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(4):iii–v, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004064.

Anonymous:2009:EBp

[Ano09p]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(5):iii–v, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109004350.

Anonymous:2009:EBq

[Ano09q]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(6):iii–v, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109005185.

Anonymous:2009:EBr

[Ano09r]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(7):iii–v, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005288.

Anonymous:2009:EBs

[Ano09s]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(8):iii–v, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109006233.

Anonymous:2009:EBt

[Ano09t]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(9):iii–v, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109006646.

Anonymous:2009:EBu

[Ano09u]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(10):iii–v, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109006841.

Anonymous:2009:EBv

[Ano09v]  Anonymous.  Editorial

Board. Computers and Mathematics with Applications, 58(11–12):iii–v, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109007007.

Antczak:2005:PIM

[Ant05] T. Antczak. γ-preinvexity and γ-invexity in mathematical programming. Computers and Mathematics with Applications, 50(3–4):551– 566, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002701.

Antczak:2008:GFM

[Ant08] Tadeusz Antczak. Generalized fractional minimax programming with B–(p, r)- invexity. Computers and Mathematics with Applications, 56(6):1505–1525, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001776.

Ashino:2000:BBM

[ANV00] R. Ashino, M. Nagase, and R. Vaillancourt. Behind and beyond the Matlab ODE suite. Computers and Mathematics with Applications, 40(4–5):491–512, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001759.

Atabaigi:2009:NLC

[ANZ09] Ali Atabaigi, Nemat Nyamoradi, and Hamid R. Z. Zangeneh. The number of limit cycles of a quintic polynomial system. Computers and Mathematics with Applications, 57(4):677–684, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006214.

Agarwal:2001:FPTb

[AO01a] R. P. Agarwal and D. O’Regan. Fixed-point theorems for countably condensing maps on Fréchet spaces. Computers and Mathematics with Applications, 42(6– 7):909–916, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002073.

Agarwal:2001:FPTa

[AO01b] R. P. Agarwal and D. O’Regan. Fixed-point theory for set valued mappings between topological vector spaces having sufficiently many linear functions. Computers and Mathematics with Applications, 41(7–8):917– 928, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003291.

Agarwal:2001:NUD

[AO01c] R. P. Agarwal and D. O’Regan.

Nonlinear Urysohn discrete equations on the infinite interval: a fixed-point approach. Computers and Mathematics with Applications, 42(3–5):273–281, August/September 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101001523.

Agarwal:2001:P

[AO01d] RaviP. Agarwal and Donal O’Regan. Preface. Computers and Mathematics with Applications, 42(3–5):xv, August/September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101001511.

Ashyralyev:2005:SDS

[AO05] A. Ashyralyev and Y. Ozdemir.

Stability of difference schemes for hyperbolic-parabolic equations. Computers and Mathematics with Applications, 50(8–9):1443–1476, October/November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105003937.

Arikoglu:2008:SII

[AO08a] Aytac Arikoğlu and Ibrahim Ozkol. Solutions of integral and integro-differential equation systems by using differential transform method. Computers and Mathematics with Applications, 56(9):2411–2417, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003805.

Atlihan:2008:SPP

[AO08b] Ö. G. Atlihan and C. Orhan. Summation process of positive linear operators. Computers and Mathematics with Applications, 56(5): 1188–1195, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210800120X.

Asinari:2009:CBK

[AO09] Pietro Asinari and Taku Ohwada. Connection between kinetic methods for fluid-dynamic equations and macroscopic finitedifference schemes. Computers and Mathematics with Applications, 58(5):841–861, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900090X.

Agarwal:2004:GUL

[AOLL04] R. P. Agarwal, D. O’Regan, V. Lakshmikantham, and S. Leela. A generalized upper and lower solution method for singular initial value problems. Computers and Mathematics with Applications, 47 (4–5):739–750, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900612.

Agarwal:2003:TPH

[AORS03] R. P. Agarwal, D. O’Regan, I. Rachunková, and S. Stanek. Two-point higher-order BVPs with singularities in phase variables. Computers and Mathematics with Applications, 46(12):1799–1826, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902380.

Altintas:2004:NCF

[AÖS04] O. Altintas, Ö. Özkan, and H. M. Srivastava. Neighborhoods of a certain family of multivalent functions with negative coefficients. Computers and Mathematics with Applications, 47(10–11):1667–1672, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104839804.

Agarwal:2007:PDC

[AOS07] Ravi P. Agarwal, Donal O’Regan, and Svatoslav Stanek. Positive and dead core solutions of singular Dirichlet boundary value problems with ϕ-Laplacian. Computers and Mathematics with Applications, 54(2):255– 266, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002118.

Agarwal:2007:CSS

[AOTW07] Ravi P. Agarwal, Donal O’Regan, Christopher C. Tisdell, and Patricia J. Y. Wong. Constant-sign solutions of a system of Volterra integral equations. Computers and Mathematics with Applications, 54(1):58– 75, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001782.

Aouf:2008:CSM

[Aou08] M. K. Aouf. Certain subclasses of meromorphically multivalent functions associated with generalized hypergeometric function. Computers and Mathematics with Applications, 55 (3):494–509, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004865.

Agarwal:2005:CSP

[AOW05] R. P. Agarwal, D. O’Regan, and Patricia J. Y. Wong. Constant-sign periodic and almost periodic solutions of a system of difference equations. Computers and Mathematics with Applications, 50(10– 12):1725–1754, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004219.

Avery:2001:TPF

[AP01] R. I. Avery and A. C. Peterson. Three positive fixed points of nonlinear operators on ordered Banach spaces. Computers and Mathematics with Applications, 42(3–5):313–322, August/September 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101001560.

Aulbach:2003:IMA

[AP03] B. Aulbach and C. Pötzsche. Invariant manifolds with asymptotic phase for nonautonomous difference equations. Computers and Mathematics with Applications, 45(6–9):1385–1398, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000932.

Aljinovic:2004:DWM

[AP04] A. A. Aljinović and J. Pecarić. Discrete weighted Montgomery identity and discrete Ostrowski type inequalities. Computers and Mathematics with Applications, 48(5–6):731–745, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003098.

Abelman:2008:CSR

[AP08] Shirley Abelman and Kailash C. Patidar. Comparison of some recent numerical methods for initial-value problems for stiff ordinary differential equations. Computers and Mathematics with Applications, 55 (4):733–744, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107005056.

Artalejo:2009:ICB

[AP09] J. R. Artalejo and V. Pla.

On the impact of customer balking, impatience and retrials in telecommunication systems. Computers and Mathematics with Applications, 57(2):217–229, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006354.

Arada:2007:VEF

[APS07] Nadir Arada, Marilia Pires, and Adélia Sequeira. Viscosity effects on flows of generalized Newtonian fluids through curved pipes. Computers and Mathematics with Applications, 53(3– 4):625–646, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107001460.

Arato:2002:EMM

[APV02] M. Arató, G. Pap, and K. Varga. Estimation of the mean of multivariate AR processes. Computers and Mathematics with Applications, 43(6–7):707–719, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101003157.

Arato:2001:AIS

[APvZ01] M. Arató, G. Pap, and M. C. A. van Zuijlen. Asymptotic inference for spatial autoregression and orthogonality of Ornstein– Uhlenbeck sheets. Computers and Mathematics with Applications, 42(1–2):219– 229, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001468.

Anguelov:2007:SLC

[AR07] Roumen Anguelov and Elemér E. Rosinger. Solving large classes of nonlinear systems of PDEs. Computers and Mathematics with Applications, 53(3–4):491– 507, February 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107001800.

Adivar:2009:SPD

[AR09] Murat Adivar and Youssef N. Raffoul. Stability and periodicity in dynamic delay equations. Computers and Mathematics with Applications, 58(2):264– 272, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002235.

See note [AR10].

Adivar:2010:NSP

[AR10] Murat Adivar and Youssef N.

Raffoul. A note on “Stability and periodicity in dynamic delay equations” [Comput. Math. Appl. 58 (2009) 264–272]. Computers and Mathematics with Applications, 59(10):3351– 3354, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122110002014.

See [AR09].

Arany:2001:NES

[Ara01] I. Arany. Numerical experiences of solving elasticity systems by PCG methods. Computers and Mathematics with Applications, 42 (8–9):1025–1033, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002188.

Argyros:2000:LCI

[Arg00a] I. K. Argyros. Local convergence of inexact Newtonlike-iterative methods and applications. Computers and Mathematics with Applications, 39(1–2):69–75, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003144.

Argyros:2000:NMB

[Arg00b] I. K. Argyros. Newton methods on Banach spaces with a convergence structure and applications. Computers and Mathematics with Applications, 40(1):37– 48, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001383.

Ariel:2007:AFD

[Ari07a]  P. Donald Ariel.  Ax-

isymmetric flow due to a stretching sheet with partial slip. Computers and Mathematics with Applications, 54(7–8):1169–1183, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002866.

Ariel:2007:TDF

[Ari07b]  P. Donald Ariel.  The

three-dimensional flow past a stretching sheet and the homotopy perturbation method. Computers and Mathematics with Applications, 54(7–8):920–925, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002532.

See [EM09b] and reply [Ari09c].

Ariel:2009:EHP

[Ari09a]  P. Donald Ariel.  Ex-

tended homotopy perturbation method and computation of flow past a stretching sheet. Computers and Mathematics with Applications, 58(11–12):2402–2409, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001680.

Ariel:2009:HPM

[Ari09b] P. Donald Ariel. The homotopy perturbation method and analytical solution of the problem of flow past a rotating disk. Computers and Mathematics with Applications, 58(11–12):2504– 2513, December 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109001953.

Ariel:2009:RCT

[Ari09c]  P. Donald Ariel.  Re-

ply to the comment on “The three-dimensional flow past a stretching sheet and the homotopy perturbation method” by P. D. Ariel, Computers and Mathematics with Applications 54 (2007) 920–925 by El-Mistikawy. Computers and Mathematics with Applications, 57(3):407–409, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005580.

See [Ari07b, EM09b].

Alonso:2008:OCN

[ARM08] María Alonso and Luis Rodríguez-Marín. Optimality conditions for a nonconvex set-valued optimization problem. Computers and Mathematics with Applications, 56(1):82– 89, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008280.

Arora:2008:NPS

[Aro08] Rajan Arora. Non-planar shock waves in a magnetic field. Computers and Mathematics with Applications, 56(10):2686–2691, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004082.

Al-Radadi:2002:FMS

[ARS02] E. Al-Radadi and P. Siy.

Four-moduli set (2, 2^{n} − 1, 2^{n} + 2^{n−1}1, 2^{n+1} + 2^{n} − 1) simplifies the residue to binary converters based on CRT II. Computers and Mathematics with Applications, 44(12):1581–1587, December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002808.

Al-Radadi:2003:RSD

[ARS03] E. Al-Radadi and P. Siy.

RNS: sign detector based on Chinese Remainder Theorem II (CRT II). Computers and Mathematics with Applications, 46(10– 11):1559–1570, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390191X.

Al-Said:2001:UCS

[AS01]  E. A. Al-Said.  The

use of cubic splines in the numerical solution of a system of second-order boundary value problems. Computers and Mathematics with Applications, 42 (6–7):861–869, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002048.

Actis:2003:ABF

[AS03a] R. L. Actis and B. A. Szabo. Analysis of bonded and fastened repairs by the p-version of the finiteelement method. Computers and Mathematics with Applications, 46(1):1– 14, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900770.

Agarwal:2003:NSS

[AS03b] R. P. Agarwal and S. Stanek. Nonnegative solutions of singular boundary value problems with sign changing nonlinearities. Computers and Mathematics with Applications, 46(12):1827– 1837, December 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103902392.

Abtahi:2005:CFR

[AS05] M. Abtahi and P. Siy. Core function of an RNS number with no ambiguity. Computers and Mathematics with Applications, 50(3–4):459– 470, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002890.

Abtahi:2007:FSD

[AS07] M. Abtahi and P. Siy. The factor-2 sign detection algorithm using a core function for RNS numbers. Computers and Mathematics with Applications, 53(9):1455– 1463, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001150.

Ashyralyev:2008:NBV

[AS08] Allaberen Ashyralyev and Ali Sirma. Nonlocal boundary value problems for the Schrödinger equation. Computers and Mathematics with Applications, 55 (3):392–407, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107003720.

Abboud:2009:UDT

[AS09a] Hyam Abboud and Toni Sayah. Upwind discretization of a time-dependent two-dimensional grade-two fluid model. Computers and Mathematics with Applications, 57(8):1249– 1264, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000510.

Anwar:2009:RST

[AS09b] Shadab Anwar and Michael C. Sukop. Regional scale transient groundwater flow modeling using lattice Boltzmann methods. Computers and Mathematics with Applications, 58(5):1015–1023, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001060.

Abu-Saris:2008:AS

[ASÇY08] R. Abu-Saris, C. Çinar, and I. Yalçinkaya. On the asymptotic stability of x_{n+1} = (a + x_{n}x_{n−k})/(x_{n} + x_{n−k}). Computers and Mathematics with Applications, 56(5):1172–1175, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001181.

Al-Seedy:2009:TSM

[ASESESA09] R. O. Al-Seedy, A. A. El-Sherbiny, S. A. El-Shehawy, and S. I. Ammar. Transient solution of the M/M/c queue with balking and reneging. Computers and Mathematics with Applications, 57(8):1280– 1285, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000376.

Ashordia:2005:LSS

[Ash05] M. Ashordia. Lyapunov stability of systems of linear generalized ordinary differential equations. Computers and Mathematics with Applications, 50(5–6):957–982, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003287.

Ahmad:2008:UAF

[ASHA08] I. Ahmad, M. Sajid, T. Hayat, and M. Ayub. Unsteady axisymmetric flow of a secondgrade fluid over a radially stretching sheet. Computers and Mathematics with Applications, 56(5):1351–1357, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001582.

Asinari:2008:AAM

[Asi08] Pietro Asinari. Asymptotic analysis of multiplerelaxation-time lattice Boltzmann schemes for mixture modeling. Computers and Mathematics with Applications, 55(7):1392– 1407, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700630X.

Ai:2006:DFE

[ASL06] X. Ai, Y. Shu, and B. Q. Li. Discontinuous finite-element phase-field modeling of polycrystalline grain growth with convection. Computers and Mathematics with Applications, 52(5):721–734, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002513.

Aslan:2011:CAE

[Asl11] Ismail Aslan. Comment on: “Application of Expfunction method for (3 + 1)- dimensional nonlinear evolution equations” [Comput. Math. Appl. 56 (2008) 1451–1456]. Computers and Mathematics with Applications, 61(6):1700–1703, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/S089812211100071X.

See [BB08b].

An:2009:CCO

[ASM09] Wei An, Fang-Ming Shao, and Huajun Meng. The coverage-control optimization in sensor network subject to sensing area. Computers and Mathematics with Applications, 57(4): 529–539, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108006767.

Aouf:2008:SFL

[ASS08] M. K. Aouf, H. Silverman, and H. M. Srivastava. Some families of linear operators associated with certain subclasses of multivalent functions. Computers and Mathematics with Applications, 55(3):535–549, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107004890.

Ayanlar:2002:OTN

[AT02] B. Ayanlar and A. Tiryaki.

Oscillation theorems for nonlinear second-order differential equations. Computers and Mathematics with Applications, 44(3–4):529– 538, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001670.

Akiyama:2005:TSF

[AT05] Shigeki Akiyama and J. M. Thuswaldner. The topological structure of fractal tilings generated by quadratic number systems. Computers and Mathematics with Applications, 49(9– 10):1439–1485, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122105001628.

Atallah:2009:EPD

[Ata09] M. A. Atallah. Erratum to: “Probability distributions conditioned by the available information: Gamma Distribution and Moments” [Comput. Math. Appl. 52 (2006) 289–304]. Computers and Mathematics with Applications, 58(9):1878–1886, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004611.

See [Kum06].

Agarwal:2003:DEB

[ATT03a] R. P. Agarwal, H. B. Thompson, and C. C. Tisdell. Difference equations in Banach spaces. Computers and Mathematics with Applications, 45 (6–9):1437–1444, March/ May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001007.

Agarwal:2003:TPB

[ATT03b] R. P. Agarwal, H. B.

Thompson, and C. C. Tisdell. Three-point boundary value problems for second-order discrete equations. Computers and Mathematics with Applications, 45(6–9):1429–1435, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000981.

Akhmet:2008:ABL

[ATY08] M. U. Akhmet, M. A.

Tleubergenova, and O. Yilmaz. Asymptotic behavior of linear impulsive integro-differential equations. Computers and Mathematics with Applications, 56(4):1071–1081, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001090.

Arnold:2003:EDD

[AU03] A. Arnold and A. Unterreiter. Entropy decay of discretized Fokker–Planck equations. I — temporal semidiscretization. Computers and Mathematics with Applications, 46(10– 11):1683–1690, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902021.

Altintan:2009:VIM

[AU09] D. Altintan and Ö. Ugur. Variational iteration method for Sturm–Liouville differential equations. Computers and Mathematics with Applications, 58(2):322– 328, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002326.

Avudainayagam:2000:DDM

[AV00] A. Avudainayagam and C. Vani. A domain decomposition method for biharmonic equation. Computers and Mathematics with Applications, 40(6– 7):865–876, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002029.

Artikis:2001:PAR

[AVA01] T. P. Artikis, A. P.

Voudouri, and P. T. Artikis. Properties and applications in risk frequency reduction operations of an integral part model. Computers and Mathematics with Applications, 42(1–2):211– 218, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001456.

Aranda:2008:MMT

[AVAGP08] Diego F. Aranda, Rafael J.

Villanueva, Abraham J. Arenas, and Gilberto C. González-Parra. Mathematical modeling of toxoplasmosis disease in varying size populations. Computers and Mathematics with Applications, 56(3):690–696, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800059X.

Arnout:2008:LBM

[AVBW08] S. Arnout, F. Verhaeghe, B. Blanpain, and P. Wollants. Lattice Boltzmann model for diffusioncontrolled indirect dissolution. Computers and Mathematics with Applications, 55(7):1377–1391, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006293.

Abbasbandy:2004:NMF

[AVLPN04] S. Abbasbandy, T. A. Viranloo, Ó. López-Pouso, and J. J. Nieto. Numerical methods for fuzzy differential inclusions. Computers and Mathematics with Applications, 48(10– 11):1633–1641, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210400416X.

Akyildiz:2008:FGD

[AVO08] F. Talay Akyildiz, K. Vajravelu, and H. Ozekes. Fourier–Galerkin domain truncation method for Stokes’ first problem with Oldroyd four-constant liquid. Computers and Mathematics with Applications, 55(11):2452–2457, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007006.

Agarwal:2001:SRD

[AW01] R. P. Agarwal and Yuan-Ming Wang. Some recent developments of Numerov’s method. Computers and Mathematics with Applications, 42(3–5):561–592, August/September 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210100178X.

Anderson:2009:PSS

[AW09] Douglas R. Anderson and Patricia J. Y. Wong. Positive solutions for secondorder semipositone problems on time scales. Computers and Mathematics with Applications, 58(2): 281–291, July 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109002338.

Ashyralyev:2006:SDS

[AY06] A. Ashyralyev and H. A.

Yurtsever. The stability of difference schemes of second-order of accuracy for hyperbolic-parabolic equations. Computers and Mathematics with Applications, 52(3–4):259–268, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001945.

Aouf:2009:CCM

[AY09] M. K. Aouf and M. F. Yassen. On certain classes of meromorphically multivalent functions associated with the generalized hypergeometric function. Computers and Mathematics with Applications, 58(3):449–463, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003253.

Aykut:2004:EPS

[Ayk04] N. Aykut. Existence of positive solutions for boundary value problems of secondorder functional difference equations. Computers and Mathematics with Applications, 48(3–4):517–527, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840732.

Agarwal:2001:PSD

[AZ01] R. P. Agarwal and Weinian Zhang. Periodic solutions of difference equations with general periodicity. Computers and Mathematics with Applications, 42(3–5):719–727, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001912.

Al-Zoubi:2008:SFF

[AZB08] Ahmad Al-Zoubi and Gunther Brenner. Simulating fluid flow over sinusoidal surfaces using the lattice Boltzmann method. Computers and Mathematics with Applications, 55 (7):1365–1376, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107006426.

Agarwal:2005:ZMT

[AZCK05] R. P. Agarwal, Haiyun Zhou, Yeol Je Cho, and Shin Min Kang. Zeros and mapping theorems for perturbations of maccretive operators in Banach spaces. Computers and Mathematics with Applications, 49(1):147–155, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000131.

Balachandran:2001:CSO

[BA01] K. Balachandran and S. Marshal Anthoni. Controllability of second-order semilinear neutral functional differential systems in Banach spaces. Computers and Mathematics with Applications, 41(10–11):1223–1235, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101000931.

Beg:2008:CMM

[BA08] Ismat Beg and Mujahid Abbas. Convergence of multivalued mappings via approximable measurable selectors. Computers and Mathematics with Applications, 55(12):2740–2745, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007730.

Biazar:2009:EEF

[BA09a] J. Biazar and Z. Ayati. Extension of the Exp-function method for systems of twodimensional Burgers equations. Computers and Mathematics with Applications, 58(11–12):2103–2106, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001333.

Biazar:2009:NTS

[BA09b] Jafar Biazar and Hossein Aminikhah. A new technique for solving nonlinear integral-differential equations. Computers and Mathematics with Applications, 58(11–12):2084–2090, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001308.

Biazar:2009:SCH

[BA09c] Jafar Biazar and Hossein Aminikhah. Study of convergence of homotopy perturbation method for systems of partial differential equations. Computers and Mathematics with Applications, 58(11–12):2221–2230, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001515.

Babovsky:2009:NMB

[Bab09] Hans Babovsky. A numerical model for the Boltzmann equation with applications to micro flows. Computers and Mathematics with Applications, 58(4):791–804, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003083.

Baesu:2007:CES

[Bae07] Eveline Baesu. A class of exact solutions for a 3D filamentary continuous medium. Computers and Mathematics with Applications, 53(2):317–323, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700199X.

Bai:2000:CAP

[Bai00] Zhong-Zhi Bai. A class of asynchronous parallel iterations for the systems of nonlinear algebraic equations. Computers and Mathematics with Applications, 39(7–8):81– 94, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000675.

Bai:2001:PIP

[Bai01] Min-Ru Bai. Perturbed iterative process for fixed points of multivalued ϕhemicontractive mappings in Banach spaces. Computers and Mathematics with Applications, 41(1–2):103– 109, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101850100.

Bai:2008:TPS

[Bai08] Chuanzhi Bai. Triple positive solutions of three-point boundary value problems for fourth-order differential equations. Computers and Mathematics with Applications, 56(5):1364–1371, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001600.

Balasubramaniam:2009:FGA

[BAK09] P. Balasubramaniam, M. Syed Ali, and J. H. Kim. Faedo–Galerkin approximate solutions for stochastic semilinear integrodifferential equations. Computers and Mathematics with Applications, 58(1):48– 57, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002466.

Balandin:2000:TAS

[Bal00] A. L. Balandin. Tomographic analysis of signaltering functions by maximum entropy method. Computers and Mathematics with Applications, 39(9–10): 15–24, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000833.

Baltensperger:2002:SRL

[Bal02] R. Baltensperger. Some results on linear rational trigonometric interpolation. Computers and Mathematics with Applications, 43(6–7):737–746, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101003170.

Balazs:2004:WSA

[Bal04] K. Balázs. Weighted simultaneous approximation by Lagrange interpolation on the real line. Computers and Mathematics with Applications, 48(9):1345–1349, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003797.

Banas:2000:CSS

[Ban00] K. Banaś. Convergence to steady-state solutions for stabilized finite element simulations of compressible flows. Computers and Mathematics with Applications, 40(4–5):625–643, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000184X.

Banerjee:2007:WSS

[Ban07] Abhijit Banerjee. Weighted sharing of a small function by a meromorphic function and its derivative. Computers and Mathematics with Applications, 53(11):1750– 1761, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001484.

Bao:2003:ACM

[Bao03] Weizhu Bao. Approximation and comparison for motion by mean curvature with intersection points. Computers and Mathematics with Applications, 46 (8–9):1211–1228, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902136.

Baran:2001:NCE

[Bar01a] S. Baran. A new consistent estimator for linear errors-in-variables models. Computers and Mathematics with Applications, 41 (7–8):821–833, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100003229.

Barnabei:2001:PSI

[Bar01b] M. Barnabei. Polynomial sequences of integral type and recursive matrices. Computers and Mathematics with Applications, 41(9):1125– 1141, May 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101000864.

Barnabei:2001:P

[Bar01c] Marilena Barnabei. Preface. Computers and Mathematics with Applications, 41 (9):xiv, May 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101000827.

Barrio:2001:SPA

[Bar01d] R. Barrio. Stability of parallel algorithms to evaluate Chebyshev series. Computers and Mathematics with Applications, 41 (10–11):1365–1377, May/ June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210100102X.

Bataller:2007:EHS

[Bat07] Rafael Cortell Bataller. Effects of heat source/sink, radiation and work done by deformation on flow and heat transfer of a viscoelastic fluid over a stretching sheet. Computers and Mathematics with Applications, 53(2):305–316, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001824.

Baxter:2002:PCG

[Bax02] B. J. C. Baxter. Preconditioned conjugate gradients, radial basis functions, and Toeplitz matrices. Computers and Mathematics with Applications, 43(3–5):305–318, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002887.

Baxter:2006:SRB

[Bax06] B. J. C. Baxter. Scaling radial basis functions via Euclidean distance matrices. Computers and Mathematics with Applications, 51(8):1163–1170, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000812.

Bishwal:2001:RCA

[BB01] J. P. N. Bishwal and A. Bose. Rates of convergence of approximate maximum likelihood estimators in the Ornstein– Uhlenbeck process. Computers and Mathematics with Applications, 42(1–2):23– 38, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001274.

Benouaz:2004:LSA

[BB04] T. Benouaz and F. Bendahmane. Least-square approximation of a nonlinear O.D.E. with excitation. Computers and Mathematics with Applications, 47(2–3):473–489, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900399.

Berezansky:2006:MGE

[BB06] L. Berezansky and E. Braverman. Mackey-glass equation with variable coefficients. Computers and Mathematics with Applications, 51(1):1–16, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004414.

Baksalary:2008:LCT

[BB08a] Oskar Maria Baksalary and Julio Benítez. On linear combinations of two commuting hypergeneralized projectors. Computers and Mathematics with Applications, 56(10):2481–2489, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003842.

Boz:2008:AEF

[BB08b] Ahmet Boz and Ahmet Bekir. Application of Expfunction method for (3 + 1)- dimensional nonlinear evolution equations. Computers and Mathematics with Applications, 56(5):1451–1456, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001703.

See [Asl11].

Bekir:2009:AHE

[BB09a] Ahmet Bekir and Ahmet Boz. Application of He’s Exp-function method for nonlinear evolution equations. Computers and Mathematics with Applications, 58(11–12):2286–2293, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001552.

Bellomo:2009:DMH

[BB09b] N. Bellomo and A. Bellouquid. On the derivation of macroscopic hyperbolic equations for binary multicellular growing mixtures. Computers and Mathematics with Applications, 57(5):744–756, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005853.

Biazar:2009:AHP

[BBA09] J. Biazar, F. Badpeima, and F. Azimi. Application of the homotopy perturbation method to Zakharov–Kuznetsov equations. Computers and Mathematics with Applications, 58(11–12):2391–2394, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001746.

Bongue-Boma:2008:LVB

[BBB08] Malika Bongué-Boma and Maurizio Brocato. Liquids with vapour bubbles. Computers and Mathematics with Applications, 55(2):268–284, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003537.

Barnett:2002:OIV

[BBCD02] N. S. Barnett, C. Buse, P. Cerone, and S. S. Dragomir. Ostrowski’s inequality for vector-valued functions and applications. Computers and Mathematics with Applications, 44 (5–6):559–572, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102001712.

Belley:2009:NSK

[BBCE09] J.-M. Belley, P. Belley, F. Colin, and R. Egli. Nonsmooth kernels for meshfree methods in fluid dynamics. Computers and Mathematics with Applications, 58(6):1253–1272, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003617.

Balbas:2008:DRM

[BBGG08] Alejandro Balbás, Beatriz Balbás, Inna Galperin, and Efim Galperin. Deterministic regression model and visual basic code for optimal forecasting of financial time series. Computers and Mathematics with Applications, 56(10):2757–2771, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004276.

Bellomo:2007:MMD

[BBH07] Nicola Bellomo, Abdelghani Bellouquid, and Miguel A. Herrero. From microscopic to macroscopic description of multicellular systems and biological growing tissues. Computers and Mathematics with Applications, 53(3–4):647–663, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001198.

Berezansky:2005:DDE

[BBI05] L. Berezansky, E. Braverman, and L. Idels. Delay differential equations with Hill’s type growth rate and linear harvesting. Computers and Mathematics with Applications, 49(4):549–563, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000490.

Barrio:2005:VFT

[BBL05] R. Barrio, F. Blesa, and M. Lara. VSVO formulation of the Taylor method for the numerical solution of ODEs. Computers and Mathematics with Applications, 50(1–2):93– 111, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002348.

Bersani:2008:DSM

[BBM08] A. M. Bersani, E. Bersani, and L. Mastroeni. Deterministic and stochastic models of enzymatic networks-applications to pharmaceutical research. Computers and Mathematics with Applications, 55 (5):879–888, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107005603.

Berezansky:2009:NMA

[BBP09] Leonid Berezansky, Elena Braverman, and Sandra Pinelas. On nonoscillation of mixed advanceddelay differential equations with positive and negative coefficients. Computers and Mathematics with Applications, 58(4):766–775, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003046.

Benko:2008:NMS

[BBRS08] David Benko, Daniel C. Biles, Mark P. Robinson, and John S. Spraker. Nyström methods and singular second-order differential equations. Computers and Mathematics with Applications, 56(8):1975–1980, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002757.

Barcelos:2000:HFR

[BC00] C. A. Z. Barcelos and Y. Chen. Heat flows and related minimization problem in image restoration. Computers and Mathematics with Applications, 39 (5–6):81–97, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100000481.

Bohner:2001:MMM

[BC01] M. Bohner and J. E. Castillo. Mimetic methods on measure chains. Computers and Mathematics with Applications, 42(3–5):705–710, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001894.

Berriochoa:2002:SAI

[BC02] E. Berriochoa and A. Cachafeiro.

Strong asymptotics inside the unit disk for Sobolev orthogonal polynomials. Computers and Mathematics with Applications, 44(1–2):253–261, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001451.

Berriochoa:2003:EID

[BC03] E. Berriochoa and A. Cachafeiro.

Extension inside the disk of asymptotics for Sobolev orthogonal polynomials. Computers and Mathematics with Applications, 46 (8–9):1263–1272, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902173.

Bonzani:2008:MSM

[BC08] I. Bonzani and L. M. Gramani Cumin. Modelling and simulations of multilane traffic flow by kinetic theory methods. Computers and Mathematics with Applications, 56(9):2418–2428, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003817.

Berriochoa:2009:CMU

[BCA09] E. Berriochoa, A. Cachafeiro, and J. García Amor. Characterizing the measures on the unit circle with exact quadrature formulas in the space of polynomials. Computers and Mathematics with Applications, 58(7):1370–1382, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004696.

Baralis:2008:ASR

[BCD08] Elena Baralis, Silvia Chiusano, and Riccardo Dutto. Applying sequential rules to protein localization prediction. Computers and Mathematics with Applications, 55(5):867–878, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S089812210700572X.

Barnett:2002:CTI

[BCDF02] N. S. Barnett, P. Cerone, S. S. Dragomir, and A. M. Fink. Comparing two integral means for absolutely continuous mappings whose derivatives are in L_{∞}[a, b] and applications. Computers and Mathematics with Applications, 44(1–2):241– 251, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210200144X.

Barnett:2009:OTT

[BCDS09] N. S. Barnett, W.-S. Cheung, S. S. Dragomir, and A. Sofo. Ostrowski and trapezoid type inequalities for the Stieltjes integral with Lipschitzian integrands or integrators. Computers and Mathematics with Applications, 57(2):195–201, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006391.

Berriochoa:2005:ASS

[BCGA05] E. Berriochoa, A. Cachafeiro, and J. Garcia-Amor. Asymptotics on the support for Sobolev orthogonal polynomials on a bounded interval. Computers and Mathematics with Applications, 50(3–4):381–391, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002828.

Berriochoa:2008:ASA

[BCGA08] E. Berriochoa, A. Cachafeiro, and J. García-Amor. About a system of anti-periodic trigonometric functions. Computers and Mathematics with Applications, 56(6): 1526–1537, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108001788.

Benzi:2003:NEN

[BCGP03] R. Benzi, C. M. Casciola, P. Gualtieri, and R. Piva. Numerical evidence of a new similarity law in shear dominated flows. Computers and Mathematics with Applications, 46(4):617–631, August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900204.

Ballester:2008:EMO

[BCJ08] Cristina Ballester, Rafael Company, and Lucas Jódar. An efficient method for option pricing with discrete dividend payment. Computers and Mathematics with Applications, 56(3):822–835, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000801.

Bretti:2004:LTE

[BCR04] G. Bretti, C. Cesarano, and P. E. Ricci. Laguerretype exponentials and generalized Appell polynomials. Computers and Mathematics with Applications, 48(5–6):833–839, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003177.

Banas:2005:MSC

[BCRS05] J. Banaś, J. Caballero, J. Rocha, and K. Sadarangani. Monotonic solutions of a class of quadratic integral equations of Volterra type. Computers and Mathematics with Applications, 49(5–6):943–952, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001033.

Belgacem:2003:MFE

[BCS03] F. Ben Belgacem, L. K.

Chilton, and P. Seshaiyer. The hp-mortar finite-element method for the mixed elasticity and Stokes problems. Computers and Mathematics with Applications, 46(1): 35–55, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900794.

Brunetto:2000:AAI

[BCT00] M. A. O. Camargo Brunetto, D. M. Claudio, and V. Trevisan. An algebraic algorithm to isolate complex polynomial zeros using Sturm sequences. Computers and Mathematics with Applications, 39(3–4):95– 105, February 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122199003363.

Baghaie:2001:SIR

[BD01] R. Baghaie and V. Dimitrov. Systolic implementation of real-valued discrete transforms via algebraic integer quantization. Computers and Mathematics with Applications, 41(10–11):1403–1416, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001055.

Barnett:2002:BTF

[BD02] N. S. Barnett and S. S.

Dragomir. Bounds in terms of the fourth derivative for the remainder in the corrected trapezoid formula. Computers and Mathematics with Applications, 44(5–6):595–605, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001748.

Bastinec:2003:AFP

[BD03] J. Bastinec and J. Diblík.

Asymptotic formulae for a particular solution of linear nonhomogeneous discrete equations. Computers and Mathematics with Applications, 45(6–9):1163–1169, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000774.

Barnett:2004:SFI

[BD04a] N. S. Barnett and S. S.

Dragomir. Some further inequalities for univariate moments and some new ones for the covariance. Computers and Mathematics with Applications, 47(1):23–36, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210490003X.

Blokhin:2004:SWS

[BD04b] A. M. Blokhin and S. V. Dorovskii. Shock waves stability in layered structures. Computers and Mathematics with Applications, 47(2–3):427–440, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900351.

Boglaev:2004:IDD

[BD04c] I. Boglaev and V. Duoba. Iterative domain decomposition algorithms for a convection–diffusion problem. Computers and Mathematics with Applications, 47(4–5):501–518, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900417.

Bhoosnurmath:2007:UVS

[BD07] Subhas S. Bhoosnurmath and Renukadevi S. Dyavanal. Uniqueness and value-sharing of meromorphic functions. Computers and Mathematics with Applications, 53(8):1191– 1205, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000764.

Barnett:2002:SIP

[BDA02] N. S. Barnett, S. S.

Dragomir, and R. P. Agarwal. Some inequalities for probability, expectation, and variance of random variables defined over a finite interval. Computers and Mathematics with Applications, 43(10–11):1319–1357, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001037.

Bellomo:2001:SNP

[BDGR01] N. Bellomo, E. De Angelis, L. Graziano, and A. Romano. Solution of nonlinear problems in applied sciences by generalized collocation methods and mathematica. Computers and Mathematics with Applications, 41(10–11):1343–1363, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001018.

Blokhin:2004:SSW

[BDO04] A. M. Blokhin, S. V.

Dorovskii, and E. V. Ovechkin. Stability of shock waves in layered structures (part 2). Computers and Mathematics with Applications, 47(8–9):1379–1387, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901307.

Byatt:2004:SPD

[BDT04] D. Byatt, M. L. Dalrymple, and R. M. Turner. Searching for primes in the digits of π. Computers and Mathematics with Applications, 48(3–4):497–504, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840719.

Bedivan:2002:EEL

[Bed02]  D. M. Bedivan.  Er-

ror estimates for least squares finite element methods. Computers and Mathematics with Applications, 43(8–9):1003–1020, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800098.

Belenky:2001:APA

[Bel01] A. S. Belenky. An approach to planning an advertising campaign of goods and services. Computers and Mathematics with Applications, 42(6–7):993–1008, September/October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002152.

Belenky:2005:CMF

[Bel05a] A. S. Belenky. Calculating the minimal fraction of the popular vote to win the U.S. Presidency in the Electoral College. Computers and Mathematics with Applications, 50(5–6):783–802, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003172.

Belenky:2005:CSU

[Bel05b] A. S. Belenky. Competitive strategies of U.S. Presidential candidates in election campaigns. Computers and Mathematics with Applications, 49(1):53–71, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000064.

Belenky:2006:TRS

[Bel06] A. S. Belenky. Two rules of a sealed ceiling bid and their analysis by mathematical programming techniques. Computers and Mathematics with Applications, 52(12):1711–1732, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003531.

Belendez:2009:HPM

[Bel09] Augusto Beléndez. Homotopy perturbation method for a conservative x^{1/3} force nonlinear oscillator. Computers and Mathematics with Applications, 58(11–12):2267–2273, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001527.

Benschop:2000:PRR

[Ben00] N. F. Benschop. Powersums representing residues mod p^{k}, from Fermat to Waring. Computers and Mathematics with Applications, 39(7–8):253–261, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000808.

See note [Cha01b] and reply [Ben01].

Benschop:2001:RDR

[Ben01] N. Benschop. Reply to Dr. Robin Chapman’s critique. Computers and Mathematics with Applications, 41(12):1475–1476, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001110.

See [Ben00, Cha01b].

Benjumea:2006:MOL

[BENT06] J. C. Benjumea, F. J.

Echarte, J. Núñez, and A. F. Tenorio. A method to obtain the lie group associated with a nilpotent lie algebra. Computers and Mathematics with Applications, 51(9–10):1493– 1506, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000964.

Berlekamp:2000:UA

[Ber00] E. Berlekamp. Unimodular arrays. Computers and Mathematics with Applications, 39(11):77– 88, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001115.

Bertolazzi:2005:CFM

[Ber05] E. Bertolazzi. A combination formula of Michaelis– Menten–Monod type. Computers and Mathematics with Applications, 50(1–2): 201–215, July 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105002439.

Bai:2003:EMP

[BF03] Chuan-Zhi Bai and Jin-Xuan Fang. Existence of multiple positive solutions for functional differential equations. Computers and Mathematics with Applications, 45(12):1797– 1806, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900010.

Bertolazzi:2008:LCF

[BFFL08] P. Bertolazzi, G. Felici, P. Festa, and G. Lancia. Logic classification and feature selection for biomedical data. Computers and Mathematics with Applications, 55(5):889–899, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005615.

Feng:2009:GTC

[bFFlT09] Wei bing Feng, Cong Feng, and Han li Tian. Growth of thrombus core in supersaturated blood. Computers and Mathematics with Applications, 58(11–12):2468–2472, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001825.

Bruni:2008:MII

[BFKL08] C. Bruni, C. Ferrone, G. Koch, and M. Lucchetti. Motion identification from image sequences: Information theory and pixel selection. Computers and Mathematics with Applications, 56(2):375– 386, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000230.

Bruni:2006:GOP

[BFL06] C. Bruni, C. Ferrone, and M. Lucchetti. Global optimization for possibly timedependent cost functions by a population set-based algorithm with births control. Computers and Mathematics with Applications, 52(6–7):1145–1164, September/October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002859.

Bini:2004:FMR

[BFM04] D. A. Bini, P. Favati, and O. Menchi. A family of modified regularizing circulant preconditioners for two-levels Toeplitz systems. Computers and Mathematics with Applications, 48(5–6):755–768, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003116.

Bosseur:2002:IBC

[BFO02] F. Bosseur, F. Flori, and P. Orenga. Identification of boundary conditions in a nonlinear shallow water flow. Computers and Mathematics with Applications, 43(12):1559– 1573, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001190.

Brassier:2001:RSB

[BG01] S. Brassier and G. Gallice. A Roe scheme for the bi-temperature model of magnetohydrodynamics. Computers and Mathematics with Applications, 41(1–2):257–267, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101850197.

Bai:2004:ETP

[BG04a] Zhanbing Bai and Weigao Ge. Existence of three positive solutions for some second-order boundary value problems. Computers and Mathematics with Applications, 48(5–6):699–707, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003050.

Bochev:2004:LSF

[BG04b] P. Bochev and M. D. Gunzburger. Least-squares finite-element methods for optimization and control problems for the Stokes equations. Computers and Mathematics with Applications, 48(7–8):1035–1057, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210400344X.

Biazar:2007:HVI

[BG07a] J. Biazar and H. Ghazvini.

He’s variational iteration method for fourth-order parabolic equations. Computers and Mathematics with Applications, 54(7–8): 1047–1054, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107002738.

Bohner:2007:DIC

[BG07b] Martin Bohner and Gusein Sh. Guseinov. Double integral calculus of variations on time scales. Computers and Mathematics with Applications, 54(1):45– 57, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001757.

Biazar:2008:HPM

[BG08] J. Biazar and H. Ghazvini. Homotopy perturbation method for solving hyperbolic partial differential equations. Computers and Mathematics with Applications, 56(2): 453–458, July 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108000291.

Biazar:2011:NIH

[BG11] Jafar Biazar and Behzad Ghanbari. Notes on “An improvement to homotopy perturbation method for solving system of linear equations”. Computers and Mathematics with Applications, 61(6):1704, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122111000125.

See [Yus09].

Baselga:2008:GRE

[BGA08] Sergio Baselga and Luis García-Asenjo. Global robust estimation and its application to GPS positioning. Computers and Mathematics with Applications, 56(3):709–714, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000692.

Belendez:2009:NOD

[BGAM09] A. Beléndez, E. Gimeno, M. L. Alvarez, and D. I. Méndez. Nonlinear oscillator with discontinuity by generalized harmonic balance method. Computers and Mathematics with Applications, 58(11–12):2117–2123, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001357.

Balbas:2004:NNG

[BGG04] A. Balbás, E. Galperin, and P. Jiménez Guerra. Nonsmooth nonconvex global optimization in a Banach space with a basis. Computers and Mathematics with Applications, 48(1–2):131– 144, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840045.

Bertolazzi:2008:SHI

[BGLT08] Paola Bertolazzi, Alessandra Godi, Martine Labbé, and Leonardo Tininini. Solving haplotyping inference parsimony problem using a new basic polynomial formulation. Computers and Mathematics with Applications, 55(5):900–911, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005627.

Bru:2002:ISN

[BGM^{+}02] R. Bru, D. Genstar, J. Marin, G. Verdú, J. Mas, and T. Manteuffel. Iterative schemes for the neutron diffusion equation. Computers and Mathematics with Applications, 44(10– 11):1307–1323, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002584.

Balbas:2002:BSA

[BGMB02] A. Balbás, P. Jiménez Guerra, and M. J. Muñoz-Bouzo. The balance space approach in optimization with Riesz spaces valued objectives. an application to financial markets. Computers and Mathematics with Applications, 44(7):887–897, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002018.

Berenhaut:2007:BFO

[BGNA07] Kenneth S. Berenhaut, Benjamin G. Gibson, Jonathan H. Newman, and Jacob F. Anderson. Bounds for fourthorder [0, 1] difference equations. Computers and Mathematics with Applications, 54(9–10):1250–1259, November 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003276.

Benchohra:2004:NIF

[BGO04] M. Benchohra, J. R. Graef, and A. Ouahabi. Nonresonance impulsive functional differential inclusions with variable times. Computers and Mathematics with Applications, 47 (10–11):1725–1737, May/ June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839865.

Benchohra:2005:ONS

[BGO05] M. Benchohra, J. R. Graef, and A. Ouahab. Oscillatory and nonoscillatory solutions of multivalued differential inclusions. Computers and Mathematics with Applications, 49(9–10):1347– 1354, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001549.

Bini:2004:IPD

[BGP04] D. A. Bini, L. Gemignani, and V. Y. Pan. Inverse power and Durand–Kerner iterations for univariate polynomial root-finding. Computers and Mathematics with Applications, 47 (2–3):447–459, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900375.

Benzoni-Gavage:2009:WNS

[BGR09] S. Benzoni-Gavage and M. D. Rosini. Weakly nonlinear surface waves and subsonic phase boundaries. Computers and Mathematics with Applications, 57(9):1463–1484, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000571.

Bolton:2004:AUB

[BGS04] H. P. J. Bolton, A. A.

Groenwold, and J. A. Snyman. The application of a unified Bayesian stopping criterion in competing parallel algorithms for global optimization. Computers and Mathematics with Applications, 48(3–4):549–560, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840768.

Butler:2000:GFG

[BH00] L. M. Butler and A. W.

Hales. Generalized flags in p-groups. Computers and Mathematics with Applications, 39(11):67– 76, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001103.

Boulbrachene:2001:FEA

[BH01] M. Boulbrachene and M. Haiour. The finite element approximation of Hamilton–Jacobi– Bellman equations. Computers and Mathematics with Applications, 41(7–8): 993–1007, April 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100003345.

Bellos:2003:NSM

[BH03] C. Bellos and V. Hrissanthou. Numerical simulation of morphological changes in rivers and reservoirs. Computers and Mathematics with Applications, 45(1–3):453–467, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800305.

Bounaim:2007:DBL

[BHCØ07] A. Bounaïm, S. Holm, Wen Chen, and Å. Ødegård. Detectability of breast lesions with CARI ultrasonography using a bioacoustic computational approach. Computers and Mathematics with Applications, 54(1):

96–106, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001861.

Behiry:2007:NAD

[BHEKE07] S. H. Behiry, H. Hashish, I. L. El-Kalla, and A. Elsaid. A new algorithm for the decomposition solution of nonlinear differential equations. Computers and Mathematics with Applications, 54(4):459–466, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002271.

Bahri:2008:UPQ

[BHHA08] Mawardi Bahri, Eckhard S. M. Hitzer, Akihisa Hayashi, and Ryuichi Ashino. An uncertainty principle for quaternion Fourier transform. Computers and Mathematics with Applications, 56(9):2398–2410, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003799.

Benchohra:2001:ERF

[BHN01] M. Benchohra, J. Henderson, and S. K. Ntouyas. An existence result for first-order impulsive functional differential equations in Banach spaces. Computers and Mathematics with Applications, 42(10– 11):1303–1310, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002413.

Benchohra:2004:IFD

[BHNO04a] M. Benchohra, J. Henderson, S. K. Ntouyas, and A. Ouahab. Impulsive functional differential equations with variable times. Computers and Mathematics with Applications, 47(10–11):1659–1665, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839798.

Benchohra:2004:ULS

[BHNO04b] M. Benchohra, J. Henderson, S. K. Ntouyas, and A. Ouahab. Upper and lower solutions method for first-order impulsive differential inclusions with nonlinear boundary conditions. Computers and Mathematics with Applications, 47(6–7):1069–1078, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900879.

Bernsdorf:2008:ALB

[BHS^{+}08] J. Bernsdorf, S. E. Harrison, S. M. Smith, P. V. Lawford, and D. R. Hose. Applying the lattice Boltzmann technique to biofluids: a novel approach to simulate blood coagulation. Computers and Mathematics with Applications, 55(7):1408– 1414, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006323.

Bigerelle:2001:NMC

[BI01] M. Bigerelle and A. Iost.

A new method to calculate the fractal dimension of surfaces: application to human cell proliferation. Computers and Mathematics with Applications, 42(1–2):241– 253, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001481.

Behrens:2002:GFA

[BI02] J. Behrens and A. Iske.

Grid-free adaptive semi-Lagrangian advection using radial basis functions. Computers and Mathematics with Applications, 43 (3–5):319–327, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002899.

Bibi:2006:ETF

[Bib06] A. Bibi. Evolutionary transfer functions of bilinear processes with time-varying coefficients. Computers and Mathematics with Applications, 52(3–4):331–338, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002008.

Bruni:2006:FIS

[BIKL06] C. Bruni, D. Iacoviello, G. Koch, and M. Lucchetti. Filtering image sequences from a moving object and the edge detection problem. Computers and Mathematics with Applications, 51(3–4):559–578, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105005225.

Bilge:2005:TCS

[Bil05] A. H. Bilge. Towards the classification of scalar nonpolynomial evolution equations: Quasilinearity. Computers and Mathematics with Applications, 49(11– 12):1837–1848, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105002130.

Butzer:2000:MFS

[BJ00] P. L. Butzer and S. Jansche.

Mellin–Fourier series and the classical Mellin transform. Computers and Mathematics with Applications, 40(1):49–62, July 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100001395.

Bartoszewski:2002:SAT

[BJ02] Z. Bartoszewski and Z. Jackiewicz. Stability analysis of two-step Runge–Kutta methods for delay differential equations. Computers and Mathematics with Applications, 44(1–2):83– 93, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001311.

Basu:2003:HOM

[BJBL03] P. K. Basu, A. B. Jorge, S. Badri, and J. Lin. Higher-order modeling of continua by finite-element, boundary-element, meshless, and wavelet methods. Computers and Mathematics with Applications, 46(1): 15–33, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900782.

Bo:2008:CVM

[BJK^{+}08] W. Bo, H. Jin, D. Kim, X. Liu, H. Lee, N. Pestieau, Y. Yu, J. Glimm, and J. W. Grove. Comparison and validation of multi phase closure models. Computers and Mathematics with Applications, 56(5):1291–1302, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800151X.

Becker:2009:CLB

[BJK^{+}09] Jürgen Becker, Michael Junk, Dirk Kehrwald, Guido Thömmes, and Zhaoxia Yang. A combined lattice BGK/level set method for immiscible two-phase flows. Computers and Mathematics with Applications, 58(5): 950–964, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210900100X.

Jian:2009:FQF

[bJQxC09] Jin bao Jian, Ran Quan, and Wei xin Cheng. A feasible QP-free algorithm combining the interiorpoint method with active set for constrained optimization. Computers and Mathematics with Applications, 58(8):1520–1533, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004738.

Baasansuren:2000:RAP

[BK00] J. Baasansuren and A. A.

Khan. Regularized auxiliary problem principle for variational inequalities. Computers and Mathematics with Applications, 40 (8–9):995–1002, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850105.

Bartoszewski:2004:DDE

[BK04] Z. Bartoszewski and M. Kwapisz.

Delay dependent estimates for waveform relaxation methods for neutral differentialfunctional systems. Computers and Mathematics with Applications, 48(12): 1877–1892, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004468.

Burcsi:2007:ICT

[BK07] Péter Burcsi and Attila Kovács. On the importance of cache tuning in a cache-aware algorithm: a case study. Computers and Mathematics with Applications, 53(6):880–885, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000405.

Brandts:2008:ERC

[BKK08] Jan Brandts, Sergey Korotov, and Michal Krízek. On the equivalence of regularity criteria for triangular and tetrahedral finite element partitions. Computers and Mathematics with Applications, 55(10):2227– 2233, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007407.

Baeumer:2008:NSF

[BKM08] Boris Baeumer, Mihály Kovács, and Mark M. Meerschaert. Numerical solutions for fractional reactiondiffusion equations. Computers and Mathematics with Applications, 55(10): 2212–2226, May 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107007420.

Bang:2009:RFN

[BKN09] Keumseong Bang, H. S.

Kim, and J. Neggers. Representation functions and the Neggers–Stanley condition for weight-shaped-posets. Computers and Mathematics with Applications, 57(1):17–20, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006123.

Burger:2005:RSC

[BKR05] R. Bürger, K. H. Karlsen, and N. H. Risebro. A relaxation scheme for continuous sedimentation in ideal clarifier–thickener units. Computers and Mathematics with Applications, 50 (7):993–1009, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105003470.

Bacchelli:2001:SPA

[BL01] S. Bacchelli and D. Lazzaro. Some practical applications of block recursive matrices. Computers and Mathematics with Applications, 41(9):1183– 1198, May 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101000906.

Balaj:2008:SFT

[BL08] M. Balaj and Lai-Jiu Lin. Selecting families and their applications. Computers and Mathematics with Applications, 55(6):1257–1261, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005937.

Blacker:2004:RPE

[Bla04] D. J. Blacker. A robust a posteriori error estimate for the Fortin–Soulie finiteelement method. Computers and Mathematics with Applications, 48(12):1863– 1876, December 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104004456.

Blasco:2007:APS

[Bla07] Jordi Blasco. An anisotropic pressure-stabilized finite element method for incompressible flow problems.

Computers and Mathematics with Applications, 53 (6):895–909, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107000429.

Bai:2007:TPS

[BLD07] Zhanbing Bai, Xiangqian Liang, and Zengji Du. Triple positive solutions for some second-order boundary value problem on a measure chain. Computers and Mathematics with Applications, 53(12):1832– 1839, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001538.

Bloom:2000:SGS

[Blo00] G. S. Bloom. Solomon Golomb — some personal reflections plus random biographical and bibliographical notes. Computers and Mathematics with Applications, 39(11):xv–xviii, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100902633.

Blows:2008:SAL

[Blo08] Terence R. Blows. Small amplitude limit cycles of symmetric cubic systems. Computers and Mathematics with Applications, 56(9): 2323–2328, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108003702.

Burns:2005:PGF

[BLS05] J. A. Burns, T. Lin, and L. G. Stanley. A Petrov Galerkin finiteelement method for interface problems arising in sensitivity computations. Computers and Mathematics with Applications, 49(11–12):1889– 1903, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002063.

Beran:2000:TMS

[BM00] J. Beran and G. Mazzola. Timing microstructure in Schumann’s “Träumerei” as an expression of harmony, rhythm, and motivic structure in music performance. Computers and Mathematics with Applications, 39(5–6):99–130, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000493.

Balasubramaniam:2001:EUF

[BM01] P. Balasubramaniam and S. Muralisankar. Existence and uniqueness of a fuzzy solution for the nonlinear fuzzy neutral functional differential equation. Computers and Mathematics with Applications, 42 (6–7):961–967, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002127.

See erratum [BM02].

Balasubramaniam:2002:EEU

[BM02] P. Balasubramaniam and S. Muralisankar. Erratum to “Existence and uniqueness of a fuzzy solution for the nonlinear fuzzy neutral functional differential equation”. Computers and Mathematics with Applications, 43(12):1631– 1632, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001244.

See [BM01].

Balasubramaniam:2004:EUF

[BM04a] P. Balasubramaniam and S. Muralisankar. Existence and uniqueness of fuzzy solution for semilinear fuzzy integrodifferential equations with nonlocal conditions. Computers and Mathematics with Applications, 47(6–7):1115–1122, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900910.

Banas:2004:MSQ

[BM04b] J. Banaś and A. Martinon. Monotonic solutions of a quadratic integral equation of Volterra type. Computers and Mathematics with Applications, 47(2–3):271–279, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900247.

Banas:2005:SPN

[BM05a] J. Banaś and J. Caballero Mena. Some properties of nonlinear Volterra– Stieltjes integral operators. Computers and Mathematics with Applications, 49(9–10):1565–1573, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001690.

Bertolazzi:2005:UTB

[BM05b] E. Bertolazzi and G. Manzini.

A unified treatment of boundary conditions in least-square based finitevolume methods. Computers and Mathematics with Applications, 49(11– 12):1755–1765, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105001938.

Balbas:2006:NOP

[BM06] A. Balbas and S. Mayoral.

Nonconvex optimization for pricing and hedging in imperfect markets. Computers and Mathematics with Applications, 52(1–2):121– 136, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001866.

Bianca:2009:MMC

[BM09] Carlo Bianca and Santo Motta. The MWF method: a convergence theorem for homogeneous one-dimensional case. Computers and Mathematics with Applications, 58(3):579–588, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002284.

Berndt:2008:ENS

[BMH08] Markus Berndt, J. David Moulton, and Glen Hansen. Efficient nonlinear solvers for Laplace–Beltrami smoothing of three-dimensional unstructured grids. Computers and Mathematics with Applications, 55(12):2791– 2806, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700778X.

Babolian:2009:NSN

[BMHV09] E. Babolian, Z. Masouri, and S. Hatamzadeh-Varmazyar. Numerical solution of nonlinear Volterra–Fredholm integro-differential equations via direct method using triangular functions. Computers and Mathematics with Applications, 58(2): 239–247, July 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109002491.

Brezzi:2006:SEA

[BMM^{+}06] F. Brezzi, L. D. Marini, S. Micheletti, P. Pietra, and R. Sacco. Stability and error analysis of mixed finite-volume methods for advection dominated problems. Computers and Mathematics with Applications, 51(5):681–696, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000113.

Benchohra:2000:EMS

[BN00] M. Benchohra and S. K.

Ntouyas. Existence of mild solutions on noncompact intervals to secondorder initial value problems for a class of differential inclusions with nonlocal conditions. Computers and Mathematics with Applications, 39(12): 11–18, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001255.

Bawa:2005:CMS

[BN05] R. K. Bawa and S. Natesan. A computational method for self-adjoint singular perturbation problems using quintic spline. Computers and Mathematics with Applications, 50 (8–9):1371–1382, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003871.

Batiha:2007:VIM

[BNH07] B. Batiha, M. S. M. Noorani, and I. Hashim. Variational iteration method for solving multispecies Lotka–Volterra equations. Computers and Mathematics with Applications, 54(7– 8):903–909, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107003069.

Bataineh:2008:AAS

[BNH08] A. Sami Bataineh, M. S. M. Noorani, and I. Hashim. Approximate analytical solutions of systems of PDEs by homotopy analysis method. Computers and Mathematics with Applications, 55 (12):2913–2923, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107008139.

Bai:2005:ASW

[BNP05] Zhong-Zhi Bai, Michael K. Ng, and Jian-Yu Pan. Alternating splitting waveform relaxation method and its successive overrelaxation acceleration. Computers and Mathematics with Applications, 49(2–3):157–170, January/February 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105000258.

Bernardini:2005:MBP

[BNR05] A. Bernardini, P. Natalini, and P. E. Ricci. Multidimensional bell polynomials of higher order. Computers and Mathematics with Applications, 50(10– 12):1697–1708, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004190.

Bogdanowicz:2009:NEA

[Bog09] Zbigniew R. Bogdanowicz. A new efficient algorithm for optimal assignment of smart weapons to targets. Computers and Mathematics with Applications, 58(10):1965–1969, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006063.

Bouchut:2007:UST

[BOP07] François Bouchut, Haythem Ounaissa, and Benoît Perthame. Upwinding of the source term at interfaces for Euler equations with high friction. Computers and Mathematics with Applications, 53(3–4):361–375, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001939.

Boulbrachene:2000:FPA

[Bou00]  M. Boulbrachene.  A

fixed-point approximation method for some noncoercive variational problems. Computers and Mathematics with Applications, 39(7– 8):17–27, April 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100000614.

Boulbrachene:2003:EES

[Bou03]  M. Boulbrachene.  L^{∞}-

error estimate for a system of elliptic quasi-variational inequalities with noncoercive operators. Computers and Mathematics with Applications, 45(6–9):983–989, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000725.

Boyd:2001:RCS

[Boy01]  J. P. Boyd.  Rational

Chebyshev spectral methods for unbounded solutions on an infinite interval using polynomialgrowth special basis functions. Computers and Mathematics with Applications, 41(10–11):1293–1315, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101000980.

Boyd:2007:CZF

[Boy07] John P. Boyd. Computing the zeros of a Fourier series or a Chebyshev series or general orthogonal polynomial series with parity symmetries. Computers and Mathematics with Applications, 54(3):336–349, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002167.

Bay:2003:SAN

[BP03] N. S. Bay and Vu Ngoc Phat. Stability analysis of nonlinear retarded difference equations in Banach spaces. Computers and Mathematics with Applications, 45(6–9):951–960, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000683.

Bajak:2009:CAS

[BP09] Szabolcs Baják and Zsolt Páles. Computer aided solution of the invariance equation for two-variable Gini means. Computers and Mathematics with Applications, 58(2):334– 340, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002879.

Balachandran:2003:ESA

[BPA03] K. Balachandran, D. G. Park, and S. Marshal Anthoni. Existence of solutions of abstract nonlinear second-order neutral functional integrodifferential equations. Computers and Mathematics with Applications, 46(8– 9):1313–1324, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902215.

See comments [HM05b].

Balachandran:2005:CSO

[BPM05] K. Balachandran, D. G. Park, and P. Manimegalai. Controllability of secondorder integrodifferential evolution systems in Banach spaces. Computers and Mathematics with Applications, 49(11–12):1623– 1642, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001847.

Bianca:2009:MMK

[BPM09] Carlo Bianca, Francesco Pappalardo, and Santo Motta. The MWF method for kinetic equations system. Computers and Mathematics with Applications, 57(5):831–840, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004756.

Banas:2004:MSU

[BPP04] J. Banaś and M. Paslawska-Poludniak. Monotonic solutions of Urysohn integral equation on unbounded interval. Computers and Mathematics with Applications, 47(12):1947– 1954, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840306.

Borisevich:2000:GMS

[BPSW00] V. D. Borisevich, V. G.

Potemkin, S. P. Strunkov, and H. G. Wood. Global methods for solving systems of nonlinear algebraic equations. Computers and Mathematics with Applications, 40(8– 9):1015–1025, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850129.

Baran:2003:EMS

[BPvZ03] S. Baran, G. Pap, and M. C. A. van Zuijlen. Estimation of the mean of stationary and nonstationary Ornstein–Uhlenbeck processes and sheets. Computers and Mathematics with Applications, 45(4–5):563– 579, February/March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103000178.

Borisevich:2000:NAF

[BPW00] V. D. Borisevich, V. G. Potemkin, and H. G. Wood. A new approach for finding all zeros for systems of nonlinear functions. Computers and Mathematics with Applications, 40(8–9):965–970, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850075.

Bozzini:2003:NDF

[BR03] M. Bozzini and M. Rossini.

Numerical differentiation of 2D functions from noisy data. Computers and Mathematics with Applications, 45(1–3):309–327, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800214.

Bogdanowicz:2005:POS

[BR05a] Z. R. Bogdanowicz and R. Ramamurthy. Properties of optimal survivable paths in a graph. Computers and Mathematics with Applications, 50(3–4):425– 432, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002865.

Bognar:2005:NAI

[BR05b] G. Bognár and M. Rontó.

Numerical–analytic investigation of the radially symmetric solutions for some nonlinear PDEs. Computers and Mathematics with Applications, 50(7):983–991, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003299.

Bridges:2006:PFC

[BR06] C. Bridges and K. R. Rajagopal. Pulsatile flow of a chemically-reacting nonlinear fluid. Computers and Mathematics with Applications, 52(6– 7):1131–1144, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002847.

Braumann:2008:GEP

[Bra08] Carlos A. Braumann. Growth and extinction of populations in randomly varying environments. Computers and Mathematics with Applications, 56(3):631–644, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000540.

Bhutani:2000:ASR

[BRC00] O. P. Bhutani and L. Roy-Chowdhury. Applications of some recent techniques for the exact solutions of the small disturbance potential flow equation of nonequilibrium transonic gas dynamics. Computers and Mathematics with Applications, 40(12):1349–1361, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002443.

Bribiesca:2000:MCS

[Bri00] E. Bribiesca. A measure of compactness for 3D shapes. Computers and Mathematics with Applications, 40 (10–11):1275–1284, November/December 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100002388.

Brandao:2001:NCT

[BRMS01] A. J. V. Brandão, M. A.

Rojas-Medar, and G. N. Silva. Nonsmooth continuoustime optimization problems: necessary conditions. Computers and Mathematics with Applications, 41(12):1477–1486, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001122.

Bruni:2008:SPS

[Bru08]  Renato Bruni.  Solv-

ing peptide sequencing as satisfiability. Computers and Mathematics with Applications, 55(5):912–923, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005639.

Barakat:2000:FTS

[BS00a] R. Barakat and B. H. Sandler. Filon trapezoidal schemes for Hankel transforms of orders zero and one. Computers and Mathematics with Applications, 40 (8–9):1037–1041, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100850142.

Bhardwaj:2000:CMR

[BS00b] D. Bhardwaj and R. Shankar. A computational method for regularized long wave equation. Computers and Mathematics with Applications, 40(12):1397–1404, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002480.

Banas:2001:SVS

[BS01a] J. Banaś and K. Sadarangani. Solvability of Volterra– Stieltjes operator-integral equations and their applications. Computers and Mathematics with Applications, 41(12):1535–1544, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001183.

Brini:2001:ECD

[BS01b] F. Brini and S. Siboni. Estimates of correlation decay in auto/endomorphisms of the n-torus. Computers and Mathematics with Applications, 42(6– 7):941–951, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002103.

Blanchet-sadri:2003:PRP

[Bs03a] F. Blanchet-sadri. A periodicity result of partial words with one hole. Computers and Mathematics with Applications, 46(5–6):813–820, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901441.

Bognar:2003:SNE

[BS03b] G. Bognár and T. Szabó.

Solving nonlinear eigenvalue problems by using pversion of FEM. Computers and Mathematics with Applications, 46(1):57– 68, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900800.

Blanchet-Sadri:2004:PPW

[BS04] F. Blanchet-Sadri. Periodicity on partial words. Computers and Mathematics with Applications, 47(1):71– 82, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900065.

Barrio:2006:MMV

[BS06] Roberto Barrio and Sergio Serrano. Modifications of the method of variation of parameters. Computers and Mathematics with Applications, 51(3–4):451–466, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105005122.

Bisi:2008:KAG

[BS08] M. Bisi and G. Spiga. Kinetic approach to granular matter in a medium. Computers and Mathematics with Applications, 55 (2):162–174, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0898122107003446.

Balachandran:2000:CNF

[BSD00] K. Balachandran, R. Sakthivel, and J. P. Dauer. Controllability of neutral functional integrodifferential systems in Banach spaces. Computers and Mathematics with Applications, 39(1–2):117–126, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003181.

See [HPT05].

Balachandran:2004:ESN

[BSK04] K. Balachandran, G. Shija, and J. H. Kim. Existence of solutions of nonlinear abstract neutral integrodifferential equations. Computers and Mathematics with Applications, 48(10– 11):1403–1414, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003967.

Blanchet-Sadri:2001:MSD

[BSM01] F. Blanchet-Sadri and C. Morgan. Multiset and set decipherable codes. Computers and Mathematics with Applications, 41(10–11):1257– 1262, May/June 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122101000967.

Balakrishnan:2002:OSR

[BSR02] K. Balakrishnan, R. Sureshkumar, and P. A. Ramachandran. An operator splittingradial basis function method for the solution of transient nonlinear Poisson problems. Computers and Mathematics with Applications, 43 (3–5):289–304, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002875.

Bai:1998:SAM

[BSS98] Zhong-Zhi Bai, Jia-Chang Sun, and D. B. Szyld. The spectral analysis of the matrix multisplitting method for the one-dimensional model problem. Computers and Mathematics with Applications, 36(2):31– 38, July 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812219800114X.

See corrigendum [Ano00-94].

Basebi:2003:SMM

[BT03] T. Basebi and R. M. Thomas. A study of moving mesh methods applied to a thin flame propagating in a detonator delay element. Computers and Mathematics with Applications, 45(1–3):131–163, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800123.

Bu:2000:GNF

[Bu00] T. Bu. Golombs Norske forbindelser. Computers and Mathematics with Applications, 39(11):135– 138, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001206.

Buhmann:2006:HPA

[Buh06] M. D. Buhmann. Halfplane approximation with radial functions. Computers and Mathematics with Applications, 51(8):1171– 1184, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000824.

Bahuguna:2009:CSN

[BUP09a] D. Bahuguna, A. Ujlayan, and D. N. Pandey. A comparative study of numerical methods for solving an integro-differential equation. Computers and Mathematics with Applications, 57(9):1485–1493, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000558.

Bahuguna:2009:MKP

[BUP09b] D. Bahuguna, A. Ujlayan, and Dwijendra N. Pandey. Method of Kronecker product to advanced type Riccati differential systems with strongly coupled quadratic terms. Computers and Mathematics with Applications, 58(8):1615–1622, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004507.

Balibrea:2001:SSU

[BV01a] F. Balibrea and J. C. Valverde. Structural stability under conditions of nonhyperbolicity. Computers and Mathematics with Applications, 41(5–6):757– 768, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003187.

Bribiesca:2001:FLA

[BV01b] E. Bribiesca and C. Velarde. A formal language approach for a 3D curve representation. Computers and Mathematics with Applications, 42(12):1571–1584, December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002632.

Balasubramaniam:2005:ESN

[BV05] P. Balasubramaniam and D. Vinayagam. Existence of solutions of nonlinear stochastic integrodifferential inclusions in a Hilbert space. Computers and Mathematics with Applications, 50(5–6):809–821, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003196.

Bombardelli:2009:PCF

[BV09] Mea Bombardelli and Sanja Varošanec. Properties of h-convex functions related to the Hermite– Hadamard–Fejér inequalities. Computers and Mathematics with Applications, 58(9):1869–1877, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005951.

Boey:2005:ETP

[BW05] K. L. Boey and Patricia J. Y. Wong. Existence of triple positive solutions of two-point right focal boundary value problems on time scales. Computers and Mathematics with Applications, 50(10– 12):1603–1620, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105004116.

Baatar:2006:AEM

[BW06a] D. Baatar and M. M. Wiecek. Advancing equitability in multiobjective programming. Computers and Mathematics with Applications, 52(1–2):225– 234, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210600191X.

Bao:2006:CSF

[BW06b] Yu-E. Bao and Cong-Xin Wu. Convexity and semicontinuity of fuzzy mappings. Computers and Mathematics with Applications, 51(12):1809–1816, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210600126X.

Boey:2006:PST

[BW06c] K. L. Boey and Patricia J. Y. Wong. Positive solutions of two-point right focal boundary value problems on time scales. Computers and Mathematics with Applications, 52(3–4):555– 576, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002185.

Bernsdorf:2009:NNB

[BW09] J. Bernsdorf and D. Wang.

Non-Newtonian blood flow simulation in cerebral aneurysms. Computers and Mathematics with Applications, 58(5): 1024–1029, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109001072.

Belytschko:2002:SAP

[BX02] T. Belytschko and Shaoping Xiao. Stability analysis of particle methods with corrected derivatives. Computers and Mathematics with Applications, 43 (3–5):329–350, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002905.

Bai:2007:PSF

[BX07] Dingyong Bai and Yuantong Xu. Periodic solutions of first order functional differential equations with periodic deviations. Computers and Mathematics with Applications, 53(9):1361– 1366, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000971.

Bnouhachem:2009:MEM

[BXFZ09] Abdellah Bnouhachem, M. H. Xu, Xiao-Ling Fu, and Sheng Zhaohan. Modified extragradient methods for solving variational inequalities. Computers and Mathematics with Applications, 57(2):230–239, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006159.

Barrio:2003:SPA

[BY03] R. Barrio and P. Yalamov. Stability of parallel algorithms for polynomial evaluation. Computers and Mathematics with Applications, 46(5–6):769–781, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901404.

Buck:2000:DLR

[BZ00] M. Buck and N. Zierler.

Decimations of linear recurring sequences. Computers and Mathematics with Applications, 39(11): 95–102, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001139.

Bai:2007:GMV

[BZN07] Min-Ru Bai, Shu-Zi Zhou, and Gu-Yan Ni. On the generalized monotonicity of variational inequalities. Computers and Mathematics with Applications, 53 (6):910–917, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107000430.

Coskun:2009:DCB

[CA09] Safa Bozkurt Coskun and Mehmet Tarik Atay. Determination of critical buckling load for elastic columns of constant and variable cross-sections using variational iteration method. Computers and Mathematics with Applications, 58(11–12):2260–2266, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001540.

Cabada:2001:ESD

[Cab01]  A. Cabada.  Extremal

solutions for the difference ϕ-Laplacian problem with nonlinear functional boundary conditions. Computers and Mathematics with Applications, 42(3–5):593–601, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001791.

Caballero-Aguila:2009:RED

[CÁHCJL^{+}09] R. Caballero-Águila, A. Hermoso-Carazo, J. D. Jiménez-López, J. Linares-Pérez, and S. Nakamori. Recursive estimation of discretetime signals from nonlinear randomly delayed observations. Computers and Mathematics with Applications, 58(6):1160–1168, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900460X.

Cancelliere:2003:DPF

[Can03] R. Cancelliere. Data processing and feature screening in function approximation: an application to neural networks. Computers and Mathematics with Applications, 46(2–3):455–461, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900381.

Candan:2008:OFO

[Can08]  T. Candan.  Oscilla-

tion of first-order neutral differential equations with distributed deviating arguments. Computers and Mathematics with Applications, 55(3):510–515, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107004877.

Capizzano:2000:STL

[Cap00] S. Serra Capizzano. Some theorems on linear positive operators and functionals and their applications. Computers and Mathematics with Applications, 39(7–8):139–167, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000729.

Capinski:2007:MCR

[Cap07] Marek Capinski. A model of credit risk based on cash flow. Computers and Mathematics with Applications, 54(4):499–506, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002337.

Cattani:2005:HWT

[Cat05]  C. Cattani.  Harmonic

wavelets towards the solution of nonlinear PDE. Computers and Mathematics with Applications, 50 (8–9):1191–1210, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003743.

Cavazzoni:2000:DAF

[Cav00] R. Cavazzoni. Diffusive approximation of Fisher’s equation. Computers and Mathematics with Applications, 39(9–10):101–114, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000900.

Chawla:2001:ETF

[CAZ01] M. M. Chawla and M. A.

Al-Zanaidi. An extended trapezoidal formula for the diffusion equation in two space dimensions. Computers and Mathematics with Applications, 42(1–2):157– 168, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001407.

Chawla:2000:EOS

[CAZAA00] M. M. Chawla, M. A. Al-Zanaidi, and M. G. Al-Aslab. Extended one-step time-integration schemes for convection–diffusion equations. Computers and Mathematics with Applications, 39(3–4):71–84, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812219900334X.

Chakrabarti:2002:NVS

[CB02] A. Chakrabarti and R. K.

Bera. Nonlinear vibration and stability of a shallow unsymmetrical orthotropic sandwich shell of double curvature with orthotropic core. Computers and Mathematics with Applications, 43(12):1617– 1630, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001232.

Cruz-Barroso:2009:QFA

[CBGVPP09] Ruymán Cruz-Barroso, Pablo González-Vera, and F. Perdomo-Pío. Quadrature formulas associated with Rogers– Szegő polynomials. Computers and Mathematics with Applications, 57(2):308–323, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006317.

Chen:2001:FCP

[CBJ01] B. Chen, R. Garcia Bolós, and L. Jódar. Frobenius– Chebyshev polynomial approximations with a priori error bounds for nonlinear initial value differential problems. Computers and Mathematics with Applications, 41(3–4):269–280, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002716.

Costabile:2002:KTE

[CC02] F. Costabile and S. Serra Capizzano. Komleva-type expansions and asymptotics for linear operators. Computers and Mathematics with Applications, 43(6–7):799–820, March/ April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003224.

Cabada:2004:NFP

[CC04a] A. Cabada and J. Á. Cid.

A note on fixed-points theorems for T -monotone operators. Computers and Mathematics with Applications, 47(6–7):853–857, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900697.

Choo:2004:NSD

[CC04b] S. M. Choo and S. K.

Chung. Numerical solutions for the damped Boussinesq equation by FD– FFT-Perturbation method. Computers and Mathematics with Applications, 47 (6–7):1135–1140, March/ April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900934.

Chen:2005:FIR

[CC05] Shi-Jay Chen and Shyi-Ming Chen. Fuzzy information retrieval based on geometric-mean averaging operators. Computers and Mathematics with Applications, 49(7–8):1213–1231, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500132X.

Chen:2008:FRA

[CC08a] Shi-Jay Chen and Shyi-Ming Chen. Fuzzy risk analysis based on measures of similarity between interval-valued fuzzy numbers. Computers and Mathematics with Applications, 55(8):1670–1685, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006062.

Cui:2008:CCI

[CC08b] Lihong Cui and Ruixue Cong. Construction for a class of interpolation multiscaling functions with dilation factor a ≥ 3. Computers and Mathematics with Applications, 56(11):2948– 2956, December 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004963.

Caglar:2009:BSM

[CC09a] Nazan Caglar and Hikmet Caglar. B-spline method for solving linear system of second-order boundary value problems. Computers and Mathematics with Applications, 57(5):757–762, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005816.

Chen:2009:AOS

[CC09b] Zhixiang Chen and Feilong Cao. The approximation operators with sigmoidal functions. Computers and Mathematics with Applications, 58(4):758–765, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003071.

Cavallini:2008:FVW

[CCC08] N. Cavallini, V. Caleffi, and V. Coscia. Finite volume and WENO scheme in onedimensional vascular system modelling. Computers and Mathematics with Applications, 56(9):2382–2397, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003775.

Cerone:2007:OTI

[CCD07] P. Cerone, W. S. Cheung, and S. S. Dragomir. On Ostrowski type inequalities for Stieltjes integrals with absolutely continuous integrands and integrators of bounded variation. Computers and Mathematics with Applications, 54(2):183– 191, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002040.

Cheng:2006:FDP

[CCF06] Chi-Bin Cheng, Ching-Lung Chen, and Chung-Jen Fu. Financial distress prediction by a radial basis function network with logit analysis learning. Computers and Mathematics with Applications, 51(3–4):579–588, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105005237.

Chiandussi:2009:TOO

[CCF09] G. Chiandussi, M. Codegone, and S. Ferrero. Topology optimization with optimality criteria and transmissible loads. Computers and Mathematics with Applications, 57(5):772–788, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800607X.

Cho:2008:DNX

[CCHK08] Sung-Jin Cho, Un-Sook Choi, Yoon-Hee Hwang, and Han-Doo Kim. Design of new XOR-based hash functions for cache memories. Computers and Mathematics with Applications, 55(9):2005–2011, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006682.

Calbo:2008:RAS

[CCJ08] Gema Calbo, Juan Carlos Cortés, and Lucas Jódar. Random analytic solution of coupled differential models with uncertain initial condition and source term. Computers and Mathematics with Applications, 56(3):785–798, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000734.

Choo:2000:CND

[CCK00] S. M. Choo, S. K. Chung, and K. I. Kim. Conservative nonlinear difference scheme for the Cahn–Hilliard equation — II. Computers and Mathematics with Applications, 39(1–2):229–243, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003260.

Cho:2003:ACC

[CCK03] Sung-Jin Cho, Un-Sook Choi, and Han-Doo Kim. Analysis of complemented CA derived from a linear TPMACA. Computers and Mathematics with Applications, 45(4–5):689– 698, February/March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103000282.

Chen-Charpentier:2008:MMB

[CCK08] Benito M. Chen-Charpentier and Hristo V. Kojouharov. Mathematical modeling of bioremediation of trichloroethylene in aquifers. Computers and Mathematics with Applications, 56(3):645–656, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000552.

Cho:2007:ACC

[CCKH07] Sung-Jin Cho, Un-Sook Choi, Han-Doo Kim, and Yoon-Hee Hwang. Analysis of complemented CA derived from linear hybrid group CA. Computers and Mathematics with Applications, 53(1):54–63, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000314.

Cheng:2002:NFG

[CCL02] Chi-Bin Cheng, C.-J. Cheng, and E. S. Lee. Neurofuzzy and genetic algorithm in multiple response optimization. Computers and Mathematics with Applications, 44(12):1503–1514, December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002742.

Cabada:2002:ERA

[CCP02] A. Cabada, J. Á. Cid, and R. L. Pouso. Existence results and approximation methods for functional ordinary differential equations with singular diffusion-type differential operators. Computers and Mathematics with Applications, 44(1–2):111– 124, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001347.

Chung:2009:SEB

[CCP09] Kuo-Liang Chung, Ping-Zen Chen, and Ying-Lin Pan. Speed up of the edge-based inverse halftoning algorithm using a finite state machine model approach. Computers and Mathematics with Applications, 58(3):484–497, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003307.

Chalco-Cano:2004:CFP

[CCRMOG04] Y. Chalco-Cano, M. A. Rojas-Medar, and R. Osuna-Gómez. s-convex fuzzy processes. Computers and Mathematics with Applications, 47(8–9):1411–1418, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901332.

Chalco-Cano:2000:CFM

[CCRMRF00] Y. Chalco-Cano, M. A. Rojas-Medar, and H. Román-Flores. M-convex fuzzy mappings and fuzzy integral mean. Computers and Mathematics with Applications, 40(10–11):1117–1126, November/December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100002261.

Chalmers:2000:MBS

[CCS00] B. Chalmers, C. Cottin, and B. Shekhtman. Minimal Boolean sum and blending-type projections and extensions. Computers and Mathematics with Applications, 40(1):63– 70, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001401.

Chan:2002:CCG

[CCS02] Whei-Ching C. Chan, Kung-Yu Chen, and H. M. Srivastava. Certain classes of generating functions for the Jacobi and related hypergeometric polynomials. Computers and Mathematics with Applications, 44(12):1539–1556, December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002778.

Canto:2008:DMM

[CCS08] Begoña Cantó, Carmen Coll, and Elena Sánchez. Discrete mathematical models in the analysis of splitting iterative methods for linear systems. Computers and Mathematics with Applications, 56(3):727–732, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000709.

Chang:2004:FPT

[CCSL04] Jung-Chan Chang, Hsinjung Chen, Shih-Min Shyu, and Wei-Cheng Lian. Fixedpoint theorems in fuzzy real line. Computers and Mathematics with Applications, 47(6–7):845–851, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900685.

Cui:2000:EDP

[CCW00] Jingan Cui, Lansun Chen, and Wendi Wang. The effect of dispersal on population growth with stagestructure. Computers and Mathematics with Applications, 39(1–2):91–102, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003168.

Crnkovic:2009:ISI

[CCZK09] Bojan Crnković, Nelida Crnjarić-Zic, and Lado Kranjcević. Improvements of semi-implicit schemes for hyperbolic balance laws applied on open channel flow equations. Computers and Mathematics with Applications, 58(2):292– 309, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002843.

Cerone:2003:DBM

[CD03] P. Cerone and S. S.

Dragomir. Differences between means with bounds from a Riemann–Stieltjes integral. Computers and Mathematics with Applications, 46(2–3):445–453, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390037X.

Cao:2005:OEE

[CD05a] W. Cao and L. Demkowicz. Optimal error estimate of a projection based interpolation for the pversion approximation in three dimensions. Computers and Mathematics with Applications, 50(3–4):359– 366, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002804.

Cerone:2005:BGM

[CD05b] P. Cerone and S. S. Dragomir. Bounds for the Gini mean difference via the sonin identity. Computers and Mathematics with Applications, 50(3–4):599– 609, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002750.

Cuevas:2005:ATR

[CD05c] C. Cuevas and L. Del Campo. An asymptotic theory for retarded functional difference equations. Computers and Mathematics with Applications, 49(5–6):841–855, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000957.

Cerone:2006:BGM

[CD06] P. Cerone and S. S. Dragomir. Bounds for the Gini mean difference of continuous distributions defined on finite intervals (II). Computers and Mathematics with Applications, 52(10– 11):1555–1562, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003397.

Cerone:2009:BCF

[CD09] P. Cerone and S. S.

Dragomir. Bounding the cebysev functional for the Riemann–Stieltjes integral via a Beesack inequality and applications. Computers and Mathematics with Applications, 58(6):1247–1252, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004647.

Company:2004:EAN

[CDJ04] R. Company, E. Defez, and L. Jódar. Exact and analytic numerical solution of coupled parabolic mixed problems in a semi-infinite medium. Computers and Mathematics with Applications, 47 (2–3):381–390, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900326.

Camacho:2003:DNS

[CDJR03] J. Camacho, E. Defez, L. Jódar, and J. V. Romero. Discrete numerical solution of coupled mixed hyperbolic problems. Computers and Mathematics with Applications, 46 (8–9):1183–1193, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902100.

Cecchi:2001:PDS

[CDM01] M. Cecchi, Z. Doslá, and M. Marini. Positive decreasing solutions of quasi-linear difference equations. Computers and Mathematics with Applications, 42(10– 11):1401–1410, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002498.

Cecchi:2003:USQ

[CDM03] M. Cecchi, Z. Doslá, and M. Marini. Unbounded solutions of quasi-linear difference equations. Computers and Mathematics with Applications, 45(6–9):1113– 1123, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103000695.

Caliari:2008:HC

[CDV08] Marco Caliari, Stefano De Marchi, and Marco Vianello. Hyperinterpolation in the cube. Computers and Mathematics with Applications, 55(11):2490– 2497, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007493.

Chikrii:2002:GPF

[CE02] A. A. Chikrii and S. D.

Eidelman. Game problems for fractional quasilinear systems. Computers and Mathematics with Applications, 44(7):835–851, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001979.

Cox:2004:D

[CEG04] Chris Cox, Vince Ervin, and Max Gunzburger. Dedication. Computers and Mathematics with Applications, 48(7–8):xiii, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003414.

Chan:2004:PEW

[CF04] D. M. Chan and J. E.

Franke. Probabilities of extinction, weak extinction permanence, and mutual exclusion in discrete, competitive, Lotka–Volterra systems. Computers and Mathematics with Applications, 47(2–3):365–379, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900314.

Csomos:2005:WSS

[CFH05] P. Csomós, I. Faragó, and Á. Havasi. Weighted sequential splittings and their analysis. Computers and Mathematics with Applications, 50(7):1017–1031, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003317.

Cabrelli:2002:FBC

[CFM02] C. A. Cabrelli, M. C. Falsetti, and U. M. Molter. Fractal block-coding: a functional approach for image and signal processing. Computers and Mathematics with Applications, 44 (8–9):1183–1200, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002250.

Calvo:2003:MSR

[CFR03] M. Calvo, J. M. Franco, and L. Rández. Minimum storage Runge–Kutta schemes for computational acoustics. Computers and Mathematics with Applications, 45(1–3):535–545, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800354.

Cheng:2008:SRD

[CFY08] Wei Cheng, Lianggui Feng, and Sen Yu. Simultaneous real diagonalization of rectangular quaternionic matrix pairs and its algorithm. Computers and Mathematics with Applications, 56(6):1538–1544, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800179X.

Caballero-Gil:2000:NUB

[CG00] P. Caballero-Gil. New upper bounds on the linear complexity. Computers and Mathematics with Applications, 39(3–4):31–38, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003314.

Chen:2001:IDN

[CG01] Yu-Fu Chen and Xiao-Shan Gao. Involutive directions and new involutive divisions. Computers and Mathematics with Applications, 41(7–8):945– 956, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003321.

Caboussat:2005:TGP

[CG05] A. Caboussat and R. Glowinski. A two-grids/projection algorithm for obstacle problems. Computers and Mathematics with Applications, 50(1–2):171–178, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002403.

Candito:2008:MSD

[CG08] Pasquale Candito and Nicola Giovannelli. Multiple solutions for a discrete boundary value problem involving the p-Laplacian. Computers and Mathematics with Applications, 56(4):959–964, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000989.

Chen:2003:TET

[CGBJR03] B. Chen, R. García-Bolós, L. Jódar, and M. D. Roselló. The truncation error of the two-variable Chebyshev series expansions. Computers and Mathematics with Applications, 45(10–11):1647–1653, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001445.

Caballero-Gil:2009:SAS

[CGFS09] P. Caballero-Gil and A. Fúster-Sabater. A simple attack on some clock-controlled generators. Computers and Mathematics with Applications, 58(1):179–188, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900251X.

Chen:2002:MCR

[CGGC02] C. S. Chen, M. Ganesh, M. A. Golberg, and A. H.- D. Cheng. Multilevel compact radial functions based computational schemes for some elliptic problems. Computers and Mathematics with Applications, 43 (3–5):359–378, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002929.

Chen:2001:AMT

[CGL01] William Y. C. Chen, H. W. Galbraith, and J. D. Louck. Angular momentum theory, umbral calculus, and combinatorics. Computers and Mathematics with Applications, 41(9):1199– 1214, May 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101000918.

Capriz:2008:MG

[CGM08] Gianfranco Capriz, Pasquale Giovine, and Paolo Maria Mariano. Modeling granularity. Computers and Mathematics with Applications, 55(2):131, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003410.

Contreras:2008:MPC

[CGPB08] Wilson A. Contreras, Damián Ginestar, Lourival C. Paraíba, and Rafael Bru. Modelling the pesticide concentration in a rice field by a level IV fugacity model coupled with a dispersion-advection equation. Computers and Mathematics with Applications, 56(3):657–669, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000564.

Ceng:2008:IAF

[CGY08a] Lu-Chuan Ceng, Sy-Ming Guu, and Jen-Chih Yao. Iterative algorithm for finding approximate solutions of mixed quasi-variational-like inclusions. Computers and Mathematics with Applications, 56(4):942–952, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000965.

Ceng:2008:IAS

[CGY08b] Lu-Chuan Ceng, Sy-Ming Guu, and Jen-Chih Yao. Iterative approximation of solutions for a class of completely generalized setvalued quasi-variational inclusions. Computers and Mathematics with Applications, 56(4):978–987, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001004.

Ceng:2009:HVL

[CGY09a] Lu-Chuan Ceng, Sy-Ming Guu, and Jen-Chih Yao. Hybrid viscosity-like approximation methods for nonexpansive mappings in Hilbert spaces. Computers and Mathematics with Applications, 58(3):605–617, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900265X.

Ceng:2009:GIV

[CGY09b] Lu-Chuan Ceng, Sy-Ming Guu, and Jen-Chih Yao. On generalized implicit vector equilibrium problems in Banach spaces. Computers and Mathematics with Applications, 57(10):1682– 1691, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002065.

Carl:2000:ERF

[CH00a] S. Carl and S. Heikkilä. Extremality results for firstorder discontinuous functional differential equations. Computers and Mathematics with Applications, 40 (10–11):1217–1232, November/December 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100002339.

Choi:2000:HBL

[CH00b] Y. K. Choi and K. S. Hwang.

How big are the lag increments of a Gaussian process? Computers and Mathematics with Applications, 40(8–9):911–919, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850026.

Chyan:2000:MSO

[CH00c] Chuan Jen Chyan and J. Henderson. Multiple solutions for 2m^{th}-order Sturm–Liouville boundary value problems. Computers and Mathematics with Applications, 40(2–3):231–237, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001565.

Cabada:2001:UCE

[CH01] A. Cabada and S. Heikkilä.

Uniqueness, comparison, and existence results for third-order initial boundary value problems. Computers and Mathematics with Applications, 41(5–6):607– 618, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003047.

Clark:2003:DCM

[CH03] S. Clark and D. Hinton. Disconjugacy criteria for matrix difference equations. Computers and Mathematics with Applications, 45(6–9):1013–1020, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000786.

Cao:2004:PSH

[CH04] Jinde Cao and Guangming He. Periodic solutions for higher-order neutral differential equations with several delays. Computers and Mathematics with Applications, 48(10– 11):1491–1503, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104004043.

Chang:2005:AFW

[CH05a] Ping-Teng Chang and Kuo-Chen Hung. Applying the fuzzy-weighted-average approach to evaluate network security systems. Computers and Mathematics with Applications, 49(11–12):1797– 1814, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001975.

Chen:2005:EGA

[CH05b] Y. Chen and L. Huang.

Existence and global attractivity of a positive periodic solution of a delayed periodic respiration model. Computers and Mathematics with Applications, 49(5–6):677–687, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000829.

Chang:2001:PMT

[Cha01a] Da-Wei Chang. The parallel multisplitting TOR (MTOR) method for linear systems. Computers and Mathematics with Applications, 41(1–2):215–227, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101850173.

Chapman:2001:RPR

[Cha01b] R. J. Chapman. Regarding “Powersums representing residues mod p^{k}, from Fermat to Waring”. Computers and Mathematics with Applications, 41(12):1473– 1474, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001109.

See [Ben00] and reply

[Ben01].

Chang:2006:CEI

[Cha06] Wei-Der Chang. Coefficient estimation of IIR filter by a multiple crossover genetic algorithm. Computers and Mathematics with Applications, 51(9–10):1437– 1444, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000903.

Chaudhry:2008:ETR

[Cha08] M. Aslam Chaudhry. Extension of the Titchmarsh representation of the Mertens function M(x), and numerical support of the Riemann hypothesis. Computers and Mathematics with Applications, 56(10):2638–2642, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004021.

Chen:2000:DQE

[Che00a] Chang-New Chen. Differential quadrature element analysis using extended differential quadrature. Computers and Mathematics with Applications, 39(5–6): 65–79, March 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210000047X.

Cheng:2000:DQF

[Che00b] Xiao-Liang Cheng. A discontinuous quadratic finite element method for onedimensional diffusion problems. Computers and Mathematics with Applications, 40(8–9):921–925, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850038.

Chen:2001:OLS

[Che01a] R. Chen. Optimal location of a single facility with circular demand areas. Computers and Mathematics with Applications, 41(7–8):1049– 1061, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003394.

Chen:2001:PHW

[Che01b] Shaozhu Chen. A problem of Hartman and Wintner: approximation for discrete perturbations. Computers and Mathematics with Applications, 42(3–5):655–669, August/September 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122101001857.

Cheng:2001:ISO

[Che01c] Xiao-Liang Cheng. Improvement of some Ostrowski-Grüss type inequalities. Computers and Mathematics with Applications, 42 (1–2):109–114, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122101001353.

Chernyavskaya:2003:SAS

[Che03] N. A. Chernyavskaya. Stability and absolute stability of a three-point difference scheme. Computers and Mathematics with Applications, 45(6–9):1181–1194, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210300083X.

Chen:2004:EGR

[Che04a] C.-N. Chen. Extended GDQ and related discrete element analysis methods for transient analyses of continuum mechanics problems. Computers and Mathematics with Applications, 47(1):91– 99, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900089.

Chen:2004:PIV

[Che04b] Hongqiu Chen. Periodic initial-value problem for BBM-equation. Computers and Mathematics with Applications, 48(9):1305–1318, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003773.

Chen:2004:AVM

[Che04c] Shih-Pin Chen. An alternating variable method with varying replications for simulation response optimization. Computers and Mathematics with Applications, 48(5–6):769–778, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003128.

Chen:2004:WLS

[Che04d] Tsu-Fen Chen. Weighted least-squares approximationsto nonlinear hyperbolic equations. Computers and Mathematics with Applications, 48(7–8):1059–1076, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003451.

Chen:2004:BSH

[Che04e] Youpeng Chen. Blow-up for a system of heat equations with nonlocal sources and absorptions. Computers and Mathematics with Applications, 48(3–4):361– 372, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840604.

Chen:2004:CNF

[Che04f] Zhangxin Chen. Characteristicnonconforming finite-element methods for advectiondominated diffusion problems. Computers and Mathematics with Applications, 48(7–8):1087–1100, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003475.

Cheng:2004:GOA

[Che04g] Chi-Bin Cheng. Group opinion aggregation based on a grading process: a method for constructing triangular fuzzy numbers. Computers and Mathematics with Applications, 48(10– 11):1619–1632, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004158.

Cheng:2006:SAS

[Che06] Daizhan Cheng. Some applications of semitensor product of matrices in algebra. Computers and Mathematics with Applications, 52(6– 7):1045–1066, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210600277X.

Cheng:2008:RCC

[Che08a] Chao-Jung Cheng. Robust control of a class of neural networks with bounded uncertainties and time-varying delays. Computers and Mathematics with Applications, 56(5):1245–1254, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800148X.

Cheng:2008:SSB

[Che08b] Chi-Bin Cheng. Solving a sealed-bid reverse auction problem by multiplecriterion decision-making methods. Computers and Mathematics with Applications, 56(12):3261–3274, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004896.

Chen:2009:MPS

[CHL09] Taolue Chen, Tingting Han, and Jian Lu. On metrics for probabilistic systems: Definitions and algorithms. Computers and Mathematics with Applications, 57(6):991–999, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005737.

Choi:2006:LRF

[CHM06] Yong-Kab Choi, Kyo-Shin Hwang, and Hee-Tin Moon. Limit results on finitedimensional Gaussian random fields. Computers and Mathematics with Applications, 52(3–4):485–496, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002148.

Chou:2003:CRM

[Cho03] Ch.-Ch Chou. The canonical representation of multiplication operation on triangular fuzzy numbers. Computers and Mathematics with Applications, 45 (10–11):1601–1610, May/ June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001391.

Chong:2008:EQ

[Cho08] Terence Tai-Leung Chong.

The empirical quest for π. Computers and Mathematics with Applications, 56 (10):2772–2778, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004306.

Cuevas:2009:ESI

[CHR09] Claudio Cuevas, Eduardo Hernández, and Marcos Rabelo. The existence of solutions for impulsive neutral functional differential equations. Computers and Mathematics with Applications, 58(4):744–757, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003034.

Campbell:2002:MFD

[CHS02] J. C. Campbell, J. M. Hyman, and M. J. Shashkov. Mimetic finite difference operators for second-order tensors on unstructured grids. Computers and Mathematics with Applications, 44(1–2):157–173, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001372.

Chun:2005:IMI

[Chu05] Changbum Chun. Iterative methods improving Newton’s method by the decomposition method. Computers and Mathematics with Applications, 50(10– 12):1559–1568, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105004074.

Chun:2007:GCI

[Chu07] Changbum Chun. A geometric construction of iterative functions of order three to solve nonlinear equations. Computers and Mathematics with Applications, 53(6):972–976, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700051X.

Chen:2007:ORP

[CHW07] Cheng-Kang Chen, Ta-Wei Hung, and Tzu-Chun Weng. Optimal replenishment policies with allowable shortages for a product life cycle. Computers and Mathematics with Applications, 53(10):1582– 1594, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001307.

Chen:2001:GCT

[CHY01] Hung-Yih Chen, Huei-Lin Hong, and Cheh-Chih Yeh. A general comparison theorem for nonlinear equations. Computers and Mathematics with Applications, 42(6–7):917–925, September/October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002085.

Chen:2009:AGP

[CHZ09] Zhe Chen, Haiqiao Huang, and Kequan Zhao. Approximate generalized proximaltype method for convex vector optimization problem in Banach spaces. Computers and Mathematics with Applications, 57(7):1196– 1203, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000601.

Cvetkovic-Iliic:2006:RSM

[CI06] D. S. Cvetković-Iliíc. The reflexive solutions of the matrix equation AXB = C. Computers and Mathematics with Applications, 51(6–7):897–902, March/ April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000459.

Cinlar:2003:CLP

[Çin03]  E. Çinlar.  Conditional

lévy processes. Computers and Mathematics with Applications, 46(7):993–997, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901131.

Cipu:2008:CAP

[Cip08] Mihai Cipu. A computeraided proof of a conjecture in Euclidean geometry. Computers and Mathematics with Applications, 56(11):2814–2818, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004574.

Chen:2004:TNC

[CJ04] Guizhi Chen and Zhong Xiao Jia. Theoretical and numerical comparisons of GM-RES and WZ–GMRES. Computers and Mathematics with Applications, 47(8–9):1335–1350, April/ May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901277.

Cai:2005:BPS

[CJ05] Ming-Chao Cai and Xiao-Qing Jin. BCCB preconditioners for solving linear systems from delay differential equations. Computers and Mathematics with Applications, 50(1–2):281– 288, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002506.

Caiazzo:2008:BFL

[CJ08] Alfonso Caiazzo and Michael Junk. Boundary forces in lattice Boltzmann: Analysis of momentum exchange algorithm. Computers and Mathematics with Applications, 55(7):1415– 1423, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006281.

Chun:2009:NFN

[CjBN09] Changbum Chun, Hwa ju Bae, and Beny Neta. New families of nonlinear third-order solvers for finding multiple roots. Computers and Mathematics with Applications, 57(9):1574– 1582, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006226.

Chun:2009:NMW

[CJK09] Changbum Chun, Hossein Jafari, and Yong-Il Kim. Numerical method for the wave and nonlinear diffusion equations with the homotopy perturbation method. Computers and Mathematics with Applications, 57(7):1226– 1231, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000315.

Cai:2004:AGG

[CJL04] Xian-Xin Cai, Bonan Jiang, and Guojun Liao. Adaptive grid generation based onthe least-squares finite-element method. Computers and Mathematics with Applications, 48(7–8):1077–1085, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003463.

Casaban:2005:DFM

[CJO05] M. C. Casabán, L. Jódar, and G. A. Ossandón. A discrete Fourier method for numerical solution of strongly coupled mixed parabolic systems. Computers and Mathematics with Applications, 49(4):611–624, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000544.

Caiazzo:2009:CAT

[CJR09] Alfonso Caiazzo, Michael Junk, and Martin Rheinländer. Comparison of analysis techniques for the lattice Boltzmann method. Computers and Mathematics with Applications, 58(5): 883–897, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109000935.

Cortes:2007:MSN

[CJV07] J. C. Cortés, L. Jódar, and L. Villafuerte. Mean square numerical solution of random differential equations: Facts and possibilities. Computers and Mathematics with Applications, 53(7):1098–1106, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000685.

Chan:2002:NCS

[CK02a] C. Y. Chan and L. Ke. Numerical computations for singular semilinear elliptic boundary value problems. Computers and Mathematics with Applications, 43 (3–5):351–358, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002917.

Clark:2002:CSR

[CK02b] D. Clark and M. R. S. Kulenović. A coupled system of rational difference equations. Computers and Mathematics with Applications, 43(6–7):849–867, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003261.

Cooke:2002:EPS

[CK02c] C. H. Cooke and J. Kroll.

The existence of periodic solutions to certain impulsive differential equations. Computers and Mathematics with Applications, 44(5–6):667–676, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001815.

Cao:2003:PLE

[CK03] T. Cao and D. W. Kelly.

Pointwise and local error estimates for the quantities of interest in twodimensional elasticity. Computers and Mathematics with Applications, 46(1):69– 79, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900812.

Choi:2004:AEB

[CK04] Sung Kyu Choi and Nam Jip Koo. Asymptotic equivalence between two linear Volterra difference systems. Computers and Mathematics with Applications, 47(2–3):461–471, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900387.

Cho:2006:IPC

[CK06] Nak Eun Cho and Ji A. Kim. Inclusion properties of certain subclasses of analytic functions defined by a multiplier transformation. Computers and Mathematics with Applications, 52(3–4):323–330, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001994.

Chang:2008:FIE

[CK08] Ick-Soon Chang and Sheon-Young Kang. Fredholm integral equation method for the integro-differential Schrödinger equation. Computers and Mathematics with Applications, 56(10): 2676–2685, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108004070.

Cao:2003:DEM

[CKCG03] Hongqing Cao, Lishan Kang, Yuping Chen, and Tao Guo. The dynamic evolutionary modeling of HODEs for time series prediction. Computers and Mathematics with Applications, 46(8– 9):1397–1411, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902288.

Cho:2006:NED

[CKK06] Chung-Ki Cho, Sungkwon Kang, and YongHoon Kwon. Numerical estimation of diffusivity in a nonhysteretic infiltration problem. Computers and Mathematics with Applications, 52(10– 11):1511–1528, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003361.

Cho:2006:PEA

[CKKK06] Chung-Ki Cho, Sunbu Kang, Taekkeun Kim, and Yonghoon Kwon. Parameter estimation approach to the free boundary for the pricing of an American call option. Computers and Mathematics with Applications, 51(5):713–720, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000198.

Chang:2006:PDP

[CKL06] Yeon-Woo Chang, Hyea Hyun Kim, and Chang-Ock Lee. Preconditioners for the dual-primal FETI methods on nonmatching grids: Numerical study. Computers and Mathematics with Applications, 51(5):697–712, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000186.

Cho:2007:CLU

[CKL07] G. M. Cho, M. K. Kim, and Y. H. Lee. Complexity of large-update interior point algorithm for P_{⋆}(κ) linear complementarity problems. Computers and Mathematics with Applications, 53(6):948–960, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000478.

Chang:2004:MQV

[CKN04] S. S. Chang, J. K. Kim, and Y. M. Nam. Multivalued quasi-variational inclusions and multivalued accretive equations. Computers and Mathematics with Applications, 48(10– 11):1441–1452, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003992.

Cho:2008:ACF

[CKQ08] Yeol Je Cho, Shin Min Kang, and Xiaolong Qin. Approximation of common fixed points of an infinite family of nonexpansive mappings in Banach spaces. Computers and Mathematics with Applications, 56(8):2058–2064, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003404.

Choi:2003:AEB

[CKR03] Sung Kyu Choi, Nam Jip Koo, and Hyun Sook Ryu. Asymptotic equivalence between two difference systems. Computers and Mathematics with Applications, 45(6–9):1327–1337, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001068.

Chadli:2004:DME

[CKY04] O. Chadli, I. V. Konnov, and J. C. Yao. Descent methods for equilibrium problems in a Banach space. Computers and Mathematics with Applications, 48(3–4):609–616, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210484080X.

Chen:2007:CRA

[CKY07] Chien-Yuan Chen, Cheng-Yuan Ku, and David C. Yen. Cryptographic relational algebra for databases using the field authenticator. Computers and Mathematics with Applications, 54(1): 38–44, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001721.

Choi:2001:OPD

[CKZ01] Sung Kyu Choi, Nam Jip Koo, and Binggen Zhang. The oscillation of partial difference equations with continuous arguments. Computers and Mathematics with Applications, 41(12): 1495–1503, June 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101001146.

Chang:2000:MMC

[CL00] Ping-Teng Chang and E. S.

Lee. A multisolution method for cell formation

— exploring practical alternatives in group technology manufacturing. Computers and Mathematics with Applications, 40(10– 11):1285–1296, November/ December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000239X.

Chen:2001:AAR

[CL01] Liang-Hsuan Chen and Hai-Wen Lu. An approximate approach for ranking fuzzy numbers based on left and right dominance. Computers and Mathematics with Applications, 41(12):1589– 1602, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001249.

Cao:2002:NES

[CL02a] Jianwen Cao and Choi-Hong Lai. Numerical experiments of some Krylov subspace methods for black oil model. Computers and Mathematics with Applications, 44(1–2):125–141, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001359.

Chen:2002:POF

[CL02b] L.-H. Chen and H.-W. Lu. The preference order of fuzzy numbers. Computers and Mathematics with Applications, 44(10– 11):1455–1465, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002705.

Cai:2007:PSJ

[CL07a] Xu-Chu Cai and Meng-Su Li. Periodic solution of Jacobi elliptic equations by He’s perturbation method. Computers and Mathematics with Applications, 54(7–8):1210–1212, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002921.

Chang:2007:ERD

[CL07b] Yong-Kui Chang and Wan-Tong Li. Existence results for dynamic inclusions on time scales with nonlocal initial conditions. Computers and Mathematics with Applications, 53(1):12–20, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000181.

Chen:2008:TPB

[CL08a] Haibo Chen and Peiluan Li. Three-point boundary value problems for second-order ordinary differential equations in Banach spaces. Computers and Mathematics with Applications, 56(7):1852–1860, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002654.

Chen:2008:BCS

[CL08b] Sun-Kang Chen and Huan-Wen Liu. A bivariate C^{1} cubic super spline space on Powell–Sabin triangulation. Computers and Mathematics with Applications, 56(5):1395–1401, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001636.

Chu:2009:EFM

[CL09] Ta-Chung Chu and Yichen Lin. An extension to fuzzy MCDM. Computers and Mathematics with Applications, 57(3):445–454, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006408.

Chen:2001:RIM

[CLC01] Tze-Jang Chen, Jenn-Tsann Lin, and C. H. Cooke. The range of the iterated matrix adjoint operator. Computers and Mathematics with Applications, 41(1–2):97–102, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101850094.

Cao:2009:IMM

[CLJ09] Fuyuan Cao, Jiye Liang, and Guang Jiang. An initialization method for the K-Means algorithm using neighborhood model. Computers and Mathematics with Applications, 58 (3):474–483, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S089812210900323X.

Chang:2009:BCL

[CLL09] Cheng Chang, Chih-Hao Liu, and Chao-An Lin. Boundary conditions for lattice Boltzmann simulations with complex geometry flows. Computers and Mathematics with Applications, 58(5):940–949, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000996.

Calvo:2003:SAG

[CLM03] M. Calvo, M. P. Laburta, and J. I. Montijano. Starting algorithms for Gauss Runge–Kutta methods for Hamiltonian systems. Computers and Mathematics with Applications, 45(1–3): 401–410, January/February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800263.

Casciola:2006:SPS

[CLMM06] G. Casciola, D. Lazzaro, L. B. Montefusco, and S. Morigi. Shape preserving surface reconstruction using locally anisotropic radial basis function interpolants. Computers and Mathematics with Applications, 51(8):1185–1198, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000721.

Choi:2007:LSP

[CLS07] Youngmi Choi, Hyung-Chun Lee, and Byeong-Chun Shin. A least-squares/ penalty method for distributed optimal control problems for Stokes equations. Computers and Mathematics with Applications, 53(11):1672–1685, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700140X.

Cheng:2008:OCG

[CLS08] Eddie Cheng, László Lipták, and Nart Shawash. Orienting Cayley graphs generated by transposition trees. Computers and Mathematics with Applications, 55(11):2662–2672, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007651.

Chen:2008:SLB

[CLT^{+}08] Sheng Chen, Zhaohui Liu, Zhiwei Tian, Baochang Shi, and Chuguang Zheng. A simple lattice Boltzmann scheme for combustion simulation. Computers and Mathematics with Applications, 55(7):1424– 1432, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700627X.

Chen:2009:HOW

[CLT09] C. R. Chen, S. J. Li, and K. L. Teo. Higher order weak epiderivatives and applications to duality and optimality conditions. Computers and Mathematics with Applications, 57(8):1389–1399, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000297.

Chen:2008:GCQ

[CLX08] Ren Chen, Xiaonan Luo, and Hao Xu. Geometric compression of a quadrilateral mesh. Computers and Mathematics with Applications, 56(6):1597–1603, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001867.

Ceng:2008:APA

[CLY08] L. C. Ceng, T. C. Lai, and J. C. Yao. Approximate proximal algorithms for generalized variational inequalities with paramonotonicity and pseudomonotonicity. Computers and Mathematics with Applications, 55(6):1262–1269, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005949.

Coles:2000:IPD

[CM00] C. Coles and D. A. Murio. Identification of parameters in the 2-D IHCP. Computers and Mathematics with Applications, 40 (8–9):939–956, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100850051.

Calio:2001:ABQ

[CM01a] F. Caliò and E. Marchetti. An algorithm based on Q.I. modified splines for singular integral models. Computers and Mathematics with Applications, 41(12):1579– 1588, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001237.

Coles:2001:SSD

[CM01b] C. Coles and D. A. Murio. Simultaneous space diffusivity and source term reconstruction in 2D IHCP. Computers and Mathematics with Applications, 42 (12):1549–1564, December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101002619.

Cong:2002:PIP

[CM02] N. H. Cong and N. T. H.

Minh. Parallel-iterated pseudo two-step Runge– Kutta–Nyström methods for nonstiff second-order IVPs. Computers and Mathematics with Applications, 44(1–2):143–155, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001360.

Choi:2003:VFC

[CM03] S. Choi and M. D. Marcozzi.

The valuation of foreign currency options under stochastic interest rates. Computers and Mathematics with Applications, 46(5–6):741–749, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901386.

Cash:2004:HOI

[CM04] J. R. Cash and D. R.

Moore. High-order interpolants for solutions of twopoint boundary value problems using MIRK methods. Computers and Mathematics with Applications, 48 (10–11):1749–1763, November/December 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004250.

Coles:2006:PED

[CM06] C. Coles and D. Murio. Parameter estimation for a drying system in a porous medium. Computers and Mathematics with Applications, 51(9–10):1519– 1528, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000988.

Cuomo:2008:NAN

[CM08] S. Cuomo and A. Marasco. A numerical approach to nonlinear two-point boundary value problems for ODEs. Computers and Mathematics with Applications, 55(11):2476–2489, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007481.

Caiazzo:2009:LBB

[CM09] Alfonso Caiazzo and Shankar Maddu. Lattice Boltzmann boundary conditions via singular forces: Irregular expansion analysis and numerical investigations. Computers and Mathematics with Applications, 58(5):930–939, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000984.

Chen:2009:CLF

[CML^{+}09] Zhumin Chen, Jun Ma, Jingsheng Lei, Bo Yuan, Li Lian, and Ling Song. A cross-language focused crawling algorithm based on multiple relevance prediction strategies. Computers and Mathematics with Applications, 57(6):1057– 1072, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005063.

Charina:2004:MVS

[CMS04] M. Charina, A. J. Meir, and P. G. Schmidt. Mixed velocity, stress, current, and potential boundary conditions for stationary MHD flow. Computers and Mathematics with Applications, 48(7–8):1181–1190, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003542.

Castaings:2001:MRS

[CN01] W. Castaings and I. M. Navon. Mesh refinement strategies for solving singularly perturbed reactiondiffusion problems. Computers and Mathematics with Applications, 41(1–2):157– 176, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101850148.

Costabile:2004:SCC

[CN04] F. Costabile and A. Napoli. Stability of Chebyshev collocation methods. Computers and Mathematics with Applications, 47(4–5):659– 666, February/March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104900533.

Chun:2008:SMN

[CN08a] Changbum Chun and Beny Neta. Some modification of Newton’s method by the method of undetermined coefficients. Computers and Mathematics with Applications, 56(10):2528–2538, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003891.

Csomos:2008:OSD

[CN08b] Petra Csomós and Gregor Nickel. Operator splitting for delay equations. Computers and Mathematics with Applications, 55(10):2234–2246, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007419.

Chowdhury:2000:HM

[CNK00] R. A. Chowdhury, S. K.

Nath, and M. Kaykobad. The heap-mergesort. Computers and Mathematics with Applications, 39(7–8): 193–197, April 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100000754.

Company:2008:NSL

[CNPP08] Rafael Company, Enrique Navarro, José Ramón Pintos, and Enrique Ponsoda. Numerical solution of linear and nonlinear Black–Scholes option pricing equations. Computers and Mathematics with Applications, 56(3):813–821, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000795.

Castillo:2000:NTT

[CO00] J. E. Castillo and J. S. Otto. Numerical techniques for the transformation to an orthogonal coordinate system aligned with a vector field. Computers and Mathematics with Applications, 40(4–5):523–535, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001772.

Chen:2007:PBV

[CO07] Jinhai Chen and Donal O’Regan. The periodic boundary value problem for semilinear elastic beam equations: the resonance case. Computers and Mathematics with Applications, 53(8):1284–1292, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000879.

Cabada:2003:FSS

[COE03] A. Cabada and V. Otero-Espinar. Fixed sign solutions of second-order difference equations with Neumann boundary conditions. Computers and Mathematics with Applications, 45(6–9):1125–1136, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000713.

Cabada:2000:EAS

[COEP00] A. Cabada, V. Otero-Espinar, and R. L. Pouso. Existence and approximation of solutions for firstorder discontinuous difference equations with nonlinear global conditions in the presence of lower and upper solutions. Computers and Mathematics with Applications, 39(1–2):21–33, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003107.

Cohen:2000:DBP

[Coh00] M. J. Cohen. The difference between the product of n consecutive integers and the n^{th} power of an integer. Computers and Mathematics with Applications, 39(11):139– 157, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001218.

Coles:2002:PED

[Col02] C. Coles. Parameter estimation in the 1-D transport equation with advection. Computers and Mathematics with Applications, 44(12):1493–1502, December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002730.

Coles:2004:PIU

[Col04] C. Coles. Parameter identification using mollification for predator–prey models in spatially heterogeneous environments. Computers and Mathematics with Applications, 48(3–4):505– 515, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840720.

Constantin:2003:FVF

[Con03] P. Constantin. Filtered viscous fluid equations. Computers and Mathematics with Applications, 46(4): 537–546, August 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0898122103900150.

Cooke:2005:BEO

[Coo05] C. H. Cooke. Bounds on element order in rings Z_{m} with divisors of zero. Computers and Mathematics with Applications, 49(11–12):1643– 1645, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001859.

Cambazard:2009:CBA

[COS09] Hadrien Cambazard, Barry O’Sullivan, and Barbara M. Smith. A constraintbased approach to enigma 1225. Computers and Mathematics with Applications, 58(8):1487–1497, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002855.

Castillo:2001:AFN

[CP01a] S. Castillo and M. Pinto.

Asymptotic formulae for nonlinear functional difference equations. Computers and Mathematics with Applications, 42(3–5):551–559, August/September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0898122101001778.

Cuevas:2001:APS

[CP01b] C. Cuevas and M. Pinto. Asymptotic properties of solutions to nonautonomous Volterra difference systems with infinite delay. Computers and Mathematics with Applications, 42(3–5):671–685, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001869.

Castillo:2002:ATN

[CP02] S. Castillo and M. Pinto. An asymptotic theory for nonlinear functional differential equations. Computers and Mathematics with Applications, 44(5–6):763–775, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210200189X.

Cercignani:2003:P

[CP03] Carlo Cercignani and Ugo Piomelli. Preface. Computers and Mathematics with Applications, 46(4):xiii–xiv, August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900137.

Chen:2004:ESC

[CP04] Z. Chen and V. Pan. An efficient solution for Cauchylikesystems of linear equations. Computers and Mathematics with Applications, 48(3–4):529–537, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840744.

Chang:2000:AEC

[CPC00] Shih-Sen Chang, Jong Yeoul Park, and Yu Qing Chen. Accretive equations in cones of Banach space. Computers and Mathematics with Applications, 39(5–6):55– 64, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000468.

Carlone:2007:PMP

[CPP07] P. Carlone, G. S. Palazzo, and R. Pasquino. Pultrusion manufacturing process development: Cure optimization by hybrid computational methods. Computers and Mathematics with Applications, 53(9):1464– 1471, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001265.

Carlone:2008:IAL

[CPP08] P. Carlone, G. S. Palazzo, and R. Pasquino. Inverse analysis of the laser forming process by computational modelling and methods. Computers and Mathematics with Applications, 55(9):2018–2032, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006839.

Costa:2005:SQB

[CPS05] C. Costa, T.-P. Azevedo Perdicoúlis, and R. Serôdio. A study on quaternion blockquasi-tridiagonal systems. Computers and Mathematics with Applications, 50(7):1011–1016, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003305.

Chartrand:2000:RUD

[CPZ00] G. Chartrand, C. Poisson, and P. Zhang. Resolvability and the upper dimension of graphs. Computers and Mathematics with Applications, 39(12): 19–28, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001267.

Chen:2003:PPS

[CQJ03] Shaozhu Chen, Jiangang Qi, and Mingzhong Jin. Pulse phenomena of secondorder impulsive differential equations with variable moments. Computers and Mathematics with Applications, 46(8– 9):1281–1287, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902197.

Chen:2002:SMS

[CQT02] Jing-Bo Chen, Meng-Zhao Qin, and Yi-Fa Tang. Symplectic and multi-symplectic methods for the nonlinear Schrödinger equation. Computers and Mathematics with Applications, 43(8–9):1095–1106, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800153.

Cherkas:2006:CCL

[CR06] L. A. Cherkas and V. G.

Romanovski. The center conditions for a Liénard system. Computers and Mathematics with Applications, 52(3–4):363–374, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002033.

Chowdhury:2002:BMM

[CRK02] R. A. Chowdhury, M. Z. Rahman, and M. Kaykobad. The bounds of min-max pair heap construction. Computers and Mathematics with Applications, 43(6–7):911–916, March/ April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101003315.

Ciric:2008:CFP

[ĆRRU08] Ljubomir Ćirić, A. Razani, Stojan Radenović, and J. S. Ume. Common fixed point theorems for families of weakly compatible maps. Computers and Mathematics with Applications, 55(11):2533– 2543, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007547.

Chalmers:2000:SEA

[CS00a] B. L. Chalmers and B. Shekhtman. Some estimates of action constants and related parameters. Computers and Mathematics with Applications, 40(1):71– 79, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001413.

Chang:2000:MSA

[CS00b] Chin-Chen Chang and Jau-Ji Shen. A multichannel storage arrangement scheme of pictorial databases for concurrent spatial match retrieval. Computers and Mathematics with Applications, 39(5–6):239–248, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000572.

Casals:2001:ELS

[CS01] J. Casals and S. Sotoca.

The exact likelihood for a state space model with stochastic inputs. Computers and Mathematics with Applications, 42(1–2):199– 209, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001444.

Cao:2002:SON

[CS02a] Yanzhao Cao and D. Stanescu.

Shape optimization for noise radiation problems. Computers and Mathematics with Applications, 44(12): 1527–1537, December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122102002766.

Chalabi:2002:CRS

[CS02b] A. Chalabi and D. Seghir. Convergence of relaxation schemes for initial boundary value problems for conservation laws. Computers and Mathematics with Applications, 43(8–9):1079–1093, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800141.

Chiarella:2002:ABD

[CS02c] C. Chiarella and F. Szidarovszky. The asymptotic behavior of dynamic rentseeking games. Computers and Mathematics with Applications, 43(1–2):169– 178, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002814.

Chen:2004:PLT

[CS04a] Fengde Chen and Jinlin Shi. Periodicity in a logistic type system with several delays. Computers and Mathematics with Applications, 48(1–2):35– 44, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839968.

Chen:2004:LCL

[CS04b] Jingnian Chen and Yuming Shi. The limit circle and limit point criteria for second-order linear difference equations. Computers and Mathematics with Applications, 47(6–7):967–976, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900806.

Calvin:2005:ODG

[CS05a] J. M. Calvin and A. Silinskas. One-dimensional global optimization for observations with noise. Computers and Mathematics with Applications, 50(1–2): 157–169, July 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105002397.

Chiarella:2005:COP

[CS05b] C. Chiarella and F. Szidarovszky. Cournot oligopolies with product differentiation under uncertainty. Computers and Mathematics with Applications, 50(3–4):413– 424, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002853.

Chauhan:2008:MEG

[CS08] Sanchita Chauhan and V. K. Srivastava. Modeling exhaust gas pollution abatement: Part I- single hydrocarbon propylene. Computers and Mathematics with Applications, 55(3):319–330, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000491.

Csallner:2001:LCT

[Csa01] A. E. Csallner. Lipschitz continuity and the termination of interval methods for global optimization. Computers and Mathematics with Applications, 42 (8–9):1035–1042, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210100219X.

Choudhury:2009:IGD

[CSC^{+}09] Pabitra Pal Choudhury, Sudhakar Sahoo, Mithun Chakraborty, Subir Kumar Bhandari, and Amita Pal. Investigation of the global dynamics of cellular automata using Boolean derivatives. Computers and Mathematics with Applications, 57(8):1337–1351, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000479.

Chen:2008:NRS

[CSdS08] Jun Chen, Wenyu Sun, and Raimundo J. B. de Sampaio. Numerical research on the sensitivity of nonmonotone trust region algorithms to their parameters. Computers and Mathematics with Applications, 56(11):2932–2940, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003994.

Csenki:2006:EPA

[Cse06] A. Csenki. Enigma 1225:

Prolog-assisted solution of a puzzle using discrete mathematics. Computers and Mathematics with Applications, 52(3–4):383–400, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002057.

Chou:2007:FFM

[CSLC07] Shuo-Yan Chou, Chun-Ying Shen, Shih-Wei Lin, and Yi-Kuang Chen. FALSCAL: a fuzzy multidimensional scaling algorithm. Computers and Mathematics with Applications, 53(5):717–728, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000247.

Cortes:2005:ANA

[CSPJ05] J. C. Cortés, P. Sevilla-Peris, and L. Jódar. Analyticnumerical approximating processes of diffusion equation with data uncertainty. Computers and Mathematics with Applications, 49(7–8):1255–1266, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001355.

Chen:2001:ACM

[CSX01] Hanlin Chen, Qiuhui Sheng, and Shaoliang Xiao. Accuracy conditions of multivariate refinable vector in frequency domain. Computers and Mathematics with Applications, 41(3–4):461–481, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002881.

Chen:2009:WAM

[CSZH09] Yidong Chen, Xiaodong Shi, Changle Zhou, and Qingyang Hong. A word alignment model based on multiobjective evolutionary algorithms. Computers and Mathematics with Applications, 57(11–12):1724– 1729, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005567.

Chen:2002:MIF

[CT02a] W. Chen and M. Tanaka.

A meshless, integrationfree, and boundary-only RBF technique. Computers and Mathematics with Applications, 43(3–5):379– 391, February/March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101002930.

Chu:2002:RFN

[CT02b] Ta-Chung Chu and Chung-Tsen Tsao. Ranking fuzzy numbers with an area between the centroid point and original point. Computers and Mathematics with Applications, 43(1–2):111– 117, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002772.

Cakmak:2005:OCC

[ÇT05] D. Çakmak and A. Tiryaki. Oscillation criteria for certain forced second-order nonlinear differential equations with delayed argument. Computers and Mathematics with Applications, 49(11–12):1647– 1653, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001860.

Chen:2008:RNN

[CT08] Chi-Chi Chen and Hui-Chin Tang. Ranking nonnormal p-norm trapezoidal fuzzy numbers with integral value. Computers and Mathematics with Applications, 56(9):2340–2346, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003726.

Carden:2009:SIT

[CT09] Russell L. Carden and Pablo Tarazaga. Sequential iterations for two diagonal preconditioners. Computers and Mathematics with Applications, 58(1):88– 94, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002077.

Chen:2005:PRS

[CTYW05] Degang Chen, E. C. C.

Tsang, Daniel S. Yeung, and Xizhao Wang. The parameterization reduction of soft sets and its applications. Computers and Mathematics with Applications, 49(5–6):757–763, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000891.

Cui:2002:EDP

[Cui02] Jingan Cui. The effect of dispersal on permanence in a predator–prey population growth model. Computers and Mathematics with Applications, 44 (8–9):1085–1097, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002171.

Cui:2009:MSF

[Cui09]  Yaqiong Cui.  Multiple

solutions to fourth-order boundary value problems. Computers and Mathematics with Applications, 57 (4):643–649, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108006275.

Chen:2000:IDS

[CVW00] Yunmei Chen, B. C. Vemuri, and Li Wang. Image denoising and segmentation via nonlinear diffusion. Computers and Mathematics with Applications, 39(5–6):131–149, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210000050X.

Chyan:2005:TSF

[CW05] Chuan Jen Chyan and Patricia J. Y. Wong. Triple solutions of focal boundary value problems on time scale. Computers and Mathematics with Applications, 49(7–8):963–979, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001148.

Chen:2006:RSB

[CW06] Chun-Bao Chen and Li-Ya Wang. Rough setbased clustering with refinement using Shannon’s entropy theory. Computers and Mathematics with Applications, 52(10– 11):1563–1576, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003403.

Chen:2008:QAD

[CW08] Bin Chen and Mingxin Wang. Qualitative analysis for a diffusive predator– prey model. Computers and Mathematics with Applications, 55(3):339–355, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003641.

Cai:2009:HFF

[CW09a] Xu-Chu Cai and Wen-Ying Wu. He’s frequency formulation for the relativistic harmonic oscillator. Computers and Mathematics with Applications, 58(11–12):2358–2359, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001710.

Wu:2009:UCS

[cW09b]  Guo cheng Wu.  Uni-

formly constructing soliton solutions and periodic solutions to Burgers–Fisher equation. Computers and Mathematics with Applications, 58(11–12):2355–2357, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001679.

Cheng:2001:QBS

[CWB01] Fuhua Cheng, Xuefu Wang, and B. A. Barsky. Quadratic B-spline curve interpolation. Computers and Mathematics with Applications, 41(1–2):39–50, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101850045.

Wu:2009:UCE

[cWcX09] Guo cheng Wu and Tie cheng Xia. Uniformly constructing exact discrete soliton solutions and periodic solutions to differentialdifference equations. Computers and Mathematics with Applications, 58(11– 12):2351–2354, December 2009. CODEN CMAPDK.

ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109001643.

Chen:2007:PGA

[CWL07] Fengde Chen, Liping Wu, and Zhong Li. Permanence and global attractivity of the discrete Gilpin–Ayala type population model. Computers and Mathematics with Applications, 53(8):1214– 1227, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000806.

Chen:2007:ETM

[CWLC07] J. T. Chen, C. S. Wu, Y. T.

Lee, and K. H. Chen. On the equivalence of the Trefftz method and method of fundamental solutions for Laplace and biharmonic equations. Computers and Mathematics with Applications, 53(6):851–879, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000193.

Chang:2009:LCB

[CWLC09] Te-Jen Chang, Chia-Long Wu, Der-Chyuan Lou, and Ching-Yin Chen. A low-complexity LUT -based squaring algorithm. Computers and Mathematics with Applications, 57(9): 1494–1501, May 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109000509.

Chen:2008:STN

[CWW08] Qiuhui Chen, Yanbo Wang, and Yi Wang. A sampling theorem for nonbandlimited signals using generalized sinc functions. Computers and Mathematics with Applications, 56(6):1650–1661, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001934.

Chen:2002:AGC

[CWY02] Juh-Yin Chen, Ngai-Ching Wong, and Jen-Chih Yao. Algorithm for generalized co-complementarity problems in Banach spaces. Computers and Mathematics with Applications, 43 (1–2):49–54, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210100270X.

Ceng:2008:FPS

[CWY08] Lu-Chuan Ceng, Ngai-Ching Wong, and Jen-Chih Yao. Fixed point solutions of variational inequalities for a finite family of asymptotically nonexpansive mappings without common fixed point assumption. Computers and Mathematics with Applications, 56(9):2312–2322, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003696.

Cheng:2003:SIC

[CWZ03] Lizhi Cheng, Hongxia Wang, and Zenghui Zhang. The solution of ill-conditioned symmetric Toeplitz systems via two-grid and wavelet methods. Computers and Mathematics with Applications, 46(5–6):793–804, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901428.

Cui:2006:OCA

[CWZ06] Hongquan Cui, Changchun Wang, and Quan Zheng. Optimality conditions and algorithms for integral global minimization. Computers and Mathematics with Applications, 52(1–2):

55–64, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001817.

Chen:2007:DPS

[CWZZ07] Haibo Chen, Haihua Wang, Qi Zhang, and Tiejun Zhou. Double positive solutions of boundary value problems for p-Laplacian impulsive functional dynamic equations on time scales. Computers and Mathematics with Applications, 53(10):1473– 1480, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000818.

Chen:2004:SEM

[CX04] Chuanmiao Chen and Ziqing Xie. Search extension method for multiple solutions of a nonlinear problem. Computers and Mathematics with Applications, 47(2–3):327–343, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900284.

Chen:2009:PMC

[CXZ09] Ruyun Chen, Shuhuang Xiang, and Yongxiong Zhou. A parameter method for computing highly oscillatory integrals. Computers and Mathematics with Applications, 58(9):1830–1837, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005987.

Chen:2000:NCO

[CY00] Tsu-Fen Chen and Haw-Dai Yang. Numerical construction of optimal adaptive grids in two spatial dimensions. Computers and Mathematics with Applications, 39(12):101– 120, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001334.

Cheng:2005:PCY

[CY05] Hon-Wing Cheng and Stephen S.-T. Yau. Parallel computations for Yau filters. Computers and Mathematics with Applications, 50(5–6):881–893, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003238.

Cai:2005:EPS

[CYG05] Xiaochun Cai, Jianshe Yu, and Zhiming Guo. Existence of periodic solutions for fourth-order difference equations. Computers and Mathematics with Applications, 50(1–2):49– 55, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002312.

Cai:2006:PSC

[CYG06] Xiaochun Cai, Jianshe Yu, and Zhiming Guo. Periodic solutions of a class of nonlinear difference equations via critical point method. Computers and Mathematics with Applications, 52(12):1639–1647, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003543.

Chang:2005:SAK

[CYH05] Ting-Yi Chang, Wei-Pang Yang, and Min-Shiang Hwang. Simple authenticated key agreement and protected password change protocol. Computers and Mathematics with Applications, 49(5–6):703–714, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000842.

Cho:2008:GNR

[CyL08] Yeol Je Cho and Heng you Lan. Generalized nonlinear random (A, η)-accretive equations with random relaxed cocoercive mappings in Banach spaces. Computers and Mathematics with Applications, 55(9):2173– 2182, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006979.

Cao:2009:MVU

[CYX09] Ting-Bin Cao, Hong-Xun Yi, and Hong-Yan Xu. On the multiple values and uniqueness of meromorphic functions on annuli. Computers and Mathematics with Applications, 58(7):1457–1465, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004830.

Ching:2000:CAP

[CZ00] Wai Ki Ching and Xun Yu Zhou. Circulant approximation for preconditioning in stochastic automata networks. Computers and Mathematics with Applications, 39(3–4):147–160, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003417.

Clausen:2002:SSI

[CZ02] J. Clausen and A. Zilinskas. Subdivision, sampling, and initialization strategies for simplical branch and bound in global optimization. Computers and Mathematics with Applications, 44(7):943–955, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002055.

Chen:2003:OCY

[CZ03a] Shaozhu Chen and Zhaowen Zheng. Oscillation criteria of Yan type for linear Hamiltonian systems. Computers and Mathematics with Applications, 46(5–6):855–862, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901489.

Chen:2003:GAO

[CZ03b] Yuming Chen and Zhan Zhou. Global attractivity and oscillation in a nonlinear periodic delay difference equation. Computers and Mathematics with Applications, 45(6–9):943–950, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000671.

Chen:2005:ERD

[CZ05a] Guanrong Chen and Zuohuan Zheng. An equivalent relationship in discrete dynamical systems. Computers and Mathematics with Applications, 49(9–10):1433– 1437, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001616.

Chen:2005:NMN

[CZ05b] Zhenmin Chen and Chi Zhang. A new measurement for network sharing fairness. Computers and Mathematics with Applications, 50(5–6):803–808, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003184.

Chen:2006:MTS

[CZ06] Yixiang Chen and Guo-Qiang Zhang. Maximality and totality of stable functions in the category of stable bifinite domains. Computers and Mathematics with Applications, 51(6–7):1011–1020, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000332.

Cui:2008:EPS

[CZ08] Yujun Cui and Yumei Zou. Existence of positive solutions for 2n thorder singular superlinear boundary value problems. Computers and Mathematics with Applications, 56(12):3195–3203, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004379.

Chen:2009:IFL

[CZ09a] Wenjuan Chen and Shunhua Zhang. Intuitionistic fuzzy Lie sub-superalgebras and intuitionistic fuzzy ideals. Computers and Mathematics with Applications, 58(8):1645–1661, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900443X.

Chen:2009:EAS

[CZ09b] Zhong Chen and YongFang Zhou. An efficient algorithm for solving Hilbert type singular integral equations of the second kind. Computers and Mathematics with Applications, 58(4):632–640, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003216.

Cui:2009:EUT

[CZ09c] Yujun Cui and Yumei Zou.

Existence and uniqueness theorems for fourth-order singular boundary value problems. Computers and Mathematics with Applications, 58(7):1449–1456, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004829.

Cho:2004:WSC

[CZG04] Yeol Je Cho, Haiyun Zhou, and Ginti Guo. Weak and strong convergence theorems for three-step iterations with errors for asymptotically nonexpansive mappings. Computers and Mathematics with Applications, 47(4–5):707–717, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900582.

Chen:2008:UEF

[CZLC08] Jun-Fan Chen, Xiao-Yu Zhang, Wei-Chuan Lin, and Sheng-Jiang Chen. Uniqueness of entire functions that share one value. Computers and Mathematics with Applications, 56(12):3000– 3014, December 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108004495.

Cao:2006:ROM

[CZLN06] Yanhua Cao, Jiang Zhu, Zhendong Luo, and I. M. Navon. Reduced-order modeling of the upper tropical Pacific ocean model using proper orthogonal decomposition. Computers and Mathematics with Applications, 52(8– 9):1373–1386, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003245.

Chen:2008:ICG

[CZW08] Yongzhu Chen, Weiyi Zhu, and Weifan Wang. Improper choosability of graphs of nonnegative characteristic. Computers and Mathematics with Applications, 56(8): 2073–2078, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108003428.

Doostparast:2006:SAG

[DA06] M. Doostparast and J. Ahmadi. Statistical analysis for geometric distribution based on records. Computers and Mathematics with Applications, 52 (6–7):905–916, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002653.

Dohi:2006:SEA

[DAKO06] Tadashi Dohi, Akira Ashioka, Naoto Kaio, and Shunji Osaki. Statistical estimation algorithms for repairs-time limit replacement scheduling under earning rate criteria. Computers and Mathematics with Applications, 51(2):345–356, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004724.

Darcy:2008:MPD

[Dar08] Isabel K. Darcy. Modeling protein–DNA complexes with tangles. Computers and Mathematics with Applications, 55(5):924–937, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005718.

Das:2004:LSG

[Das04] K. Ch. Das. The Laplacian spectrum of a graph. Computers and Mathematics with Applications, 48 (5–6):715–724, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104003074.

Das:2009:ASF

[Das09] S. Das. Analytical solution of a fractional diffusion equation by variational iteration method. Computers and Mathematics with Applications, 57(3):483–487, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006111.

Diaz-Barrero:2008:KFI

[DBEGB08] J. L. Díaz-Barrero, J. J. Egozcue, and J. Gibergans-Báguena. On the Ky fan inequality and some of its applications. Computers and Mathematics with Applications, 56(9):2279–2284, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003611.

Diaz-Barrero:2008:RAP

[DBGSP08] J. L. Díaz-Barrero, M. Grau-Sánchez, and P. G. Popescu. Refinements of Aczél, Popoviciu and Bellman’s inequalities. Computers and Mathematics with Applications, 56(9):2356–2359, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800374X.

Ding:2004:SPE

[DBX04] Shusen Ding, Gejun Bao, and Yuming Xing. Sobolev– Poincaré embeddings for operators on harmonic forms on manifolds. Computers and Mathematics with Applications, 47(2–3):259–270, January/February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104900235.

Davvaz:2009:FAS

[DCLF09] B. Davvaz, P. Corsini, and V. Leoreanu-Fotea. Fuzzy nary subpolygroups. Computers and Mathematics with Applications, 57(1):141–152, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800429X.

Dutta:2007:IMS

[DCR07] Pankaj Dutta, Debjani Chakraborty, and A. R. Roy. An inventory model for single-period products with reordering opportunities under fuzzy demand. Computers and Mathematics with Applications, 53(10):1502– 1517, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001186.

Deng:2009:SSH

[DCS09] Xiaoqing Deng, Gong Cheng, and Haiping Shi. Subharmonic solutions and homoclinic orbits of second order discrete Hamiltonian systems with potential changing sign. Computers and Mathematics with Applications, 58(6):1198–1206, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004593.

Dai:2008:ESD

[DCW08] Chao-Qing Dai, Xu Cen, and Sheng-Sheng Wu. Exact solutions of discrete complex cubic Ginzburg–Landau equation via extended tanhfunction approach. Computers and Mathematics with Applications, 56(1):55– 62, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008243.

Das:2000:SSC

[DD00] S. Das and L. Debnath.

Study of a static cruciform crack problem in an orthotropic elastic plane. Computers and Mathematics with Applications, 40(4–5):569–575, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001802.

DAcunto:2003:EIE

[DD03] B. D’Acunto and M. De Angelis. Evolution of the interface for an extended melting model. Computers and Mathematics with Applications, 46(5–6):971–976, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390157X.

DeLeo:2004:RLQ

[DD04] S. De Leo and G. C. Ducati. Real linear quaternionic differential operators. Computers and Mathematics with Applications, 48(12):1893–1903, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210400447X.

Ding:2008:RBE

[DD08] Jie Ding and Feng Ding. The residual based extended least squares identification method for dualrate systems. Computers and Mathematics with Applications, 56(6):1479–1487, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001740.

Dragomir:2004:AFR

[DDCB04] N. M. Dragomir, S. S. Dragomir, Y. J. Cho, and G. W. Baxter. Approximations for the fiber refractive index profile. Computers and Mathematics with Applications, 48(3–4):481– 495, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840707.

Dragomir:2002:AFH

[DDF02] N. M. Dragomir, S. S.

Dragomir, and P. Farrell. Approximating the finite Hilbert transform via trapezoid type inequalities. Computers and Mathematics with Applications, 43(10–11):1359–1369, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001049.

Douraki:2008:DDE

[DDM08] M. Jaberi Douraki, Mehdi Dehghan, and Javad Mashreghi. Dynamics of the difference equation \frac{x_{n+1}=x_{n}+px_{n−k}}{x_{n}+q}. Computers and Mathematics with Applications, 56(1): 186–198, July 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108000047.

Duan:2004:UBP

[DDX04] Zhiwen Duan, Weibing Deng, and Chunhong Xie. Uniform blow-up profile for a degenerate parabolic system with nonlocal source. Computers and Mathematics with Applications, 47(6–7):977–995, March/ April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900818.

DeChant:2002:IDQ

[De 02] L. J. De Chant. Impulsive displacement of a quasilinear viscoelastic material through accurate numerical inversion of the Laplace transform. Computers and Mathematics with Applications, 43(8–9):1161–1170, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800207.

DeChant:2004:ASU

[De 04] L. J. De Chant. An analytical solution for unconfined, unsteady, inviscid jets; with applications to penetration problem debris cloud formation. Computers and Mathematics with Applications, 48(1–2):201– 213, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840100.

Douglas:2006:DDD

[DE06a] C. C. Douglas and Y. Efendiev. A dynamic data-driven application simulation framework for contaminant transport problems. Computers and Mathematics with Applications, 51(11):1633– 1646, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001325.

Duman:2006:ACP

[DE06b] O. Duman and E. Erkus.

Approximation of continuous periodic functions via statistical convergence. Computers and Mathematics with Applications, 52 (6–7):967–974, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002707.

Dehghan:2002:NAT

[Deh02]  M. Dehghan.  A new

ADI technique for twodimensional parabolic equation with an integral condition. Computers and Mathematics with Applications, 43(12):1477–1488, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210200113X.

Dementiev:2000:SRM

[Dem00]  O. Dementiev.  Stabil-

ity of a rotor motion in a chamber filled with viscous gas. Computers and Mathematics with Applications, 39(7–8):183–191, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000742.

Deng:2007:EUP

[Den07] Jiqin Deng. Existence for unbounded positive solutions of Schrödinger equations in two-dimensional exterior domains. Computers and Mathematics with Applications, 54(9–10):1276– 1284, November 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107003392.

Deng:2008:LHI

[Den08] Hanyuan Deng. The largest Hosoya index of (n, n + 1)- graphs. Computers and Mathematics with Applications, 56(10):2499–2506, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003866.

Deutsch:2001:MEN

[Deu01] J. I. Deutsch. The modular equations of norm 3 for Q(\sqrt{3}). Computers and Mathematics with Applications, 42(10– 11):1293–1301, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002401.

Dey:2000:D

[Dey00a]  S. K. Dey.  Dedica-

tion. Computers and Mathematics with Applications, 40(12):xv, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002431.

Dey:2000:P

[Dey00b] S. K. Dey. Preface. Computers and Mathematics with Applications, 40(12):xiii, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000242X.

Driscoll:2002:ILI

[DF02] T. A. Driscoll and B. Fornberg. Interpolation in the limit of increasingly flat radial basis functions. Computers and Mathematics with Applications, 43 (3–5):413–422, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002954.

Diethelm:2006:EAE

[DF06] K. Diethelm and A. D. Freed. An efficient algorithm for the evaluation of convolution integrals. Computers and Mathematics with Applications, 51(1):51– 72, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105004438.

DeSalvo:2009:ICH

[DFFL09] Mario De Salvo, Dario Fasino, Domenico Freni, and Giovanni Lo Faro. Isomorphism classes of the hypergroups of type U on the right of size five. Computers and Mathematics with Applications, 58(2):390– 402, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002533.

Domingo-Ferrer:2008:PTA

[DFSS08] Josep Domingo-Ferrer, Francesc Sebé, and Agusti Solanas. A polynomial-time approximation to optimal multivariate microaggregation. Computers and Mathematics with Applications, 55 (4):714–732, February 2008.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107005044.

Duan:2002:LON

[DFY02] Yongrui Duan, Wei Feng, and Jurang Yan. Linearized oscillation of nonlinear impulsive delay differential equations. Computers and Mathematics with Applications, 44(8– 9):1267–1274, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002316.

Dillon:2006:SMM

[DFY06] R. H. Dillon, L. J. Fauci, and Xingzhou Yang. Sperm motility and multiciliary beating: an integrative mechanical model. Computers and Mathematics with Applications, 52(5):749–758, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002537.

Dong:2005:PSP

[DG05] Shijie Dong and Weigao Ge. Positive solutions of an m-point boundary value problem with sign changing nonlinearities. Computers and Mathematics with Applications, 49(4):589–598, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000520.

dHumieres:2009:VIN

[dG09] Dominique d’Humières and Irina Ginzburg. Viscosity independent numerical errors for lattice Boltzmann models: From recurrence equations to “magic” collision numbers. Computers and Mathematics with Applications, 58(5):823–840, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000893.

Dutta:2004:ECT

[DGG^{+}04] S. Dutta, J. Glimm, J. W. Grove, D. H. Sharp, and Yongmin Zhang. Error comparison in tracked and untracked spherical simulations. Computers and Mathematics with Applications, 48(10–11):1733–1747, November/December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004249.

Diaz:2008:LPG

[DGPT08] A. Díaz, M. I. García-Planas, and S. Tarragona. Local perturbations of generalized systems under feedback and derivative feedback. Computers and Mathematics with Applications, 56(4):988–1000, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001016.

Dattoli:2004:MEE

[DGR04] G. Dattoli, B. Germano, and P. E. Ricci. Matrix evolution equations and special functions. Computers and Mathematics with Applications, 48(10– 11):1611–1617, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004146.

Darius:2006:QFA

[DGVP06] L. Darius, P. González-Vera, and M. Jiménez Paiz. Quadrature formulas associated with rational modifications of the Chebyshev weight functions. Computers and Mathematics with Applications, 51(3–4):419–430, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105005092.

Dosly:2003:CSL

[DH03] O. Doslý and R. Hilscher. A class of Sturm–Liouville difference equations: (non)oscillation constants and property BD. Computers and Mathematics with Applications, 45(6–9):961–981, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000701.

Driscoll:2007:ARS

[DH07] Tobin A. Driscoll and Alfa R. H. Heryudono. Adaptive residual subsampling methods for radial basis function interpolation and collocation problems. Computers and Mathematics with Applications, 53(6):927–939, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000454.

Dai:2008:SWS

[DH08a] Hui-Hui Dai and Yi Huo. Solitary waves in a slender tube composed of an incompressible nonlinear elastic material. Computers and Mathematics with Applications, 55(4):620–635, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107004968.

Dehghan:2008:IAS

[DH08b] Mehdi Dehghan and Masoud Hajarian. An iterative algorithm for solving a pair of matrix equations AY B = E, CY D = F over generalized centro-symmetric matrices. Computers and Mathematics with Applications, 56(12):3246–3260, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004318.

Ding:2008:DNA

[DH08c] Wei Ding and Maoan Han. Dynamic of a nonautonomous predator–prey system with infinite delay and diffusion. Computers and Mathematics with Applications, 56(5):1335–1350, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001570.

Dhage:2003:RTF

[Dha03] B. C. Dhage. Remarks on two fixed-point theorems involving the sum and the product of two operators. Computers and Mathematics with Applications, 46(12):1779–1785, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902367.

Dhage:2004:MOF

[Dha04] B. C. Dhage. Multivalued operators and fixedpoint theorems in Banach algebras II. Computers and Mathematics with Applications, 48(10– 11):1461–1476, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004018.

Dhage:2006:FPT

[Dha06] B. C. Dhage. Fixed-point theorems for discontinuous multivalued operators on ordered spaces with applications. Computers and Mathematics with Applications, 51(3–4):589–604, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005249.

Dhage:2007:HFP

[Dha07] B. C. Dhage. Hybrid fixed point theory for strictly monotone increasing multivalued mappings with applications. Computers and Mathematics with Applications, 53(5):803–824, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000338.

Ding:2005:PBV

[DHM05] Wei Ding, Maoan Han, and Junrong Mi. Periodic boundary value problem for the second-order impulsive functional differential equations. Computers and Mathematics with Applications, 50(3–4):491– 507, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002919.

Dutta:2006:TCM

[DHY06] P. Dutta, Keisuke Horiuchi, and Hong-Ming Yin. Thermal characteristics of mixed electroosmotic and pressuredriven microflows. Computers and Mathematics with Applications, 52(5):651–670, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002434.

Darania:2008:NSN

[DI08] P. Darania and K. Ivaz. Numerical solution of nonlinear Volterra–Fredholm integro-differential equations. Computers and Mathematics with Applications, 56(9):2197–2209, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003544.

Diblik:2001:DRP

[Dib01]  J. Diblík.  Discrete re-

tract principle for systems of discrete equations. Computers and Mathematics with Applications, 42(3–5):515–528, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001742.

Diblik:2003:ABS

[Dib03] J. Diblík. Asymptotic behaviour of solutions of systems of discrete equations via Liapunov type technique. Computers and Mathematics with Applications, 45(6–9):1041–1057, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000828.

Dimitriu:2008:NCS

[Dim08]  G. Dimitriu.  A nu-

merical comparative study on data assimilation using Kalman filters. Computers and Mathematics with Applications, 55(10):2247– 2265, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007390.

Ding:2000:CMG

[Din00a] Xie Ping Ding. Constrained multiobjective games in general topological space. Computers and Mathematics with Applications, 39(3–4): 23–30, February 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122199003302.

Ding:2000:EPE

[Din00b]  Xie Ping Ding.  Exis-

tence of Pareto equilibria for constrained multiobjective games in H-space. Computers and Mathematics with Applications, 39(9–10):125– 134, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000924.

Ding:2000:ESQ

[Din00c]  Xie Ping Ding.  Ex-

istence of solutions for quasi-equilibrium problems in noncompact topological spaces. Computers and Mathematics with Applications, 39(3–4):13–21, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003296.

Ding:2002:GKT

[Din02] Xie Ping Ding. Generalized L-KKM type theorems in L-convex spaces with applications. Computers and Mathematics with Applications, 43(10–11):1249–1256, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102000962.

Ding:2004:LSU

[Din04a] Changming Ding. The limit sets of uniformly asymptotically zhukovskij stable orbits. Computers and Mathematics with Applications, 47(6–7):859–862, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900703.

Ding:2004:ADQ

[Din04b]  Shusen Ding.  L^{ϕ}(μ)-

averaging domains and the quasi-hyperbolic metric. Computers and Mathematics with Applications, 47(10–11):1611–1618, May/ June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839749.

Dix:2005:ABS

[Dix05] J. G. Dix. Asymptotic behavior of solutions to a firstorder differential equation with variable delays. Computers and Mathematics with Applications, 50(10– 12):1791–1800, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004244.

Debnath:2008:BOC

[DJ08] L. Debnath and J. C. Jiang.

Bounded oscillation criteria for second-order nonlinear delay difference equations of unstable type. Computers and Mathematics with Applications, 56(7):1797–1807, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002605.

Defez:2004:JMD

[DJL04] E. Defez, L. Jódar, and A. Law. Jacobi matrix differential equation, polynomial solutions, and their properties. Computers and Mathematics with Applications, 48(5–6):789–803, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210400313X.

Dosla:2004:GAP

[DK04] Z. Doslá and A. Kobza. Global asymptotic properties of third-order difference equations. Computers and Mathematics with Applications, 48(1–2):191– 200, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840094.

DelMedico:2005:NKT

[DK05] A. Del Medico and Qingkai Kong. New Kamenevtype oscillation criteria for second-order differential equations on a measure chain. Computers and Mathematics with Applications, 50(8–9):1211–1230, October/November 2005.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105003755.

Darvishi:2009:SSF

[DK09] M. T. Darvishi and F. Khani.

A series solution of the foam drainage equation. Computers and Mathematics with Applications, 58(2):360– 368, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900282X.

dHumieres:2009:DPL

[dKLR09] Dominique d’Humières, Manfred Krafczyk, Li-Shi Luo, and Robert Rubinstein. Dedication to Pierre Lallemand on the occasion of his retirement. Computers and Mathematics with Applications, 58(5):821–822, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000881.

Dohi:2003:NGM

[DKO03] T. Dohi, N. Kaio, and S. Osaki. A new graphical method to estimate the optimal repair-time limit with incomplete repair and discounting. Computers and Mathematics with Applications, 46(7):999–1007, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901143.

Darvishi:2007:NSS

[DKS07] M. T. Darvishi, F. Khani, and A. A. Soliman. The numerical simulation for stiff systems of ordinary differential equations. Computers and Mathematics with Applications, 54(7–8):1055– 1063, October 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210700274X.

Ding:2005:BSG

[DL05] Juntang Ding and Shengjia Li. Blow-up solutions and global solutions for a class of quasilinear parabolic equations with robin boundary conditions. Computers and Mathematics with Applications, 49(5–6):689–701, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000830.

Dshalalow:2006:LCO

[DL06] J. H. Dshalalow and A. Liew. Level crossings of an oscillating marked random walk. Computers and Mathematics with Applications, 52 (6–7):917–932, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002665.

Du:2008:GCC

[DL08] Chaoxiong Du and Yirong Liu. General center conditions and bifurcation of limit cycles for a quasisymmetric seventh degree system. Computers and Mathematics with Applications, 56(11):2957–2969, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004537.

DeBarros:2003:SEM

[DLB03] L. C. De Barros, M. B. Ferreira Leite, and R. C. Bassanezi. The SI epidemiological models with a fuzzy transmission parameter. Computers and Mathematics with Applications, 45(10–11):1619–1628, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210300141X.

Davvaz:2009:AIV

[DLF09] B. Davvaz and V. Leoreanu-Fotea. Applications of interval valued fuzzy n-ary polygroups with respect to tnorms (t-conorms). Computers and Mathematics with Applications, 57(8): 1413–1424, April 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109000340.

Dai:2008:LTO

[DLL08] Tian-Shyr Dai, Li-Min Liu, and Yuh-Dauh Lyuu. Linear-time option pricing algorithms by combinatorics. Computers and Mathematics with Applications, 55(9):2142–2157, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006931.

Du:2008:CND

[DLM08] Chaoxiong Du, Yirong Liu, and Heilong Mi. A class of ninth degree system with four isochronous centers. Computers and Mathematics with Applications, 56(10):2609–2620, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003982.

Dohi:2006:P

[DLO06] Tadashi Dohi, Nikolaos Limnios, and Shunji Osaki. Preface. Computers and Mathematics with Applications, 51(2):xiii, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004529.

Das:2008:RGT

[DLSB08] N. C. Das, A. Lahiri, S. Sarkar, and S. Basu. Reflection of generalized thermoelastic waves from isothermal and insulated boundaries of a half space. Computers and Mathematics with Applications, 56 (11):2795–2805, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004525.

deMalafosse:2006:AIM

[dM06] Bruno de Malafosse. An application of the infinite matrix theory to Mathieu equation. Computers and Mathematics with Applications, 52(10– 11):1439–1452, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003312.

Dosly:2007:EOL

[DM07] Ondrej Doslý and Daniel Marek. Even-order linear dynamic equations with mixed derivatives. Computers and Mathematics with Applications, 53(7):1140– 1152, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000739.

Djebali:2008:MPS

[DM08] Smaïl Djebali and Karima Mebarki. Multiple positive solutions for singular BVPs on the positive half-line. Computers and Mathematics with Applications, 55(12):2940– 2952, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107008164.

Damavandi:2009:FTC

[DM09] Yasser Baleghi Damavandi and Karim Mohammadi. Fault tolerance in coevolutionary communication of EHW modules. Computers and Mathematics with Applications, 57(11– 12):1730–1735, June 2009.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108005579.

Dedic:2001:SIE

[DMP01] Lj. Dedić, M. Matić, and J. Pecarić. Some inequalities of Euler–Grüss type. Computers and Mathematics with Applications, 41(7–8):843–856, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003242.

Dimas:2003:LES

[DMP03] A. A. Dimas, B. M. Mowli, and U. Piomelli. Largeeddy simulation of subcritical transition in an attachment-line boundary layer. Computers and Mathematics with Applications, 46(4):571–589, August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900174.

Demkowicz:2000:MED

[DMSV00] L. Demkowicz, P. Monk, Ch. Schwab, and L. Vardapetyan. Maxwell eigenvalues and discrete compactness in two dimensions. Computers and Mathematics with Applications, 40(4–5):589–605, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001826.

Demkowicz:2000:RDF

[DMVR00] L. Demkowicz, P. Monk, L. Vardapetyan, and W. Rachowicz. De Rham diagram for hp finite element spaces. Computers and Mathematics with Applications, 39(7–8):29– 38, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000626.

Diaz:2003:BSN

[DN03] L. Díaz and R. Naulin. Bounded solutions of nonlinear difference equations with advanced arguments. Computers and Mathematics with Applications, 45 (6–9):1021–1032, March/ May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000804.

Dogbe:2001:RTE

[Dog01] C. Dogbe. The radiative transfer equations: diffusion approximation under accretiveness and compactness assumptions. Computers and Mathematics with Applications, 42(6– 7):783–791, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001985.

Dogbe:2005:ADI

[Dog05] C. Dogbe. Approximation of a diffusion induced on a kinetic equation by a dynamical system. Computers and Mathematics with Applications, 49(9–10):1303– 1325, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001513.

Dogbe:2008:NSS

[Dog08] C. Dogbé. On the numerical solutions of second order macroscopic models of pedestrian flows. Computers and Mathematics with Applications, 56(7):1884–1898, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800268X.

Dong:2004:RDG

[Don04]  Shi-Hai Dong.  Realiza-

tion of the dynamical group for the generalized Laguerre functions. Computers and Mathematics with Applications, 47(6–7):1035–1039, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900843.

Dong:2009:OSF

[Don09] Jiu-Gang Dong. Oscillation of solutions for first order neutral differential equations with distributed deviating arguments. Computers and Mathematics with Applications, 58(4):784–790, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003708.

Dorfmann:2000:SWW

[Dor00] A. A. Dorfmann. Scattering of water waves by a submerged torus. Computers and Mathematics with Applications, 40(6– 7):813–833, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001991.

Dorovsky:2001:OND

[Dor01] V. N. Dorovsky. One nonlinear differential equation and operation of hydraulic dissipative pendulum. Computers and Mathematics with Applications, 41(5–6):697– 702, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000314X.

deOliveira:2008:MOI

[dORM08] Valeriano Antunes de Oliveira and Marko Antonio Rojas-Medar. Multi-objective infinite programming. Computers and Mathematics with Applications, 55(9):1907– 1922, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006591.

Ding:2002:CST

[DP02] Xie Ping Ding and J. Y.

Park. Continuous selection theorem, coincidence theorem, and generalized equilibrium in Lconvex spaces. Computers and Mathematics with Applications, 44(1–2):95– 103, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001323.

Dosla:2007:CPS

[DP07] Zuzana Doslá and Sárka Pechancová. Conjugacy and phases for second order linear difference equation. Computers and Mathematics with Applications, 53(7):1129–1139, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000727.

Dhompongsa:2008:CTC

[DP08] S. Dhompongsa and B. Panyanak. On △-convergence theorems in CAT(0) spaces. Computers and Mathematics with Applications, 56 (10):2572–2579, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108003945.

Ding:2003:PEC

[DPJ03] Xie Ping Ding, Jong Yeoul Park, and Il Hyo Jung. Pareto equilibria for constrained multiobjective games in locally L-convex spaces. Computers and Mathematics with Applications, 46 (10–11):1589–1599, November/December 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0898122103901945.

DAcunto:2007:NMF

[DPU07] B. D’Acunto, F. Parente, and G. Urciuoli. Numerical models for 2D free boundary analysis of groundwater in slopes stabilized by drain trenches. Computers and Mathematics with Applications, 53(10):1615– 1626, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001356.

Deakin:2001:NBC

[DR01a] A. S. Deakin and H. Rasmussen. Nonreflecting boundary condition for the Helmholtz equation. Computers and Mathematics with Applications, 41(3–4): 307–318, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100002753.

Dosly:2001:NCH

[DR01b] O. Došlý and P. Rehák.

Nonoscillation criteria for half-linear second-order difference equations. Computers and Mathematics with Applications, 42(3–5):453– 464, August/September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122101001699.

Dodu:2002:VIU

[DR02] F. Dodu and C. Rabut. Vectorial interpolation using radial-basis-like functions. Computers and Mathematics with Applications, 43(3–5):393–411, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002942.

deBruynKops:2003:LES

[dR03] S. M. deBruynKops and J. J. Riley. Large-eddy simulation of a reacting scalar mixing layer with Arrhenius chemistry. Computers and Mathematics with Applications, 46(4):547–569, August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900162.

Diaz:2004:NED

[DR04] J. I. Díaz and Á. M. Ramos. Numerical experimentsregarding the distributed control of semilinear parabolic problems. Computers and Mathematics with Applications, 48(10–11):1575–1586, November/December 2004.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004110.

Dragomir:2003:ILM

[Dra03a] S. S. Dragomir. An inequality for logarithmic mapping and applications for the Shannon entropy. Computers and Mathematics with Applications, 46 (8–9):1273–1279, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902185.

Drakopoulos:2003:TJS

[Dra03b] V. Drakopoulos. Are there any Julia sets for the Laguerre iteration function? Computers and Mathematics with Applications, 46 (8–9):1201–1210, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902124.

Dragut:2006:MRO

[Dra06]  A. B. Dragut.  Mono-

tonic robust optimal control policies for the timequality trade-offs in concurrent new product development (NPD). Computers and Mathematics with Applications, 51(2):229–238, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0898122105004608.

Dragomir:2008:ARF

[Dra08] Sever S. Dragomir. Approximating real functions which possess n-th derivatives of bounded variation and applications. Computers and Mathematics with Applications, 56(9):2268–2278, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800360X.

DeAngelis:2009:TDM

[DRR09] E. De Angelis, R. Revelli, and L. Ridolfi. Transportdiffusion models with nonlinear boundary conditions and solution by generalized collocation methods. Computers and Mathematics with Applications, 58(3):558–565, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002375.

DeChant:2002:MSS

[DS02] L. J. De Chant and J. A. Seidel. Multiple stream superposition of a two-stream combined analytical numerical initial value problem solver for aerodynamic mixing. Computers and Mathematics with Applications, 43(10–11):1451–1468, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001116.

Dehghan:2004:BSS

[DS04] M. Dehghan and A. Saadatmandi. Bounds for solutions of a six-point partialdifference scheme. Computers and Mathematics with Applications, 47(1):83–89, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900077.

DeLillo:2005:ESP

[DS05a] S. De Lillo and G. Sanchini. The eckhaus soliton propagation in a random medium. Computers and Mathematics with Applications, 50(1–2):255– 261, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002476.

Ding:2005:UAZ

[DS05b] Changming Ding and J. M. Soriano. Uniformly asymptotically zhukovskij stable orbits. Computers and Mathematics with Applications, 49(1):81–84, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000088.

Dehghan:2006:TMO

[DS06a] M. Dehghan and A. Saadatmandi. A Tau method for the one-dimensional parabolic inverse problem subject to temperature overspecification. Computers and Mathematics with Applications, 52(6– 7):933–940, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002677.

Dragomir:2006:IMF

[DS06b] S. S. Dragomir and A. Sofo. An inequality for monotonic functions generalizing Ostrowski and related results. Computers and Mathematics with Applications, 51(3–4):497–506, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500516X.

Dehghan:2007:NMT

[DS07] Mehdi Dehghan and Ali Shokri. A numerical method for two-dimensional Schrödinger equation using collocation and radial basis functions. Computers and Mathematics with Applications, 54(1):136– 146, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002027.

Du:2008:LBM

[DS08] Rui Du and Baochang Shi. The lattice Boltzmann method for the thermocapillary flow in a cavity under microgravity condition. Computers and Mathematics with Applications, 55(7):1433–1440, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006347.

Dudek:2009:FIH

[DSA09] W. A. Dudek, M. Shabir, and M. Irfan Ali. (α, β)- fuzzy ideals of hemirings. Computers and Mathematics with Applications, 58(2): 310–321, July 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109002636. See

note [Jun10].

Defez:2005:NSM

[DSHS05] E. Defez, L. Soler, A. Hervás, and C. Santamaría. Numerical solution of matrix differential models using cubic matrix splines. Computers and Mathematics with Applications, 50(5–6):693–699, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003093.

deSocio:2002:RFI

[dSM02] L. M. de Socio and L. Marino. Reference frame influence on transport phenomena in gases. a direct simulations approach. Computers and Mathematics with Applications, 44(8– 9):1201–1206, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002262.

dosSantos:2004:PFW

[dSOGRMRL04] L. Batista dos Santos, R. Osuna-Gómez, M. A. Rojas-Medar, and A. Rufián-Lizana. Preinvex functions and weak efficient solutions for some vectorial optimization problem in Banach spaces. Computers and Mathematics with Applications, 48(5–6):885–895, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003219.

dosSantos:2008:SRB

[dSRGRMRL08] Lucelina Batista dos Santos, Gabriel Ruiz-Garzón, Marko A. Rojas-Medar, and Antonio Rufián-Lizana. Some relations between variational-like inequality problems and vectorial optimization problems in Banach spaces. Computers and Mathematics with Applications, 55(8):1808– 1814, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006530.

Dragomir:2000:SUB

[DSS00] S. S. Dragomir, M. L.

Scholz, and J. Sunde. Some upper bounds for relative entropy and applications. Computers and Mathematics with Applications, 39(9–10):91–100, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000894.

Diblik:2008:ECE

[DSS08] Josef Diblík, Zdenek Svoboda, and Zdenek Smarda. Explicit criteria for the existence of positive solutions for a scalar differential equation with variable delay in the critical case. Computers and Mathematics with Applications, 56(2):556– 564, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000849.

Dong:2005:ASO

[DTXW05] Ying Dong, Jiafu Tang, Baodong Xu, and Dingwei Wang. An application of swarm optimization to nonlinear programming. Computers and Mathematics with Applications, 49(11–12):1655– 1668, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001872.

Duan:2003:GOD

[DTZF03] Yongrui Duan, Peng Tian, Weiping Zhang, and Wei Feng. Generic oscillations of delay differential system with impulses.

Computers and Mathematics with Applications, 45 (4–5):615–622, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210300021X.

Du:2008:SFD

[Du08]  Zengji Du.  Solvability

of functional differential equations with multi-point boundary value problems at resonance. Computers and Mathematics with Applications, 55(11):2653– 2661, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700764X.

Duan:2008:NMM

[Dua08] Yong Duan. A note on the meshless method using radial basis functions. Computers and Mathematics with Applications, 55(1):66– 75, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700332X.

Duan:2009:ASS

[Dua09] Zhiwen Duan. The asymptotic self-similar behavior for the quasilinear heat equation with nonlinear boundary condition. Computers and Mathematics with Applications, 58(10):2005–2021, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109005859.

Dubois:2008:EPD

[Dub08] François Dubois. Equivalent partial differential equations of a lattice Boltzmann scheme. Computers and Mathematics with Applications, 55(7):1441– 1449, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006268.

Ding:2002:REV

[DW02a] Jiu Ding and Yimin Wei. Relative errors versus residuals of approximate solutions of weighted least squares problems in Hilbert space. Computers and Mathematics with Applications, 44(3–4):407–411, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210200158X.

Du:2002:NSM

[DW02b] Qiang Du and Tak-Win Wong. Numerical studies of MacQueen’s k-means algorithm for computing the centroidal Voronoi tessellations. Computers and Mathematics with Applications, 44(3–4):511–523, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001657.

Du:2005:OCV

[DW05] Qiang Du and Desheng Wang. The optimal centroidal Voronoi tessellations and the Gersho’s conjecture in the three-dimensional space. Computers and Mathematics with Applications, 49(9–10):1355–1373, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001550.

Dasgupta:2008:AAF

[DW08a] G. Dasgupta and E. Wachspress. The adjoint for an algebraic finite element. Computers and Mathematics with Applications, 55(9):1988–1997, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006669.

Dasgupta:2008:BFC

[DW08b] Gautam Dasgupta and Eugene L. Wachspress. Basis functions for concave polygons. Computers and Mathematics with Applications, 56(2):459– 468, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000308.

Duan:2009:NDC

[DW09] Shukai Duan and Lidan Wang. A novel delayed chaotic neural model and its circuitry implementation. Computers and Mathematics with Applications, 57(11–12):1736– 1742, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005555.

Deng:2007:AOK

[DWL07] Weihua Deng, Yujiang Wu, and Changpin Li. Attractors for one kind of lattice dynamical system. Computers and Mathematics with Applications, 54(5):617–626, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001642.

Deng:2008:PCM

[DWL08] Xiaoyan Deng, Bangju Wang, and Guangqing Long. The Picard contraction mapping method for the parameter inversion of reaction-diffusion systems. Computers and Mathematics with Applications, 56(9):2347–2355, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003738.

Duan:2005:BSD

[DXZ05] Zhi-Wen Duan, Chun-Hong Xie, and Li Zhou. Blow-up solution for the degenerate quasilinear parabolic system with nonlinear boundary condition. Computers and Mathematics with Applications, 50(5–6):895–917, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500324X.

Ding:2005:EAS

[DY05] Xie Ping Ding and Jen-Chih Yao. Existence and algorithm of solutions for mixed quasi-variational-like inclusions in Banach spaces. Computers and Mathematics with Applications, 49(5–6):857–869, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000969.

Ding:2009:TID

[DY09] Junfeng Ding and Stephen S. T. Yau. TCOM, an innovative data structure for mining association rules among infrequent items. Computers and Mathematics with Applications, 57(2):290–301, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006470.

Ding:2008:EAS

[DYZ08] Xie Ping Ding, Jen-Chih Yao, and Lu-Chuan Zeng. Existence and algorithm of solutions for generalized strongly nonlinear mixed variational-like inequalities in Banach spaces. Computers and Mathematics with Applications, 55(4):669–679, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005007.

Du:2005:MMH

[DZ05] Zhengdong Du and Weinian Zhang. Melnikov method for homoclinic bifurcation in nonlinear impact oscillators. Computers and Mathematics with Applications, 50(3–4):445–458, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002889.

Ding:2008:SSA

[DZ08] Deng Ding and Ying Ying Zhang. A splitting-step algorithm for reflected stochastic differential equations in R_{+}^{1}. Computers and Mathematics with Applications, 55(11):2413–2425, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006955.

Zhu:2009:HPM

[dZmClQ09] Shun dong Zhu, Yu ming Chu, and Song liang Qiu. The homotopy perturbation method for discontinued problems arising in nanotechnology. Computers and Mathematics with Applications, 58(11–12):2398–2401, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001813.

Deng:2006:RFN

[DZQ06] Yong Deng, Zhu Zhenfu, and Liu Qi. Ranking fuzzy numbers with an area method using radius of gyration. Computers and Mathematics with Applications, 51(6–7):1127–1136, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000423.

Duan:2006:BPP

[DZZT06] Qi Duan, Huanling Zhang, Yunfeng Zhang, and E. H. Twizell. Bounded property and point control of a bivariate rational interpolating surface. Computers and Mathematics with Applications, 52(6– 7):975–984, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002719.

El-Azab:2003:ELS

[EA03a] M. S. El-Azab. An effective linear scheme to solve nonlinear degenerate parabolic differential equations. Computers and Mathematics with Applications, 46(8– 9):1435–1442, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902318.

Engelstad:2003:DVH

[EA03b] S. P. Engelstad and R. L.

Actis. Development of p-version handbook solutions for analysis of composite bonded joints. Computers and Mathematics with Applications, 46(1):81– 94, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900824.

El-Ashwah:2009:HPC

[EAA09] R. M. El-Ashwah and M. K. Aouf. Hadamard product of certain meromorphic starlike and convex functions. Computers and Mathematics with Applications, 57(7):1102– 1106, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006597.

Elabbasy:2009:ANT

[EAE09] E. M. Elabbasy, H. N. Agiza, and A. A. Elsadany. Analysis of nonlinear triopoly game with heterogeneous players. Computers and Mathematics with Applications, 57(3):488–499, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006469.

Esteban-Bravo:2006:VBL

[EBVS06] M. Esteban-Bravo and J. M. Vidal-Sanz. Valuation of boundary-linked assets by stochastic boundary value problems solved with a wavelet-collocation algorithm. Computers and Mathematics with Applications, 52(1–2):137–160, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001878.

Eslahchi:2009:QRU

[ED09] M. R. Eslahchi and Mehdi Dehghan. Quadrature rules using an arbitrary fixed order of derivatives. Computers and Mathematics with Applications, 57(7):1212– 1225, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000285.

Ehrgott:2002:MMF

[EG02] M. Ehrgott and E. A. Galperin. Min-max formulation of the balance number in multiobjective global optimization. Computers and Mathematics with Applications, 44(7):899–907, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210200202X.

El-Gebeily:2003:ISM

[EGA03] M. A. El-Gebeily and B. S.

Attili. An iterative shooting method for a certain class of singular twopoint boundary value problems. Computers and Mathematics with Applications, 45(1–3):69–76, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800081.

Eklund:2002:SFF

[EGM^{+}02] P. Eklund, M. A. Galán, J. Medina, M. Ojeda-Aciego, and A. Valverde. Set functors, L-fuzzy set categories, and generalized terms. Computers and Mathematics with Applications, 43(6–7):693–705, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003145.

Elhay:2003:SML

[EGR03] S. Elhay, G. H. Golub, and Y. M. Ram. The spectrum of a modified linear pencil. Computers and Mathematics with Applications, 46 (8–9):1413–1426, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210390229X.

El-Gamel:2004:SGM

[EGZ04] M. El-Gamel and A. I. Zayed. Sinc–Galerkin method for solving nonlinear boundary-value problems. Computers and Mathematics with Applications, 48(9):1285–1298, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210400375X.

Ehrgott:2002:BSA

[Ehr02] M. Ehrgott. The balance space approach to multicriteria decision makinginvolving the decision maker. Computers and Mathematics with Applications, 44 (7):909–923, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122102002031.

Eloe:2003:UAS

[EIR03] P. W. Eloe, M. N. Islam, and Y. N. Raffoul. Uniform asymptotic stability in nonlinear Volterra discrete systems. Computers and Mathematics with Applications, 45(6–9):1033–1039, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000816.

Ekel:2002:FSM

[Eke02] P. Ya. Ekel. Fuzzy sets and models of decision making. Computers and Mathematics with Applications, 44(7):863–875, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001992.

ElSheikh:2003:EER

[El 03] M. G. El Sheikh. The estimation of the error resulting due to the truncation applied to homogeneous integral equations with Cauchy’s kernel. Computers and Mathematics with Applications, 46(10– 11):1741–1748, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902082.

Elkahlout:2006:BEE

[Elk06] G. R. Elkahlout. Bayes estimators for the extremevalue reliability function. Computers and Mathematics with Applications, 51(3– 4):673–679, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122105005316.

Elyseeva:2004:TSS

[Ely04] J. Elyseeva. A transformation for symplectic systems and the definition of a focal point. Computers and Mathematics with Applications, 47(1):123–134, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900119.

El-Morshedy:2003:GAD

[EM03] H. A. El-Morshedy. The global attractivity of difference equations of nonincreasing nonlinearities with applications. Computers and Mathematics with Applications, 45(4–5):749–758, February/March 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210300035X.

El-Metwally:2009:QPS

[EM09a]  H. El-Metwally.  Quali-

tative properties of some higher order difference equations. Computers and Mathematics with Applications, 58(4):686–692, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002557.

El-Mistikawy:2009:CTD

[EM09b] Tarek M. A. El-Mistikawy.

Comment on: “The threedimensional flow past a stretching sheet and the homotopy perturbation method”, by P. D. Ariel, Computers and Mathematics with Applications 54 (2007) 920–925. Computers and Mathematics with Applications, 57(3):404–406, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005609.

See [Ari07b] and reply [Ari09c].

El-Morshedy:2009:NOC

[EM09c] Hassan A. El-Morshedy.

New oscillation criteria for second order linear difference equations with positive and negative coefficients. Computers and Mathematics with Applications, 58(10):1988–1997, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006014.

Encinas:2008:SPC

[EMD08] L. Hernández Encinas, J. Muñoz Masqué, and A. Queiruga Dios. Safer parameters for the Chor– Rivest cryptosystem. Computers and Mathematics with Applications, 56(11): 2883–2886, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004835.

ElAmrouss:2009:EMN

[EMM09] A. R. El Amrouss, F. Moradi, and M. Moussaoui. Existence of multiple nontrivial solutions for semilinear elliptic problems. Computers and Mathematics with Applications, 57(7):1115– 1126, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000467.

Ekel:2006:FSB

[EMP^{+}06] P. Ya. Ekel, C. A. P. S.

Martins, J. G. Pereira, Jr., R. M. Palhares, and L. N. Canha. Fuzzy set based multiobjective allocation of resources and its applications. Computers and Mathematics with Applications, 52(1–2):197– 210, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001908.

Engelberg:2004:EOP

[Eng04] S. Engelberg. Errata to “The operation of the phaselock loop”. Computers and Mathematics with Applications, 47(6–7):1141–1142, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900946.

See [EW00a].

Eastham:2004:FEM

[EP04] J. F. Eastham and J. S.

Peterson. The finite element method in anisotropic Sobolev spaces. Computers and Mathematics with Applications, 47(10–11):1775– 1786, May/June 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104839907.

Erbe:2005:SIG

[EPS05] L. Erbe, A. Peterson, and M. Simon. Square integrability of Gaussian bells on time scales. Computers and Mathematics with Applications, 49(5–6):871–883, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000970.

Ettwein:2007:ELS

[ER07] F. Ettwein and M. Ruzicka. Existence of local strong solutions for motions of electrorheological fluids in three dimensions. Computers and Mathematics with Applications, 53(3–4):595–604, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001563.

Eloe:2001:PSN

[ERRY01] P. W. Eloe, Y. Raffoul, D. T. Reid, and K. C. Yin. Positive solutions of nonlinear functional difference equations. Computers and Mathematics with Applications, 42(3–5):639–646, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001833.

El-Sayed:2001:TIP

[ES01] S. M. El-Sayed. Two iteration processes for computing positive definite solutions of the equation X − A^{∗}X^{−n}A = Q. Computers and Mathematics with Applications, 41(5–6):579– 588, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003011.

Ervin:2005:NAN

[ES05] V. J. Ervin and J. J.

Shepherd. Numerical approximation of the Newtonian film blowing problem. Computers and Mathematics with Applications, 49(11–12):1687– 1707, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001902.

Ell:2007:QIA

[ES07] Todd A. Ell and Stephen J.

Sangwine. Quaternion involutions and anti-involutions. Computers and Mathematics with Applications, 53 (1):137–143, January 2007.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107001836.

El-Saify:2008:OCD

[ES08a] H. A. El-Saify. Optimality conditions for differential system of Petrowsky type with infinite number of variables and boundary control. Computers and Mathematics with Applications, 55(4):776–787, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005081.

Erdik:2008:CAA

[ES08b] Tarkan Erdik and Zekai Sen. Comment on: “Application of an adaptive neural fuzzy inference system to thermal comfort and group technology problems” [Comput. Math. Appl. 54 (2007) 1395–1402]. Computers and Mathematics with Applications, 56(7):1901–1902, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003131.

See [NLL07, NLL08].

Ekel:2006:FPM

[ESNP06] P. Ya. Ekel, M. R. Silva, F. Schuffner Neto, and R. M. Palhares. Fuzzy preference modeling and its application to multiobjective decision making. Computers and Mathematics with Applications, 52(1–2):179– 196, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001891.

El-Sayed:2006:NSQ

[ESR06] W. G. El-Sayed and B. Rzepka.

Nondecreasing solutions of a quadratic integral equation of Urysohn type. Computers and Mathematics with Applications, 51 (6–7):1065–1074, March/ April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210600037X.

El-Taha:2009:DPM

[ET09] Muhammad El-Taha. Departure process in a mixed fork-join synchronization network. Computers and Mathematics with Applications, 57(8):1272–1279, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000388.

El-Tawil:2009:ASM

[ETAJ09] M. A. El-Tawil and Amna S. Al-Johani. Approximate solution of a mixed nonlinear stochastic oscillator. Computers and Mathematics with Applications, 58(11–12):2236–2259, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900159X.

El-Taha:2005:QNM

[ETH05] M. El-Taha and J. R. Heath. Queueing network models of credit-based flow control. Computers and Mathematics with Applications, 50(3–4):393–398, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210500283X.

Engelberg:2000:OPL

[EW00a] S. Engelberg and M. Wexler. The operation of the phaselock loop. Computers and Mathematics with Applications, 40(10–11):1253–1261, November/December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100002364. See

errata [Eng04].

Estes:2000:AAB

[EW00b] D. R. Estes and C. Waid.

An algebraic analysis of bore hole samples. Computers and Mathematics with Applications, 39(11): 25–30, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001036.

Elek:2008:TBE

[EZ08] Péter Elek and András Zempléni. Tail behaviour and extremes of two-state Markov-switching autoregressive models. Computers and Mathematics with Applications, 55(12):2839– 2855, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007821.

Fetecau:2009:UFG

[FAF09] Corina Fetecau, M. Athar, and C. Fetecau. Unsteady flow of a generalized Maxwell fluid with fractional derivative due to a constantly accelerating plate. Computers and Mathematics with Applications, 57(4):596–603, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006603.

Fan:2002:TWS

[Fan02a]  En-Gui Fan.  Traveling

wave solutions for nonlinear equations using symbolic computation. Computers and Mathematics with Applications, 43(6–7):671–680, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101003121.

Fang:2002:UVS

[Fan02b] Ming-Liang Fang. Uniqueness and value-sharing of entire functions. Computers and Mathematics with Applications, 44(5–6):823–831, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001943.

Fan:2009:HFA

[Fan09] Jie Fan. He’s frequencyamplitude formulation for the Duffing harmonic oscillator. Computers and Mathematics with Applications, 58(11–12):2473–2476, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001849.

Farkas:2001:UMR

[Far01]  G. Farkas.  Unstable

manifolds for RFDEs under discretization: the Euler method. Computers and Mathematics with Applications, 42(8– 9):1069–1081, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210100222X.

Farkas:2002:CDF

[Far02] G. Farkas. Conjugacy in the discretized fold bifurcation. Computers and Mathematics with Applications, 43(8–9):1027–1033, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800116.

Fasshauer:2002:NIM

[Fas02] G. E. Fasshauer. Newton iteration with multiquadrics for the solution of nonlinear PDEs. Computers and Mathematics with Applications, 43 (3–5):423–438, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002966.

Fasshauer:2006:P

[Fas06] Greg Fasshauer. Preface. Computers and Mathematics with Applications, 51(8): xiii–xv, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210600071X.

Fan:2005:BLR

[FB05] Ying Wai Fan and K. S. Berenhaut. A bound for linear recurrence relations with unbounded order. Computers and Mathematics with Applications, 50(3–4):509– 518, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002920.

Furkan:2007:FSO

[FBA07] Hasan Furkan, Hüseyin Bilgiç, and Bilâl Altay. On the fine spectrum of the operator B(r, s, t) over C_{0} and C. Computers and Mathematics with Applications, 53(6):989–998, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000533.

Fornberg:2002:OBR

[FDWC02] B. Fornberg, T. A. Driscoll, G. Wright, and R. Charles.

Observations on the behavior of radial basis function approximations near boundaries. Computers and Mathematics with Applications, 43(3–5):473–490, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002991.

Ding:2011:NIF

[fDxZ11] Heng fei Ding and Yu xin Zhang. Notes on “Implicit finite difference approximation for time fractional diffusion equations” [Comput. Math. Appl. 56 (2008) 1138–1145]. Computers and Mathematics with Applications, 61(9):2924– 2928, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122111001520.

See [Mur08a].

Ficarra:2008:JCC

[FDY^{+}08] Elisa Ficarra, Giovanni De Micheli, Sungroh Yoon, Luca Benini, and Enrico Macii. Joint coclustering: Co-clustering of genomic and clinical bioimaging data. Computers and Mathematics with Applications, 55(5):938–949, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005652.

Feireisl:2007:MTC

[Fei07] Eduard Feireisl. Mathematical theory of compressible, viscous, and heat conducting fluids. Computers and Mathematics with Applications, 53(3–4):461–490, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001848.

Feng:2008:EUR

[Fen08]  Yuqiang Feng.  Exis-

tence and uniqueness results for a third-order implicit differential equation. Computers and Mathematics with Applications, 56(10):2507–2514, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003878.

Ferrari:2001:UCI

[Fer01]  L. Ferrari.  An umbral

calculus over infinite coefficient fields of positive characteristic. Computers and Mathematics with Applications, 41(9):1099– 1108, May 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101000840.

Ferziger:2003:SEF

[Fer03]  J. H. Ferziger.  Sim-

ulation of environmental flows: Similarities and differences from technological flows. Computers and Mathematics with Applications, 46(4):591–602, August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900186.

Feinsilver:2001:AGG

[FF01] Ph. Feinsilver and U. Franz.

The affine group and generalized Gegenbauer polynomials. Computers and Mathematics with Applications, 41(9):1173–1182, May 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210100089X.

Fang:2002:UMF

[FF02] Cai-Yun Fang and Ming-Liang Fang. Uniqueness of meromorphic functions and differential polynomials. Computers and Mathematics with Applications, 44(5–6):607–617, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210200175X.

Fusi:2007:MBL

[FF07] Lorenzo Fusi and Angiolo Farina. Modelling of Bingham-like fluids with deformable core. Computers and Mathematics with Applications, 53(3–4):583–594, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001617.

Fedoseyev:2002:IMM

[FFK02] A. I. Fedoseyev, M. J. Friedman, and E. J. Kansa. Improved multiquadric method for elliptic partial differential equations via PDE collocation on the boundary. Computers and Mathematics with Applications, 43(3–5):439–455, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002978.

Fu:2005:TRM

[FFXL05] Peng Fu, Chu-Li Fu, Xiang-Tuan Xiong, and Hong-Fang Li. Two regularization methods and the order optimal error estimates for a sideways parabolic equation. Computers and Mathematics with Applications, 49(5–6):777–788, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500091X.

Frontini:2000:VSR

[FG00] M. Frontini and L. Gotusso. The vibrating string and rod free to slide, at both end points, on a vertical guide. Computers and Mathematics with Applications, 40(2–3):277–290, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001607.

Farago:2008:AIO

[FGH08] István Faragó, Boglárka Gnandt, and Ágnes Havasi. Additive and iterative operator splitting methods and their numerical investigation. Computers and Mathematics with Applications, 55(10):2266–2279, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007389.

Fang:2005:DEU

[FGX05] Ming Fang, R. P. Gilbert, and Xuming Xie. Deriving the effective ultrasound equations for soft tissue interrogation. Computers and Mathematics with Applications, 49(7–8):1069–1080, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001239.

Freiling:2003:PSR

[FH03] G. Freiling and A. Hochhaus. Properties of the solutions of rational matrix difference equations. Computers and Mathematics with Applications, 45(6–9):1137–1154, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000737.

Fang:2007:PWP

[FH07] Ya-Ping Fang and Rong Hu. Parametric wellposedness for variational inequalities defined by bifunctions. Computers and Mathematics with Applications, 53(8):1306–1316, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000909.

He:2009:GPM

[fHCH09] Xin feng He, Junmin Chen, and Zhen He. Generalized projection method for a variational inequality system with different mapping in Banach spaces. Computers and Mathematics with Applications, 58(7):1391–1396, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004489.

Fang:2008:WPE

[FHH08] Ya-Ping Fang, Rong Hu, and Nan-Jing Huang. Wellposedness for equilibrium problems and for optimization problems with equilibrium constraints. Computers and Mathematics with Applications, 55(1):89–100, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003379.

Fang:2005:NSV

[FHT05] Ya-Ping Fang, Nan-Jing Huang, and H. B. Thompson. A new system of variational inclusions with (H, η)-monotone operators in Hilbert spaces. Computers and Mathematics with Applications, 49 (2–3):365–374, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000398.

Farago:2008:SIA

[FHZ08] István Faragó, Ágnes Havasi, and Zahari Zlatev. Special issue on advanced numerical algorithms for large-scale computations: Introduction. Computers and Mathematics with Applications, 55(10):2183– 2184, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007456.

Feng:2005:RJD

[FJ05] Shaoqiang Feng and Zhongxiao Jia. A refined Jacobi– Davidson method and its correction equation. Computers and Mathematics with Applications, 49(2–3): 417–427, January/February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S089812210500043X.

Fang:2007:GKT

[FjH07] Min Fang and Nan jing Huang. Generalized L-KKM type theorems in topological spaces with an application. Computers and Mathematics with Applications, 53(12):1896– 1903, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001599.

Feng:2008:SS

[FJZ08] Feng Feng, Young Bae Jun, and Xianzhong Zhao. Soft semirings. Computers and Mathematics with Applications, 56(10):2621–2628, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004008.

Farago:2001:GFE

[FK01] I. Faragó and J. Karátson.

The gradient-finite element method for elliptic problems. Computers and Mathematics with Applications, 42(8– 9):1043–1053, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002206.

Fan:2005:SPS

[FKM05] Y. Y. Fan, R. E. Kalaba, and J. E. Moore II. Shortest paths in stochastic networks with correlated link costs. Computers and Mathematics with Applications, 49(9–10):1549– 1564, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001689.

Fazio:2003:MMM

[FL03] Riccardo Fazio and Randall J. LeVeque. Movingmesh methods for onedimensional hyperbolic problems using CLAWPACK. Computers and Mathematics with Applications, 45 (1–3):273–298, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800196.

Numerical methods in physics, chemistry, and engineering.

Farrag:2005:AFT

[FL05] A. A. Farrag and S. Lou. Applying fault-tolerant solutions of circulant graphs to multidimensional meshes. Computers and Mathematics with Applications, 50 (8–9):1383–1394, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003883.

Fang:2008:NSB

[FLfFL08] Tiegang Fang, Wei Liang, and Chia fon F. Lee. A new solution branch for the Blasius equation — a shrinking sheet problem. Computers and Mathematics with Applications, 56(12):3088–3095, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004410.

Farkas:2005:APM

[FLR05] A. Farkas, P. Lancaster, and P. Rózsa. Approximation of positive matrices by transitive matrices. Computers and Mathematics with Applications, 50(7):1033–1039, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003329.

Fornberg:2006:NCO

[FLW06] B. Fornberg, E. Larsson, and G. Wright. A new class of oscillatory radial basis functions. Computers and Mathematics with Applications, 51(8):1209– 1222, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000745.

Flyer:2006:EPR

[Fly06] N. Flyer. Exact polynomial reproduction for oscillatory radial basis functions on infinite lattices. Computers and Mathematics with Applications, 51(8):1199– 1208, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000733.

Franchini:2003:FIP

[FM03] E. Franchini and M. Maioli. Feynman integrals with point interactions. Computers and Mathematics with Applications, 46(5–6):685– 694, September 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103901349.

Ford:2005:UAR

[FM05] J. A. Ford and I. A. Moghrabi. The use of alternation and recurrences in two-step quasi-Newton methods. Computers and Mathematics with Applications, 50(7):1041–1050, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003330.

Feddag:2006:AEG

[FM06] M. L. Feddag and M. Mesbah. Approximate estimation in generalized linear mixed models with applications to the Rasch model. Computers and Mathematics with Applications, 51(2):269–278, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004645.

Fasano:2007:MMC

[FM07] Antonio Fasano and Alberto Mancini. A mathematical model for a class of frying processes. Computers and Mathematics with Applications, 53(3–4):395–412, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001903.

Fasano:2008:MFN

[FM08a] A. Fasano and A. Mancini.

Modelling the frying of a non-deformable specimen by immersion in edible oil. Computers and Mathematics with Applications, 55(2):149–161, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003434.

Feuer:2008:EPP

[FM08b] J. Feuer and K. T. Mc-Donnell. On the eventual periodicity of x_{n+1} = max{1/x_{n}, A_{n}/x_{n−1}} with a period-five parameter. Computers and Mathematics with Applications, 56 (4):883–890, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108000898.

Fragnelli:2008:GTA

[FM08c] Vito Fragnelli and Stefano Moretti. A game theoretical approach to the classification problem in gene expression data analysis. Computers and Mathematics with Applications, 55(5):950–959, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005548.

Fetecau:2008:SES

[FMFV08] C. Fetecau, A. Mahmood, Corina Fetecau, and D. Vieru. Some exact solutions for the helical flow of a generalized Oldroyd-b fluid in a circular cylinder. Computers and Mathematics with Applications, 56(12):3096–3108, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800446X.

Ford:2001:IUM

[For01] J. A. Ford. Implicit updates in multistep quasi-Newton methods. Computers and Mathematics with Applications, 42(8– 9):1083–1091, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002231.

Fotea:2009:FH

[Fot09] Violeta Leoreanu Fotea.

Fuzzy hypermodules. Computers and Mathematics with Applications, 57(3): 466–475, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108006779.

Florez:2002:DMM

[FP02] W. F. Florez and H. Power.

DRM multidomain mass conservative interpolation approach for the BEM solution of the two-dimensional Navier–Stokes equations. Computers and Mathematics with Applications, 43 (3–5):457–472, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210100298X.

Ferreira:2006:AGM

[FP06] F. Ferreira and A. Pacheco. Analysis of GI/M/s/c queues using uniformisation. Computers and Mathematics with Applications, 51(2): 291–304, January 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0898122105004670.

Fu:2002:NSH

[FQZ02] Chu-Li Fu, Chun-Yu Qiu, and You-Bin Zhu. A note on “sideways heat equation and wavelets” and constant e. Computers and Mathematics with Applications, 43(8–9):1125–1134, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800177.

Fix:2004:LSF

[FR04] G. J. Fix and J. P. Roof. Least squares finite-element solution of a fractional order two-point boundary value problem. Computers and Mathematics with Applications, 48(7–8):1017–1033, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003438.

From:2008:DBV

[Fro08] Steven G. From. Distance between vertical asymptotes of solutions of certain nonlinear differential equations. Computers and Mathematics with Applications, 55(11):2444– 2451, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006992.

Fernandez:2004:NAF

[FS04] J. R. Fernández and M. Sofonea. Numerical analysis of a frictionless viscoelastic contact problem with normal damped response. Computers and Mathematics with Applications, 47 (4–5):549–568, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900454.

Fuchs:2007:GNE

[FS07] M. Fuchs and G. Seregin.

A global nonlinear evolution problem for generalized Newtonian fluids: Local initial regularity of the strong solution. Computers and Mathematics with Applications, 53(3–4):509–520, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001794.

Farago:2001:QAM

[FT01] I. Faragó and P. Tarvainen. Qualitative analysis of matrix splitting methods. Computers and Mathematics with Applications, 42 (8–9):1055–1067, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002218.

Fu:2003:OTF

[FT03] S. C. Fu and L. Y. Tsai. Oscillation theorems for fourth-order partial difference equations. Computers and Mathematics with Applications, 46(2–3):253–262, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900290.

Ford:2005:TSF

[FT05] J. A. Ford and S. Tharmlikit. Three-step fixedpoint quasi-Newton methods for unconstrained optimisation. Computers and Mathematics with Applications, 50(7):1051–1060, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003342.

Francomano:2003:ROF

[FTC03] E. Francomano, A. Tortorici, and V. Calderone. Regularization of optical flow with M-band wavelet transform. Computers and Mathematics with Applications, 45(1–3):437–452, January/February 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800299.

Fan:2000:GPS

[FW00] Meng Fan and Ke Wang.

Global periodic solutions of a generalized n-species Gilpin–Ayala competition model. Computers and Mathematics with Applications, 40(10–11):1141–1151, November/December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100002285.

Feng:2004:EAP

[FW04a] Chunhua Feng and Peiguang Wang. The existence of almost periodic solutions of some delay differential equations. Computers and Mathematics with Applications, 47(8–9):1225–1231, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901162.

Fornberg:2004:SCM

[FW04b] B. Fornberg and G. Wright. Stable computation of multiquadric interpolants for all values of the shape parameter. Computers and Mathematics with Applications, 48(5–6):853–867, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003190.

Fan:2008:NSC

[FW08] Jianghua Fan and Wenhong Wei. Nonzero solutions for a class of set-valued variational inequalities in reflexive Banach spaces. Computers and Mathematics with Applications, 56(1):233– 241, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000114.

Fornberg:2004:SOR

[FWL04] B. Fornberg, G. Wright, and E. Larsson. Some observations regarding interpolants in the limit of flat radial basis functions. Computers and Mathematics with Applications, 47(1):37–55, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900041.

Frangos:2001:CTL

[FY01] C. Frangos and Y. Yavin.

Control of a three-link manipulator with inequality constraints on the trajectories of its joints. Computers and Mathematics with Applications, 41(12):1561– 1574, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001213.

Frangos:2003:MCR

[FY03a] C. Frangos and Y. Yavin.

Maneuvering control of a rolling four-links robot. Computers and Mathematics with Applications, 45 (12):1905–1926, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103900113. See

errata [Ano05g].

Frangos:2003:PPC

[FY03b] C. Frangos and Y. Yavin. Point-to-point control for nonlinear systems. Computers and Mathematics with Applications, 46(8– 9):1373–1385, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902264.

Frangos:2004:CMS

[FY04] C. Frangos and Y. Yavin. Control of the motion of a small-scale car with four steerable wheels. Computers and Mathematics with Applications, 48(3–4):437– 453, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840677.

Fornberg:2007:RPS

[FZ07a] Bengt Fornberg and Julia Zuev. The Runge phenomenon and spatially variable shape parameters in RBF interpolation. Computers and Mathematics with Applications, 54(3):379–398, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002210.

Fu:2007:PRL

[FZ07b] Ke-Ang Fu and Li-Xin Zhang. Precise rates in the law of the logarithm for negatively associated random variables. Computers and Mathematics with Applications, 54(5):687–698, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002477.

Feng:2007:PSC

[FZG07] Meiqiang Feng, Xuemei Zhang, and Weigao Ge. Positive solutions for a class of boundary value problems on time scales. Computers and Mathematics with Applications, 54(4):467–475, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002295.

Gulati:2006:SDN

[GA06] T. R. Gulati and D. Agarwal. Sufficiency and duality in nonsmooth multiobjective optimization involving generalized (F, α, π, d)– Type I functions. Computers and Mathematics with Applications, 52(1–2):81– 94, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001830.

Gadjiev:2007:EAU

[GA07] A. D. Gadjiev and A. Aral. The estimates of approximation by using a new type of weighted modulus of continuity. Computers and Mathematics with Applications, 54(1):127– 135, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002015.

Gordeziani:2006:HME

[GAA06] D. Gordeziani, G. Avalishvili, and M. Avalishvili. Hierarchical models of elastic shells in curvilinear coordinates. Computers and Mathematics with Applications, 51(12):1789–1808, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001234.

Galantai:2001:SAE

[Gal01] A. Galántai. A study of Auchmuty’s error estimate. Computers and Mathematics with Applications, 42 (8–9):1093–1102, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002243.

Galperin:2002:GOP

[Gal02a] E. A. Galperin. Global optimization in problems with uncertainties: the gamma algorithm. Computers and Mathematics with Applications, 44(7):853–862, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001980.

Galperin:2002:OID

[Gal02b]  E. A. Galperin.  Opti-

mization of implicitly defined functions. Computers and Mathematics with Applications, 44(7):969–977, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002079.

Galperin:2002:P

[Gal02c] Efim A. Galperin. Preface. Computers and Mathematics with Applications, 44(7):xiii–xiv, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001967.

Galantai:2005:PBT

[Gal05]  A. Galántai.  Perturba-

tion bounds for triangular and full rank factorizations. Computers and Mathematics with Applications, 50(7):1061–1068, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003354.

Galperin:2006:ROD

[Gal06a] E. A. Galperin. Reflections on optimality and dynamic programming. Computers and Mathematics with Applications, 52(1–2):235– 257, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001921.

Galperin:2006:P

[Gal06b] Efim A. Galperin. Preface. Computers and Mathematics with Applications, 52(1–2):xiii–xv, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001787.

Gal:2008:AGP

[Gal08a] Sorin G. Gal. Approximation and geometric properties of complex Favard– Szász–Mirakjan operators in compact disks. Computers and Mathematics with Applications, 56(4):1121– 1127, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001132.

Galperin:2008:ITT

[Gal08b] E. A. Galperin. Information transmittal and time uncertainty, measuring the speed of light and time of reflection, representations of Newton’s second law and related problems. Computers and Mathematics with Applications, 56(5):1271–1290, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001508.

Galperin:2008:IED

[Gal08c]  E. A. Galperin.  The

Isaacs equation for differential games, totally optimal fields of trajectories and related problems. Computers and Mathematics with Applications, 55(6):1333– 1362, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006049.

Galperin:2009:ITT

[Gal09a] E. A. Galperin. Information transmittal, time uncertainty and special relativity. Computers and Mathematics with Applications, 57(9):1554–1573, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006615.

Galperin:2009:SRM

[Gal09b] E. A. Galperin. Special relativity for motions with variable speed. Computers and Mathematics with Applications, 58(3):536–547, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002545.

Ganzburg:2000:MSB

[Gan00] M. Ganzburg. Moduli of smoothness and best approximation of functions with singularities. Computers and Mathematics with Applications, 40(1):81– 93, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001425.

Gan:2006:DLM

[Gan06] Siqing Gan. Dissipativity of linear θ-methods for integro-differential equations. Computers and Mathematics with Applications, 52(3–4):449–458, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002112.

Gardner:2000:D

[Gar00] M. Gardner. Dominono.

Computers and Mathematics with Applications, 39 (11):55–56, June 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100902657.

Garay:2001:EDN

[Gar01]  B. M. Garay.  Esti-

mates in discretizing normally hyperbolic compact invariant manifolds of ordinary differential equations. Computers and Mathematics with Applications, 42 (8–9):1103–1122, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002255.

Ganji:2007:AHV

[GAT07] D. D. Ganji, G. A. Afrouzi, and R. A. Talarposhti. Application of He’s variational iteration method for solving the reaction-diffusion equation with ecological parameters. Computers and Mathematics with Applications, 54(7–8):1010–1017, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002647.

Gauchman:2004:IIC

[Gau04]  H. Gauchman.  Integral

inequalities in q-calculus. Computers and Mathematics with Applications, 47 (2–3):281–300, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900259.

Gomez-Corral:2002:ASS

[GC02] A. Gómez-Corral. Analysis of a single-server retrial queue with quasi-random input and nonpreemptive priority. Computers and Mathematics with Applications, 43(6–7):767–782, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003200.

Gejadze:2006:OBCa

[GC06] I. Yu. Gejadze and G. J. M.

Copeland. Open boundary control problem for Navier–Stokes equations including a free surface: Adjoint sensitivity analysis. Computers and Mathematics with Applications, 52 (8–9):1243–1268, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003166.

Gorguis:2008:HEC

[GC08] Alice Gorguis and Wai Kit Benny Chan. Heat equation and its comparative solutions. Computers and Mathematics with Applications, 55(12):2973– 2980, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700819X.

Gejadze:2006:OBCb

[GCN06] I. Yu. Gejadze, G. J. M.

Copeland, and I. M. Navon. Open boundary control problem for Navier–Stokes equations including a free surface: Data assimilation. Computers and Mathematics with Applications, 52 (8–9):1269–1288, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003178.

Gimenes:2006:OIS

[GF06] L. P. Gimenes and M. Federson. Oscillation by impulses for a second-order delay differential equation. Computers and Mathematics with Applications, 52 (6–7):819–828, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002616.

Galperin:2006:SSU

[GG06] E. A. Galperin and I. Galperin. Small sample uniformity in random number generation. Computers and Mathematics with Applications, 52(1–2):95–108, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001842.

Ganji:2009:FGE

[GGA09] Z. Z. Ganji, D. D. Ganji, and A. Asgari. Finding general and explicit solutions of high nonlinear equations by the Exp-function method. Computers and Mathematics with Applications, 58(11–12):2124–2130, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001369.

Gopalakrishnan:2005:NSA

[GGCD05] J. Gopalakrishnan, L. E.

García-Castillo, and L. F. Demkowicz. Nédélec spaces in affine coordinates. Computers and Mathematics with Applications, 49(7–8): 1285–1294, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105001380.

Ghosh:2000:HFD

[GGD00] N. C. Ghosh, B. C. Ghosh, and L. Debnath. The hydromagnetic flow of a dusty visco-elastic fluid between two infinite parallel plates. Computers and Mathematics with Applications, 39(1–2):103–116, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812219900317X.

Ganji:2009:SNJ

[GGE09] Z. Z. Ganji, D. D. Ganji, and M. Esmaeilpour. Study on nonlinear Jeffery–Hamel flow by He’s semi-analytical methods and comparison with numerical results. Computers and Mathematics with Applications, 58(11–12):2107–2116, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001345.

Galperin:2003:PLG

[GGG03] E. A. Galperin, Alexandre Guinzbourg, and Anna Guinzbourg. Pseudorandom lattices for global optimization. Computers and Mathematics with Applications, 45(10–11):1537–1549, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001354.

Graef:2003:DSB

[GH03] J. R. Graef and J. Henderson. Double solutions of boundary value problems for 2m^{th}-order differential equations and difference equations. Computers and Mathematics with Applications, 45(6–9):873–885, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000634.

Guo:2009:DAR

[GHC09] Jinchao Guo, Xinhan Huang, and Yu Cui. Design and analysis of robust fault detection filter using LMI tools. Computers and Mathematics with Applications, 57(11–12):1743– 1747, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005488.

Ghatee:2008:SDI

[GHHD08] Mehdi Ghatee, S. Mehdi Hashemi, Behnam Hashemi, and Mehdi Dehghan. The solution and duality of imprecise network problems. Computers and Mathematics with Applications, 55(12):2767–2790, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007778.

Gunzburger:2006:NME

[GHJ06] M. Gunzburger, L. S. Hou, and L. Ju. A numerical method for exact boundary controllability problems for the wave equation. Computers and Mathematics with Applications, 51(5):721–750, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000137.

Guo:2004:ESD

[GHW04] Shangjiang Guo, Lihong Huang, and Lin Wang. Exponential stability of discrete-time Hopfield neural networks. Computers and Mathematics with Applications, 47(8–9):1249– 1256, April/May 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104901198.

Gille:2000:ACL

[Gil00] W. Gille. Analysis of the chord length distribution of the right circular cone for small chord lengths. Computers and Mathematics with Applications, 40 (8–9):1027–1035, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850130.

Gine:2002:CEC

[Gin02] J. Giné. Conditions for the existence of a center for the Kukles homogeneous systems. Computers and Mathematics with Applications, 43(10–11):1261–1269, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102000986.

Giraldo:2003:SWL

[Gir03] F. X. Giraldo. Strong and weak Lagrange–Galerkin spectral element methods for the shallow water equations. Computers and Mathematics with Applications, 45(1–3):97–121, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210380010X.

Giraldo:2006:HEL

[Gir06] F. X. Giraldo. Hybrid Eulerian–Lagrangian semiimplicit time-integrators. Computers and Mathematics with Applications, 52 (8–9):1325–1342, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210600321X.

Golbabai:2007:VIM

[GJ07] A. Golbabai and M. Javidi. A variational iteration method for solving parabolic partial differential equations. Computers and Mathematics with Applications, 54(7–8):987– 992, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002593.

Gonzalez:2003:SCM

[GJM03] C. González and A. Jiménez-Melado. Set-contractive mappings and difference equations in Banach spaces. Computers and Mathematics with Applications, 45 (6–9):1235–1243, March/ May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000944.

Guo:2007:ABT

[GJQ07] Shunsheng Guo, Hongbiao Jiang, and Qiulan Qi. Approximation by Bézier type of Meyer-König and zeller operators. Computers and Mathematics with Applications, 54(11–12):1387–1394, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003719.

Gyorfi:2001:STS

[GJS01] L. Győrfi, I. Joó, and A. Sövegjártó. Spline techniques for solving firstorder evolution equations in Banach spaces. Computers and Mathematics with Applications, 42(8– 9):1123–1133, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002267.

Guo:2000:PMP

[GK00] Xian-Zhong Guo and I. N.

Katz. A parallel multip method. Computers and Mathematics with Applications, 39(9–10):115– 123, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000912.

Galperin:2002:AGO

[GK02] E. A. Galperin and E. J.

Kansa. Application of global optimization and radial basis functions to numerical solutions of weakly singular Volterra integral equations. Computers and Mathematics with Applications, 43(3–5):491–499, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003005.

Guseinov:2003:LID

[GK03] G. Sh. Guseinov and B. Kaymakçalan. Lyapunov inequalities for discrete linear Hamiltonian systems. Computers and Mathematics with Applications, 45 (6–9):1399–1416, March/ May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000956.

Geum:2004:ACC

[GK04a] Young Hee Geum and Young Ik Kim. Accurate computation of component centers in the degree- n bifurcation set. Computers and Mathematics with Applications, 48(1–2):163– 175, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840070.

Guseinov:2004:IIH

[GK04b] G. Sh. Guseinov and I. Y. Karaca. Instability intervals of a Hill’s equation with piecewise constant and alternating coefficient. Computers and Mathematics with Applications, 47(2–3):319–326, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900272.

Gunzburger:2006:P

[GKKL06] Max Gunzburger, Kwang Ik Kim, Sang Dong Kim, and Hyung-Chun Lee. Preface. Computers and Mathematics with Applications, 51(5): i, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000241.

Georgiev:2008:MPR

[GKM08] I. Georgiev, J. Kraus, and S. Margenov. Multilevel preconditioning of rotated bilinear non-conforming FEM problems. Computers and Mathematics with Applications, 55(10):2280–2294, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007377.

Gunzburger:2000:DDM

[GL00] M. D. Gunzburger and Jeehyun Lee. A domain decomposition method for optimization problems for partial differential equations. Computers and Mathematics with Applications, 40(2–3):177–192, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001528.

Garcia-Lopez:2003:PLL

[GL03a] C. M. García-López. Piecewiselinearized and linearized θmethods for ordinary and partial differential equations. Computers and Mathematics with Applications, 45(1–3):351–381, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103800238.

Girardin:2003:SDE

[GL03b] V. Girardin and M. Rachdi Labsad. Spectral density estimation from random sampling for multiplicative stationary processes. Computers and Mathematics with Applications, 46(7):1009–1021, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901155.

Gao:2005:FMP

[GL05] Jinwu Gao and Baoding Liu. Fuzzy multilevel programming with a hybrid intelligent algorithm. Computers and Mathematics with Applications, 49(9–10):1539– 1548, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001677.

Gine:2006:IDC

[GL06] J. Giné and J. Llibre. Integrability, degenerate centers, and limit cycles for a class of polynomial differential systems. Computers and Mathematics with Applications, 51(9–10):1453– 1462, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000927.

Gang:2008:MVV

[GL08a] Xiao Gang and Sanyang Liu.

On Minty vector variationallike inequality. Computers and Mathematics with Applications, 56(2):311– 323, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000175.

Gu:2008:EMC

[GL08b] Wen-Xiang Gu and Bing Li. Effective method for constrained minimum

— reverse bridge theorem. Computers and Mathematics with Applications, 56(10):2629–2637, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800401X.

Gang:2009:ESG

[GL09a] Xiao Gang and Sanyang Liu. Existence of solutions for generalized vector quasi-variational-like inequalities without monotonicity. Computers and Mathematics with Applications, 58(8):1550–1557, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004933.

Gao:2009:DBS

[GL09b] Hong Gao and Huiping Long. Dynamic behavior for a system of neutral functional differential equations. Computers and Mathematics with Applications, 58(10):2046–2050, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109005720.

Geng:2009:AVI

[GL09c] Fazhan Geng and Yingzhen Lin. Application of the variational iteration method to inverse heat source problems. Computers and Mathematics with Applications, 58(11–12):2098–2102, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001321.

Glaister:2000:CUD

[Gla00] P. Glaister. Conservative upwind difference schemes for the shallow water equations. Computers and Mathematics with Applications, 39(5–6):189–199, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000535.

Glaister:2003:CUD

[Gla03] P. Glaister. Conservative upwind difference schemes for the Euler equations. Computers and Mathematics with Applications, 45 (10–11):1673–1682, May/ June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001469.

Glaister:2004:CUD

[Gla04] P. Glaister. Conservative upwind difference schemes for compressible flows of a real gas. Computers and Mathematics with Applications, 48(3–4):469–480, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840690.

Glaister:2005:CCU

[Gla05a] P. Glaister. A comparison of conservative upwind difference schemes for the shallow water equations. Computers and Mathematics with Applications, 49(11–12):1773– 1786, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001951.

Glaister:2005:CUD

[Gla05b]  P. Glaister.  Conser-

vative upwind difference schemes for open channel flows-theory and applications. Computers and Mathematics with Applications, 50(1–2):57–72, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002324.

Glaister:2006:CCU

[Gla06] P. Glaister. A comparison of conservative upwind difference schemes for the Euler equations. Computers and Mathematics with Applications, 51(6–7):879–888, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000447.

Glaister:2007:ACC

[Gla07a]  P. Glaister.  An anal-

ysis and comparison of conservative upwind difference schemes for ideal and non-ideal gases. Computers and Mathematics with Applications, 53(12):1921– 1932, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001666.

Glaister:2007:PSW

[Gla07b]  P. Glaister.  Prediction

of shallow water flows over dry beds. Computers and Mathematics with Applications, 54(2):296– 309, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002155.

Glaister:2008:CUD

[Gla08] P. Glaister. Conservative upwind difference schemes for compressible flows in a duct. Computers and Mathematics with Applications, 56(7):1787–1796, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002587.

Glaister:2009:CUD

[Gla09]  P. Glaister.  Conser-

vative upwind difference schemes for the Euler equations for real gas flows in a duct. Computers and Mathematics with Applications, 57(9):1432– 1437, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000662.

Geng:2009:PVI

[GLC09] Fazhan Geng, Yingzhen Lin, and Minggen Cui. A piecewise variational iteration method for Riccati differential equations. Computers and Mathematics with Applications, 58(11–12):2518–2522, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001941.

Giraud:2005:LOG

[GLR05] L. Giraud, J. Langou, and M. Rozloznik. The loss of orthogonality in the Gram– Schmidt orthogonalization process. Computers and Mathematics with Applications, 50(7):1069–1075, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003366.

Guo:2008:OBT

[GLTS08] S. M. Guo, K. T. Liu, J. S. H. Tsai, and L. S. Shieh. An observer-based tracker for hybrid interval chaotic systems with saturating inputs: the chaosevolutionary-programming approach. Computers and Mathematics with Applications, 55(6):1225–1249, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005913.

Gupur:2000:EUN

[GLZ00] G. Gupur, Xue-Zhi Li, and Guang-Tian Zhu. Existence and uniqueness of nonnegative solution of M/G^{B} /1 queueing model. Computers and Mathematics with Applications, 39(7–8):199– 209, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000766.

Gleyse:2000:ECN

[GM00] B. Gleyse and M. Moflih.

The exact computation of the number of zeros of a real polynomial in the open unit disk: an algebraic approach. Computers and Mathematics with Applications, 39(1–2):209– 214, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003247.

Garza:2009:STN

[GM09a] L. Garza and F. Marcellán. Szegő transformations and nth order associated polynomials on the unit circle. Computers and Mathematics with Applications, 57(10):1659– 1671, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002090.

Goswami:2009:MBS

[GM09b] V. Goswami and G. B. Mund. Multiserver bulk service discrete-time queue with finite buffer and renewal input. Computers and Mathematics with Applications, 57(8):1377– 1388, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000674.

Graef:2009:CTO

[GM09c] John R. Graef and Toufik Moussaoui. A class of n th-order BVPs with nonlocal conditions. Computers and Mathematics with Applications, 58(8):1662–1671, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004222.

Gomez-Mourelo:2009:DEC

[GMG09] Pablo Gómez-Mourelo and Marta Ginovart. The differential equation counterpart of an individual-based model for yeast population growth. Computers and Mathematics with Applications, 58(7):1360–1369, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005008.

Guerra:2006:SAM

[GMMB06] P. Jiménez Guerra, M. A.

Melguizo, and M. J. Munoz-Bouzo. Sensitivity analysis in multiobjective differential programming. Computers and Mathematics with Applications, 52(1–2):109– 120, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001854.

Guerra:2009:SAC

[GMMB09] P. Jiménez Guerra, M. A. Melguizo, and M. J. Muñoz-Bouzo. Sensitivity analysis in convex programming. Computers and Mathematics with Applications, 58(6):1239–1246, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004477.

Gwiazda:2007:FIF

[GMŚ07] Piotr Gwiazda, Josef Málek, and Agnieszka Świerczewska. On flows of an incompressible fluid with a discontinuous power-law-like rheology. Computers and Mathematics with Applications, 53(3–4):531–546, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001769.

Giovine:2008:WPP

[GMS08] P. Giovine, L. Margheriti, and M. P. Speciale. On wave propagation in porous media with strain gradient effects. Computers and Mathematics with Applications, 55(2):307–318, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003562.

Golbabai:2009:SSN

[GMS09] A. Golbabai, M. Mammadov, and S. Seifollahi. Solving a system of nonlinear integral equations by an RBF network. Computers and Mathematics with Applications, 57(10):1651– 1658, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002089.

Garcia-Mora:2008:MRP

[GMSRP08] Belén García-Mora, Cristina Santamaría, Gregorio Rubio, and José Luis Pontones. Modeling the recurrenceprogression process in bladder carcinoma. Computers and Mathematics with Applications, 56(3):619–630, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000539.

Guo:2009:STE

[GN09] Qianqiao Guo and Pengcheng Niu. Some theorems on existence and uniqueness of fixed points for decreasing operators. Computers and Mathematics with Applications, 57(9):1515– 1521, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000339.

Ganji:2007:AHM

[GNM07] D. D. Ganji, M. Nourollahi, and E. Mohseni. Application of He’s methods to nonlinear chemistry problems. Computers and Mathematics with Applications, 54(7–8):1122–1132, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002994.

Gordeziani:2005:FDM

[GNR05] N. Gordeziani, P. Natalini, and P. E. Ricci. Finite-difference methods for solution of nonlocal boundary value problems. Computers and Mathematics with Applications, 50 (8–9):1333–1344, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003846.

Ghasabi-Oskoei:2007:NAN

[GOMA07] H. Ghasabi-Oskoei, A. Malek, and A. Ahmadi. Novel artificial neural network with simulation aspects for solving linear and quadratic programming problems. Computers and Mathematics with Applications, 53(9): 1439–1454, May 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107001149.

Gong:2009:APS

[Gon09]  Shuhua Gong.  Anti-

periodic solutions for a class of Cohen–Grossberg neural networks. Computers and Mathematics with Applications, 58(2):341– 347, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002776.

Gordon:2000:MTE

[Gor00]  B. Gordon.  Multiple

tilings of Euclidean space by unit cubes. Computers and Mathematics with Applications, 39(11): 49–53, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001073.

Gosse:2000:WBF

[Gos00] L. Gosse. A well-balanced flux-vector splitting scheme designed for hyperbolic systems of conservation laws with source terms. Computers and Mathematics with Applications, 39(9–10):135– 159, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000936.

Gourley:2001:WFS

[Gou01] S. A. Gourley. Wave front solutions of a diffusive delay model for populations of Daphnia magna. Computers and Mathematics with Applications, 42(10– 11):1421–1430, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002516.

Gout:2002:SAS

[Gou02] C. Gout. C^{k} surface approximation from surface patches. Computers and Mathematics with Applications, 44(3–4):389–406, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001578.

Gonzalez-Parra:2009:DMT

[GPAA^{+}09] Gilberto C. González-Parra, Abraham J. Arenas, Diego F.

Aranda, Rafael J. Villanueva, and Lucas Jódar. Dynamics of a model of toxoplasmosis disease in human and cat populations. Computers and Mathematics with Applications, 57(10):1692–1700, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004902.

Guh:2008:FWA

[GPL08] Yuh-Yuan Guh, Rung-Wei Po, and E. Stanley Lee. The fuzzy weighted average within a generalized means function. Computers and Mathematics with Applications, 55(12):2699– 2706, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007699.

Gonzalez-Pinto:2003:SSA

[GPMPR03] S. González-Pinto, J. I.

Montijano, and S. Pérez-Rodríguez. Stabilized starting algorithms for collocation Runge–Kutta methods. Computers and Mathematics with Applications, 45(1–3):411–428, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103800275.

Gonzalez-Pinto:2001:IHO

[GPPRM01] S. González-Pinto, S. Pérez-Rodríguez, and J. I. Montijano. Implementation of high-order implicit Runge– Kutta methods. Computers and Mathematics with Applications, 41(7–8):1009– 1024, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003357.

Goseva-Popstojanova:2003:ABA

[GPT03] K. Goseva-Popstojanova and K. S. Trivedi. Architecturebased approaches to software reliability prediction. Computers and Mathematics with Applications, 46(7):1023–1036, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901167.

Gall:2006:FIR

[GPvZ06] J. Gáll, G. Pap, and M. C. A. van Zuijlen. Forward interest rate curves in discrete time settings driven by random fields. Computers and Mathematics with Applications, 51(3–4):387– 396, February 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105005079.

Guo:2005:PAB

[GQL05] Shunsheng Guo, Qiulan Qi, and Lixia Liu. Pointwise approximation by Baskakov quasi-interpolants. Computers and Mathematics with Applications, 49(7–8):1011– 1020, May/June 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105001173.

Gravvanis:2003:EPD

[Gra03] G. A. Gravvanis. Explicit preconditioned domain decomposition schemes for solving nonlinear boundary value problems. Computers and Mathematics with Applications, 45(1–3): 263–272, January/February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800184.

Graillat:2008:ASZ

[Gra08]  Stef Graillat.  Accurate

simple zeros of polynomials in floating point arithmetic. Computers and Mathematics with Applications, 56(4):1114–1120, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001120.

Greenspan:2001:EAC

[Gre01]  D. Greenspan.  Elec-

tron attraction and the chemical bond. Computers and Mathematics with Applications, 41(12):1515– 1520, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210100116X.

Greenspan:2003:PMC

[Gre03]  D. Greenspan.  Parti-

cle modelling of the cavity problem for liquids. Computers and Mathematics with Applications, 45 (4–5):715–722, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000312.

Greenspan:2004:ELF

[Gre04] D. Greenspan. The extraction of laminar flow from certain Brownian motions. Computers and Mathematics with Applications, 48(12):1915–1928, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004493.

Greenspan:2009:CSM

[Gre09] Donald Greenspan. Computer studies of molecular air flow around a circular cylinder. Computers and Mathematics with Applications, 58(3):414–421, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006457.

Garey:2001:PML

[GS01] L. E. Garey and R. E.

Shaw. A parallel method for linear equations with tridiagonal Toeplitz coefficient matrices. Computers and Mathematics with Applications, 42(1–2): 1–11, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001250.

Grinfeld:2004:LEP

[GS04] P. Grinfeld and G. Strang.

The Laplacian eigenvalues of a polygon. Computers and Mathematics with Applications, 48(7–8):1121– 1133, October 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104003505.

Galantai:2005:DKB

[GS05] Aurél Galántai and György Szeidl. Dedication katalin balla 1947-2005. Computers and Mathematics with Applications, 50(7):xv, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003469.

Grzymkowski:2006:OPI

[GS06] R. Grzymkowski and D. Slota. One-phase inverse Stefan problem solved by Adomian decomposition method. Computers and Mathematics with Applications, 51(1):33–40, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105004475.

Girault:2007:TDG

[GS07a] V. Girault and M. Saadouni. On a time-dependent gradetwo fluid model in two dimensions. Computers and Mathematics with Applications, 53(3–4):347–360, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001952.

Goswami:2007:PAG

[GS07b] V. Goswami and S. K.

Samanta. Performance analysis of a GI^{[X]} /Geo/m/N queue with partial- and total-batch rejection. Computers and Mathematics with Applications, 54(2): 203–219, July 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107002064.

Gong:2008:GEU

[GS08] Zengtai Gong and Yabin Shao. Global existence and uniqueness of solutions for fuzzy differential equations under dissipative-type conditions. Computers and Mathematics with Applications, 56(10):2716–2723, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004112.

Guan:2009:OCF

[GS09] Kaizhong Guan and Jianhua Shen. Oscillation criteria for a first-order impulsive neutral differential equation of Euler form. Computers and Mathematics with Applications, 58(4):670–677, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003733.

Ghorbani:2008:ESN

[GSN08] Asghar Ghorbani and Jafar Saberi-Nadjafi. Exact solutions for nonlinear integral equations by a modified homotopy perturbation method. Computers and Mathematics with Applications, 56(4):1032–1039, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001053.

Grau-Sanchez:2008:AAL

[GSNDB^{+}08] M. Grau-Sánchez, M. Noguera, J. L. Díaz-Barrero, V. D. Chamizo, and T. Rodrigo. An application of an associative learning model to a Morris pool with a single landmark. Computers and Mathematics with Applications, 56(1):212– 217, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000072.

Gangadharan:2002:GHF

[GSS02] A. Gangadharan, T. N.

Shanmugam, and H. M. Srivastava. Generalized hypergeometric functions associated with k- uniformly convex functions. Computers and Mathematics with Applications, 44(12):1515– 1526, December 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102002754.

Gupta:2004:CQL

[GSS04] U. C. Gupta, S. K. Samanta, and R. K. Sharma. Computing queue length and waiting-time distributions in finite-buffer discrete-time multiserver queues with late and early arrivals. Computers and Mathematics with Applications, 48(10– 11):1557–1573, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004109.

Gan:2002:ELM

[GSZ02] Siqing Gan, Geng Sun, and Weimin Zheng. Errors of linear multistep methods and Runge–Kutta methods for singular perturbation problems with delays. Computers and Mathematics with Applications, 44 (8–9):1157–1173, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002237.

Gao:2000:CAP

[GT00] Yi-Tian Gao and Bo Tian. Computer algebra, Painlevé analysis and the timedependent-coefficient nonlinear Schrödinger equation. Computers and Mathematics with Applications, 40 (8–9):1107–1108, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100850208.

Ghahramani:2008:NGM

[GT08] M. Ghahramani and A. Thavaneswaran. A note on GARCH model identification. Computers and Mathematics with Applications, 55(11):2469–2475, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700747X.

Ganji:2007:VIM

[GTJ07] D. D. Ganji, Hafez Tari, and M. Bakhshi Jooybari. Variational iteration method and homotopy perturbation method for nonlinear evolution equations. Computers and Mathematics with Applications, 54(7–8):1018–1027, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002660.

Gu:2007:AFP

[Gu07]  Feng Gu.  Approximate

fixed point sequence for a finite family of asymptotically pseudocontractive mappings. Computers and Mathematics with Applications, 53(12):1792–1799, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001381.

Gu:2008:ESA

[Gu08] Rongbao Gu. On ergodicity of systems with the asymptotic average shadowing property. Computers and Mathematics with Applications, 55(6):1137– 1141, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005792.

Guo:2001:ESO

[Guo01] Dajun Guo. Existence of solutions for n^{th}-order integrodifferential equations in Banach spaces. Computers and Mathematics with Applications, 41(5–6):597– 606, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003035.

Guo:2002:COI

[Guo02] Dajun Guo. A class of n^{th}-order impulsive integrodifferential equations in Banach spaces. Computers and Mathematics with Applications, 44(10– 11):1339–1356, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002614.

Guo:2003:JSM

[Guo03]  Ben-Yu Guo.  Jacobi

spectral method for differential equations with rough asymptotic behaviors at infinity. Computers and Mathematics with Applications, 46(1):95– 104, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900836.

Guo:2007:LSR

[Guo07] Ji-Ming Guo. The Laplacian spectral radius of a graph under perturbation. Computers and Mathematics with Applications, 54(5): 709–720, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107002490.

Gupur:2002:WPM

[Gup02a] G. Gupur. Well-posedness of M/G/1 queueing model with single vacations. Computers and Mathematics with Applications, 44(8– 9):1041–1056, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002134.

Gupur:2002:RSM

[Gup02b]  Geni Gupur.  Resol-

vent set of the M/M^{B} /1 operator. Computers and Mathematics with Applications, 44(1–2):67– 82, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210200130X.

Gupta:2003:RAN

[Gup03] V. Gupta. Rate of approximation by a new sequence of linear positive operators. Computers and Mathematics with Applications, 45(12):1895– 1904, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900101.

Gupta:2004:BVK

[Gup04] V. Gupta. The Bézier variant of Kantorovitch operators. Computers and Mathematics with Applications, 47(2–3):227–232, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900193.

Guy:2000:PS

[Guy00]  R. K. Guy.  Problems

sessions 92-05-29 and 92-05-31. Computers and Mathematics with Applications, 39(11):169–171, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001231.

Ganesh:2001:DGM

[GW01a] M. Ganesh and D. J. Worth. A discrete Galerkin method for a catalytic combustion model. Computers and Mathematics with Applications, 41(12):1545– 1557, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001195.

Grzegorczyk:2001:OHO

[GW01b] G. Grzegorczyk and J. Werbowski. Oscillation of higher-order linear difference equations. Computers and Mathematics with Applications, 42(3–5):711–717, August/September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101001900.

Gerisch:2003:PLO

[GW03a] A. Gerisch and R. Weiner.

The positivity of loworder explicit Runge–Kutta schemes applied in splitting methods. Computers and Mathematics with Applications, 45(1–3):53–67, January/February 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210380007X.

Gopalsamy:2003:GAC

[GW03b] K. Gopalsamy and Peixuan Weng. Global attractivity in a competition system with feedback controls. Computers and Mathematics with Applications, 45 (4–5):665–676, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000269.

Guo:2006:MLH

[GW06] Ben-Yu Guo and Tian-Jun Wang. Mixed Legendre– Hermite spectral method for heat transfer in an infinite plate. Computers and Mathematics with Applications, 51(5):751–768, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000149.

Guo:2004:BPS

[GX04] Bao-Zhu Guo and Yu Xie. Basis property and stabilization of a translating tensioned beam through a pointwise control force. Computers and Mathematics with Applications, 47(8–9):1397–1409, April/ May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901320.

Gui:2006:SEP

[GY06] Z. Gui and X.-S. Yang. Stability and existence of periodic solutions of periodic cellular neural networks with time-vary delays. Computers and Mathematics with Applications, 52(12):1657–1670, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210600349X.

Gao:2008:CFV

[GY08] Fuzheng Gao and Yirang Yuan. The characteristic finite volume element method for the nonlinear convection-dominated diffusion problem. Computers and Mathematics with Applications, 56(1):71– 81, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008267.

Gyllenberg:2009:FLC

[GY09] Mats Gyllenberg and Ping Yan. Four limit cycles for a three-dimensional competitive Lotka–Volterra system with a heteroclinic cycle. Computers and Mathematics with Applications, 58(4):649–669, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003344.

Gyires:2000:SPE

[Gyi00] B. Gyires. Solutions of permanental equations regarding stochastic matrices. Computers and Mathematics with Applications, 40(4–5):421–431, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210000170X.

Gyires:2001:CVW

[Gyi01]  B. Gyires.  Contribu-

tion to van der Waerden’s conjecture. Computers and Mathematics with Applications, 42(10– 11):1431–1437, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002528.

Guh:2008:EPE

[GYPL08] Yuh-Yuan Guh, Miin-Shen Yang, Rung-Wei Po, and E. Stanley Lee. Establishing performance evaluation structures by fuzzy relation-based cluster analysis. Computers and Mathematics with Applications, 56(2):572–582, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0898122108000916.

Guh:2009:IVF

[GYPL09] Yuh-Yuan Guh, Miin-Shen Yang, Rung-Wei Po, and E. Stanley Lee. Intervalvalued fuzzy relation-based clustering with its application to performance evaluation. Computers and Mathematics with Applications, 57(5):841–849, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006639.

Gao:2005:SFI

[GYS05] Chun-Yi Gao, Shao-Mou Yuan, and H. M. Srivastava. Some functional inequalities and inclusion relationships associated with certain families of integral operators. Computers and Mathematics with Applications, 49(11–12):1787– 1795, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001963.

Guvenilir:2006:SOO

[GZ06] A. F. Güvenilir and A. Zafer.

Second-order oscillation of forced functional differential equations with oscillatory potentials. Computers and Mathematics with Applications, 51(9–10):1395– 1404, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000873.

Gao:2008:WCA

[GZ08] Bo Gao and Weinian Zhang. A weaker criterion of asymptotic stability. Computers and Mathematics with Applications, 56(7):1693–1698, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002460.

Gu:2009:MVA

[GZZ09] Ping Gu, QingSheng Zhu, and Cheng Zhang. A multiview approach to semisupervised document classification with incremental naive Bayes. Computers and Mathematics with Applications, 57(6):1030– 1036, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005610.

Ha:2001:NSM

[Ha01]  Sung N. Ha.  A non-

linear shooting method for two-point boundary value problems. Computers and Mathematics with Applications, 42(10– 11):1411–1420, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002504.

Hadizadeh:2003:PEE

[Had03]  M. Hadizadeh.  Poste-

riori error estimates for the nonlinear Volterra– Fredholm integral equations. Computers and Mathematics with Applications, 45(4–5):677–687, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000270.

Haidar:2003:AST

[Hai03] N. H. S. Haidar. An amplitude spectral tonometer. Computers and Mathematics with Applications, 46(5–6):959–969, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901568.

Hayat:2007:PTT

[HAKA07] T. Hayat, Ambreen Afsar, M. Khan, and S. Asghar. Peristaltic transport of a third order fluid under the effect of a magnetic field. Computers and Mathematics with Applications, 53(7):1074–1087, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700065X.

Han:2008:NIS

[HaLC08] Xiuping Han, Jun an Lu, and Guanrong Chen. Nonlinear integral synchronization of ring networks. Computers and Mathematics with Applications, 55(4): 808–818, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107005123.

Han:2004:RSC

[Han04] Qing-Long Han. Robust stability for a class of linear systems with time-varying delay and nonlinear perturbations. Computers and Mathematics with Applications, 47(8–9):1201–1209, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901149.

Han:2006:NCP

[Han06]  D. Han.  A new class

of projection and contraction methods for solving variational inequality problems. Computers and Mathematics with Applications, 51(6–7):937–950, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000289.

Han:2008:HEP

[Han08] Deren Han. A hybrid entropic proximal decomposition method with selfadaptive strategy for solving variational inequality problems. Computers and Mathematics with Applications, 55(1):101–115, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003380.

Hardy:2001:HMB

[Har01]  R. L. Hardy.  Hardy’s

multiquadric-biharmonic method for gravity field predictions II. Computers and Mathematics with Applications, 41(7–8):1043–1048, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003382.

Hassenpflug:2002:EAS

[Has02] W. C. Hassenpflug. Error analysis in the series evaluation of the exponential type integral e^{z}E_{1}(z). Computers and Mathematics with Applications, 43(1–2):207– 266, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002838.

Hassenpflug:2003:TUB

[Has03] W. C. Hassenpflug. Torsion of uniform bars with polygon cross-section. Computers and Mathematics with Applications, 46(2–3):313– 392, July/August 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103900332.

Hasanov:2004:IMS

[Has04] V. I. Hasanov. An iterative method for solving the spectral problem of complex symmetric matrices. Computers and Mathematics with Applications, 47(4–5):529–540, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900430.

Hayat:2007:MFS

[HASA07] T. Hayat, Naveed Ahmed, M. Sajid, and S. Asghar. On the MHD flow of a second grade fluid in a porous channel. Computers and Mathematics with Applications, 54(3):407–414, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002234.

Haupt:2006:QEM

[Hau06] S. E. Haupt. A quadratic empirical model formulation for dynamical systems using a genetic algorithm. Computers and Mathematics with Applications, 51(3–4):431–440, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005109.

Havasi:2008:WAD

[Hav08] Ágnes Havasi. Wave analysis for different splittings of the shallow water equations on the β-plane. Computers and Mathematics with Applications, 55(10):2295– 2305, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007365.

Hayat:2006:ESR

[Hay06]  T. Hayat.  Exact solu-

tions to rotating flows of a Burgers’ fluid. Computers and Mathematics with Applications, 52(10– 11):1413–1424, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003294.

Hilgers:2004:CDT

[HB04] J. W. Hilgers and B. S. Bertram. Comparing different types of approximators for choosing the parameters in the regularization of ill-posed problems. Computers and Mathematics with Applications, 48(10– 11):1779–1790, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104004274.

Huang:2000:GNM

[HBCK00] Nan-Jing Huang, Min-Ru Bai, Yeol Je Cho, and Shin Min Kang. Generalized nonlinear mixed quasi-variational inequalities. Computers and Mathematics with Applications, 40(2–3):205–215, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001541.

He:2006:HGL

[HBP06] Haiwu He, G. Bergere, and S. Petiton. A hybrid GMRES/LS-Arnoldi method to accelerate the parallel solution of linear systems. Computers and Mathematics with Applications, 51(11):1647– 1662, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001337.

Hu:2001:NSD

[HC01] Guang-Da Hu and B. Cahlon.

The numerical solution of discrete-delay systems. Computers and Mathematics with Applications, 42(3–5):735–741, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001936.

Huang:2002:SAN

[HC02] Chengming Huang and Qianshun Chang. Stability analysis of numerical methods for systems of functional-differential and functional equations. Computers and Mathematics with Applications, 44(5–6): 717–729, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102001852.

He:2009:DBI

[HC09] Mengxin He and Fengde Chen. Dynamic behaviors of the impulsive periodic multi-species predator– prey system. Computers and Mathematics with Applications, 57(2):248–265, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800610X.

He:2008:AEM

[HCG08] Daan He, Nick Cercone, and Zhenmei Gu. Applying the extended massconstraint EM algorithm to image retrieval. Computers and Mathematics with Applications, 56(4):1010– 1024, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800103X.

Hernandez:2007:LSL

[HCL07] F. F. Hernández, J. E. Castillo, and G. A. Larrazábal. Large sparse linear systems arising from mimetic discretization. Computers and Mathematics with Applications, 53(1):1– 11, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000156.

Huang:2001:CGL

[HCLF01] Chengming Huang, Guangnan Chen, Shoufu Li, and Hongyuan Fu. Dconvergence of general linear methods for stiff delay differential equations. Computers and Mathematics with Applications, 41(5–6):627– 639, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003060.

Huang:2000:CIP

[HCLJ00] Nan Jing Huang, Y. J. Cho, B. S. Lee, and J. S. Jung. Convergence of iterative processes with errors for setvalued pseudocontractive and accretive type mappings in Banach spaces. Computers and Mathematics with Applications, 40(10– 11):1127–1139, November/ December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002273.

Hu:2003:ACS

[HCTC03] Yi-Chung Hu, Ruey-Shun Chen, Gwo-Hshiung Tzeng, and Yu-Jing Chiu. Acquisition of compound skills and learning costs for expanding competence sets. Computers and Mathematics with Applications, 46(5–6):831–848, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901465.

Hong:2007:NAF

[HCWC07] Won-Sin Hong, Shi-Jay Chen, Li-Hui Wang, and Shyi-Ming Chen. A new approach for fuzzy information retrieval based on weighted power-mean averaging operators. Computers and Mathematics with Applications, 53(12):1800– 1819, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001496.

Hsu:2000:TPR

[HCY00] Steen J. Hsu, Yu G. Chen, and Maria C. Yuang. Terminal-pair reliability in ATM virtual path networks. Computers and Mathematics with Applications, 40 (6–7):885–894, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002042.

Hollander:2008:BEP

[HD08] Y. Hollander and D. Durban. Bifurcation of elastoplastic pressure-sensitive spheres. Computers and Mathematics with Applications, 55(2):257–267, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003525.

Han:2009:IMM

[HD09] Lili Han and Feng Ding.

Identification for multirate multi-input systems using the multi-innovation identification theory. Computers and Mathematics with Applications, 57(9):1438– 1449, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000650.

Hsieh:2003:SRB

[HdLS03] Tsai-Ming Hsieh, Shy der Lin, and H. M. Srivastava. Some relationships between certain families of ordinary and fractional differential equations. Computers and Mathematics with Applications, 46(10– 11):1483–1492, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901854.

He:2000:QMO

[He00a] Tian-Xiao He. C^{1} quadratic macroelements and C^{1} orthogonal multiresolution analyses in 2D. Computers and Mathematics with Applications, 40(1):95– 106, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001437.

He:2000:OLP

[He00b]  Yong He.  The opti-

mal on-line parallel machine scheduling. Computers and Mathematics with Applications, 39(7–8):117– 121, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000705.

He:2002:CBI

[He02]  Yinnian He.  Coupling

boundary integral and finite element methods for the Oseen coupled problem. Computers and Mathematics with Applications, 44(10– 11):1413–1429, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002663.

He:2003:LPT

[He03] Ji-Huan He. Linearized perturbation technique and its applications to strongly nonlinear oscillators. Computers and Mathematics with Applications, 45(1– 3):1–8, January/February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800020.

He:2004:BST

[He04] Tian-Xiao He. Biorthogonal spline type wavelets. Computers and Mathematics with Applications, 48(9): 1319–1334, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104003839.

He:2009:ACI

[He09a] Ji-Huan He. Application of He Chengtian’s interpolation to Bethe equation. Computers and Mathematics with Applications, 58(11–12):2427–2430, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001837.

He:2009:EIH

[He09b]  Ji-Huan He.  An ele-

mentary introduction to the homotopy perturbation method. Computers and Mathematics with Applications, 57(3):410–412, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005592.

Hemeda:2008:VIM

[Hem08] A. A. Hemeda. Variational iteration method for solving wave equation. Computers and Mathematics with Applications, 56(8):1948–1953, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002721.

Hernandez:2001:CAA

[Her01] M. A. Hernández. Chebyshev’s approximation algorithms and applications. Computers and Mathematics with Applications, 41(3– 4):433–445, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0898122100002868.

Hu:2000:SSI

[HF00] Cheng-Feng Hu and Shu-Cherng Fang. Solving a system of infinitely many fuzzy inequalities with piecewise linear membership functions. Computers and Mathematics with Applications, 40(6–7):721–733, September/October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001917.

Huang:2004:IPE

[HF04] Nan-Ding Huang and Ya-Ping Fang. Iterative processes with errors for nonlinear set-valued variational inclusions involving accretive type mappings. Computers and Mathematics with Applications, 47 (4–5):727–738, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900600.

He:2003:ETS

[HG03] Xiaoming He and Weigao Ge. Existence of three solutions for a quasilinear twopoint boundary value problem. Computers and Mathematics with Applications, 45(4–5):765–769, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000373.

Huang:2002:NIP

[HGH02] Nan-Jing Huang, Cheng-Jia Gao, and Xiao-Ping Huang. New iteration procedures with errors for multivalued ϕ-strongly pseudocontractive and ϕ-strongly accretive mappings. Computers and Mathematics with Applications, 43(10–11):1381–1390, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001062.

Huang:2007:FEM

[HGS07] Jianguo Huang, Ling Guo, and Zhongci Shi. A finite element method for investigating general elastic multi-structures. Computers and Mathematics with Applications, 53(12):1867– 1895, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001587.

Hamada:2000:ABS

[HH00] N. Hamada and T. Helleseth. Arcs, blocking sets, and minihypers. Computers and Mathematics with Applications, 39(11):159– 168, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000122X.

Han:2002:CAB

[HH02a] Houde Han and Zhongyi Huang. A class of artificial boundary conditions for heat equation in unbounded domains. Computers and Mathematics with Applications, 43(6–7):889–900, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003297.

Han:2002:EAB

[HH02b] Houde Han and Zhongyi Huang. Exact and approximating boundary conditions for the parabolic problems on unbounded domains. Computers and Mathematics with Applications, 44(5–6):655–666, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001803.

Hirano:2003:CTM

[HH03] N. Hirano and Zhenyu Huang. Convergence theorems for multivalued Φhemicontractive operators and Φ-strongly accretive operators. Computers and Mathematics with Applications, 46(10–11):1461–1471, November/December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103901830.

Hahm:2004:ANN

[HH04a] Nahmwoo Hahm and Bum Il Hong. An approximation by neural networks with a fixed weight. Computers and Mathematics with Applications, 47(12):1897– 1903, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840252.

He:2004:MIT

[HH04b] Zhimin He and Xiaoming He. Monotone iterative technique for impulsive integro-differential equations with periodic boundary conditions. Computers and Mathematics with Applications, 48(1–2):73– 84, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839993.

Hecht:2009:CSC

[HH09] Martin Hecht and Jens Harting. Computational steering of cluster formation in Brownian suspensions. Computers and Mathematics with Applications, 58(5):995–1002, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001047.

Huang:2008:EAP

[HHC08] Xuehai Huang, Jianguo Huang, and Yu Chen. Error analysis of a parameter expansion method for corrosion detection in a pipe. Computers and Mathematics with Applications, 56(10):2539–2549, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003908.

Hedetniemi:2003:SSA

[HHJS03] S. M. Hedetniemi, S. T.

Hedetniemi, D. P. Jacobs, and P. K. Srimani. Self-stabilizing algorithms for minimal dominating sets and maximal independent sets. Computers and Mathematics with Applications, 46(5–6):805–811, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210390143X.

Hayat:2006:HEF

[HHK06] T. Hayat, M. Hussain, and M. Khan. Hall effect on flows of an Oldroydb fluid through porous medium for cylindrical geometries. Computers and Mathematics with Applications, 52(3–4):269–282, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001957.

Hoffman:2006:UVM

[HHL^{+}06] R. N. Hoffman, J. M. Henderson, S. M. Leidner, C. Grassotti, and T. Nehrkorn. Using 4d-VAR to move a simulated tropical cyclone in a mesoscale model. Computers and Mathematics with Applications, 52(8– 9):1193–1204, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003257.

Huang:2008:ASA

[HHL08] Li Huang, Yong Huang, and Xian-Fang Li. Approximate solution of Abel integral equation. Computers and Mathematics with Applications, 56(7):1748–1757, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002538.

Hasson:2007:SDE

[HHR07] Maurice Hasson, James M.

Hyman, and Juan M. Restrepo. A strategy for detecting extreme eigenvalues bounding gaps in the discrete spectrum of selfadjoint operators. Computers and Mathematics with Applications, 53(8):1271– 1283, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000867.

Hu:2002:MRB

[HHRA02] Xu-Guang Hu, Tak-San Ho, H. Rabitz, and A. Askar. Multivariate radial basis interpolation for solving quantum fluid dynamical equations. Computers and Mathematics with Applications, 43(3–5):525–537, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003030.

He:2006:CSF

[HHS06] Tian-Xiao He, Leetsch C. Hsu, and Peter J.-S. Shiue. Convergence of the summation formulas constructed by using a symbolic operator approach. Computers and Mathematics with Applications, 51(3–4):441–450, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105005110.

He:2007:SGF

[HHS07] Tian-Xiao He, Leetsch C. Hsu, and Peter J.-S. Shiue. Symbolization of generating functions; an application of the Mullin– Rota theory of binomial enumeration. Computers and Mathematics with Applications, 54(5):664–678, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002441.

Huang:2008:MSE

[HHW08] Chuangxia Huang, Yigang He, and Hainu Wang. Mean square exponential stability of stochastic recurrent neural networks with timevarying delays. Computers and Mathematics with Applications, 56(7):1773–1778, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002563.

He:2009:POS

[HHY09] Tian-Xiao He, Leetsch C.

Hsu, and Dongsheng Yin. A pair of operator summation formulas and their applications. Computers and Mathematics with Applications, 58(7):1340–1348, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004702.

Hilscher:2001:DSC

[Hil01]  R. Hilscher.  Discrete

spectra criteria for certain classes of singular differential and difference operators. Computers and Mathematics with Applications, 42(3–5):465–476, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001705.

Hiptmair:2006:OP

[Hip06]  R. Hiptmair.  Operator

preconditioning. Computers and Mathematics with Applications, 52(5):699–706, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002495.

Huang:2005:CPS

[HJL05] Chih-Peng Huang, Yau-Tarng Juang, and Hui-Ling Lin. A complex projection scheme and applications. Computers and Mathematics with Applications, 49(4):515–524, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000465.

Han:2005:FDM

[HJW05] Houde Han, Jicheng Jin, and Xiaonan Wu. A finite-difference method for the one-dimensional timedependent schrödinger equation on unbounded domain. Computers and Mathematics with Applications, 50 (8–9):1345–1362, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003858.

Han:2003:TOM

[HJWL03] Soowhan Han, Seungju Jang, Youngwoon Woo, and Jungsik Lee. Thirdorder moment spectrum and weighted fuzzy classifier for robust 2-D object recognition. Computers and Mathematics with Applications, 45(4–5):699–707, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000294.

Hsieh:2007:HCS

[HK07] Sun-Yuan Hsieh and Che-Nan Kuo. Hamiltonianconnectivity and strongly Hamiltonian-laceability of folded hypercubes. Computers and Mathematics with Applications, 53(7):1040– 1044, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000594.

Hayashi:2008:CSS

[HK08] Hidemitsu Hayashi and Shuichi Kubo. Computer simulation study on filtration of soot particles in diesel particulate filter. Computers and Mathematics with Applications, 55(7):1450–1460, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006396.

Hamza:2008:RS

[HKA08] Alaa. E. Hamza and R. Khalaf-Allah. On the recursive se-

quence \frac{x_{n+1}=A ∏_{i=l}^{k} x_{n−2i−1}}{B+C ∏_{i=l}^{k−1} x_{n−2i}}.

Computers and Mathematics with Applications, 56(7): 1726–1731, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108002502.

Halim:2003:NIC

[HKL03] A. A. Halim, S. P. Kshevetskii, and S. B. Leble. Numerical integration of a coupled Korteweg–de Vries system. Computers and Mathematics with Applications, 45(4–5):581–591, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210300018X.

Heuveline:2009:THP

[HKL09] Vincent Heuveline, Mathias J. Krause, and Jonas Latt. Towards a hybrid parallelization of lattice Boltzmann methods. Computers and Mathematics with Applications, 58(5):1071–1080, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002454.

Hong:2000:SNF

[HL00] Don Hong and Huan-Wen Liu. Some new formulations of smoothness conditions and conformality conditions for bivariate splines. Computers and Mathematics with Applications, 40(1):117– 125, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001450.

Ha:2002:NST

[HL02a] S. N. Ha and C. R. Lee.

Numerical study for twopoint boundary value problems using Green’s functions. Computers and Mathematics with Applications, 44(12):1599–1608, December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002821.

Han:2002:TNS

[HL02b] Deren Han and Hong K.

Lo. Two new selfadaptive projection methods for variational inequality problems. Computers and Mathematics with Applications, 43(12):1529– 1537, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001165.

Huang:2002:SSA

[HL02c] Tetz C. Huang and Ji-Cherng Lin. A selfstabilizing algorithm for the shortest path problem in a distributed system. Computers and Mathematics with Applications, 43(1–2):103–109, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002760.

Huang:2003:EBM

[HL03a] Ting-Zhu Huang and Fu-Ti Liu. An error bound for the MAOR method. Computers and Mathematics with Applications, 45 (10–11):1739–1748, May/ June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001524.

Huo:2003:OCT

[HL03b] Hai-Feng Huo and Wan-Tong Li. Oscillation of certain two-dimensional nonlinear difference systems. Computers and Mathematics with Applications, 45 (6–9):1221–1226, March/ May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000890.

Huang:2004:PDS

[HL04] Wentao Huang and Yirong Liu. A polynomial differential system with nine limit cycles at infinity. Computers and Mathematics with Applications, 48(3–4):577– 588, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840781.

Huang:2005:ECD

[HL05] Ting-Zhu Huang and Xing-Ping Liu. Estimations for certain determinants. Computers and Mathematics with Applications, 50(10– 12):1677–1684, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004189.

Hu:2006:CCF

[HL06] Hsin-Yun Hu and Zi-Cai Li. Combinations of collocation and finiteelement methods for Poisson’s equation. Computers and Mathematics with Applications, 51(12):1831– 1853, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001283.

Hwang:2007:FLT

[HL07] Kyo-Shin Hwang and Sungchul Lee. Functional limit theorems for the increments of d-dimensional Gaussian processes in a Hölder type norm. Computers and Mathematics with Applications, 54(5):651–663, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700243X.

He:2008:NNP

[HL08a] Anqi He and Jianzhou Liu. A new non-parameter criterion for H-matrices. Computers and Mathematics with Applications, 55(6): 1148–1153, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107005810.

Huang:2008:SML

[HL08b] Kuo-Nan Huang and Jui-Jung Liao. A simple method to locate the optimal solution for exponentially deteriorating items under trade credit financing. Computers and Mathematics with Applications, 56(4):965–977, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000990.

Huang:2008:RFB

[HL08c] Ting-Zhu Huang and Liang Li. Relaxed forms of BBK algorithm and FBP algorithm for symmetric indefinite linear systems. Computers and Mathematics with Applications, 55(4): 801–807, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210700510X.

He:2009:NAM

[HL09a] J.-H. He and E. W. M.

Lee. New analytical methods for cleaning up the solution of nonlinear equations. Computers and Mathematics with Applications, 58(11–12):2081–2083, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001291.

Hu:2009:SMP

[HL09b] Cheng-Feng Hu and Fung-Bao Liu. Solving mathematical programs with fuzzy equilibrium constraints. Computers and Mathematics with Applications, 58(9): 1844–1851, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109005732.

Huang:2000:SSA

[HLC00] Tetz C. Huang, Ji-Cherng Lin, and Hsueh-Jen Chen. A self-stabilizing algorithm which finds a 2-center of a tree. Computers and Mathematics with Applications, 40(4–5):607–624, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001838.

Hu:2005:RBC

[HLC05] Hsin-Yun Hu, Zi-Cai Li, and Alexander H.-D. Cheng. Radial basis collocation methods for elliptic boundary value problems. Computers and Mathematics with Applications, 50(1–2):289– 320, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002518.

He:2000:EPF

[HLCL00] Qing He, Hong-Xing Li, C. L. P. Chen, and E. S. Lee. Extension principles and fuzzy set categories. Computers and Mathematics with Applications, 39(1–2):45–53, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003120.

Huang:2007:SSA

[HLCW07] Tetz C. Huang, Ji-Cherng Lin, Chih-Yuan Chen, and Cheng-Pin Wang. A self-stabilizing algorithm for finding a minimal 2-dominating set assuming the distributed demon model. Computers and Mathematics with Applications, 54(3):350–356, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002180.

Huang:2004:SSA

[HLM04] Tetz C. Huang, Ji-Cherng Lin, and Nathan Mou. A self-stabilizing algorithm for the center-finding problem assuming read/write separate atomicity. Computers and Mathematics with Applications, 48(10– 11):1667–1676, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104004201.

Hron:2008:FIF

[HLMR08] J. Hron, C. Le Roux, J. Málek, and K. R. Rajagopal. Flows of incompressible fluids subject to Navier’s slip on the boundary. Computers and Mathematics with Applications, 56(8):2128–2143, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003489.

Hyman:2003:AMM

[HLP03] J. M. Hyman, Shengtai Li, and L. R. Petzold. An adaptive moving mesh method with static rezoning for partial differential equations. Computers and Mathematics with Applications, 46(10– 11):1511–1524, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901878.

He:2005:SAI

[HLP05] Suyan He, Xingsi Li, and Shaohua Pan. A selfadjusting interior point algorithm for linear complementarity problems. Computers and Mathematics with Applications, 50(1–2): 33–40, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002294.

Hong:2005:SBS

[HLS05] Jialin Hong, Hongyu Liu, and Geng Sun. Spurious behavior of a symplectic integrator. Computers and Mathematics with Applications, 50(3–4):519– 528, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002932.

He:2003:FCM

[HLSL03] Qing He, Hong-Xing Li, Zhong-Zhi Shi, and E. S. Lee. Fuzzy clustering method based on perturbation. Computers and Mathematics with Applications, 46(5–6):929–946, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901544.

Han:2007:SSA

[HLW07] Deren Han, Hong K. Lo, and Zhiwei Wang. A simple self-adaptive alternating direction method for linear variational inequality problems. Computers and Mathematics with Applications, 53(10):1595– 1604, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001320.

Huang:2004:CM

[HLWT04] Ting-Zhu Huang, Jin-Song Leng, E. L. Wachspress, and Yuan Yan Tang. Characterization of H-matrices. Computers and Mathematics with Applications, 48(10– 11):1587–1601, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004122.

Hu:2007:GAS

[HLX07] Lin-Xia Hu, Wan-Tong Li, and Hong-Wu Xu. Global asymptotical stability of a second order rational difference equation. Computers and Mathematics with Applications, 54(9–10):1260–1266, November 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003343.

Hwang:2005:ABN

[HLY05] Tsang-Hwai Hwang, Horng-Jaan Li, and Cheh-Chih Yeh. Asymptotic behavior of nonoscillatory solutions of second-order differential equations. Computers and Mathematics with Applications, 50(1–2):271– 280, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500249X.

Hu:2006:EMP

[HLY06] Xlaoling Hu, Guirong Liu, and Jurang Yan. Existence of multiple positive periodic solutions of delayed predator–prey models with functional responses. Computers and Mathematics with Applications, 52(10– 11):1453–1462, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003324.

Han:2008:PAN

[HLY08] Deren Han, Hong K. Lo, and Hai Yang. On the price of anarchy for non-atomic congestion games under asymmetric cost maps and elastic demands. Computers and Mathematics with Applications, 56(10):2737–2743, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800415X.

Hu:2009:ICF

[HLY09] Dan Hu, Hongxing Li, and Xianchuan Yu. The information content of fuzzy relations and fuzzy rules. Computers and Mathematics with Applications, 57(2):202–216, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006482.

Huang:2009:DAC

[HLZ09] Tingwen Huang, Chuandong Li, and Zhigang Zeng. A domain attraction criterion for interval fuzzy neural networks. Computers and Mathematics with Applications, 58(3):508–513, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003265.

Hong:2000:OOA

[HM00] Don Hong and R. N. Mohapatra. Optimal-order approximation by mixed threedirectional spline elements. Computers and Mathematics with Applications, 40(1):

127–135, July 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100001462.

Huy:2001:EDD

[HM01] Nguyen Thieu Huy and Nguyen Van Minh. Exponential dichotomy of difference equations and applications to evolution equations on the halfline. Computers and Mathematics with Applications, 42(3–5):301–311, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001559.

Harko:2003:RDC

[HM03a] T. Harko and M. K.

Mak. Relativistic dissipative cosmological models and Abel differential equation. Computers and Mathematics with Applications, 46(5–6):849–853, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901477.

Henrion:2003:OCD

[HM03b] R. Henrion and A. Möller.

Optimization of a continuous distillation process under random inflow rate. Computers and Mathematics with Applications, 45 (1–3):247–262, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800172.

Higueras:2004:DAE

[HM04] I. Higueras and R. März. Differential algebraic equations with properly stated leading terms. Computers and Mathematics with Applications, 48(1–2):215– 235, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840112.

Hayat:2005:ROP

[HM05a] T. Hayat and S. Mumtaz. Resonant oscillations of a plate in an electrically conducting rotating Johnson– Segalman fluid. Computers and Mathematics with Applications, 50(10– 12):1669–1676, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004177.

Hernandez:2005:SCE

[HM05b] E. Hernández and M. A.

McKibben. Some comments on: Existence of solutions of abstract nonlinear second-order neutral functional integrodifferential equations. Computers and Mathematics with Applications, 50(5–6):655–669, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003068.

See [BPA03].

Hasanov:2000:NSN

[HMCR00] A. Hasanov, J. L. Mueller, S. Cohn, and J. Redepenning. Numerical solution of a nonlocal identification problem for nonlinear ion transport. Computers and Mathematics with Applications, 39(7–8):225– 235, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000078X.

Hajji:2004:RDO

[HMV04] M. A. Hajji, S. Melkonian, and R. Vaillancourt. Representation of differential operators in wavelet basis. Computers and Mathematics with Applications, 47(6–7):1011–1033, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900831.

Huang:2007:ECB

[HN07] Zhenyu Huang and Muhammad Aslam Noor. Equivalency of convergence between one-step iteration algorithm and two-step iteration algorithm of variational inclusions for H-monotone mappings. Computers and Mathematics with Applications, 53(10):1567–1571, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001277.

Hariri:2008:NHD

[HNR08] Arash Hariri, Keivan Navi, and Reza Rastegar. A new high dynamic range moduli set with efficient reverse converter. Computers and Mathematics with Applications, 55(4):660–668, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107004993.

Hayat:2009:ETR

[HNSA09] T. Hayat, M. Nawaz, M. Sajid, and S. Asghar. The effect of thermal radiation on the flow of a second grade fluid. Computers and Mathematics with Applications, 58(2):369– 379, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002314.

Hong:2000:NSS

[HO00] Woopyo Hong and Kwang-Sik Oh. New solitonic solutions to a (3 + 1)- dimensional Jimbo–Miwa equation. Computers and Mathematics with Applications, 39(5–6):29–31, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000432.

Hoffacker:2003:GFE

[Hof03]  J. Hoffacker.  Green’s

functions and eigenvalue comparisons for a focal problem on time scales. Computers and Mathematics with Applications, 45 (6–9):1339–1368, March/ May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001081.

Hohjo:2001:CIM

[Hoh01] H. Hohjo. A competitive inventory model with the customer’s general choice probability. Computers and Mathematics with Applications, 41(3–4):523–530, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002947.

Hon:2002:QRB

[Hon02] Y. C. Hon. A quasi-radial basis functions method for American options pricing. Computers and Mathematics with Applications, 43 (3–5):513–524, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101003029.

Horvath:2001:SIP

[Hor01] R. Horváth. Some integral properties of the heat equation. Computers and Mathematics with Applications, 42 (8–9):1135–1141, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002279.

Horvath:2008:SCD

[Hor08] Róbert Horváth. Sufficient conditions of the discrete maximum-minimum principle for parabolic problems on rectangular meshes. Computers and Mathematics with Applications, 55 (10):2306–2317, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107007341.

Hou:2002:ICV

[HP02] Shui-Hung Hou and Wan-Kai Pang. Inversion of confluent Vandermonde matrices. Computers and Mathematics with Applications, 43(12):1539–1547, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001177.

Hernandez:2006:ERI

[HPG06] E. Hernández, M. Pierri, and G. Goncalves. Existence results for an impulsive abstract partial differential equation with statedependent delay. Computers and Mathematics with Applications, 52(3–4):411– 420, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002070.

Hernandez:2005:CPS

[HPT05] E. Hernández, M. Pierri, and P. Táboas. A comment on the papers “A study on controllability of semilinear integrodifferential systems in Banach spaces” and “Controllability of neutral functional integrodifferential systems in Banach spaces”. Computers and Mathematics with Applications, 50 (8–9):1291–1292, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003809.

See [BSD00, SC04].

Hernandez:2002:SCS

[HR02] M. A. Hernández and M. J. Rubio. Semilocal convergence of the secant method under mild convergence conditions of differentiability. Computers and Mathematics with Applications, 44(3–4):277–285, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001475.

Higueras:2005:SAC

[HR05] I. Higueras and T. Roldán. Starting algorithms for a class of RK methods for index-2 DAEs. Computers and Mathematics with Applications, 49(7–8):1081–1099, May/ June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001240.

Haji:2007:ASS

[HR07] Abdukerim Haji and Agnes Radl. Asymptotic stability of the solution of the M/M^{B}/1 queueing model. Computers and Mathematics with Applications, 53(9):1411–1420, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700106X.

He:2008:TFR

[HR08] Matthew X. He and Paolo E.

Ricci. On Taylor’s formula for the resolvent of a complex matrix. Computers and Mathematics with Applications, 56(9):2285–2288, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003623.

Hosseinnia:2008:UEH

[HRM08] S. H. Hosseinnia, A. Ranjbar, and S. Momani. Using an enhanced homotopy perturbation method in fractional differential equations via deforming the linear part. Computers and Mathematics with Applications, 56(12):3138–3149, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004343.

Hernandez:2009:SMR

[HRMS09] Elvira Hernández, Luis Rodríguez-Marín, and Miguel Sama. Scalar multiplier rules in set-valued optimization. Computers and Mathematics with Applications, 57(8):1286–1293, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000236.

Ha:2001:NSO

[HRP01] S. N. Ha, S. W. Roh, and J. Park. Numerical study of an optimal control problem for a catalytic converter function. Computers and Mathematics with Applications, 41(7–8):893– 902, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003278.

Heath:2004:PES

[HRP04] L. S. Heath, C. J. Ribbens, and S. V. Pemmaraju. Processor-efficient sparse matrix–vector multiplication. Computers and Mathematics with Applications, 48(3–4):589–608, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840793.

Hu:2008:RCH

[HRS08] Zhaoping Hu, Valery G. Romanovski, and Douglas S. Shafer. 1: −3 resonant centers on C^{2} with homogeneous cubic nonlinearities. Computers and Mathematics with Applications, 56(8):1927–1940, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002708.

Hettich:2000:ABF

[HS00] U. Hettich and R. L. Stens.

Approximating a bandlimited function in terms of its samples. Computers and Mathematics with Applications, 40(1):107– 116, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001449.

Hay:2002:CMD

[HS02] A. M. Hay and J. A. Snyman. The chord method for the determination of nonconvex workspaces of planar parallel manipulators. Computers and Mathematics with Applications, 43(8–9):1135–1151, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800189.

Han:2004:NMG

[HS04a] Derek Han and Wenyu Sun. A new modified Goldstein-Levitin–Polyak projection method for variational inequality problems. Computers and Mathematics with Applications, 47(12):1817– 1825, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840197.

Hyman:2004:CMD

[HS04b] J. M. Hyman and S. Steinberg. The convergence of mimetic discretization for rough grids. Computers and Mathematics with Applications, 47(10–11):1565–1610, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839737.

Hasanov:2007:DFA

[HS07] Anvar Hasanov and H. M.

Srivastava. Decomposition formulas associated with the Lauricella multivariable hypergeometric functions. Computers and Mathematics with Applications, 53(7):1119–1128, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000715.

Huang:2009:HNI

[HS09] Jianhua Huang and Jinliang Shen. Hybrid Noor iteration method for nonexpansive mappings in Hilbert spaces. Computers and Mathematics with Applications, 58(9):1755–1761, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900604X.

He:2006:CCE

[HSG^{+}06] Fei He, Xiaoyu Song, Ming Gu, Lerong Cheng, Guowu Yang, Zhiwei Tang, and Jiaguang Sun. A combinatorial congestion estimation approach with generalized detours. Computers and Mathematics with Applications, 51(6–7):1113–1126, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000411.

Hay:2000:DNW

[HSM00] A. M. Hay, J. A. Snyman, and Multidisciplinary Design Optimization Group (MDOG). The determination of nonconvex workspaces of generally constrained planar Stewart platforms. Computers and Mathematics with Applications, 40(8– 9):1043–1060, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850154.

Hyman:2001:EIP

[HSS01] J. Hyman, M. Shashkov, and S. Steinberg. The effect of inner products for discrete vector fields on the accuracy of mimetic finite difference methods. Computers and Mathematics with Applications, 42(12):1527– 1547, December 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101002607.

Hu:2008:NCP

[HSWL08] Minghan Hu, Xiquan Shi, Tianjun Wang, and Fengshan Liu. A note on cubic polynomial interpolation. Computers and Mathematics with Applications, 56(5):1358–1363, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001594.

Henderson:2002:EMS

[HT02] J. Henderson and H. B.

Thompson. Existence of multiple solutions for second-order discrete boundary value problems. Computers and Mathematics with Applications, 43(10– 11):1239–1248, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102000950.

Hassenpflug:2004:CNE

[HT04] W. C. Hassenpflug and G. D. Thiart. Convolution number evaluation of wavemaking velocity. Computers and Mathematics with Applications, 48(5–6):665–689, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003037.

Hoffacker:2005:SID

[HT05] J. Hoffacker and C. C. Tisdell. Stability and instability for dynamic equations on time scales. Computers and Mathematics with Applications, 49(9–10):1327– 1334, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001525.

Han:2009:SUM

[HTKM09] Yi Han, Jiafu Tang, Iko Kaku, and Lifeng Mu. Solving uncapacitated multilevel lot-sizing problems using a particle swarm optimization with flexible inertial weight. Computers and Mathematics with Applications, 57(11–12):1748– 1755, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005506.

He:2003:PPM

[HTZY03] Yong He, Zhiyi Tan, Jing Zhu, and Enyu Yao. κ-Partitioning problems for maximizing the minimum load. Computers and Mathematics with Applications, 46 (10–11):1671–1681, November/December 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210390201X.

Hu:2008:SBE

[Hu08]  Shengbiao Hu.  Sharp

bounds on the eigenvalues of trees. Computers and Mathematics with Applications, 56(4):909–917, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800093X.

Huang:2000:IPE

[Hua00] Zhenyu Huang. Iterative process with errors for fixed points of multivalued Φhemicontractive operators in uniformly smooth Banach spaces. Computers and Mathematics with Applications, 39(3–4):137–145, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003405.

Huang:2001:NCG

[Hua01]  Nan-Jing Huang.  A

new class of generalized set-valued implicit variational inclusions in Banach spaces with an application. Computers and Mathematics with Applications, 41(7–8):937–943, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000331X.

Huang:2004:CBN

[Hua04] Zhengda Huang. The convergence ball of Newton’s method and the uniqueness ball of equations under Hölder-type continuous derivatives. Computers and Mathematics with Applications, 47(2–3):247–251, January/February 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104900211.

Huang:2005:SSA

[Hua05] Tetz C. Huang. A selfstabilizing algorithm for the shortest path problem assuming the distributed demon. Computers and Mathematics with Applications, 50(5–6):671–681, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500307X.

Huang:2006:ETM

[Hua06] Po-Hsian Huang. The extenics theory for a matching evaluation system. Computers and Mathematics with Applications, 52(6– 7):997–1010, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002732.

Huang:2008:SLM

[Hua08a] Chengming Huang. Stability of linear multistep methods for delay integro-differential equations. Computers and Mathematics with Applications, 55(12):2830–2838, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700781X.

Huang:2008:ESS

[Hua08b]  Ting-Zhu Huang.  Esti-

mation of ||A^{−1}||_{∞} and the smallest singular value. Computers and Mathematics with Applications, 55 (6):1075–1080, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107005391.

Huang:2008:RCF

[Hua08c]  Xiaoxia Huang.  Risk

curve and fuzzy portfolio selection. Computers and Mathematics with Applications, 55(6):1102–1112, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005755.

Huo:2004:EPP

[Huo04] Hai-Feng Huo. Existence of positive periodic solutions of a neutral delay Lotka–Volterra system with impulses. Computers and Mathematics with Applications, 48(12):1833–1846, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104004432.

Huo:2005:PGA

[Huo05] Hai-Feng Huo. Permanence and global attractivity of delay diffusive prey–predator systems with the Michaelis– Menten functional response. Computers and Mathematics with Applications, 49 (2–3):407–416, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000428.

Hurd:2000:MLG

[Hur00] W. J. Hurd. Maximum likelihood global positioning system receiver. Computers and Mathematics with Applications, 39(11):121– 125, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001176.

Hong:2000:OMM

[HW00] Don Hong and Ai-Di Wu.

Orthogonal multiwavelets of multiplicity four. Computers and Mathematics with Applications, 40(10– 11):1153–1169, November/ December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002297.

Han:2005:GAB

[HW05a] Houde Han and Xin Wen.

The global artificial boundary conditions for numerical simulations of the flow around a submerged body. Computers and Mathematics with Applications, 50 (8–9):1499–1512, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003950.

Henderson:2005:NEP

[HW05b] J. Henderson and Haiyan Wang. Nonlinear eigenvalue problems for quasilinear systems. Computers and Mathematics with Applications, 49(11–12):1941– 1949, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002105.

Hwang:2005:ESV

[HW05c] Tsung-Min Hwang and Weichung Wang. Energy states of vertically aligned quantum dot array with nonparabolic effective mass. Computers and Mathematics with Applications, 49(1):39–51, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000052.

He:2007:VIMb

[HW07] Ji-Huan He and Xu-Hong Wu. Variational iteration method: New development and applications. Computers and Mathematics with Applications, 54(7–8):881– 894, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005494.

Hsu:2008:EFP

[HW08] Hui-Li Hsu and Berlin Wu. Evaluating forecasting performance for interval data. Computers and Mathematics with Applications, 56(9):2155–2163, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003507.

He:2009:EMS

[HW09] Xiumei He and Xian Wu.

Existence and multiplicity of solutions for nonlinear second order difference boundary value problems. Computers and Mathematics with Applications, 57(1): 1–8, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005828.

Huang:2004:ICN

[HWC04] Cheng-I. Huang, Yue-Li Wang, and Sheng-Shiung Chung. The incidence coloring numbers of meshes. Computers and Mathematics with Applications, 48 (10–11):1643–1649, November/December 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004171.

Huang:2008:PIG

[HWHG08] Zhaohui Huang, Chi Wang, Youqiu Hu, and Xiaocheng Guo. Parallel implementation of 3D global MHD simulations for Earth’s magnetosphere. Computers and Mathematics with Applications, 55(6):1094–1101, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005743.

He:2007:VIMa

[HWX07] Ji-Huan He, Abdul-Majid Wazwaz, and Lan Xu. The variational iteration method: Reliable, efficient, and promising. Computers and Mathematics with Applications, 54(7–8):879– 880, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002970.

Huang:2003:GDD

[HX03] Ting-Zhu Huang and Cheng-Xian Xu. Generalized αdiagonal dominance. Computers and Mathematics with Applications, 45(10– 11):1721–1727, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0898122103001500.

Huang:2006:BNE

[HX06] Jia Huang and Jun-Ming Xu. The bondage numbers of extended de Bruijn and Kautz digraphs. Computers and Mathematics with Applications, 51(6–7):1137– 1147, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106000435.

Huang:2007:TDT

[HX07] Jia Huang and Jun-Ming Xu. The total domination and total bondage numbers of extended de Bruijn and Kautz digraphs. Computers and Mathematics with Applications, 53(8):1206– 1213, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000788.

He:2009:NSI

[HX09] Ji-Huan He and Fei Xu.

Nonlinearity as a sensitive informative marker in the ENSO model. Computers and Mathematics with Applications, 58(11–12):2431– 2443, December 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109001862.

Ho:2003:SCM

[HXG03] S. L. Ho, M. Xie, and T. N. Goh. A study of the connectionist models for software reliability prediction. Computers and Mathematics with Applications, 46(7):1037–1045, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901179.

Hsu:2000:NRA

[HY00] S. J. Hsu and M. C. Yuang. A network reduction axion for efficient computation of terminal-pair reliability. Computers and Mathematics with Applications, 40(2–3):359–372, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001668.

Henderson:2003:EST

[HY03] J. Henderson and W. K. C. Yin. Existence of solutions for third-order boundary value problems on a time scale. Computers and Mathematics with Applications, 45(6–9):1101–1111, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000919.

Huang:2006:NAO

[HY06] Jia-Yen Huang and Ming-Jong Yao. A new algorithm for optimally determining lot-sizing policies for a deteriorating item in an integrated productioninventory system. Computers and Mathematics with Applications, 51(1):83–104, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004487.

Huang:2008:CMS

[HY08] Jia-Yen Huang and Ming-Jong Yao. On the coordination of maintenance scheduling for transportation fleets of many branches of a logistic service provider. Computers and Mathematics with Applications, 56(5):1303–1313, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001521.

Herbert:2009:CCR

[HY09] Joseph P. Herbert and JingTao Yao. Criteria for choosing a rough set model. Computers and Mathematics with Applications, 57(6):908–918, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005786.

Yuan:2009:CNT

[hYL09] Xue hai Yuan and E. Stanley Lee. Comments on “Notes on triple I method of fuzzy reasoning”. Computers and Mathematics with Applications, 58(8):1604– 1607, October 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109005021. See

[SFL02, Liu09a].

Yuan:2009:TNC

[hYLL09] Xue hai Yuan, Hongxing Li, and E. Stanley Lee. Three new cut sets of fuzzy sets and new theories of fuzzy sets. Computers and Mathematics with Applications, 57(5):691–701, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800672X.

Hong:2000:EPS

[HYLW00] Chen-Huang Hong, Cheh-Chih Yeh, Chung-Fen Lee, and Fu-Hsiang Wong. Existence of positive solutions for functional differential equations. Computers and Mathematics with Applications, 40(6– 7):783–792, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001966.

Hou:2002:BPD

[HYZ02] Xiaorong Hou, Rui Yan, and Weinian Zhang. Bifurcations of a polynomial differential system of degree n in biochemical reactions. Computers and Mathematics with Applications, 43(10–11):1407–1423, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001086.

Han:2006:EOA

[HYZ06] Boshun Han, Yirong Yao, and Quan Zheng. Existence and optimality of accessible and approximatable minimizers of quasi upper robust functions. Computers and Mathematics with Applications, 52(1–2):65– 80, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001829.

Hilscher:2003:NDQ

[HZ03] R. Hilscher and V. Zeidan. Nonnegativity of a discrete quadratic functional in terms of the (strengthened) Legendre and Jacobi conditions. Computers and Mathematics with Applications, 45(6–9):1369–1383, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001093.

Hui:2005:GSP

[HZ05] Jing Hui and Deming Zhu. Global stability and periodicity on SIS epidemic models with backward bifurcation. Computers and Mathematics with Applications, 50 (8–9):1271–1290, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003792.

Huang:2007:PAM

[HZ07] Y. S. Huang and Y. Y. Zhou. Penalty approximation method for a class of elliptic variational inequality problems. Computers and Mathematics with Applications, 53(11):1665– 1671, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001393.

He:2008:MSB

[HZ08a] Xiaoming He and Wenming Zou. Multiple solutions for the Brezis–Nirenberg problem with a Hardy potential and singular coefficients. Computers and Mathematics with Applications, 56(4):1025–1031, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001041.

Hu:2008:FPP

[HZ08b] Daowei Hu and Zhengqiu Zhang. Four positive periodic solutions of a discrete time delayed predator–prey system with nonmonotonic functional response and harvesting. Computers and Mathematics with Applications, 56(12):3015–3022, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004781.

Huang:2009:GSM

[HZ09a] Ting-Zhu Huang and Li-Tao Zhang. GRPM-style methods and comparisons of convergent and divergent rates. Computers and Mathematics with Applications, 57(4):550–559, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006743.

Huang:2009:ESC

[HZ09b] Yisheng Huang and Yuying Zhou. Existence of solutions for a class of hemivariational inequality problems. Computers and Mathematics with Applications, 57(9):1456– 1462, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000637.

Hon:2002:TGE

[HZZ02] Yiu-Chung Hon, Jian Zhang, and Xu Zhou. Threshold of global existence for 3-D attractive nonlinear Schrödinger equations under confined potentials. Computers and Mathematics with Applications, 44(3–4):317–322, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001505.

Iacoviello:2007:APF

[Iac07] Daniela Iacoviello. Analysis of pupil fluctuations after a light stimulus by image processing and neural network. Computers and Mathematics with Applications, 53(8):1260– 1270, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000843.

Ibrahim:2009:ESF

[Ibr09] Ibrahim Abouelfarag Ibrahim.

On the existence of solutions of functional integral equation of Urysohn type. Computers and Mathematics with Applications, 57(10):1609–1614, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005944.

Ikbal:2009:TLM

[ICM09] Md. A. Ikbal, S. Chakravarty, and P. K. Mandal. Twolayered micropolar fluid flow through stenosed artery: Effect of peripheral layer thickness. Computers and Mathematics with Applications, 58(7):1328–1339, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004519.

Iwaniec:2005:RGE

[ID05] T. Iwaniec and Shusen Ding. RETRACTED: Global estimates for the Jacobians of orientation preserving mappings. Computers and Mathematics with Applications, 50(5–6):707–718, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003111.

See [ID08].

Iwaniec:2008:RNG

[ID08] T. Iwaniec and Shushen Ding. Retraction notice to “Global estimates for the Jacobians of orientation preserving mappings” [Comput. Math. Appl. 50 (2005) 707–718]. Computers and Mathematics with Applications, 56(10):2779, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004173.

See [ID05].

Itsariyawanich:2010:ECE

[ID10] Kan Itsariyawanich and Natasha Dejdumrong. Erratum to: “Conversion and evaluation for two types of parametric surface constructed by NTP bases” [Comput. Math. Appl. 49(2) (2005) 321– 329]. Computers and Mathematics with Applications, 59(1):595–601, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005884.

See [JW05].

Irishina:2008:LSE

[IDM08] Natalia Irishina, Oliver Dorn, and Miguel Moscoso. A level set evolution strategy in microwave imaging for early breast cancer detection. Computers and Mathematics with Applications, 56(3):607–618, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000527.

Iwamoto:2006:DTC

[IDOK06] Kazuki Iwamoto, Tadashi Dohi, Hiroyuki Okamura, and Naoto Kaio. Discretetime cost analysis for a telecommunication billing application with rejuvenation. Computers and Mathematics with Applications, 51(2):335–344, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004712.

Iyengar:2000:VMD

[IJB00] S. R. K. Iyengar, G. Jayaraman, and V. Balasubramanian. Variable mesh difference schemes for solving a nonlinear Schrödinger equation with a linear damping term. Computers and Mathematics with Applications, 40(12):1375–1385, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002467.

Imdad:2003:RTF

[IK03] M. Imdad and S. Kumar. Rhoades-type fixed-point theorems for a pair of nonself mappings. Computers and Mathematics with Applications, 46(5–6):919–927, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901532.

Imam:2004:SSB

[IK04] T. Imam and M. Kaykobad.

Symbolic substitution-based canonical recoding algorithms. Computers and Mathematics with Applications, 48(10–11):1541–1548, November/December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004080.

Imam:2005:NSS

[IK05] T. Imam and M. Kaykobad.

A new symbolic substitution based addition algorithm. Computers and Mathematics with Applications, 50 (8–9):1303–1310, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003822.

Ikhile:2001:CSO

[Ikh01] M. N. O. Ikhile. Coefficients for studying one-step rational schemes for IVPs in ODEs: I. Computers and Mathematics with Applications, 41(5–6):769– 781, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003199.

Ikhile:2002:CSO

[Ikh02] M. N. O. Ikhile. Coefficients for studying one-step rational schemes for IVPs in ODEs: II. Computers and Mathematics with Applications, 44(3–4):545– 557, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001694.

Ikhile:2004:CSO

[Ikh04] M. N. O. Ikhile. Coefficients for studying one-step rational schemes for IVPs in ODEs: III. Extrapolation methods. Computers and Mathematics with Applications, 47(8–9):1463–1475, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210490137X.

Ikhile:2008:RBD

[Ikh08] M. N. O. Ikhile. The Root and Bell’s disk iteration methods are of the same error propagation characteristics in the simultaneous determination of the zeros of a polynomial, Part I: Correction methods. Computers and Mathematics with Applications, 56(2):411– 430, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000266.

Imaizumi:2006:RAD

[IKY06] Mitsuhiro Imaizumi, Mitsutaka Kimura, and Kazumi Yasui. Reliability analysis of a data transmission model for multicast. Computers and Mathematics with Applications, 51(2):357–362, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004736.

Isac:2004:CPI

[IL04] G. Isac and Jinlu Li. The convergence property of Ishikawa iteration schemes in noncompact subsets of Hilbert spaces and its applications to complementarity theory. Computers and Mathematics with Applications, 47(10–11):1745–1751, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839889.

Iliescu:2004:GNM

[Ili04] T. Iliescu. Genuinely nonlinear models for convectiondominated problems. Computers and Mathematics with Applications, 48(10– 11):1677–1692, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104004213.

Izquierdo:2009:AOB

[IMLF09] Salvador Izquierdo, Paula Martínez-Lera, and Norberto Fueyo. Analysis of open boundary effects in unsteady lattice Boltzmann simulations. Computers and Mathematics with Applications, 58(5):914–921, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000960.

Izquierdo:2008:DOW

[IMPF08] Joaquín Izquierdo, Idel Montalvo, Rafael Pérez, and Vicente S. Fuertes. Design optimization of wastewater collection networks by PSO. Computers and Mathematics with Applications, 56(3):777–784, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000722.

Iovane:2007:STA

[IN07] G. Iovane and A. V. Nasedkin. Some theorems about spectrum and finite element approach for eigenvalue problems for elastic bodies with voids. Computers and Mathematics with Applications, 53(5):789–802, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000326.

Inchan:2009:SCT

[Inc09] Issara Inchan. Strong convergence theorems for a new iterative method of kstrictly pseudo-contractive mappings in Hilbert spaces. Computers and Mathematics with Applications, 58(7): 1397–1407, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109004714.

Indlekofer:2002:NTP

[Ind02] K.-H Indlekofer. Number theory-probabilistic, heuristic, and computational approaches. Computers and Mathematics with Applications, 43(8–9):1035–1061, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800128.

Ishak:2008:MMF

[INP08] Anuar Ishak, Roslinda Nazar, and Ioan Pop. Magnetohydrodynamic (MHD) flow of a micropolar fluid towards a stagnation point on a vertical surface. Computers and Mathematics with Applications, 56(12):3188– 3194, December 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108004914.

Inchan:2007:ASS

[IP07] Issara Inchan and Somyot Plubtieng. Approximating solutions for the systems of strongly accretive operator equations. Computers and Mathematics with Applications, 53(9):1317– 1324, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000934.

Ibrahim:2008:CRT

[IR08] Adel K. Ibrahim and Medhat A. Rakha. Contiguous relations and their computations for _{2}F_{1} hypergeometric series. Computers and Mathematics with Applications, 56(8):1918–1926, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002691.

Ilicasu:2004:HOF

[IS04] F. O. Ilicasu and D. H.

Schultz. High-order finitedifference techniques for linear singular perturbation boundary value problems. Computers and Mathematics with Applications, 47(2–3):391–417, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900338.

Iga:2006:CAA

[IS06] Tatsunari Iga and Takashi Suzuki. Clustering applied to the acoustic source localization. Computers and Mathematics with Applications, 52(5):671–676, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002446.

Imani:2009:CSP

[ISAAM09] N. Imani, H. Sarbazi-Azad, S. G. Akl, and P. Moinzadeh. Chromatic sets of power graphs and their application to resource placement in multicomputer networks. Computers and Mathematics with Applications, 58(3):403–413, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003198.

Ishida:2000:SSS

[Ish00] Y. Ishida. Symbol system for symmetry of physical systems. Computers and Mathematics with Applications, 39(9–10):221– 240, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000973.

Ilinski:2000:MME

[ISK00] A. S. Ilinski, A. B. Samokhin, and U. U. Kapustin. Mathematical modelling of 2D electromagnetic scattering. Computers and Mathematics with Applications, 40(12):1363–1373, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002455.

Iskander:2002:CFN

[Isk02] M. G. Iskander. Comparison of fuzzy numbers using possibility programming: Comments and new concepts. Computers and Mathematics with Applications, 43(6–7):833–840, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003248.

Iskander:2004:PPA

[Isk04] M. G. Iskander. A possibility programming approach for stochastic fuzzy multiobjective linear fractional programs. Computers and Mathematics with Applications, 48(10– 11):1603–1609, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004134.

Iskander:2008:CCB

[Isk08] Maged George Iskander.

A computational comparison between two evaluation criteria in fuzzy multiobjective linear programs using possibility programming. Computers and Mathematics with Applications, 55(11):2506–2511, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007511.

Inoue:2008:SMA

[ITM08] Yasuhiro Inoue, Shu Takagi, and Yoichiro Matsumoto. A simulation model for amphiphilic molecules in a mesoscale solvent. Computers and Mathematics with Applications, 55(7):1469– 1480, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006384.

Iglberger:2008:SMP

[ITR08] Klaus Iglberger, Nils Thürey, and Ulrich Rüde. Simulation of moving particles in 3D with the lattice Boltzmann method. Computers and Mathematics with Applications, 55(7):1461– 1468, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006256.

Itoh:2007:SNS

[ITT07] Shigeharu Itoh, Naoto Tanaka, and Atusi Tani. Steady Navier–Stokes equations in a domain with piecewise smooth boundary. Computers and Mathematics with Applications, 53(3–4):685–691, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000545.

Ionescu:2006:RMM

[IUCR06] D. C. Ionescu, A. P. Ulmeanu, A. C. Constantinescu, and I. Rotaru. Reliability modelling of medium voltage distribution systems of nuclear power plants using generalized stochastic petri nets. Computers and Mathematics with Applications, 51(2):285–290, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004669.

Ivanov:2007:ISR

[Iva07] Ivan Ganchev Ivanov. Iterations for solving a rational Riccati equation arising in stochastic control. Computers and Mathematics with Applications, 53(6):977–988, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000521.

Indlekofer:2006:MBM

[IW06] K.-H. Indlekofer and S. Wehmeier.

Mean behaviour of multiplicative functions on additive arithmetical semigroups. Computers and Mathematics with Applications, 52(3–4):577–592, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002197.

Indlekofer:2004:MBD

[IWL04] K.-H. Indlekofer, S. Wehmeier, and L. G. Lucht. Mean behaviour and distribution properties of multiplicative functions. Computers and Mathematics with Applications, 48(12):1947, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104004523.

Inoue:2006:DSR

[IY06] Shinji Inoue and Shigeru Yamada. Discrete software reliability assessment with discretized NHPP models. Computers and Mathematics with Applications, 51(2):161–170, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0898122105004530.

Imaizumi:2003:ORN

[IYN03] M. Imaizumi, K. Yasui, and T. Nakagawa. Optimal reset number of a microprocessor system with network processing. Computers and Mathematics with Applications, 46(7):1047–1054, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901180.

Jiang:2003:EPP

[JA03] Daqing Jiang and R. P.

Agarwal. Existence of positive periodic solutions for a class of difference equations with several deviating arguments. Computers and Mathematics with Applications, 45(6–9):1303–1309, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001032.

Jaheen:2004:EBA

[Jah04] Z. F. Jaheen. Empirical Bayes analysis of record statistics based on linex and quadratic loss functions. Computers and Mathematics with Applications, 47(6–7):947–954, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900788.

Jain:2005:FCM

[Jai05] M. Jain. Finite capacity M/M/r queueing system with queue-dependent servers. Computers and Mathematics with Applications, 50(1–2):187–199, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002427.

Jaheen:2002:BPB

[JAM02] Z. F. Jaheen and B. N. Al-Matrafi. Bayesian prediction bounds from the scaled Burr type X model. Computers and Mathematics with Applications, 44(5–6):587–594, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001736.

Jankowski:2003:MNA

[Jan03]  T. Jankowski.  Mono-

tone and numerical–analytic methods for differential equations. Computers and Mathematics with Applications, 45(12):1823–1828, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900034.

Jankowski:2004:ESD

[Jan04a] T. Jankowski. Existence of solutions of differential equations with nonlinear multipoint boundary conditions. Computers and Mathematics with Applications, 47(6–7):1095–1103, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900892.

Jankowski:2004:ODE

[Jan04b]  T. Jankowski.  Ordi-

nary differential equations with nonlinear boundary conditions of antiperiodic type. Computers and Mathematics with Applications, 47(8–9):1419–1428, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901344.

Jankowski:2004:QAS

[Jan04c] T. Jankowski. Quadratic approximation of solutions for differential equations with nonlinear boundary conditions. Computers and Mathematics with Applications, 47(10–11):1619–1626, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839750.

Jankowski:2005:WRC

[Jan05]  T. Jankowski.  Wave-

form relaxation and comparison methods for multipoint boundary value problems. Computers and Mathematics with Applications, 50(5–6):753–768, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003159.

Jankowski:2007:ESB

[Jan07] Tadeusz Jankowski. Existence of solutions of boundary value problems for differential equations in which deviated arguments depend on the unknown solution. Computers and Mathematics with Applications, 54(3):357–363, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002192.

Jankowski:2009:EPS

[Jan09] Tadeusz Jankowski. Existence of positive solutions to second order fourpoint impulsive differential problems with deviating arguments. Computers and Mathematics with Applications, 58(4):805–817, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003605.

Jodar:2008:P

[JAV08] Lucas Jódar, Luis Acedo, and Rafael J. Villanueva. Preface. Computers and Mathematics with Applications, 56(3):593, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000503.

Jay:2004:PPS

[Jay04]  L. O. Jay.  Preserv-

ing Poisson structure and orthogonality in numerical integration of differential equations. Computers and Mathematics with Applications, 48(1–2):237– 255, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840124.

Jang:2005:NPM

[JB05] S. R.-J. Jang and J. Baglama.

Nutrient-plankton models with nutrient recycling. Computers and Mathematics with Applications, 49 (2–3):375–387, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000404.

Jodar:2000:DNS

[JC00] L. Jódar and J. A. Sánchez Cano. Discrete numerical solution of strongly coupled mixed diffusion problems. Computers and Mathematics with Applications, 40(4–5):471–489, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001747.

Jin:2005:GKT

[JC05] Cai-Yun Jin and Cao-Zong Cheng. S-G-L-KKM theorems in L-convex space and their applications to minimax inequalities. Computers and Mathematics with Applications, 50(1–2):123– 131, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002361.

Jia:2009:MOO

[JCC^{+}09] Jie Jia, Jian Chen, Guiran Chang, Yingyou Wen, and Jingping Song. Multiobjective optimization for coverage control in wireless sensor network with adjustable sensing radius. Computers and Mathematics with Applications, 57 (11–12):1767–1775, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108005464.

Jia:2009:EEC

[JCCT09] Jie Jia, Jian Chen, Guiran Chang, and Zhenhua Tan. Energy efficient coverage control in wireless sensor networks based on multi-objective genetic algorithm. Computers and Mathematics with Applications, 57(11–12):1756– 1766, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800552X.

Jia:2009:NMR

[JCZ09] Zhigang Jia, Xuehan Cheng, and Meixiang Zhao. A new method for roots of monic quaternionic quadratic polynomial. Computers and Mathematics with Applications, 58(9):1852–1858, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005690.

Jang:2007:HPI

[JD07] Sophia R.-J. Jang and Sandra L. Diamond. A host-parasitoid interaction with Allee effects on the host. Computers and Mathematics with Applications, 53(1):89–103, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700048X.

Jiao:2003:SOD

[JDH03] Yong-Chang Jiao, Chuangyin Dang, and Yue Hao. The solution of the one-dimensional nonlinear Poisson’s equations by the decomposition method. Computers and Mathematics with Applications, 46(10– 11):1645–1656, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901994.

Jiao:2008:EDM

[JDY08] Yong-Chang Jiao, Chuangyin Dang, and Yoshitsugu Yamamoto. An extension of the decomposition method for solving nonlinear equations and its convergence. Computers and Mathematics with Applications, 55(4):760–775, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700507X.

Jeong:2004:GSV

[Jeo04]  Jae Ug Jeong.  Gen-

eralized set-valued variational inclusions and resolvent equations in Banach spaces. Computers and Mathematics with Applications, 47(8–9):1241–1247, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901186.

Jia:2008:NCO

[JEP08] Baoguo Jia, Lynn Erbe, and Allan Peterson. New comparison and oscillation theorems for second-order halflinear dynamic equations on time scales. Computers and Mathematics with Applications, 56(10):2744–2756, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004161.

Jiang:2004:FEN

[JG04] Shidong Jiang and L. Greengard. Fast evaluation of nonreflecting boundary conditions for the Schrödinger equation in one dimension. Computers and Mathematics with Applications, 47(6–7):955–966, March/ April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210490079X.

Javidi:2007:CSW

[JG07] M. Javidi and A. Golbabai. Construction of a solitary wave solution for the generalized Zakharov equation by a variational iteration method. Computers and Mathematics with Applications, 54(7–8):1003–1009, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002623.

Jing:2009:RWF

[JH09] Yan-Fei Jing and Ting-Zhu Huang. Restarted weighted full orthogonalization method for shifted linear systems. Computers and Mathematics with Applications, 57(9):1583– 1591, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006160.

Ji:2002:NMS

[Ji02] H. B. Ji. Numerical methods for static and periodic bifurcation problems. Computers and Mathematics with Applications, 44(1–2):105– 110, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001335.

Jankovic:2004:AAS

[JI04] S. Janković and D. Ilić. An analytic approximation of solutions of stochastic differential equations. Computers and Mathematics with Applications, 47(6–7):903–912, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900740.

Jia:2000:SRA

[Jia00a] Z. Jia. Some recursions on Arnoldi’s method and IOM for large non-Hermitian linear systems. Computers and Mathematics with Applications, 39(3–4):125–129, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003387.

Jiang:2000:MPS

[Jia00b] Daqing Jiang. Multiple positive solutions to singular boundary value problems for superlinear higherorder ODEs. Computers and Mathematics with Applications, 40(2–3):249–259, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001589.

Jia:2001:RRP

[Jia01a]  Zhongxiao Jia.  Residu-

als of refined projection methods for large matrix eigenproblems. Computers and Mathematics with Applications, 41(7–8):813– 820, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003217.

Jiang:2001:ULSa

[Jia01b] Daqing Jiang. Upper and lower solutions for a superlinear singular boundary value problem. Computers and Mathematics with Applications, 41(5–6):563– 569, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002996.

Jiang:2001:ULSb

[Jia01c]  Daqing Jiang.  Upper

and lower solutions method and a singular superlinear boundary value problem for the one-dimensional p-Laplacian. Computers and Mathematics with Applications, 42(6–7):927–940, September/October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101002097.

Jiang:2002:ULS

[Jia02]  Daqing Jiang.  Upper

and lower solutions method and a superlinear singular boundary value problem. Computers and Mathematics with Applications, 44(3–4):323–337, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001517.

Jia:2009:MDS

[Jia09]  Baoguo Jia.  Maximum

density for the Sierpinski carpet. Computers and Mathematics with Applications, 57(10):1615– 1621, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002259.

Jin:2007:IAN

[Jin07] Mao-Ming Jin. Iterative algorithms for a new system of nonlinear variational inclusions with (A, η)- accretive mappings in Banach spaces. Computers and Mathematics with Applications, 54(4):579–588, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002404.

Jin:2008:CSI

[Jin08] Mao-Ming Jin. Convergence and stability of iterative algorithm for a new system of (A, η)-accretive mapping inclusions in Banach spaces. Computers and Mathematics with Applications, 56(9):2305–2311, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003659.

Jin:2009:PRS

[Jin09] Meng Jin. Properties of regular systems and algorithmic improvements for regular decomposition. Computers and Mathematics with Applications, 57(7):1179– 1186, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000303.

Jankovic:2001:PSH

[JJ01] S. Janković and M. Jovanović. Perturbed stochastic hereditary differential equations with integral contractors. Computers and Mathematics with Applications, 42(6– 7):871–881, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210100205X.

Ji:2001:EMR

[JK01] Jun Ji and C. Kicey.

An efficient methods for l^{∞} regression. Computers and Mathematics with Applications, 42(10– 11):1311–1317, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002425.

Janowski:2009:AGS

[JK09] E. J. Janowski and M. R. S.

Kulenović. Attractivity and global stability for linearizable difference equations. Computers and Mathematics with Applications, 57(9):1592–1607, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006147.

Jangveladze:2009:LTB

[JKN09] Temur Jangveladze, Zurab Kiguradze, and Beny Neta. Large time behavior of solutions and finite difference scheme to a nonlinear integro-differential equation. Computers and Mathematics with Applications, 57(5):799–811, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006640.

Jeong:2002:RPS

[JKP02a] Jin-Mun Jeong, Weon-Kee Kang, and Dong-Gun Park. Regular problem for solutions of a retarded semilinear differential nonlocal equation. Computers and Mathematics with Applications, 43(6–7):869–876, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101003273.

Jeong:2002:RSN

[JKP02b] Jin-Mun Jeong, Young-Chel Kwun, and Jong-Yeoul Park. Regularity for solutions of nonlinear evolution equations with nonlinear perturbations. Computers and Mathematics with Applications, 43(10–11):1231–1237, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102000949.

Jadamba:2008:IPI

[JKR08] B. Jadamba, A. A. Khan, and F. Raciti. On the inverse problem of identifying Lamé coefficients in linear elasticity. Computers and Mathematics with Applications, 56(2):431– 443, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000278.

Jeon:2006:NAB

[JKY06] Jun-Cheol Jeon, Kee-Won Kim, and Kee-Young Yoo. A novel approach for bitserial AB^{2} multiplication in finite fields GF(2^{m}). Computers and Mathematics with Applications, 51 (6–7):1103–1112, March/ April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210600040X.

Jin:2009:SID

[JL09a] Chuhua Jin and Jiaowan Luo. Stability of an integro-differential equation. Computers and Mathematics with Applications, 57(7):1080–1088, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000224.

Jin:2009:BPC

[JL09b] Xiao-Qing Jin and Fu-Rong Lin. Block preconditioners with circulant blocks for general linear systems. Computers and Mathematics with Applications, 58(7):1309–1319, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004544.

Jun:2009:SST

[JLP09] Young Bae Jun, Kyoung Ja Lee, and Chul Hwan Park. Soft set theory applied to ideals in d-algebras. Computers and Mathematics with Applications, 57(3):367–378, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006706.

Jiang:2004:NOC

[JLT04] Jianchu Jiang, Xiaoping Li, and Xianhua Tang. New oscillation criteria for first-order delay difference equations. Computers and Mathematics with Applications, 47(12):1875– 1884, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840239.

Jin:2004:CPS

[JLW04] Xiao-Qing Jin, Siu-Long Lei, and Yi-Min Wei. Circulant preconditioners for solving differential equations with multidelays. Computers and Mathematics with Applications, 47(8–9):1429–1436, April/ May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901356.

Jun:2009:SIS

[JLZ09] Young Bae Jun, Kyoung Ja Lee, and Jianming Zhan. Soft p-ideals of soft BCIalgebras. Computers and Mathematics with Applications, 58(10):2060–2068, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900594X.

Jodar:2000:CEB

[JMR00] L. Jódar, J. L. Morera, and G. Rubio. Computable explicit bounds for the discretization error of variable stepsize multistep methods. Computers and Mathematics with Applications, 39(12): 63–77, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001309.

Jain:2007:RDP

[JN07] Rahul Jain and B. R. Nagaraj. C^{(1,1/3)}-regularity in the Dirichlet problem for Δ_{∞}. Computers and Mathematics with Applications, 53(3–4):377–394, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001927.

Jog:2004:AIV

[Jog04]  C. S. Jog.  The accu-

rate inversion of Vandermonde matrices. Computers and Mathematics with Applications, 47(6–7):921–929, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900764.

Jones:2007:ACV

[Jon07] Andrew D. Jones, Jr.

Analysis of the chemical vapor infiltration process computational experiments. Computers and Mathematics with Applications, 54(9–10):1213–1227, November 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001125.

Jodar:2001:ESM

[JP01] L. Jódar and J. Pérez. Exact solution of mixed problems for variable coefficient one-dimensional diffusion equation. Computers and Mathematics with Applications, 41(5–6):689– 696, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003138.

Jeronimo:2009:PRT

[JPS09] Gabriela Jeronimo, Daniel Perrucci, and Juan Sabia. A parametric representation of totally mixed Nash equilibria. Computers and Mathematics with Applications, 58(6):1126–1141, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004568.

Jodar:2001:ANS

[JPV01] L. Jódar, J. Pérez, and R. J. Villanueva. Analytic and numerical solution of coupled implicit semi-infinite diffusion problems. Computers and Mathematics with Applications, 41(3–4):447–459, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210000287X.

Jodar:2002:EST

[JPV02] L. Jódar, J. Pérez, and R. J. Villanueva. Explicit solution of time dependent diffusion problems in a semiinfinite medium. Computers and Mathematics with Applications, 43(1–2):157– 167, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002802.

Juvigny:2001:PDD

[JR01] X. Juvigny and J. Ryan. Preconditioners for domain decomposition methods. Computers and Mathematics with Applications, 42 (8–9):1143–1155, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002280.

Jakszto:2003:EOC

[JS03] M. Jakszto and A. Skowron.

Existence of optimal controls via continuous dependence on parameters. Computers and Mathematics with Applications, 46(10– 11):1657–1669, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902008.

John:2004:DAN

[JS04] V. John and N. Sahin.

Derivation and analysis of near wall models for channel and recirculating flows. Computers and Mathematics with Applications, 48(7–8):1135–1151, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003517.

Jana:2008:PAO

[JS08] Prasanta K. Jana and Koushik Sinha. Permutation algorithms on optical multi-trees. Computers and Mathematics with Applications, 56(10):2656–2665, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004057.

Jodar:2000:SDN

[JSF00] L. Jódar, J. A. Sánchez, and M. V. Ferrer. The stable discrete numerical solution of strongly coupled mixed partial differential systems. Computers and Mathematics with Applications, 39(1–2):169– 182, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003211.

Jodar:2008:MDI

[JSGP08] Lucas Jódar, Francisco J. Santonja, and Gilberto González-Parra. Modeling dynamics of infant obesity in the region of valencia, Spain. Computers and Mathematics with Applications, 56(3):679–689, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000588.

Jimenez:2004:IEP

[JSK^{+}04] R. D. Jiménez, L. C. Santos, N. M. Kuhl, J. C. Egaña, and R. L. Soto. An inverse eigenvalue procedure for damage detection in rods. Computers and Mathematics with Applications, 47(4–5):643–657, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900521.

Jimenez:2005:RSS

[JSKE05] R. Jiménez, L. Santos, N. Kuhl, and J. Egaña. The reconstruction of a specially structured Jacobi matrix with an application to damage detection in rods. Computers and Mathematics with Applications, 49(11–12):1815– 1823, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001987.

Jaimini:2001:IAL

[JSS01] B. B. Jaimini, N. Shrivastava, and H. M. Srivastava. The integral analogue of the Leibniz rule for fractional calculus and its applications involving functions of several variables. Computers and Mathematics with Applications, 41(1–2):149–155, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101850136.

Jian:2005:SFD

[JT05] Jin-Bao Jian and Chun-Ming Tang. An SQP feasible descent algorithm for nonlinear inequality constrained optimization without strict complementarity. Computers and Mathematics with Applications, 49(2–3):223–238, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000283.

Jiang:2007:OAB

[JT07a] Jianchu Jiang and Xianhua Tang. Oscillation and asymptotic behavior of two-dimensional difference systems. Computers and Mathematics with Applications, 54(9–10):1240–1249, November 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003264.

Jiang:2007:OCT

[JT07b] Jianchu Jiang and Xianhua Tang. Oscillation criteria for two-dimensional difference systems of first order linear difference equations. Computers and Mathematics with Applications, 54(6): 808–818, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107002775.

Jain:2009:LBF

[JTRu09] Prashant K. Jain, Adrian Tentner, and Rizwan-uddin. A lattice Boltzmann framework to simulate boiling water reactor core hydrodynamics. Computers and Mathematics with Applications, 58(5):975–986, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001023.

Juang:2003:PAP

[Jua03] Wen-Shenq Juang. A practical anonymous payment scheme for electronic commerce. Computers and Mathematics with Applications, 46(12):1787–1798, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902379.

Jumarie:2001:RTF

[Jum01]  G. Jumarie.  Random-

ization of transfer functions in control systems via computer vision with pixel noises of order n. Computers and Mathematics with Applications, 41(7–8):1063– 1076, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003400.

Jumarie:2006:MRL

[Jum06]  G. Jumarie.  Modified

Riemann–Liouville derivative and fractional Taylor series of nondifferentiable functions further results. Computers and Mathematics with Applications, 51(9–10):1367–1376, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000861.

Jun:2001:NFD

[Jun01] S. C. Jun. A note on fractional differences based on a linear combination between forward and backward differences. Computers and Mathematics with Applications, 41(3–4):373–378, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002807.

Jun:2008:SBB

[Jun08] Young Bae Jun. Soft BCK/ BCI-algebras. Computers and Mathematics with Applications, 56(5):1408–1413, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800165X.

Jun:2009:GFS

[Jun09]  Young Bae Jun.  Gen-

eralizations of (ϵ, ϵ ∨ q)- fuzzy subalgebras in BCK/ BCI-algebras. Computers and Mathematics with Applications, 58(7):1383–1390, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004866.

Jun:2010:NFI

[Jun10]  Young Bae Jun.  Note

on “(α, β)-fuzzy ideals of hemirings”. Computers and Mathematics with Applications, 59(8):2582–2586, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122110000398.

See [DSA09].

Janglajew:2001:FDO

[JV01] K. Janglajew and K. Valeev.

The factorization of the difference operator. Computers and Mathematics with Applications, 42(3–5):729–733, August/September 2001.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101001924.

Jiang:2005:CET

[JW05] Su-Rong Jiang and Guo-Jin Wang. Conversion and evaluation for two types of parametric surfaces constructed by NTP bases. Computers and Mathematics with Applications, 49(2–3):321–329, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000362.

See erratum [ID10].

Jea:2008:LCM

[JWC08] Kuen-Fang Jea, Jen-Ya Wang, and Shih-Ying Chen. A linearly convergent method for broadcast data allocation. Computers and Mathematics with Applications, 56(2):324–338, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108000187.

Jiang:2004:CSO

[JWSS04] Yingtao Jiang, Yuke Wang, Xiaoyu Song, and Y. Savaria. Computation of signal output probability for Boolean functions represented by OBDD. Computers and Mathematics with Applications, 47(12):1865–1874, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840227.

Jiang:2004:MPS

[JX04] Daqing Jiang and Xiaojie Xu. Multiple positive solutions to a class of singular boundary value problems for the onedimensional p-Laplacian. Computers and Mathematics with Applications, 47 (4–5):667–681, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900545.

Junk:2008:CLB

[JY08] Michael Junk and Zhaoxia Yang. Convergence of lattice Boltzmann methods for Stokes flows in periodic and bounded domains. Computers and Mathematics with Applications, 55(7):1481– 1491, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006244.

Junk:2009:PBC

[JY09] Michael Junk and Zhaoxia Yang. Pressure boundary condition for the lattice Boltzmann method. Computers and Mathematics with Applications, 58(5): 922–929, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109000972.

Jun:2011:NPC

[JY11] Young Bae Jun and Xibei Yang. A note on the paper “Combination of interval-valued fuzzy set and soft set” [Comput. Math. Appl. 58 (2009) 521–527]. Computers and Mathematics with Applications, 61(5):1468–1470, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122111000034.

Jiao:2002:ATT

[JYDH02] Y. C. Jiao, Y. Yamamoto, C. Dang, and Y. Hao. An after-treatment technique for improving the accuracy of Adomian’s decomposition method. Computers and Mathematics with Applications, 43(6–7):783–798, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003212.

Jia:2002:RSI

[JZ02] Zhongxiao Jia and Yong Zhang. A refined shiftand-invert Arnoldi algorithm for large unsymmetric generalized eigenproblems. Computers and Mathematics with Applications, 44 (8–9):1117–1127, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002201.

Jing:2003:PSD

[JZ03] Zeng Jing and Binggen Zhang. The positive solutions of difference equations relevant to an economic problem. Computers and Mathematics with Applications, 45(4–5):835– 842, February/March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103000440.

Jiang:2008:SWR

[JZ08] Yao-Lin Jiang and Hui Zhang. Schwarz waveform relaxation methods for parabolic equations in space-frequency domain.

Computers and Mathematics with Applications, 55 (12):2924–2939, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107008152.

Jiang:2006:NSF

[JZC06] X. Jiang, H. Zhao, and L. Cao. Numerical simulations of the flow and sound fields of a heated axisymmetric pulsating jet. Computers and Mathematics with Applications, 51(3– 4):643–660, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105005286.

Ji:2007:TRM

[JZQZ07] Ying Ji, Ke-Cun Zhang, Shao-Jian Qu, and Ying Zhou. A trust-region method by active-set strategy for general nonlinear optimization. Computers and Mathematics with Applications, 54(2):229– 241, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700209X.

Jian:2007:GSM

[JZT07] Jin-Bao Jian, You-Fang Zeng, and Chun-Ming Tang. A generalized super-memory gradient projection method of strongly sub-feasible directions with strong convergence for nonlinear inequality constrained optimization. Computers and Mathematics with Applications, 54(4):507–524, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002349.

Kim:2000:GDR

[KA00a] H. J. Kim and Y. J. Ahn.

Good degree reduction of Bézier curves using Jacobi polynomials. Computers and Mathematics with Applications, 40(10– 11):1205–1215, November/ December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002327.

Kumar:2000:TSM

[KA00b] B. Krishna Kumar and D. Arivudainambi. Transient solution of an M/M/1 queue with catastrophes. Computers and Mathematics with Applications, 40 (10–11):1233–1240, November/December 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100002340.

Kumar:2002:MGR

[KA02] B. Krishna Kumar and D. Arivudainambi. The M/G/1 retrial queue with Bernoulli schedules and general retrial times. Computers and Mathematics with Applications, 43(1–2):15–30, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210100267X.

Keskin:2007:IDD

[KA07] Tuba Keskin and Kemal Aydin. Iterative decreasing dimension algorithm. Computers and Mathematics with Applications, 53(7):1153–1158, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000922.

Kadalbajoo:2009:BSC

[KA09a] M. K. Kadalbajoo and Puneet Arora. B-spline collocation method for the singular-perturbation problem using artificial viscosity. Computers and Mathematics with Applications, 57 (4):650–663, February 2009.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004872.

Kanth:2009:VIM

[KA09b] A. S. V. Ravi Kanth and K. Aruna. Variational iteration method for twelfthorder boundary-value problems. Computers and Mathematics with Applications, 58(11–12):2360–2364, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001734.

Khusainov:2000:SCL

[KAKK00] D. Khusainov, R. P. Agarwal, N. Kosarevskaya, and A. Kojametov. Spectrum control in linear stationary systems with delay. Computers and Mathematics with Applications, 39 (3–4):39–55, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122199003326.

Kampke:2000:ONO

[Käm00] T. Kämpke. Optimal and near optimal quantization of integrable functions. Computers and Mathematics with Applications, 40(10– 11):1315–1347, November/ December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002418.

Kamran:2009:MTT

[Kam09] Tayyab Kamran. Mizoguchi– Takahashi’s type fixed point theorem. Computers and Mathematics with Applications, 57(3):507–511, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108006378.

Kansa:2002:LPW

[Kan02]  E. J. Kansa.  Local,

point-wise rotational transformations of the conservation equations into stream-wise coordinates. Computers and Mathematics with Applications, 43 (3–5):501–511, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101003017.

Kang:2004:VTI

[Kan04] Sheon-Young Kang. Volterra type integral equation method for the radial Schrödinger equation: Single channel case. Computers and Mathematics with Applications, 48 (10–11):1425–1440, November/December 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104003980.

Kannan:2007:NAV

[Kan07]  K. Kannan.  A note

on aging of a viscoelastic cylinder. Computers and Mathematics with Applications, 53(2):324–328, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001988.

Kang:2009:QEP

[Kan09] Dongsheng Kang. Quasilinear elliptic problems involving multiple critical Hardy–Sobolev exponents. Computers and Mathematics with Applications, 57(1):107–116, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006068.

Karbassi:2001:AMN

[Kar01] S. M. Karbassi. An algorithm for minimizing the norm of state feedback controllers in eigenvalue assignment. Computers and Mathematics with Applications, 41(10–11):1317–1326, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101000992.

Karoui:2003:WPA

[Kar03] A. Karoui. Wavelets: Properties and approximate solution of a second kind integral equation. Computers and Mathematics with Applications, 46(2–3):263–277, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900307.

Karbassi:2005:TOC

[Kar05] S. M. Karbassi. Timeoptimal control of disturbancerejection tracking systems for discrete-time time-delayed systems by state feedback. Computers and Mathematics with Applications, 50 (8–9):1415–1424, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003913.

Karaca:2008:NTP

[Kar08] Ilkay Yaslan Karaca. Nonlinear triple-point problems with change of sign. Computers and Mathematics with Applications, 55 (4):691–703, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107005020.

Karpuz:2009:SON

[Kar09] Basak Karpuz. Some oscillation and nonoscillation criteria for neutral delay difference equations with positive and negative coefficients. Computers and Mathematics with Applications, 57(4):633–642, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006445.

Kawanishi:2006:WTD

[Kaw06] Ken’ichi Kawanishi. Waiting time distribution of a queueing system with postservice activity. Computers and Mathematics with Applications, 51(2):209–218, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/089812210500458X.

Kaya:2004:ADM

[Kay04]  D. Kaya.  An applica-

tion of the decomposition method for the twodimensional KdV–Burgers equation. Computers and Mathematics with Applications, 48(10–11):1659–1665, November/December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004195.

Kazmi:2004:IAS

[KB04] K. R. Kazmi and M. I. Bhat. Iterative algorithm for a system of nonlinear variational-like inclusions. Computers and Mathematics with Applications, 48(12):1929–1935, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210400450X.

Kumaresan:2008:OCS

[KB08] N. Kumaresan and P. Balasubramaniam. Optimal control for stochastic nonlinear singular system using neural networks. Computers and Mathematics with Applications, 56(9):2145–2154, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003490.

Kannan:2002:FDA

[KC02] R. Kannan and S. K.

Chung. Finite difference approximate solutions for the two-dimensional Burgers’ system. Computers and Mathematics with Applications, 44(1–2):193– 200, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001402.

Keselj:2003:GUM

[KC03a] V. Keselj and N. Cercone.

A graph unification machine for NL parsing. Computers and Mathematics with Applications, 46(2–3):393–419, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900344.

Kreng:2003:BNB

[KC03b] V. B. Kreng and Chia-Hua Chang. Bayesian network based multiagent system-application in emarketplace. Computers and Mathematics with Applications, 46(2–3):429–444, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900368.

Kumar:2005:SID

[KC05a] P. Kumar and S. Chhina. A symmetric information divergence measure of the Csiszár’s f-divergence class and its bounds. Computers and Mathematics with Applications, 49(4):575–588, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000519.

Kung:2005:SPP

[KC05b] Jung-Yuan Kung and Tzung-Nan Chuang. The shortest path problem with discrete fuzzy arc lengths. Computers and Mathematics with Applications, 49 (2–3):263–270, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000313.

Klimenok:2005:AAM

[KCD05] V. I. Klimenok, S. R. Chakravarthy, and A. N. Dudin. Algorithmic analysis of a multiserver Markovian queue with primary and secondary services. Computers and Mathematics with Applications, 50 (8–9):1251–1270, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003780.

Korkmaz:2008:DQA

[KD08] Alper Korkmaz and Idris Dag. A differential quadrature algorithm for simulations of nonlinear Schrödinger equation. Computers and Mathematics with Applications, 56(9):2222–2234, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003568.

Kikianty:2008:SIO

[KDC08] Eder Kikianty, S. S. Dragomir, and P. Cerone. Sharp inequalities of Ostrowski type for convex functions defined on linear spaces and application. Computers and Mathematics with Applications, 56(9):2235–2246, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800357X.

Kobayashi:2001:ESV

[KDKS01] M. Kobayashi, G. Dupret, O. King, and H. Samukawa. Estimation of singular values of very large matrices using random sampling. Computers and Mathematics with Applications, 42(10– 11):1331–1352, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002449.

Kazanci:2009:FAP

[KDY09] O. Kazancı, B. Davvaz, and S. Yamak. Fuzzy nary polygroups related to fuzzy points. Computers and Mathematics with Applications, 58(7):1466–1474, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004878.

Kellogg:2006:SSB

[Kel06] B. Kellogg. Some simple boundary value problems with corner singularities and boundary layers. Computers and Mathematics with Applications, 51(5):783–792, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000204.

Karasz:2005:EST

[KF05a] P. Kárász and G. Farkas. Exact solution for a twotype customers retrial system. Computers and Mathematics with Applications, 49(1):95–102, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000106.

Karatson:2005:POS

[KF05b] J. Karátson and I. Faragó.

Preconditioning operators and Sobolev gradients for nonlinear elliptic problems. Computers and Mathematics with Applications, 50(7):1077–1092, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003378.

Karcanias:2006:AGC

[KFMH06] N. Karcanias, S. Fatouros, M. Mitrouli, and G. H. Halikias. Approximate greatest common divisor of many polynomials, generalised resultants, and strength of approximation. Computers and Mathematics with Applications, 51(12):1817– 1830, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001271.

Kajani:2007:CBH

[KGB07] M. Tavassoli Kajani, M. Ghasemi, and E. Babolian. Comparison between the homotopy perturbation method and the sine-cosine wavelet method for solving linear integro-differential equations. Computers and Mathematics with Applications, 54(7–8):1162–1168, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002854.

Kong:2008:NPR

[KGWL08] Zhi Kong, Liqun Gao, Lifu Wang, and Steven Li. The normal parameter reduction of soft sets and its algorithm. Computers and Mathematics with Applications, 56(12):3029–3037, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004483.

Kansa:2000:CIC

[KH00] E. J. Kansa and Y. C. Hon. Circumventing the illconditioning problem with multiquadric radial basis functions: Applications to elliptic partial differential equations. Computers and Mathematics with Applications, 39(7–8):123– 137, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000717.

Kansa:2002:P

[KH02a] Edward J. Kansa and Yiu-Chung Hon. Preface. Computers and Mathematics with Applications, 43(3–5):xiii, February/March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002851.

Kao:2002:LSR

[KH02b] Chiang Kao and Wen-Kai Hsu. Lot size-reorder point inventory model with fuzzy demands. Computers and Mathematics with Applications, 43(10–11):1291–1302, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001013.

Kao:2002:SPI

[KH02c] Chiang Kao and Wen-Kai Hsu. A single-period inventory model with fuzzy demand. Computers and Mathematics with Applications, 43(6–7):841–848, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210100325X.

Khan:2008:CTN

[KH08] Safeer Hussain Khan and Nawab Hussain. Convergence theorems for nonself asymptotically nonexpansive mappings. Computers and Mathematics with Applications, 55(11):2544– 2553, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007559.

Kuo:2009:APS

[KH09] R. J. Kuo and C. C. Huang. Application of particle swarm optimization algorithm for solving bi-level linear programming problem. Computers and Mathematics with Applications, 58(4):678–685, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002260.

Khan:2001:RAV

[Kha01] A. A. Khan. A regularization approach for variational inequalities. Computers and Mathematics with Applications, 42(1–2):65– 74, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001316.

Khajah:2005:TMT

[Kha05]  H. G. Khajah.  Tau

method treatment of a delayed negative feedback equation. Computers and Mathematics with Applications, 49(11–12):1767– 1772, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500194X.

Khan:2007:NSP

[KHA07] M. Khan, T. Hayat, and M. Ayub. Numerical study of partial slip on the MHD flow of an Oldroyd 8constant fluid. Computers and Mathematics with Applications, 53(7):1088– 1097, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000673.

Khatibzadeh:2009:OCD

[Kha09] Hadi Khatibzadeh. An oscillation criterion for a delay difference equation. Computers and Mathematics with Applications, 57(1):37– 41, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005920.

Khani:2009:SNE

[KHN09] F. Khani and S. Hamedi-Nezhad. Some new exact solutions of the (2 + 1)-dimensional variable coefficient Broer–Kaup system using the Exp-function method. Computers and Mathematics with Applications, 58(11–12):2325–2329, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001667.

Kvernadze:2000:DSF

[KHS00] G. Kvernadze, T. Hagstrom, and H. Shapiro. Detecting the singularities of a function of V_{p} class by its integrated Fourier series. Computers and Mathematics with Applications, 39(9–10): 25–43, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000845.

Kyurkchiev:2004:FCN

[KI04] N. Kyurkchiev and A. Iliev. Failure of convergence of the Newton–Weierstrass iterative method for simultaneous root finding of generalized polynomials. Computers and Mathematics with Applications, 47(2–3):441–446, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900363.

Kilgore:2004:RAI

[Kil04] T. Kilgore. Rational approximation on infinite intervals. Computers and Mathematics with Applications, 48(9):1335–1343, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003785.

Kirmaci:2008:IFG

[Kir08] Ugur S. Kirmaci. Improvement and further generalization of inequalities for differentiable mappings and applications. Computers and Mathematics with Applications, 55(3):485–493, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107004853.

Kareem:2009:LBS

[KIXF09] Waleed Abdel Kareem, Seiichiro Izawa, Ao-Kui Xiong, and Yu Fukunishi. Lattice Boltzmann simulations of homogeneous isotropic turbulence. Computers and Mathematics with Applications, 58(5):1055–1061, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001114.

Kim:2001:MII

[KK01] Gang Eun Kim and Tae Hwa Kim. Mann and Ishikawa iterations with errors for nonlipschitzian mappings in Banach spaces. Computers and Mathematics with Applications, 42(12):1565–1570, December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002620.

Kang:2002:CMN

[KK02] Kab Seok Kang and Do Young Kwak. Covolume method for new nonconforming rectangle element for the Stokes problem. Computers and Mathematics with Applications, 43(8–9):1063–1078, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210280013X.

Koyanagi:2003:OPJ

[KK03] J. Koyanagi and H. Kawai. An optimal policy for joining a queue in processing two kinds of jobs. Computers and Mathematics with Applications, 46(7):1081– 1087, October 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103901222.

Kaufmann:2004:SOS

[KK04] E. R. Kaufmann and N. Kosmatov. A secondorder singular boundary value problem. Computers and Mathematics with Applications, 47(8–9):1317– 1326, April/May 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104901253.

Kiguradze:2005:BPS

[KK05] Tariel Kiguradze and Takaši Kusano. Bounded and periodic in a strip solutions of nonlinear hyperbolic systems with two independent variables. Computers and Mathematics with Applications, 49(2–3):335–364, January/February 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105000386.

Kang:2006:AEC

[KK06a] Tong Kang and Kwang Ik Kim. A − θ approaches of an eddy current problem based on solving certain potentials. Computers and Mathematics with Applications, 51(5):769–782, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000150.

Karakashian:2006:NSI

[KK06b] O. Karakashian and Th. Katsaounis. Numerical simulation of incompressible fluid flow using locally solenoidal elements. Computers and Mathematics with Applications, 51(9–10):1551–1570, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001015.

Katayama:2006:PLT

[KK06c] Tsuyoshi Katayama and Kaori Kobayashi. A pseudoconservation law for a timelimited service polling system with structured batch Poisson arrivals. Computers and Mathematics with Applications, 51(2):171–178, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004542.

Katti:2006:NPA

[KK06d] C. P. Katti and R. Kumari.

A new parallel algorithm for Lagrange interpolation on a hypercube. Computers and Mathematics with Applications, 51(6–7):1057–1064, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000368.

Kazmi:2008:EIA

[KK08a] K. R. Kazmi and F. A.

Khan. Existence and iterative approximation of solutions of generalized mixed equilibrium problems. Computers and Mathematics with Applications, 56(5):1314–1321, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001533.

Kong:2008:RIH

[KK08b] Lingju Kong and Qingkai Kong. Right-indefinite half-linear Sturm–Liouville problems. Computers and Mathematics with Applications, 55(11):2554–2564, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007560.

Kadalbajoo:2009:IVT

[KK09] Mohan K. Kadalbajoo and Devendra Kumar. Initial value technique for singularly perturbed two point boundary value problems using an exponentially fitted finite difference scheme. Computers and Mathematics with Applications, 57(7):1147– 1156, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000273.

Krutitskii:2000:SDW

[KKM00] P. A. Krutitskii, N. Ch. Krutitskaya, and G. Yu. Malysheva. Scattering of 2-D waves by several obstacles and screens in case of Neumann boundary condition. Computers and Mathematics with Applications, 39(7–8):49– 57, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210000064X.

Kee:2004:EEL

[KKS^{+}04a] C. Kee, J. Kim, H. So, H. Jun, B. Parkinson, and W. W. Hansen. Effect of the error in lineof sight unit vector on the accuracy of GPS and pseudolite navigation system. Computers and Mathematics with Applications, 48(5–6):779–787, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003025.

Kim:2004:CDA

[KKS04b] Hyoung Seok Kim, Hong Oh Kim, and Sung Yong Shin. Collision detection algorithm of a continuous type using spherical extreme vertex diagrams. Computers and Mathematics with Applications, 47(8–9):1363–1378, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901290.

Kadota:2006:DMD

[KKY06] Yoshinobu Kadota, Masami Kurano, and Masami Yasuda. Discounted Markov decision processes with utility constraints. Computers and Mathematics with Applications, 51(2):279–284, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0898122105004657.

Kong:2002:PSB

[KKZ02] Lingju Kong, Qingkai Kong, and Binggen Zhang. Positive solutions of boundary value problems for thirdorder functional difference equations. Computers and Mathematics with Applications, 44(3–4):481–489, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001700.

Kuo:2001:NNC

[KL01] Chung-Feng Jeffrey Kuo and Ching-Jenq Lee. Neural network control of a rotating elastic manipulator. Computers and Mathematics with Applications, 42(6–7):1009–1023, September/October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002164.

Kim:2002:ENE

[KL02] W. K. Kim and K. H. Lee. The existence of Nash equilibrium in n-person games with C-concavity. Computers and Mathematics with Applications, 44 (8–9):1219–1228, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002286.

Korolyuk:2003:IPS

[KL03] V. S. Korolyuk and N. Limnios.

Increment processes and its stochastic exponential with Markov switching in Poisson approximation scheme. Computers and Mathematics with Applications, 46(7):1073–1079, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901210.

Karatson:2005:SGP

[KL05] J. Karátson and L. Lóczi.

Sobolev gradient preconditioning for the electrostatic potential equation. Computers and Mathematics with Applications, 50(7):1093–1104, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500338X.

Komargodski:2006:HTM

[KL06] Z. Komargodski and D. Levin.

Hermite type moving-least-squares approximations. Computers and Mathematics with Applications, 51 (8):1223–1232, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122106000757.

Kim:2008:TND

[KL08a] Bara Kim and Inkyu Lee. Tests for nonergodicity of denumerable continuous time Markov processes. Computers and Mathematics with Applications, 55(6):1310–1321, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006001.

Kim:2008:EMR

[KL08b] Chan-Gyun Kim and Yong-Hoon Lee. Existence and multiplicity results for nonlinear boundary value problems. Computers and Mathematics with Applications, 55(12):2870–2886, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007845.

Krafczyk:2009:MME

[KL09] Manfred Krafczyk and Li-Shi Luo. Mesoscopic methods in engineering and science. Computers and Mathematics with Applications, 58(5):819–820, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900087X.

Ke:2008:RSU

[KLHC08] Jau-Chuan Ke, Ssu-Lang Lee, Ying-Lin Hsu, and Yu-Tzu Chen. On a repairable system with an unreliable service station–Bayesian approach. Computers and Mathematics with Applications, 56(7):1668–1683, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002265.

Kang:2009:SED

[KLJ09] Yong Han Kang, Mi Jin Lee, and Il Hyo Jung. Sharp energy decay estimates for the wave equation with a local degenerate dissipation. Computers and Mathematics with Applications, 57(1):21–27, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006135.

Kulenovic:2003:DFC

[KLMR03] M. R. S. Kulenović, G. Ladas, L. F. Martins, and I. W. Rodrigues. The dynamics of χ_{n} + 1 = α + βχ_{n}A + Bχ_{n} +Cx_{n−1} facts and conjectures. Computers and Mathematics with Applications, 45(6–9):1087–1099, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000907.

Kulenovic:2001:RDE

[KLP01] M. R. S. Kulenović, G. Ladas, and N. R. Prokup. A rational difference equation. Computers and Mathematics with Applications, 41(5–6):671–678, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003114.

Kim:2009:SFC

[KLP09] Sung Hyun Kim, Chang Hee Lee, and PooGyeon Park. H_{∞} state-feedback control for fuzzy systems with input saturation via fuzzy weighting-dependent Lyapunov functions. Computers and Mathematics with Applications, 57(6):981–990, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005750.

Kim:2006:NMN

[KLS06] Sang Dong Kim, Yong Hun Lee, and Byeong Chun Shin. Newton’s method for the Navier–Stokes equations with finite-element initial guess of Stokes equations. Computers and Mathematics with Applications, 51(5):805–816, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000174.

Kong:2009:TPS

[KLW09] Debin Kong, Lishan Liu, and Yonghong Wu. Triple positive solutions of a boundary value problem for nonlinear singular secondorder differential equations of mixed type with p-Laplacian. Computers and Mathematics with Applications, 58(7):1425–1432, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004751.

Kewen:2008:HCG

[KLZ08] Zhao Kewen, Hong-Jian Lai, and Ju Zhou. Hamiltonianconnected graphs. Computers and Mathematics with Applications, 55(12):2707– 2714, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007705.

Kou:2009:CEP

[KLZZ09] Xiaoli Kou, Sanyang Liu, Jianke Zhang, and Wei Zheng. Co-evolutionary particle swarm optimization to solve constrained optimization problems. Computers and Mathematics with Applications, 57(11–12):1776– 1784, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005439.

Kobeissi:2001:CEL

[KM01] M. Kobeissi and M. Mollard. A counter-example to a lemma about embedding starlike trees into hypercube. Computers and Mathematics with Applications, 41(5–6):669–670, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003102.

Karcanias:2003:NFP

[KM03a] N. Karcanias and M. Mitrouli.

Normal factorisation of polynomials and computational issues. Computers and Mathematics with Applications, 45(1–3): 229–245, January/February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800160.

Kim:2003:DLS

[KM03b] H. J. Kim and V. Mani.

Divisible load scheduling in single-level tree networks: Optimal sequencing and arrangement in the nonblocking mode of communication. Computers and Mathematics with Applications, 46(10– 11):1611–1623, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901969.

Kadakal:2007:SLP

[KM07] M. Kadakal and O. Sh.

Mukhtarov. Sturm–Liouville problems with discontinuities at two points. Computers and Mathematics with Applications, 54(11–12):1367–1379, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003240.

Kupershtokh:2009:ESL

[KMK09] A. L. Kupershtokh, D. A. Medvedev, and D. I. Karpov. On equations of state in a lattice Boltzmann method. Computers and Mathematics with Applications, 58(5):965–974, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001011.

Kalaba:2004:DPM

[KN04] R. Kalaba and H. Natsuyama. Dynamic programming and minimal norm solutions of least squares problems. Computers and Mathematics with Applications, 48(3–4):419–427, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840653.

Kim:2008:FMG

[KN08] Hee Sik Kim and J. Neggers. Fibonacci mean and golden section mean. Computers and Mathematics with Applications, 56(1):228– 232, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000096.

Kwiatkowska:2006:MCE

[KNP06] M. Kwiatkowska, G. Norman, and A. Pacheco. Model checking expected time and expected reward formulae with random time bounds. Computers and Mathematics with Applications, 51(2):305–316, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004682.

Karpuz:2008:EOF

[KÖ08a] Basak Karpuz and Özkan Öcalan. Erratum to: “Oscillation of forced neutral differential equations with positive and negative coefficients” [Comput. Math. Appl. 54 (2007) 1411–1421]. Computers and Mathematics with Applications, 56(2):590–591, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000060.

See [Öca07].

Karpuz:2008:ESF

[KÖ08b] Basak Karpuz and Özkan Öcalan. Erratum to: “Stability for first-order delay dynamic equations on time scales” [Comput. Math. Appl. 53 (2007) 1820– 1831]. Computers and Mathematics with Applications, 56(4):1157–1158, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001569.

See [Zho08].

Karaa:2009:TLC

[KO09a] Samir Karaa and Mohamed Othman. Two-level compact implicit schemes for three-dimensional parabolic problems. Computers and Mathematics with Applications, 58(2):257–263, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002594.

Koroglu:2009:NTW

[KÖ09b] Canan Köroğlu and Turgut Özis. A novel traveling wave solution for Ostrovsky equation using Exp-function method. Computers and Mathematics with Applications, 58(11–12):2142–2146, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001394.

Kawasaki:2008:LBM

[KOCO08] A. Kawasaki, J. Onishi, Y. Chen, and H. Ohashi. A lattice Boltzmann model for contact-line motions. Computers and Mathematics with Applications, 55 (7):1492–1502, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107006360.

Kohler:2000:PFT

[Köh00]  A. E. Köhler.  Polygo-

nal figures and their form spaces II. Angle permutations in equilateral polygons. Computers and Mathematics with Applications, 40(2–3):373–403, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000167X.

Koikari:2007:EAM

[Koi07] S. Koikari. An error analysis of the modified scaling and squaring method. Computers and Mathematics with Applications, 53(8):1293– 1305, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000892.

Konuralp:2009:STD

[Kon09] Ali Konuralp. The steady temperature distributions with different types of nonlinearities. Computers and Mathematics with Applications, 58(11–12):2152–2159, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001412.

Koplatadze:2001:OLD

[Kop01] R. Koplatadze. Oscillation of linear difference equations with deviating arguments. Computers and Mathematics with Applications, 42(3–5):477–486, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001717.

Korlie:2000:DPM

[Kor00a] M. S. Korlie. 3-D particle modeling of gas bubbles in a liquid. Computers and Mathematics with Applications, 39(3–4):235–246, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812219900348X.

Korlie:2000:TDC

[Kor00b]  M. S. Korlie.  Three-

dimensional computer simulation of liquid drop evaporation. Computers and Mathematics with Applications, 39(12):43– 52, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001280.

Koren:2005:DGE

[Kor05] Y. Koren. Drawing graphs by eigenvectors: theory and practice. Computers and Mathematics with Applications, 49(11–12):1867– 1888, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500204X.

Korlie:2007:SCF

[Kor07]  Mark S. Korlie.  3D

simulation of cracks and fractures in a molecular solid under stress and compression. Computers and Mathematics with Applications, 54(5):638–650, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002428.

Koto:2001:POE

[Kot01] T. Koto. Periodic orbits in the Euler method for a class of delay differential equations. Computers and Mathematics with Applications, 42(12):1597–1608, December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002656.

Koto:2004:MLA

[Kot04] Toshiyuki Koto. Method of lines approximations of delay differential equations. Computers and Mathematics with Applications, 48(1– 2):45–59, July 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210483997X.

Kovacevic:2000:CGG

[Kov00] M. A. Kovacević. Construction of generalized Gauss– Turán quadrature rules for systems of arbitrary functions. Computers and Mathematics with Applications, 40(8–9):895–909, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850014.

Kovacs:2005:MNR

[Kov05] M. Kovács. The maximumnorm of the restricted denominator approximations. Computers and Mathematics with Applications, 50(7):1115–1123, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003408.

Koca:2008:NAN

[KOV08] Ahmet Koca, Hakan F. Oztop, and Yasin Varol. Numerical analysis of natural convection in shed roofs with eave of buildings for cold climates. Computers and Mathematics with Applications, 56(12):3165– 3174, December 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004458.

Kadalbajoo:2001:VMS

[KP01] M. K. Kadalbajoo and K. C. Patidar. Variable mesh spline approximation method for solving singularly perturbed turning point problems having boundary layer(s). Computers and Mathematics with Applications, 42(10– 11):1439–1453, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210100253X.

Kadalbajoo:2003:EFS

[KP03] M. K. Kadalbajoo and K. C. Patidar. Exponentially fitted spline in compression for the numerical solution of singular perturbation problems. Computers and Mathematics with Applications, 46(5–6):751–767, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901398.

Kalogeropoulos:2005:PDB

[KP05a] G. Kalogeropoulos and P. Psarrakos. The polar decomposition of block companion matrices. Computers and Mathematics with Applications, 50(3–4):529– 537, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002944.

Kinyon:2005:RQR

[KP05b] M. K. Kinyon and J. D. Phillips. Rectangular quasigroups and rectangular loops. Computers and Mathematics with Applications, 49(11–12):1679– 1685, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001896.

Kim:2006:PMF

[KP06] Dongho Kim and Eun-Jae Park. Primal mixed finiteelement approximation of elliptic equations with gradient nonlinearities. Computers and Mathematics with Applications, 51(5):793–804, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000162.

Karamolegkos:2009:ESC

[KPD^{+}09] Pantelis N. Karamolegkos, Charalampos Z. Patrikakis, Nikolaos D. Doulamis, Panagiotis T. Vlacheas, and Ioannis G. Nikolakopoulos. An evaluation study of clustering algorithms in the scope of user communities assessment. Computers and Mathematics with Applications, 58(8):1498–1519, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004970.

Kim:2000:MFE

[KPP00] M.-Y. Kim, E.-J. Park, and J. Park. Mixed finite element domain decomposition for nonlinear parabolic problems. Computers and Mathematics with Applications, 40 (8–9):1061–1070, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850166.

Kwon:2006:MFB

[KPSW06] Ohin Kwon, Hyunchan Pyo, Jin Keun Seo, and Eung Je Woo. Mathematical framework for B_{z}- based MREIT model in electrical impedance imaging. Computers and Mathematics with Applications, 51 (5):817–828, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106000125.

Kolm:2001:NQS

[KR01] P. Kolm and V. Rokhlin. Numerical quadratures for singular and hypersingular integrals. Computers and Mathematics with Applications, 41(3–4):327–352, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002777.

Korotov:2005:GLR

[Kr05] S. Korotov and M. K.

ríek. Global and local refinement techniques yielding nonobtuse tetrahedral partitions. Computers and Mathematics with Applications, 50(7):1105–1113, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003391.

Kim:2006:MMA

[KR06] Seong-Hwan Kim and T. G.

Robertazzi. Modeling mobile agent behavior. Computers and Mathematics with Applications, 51(6–7):951–966, March/ April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000290.

Khan:2007:IRE

[KR07a] A. A. Khan and B. D.

Rouhani. Iterative regularization for elliptic inverse problems. Computers and Mathematics with Applications, 54(6):850–860, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002933.

Kim:2007:NCE

[KR07b] Taekyun Kim and Seog-Hoon Rim. New Changhee q-Euler numbers and polynomials associated with padic q-integrals. Computers and Mathematics with Applications, 54(4):484–489, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002313.

Krause:2001:ABM

[Kra01] U. Krause. The asymptotic behavior of monotone difference equations of higher order. Computers and Mathematics with Applications, 42(3–5):647–654, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001845.

Krutitskii:2003:HIE

[Kru03a] P. A. Krutitskii. Hypersingular integral equations not needed in the impedance problem in scattering theory. Computers and Mathematics with Applications, 45(1–3):391–399, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800251.

Krutitskii:2003:IPP

[Kru03b]  P. A. Krutitskii.  The

impedance problem for the propagative Helmholtz equation in interior multiply connected domain. Computers and Mathematics with Applications, 46(10– 11):1601–1610, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901957.

Krutitskii:2005:MHP

[Kru05]  P. A. Krutitskii.  The

mixed harmonic problem in an exterior cracked domain with Dirichlet condition on cracks. Computers and Mathematics with Applications, 50(5–6):769–782, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003160.

Krutitskii:2008:DPL

[Kru08]  P. A. Krutitskii.  The

Dirichlet problem for the Laplacian with discontinuous boundary data in a 2D multiply connected exterior domain. Computers and Mathematics with Applications, 56(12):3221–3235, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004823.

Kalpakam:2000:PSM

[KS00a] S. Kalpakam and S. Shanthi. A perishable system with modified base stock policy and random supply quantity. Computers and Mathematics with Applications, 39(12): 79–89, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001310.

Kim:2000:MBV

[KS00b] Myoung-Nyoun Kim and Dongwoo Sheen. A mixed boundary value problem for TFT/LCD: Analysis and numerical methods. Computers and Mathematics with Applications, 39(3–4):107–123, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003375.

Kawai:2003:EBW

[KS03a] H. Kawai and H. Sandoh. An efficient backup warning policy for a hard disk. Computers and Mathematics with Applications, 46(7):1055–1063, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901192.

Konguetsof:2003:EFT

[KS03b] A. Konguetsof and T. E.

Simos. An exponentiallyfitted and trigonometricallyfitted method for the numerical solution of periodic initial-value problems. Computers and Mathematics with Applications, 45 (1–3):547–554, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800366.

Kuo:2007:ARM

[KS07] R. J. Kuo and C. W.

Shih. Association rule mining through the ant colony system for national health insurance research database in Taiwan. Computers and Mathematics with Applications, 54(11–12):1303–1318, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000910.

Kong:2002:NCN

[KSZ02] Qingkai Kong, Yijun Sun, and Binggen Zhang. Nonoscillation of a class of neutral differential equations. Computers and Mathematics with Applications, 44 (5–6):643–654, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102001797.

Karbassi:2002:PSF

[KT02] S. M. Karbassi and H. A. Tehrani. Parameterizations of the state feedback controllers for linear multivariable systems. Computers and Mathematics with Applications, 44 (8–9):1057–1065, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002146.

Kawai:2003:CEB

[KT03a] M. Kawai and M. Tamaki. Choosing either the best or the second best when the number of applicants is random. Computers and Mathematics with Applications, 46(7):1065–1071, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901209.

Kuk:2003:ODN

[KT03b] H. Kuk and T. Tanino.

Optimality and duality in nonsmooth multiobjective optimization involving generalized type I functions. Computers and Mathematics with Applications, 45(10–11):1497–1506, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001330.

Khater:2008:NSG

[KT08] A. H. Khater and R. S. Temsah. Numerical solutions of the generalized Kuramoto– Sivashinsky equation by Chebyshev spectral collocation methods. Computers and Mathematics with Applications, 56(6):1465–1472, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001727.

Kayasith:2009:SCI

[KT09] Prakasith Kayasith and Thanaruk Theeramunkong. Speech confusion index ϕ: a confusion-based speech quality indicator and recognition rate prediction for dysarthria. Computers and Mathematics with Applications, 58(8):1534–1549, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109005033.

Kundu:2001:FLS

[KU01] S. Kundu and V. A. Ubhaya. Fitting a least squares piecewise linear continuous curve in two dimensions. Computers and Mathematics with Applications, 41(7–8):1033– 1041, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003370.

Kuang:2008:NIW

[Kua08]  Jichang Kuang.  The

norm inequalities for the weighted Cesaro mean operators. Computers and Mathematics with Applications, 56(10):2588–2595, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003969.

Kumar:2004:IIM

[Kum04]  P. Kumar.  Inequalities

involving moments of a continuous random variable defined over a finite interval. Computers and Mathematics with Applications, 48(1–2):257– 273, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840136.

Kumar:2005:OTM

[Kum05] P. Kumar. The Ostrowski type moment integral inequalities and momentbounds for continuous random variables. Computers and Mathematics with Applications, 49(11–12):1929– 1940, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002099.

Kumar:2006:PDC

[Kum06] P. Kumar. Probability distributions conditioned by the available information: Gamma distribution and moments. Computers and Mathematics with Applications, 52(3–4):289–304, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001970.

See erratum [Ata09].

Kurics:2008:SCP

[Kur08] Tamás Kurics. On the superlinear convergence of PCG algorithms: Numerical experiments for convection–diffusion equations. Computers and Mathematics with Applications, 55(10):2318–2328, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007353.

Kraft:2003:NSS

[KW03] M. Kraft and W. Wagner. Numerical study of a stochastic particle method for homogeneous gas-phase reactions. Computers and Mathematics with Applications, 45(1–3):329–349, January/February 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800226.

Kuo:2005:AAM

[KWHC05] R. J. Kuo, H. S. Wang, Tung-Lai Hu, and S. H. Chou. Application of ant K-means on clustering analysis. Computers and Mathematics with Applications, 50(10– 12):1709–1724, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004207.

Kryazhimskii:2002:DMM

[KWT02] A. Kryazhimskii, C. Watanabe, and Y. Tou. Dynamic model of market of patents and equilibria in technology stocks. Computers and Mathematics with Applications, 44(7):979–995, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002080.

Kunisch:2005:PBF

[KX05] K. Kunisch and L. Xie.

POD-based feedback control of the Burgers equation by solving the evolutionary HJB equation. Computers and Mathematics with Applications, 49(7–8):1113–1126, May/ June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001264.

Kemp:2000:SCM

[KY00] P. D. Kemp and Y. Yavin.

Stabilization and control of the motion of a rolling disk by using two overhead rotors. Computers and Mathematics with Applications, 39(7–8):237– 252, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000791.

Kim:2002:NEE

[KY02] Won Kyu Kim and George Xian-Zhi Yuan. New existence of equilibria in generalized SSC. Computers and Mathematics with Applications, 43(1–2):55–63, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002711.

Kuo:2007:SMSa

[KY07a] Wen-Hung Kuo and Dar-Li Yang. Single-machine scheduling problems with start-time dependent processing time. Computers and Mathematics with Applications, 53(11):1658– 1664, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700137X.

Kuo:2007:SMSb

[KY07b] Wen-Hung Kuo and Dar-Li Yang. Single-machine scheduling problems with the time-dependent learning effect. Computers and Mathematics with Applications, 53(11):1733–1739, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001459.

Khalifeh:2008:HWI

[KYAA08] M. H. Khalifeh, H. Yousefi-Azari, and A. R. Ashrafi. The hyper-Wiener index of graph operations. Computers and Mathematics with Applications, 56(5):1402– 1407, September 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108001648.

Kee:2003:PCM

[KYJ03] Changdon Kee, Doohee Yun, and Haeyoung Jun. Precise calibration method of pseudolite positions in indoor navigation systems. Computers and Mathematics with Applications, 46 (10–11):1711–1724, November/December 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103902057.

Kimura:2006:RCM

[KYNI06] Mitsutaka Kimura, Kazumi Yasui, Toshio Nakagawa, and Naohiro Ishii. Reliability consideration of a mobile communication system with network congestion. Computers and Mathematics with Applications, 51(2):377–386, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0898122105004761.

Karaa:2002:CPI

[KZ02] S. Karaa and Jun Zhang. Convergence and performance of iterative methods for solving variable coefficient convection–diffusion equation with a fourthorder compact difference scheme. Computers and Mathematics with Applications, 44(3–4):457–479, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001621.

Kou:2003:VLM

[KZD03] Chunhai Kou, Shunian Zhang, and Yongrui Duan. Variational Lyapunov method and stability analysis for impulsive delay differential equations. Computers and Mathematics with Applications, 46(12):1761–1777, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902355.

Li:2000:IOC

[LA00] Wan-Tong Li and R. P. Agarwal. Interval oscillation criteria for secondorder nonlinear differential equations with damping. Computers and Mathematics with Applications, 40(2–3):217–230, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001553.

Li:2003:PSH

[LA03] Wan-Tong Li and R. P.

Agarwal. Positive solutions of higher-order nonlinear delay difference equations. Computers and Mathematics with Applications, 45(6–9):1203–1211, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000865.

Lai:2008:CPS

[Lai08] Yisheng Lai. Counting positive solutions for polynomial systems with real coefficients. Computers and Mathematics with Applications, 56(6):1587–1596, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001855.

Lamour:2005:IDC

[Lam05] R. Lamour. Index determination and calculation of consistent initial values for DAEs. Computers and Mathematics with Applications, 50(7):1125–1140, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210500341X.

Lan:2004:ITM

[Lan04a] K. Q. Lan. An intersection theorem for multivalued maps and applications. Computers and Mathematics with Applications, 48(5–6):725–729, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003086.

Lan:2004:NFP

[Lan04b] K. Q. Lan. New fixedpoint theorems for two maps and applications to problems on sets with convex sections and minimax inequalities. Computers and Mathematics with Applications, 47(2–3):195–205, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900168.

Lancia:2008:PHT

[Lan08]  Giuseppe Lancia.  The

phasing of heterozygous traits: Algorithms and complexity. Computers and Mathematics with Applications, 55(5):960–969, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700555X.

Lasser:2008:TSR

[Las08] Dieter Lasser. Triangular subpatches of rectangular Bézier surfaces. Computers and Mathematics with Applications, 55(8):1706– 1719, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006190.

Liu:2001:UAS

[LB01] Xinzhi Liu and G. Ballinger.

Uniform asymptotic stability of impulsive delay differential equations. Computers and Mathematics with Applications, 41(7–8):903– 915, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000328X.

Livne:2004:LMA

[LB04] O. E. Livne and A. Brandt. Local mode analysis of multicolor and composite relaxation schemes. Computers and Mathematics with Applications, 47 (2–3):301–317, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900260.

Lin:2005:TPB

[LBW05] Hong-Wei Lin, Hu-Jun Bao, and Guo-Jin Wang. Totally positive bases and progressive iteration approximation. Computers and Mathematics with Applications, 50(3–4):575–586, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002671.

Lin:2009:ORB

[LBW09] Yiqin Lin, Liang Bao, and Yimin Wei. Order reduction of bilinear MIMO dynamical systems using new block Krylov subspaces. Computers and Mathematics with Applications, 58(6):1093–1102, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004787.

Li:2009:EGU

[LBZ^{+}09] Jinhua Li, Wensheng Bao, Yun Zhao, Zhibing Ma, and Huangzhen Dong. Evolutionary generation of unique input/output sequences for class behavioral testing. Computers and Mathematics with Applications, 57(11–12):1800– 1807, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800549X.

Liu:2002:SBL

[LC02] Yirong Liu and Haibo Chen. Stability and bifurcations of limit cycles of the equator in a class of cubic polynomial systems. Computers and Mathematics with Applications, 44 (8–9):997–1005, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002092.

Li:2003:IIS

[LC03] Wen Li and Qianshun Chang. Inclusion intervals of singular values and applications. Computers and Mathematics with Applications, 45(10–11):1637–1646, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001433.

Liu:2004:OTD

[LC04] Shu Tang Liu and Guanrong Chen. Oscillations of two-dimensional nonlinear partial difference systems. Computers and Mathematics with Applications, 47(4–5):621–629, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900508.

Lai:2006:DCL

[LC06a] D. Lai and G. Chen. Distribution of Controlled Lyapunov Exponents via the Lai–Chen algorithm. Computers and Mathematics with Applications, 52(12): 1649–1656, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122106003476.

Lee:2006:AAL

[LC06b] Hyung-Chun Lee and Youngmi Choi. Analysis and approximation of linear feedback control problems for the Boussinesq equations.

Computers and Mathematics with Applications, 51 (5):829–848, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106000228.

Lou:2006:ASC

[LC06c] Xu-Yang Lou and Bao-Tong Cui. Asymptotic synchronization of a class of neural networks with reaction-diffusion terms and time-varying delays. Computers and Mathematics with Applications, 52(6– 7):897–904, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002641.

Li:2007:DFT

[LC07a] Sheng-Tun Li and Yi-Chung Cheng. Deterministic fuzzy time series model for forecasting enrollments. Computers and Mathematics with Applications, 53(12):1904– 1920, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001630.

Lou:2007:BES

[LC07b] Xuyang Lou and Baotong Cui. Boundedness and exponential stability for nonautonomous RCNNs with distributed delays. Computers and Mathematics with Applications, 54(4):589–598, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002465.

Liu:2008:CAP

[LCC08] Qingbing Liu, Guoliang Chen, and Jing Cai. Convergence analysis of the preconditioned Gauss–Seidel method for H-matrices. Computers and Mathematics with Applications, 56(8): 2048–2053, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108003386.

Li:2006:SBA

[LCGX06] Z. C. Li, Y. L. Chan, G. C. Georgiou, and C. Xenophontos. Special boundary approximation methods for Laplace equation problems with boundary singularities — applications to the motz problem. Computers and Mathematics with Applications, 51(1):115–142, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004426.

Leng:2006:CPM

[LCHL06] Jin-Song Leng, Zheng-Xing Cheng, Ting-Zhu Huang, and Choi-Hong Lai. Construction and properties of multiwavelet packets with arbitrary scale and the related algorithms of decomposition and reconstruction. Computers and Mathematics with Applications, 51(11):1663–1676, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001349.

Lee:2003:GVV

[LCJL03] Byung-Soo Lee, Shih-Sen Chang, Jong Soo Jung, and Suk-Jin Lee. Generalized vector version of Minty’s lemma and applications. Computers and Mathematics with Applications, 45(4–5):647–653, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000245.

Li:2000:FSTc

[LCL00] Hongxing Li, C. L. Philip Chen, and E. S. Lee. Factor space theory and fuzzy information processing — fuzzy decision making based on the concepts of feedback extension. Computers and Mathematics with Applications, 40 (6–7):845–864, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002017.

Li:2006:HOC

[LCL06] Jichun Li, Yitung Chen, and Guoqing Liu. Highorder compact ADI methods for parabolic equations. Computers and Mathematics with Applications, 52 (8–9):1343–1356, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003221.

Li:2008:FBD

[LCL08] Sheng-Tun Li, Yi-Chung Cheng, and Su-Yu Lin. A FCM-based deterministic forecasting model for fuzzy time series. Computers and Mathematics with Applications, 56(12):3052–3063, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004409.

Li:2009:DOR

[LCL^{+}09] Ying Li, Yijia Cao, Zhaoyan Liu, Yi Liu, and Quanyuan Jiang. Dynamic optimal reactive power dispatch based on parallel particle swarm optimization algorithm. Computers and Mathematics with Applications, 57(11–12):1835– 1842, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005403.

Lin:2008:NOR

[LCT08] Hong-Dar Lin, Yuan-Shyi Peter Chiu, and Chia-Kuan Ting. A note on optimal replenishment policy for imperfect quality EMQ model with rework and backlogging. Computers and Mathematics with Applications, 56(11):2819–2824, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004550.

Lan:2006:NRC

[LCV06] Heng-You Lan, Yeol Je Cho, and R. U. Verma. Nonlinear relaxed cocoercive variational inclusions involving (A, η)-accretive mappings in Banach spaces. Computers and Mathematics with Applications, 51(9–10):1529– 1538, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210600099X.

Liu:2006:ATF

[LCWL06a] Min Liu, Degang Chen, Cheng Wu, and Hongxing Li. Approximation theorem of the fuzzy transform in fuzzy reasoning and its application to the scheduling problem. Computers and Mathematics with Applications, 51(3–4):515–526, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105005183.

Liu:2006:FRB

[LCWL06b] Min Liu, Degang Chen, Cheng Wu, and Hongxing Li. Fuzzy reasoning based on a new fuzzy rough set and its application to scheduling problems. Computers and Mathematics with Applications, 51(9–10):1507– 1518, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000976.

Liu:2006:RMB

[LCWL06c] Min Liu, Degang Chen, Cheng Wu, and Hongxing Li. Reduction method based on a new fuzzy rough set in fuzzy information system and its applications to scheduling problems. Computers and Mathematics with Applications, 51(9–10):1571– 1584, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001027.

Lian:2005:EAF

[LCY05] Wang Lian, David W. Cheung, and S. M. Yiu. An efficient algorithm for finding dense regions for mining quantitative association rules. Computers and Mathematics with Applications, 50(3–4):471–490, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002907.

Li:2000:FSTb

[LCYL00] Hongxing Li, C. L. Phillip Chen, Vincent C. Yen, and E. S. Lee. Factor spaces theory and its applications to fuzzy information processing: Two kinds of factor space canes. Computers and Mathematics with Applications, 40 (6–7):835–843, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002005.

Luo:2002:OCSb

[LD02] Jiaowan Luo and L. Debnath. Oscillation criteria for second-order quasilinear functional differential equations. Computers and Mathematics with Applications, 44(5–6):731–739, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001864.

Lai:2006:P

[LD06] C.-H. Lai and A. J. Davies. Preface. Computers and Mathematics with Applications, 51(11): xiii, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001301.

Li:2008:TIC

[LD08] Zhu Li and Huanhe Dong. Two integrable couplings of the Tu hierarchy and their Hamiltonian structures. Computers and Mathematics with Applications, 55(11):2643–2652, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007638.

Lin:2005:SMB

[LDG05] Xiaojie Lin, Zengji Du, and Weigao Ge. Solvability of multipoint boundary value problems at resonance for higher-order ordinary differential equations. Computers and Mathematics with Applications, 49(1):1– 11, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000027.

Liu:2005:OES

[LDHW05] Xiaoyan Liu, Peizhu Ding, Jialin Hong, and Lijin Wang. Optimization of explicit symplectic schemes for time-dependent schrödinger equations. Computers and Mathematics with Applications, 50(3–4):637–644, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002762.

Li:2006:NPE

[LDK^{+}06] Deng Li, C. C. Douglas, Takashi Kako, Suzuki Masabumi, and Hagiwara Ichiro. A novel perturbation expansion method for coupled system of acoustics and structures. Computers and Mathematics with Applications, 51(11):1689– 1704, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001362.

Liu:2008:EMP

[LDK08] Zeqing Liu, Lokenath Debnath, and Shin Min Kang. Existence of monotone positive solutions to a third order two-point generalized right focal boundary value problem. Computers and Mathematics with Applications, 55(3):356–367, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003653.

Llanas:2000:IAF

[LdSF00] B. Llanas, M. Fernandez de Sevilla, and V. Feliu. An iterative algorithm for finding a nearest pair of points in two convex subsets of R^{n}. Computers and Mathematics with Applications, 40(8–9):971–983, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850087.

Li:2005:EPP

[LDZW05] Meili Li, Yongrui Duan, Weiping Zhang, and Miansen Wang. The existence of positive periodic solutions of a class of Lotka–Volterra type impulsive systems with infinitely distributed delay. Computers and Mathematics with Applications, 49(7–8):1037–1044, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001203.

Lee:2000:NSP

[Lee00] Jiann-Der Lee. A new scheme for planar shape recognition using wavelets. Computers and Mathematics with Applications, 39 (5–6):201–216, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100000547.

Lee:2001:TVM

[Lee01a]  Jiann-Der Lee.  Three-

view matching algorithm for multipolyhedron reconstruction using genetic algorithm. Computers and Mathematics with Applications, 42(10–11):1465–1483, November/December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122101002553.

Lee:2001:SSF

[Lee01b] Min Ho Lee. The stability of spiral flow between coaxial cylinders. Computers and Mathematics with Applications, 41(3–4):289–300, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210000273X.

Lee:2004:ADC

[Lee04]  Hyesuk K. Lee.  Anal-

ysis of a defect correction method for viscoelastic fluid flow. Computers and Mathematics with Applications, 48(7–8):1213–1229, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003566.

Lee:2009:EIE

[Lee09] Taehun Lee. Effects of incompressibility on the elimination of parasitic currents in the lattice Boltzmann equation method for binary fluids. Computers and Mathematics with Applications, 58(5):987–994, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001035.

Lei:2002:ARS

[Lei02] Guiyuan Lei. Adaptive random search in quasi-Monte Carlo methods for global optimization. Computers and Mathematics with Applications, 43(6–7):747–754, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003182.

Lesnic:2002:CAD

[Les02] D. Lesnic. Convergence of Adomian’s decomposition method: periodic temperatures. Computers and Mathematics with Applications, 44(1–2):13– 24, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210200127X.

Lesnic:2005:DMC

[Les05]  D. Lesnic.  The de-

composition method for Cauchy advection-diffusion problems. Computers and Mathematics with Applications, 49(4):525–537, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000477.

Liu:2006:NMC

[LET06] Yanqun Liu, A. Eberhard, and Kok Lay Teo. A numerical method for a class of mixed switching and impulsive optimal control problems. Computers and Mathematics with Applications, 52(5):625–636, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002422.

Leung:2000:UPE

[Leu00]  Man-Chun Leung.  Un-

bounded positive entire solutions of rotationally symmetric harmonic map equations. Computers and Mathematics with Applications, 39(1–2):9–20, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003090.

Levitskaya:2006:SDL

[Lev06]  I. S. Levitskaya.  Sta-

bility domain of a linear differential equation with two delays. Computers and Mathematics with Applications, 51(1):153–159, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004463.

Lin:2002:NTE

[LF02] Lai-Jiu Lin and Cheng Shih Feng. Nash-type equilibrium theorems and competitive Nash-type equilibrium theorems. Computers and Mathematics with Applications, 44(10– 11):1369–1378, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002638.

Larsson:2003:NSS

[LF03] E. Larsson and B. Fornberg. A numerical study of some radial basis function based solution methods for elliptic PDEs. Computers and Mathematics with Applications, 46(5–6):891–902, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901519.

Larsson:2005:TCA

[LF05] E. Larsson and B. Fornberg. Theoretical and computational aspects of multivariate interpolation with increasingly flat radial basis functions. Computers and Mathematics with Applications, 49(1):103–130, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000118.

Liu:2009:LIU

[LF09] Huan-Wen Liu and Le-Le Fan. Lagrange interpolations using bivariate C^{2} quintic supersplines on double Clough—Tocher refinements. Computers and Mathematics with Applications, 58(8):1636–1644, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005094.

Li:2000:UPS

[LFZ00] Wan-Tong Li, Xian-Ling Fan, and Cheng-Kui Zhong. Unbounded positive solutions of higher-order difference equations with singular nonlinear term. Computers and Mathematics with Applications, 39(3–4):177–184, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003430.

Li:2008:PPSb

[LFZ08] Yongkun Li, Xuanlong Fan, and Lili Zhao. Positive periodic solutions of functional differential equations with impulses and a parameter. Computers and Mathematics with Applications, 56(10):2556–2560, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003921.

Li:2000:ICG

[LG00] Qingde Li and J. G. Griffiths. Iterative closest geometric objects registration. Computers and Mathematics with Applications, 40 (10–11):1171–1188, November/December 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100002303.

Liu:2004:GAB

[LG04a] Yuji Liu and Weigao Ge.

Global asymptotic behavior of solutions of a forced delay difference equation. Computers and Mathematics with Applications, 47(8–9):1211–1224, April/ May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901150.

Liu:2004:PSNb

[LG04b] Yuji Liu and Weigao Ge. The positive solutions of nonautonomous hyperlogistic delay difference equations. Computers and Mathematics with Applications, 47(8–9):1177–1193, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901125.

Lara:2008:ASL

[LG08] L. P. Lara and M. Gadella. An approximation to solutions of linear ODE by cubic interpolation. Computers and Mathematics with Applications, 56(6):1488–1495, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001752.

Lian:2009:CVB

[LG09a] Hairong Lian and Weigao Ge. Calculus of variations for a boundary value problem of differential system on the half line. Computers and Mathematics with Applications, 58(1):58– 64, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002508.

Lu:2009:NAB

[LG09b] Xin Lu and Jihong Guan.

A new approach to building histogram for selectivity estimation in query processing optimization. Computers and Mathematics with Applications, 57(6):1037– 1047, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005634.

Lee:2000:UIT

[LGBS00] Te-Won Lee, M. Girolami, A. J. Bell, and T. J. Sejnowski. A unifying information-theoretic framework for independent component analysis. Computers and Mathematics with Applications, 39(11): 1–21, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001012.

Li:2006:BDA

[LGGSP06] Wei Li, P. Guerrero-García, and A. Santos-Palomo. A basis-deficiency-allowing primal phase-I algorithm using the most-obtuse-angle column rule. Computers and Mathematics with Applications, 51(6–7):903–914, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000460.

Liu:2009:MFC

[LGH09] Xuejun Liu, Jihong Guan, and Ping Hu. Mining frequent closed itemsets from a landmark window over online data streams. Computers and Mathematics with Applications, 57(6):927–936, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005683.

Liu:2009:POR

[LGY09] M. Z. Liu, J. F. Gao, and Z. W. Yang. Preservation of oscillations of the Runge–Kutta method for equation x′(t) + ax(t) + a_{1}x([t − 1]) = 0. Computers and Mathematics with Applications, 58(6):1113–1125, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004660.

Li:2001:TSR

[LGZ01] Xue-Zhi Li, Geni Gupur, and Guang-Tian Zhu. Threshold and stability results for an age-structured SEIR epidemic model. Computers and Mathematics with Applications, 42(6– 7):883–907, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002061.

Lee:2000:ETR

[LH00a] Jiann-Der Lee and Yu-Lin Hsiao. Extraction of tumor region in color images using wavelets. Computers and Mathematics with Applications, 40(6– 7):793–803, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001978.

Li:2000:BCL

[LH00b] Zi-Cai Li and Hung-Tsai Huang. Blending curves for landing problems by numerical differential equations II. Numerical methods. Computers and Mathematics with Applications, 39(5–6):165–187, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000523.

Li:2005:IKF

[LH05a] Hou-Biao Li and Ting-Zhu Huang. An improvement of Ky fan theorem for matrix eigenvalues. Computers and Mathematics with Applications, 49(5–6):789–803, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000921.

Liu:2005:RFN

[LH05b] Xin-Wang Liu and Shi-Lian Han. Ranking fuzzy numbers with preference weighting function expectations. Computers and Mathematics with Applications, 49(11–12):1731– 1753, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001926.

Liu:2006:PST

[LH06] Bingwen Liu and Lihong Huang. Periodic solutions for a two-neuron network with delays. Computers and Mathematics with Applications, 52(3–4):429–438, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002094.

Liu:2008:CAL

[LH08] Bolian Liu and Fengying Huang. A conjecture about lower multiexponent of primitive matrices. Computers and Mathematics with Applications, 56(8):2100–2106, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003453.

Liu:2005:BCS

[LHC05] Huan-Wen Liu, Don Hong, and Dun-Qian Cao. Bivariate C^{1} cubic spline space over a nonuniform type-2 triangulation and its subspaces with boundary conditions. Computers and Mathematics with Applications, 49(11–12):1853– 1865, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002026.

Lai:2008:IEP

[LHC08] Hsin-Yi Lai, Jung-Chang Hsu, and Cha’o-Kuang Chen. An innovative eigenvalue problem solver for free vibration of Euler– Bernoulli beam by using the Adomian decomposition method. Computers and Mathematics with Applications, 56(12):3204–3220, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004331.

Lou:2008:IMS

[LHH08] Jian Lou, Xin-Feng He, and Zhen He. Iterative methods for solving a system of variational inclusions involving H − η-monotone operators in Banach spaces. Computers and Mathematics with Applications, 55(8):1832– 1841, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006554.

Li:2007:AHS

[LHL07] Liang Li, Ting-Zhu Huang, and Xing-Ping Liu. Asymmetric Hermitian and skew-Hermitian splitting methods for positive definite linear systems. Computers and Mathematics with Applications, 54(1):147– 159, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002076.

Liu:2008:NSS

[LHLD08] Duo Liu, Dongping Huang, Ping Luo, and Yiqi Dai. New schemes for sharing points on an elliptic curve. Computers and Mathematics with Applications, 56(6):1556–1561, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001818.

Luo:2007:FNN

[LHLW07] Xiaonan Luo, Wenbang Hou, Yi Li, and Zhong Wang. A fuzzy neural network model for predicting clothing thermal comfort. Computers and Mathematics with Applications, 53(12):1840– 1846, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700154X.

Lin:2007:FLT

[LHP07] Zhengyan Lin, Kyo-Shin Hwang, and Tian Xiao Pang. Functional limit theorems for the infinite series of OU processes in Hölder norm. Computers and Mathematics with Applications, 53(11):1644– 1657, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001368.

Lv:2009:MCP

[LHR09] Xiao-Guang Lv, Ting-Zhu Huang, and Zhi-Gang Ren. A modified T. Chan’s preconditioner for Toeplitz systems. Computers and Mathematics with Applications, 58(4):693–699, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003332.

Low:2008:MMA

[LHS08] Chinyao Low, Chou-Jung Hsu, and Chwen-Tzeng Su. Minimizing the makespan with an availability constraint on a single machine under simple linear deterioration. Computers and Mathematics with Applications, 56(1):257– 265, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000138.

Lu:2007:NEE

[LHT07] Dianchen Lu, Baojian Hong, and Lixin Tian. New explicit exact solutions for the generalized coupled Hirota–Satsuma KdV system. Computers and Mathematics with Applications, 53(8):1181–1190, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000752.

Lin:2004:MCS

[LHW04] Yung-Chien Lin, Kao-Shing Hwang, and Feng-Sheng Wang. A mixedcoding scheme of evolutionary algorithms to solve mixed-integer nonlinear programming problems. Computers and Mathematics with Applications, 47(8–9):1295–1307, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210490123X.

Lv:2008:NNA

[LHWW08] Yibing Lv, Tiesong Hu, Guangmin Wang, and Zhongping Wan. A neural network approach for solving nonlinear bilevel programming problem. Computers and Mathematics with Applications, 55(12): 2823–2829, June 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107007808.

Lin:2009:QSS

[LHYM09] Ji-Cherng Lin, Tetz C.

Huang, Cheng-Zen Yang, and Nathan Mou. Quasiself-stabilization of a distributed system assuming read/write atomicity. Computers and Mathematics with Applications, 57(2): 184–194, January 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108006366.

Li:2000:CAF

[Li00a] Jichun Li. Convergence analysis of finite element methods for singularly perturbed problems. Computers and Mathematics with Applications, 40(6– 7):735–745, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001929.

Li:2000:ARF

[Li00b]  Xin Li.  Approximation

by rational functions in hardy space. Computers and Mathematics with Applications, 40(1):137– 143, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001474.

Liu:2000:TOM

[LI00c] Baoding Liu and K. Iwamura. Topological optimization models for communication network with multiple reliability goals. Computers and Mathematics with Applications, 39(7–8):59– 69, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000651.

Li:2001:LBS

[Li01a] Luoluo Li. Lower bounds for the smallest singular value. Computers and Mathematics with Applications, 41(3–4):483–487, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002893.

Li:2001:CSN

[Li01b] Wan-Tong Li. Classification schemes for nonoscillatory solutions of two-dimensional nonlinear difference systems. Computers and Mathematics with Applications, 42(3–5):341–355, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001596.

Li:2001:HCR

[Li01c] Zi-Cai Li. High convergence rates of digital image transformation by numerical integration using spline functions. Computers and Mathematics with Applications, 41(1–2):229– 255, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101850185.

Li:2001:SCT

[Li01d] Zi-Cai Li. Superconvergence of coupling techniques in combined methods for elliptic equations with singularities. Computers and Mathematics with Applications, 41(3–4):379–398, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002819.

Li:2002:UCD

[Li02] Jichun Li. Uniform convergence of discontinuous finite element methods for singularly perturbed reactiondiffusion problems. Computers and Mathematics with Applications, 44(1–2):231– 240, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001438.

Li:2004:CMT

[Li04a]  Sheng-Gang Li.  Char-

acterizations of minimal T_{31/2}L-topological spaces. Computers and Mathematics with Applications, 47 (10–11):1739–1743, May/ June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839877.

Li:2004:PSH

[Li04b]  Yongxiang Li.  Posi-

tive solutions of higherorder periodic boundary value problems. Computers and Mathematics with Applications, 48(1–2):153– 161, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840069.

Li:2005:SSN

[Li05] Ming-Chia Li. Stability of a saddle node bifurcation under numerical approximations. Computers and Mathematics with Applications, 49(11–12):1849– 1852, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002014.

Li:2007:VPS

[Li07]  S. J. Li.  On the ϵ-

variational principle for setvalued mappings. Computers and Mathematics with Applications, 53(10):1627– 1632, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001435.

Li:2008:DDS

[Li08a] Kelin Li. Delay-dependent stability analysis for impulsive BAM neural networks with time-varying delays. Computers and Mathematics with Applications, 56(8):2088–2099, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003441.

Li:2008:PPSa

[Li08b] Yongkun Li. Positive periodic solutions of a periodic neutral delay logistic equation with impulses. Computers and Mathematics with Applications, 56(9):2189–2196, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003532.

Li:2009:SPT

[Li09a] Wei Nian Li. Some Pachpatte type inequalities on time scales. Computers and Mathematics with Applications, 57(2):275–282, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006494.

Li:2009:VIM

[Li09b] Xiuying Li. Variational iteration method for nonlinear age-structured population models. Computers and Mathematics with Applications, 58(11–12):2177–2181, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001461.

Liakos:2004:DNS

[Lia04] A. Liakos. Discretization of the Navier–Stokes equations with slip boundary condition II. Computers and Mathematics with Applications, 48(7–8):1153–1166, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003529.

Liakos:2005:FEA

[Lia05] A. Liakos. Finite-element approximation of viscoelastic fluid flow with slip boundary condition. Computers and Mathematics with Applications, 49(2–3): 281–294, January/February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105000337.

Lin:2000:NWL

[Lin00] Fu-Rong Lin. Notes on wavelet-like basis matrices. Computers and Mathematics with Applications, 40 (6–7):761–769, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001942.

Lin:2001:SMR

[Lin01] Yi-Kuei Lin. Study on the multicommodity reliability of a capacitatedflow network. Computers and Mathematics with Applications, 42(1–2):255– 264, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001493.

Lin:2004:AES

[Lin04a] Yi-Kuei Lin. An algorithm to evaluate the system reliability for multicommodity case under cost constraint. Computers and Mathematics with Applications, 48(5–6):805–812, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003141.

Lin:2004:LIL

[Lin04b] Zhengyan Lin. The law of the iterated logarithm for the rescaled RS statistics without the second moment. Computers and Mathematics with Applications, 47(8–9):1389–1396, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901319.

Ling:2004:UQM

[Lin04c]  Leevan Ling.  A uni-

variate quasi-multiquadric interpolation with better smoothness. Computers and Mathematics with Applications, 48(5–6):897–912, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003220.

Lin:2005:OHO

[Lin05]  Xiaoyan Lin.  Oscilla-

tion for higher-order neutral superlinear delay difference equations with unstable type. Computers and Mathematics with Applications, 50(5–6):683–691, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003081.

Lin:2006:OSN

[Lin06a]  Xiaoyan Lin.  Oscilla-

tion of solutions of neutral difference equations with a nonlinear neutral term. Computers and Mathematics with Applications, 52(3–4):439–448, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002100.

Lin:2006:UEL

[Lin06b] Yi-Kuel Lin. Unreliability evaluation for a limitedflow network with failed nodes subject to the budget constraint. Computers and Mathematics with Applications, 51(1):73–82, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004451.

Lin:2007:AGA

[Lin07] Yi-Kuei Lin. An algorithm to generate all upper boundary points for (d, B) in terms of minimal cuts. Computers and Mathematics with Applications, 53(12):1785– 1791, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001319.

Lin:2008:FDF

[Lin08a] Chia-Hung Lin. Frequencydomain features for ECG beat discrimination using grey relational analysisbased classifier. Computers and Mathematics with Applications, 55(4):680–690, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005019.

Lin:2008:PMA

[Lin08b] Yi-Kuei Lin. Project management for arbitrary random durations and cost attributes by applying network approaches. Computers and Mathematics with Applications, 56(10):2650– 2655, November 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108004045.

Lin:2008:CIM

[Lin08c] Yiqin Lin. A class of iterative methods for solving nonsymmetric algebraic Riccati equations arising in transport theory. Computers and Mathematics with Applications, 56(12):3046–3051, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004434.

Liu:2000:DRB

[Liu00] K. M. Liu. Dimensional reduction for the beam in elasticity on a bounded domain. Computers and Mathematics with Applications, 39(1–2):145–168, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812219900320X.

Liu:2002:PSN

[Liu02] Bing Liu. Positive solutions of a nonlinear threepoint boundary value problem. Computers and Mathematics with Applications, 44(1–2):201–211, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001414.

Liu:2004:EUS

[Liu04a] Bing Liu. Existence and uniqueness of solutions for nonlocal boundary vector value problems of ordinary differential systems with higher order. Computers and Mathematics with Applications, 48(5–6):841–851, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003189.

Liu:2004:PSSa

[Liu04b]  Bing Liu.  Positive so-

lutions of second-order three-point boundary value problems with change of sign. Computers and Mathematics with Applications, 47(8–9):1351–1361, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901289.

Liu:2004:PSSb

[Liu04c]  Bing Liu.  Positive so-

lutions of singular threepoint boundary value problems for the one-dimensional p-Laplacian. Computers and Mathematics with Applications, 48(5–6):913–925, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003232.

Liu:2006:RSL

[Liu06a] Yong Hui Liu. Ranks of solutions of the linear matrix equation AX + Y B = C. Computers and Mathematics with Applications, 52 (6–7):861–872, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002604.

Liu:2006:SPB

[Liu06b] Yuji Liu. Solvability of periodic boundary value problems for n^{th}-order ordinary differential equations. Computers and Mathematics with Applications, 52(6– 7):1165–1182, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002860.

Liu:2007:MSP

[Liu07a] Yansheng Liu. Multiple solutions of periodic boundary value problems for first order differential equations.

Computers and Mathematics with Applications, 54(1): 1–8, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001605.

Liu:2007:SOG

[Liu07b] Zheng Liu. Some Ostrowski-Grüss type inequalities and applications. Computers and Mathematics with Applications, 53(1):73–79, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000363.

Liu:2008:EUP

[Liu08a] Bingwen Liu. Existence and uniqueness of periodic solutions for a kind of Rayleigh equation with two deviating arguments. Computers and Mathematics with Applications, 55(9):2108– 2117, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006906.

Liu:2008:CPI

[Liu08b] Haidong Liu. Comment on the paper: “Interval criteria for oscillation of secondorder nonlinear neutral differential equations” [Comput. Math. Appl. 46 (5–6) (2003) 903–918]. Computers and Mathematics with Applications, 56(6):1662–1663, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001922.

See [YYZ03].

Liu:2008:UMF

[Liu08c]  Lipei Liu.  Uniqueness

of meromorphic functions and differential polynomials. Computers and Mathematics with Applications, 56(12):3236–3245, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004938.

Liu:2008:SEI

[Liu08d]  Wenjun Liu.  Several

error inequalities for a quadrature formula with a parameter and applications. Computers and Mathematics with Applications, 56(7):1766–1772, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002551.

Liu:2008:RLS

[Liu08e]  Yong Hui Liu.  Ranks

of least squares solutions of the matrix equation AXB = C. Computers and Mathematics with Applications, 55(6):1270– 1278, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005950.

Liu:2009:ENT

[Liu09a] Hua-Wen Liu. Erratum to:

“Notes on Triple I method of fuzzy reasoning” [Comput. Math. Appl. 44 (2002) 1567–1579]. Computers and Mathematics with Applications, 58(8):1598–1603, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004982.

See [SFL02, hYL09].

Liu:2009:HVA

[Liu09b] Jun-Fang Liu. He’s variational approach for nonlinear oscillators with high nonlinearity. Computers and Mathematics with Applications, 58(11–12):2423–2426, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001801.

Liu:2009:LTP

[Liu09c] Ming-Sheng Liu. Landau’s theorem for planar harmonic mappings. Computers and Mathematics with Applications, 57(7):1142– 1146, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900025X.

Li:2008:SFO

[LJ08a] Fa-Chao Li and Chen-Xia Jin. Study on fuzzy optimization methods based on principal operation and inequity degree. Computers and Mathematics with Applications, 56(6):1545–1555, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001806.

Luo:2008:PBV

[LJ08b] Zhiguo Luo and Zhujun Jing. Periodic boundary value problem for first-order impulsive functional differential equations. Computers and Mathematics with Applications, 55(9):2094– 2107, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700689X.

Lin:2006:EUS

[LJL06] Xiaoning Lin, Daqing Jiang, and Xiaoyue Li. Existence and uniqueness of solutions for singular (k, n−k) conjugate boundary value problems. Computers and Mathematics with Applications, 52(3–4):375–382, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002045.

Liu:2009:DTI

[LJSH09] De-Kai Liu, Chaiporn Jaikaeo, Chien-Chung Shen, and Ren-Hung Hwang. Discovering timely information in MANETs. Computers and Mathematics with Applications, 57(11–12):1808– 1821, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005476.

Liu:2003:ISN

[LK03] Zeqing Liu and S. M.

Kang. Iterative solutions of nonlinear equations with ϕ-strongly accretive operators in uniformly smooth Banach spaces. Computers and Mathematics with Applications, 45 (4–5):623–634, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000221.

Liang:2004:CPD

[LK04] Jihua Liang and Hui Kou. Convex power domain and Vietoris space. Computers and Mathematics with Applications, 47(4–5):541– 548, February/March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104900442.

Liu:2004:CAS

[LKC04] Zeqing Liu, Shin Min Kang, and Yeol Je Cho. Convergence and almost stability of Ishikawa iterative scheme with errors for m-accretive operators. Computers and Mathematics with Applications, 47 (4–5):767–778, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900636.

Lee:2003:DTG

[LKH03] Yutae Lee, Young Hwa Kim, and Jae Doo Huh. Discrete-time Geo^{X} /G/1 queue with non-preemptive priority. Computers and Mathematics with Applications, 46(10–11):1625–1632, November/December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103901970.

Luo:2008:MME

[LKL08] Li-Shi Luo, Manfred Krafczyk, and Yang Liu. Mesoscopic methods in engineering and science. Computers and Mathematics with Applications, 55(7):1363– 1364, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006207.

Lasater:2006:PPS

[LKS^{+}06] M. S. Lasater, C. T. Kelley, A. G. Salinger, D. L. Woolard, and P. Zhao. Parallel parameter study of the Wigner–Poisson equations for RTDs. Computers and Mathematics with Applications, 51(11):1677– 1688, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001350.

Lee:2008:GNQ

[LKS08a] Byung-Soo Lee, M. Firdosh Khan, and Salahuddin. Generalized nonlinear quasivariational inclusions in Banach spaces. Computers and Mathematics with Applications, 56(5):1414–1422, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001661.

Lee:2008:GVV

[LKS08b] Byung-Soo Lee, M. Firdosh Khan, and Salahuddin. Generalized vector variational-type inequalities. Computers and Mathematics with Applications, 55(6):1164–1169, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005834.

Li:2002:GSF

[LL02a] Z. C. Li and T. T. Lu. Global superconvergence of finite element methods for biharmonic equations and blending surfaces. Computers and Mathematics with Applications, 44(3–4):413– 437, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001591.

Lu:2002:TLC

[LL02b] Zhengyi Lu and Yong Luo. Two limit cycles in three-dimensional Lotka– Volterra systems. Computers and Mathematics with Applications, 44(1–2):51– 66, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001293.

Li:2003:IFF

[LL03a] Hong-Xing Li and E. S.

Lee. Interpolation functions of feedforward neural networks. Computers and Mathematics with Applications, 46(12):1861–1874, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902422.

Li:2003:IRF

[LL03b] Hong-Xing Li and E. S.

Lee. Interpolation representations of fuzzy logic systems. Computers and Mathematics with Applications, 45(10–11):1683–1693, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001470.

Lu:2003:TLC

[LL03c] Zhengyi Lu and Yong Luo.

Three limit cycles for a three-dimensional Lotka– Volterra competitive system with a heteroclinic cycle. Computers and Mathematics with Applications, 46(2–3):231–238, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900277.

Liao:2005:LSS

[LL05] An-Ping Liao and Yuan Lei. Least-squares solution with the minimum-norm for the matrix equation ( AXB, GXH ) = ( c, D ). Computers and Mathematics with Applications, 50(3–4):539– 549, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210500266X.

Li:2006:CTQ

[LL06a] Quan-Lin Li and Chuang Lin. Continuous-time QBD processes with continuous phase variable. Computers and Mathematics with Applications, 52(10– 11):1483–1510, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210600335X.

Li:2006:MGP

[LL06b] Quan-Lin Li and Chuang Lin. The M/G/1 processorsharing queue with disasters. Computers and Mathematics with Applications, 51(6–7):987–998, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000319.

Li:2007:NSN

[LL07] Fengwei Li and Xueliang Li.

The neighbour-scattering number can be computed in polynomial time for interval graphs. Computers and Mathematics with Applications, 54(5):679–686, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002453.

Li:2008:HCC

[LL08a] Ping Li and Guizhen Liu.

Hamilton cycles in circuit graphs of matroids. Computers and Mathematics with Applications, 55 (4):654–659, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004981.

Liu:2008:PPS

[LL08b] Xingyuan Liu and Yuji Liu. On positive periodic solutions of functional difference equations with forcing term and applications. Computers and Mathematics with Applications, 56(9):2247–2255, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003581.

Liang:2005:NPR

[LLB05] Heng Liang, Qinghua Liu, and Fengshan Bai. A note on a pseudo-random number generator for personal computers. Computers and Mathematics with Applications, 49(2–3):331–333, January/February 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105000374.

Lois:2008:MTG

[LLC08] Gregg Lois, Anaël Lemaître, and Jean M. Carlson. Momentum transport in granular flows. Computers and Mathematics with Applications, 55(2):175–183, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003458.

Little:2002:PPS

[LLCS02] L. Little, Z. Li, H. G.

Choi, and Y. Saad. Particle partitioning strategies for the parallel computation of solid–liquid flows. Computers and Mathematics with Applications, 43(12):1591– 1616, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001220.

Liu:2002:OND

[LLD02] Shu Tang Liu, Yong Qing Liu, and Fei Qi Deng. Oscillation for nonlinear delay partial difference equations with positive and negative coefficients. Computers and Mathematics with Applications, 43(10–11):1219–1230, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102000937.

Li:2009:SFO

[LLJ09a] Fa-Chao Li, Li-Min Liu, and Chen-Xia Jin. Study on fuzzy optimization methods based on quasi-linear fuzzy number and genetic algorithm. Computers and Mathematics with Applications, 57(1):67–78, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004793.

Li:2009:PST

[LLJ09b] Gaoshang Li, Xiping Liu, and Mei Jia. Positive solutions to a type of nonlinear three-point boundary value problem with sign changing nonlinearities. Computers and Mathematics with Applications, 57(3):348–355, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006780.

Lyu:2002:PAP

[LLL02] Jung Lyu, Jr., Hsing Luh, and Ming-Chang Lee. Performance analysis of a parallel Dantzig–Wolfe decomposition algorithm for linear programming. Computers and Mathematics with Applications, 44(10– 11):1431–1437, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002675.

Li:2008:SFS

[LLL08a] Yajuan Li, Chuang Lin, and Quan-Lin Li. A simplified framework for stochastic workflow networks. Computers and Mathematics with Applications, 56(10):2700–2715, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004100.

Liang:2008:OOQ

[LLL08b] Gin-Shuh Liang, Ling-Yuan Lin, and Chin-Feng Liu. The optimum output quantity of a duopoly market under a fuzzy decision environment. Computers and Mathematics with Applications, 56(5):1176–1187, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001193.

Li:2003:CFF

[LLML03] Hong-Xing Li, Yan-Da Li, Zhi-Hong Miao, and E. S. Lee. Control functions of fuzzy controllers. Computers and Mathematics with Applications, 46(5–6):875–890, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901507.

Lin:2005:SFC

[LLNS05] Shy-Der Lin, Wei-Chich Ling, Katsuyuki Nishimoto, and H. M. Srivastava. A simple fractionalcalculus approach to the solutions of the Bessel differential equation of general order and some of its applications. Computers and Mathematics with Applications, 49(9–10):1487– 1498, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500163X.

Lee:2009:CET

[LLS^{+}09a] Tian-Fu Lee, Jenn-Long Liu, Mei-Jiun Sung, Shiueng-Bien Yang, and Chia-Mei Chen. Communicationefficient three-party protocols for authentication and key agreement. Computers and Mathematics with Applications, 58(4):641–648, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003757.

Li:2009:FAS

[LLS09b] Qingyong Li, Siwei Luo, and Zhongzhi Shi. Fuzzy aesthetic semantics description and extraction for art image retrieval. Computers and Mathematics with Applications, 57(6):1000– 1009, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005701.

Liu:2009:ENN

[LLS^{+}09c] Mei Liu, Haihao Li, Yi Shen, Jianfeng Fan, and Shuangning Huang. Elastic neural network method for multi-target tracking task allocation in wireless sensor network. Computers and Mathematics with Applications, 57(11–12):1822– 1828, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005440.

Liao:2008:GAB

[LLT08] Maoxin Liao, Xianyi Li, and Xianhua Tang. Global asymptotic behavior and boundedness of positive solutions to an odd-order rational difference equation. Computers and Mathematics with Applications, 56(2): 305–310, July 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108000102.

Liu:2009:RMC

[LLT09] Zhen Liu, Liang Lu, and Sengyan Tsai. The research on mechanism of color management system based on iCAM color appearance model. Computers and Mathematics with Applications, 57(11–12):1829– 1834, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005373.

Liu:2005:RPI

[LLTL05] X. Liu, G. R. Liu, K. Tai, and K. Y. Lam. Radial point interpolation collocation method (RPICM) for partial differential equations. Computers and Mathematics with Applications, 50 (8–9):1425–1442, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003925.

Liu:2007:PSS

[LLW07] Bingmei Liu, Lishan Liu, and Yonghong Wu. Positive solutions for singular systems of threepoint boundary value problems. Computers and Mathematics with Applications, 53(9):1429–1438, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001137.

Lin:2008:LCN

[LLW08] Lijing Lin, Tzon-Tzer Lu, and Yimin Wei. On level-2 condition number for the weighted Moore– Penrose inverse. Computers and Mathematics with Applications, 55(4):788–800, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005093.

Lee:2000:UPS

[LLWY00] Chung-Fen Lee, Wei-Cheng Lian, Fu-Hsiang Wong, and Cheh-Chih Yeh. Uniqueness of positive solutions for generalized Laplacian boundary value problems. Computers and Mathematics with Applications, 40 (8–9):985–993, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850099.

Li:2005:PFP

[LLX05] Ke Li, Jin Liang, and Ti-Jun Xiao. Positive fixed points for nonlinear operators. Computers and Mathematics with Applications, 50(10– 12):1569–1578, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105004086.

Lu:2003:RLC

[LLXC03] Junan Lu, Jinhu Lü, Jin Xie, and Guanrong Chen. Reconstruction of the Lorenz and Chen systems with noisy observations. Computers and Mathematics with Applications, 46 (8–9):1427–1434, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902306.

Lee:2001:DLS

[LLY01] Won-Ho Lee, Keon-Jik Lee, and Kee-Young Yoo. Design of a linear systolic array for computing modular multiplication and squaring in GF(2^{m}). Computers and Mathematics with Applications, 42(1–2):231– 240, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210100147X.

Luo:2003:ARR

[LM03] A. C. J. Luo and C. D. Mote, Jr. Asymmetric responses of rotating, thin disks experiencing large deflections. Computers and Mathematics with Applications, 45 (1–3):217–228, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800159.

Lanzeni:2008:GMM

[LMA08] Stefano Lanzeni, Enza Messina, and Francesco Archetti. Graph models and mathematical programming in biochemical network analysis and metabolic engineering design. Computers and Mathematics with Applications, 55(5):970–983, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005706.

Li:2005:NKF

[LMHW05] Hong-Xing Li, Zhi-Hong Miao, Song-Chol Han, and Jia-Yin Wang. A new kind of fuzzy relation equations based on inner transformation. Computers and Mathematics with Applications, 50(3–4):623–636, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002725.

Li:2002:VUS

[LML02] Hong-Xing Li, Zhi-Hong Miao, and E. S. Lee. Variable universe stable adaptive fuzzy control of a nonlinear system. Computers and Mathematics with Applications, 44(5–6):799–815, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210200192X.

Lou:2004:MIC

[LMSH04] Jie Lou, Zhien Ma, Yiming Shao, and Litao Han. Modelling the interaction of T-cells, antigen presenting cells, and HIV-1 in vivo. Computers and Mathematics with Applications, 48(1– 2):9–33, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104839956.

Long:2007:IMF

[LN07] Guangqing Long and Gnaneshwar Nelakanti. Iteration methods for Fredholm integral equations of the second kind. Computers and Mathematics with Applications, 53(6):886–894, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000417.

Liu:2008:IJN

[LN08] Hao Liu and Qin Ni. Incomplete Jacobian Newton method for nonlinear equations. Computers and Mathematics with Applications, 56(1):218– 227, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000084.

Liu:2009:OGT

[LN09] Wenjun Liu and Quô’c-Anh Ngô. An Ostrowski-Grüss type inequality on time scales. Computers and Mathematics with Applications, 58(6):1207–1210, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004623.

Liu:2008:TTA

[LNL08] D. M. Liu, G. Naadimuthu, and E. S. Lee. Trajectory tracking in aircraft landing operations management using the adaptive neural fuzzy inference system. Computers and Mathematics with Applications, 56(5):1322–1327, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001545.

LoSchiavo:2002:PKM

[Lo 02] M. Lo Schiavo. Population kinetic models for social dynamics: Dependence on structural parameters. Computers and Mathematics with Applications, 44 (8–9):1129–1146, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002213.

Long:2003:MGF

[Lon03] Yong-Hong Long. A multiunit generalization of a first price auction. Computers and Mathematics with Applications, 46(5–6):681–684, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901337.

Longani:2008:FNL

[Lon08]  Vites Longani.  A for-

mula for the number of labelled trees. Computers and Mathematics with Applications, 56(11):2786–2788, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004604.

Lopez:2009:SRP

[Lóp09]  Rubén López.  Stabil-

ity results for polyhedral complementarity problems. Computers and Mathematics with Applications, 58(7):1475–1486, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900474X.

Lenin:2000:BDP

[LP00] R. B. Lenin and P. R.

Parthasarathy. A birthdeath process suggested by a chain sequence. Computers and Mathematics with Applications, 40(2–3):239–247, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001577.

Lin:2004:CGE

[LP04] Zhigui Lin and M. Pedersen. Coexistence of a general elliptic system in population dynamics. Computers and Mathematics with Applications, 48(3–4):617– 628, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840811.

Lefevre:2006:NRM

[LP06] Cl. Lefèvre and Ph. Picard. A nonhomogeneous risk model for insurance. Computers and Mathematics with Applications, 51(2):325–334, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004700.

LaRocca:2008:DBC

[LP08a] A. La Rocca and H. Power. A double boundary collocation Hermitian approach for the solution of steady state convection– diffusion problems. Computers and Mathematics with Applications, 55(9):1950– 1960, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006633.

Lakshminarayanan:2008:EEC

[LP08b] E. S. Lakshminarayanan and M. Pitchaimani. Existence and estimation of critical negative Gompertz parameter. Computers and Mathematics with Applications, 55(8):1686–1692, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006074.

Lakshminarayanan:2008:ECG

[LP08c] E. S. Lakshminarayanan and M. Pitchaimani. Existence of a critical Gompertz parameter and its asymptotic expression. Computers and Mathematics with Applications, 55(6):1302– 1309, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005998.

Lai:2005:DAE

[LPRH05] C.-H. Lai, A. K. Parrott, S. Rout, and M. E. Honnor. A distributed algorithm for European options with nonlinear volatility. Computers and Mathematics with Applications, 49(5–6):885–894, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000982.

Lloyd:2002:CKS

[LPSS02] N. G. Lloyd, J. M. Pearson, E. Saéz, and I. Szántó. A cubic Kolmogorov system with six limit cycles. Computers and Mathematics with Applications, 44(3–4):445– 455, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210200161X.

Lotstedt:2007:STA

[LPvST07] Per Lötstedt, Jonas Persson, Lina von Sydow, and Johan Tysk. Space-time adaptive finite difference method for European multiasset options. Computers and Mathematics with Applications, 53(8):1159– 1180, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000740.

Liu:2004:PSNa

[LQ04] Yansheng Liu and Aiqin Qi. Positive solutions of nonlinear singular boundary value problem in abstract space. Computers and Mathematics with Applications, 47 (4–5):683–688, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900557.

Lawrence:2003:CES

[LR03] B. A. Lawrence and D. T. Reid. Comparison of eigenvalues for Sturm– Liouville boundary value problems on a measure chain. Computers and Mathematics with Applications, 45(6–9):1319–1326, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001056.

Lenbury:2006:SSK

[LRL06] Yongwimon Lenbury, C. Rattanakul, and Dening Li. Stability of solution of Kuramoto–Sivashinsky–Korteweg– de Vries system. Computers and Mathematics with Applications, 52(3–4):497– 508, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210600215X.

Li:2009:CPC

[LRL09] Zhiyong Li, Günter Rudolph, and Kenli Li. Convergence performance comparison of quantum-inspired multi-objective evolutionary algorithms. Computers and Mathematics with Applications, 57(11–12):1843– 1854, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005361.

Laita:2001:ESM

[LRLM^{+}01] L. M. Laita, E. Roanes-Lozano, V. Maojo, E. Roanes-Macías, L. de Ledesma, and L. Laita. An expert system for managing medical appropriateness criteria based on computer algebra techniques. Computers and Mathematics with Applications, 42(12):1505–1522, December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002589.

Luo:2001:NRO

[LS01a] Zhiguo Luo and Jianhua Shen. New results for oscillation of delay difference equations. Computers and Mathematics with Applications, 41(5–6):553– 561, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002984.

Luo:2001:OSN

[LS01b] Zhiguo Luo and Jianhua Shen. Oscillation for solutions of nonlinear neutral differential equations with impulses. Computers and Mathematics with Applications, 42(10– 11):1285–1292, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002395.

Lu:2002:IBM

[LS02] Tzon-Tzer Lu and Sheng-Hua Shiou. Inverses of 2 × 2 block matrices. Computers and Mathematics with Applications, 43(1–2):119– 129, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002784.

Luo:2003:NRO

[LS03] Zhiguo Luo and Jianhua Shen. New results on oscillation for delay differential equations with piecewise constant argument. Computers and Mathematics with Applications, 45(12):1841–1848, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900058.

Lee:2004:ANI

[LS04a] Jinwoo Lee and Dongwoo Sheen. An accurate numerical inversion of Laplace transforms based on the location of their poles. Computers and Mathematics with Applications, 48(10– 11):1415–1423, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003979.

Li:2004:PSB

[LS04b] Wan-Tong Li and Jian-Ping Sun. Positive solutions of BVPs for third-order nonlinear difference systems. Computers and Mathematics with Applications, 47 (4–5):611–620, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900491.

Liu:2006:EFN

[LS06a] Guo-Dong Liu and H. M. Srivastava. Explicit formulas for the Nörlund polynomials B_{n}^{(χ)} and b_{n}^{(χ)}. Computers and Mathematics with Applications, 51(9–10): 1377–1384, May 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122106001064.

Luo:2006:SRB

[LS06b] Qiu-Ming Luo and H. M. Srivastava. Some relationships between the Apostol– Bernoulli and Apostol– Euler polynomials. Computers and Mathematics with Applications, 51(3–4):631– 642, February 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105005274.

Li:2008:NBV

[LS08] Wenjuan Li and Guangxing Song. Nonlinear boundary value problem for second order impulsive integrodifferential equations of mixed type in Banach space. Computers and Mathematics with Applications, 56(5):1372–1381, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001612.

Li:2009:PSS

[LS09] Hongyu Li and Jingxian Sun. Positive solutions of sublinear Sturm–Liouville problems with changing sign nonlinearity. Computers and Mathematics with Applications, 58(9):1808–1815, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005793.

Li:2009:NRF

[LSL09] Jinquan Li, Kaibiao Sun, and Hongxing Li. A note on “Resolution of fuzzy relation equations (I) based on Boolean-type implications”. Computers and Mathematics with Applications, 57(8):1365–1368, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000583.

See [Luo06].

Luo:2005:ABV

[LSN05] Zhiguo Luo, Jianhua Shen, and J. J. Nieto. Antiperiodic boundary value problem for first-order impulsive ordinary differential equations. Computers and Mathematics with Applications, 49(2–3):253–261, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000301.

Lopez:2002:TRA

[LST02] J. L. López, J. Soler, and G. Toscani. Time rescaling and asymptotic behavior of some fourth-order degenerate diffusion equations. Computers and Mathematics with Applications, 43(6–7):721–736, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003169.

Lin:2008:NSL

[LSZ08] Tzu-Chu Lin, David H.

Schultz, and Weiqun Zhang. Numerical solutions of linear and nonlinear singular perturbation problems. Computers and Mathematics with Applications, 55(11):2574–2592, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007584.

Liu:2007:RMC

[LSZW07] Min Liu, Mingwen Shao, Wenxiu Zhang, and Cheng Wu. Reduction method for concept lattices based on rough set theory and its application. Computers and Mathematics with Applications, 53(9):1390– 1410, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001022.

Liu:2001:SDM

[LT01] Chin-Sung Liu and Ching-Huan Tseng. Spacedecomposition multiplier method for constrained minimization problems. Computers and Mathematics with Applications, 41(1–2):

51–62, January 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101850057.

Ling:2004:MCM

[LT04] Leevan Ling and M. R. Trümmer. Multiquadric collocation method with integral formulation for boundary layer problems. Computers and Mathematics with Applications, 48(5–6):927– 941, September 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104003244.

Ladopoulos:2005:ASI

[LT05] E. G. Ladopoulos and G. Tsamasphyros. Approximations of singular integral equations on Lyapunov contours in Banach spaces. Computers and Mathematics with Applications, 50(3–4):567–573, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002749.

Li:2009:FAM

[LT09a] Changpin Li and Chunxing Tao. On the fractional Adams method. Computers and Mathematics with Applications, 58(8):1573–1588, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004581.

Liu:2009:COS

[LT09b] Yan-Kui Liu and Miao Tian.

Convergence of optimal solutions about approximation scheme for fuzzy programming with minimumrisk criteria. Computers and Mathematics with Applications, 57(6):867–884, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005749.

Lin:2003:UPC

[LTS03] Shy-Der Lin, Shih-Tong Tu, and H. M. Srivastava. A unified presentation of certain families of non-Fuchsian differential equations via fractional calculus operators. Computers and Mathematics with Applications, 45(12):1861– 1870, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900071.

Lin:2002:COF

[LTSW02] Shy-Der Lin, Shih-Tong Tu, H. M. Srivastava, and Pin-Yu Wang. Certain operators of fractional calculus and their applications to differential equations. Computers and Mathematics with Applications, 44(12):1557–1565, December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210200278X.

Leong:2002:IMM

[LTT02] P. C. Leong, E. C. Tan, and P. C. Tan. An iterative modular multiplication algorithm. Computers and Mathematics with Applications, 44(1–2):175– 180, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001384.

Lee:2009:EHP

[LTTH09] Chung-Meng Lee, Yuan-Hsiang Teng, Jimmy J. M. Tan, and Lih-Hsing Hsu. Embedding Hamiltonian paths in augmented cubes with a required vertex in a fixed position. Computers and Mathematics with Applications, 58(9):1762–1768, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006038.

Li:2008:AAA

[LTWM08] Zi-Cai Li, Heng-Shuing Tsai, Song Wang, and John J. H. Miller. Accurate and approximate analytic solutions of singularly perturbed differential equations with two-dimensional boundary layers. Computers and Mathematics with Applications, 55(11):2602– 2622, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007602.

Li:2005:SOD

[LTY05] S. J. Li, K. L. Teo, and X. Q.

Yang. Second-order directional derivatives of spectral functions. Computers and Mathematics with Applications, 50(5–6):947–955, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003275.

Lai:2009:DBA

[LTYB09] Mingyong Lai, Xiaojiao Tong, Hongming Yang, and Pingping Bing. Dynamic bidding analysis in power market based on the supply function. Computers and Mathematics with Applications, 58(1):25– 38, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800583X.

Lu:2007:VIM

[Lu07] Junfeng Lu. Variational iteration method for solving a nonlinear system of secondorder boundary value problems. Computers and Mathematics with Applications, 54(7–8):1133–1138, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700301X.

Lu:2008:SIS

[Lü08] Zhongxue Lü. On sharp inequalities of Simpson type and Ostrowski type in two independent variables. Computers and Mathematics with Applications, 56(8):2043–2047, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003362.

Liu:2009:AWL

[LU09a] Ming-Hong Liu and Vasant A. Ubhaya. An O(n) algorithm for weighted least squares regression by integer quasi-convex and unimodal or umbrella functions. Computers and Mathematics with Applications, 58(4):776–783, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002831.

Lu:2009:AAS

[Lu09b] Junfeng Lu. An analytical approach to the sine-Gordon equation using the modified homotopy perturbation method. Computers and Mathematics with Applications, 58(11–12):2313–2319, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001606.

Lu:2009:CBT

[Lu09c]  Qiuying Lu.  Codimen-

sion 2 bifurcation of twisted double homoclinic loops. Computers and Mathematics with Applications, 57 (7):1127–1141, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109000534.

Lugojan:2001:M

[Lug01] S. Lugojan. (α, A)-manifolds. Computers and Mathematics with Applications, 41(3– 4):423–431, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100002856.

Lund:2003:UEF

[Lun03]  T. S. Lund.  The use

of explicit filters in large eddy simulation. Computers and Mathematics with Applications, 46(4):603–616, August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900198.

Luo:2002:OCSa

[Luo02] Jiaowan Luo. Oscillation criteria for second-order quasi-linear neutral difference equations. Computers and Mathematics with Applications, 43(12):1549– 1557, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001189.

Luo:2003:SOQ

[Luo03]  Jiaowan Luo.  Second-

order quasilinear oscillation with impulses. Computers and Mathematics with Applications, 46(2–3):279–291, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900319.

Luo:2006:RFR

[Luo06] Yanbin Luo. Resolution of fuzzy relation equations (I) based on Boolean-type implications. Computers and Mathematics with Applications, 52(3–4):421–428, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002082.

See note [LSL09].

Li:2005:FPM

[LUZ05] Kuiyuan Li, J. Uvah, and Shengbian Zhao. A fully parallel method for the singular eigenvalue problem. Computers and Mathematics with Applications, 49(7–8):1279–1284, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001379.

Lin:2003:RTF

[LV03] E.-B. Lin and R. U. Verma.

Role of twice fréchetdifferentiable mappings in general auxiliary problem principle. Computers and Mathematics with Applications, 45(10–11):1749–1756, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001536.

Liao:2002:SAP

[LW02] Li-Zhi Liao and Shengli Wang. A self-adaptive projection and contraction method for monotone symmetric linear variational inequalities. Computers and Mathematics with Applications, 43(1–2):41–48, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002693.

Lin:2003:CSF

[LW03] Ying-Fen Lin and Ngai-Ching Wong. Constructing space-filling curves of compact connected manifolds. Computers and Mathematics with Applications, 45(12):1871–1881, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900083.

Lee:2004:RNG

[LW04] Lap-Piu Lee and Kwok-Wo Wong. A random number generator based on elliptic curve operations. Computers and Mathematics with Applications, 47(2–3):217–226, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900181.

Li:2005:EGA

[LW05] J. W. Li and Z. C. Wang.

Existence and global attractivity of positive periodic solutions of a survival model of red blood cells. Computers and Mathematics with Applications, 50(1–2):41– 47, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002300.

Liu:2006:NUF

[LW06] Hua-Wen Liu and Guo-Jun Wang. A note on the unified forms of triple I method. Computers and Mathematics with Applications, 52(10– 11):1609–1613, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003440.

Liu:2008:CTM

[LW08a] Hua-Wen Liu and Guo-Jun Wang. Continuity of triple I methods based on several implications. Computers and Mathematics with Applications, 56(8):2079–2087, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800343X.

Liu:2008:TMB

[LW08b] Hua-Wen Liu and Guo-Jun Wang. Triple I method based on pointwise sustaining degrees. Computers and Mathematics with Applications, 55(11):2680– 2688, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007675.

Lang:2009:BCS

[LW09a] Feng-Gong Lang and Ren-Hong Wang. Bivariate C^{1} cubic spline spaces with homogeneous boundary conditions over FVS triangulation. Computers and Mathematics with Applications, 58(3):464–473, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003320.

Lazarev:2009:GRD

[LW09b] Alexander A. Lazarev and Frank Werner. A graphical realization of the dynamic programming method for solving NP-hard combinatorial problems. Computers and Mathematics with Applications, 58(4):619–631, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003794.

Li:2009:NAB

[LW09c] Changpin Li and Yihong Wang. Numerical algorithm based on Adomian decomposition for fractional differential equations. Computers and Mathematics with Applications, 57(10):1672– 1681, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002028.

Li:2009:RET

[LW09d] Cunbin Li and Kecheng Wang. The risk element transmission theory research of multi-objective risk-time-cost trade-off. Computers and Mathematics with Applications, 57(11–12):1792– 1799, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005518.

Lin:2009:SVR

[LW09e] Lai-Jiu Lin and Sung-Yu Wang. Simultaneous variational relation problems and related applications. Computers and Mathematics with Applications, 58(9):1711–1721, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109006130.

Liu:2009:MIS

[LW09f] Wen Liu and Lipo Wang. Minimizing interference in satellite communications using transiently chaotic neural networks. Computers and Mathematics with Applications, 57(6):1024– 1029, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005646.

Liu:2004:ETG

[LWG04] Lishan Liu, Congxin Wu, and Fei Guo. Existence theorems of global solutions of initial value problems for nonlinear integrodifferential equations of mixed type in Banach spaces and applications. Computers and Mathematics with Applications, 47(1):13–22, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900028.

Liu:2008:IPO

[LWL08] Hsuan-Ku Liu, Berlin Wu, and Ming Long Liu. Investors’ preference order of fuzzy numbers. Computers and Mathematics with Applications, 55(11):2623– 2630, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007614.

Li:2009:AIA

[LWLJ09] Anqiang Li, Liping Wang, Jiqing Li, and Changming Ji. Application of immune algorithm-based particle swarm optimization for optimized load distribution among cascade hydropower stations. Computers and Mathematics with Applications, 57(11–12):1785– 1791, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005531.

Lin:2002:NSE

[LWY02] Tzu-Chu Lin and Y. Warnapala-Yehiya. The numerical solution of the exterior Dirichlet problem for Helmholtz’s equation via modified Green’s functions approach. Computers and Mathematics with Applications, 44(8– 9):1229–1248, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002298.

Lin:2004:NSE

[LWY04] Tzu-Chu Lin and Y. Warnapala-Yehiya. The numerical solution of exterior Neumann problem for Helmholtz’s equation via modified Green’s functions approach. Computers and Mathematics with Applications, 47(4–5): 593–609, February/March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S089812210490048X.

Li:2006:LLT

[LWY06] Jingna Li, Xiaofeng Wang, and Zheng-An Yao. A Landau–Lifshitz-type equation in studying thin film micromagnetics. Computers and Mathematics with Applications, 52(6– 7):873–882, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002884.

Liu:2009:MGR

[LWY09] Zaiming Liu, Jinbiao Wu, and Gang Yang. An M/G/1 retrial G-queue with preemptive resume and feedback under N-policy subject to the server breakdowns and repairs. Computers and Mathematics with Applications, 58(9):1792–1807, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006002.

Li:2004:SPE

[LWZ04] Quanlin Li, Guanghui Wang, and Yuan Zhou. Symmetric PH and EP distributions and their applications to the probability Hough transform. Computers and Mathematics with Applications, 47(6–7):823–844, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900673.

Lai:2008:SPS

[LWZ08] Shaoyong Lai, YongHong Wu, and Yuan Zhou. Some physical structures for the (2 + 1)-dimensional Boussinesq water equation with positive and negative exponents. Computers and Mathematics with Applications, 56(2):339– 345, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000199.

Li:2009:SIS

[LWZ09] Ming Li, Yan Wu, and Qiang Zhang. SAR image segmentation based on mixture context and wavelet hidden-class-label Markov random field. Computers and Mathematics with Applications, 57(6):961–969, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005774.

Liang:2000:CPN

[LX00] Jin Liang and Ti-Jun Xiao. The Cauchy problem for nonlinear abstract functional differential equations with infinite delay. Computers and Mathematics with Applications, 40 (6–7):693–703, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001887.

Li:2003:GBS

[LX03] Fu-Cai Li and Chun-Hong Xie. Global and blow-up solutions to a p-Laplacian equation with nonlocal source. Computers and Mathematics with Applications, 46(10– 11):1525–1533, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390188X.

Liang:2004:SIE

[LX04] Jin Liang and Ti-Jun Xiao. Semilinear integrodifferential equations with nonlocal initial conditions. Computers and Mathematics with Applications, 47(6–7):863–875, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900715.

Li:2006:CTP

[LX06] Min Li and Minghua Xu.

Comparison of two proximal point algorithms for monotone variational inequalities. Computers and Mathematics with Applications, 52 (10–11):1543–1554, November/December 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122106003385.

Liu:2007:OFS

[LX07] Xiuxiang Liu and Zhiting Xu. Oscillation of a forced super-linear second order differential equation with impulses. Computers and Mathematics with Applications, 53(11):1740– 1749, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001472.

Liang:2005:PSS

[LXH05] Jin Liang, Ti-Jun Xiao, and Zhao-Cai Hao. Positive solutions of singular differential equations on measure chains. Computers and Mathematics with Applications, 49(5–6):651–663, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000805.

Liu:2009:GSS

[LXK09] Zeqing Liu, Yuguang Xu, and Shin Min Kang. Global solvability for a second order nonlinear neutral delay difference equation. Computers and Mathematics with Applications, 57(4):587–595, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006688.

Liu:2008:DSL

[LXZ08] Dan Liu, Guoliang Xu, and Qin Zhang. A discrete scheme of Laplace– Beltrami operator and its convergence over quadrilateral meshes. Computers and Mathematics with Applications, 55(6):1081– 1093, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005731.

Limayem:2004:GRL

[LY04] F. Limayem and B. Yannou. Generalization of the RCGM and LSLR pairwise comparison methods. Computers and Mathematics with Applications, 48(3–4):539–548, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840756.

Liou:2005:BDV

[LY05a] Yeong-Cheng Liou and Jen-Chih Yao. Bilevel decision via variational inequalities. Computers and Mathematics with Applications, 49(7–8):1243–1253, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001343.

Liu:2005:EUP

[LY05b] Yansheng Liu and Huimin Yu. Existence and uniqueness of positive solution for singular boundary value problem. Computers and Mathematics with Applications, 50(1–2):133–143, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002373.

Liu:2006:PPS

[LY06a] Guirong Liu and Jurang Yan. Positive periodic solutions for a neutral differential system with feedback control. Computers and Mathematics with Applications, 52(3–4):401–410, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002069.

Lv:2006:GCP

[LY06b] Jian Cheng Lv and Zhang Yi. Global convergence of a PCA learning algorithm with a constant learning rate. Computers and Mathematics with Applications, 52(10– 11):1425–1438, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003300.

Li:2008:EID

[LY08a] Changfeng Li and Yirang Yuan. Explicit/ implicit domain decomposition method with modified upwind differences for convection–diffusion equations. Computers and Mathematics with Applications, 55(11):2565–2573, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007572.

Li:2008:PPE

[LY08b] Jinxian Li and Jurang Yan. Partial permanence and extinction in an N-species nonautonomous Lotka–Volterra competitive system. Computers and Mathematics with Applications, 55(1):76–88, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003367.

Li:2008:SRR

[LY08c] Xiao-Min Li and Hong-Xun Yi. Some results on the regular solutions of a linear differential equation. Computers and Mathematics with Applications, 56(9):2210–2221, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003556.

Liu:2008:FDS

[LY08d] Wei Liu and Yirang Yuan. Finite difference schemes for two-dimensional miscible displacement flow in porous media on composite triangular grids. Computers and Mathematics with Applications, 55(3):470–484, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700483X.

Lai:2009:GAD

[LY09a] Xiang Lai and Yirang Yuan. Galerkin alternatingdirection method for a kind of three-dimensional nonlinear hyperbolic problems.

Computers and Mathematics with Applications, 57 (3):384–403, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210800638X.

Liu:2009:NMU

[LY09b] Hao Liu and Yongxin Yuan. New model updating method for damped structural systems. Computers and Mathematics with Applications, 57(5):685–690, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006202.

Lu:2009:EOP

[LY09c] Hengyun Lu and Genke Yang. Extremal optimization for protein folding simulations on the lattice. Computers and Mathematics with Applications, 57(11–12):1855– 1861, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005427.

Liu:2007:RNN

[LYC07] Yiguang Liu, Zhisheng You, and Liping Cao. A recurrent neural network computing the largest imaginary or real part of eigenvalues of real matrices. Computers and Mathematics with Applications, 53(1):41–53, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000259.

Li:2009:TSM

[LYF09] Weiguo Li, Fan Yang, and Daining Fang. Thermal shock modeling of ultra-high temperature ceramics under active cooling. Computers and Mathematics with Applications, 58(11–12):2373–2378, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002430.

Long:2005:DDS

[LYG05] Y. Long, J. Yu, and Zh. Guo. Disconjugacy and C-disfocality of secondorder self-adjoint difference equation. Computers and Mathematics with Applications, 49(5–6):933–942, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001021.

Li:2009:DAC

[LYJ09] Jinxian Li, Jurang Yan, and Xinchun Jia. Dynamical analysis of Cohen– Grossberg neural networks with time-delays and impulses. Computers and Mathematics with Applications, 58(6):1142–1151, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003769.

Li:2000:FSTa

[LYL00a] Hongxing Li, V. C. Yen, and E. S. Lee. Factor space theory in fuzzy information processing — composition of states of factors and multifactorial decision making. Computers and Mathematics with Applications, 39(1–2):245– 265, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003272.

Li:2000:MNB

[LYL00b] Hongxing Li, V. C. Yen, and E. S. Lee. Models of neurons based on factor space. Computers and Mathematics with Applications, 39(12): 91–100, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001322.

Lv:2007:SRK

[LYL07] W. J. Lv, Z. W. Yang, and M. Z. Liu. Stability of Runge–Kutta methods for the alternately advanced and retarded differential equations with piecewise continuous arguments. Computers and Mathematics with Applications, 54(3):326–335, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001708.

Yang:2009:EES

[lYsT09] Xian lin Yang and Jia shi Tang. Explicit exact solutions for the generalized Zakharov equations with nonlinear terms of any order. Computers and Mathematics with Applications, 57(10):1622– 1629, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000443.

Liao:2002:GAL

[LYW02] Liusheng Liao, Jianshe Yu, and Lin Wang. Global attractivity in a logistic difference model with a feedback control. Computers and Mathematics with Applications, 44(10– 11):1403–1411, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002651.

Lai:2004:ATG

[LYW04] Shaoyong Lai, Xu Yang, and Yong Hong Wu. The asymptotic theory of global solutions for semilinear wave equations in three space dimensions. Computers and Mathematics with Applications, 47(10–11):1535–1543, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839701.

Lu:2009:SFR

[LYX09] Feng Lü, Hongxun Yi, and Junfeng Xu. Some further results on weighted sharing three values and Brosch’s theorem. Computers and Mathematics with Applications, 58(1):11– 24, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000595.

Liu:2002:OBD

[LZ02] Shu Tang Liu and Bing Gen Zhang. Oscillatory behavior of delay partial difference equations with positive and negative coefficients. Computers and Mathematics with Applications, 43(8–9):951–964, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800050.

Li:2004:SNR

[LZ04a] Xianyi Li and Deming Zhu. Some new results for bounded and monotone properties of solutions of second-order quasilinear forced difference equations. Computers and Mathematics with Applications, 47(6–7):997–1009, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210490082X.

Li:2004:TRR

[LZ04b] Xianyi Li and Deming Zhu. Two rational recursive sequences. Computers and Mathematics with Applications, 47(10–11):1487–1494, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210483966X.

Li:2005:RSG

[LZ05] Xianyi Li and Deming Zhu. The rule of semicycle and global asymptotic stability for a fourthorder rational difference equation. Computers and Mathematics with Applications, 49(5–6):723–730, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000866.

Lee:2009:TPP

[LZ09] Eun-Joo Lee and Jun Zhang. A two-phase preconditioning strategy of sparse approximate inverse for indefinite matrices. Computers and Mathematics with Applications, 58(6):1152–1159, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004532.

Li:2004:CUP

[LZJ04] Chong Li, Wen-Hong Zhang, and Xiao-Qing Jin. Convergence and uniqueness properties of Gauss–Newton’s method. Computers and Mathematics with Applications, 47(6–7):1057–1067, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900867.

Li:2006:NET

[LZJ06] Xiaoyue Li, Xiaoying Zhang, and Daqing Jiang. A new existence theory for positive periodic solutions to functional differential equations with impulse effects. Computers and Mathematics with Applications, 51(12):1761–1772, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001258.

Li:2004:PPS

[LZL04] Yongkun Li, Lifei Zhu, and Ping Liu. Positive periodic solutions of nonlinear functional difference equations depending on a parameter. Computers and Mathematics with Applications, 48(10– 11):1453–1459, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104004006.

Li:2006:PPF

[LZL06] Chunguang Li, Hongbin Zhang, and Xiaofeng Liao. Passivity and passification of fuzzy systems with time delays. Computers and Mathematics with Applications, 52(6– 7):1067–1078, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002781.

Liu:2007:MDF

[LZS07] Guoqiang Liu, Zongfang Zhou, and Yong Shi. A multi-dimensional forward selection method for firms’ credit sale. Computers and Mathematics with Applications, 54(9–10):1228–1233, November 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001241.

Li:2004:CAG

[LZZ04] Chenliang Li, Jinping Zeng, and Shuzi Zhou. Convergence analysis of generalized Schwarz algorithms for solving obstacle problems with T -monotone operator. Computers and Mathematics with Applications, 48(3–4):373–386, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840616.

Ma:2000:MPS

[Ma00]  Ruyun Ma.  Multiplic-

ity of positive solutions for second-order threepoint boundary value problems. Computers and Mathematics with Applications, 40(2–3):193–204, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210000153X.

Ma:2001:PSN

[Ma01] Ruyun Ma. Positive solutions of a nonlinear mpoint boundary value problem. Computers and Mathematics with Applications, 42 (6–7):755–765, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210100195X.

Ma:2004:PSN

[Ma04] Ruyun Ma. Positive solutions for nonhomogeneous m-point boundary value problems. Computers and Mathematics with Applications, 47(4–5):689–698, February/March 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104900569.

Ma:2007:BIM

[Ma07] Ruyun Ma. Bifurcation from infinity and multiple solutions for some discrete Sturm–Liouville problems. Computers and Mathematics with Applications, 54(4):535–543, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002362.

Ma:2009:ASS

[Ma09a] Zheng-Yi Ma. Approximate soliton solutions for a (2 + 1)-dimensional Broer– Kaup system by He’s methods. Computers and Mathematics with Applications, 58(11–12):2410–2415, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001722.

Mostafa:2009:NCV

[MA09b] A. O. Mostafa and M. K.

Aouf. Neighborhoods of certain p-valent analytic functions with complex order. Computers and Mathematics with Applications, 58(6):1183–1189, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004957.

Machihara:2006:FPP

[Mac06]  Fumiaki Machihara.  A

fractal Poisson process and its input queue. Computers and Mathematics with Applications, 51(2):199–208, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0898122105004578.

Maesumi:2000:JSR

[Mae00] M. Maesumi. Joint spectral radius and Hölder regularity of wavelets. Computers and Mathematics with Applications, 40(1):145– 155, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001486.

M:2008:GSI

[MAH08] Eduardo Hernández M., Sueli M. Tanaka Aki, and Hernán Henríquez. Global solutions for impulsive abstract partial differential equations. Computers and Mathematics with Applications, 56(5):1206–1215, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001223.

Mahmoud:2009:SDT

[Mah09] Magdi S. Mahmoud. Switched discrete-time systems with time-varying delays: a generalized H_{2} approach. Computers and Mathematics with Applications, 57(1):79– 95, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006184.

Mittal:2002:DNS

[MAK02] R. C. Mittal and A. Al-Kurdi. LU-decomposition and numerical structure for solving large sparse nonsymmetric linear systems. Computers and Mathematics with Applications, 43(1–2):131–155, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002796.

Mittal:2003:EMC

[MAK03] R. C. Mittal and A. H. Al-Kurdi. An efficient method for constructing an ILU preconditioner for solving large sparse nonsymmetric linear systems by the GMRES method. Computers and Mathematics with Applications, 45(10–11):1757–1772, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001548.

Makinde:2009:HPA

[Mak09] Oluwole Daniel Makinde. Hermite–Padé approach to thermal radiation effect on inherent irreversibility in a variable viscosity channel flow. Computers and Mathematics with Applications, 58(11–12):2330–2338, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001618.

Malits:2004:DTF

[Mal04] P. Malits. Dual and triple Fourier–Bessel series equations. Computers and Mathematics with Applications, 48(5–6):823–831, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003165.

Manojlovic:2000:OCS

[Man00] J. V. Manojlović. Oscillation criteria for secondorder sublinear differential equation. Computers and Mathematics with Applications, 39(9–10):161– 172, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000948.

Manojlovic:2001:IAO

[Man01] J. V. Manojlović. Integral averages and oscillation of second-order nonlinear differential equations. Computers and Mathematics with Applications, 41(12):1521– 1534, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001171.

Manouzi:2005:MVF

[Man05] H. Manouzi. A mixed variational formulation for the Navier–Stokes problem with hyper-dissipation. Computers and Mathematics with Applications, 50(10– 12):1639–1646, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004141.

Marasco:2000:LPM

[Mar00] A. Marasco. Lindstedt– Poincaré method and Mathematica applied to the motion of a solid with a fixed point. Computers and Mathematics with Applications, 40(2–3):333–343, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001644.

Marin:2005:MMS

[Mar05a]  L. Marin.  A meshless

method for solving the Cauchy problem in threedimensional elastostatics. Computers and Mathematics with Applications, 50(1– 2):73–92, July 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105002336.

Marszalek:2005:SCP

[MAR05b] W. Marszalek, T. Amdeberhan, and R. Riaza. Singularity crossing phenomena in DAEs: a two-phase fluid flow application case study. Computers and Mathematics with Applications, 49 (2–3):303–319, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000350.

Marz:2005:CDA

[Mär05c] R. März. Characterizing differential algebraic equations without the use of derivative arrays. Computers and Mathematics with Applications, 50(7):1141–1156, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003421.

Marhoune:2007:TPB

[Mar07] A. L. Marhoune. A threepoint boundary value problem with an integral twospace-variables condition for parabolic equations. Computers and Mathematics with Applications, 53(6): 940–947, March 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107000466.

Marsan:2009:MSC

[Mar09] G. Ajmone Marsan. On the modelling and simulation of the competition for a secession under media influence by active particles methods and functional subsystems decomposition. Computers and Mathematics with Applications, 57(5):710–728, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004768.

Massoudi:2007:BCM

[Mas07] Mehrdad Massoudi. Boundary conditions in mixture theory and in CFD applications of higher order models. Computers and Mathematics with Applications, 53(2):156–167, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000661.

Matic:2003:ISI

[Mat03]  M. Matić.  Improve-

ment of some inequalities of Euler-Grüss type. Computers and Mathematics with Applications, 46 (8–9):1325–1336, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902227.

Mayhew:2000:CHN

[May00] G. L. Mayhew. Clues to the hidden nature of de Bruijn sequences. Computers and Mathematics with Applications, 39(11): 57–65, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001097.

Mollaioli:2008:ISE

[MB08] F. Mollaioli and S. Bruno. Influence of site effects on inelastic displacement ratios for SDOF and MDOF systems. Computers and Mathematics with Applications, 55(2):184–207, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700346X.

Mahmoud:2008:RAC

[MBI08] Magdi S. Mahmoud, El-Kébir Boukas, and Abdulla Ismail. Robust adaptive control of uncertain discrete-time statedelay systems. Computers and Mathematics with Applications, 55(12):2887– 2902, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008115.

Maji:2003:SST

[MBR03] P. K. Maji, R. Biswas, and A. R. Roy. Soft set theory. Computers and Mathematics with Applications, 45(4–5):555–562, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000166.

See note [Yan08].

Medina:2002:ABS

[MC02] R. Medina and Sui Sun Cheng. The asymptotic behavior of the solutions of a discrete reaction-diffusion equation. Computers and Mathematics with Applications, 43(8–9):917–925, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800025.

Milovanovic:2007:GHI

[MC07] Gradimir V. Milovanović and Aleksandar S. Cvetković. Gauss–Hermite interval quadrature rule. Computers and Mathematics with Applications, 54(4):544–555, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002374.

Mani:2009:BCS

[MCC09] V. Mani, Pei Chann Chang, and Shih Hsin Chen. Bicriteria single machine scheduling problem with a learning effect: Aneja– Nair method to obtain the set of optimal sequences. Computers and Mathematics with Applications, 58(1): 39–47, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002818.

McDonald:2007:SBC

[McD07] Terry McDonald. Splines with boundary conditions. Computers and Mathematics with Applications, 54(9– 10):1234–1239, November 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107001940.

McGarry:2003:EDT

[McG03]  J. S. McGarry.  The

exact discrete time representation of a system of fourth-order differential equations. Computers and Mathematics with Applications, 46(2–3):213–230, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900265.

Mak:2001:SGT

[MCH01] M. K. Mak, H. W. Chan, and T. Harko. Solutions generating technique for Abel-type nonlinear ordinary differential equations. Computers and Mathematics with Applications, 41(10–11):1395–1401, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001043.

Marino:2009:SCM

[MCQK09] Giuseppe Marino, Vittorio Colao, Xiaolong Qin, and Shin Min Kang. Strong convergence of the modified Mann iterative method for strict pseudo-contractions. Computers and Mathematics with Applications, 57 (3):455–465, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108006330.

Moreno:2005:AIS

[MCRÁ05] J. Moreno, M. C. Casabán, and M. J. Rodríguez-Álvarez. An algorithm to initialize the search of solutions of polynomial systems. Computers and Mathematics with Applications, 50(5–6):919–933, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003251.

Milovanovic:2008:TOS

[MCS08] Gradimir V. Milovanović, Aleksandar S. Cvetković, and Marija P. Stanić. Trigonometric orthogonal systems and quadrature formulae. Computers and Mathematics with Applications, 56(11):2915–2931, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004598.

Mora:2000:FEG

[MCZ00] G. Mora, Y. Cherruault, and A. Ziadi. Functional equations generating spacedensifying curves. Computers and Mathematics with Applications, 39(9–10):45– 55, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000857.

Misra:2006:MMS

[MD06] J. C. Misra and B. Dravid.

A mathematical model in the study of genes for identifying transcription factor binding sites. Computers and Mathematics with Applications, 51(3–4):621–630, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005262.

Ma:2005:EIM

[MDG05] De-Xiang Ma, Zeng-Ji Du, and Wei-Gao Ge. Existence and iteration of monotone positive solutions for multipoint boundary value problem with p-Laplacian operator. Computers and Mathematics with Applications, 50(5–6):729–739, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003135.

Myers:2002:STR

[MDPD02] D. E. Myers, S. De Iaco, D. Posa, and L. De Cesare. Space-time radial basis functions. Computers and Mathematics with Applications, 43(3–5):539–549, February/March 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101003042.

Marinescu:2000:TSM

[ME00] D. Marinescu and A. Espesset. Tests of a simulation method for a system of Boltzmann equations. Computers and Mathematics with Applications, 40(6–7):805–812, September/October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000198X.

Muslu:2003:SSF

[ME03] G. M. Muslu and H. A. Erbay. A split-step Fourier method for the complex modified Korteweg–de Vries equation. Computers and Mathematics with Applications, 45(1–3):503–514, January/February 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800330.

Momani:2008:SNL

[ME08] Shaher Momani and Vedat Suat Ertürk. Solutions of non-linear oscillators by the modified differential transform method. Computers and Mathematics with Applications, 55(4):833–842, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005147.

Momani:2009:MHP

[MEA09] Shaher Momani, G. H. Erjaee, and M. H. Alnasr. The modified homotopy perturbation method for solving strongly nonlinear oscillators. Computers and Mathematics with Applications, 58(11–12):2209–2220, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900145X.

Medina:2001:ABV

[Med01] R. Medina. Asymptotic behavior of Volterra difference equations. Computers and Mathematics with Applications, 41(5–6):679– 687, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003126.

Mohsen:2008:GCM

[MEG08] Adel Mohsen and Mohamed El-Gamel. On the Galerkin and collocation methods for two-point boundary value problems using sinc bases. Computers and Mathematics with Applications, 56(4):930–941, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000953.

Merkle:2005:GRG

[Mer05] M. Merkle. Gurland’s ratio for the gamma function. Computers and Mathematics with Applications, 49(2–3):389–406, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000416.

Meszaros:2005:CFI

[Més05] C. Mészáros. The Cholesky factorization in interior point methods. Computers and Mathematics with Applications, 50(7):1157–1166, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003433.

Moghrabi:2001:NMF

[MF01] I. A. Moghrabi and J. A.

Ford. A nonlinear model for function-value multistep methods. Computers and Mathematics with Applications, 42(8– 9):1157–1164, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002292.

Misra:2001:MMS

[MG01] J. C. Misra and S. K.

Ghosh. A mathematical model for the study of interstitial fluid movement vis-a-vis the non-Newtonian behaviour of blood in a constricted artery. Computers and Mathematics with Applications, 41(5–6):783– 811, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003205.

Misra:2003:PFV

[MG03] J. C. Misra and S. K. Ghosh. Pulsatile flow of a viscous fluid through a porous elastic vessel of variable cross-section — a mathematical model for haemodynamic flows. Computers and Mathematics with Applications, 46(5–6):947–957, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901556.

Massouros:2005:AIL

[MG05] P. G. Massouros and G. M. Genin. Algebraic inversion of the Laplace transform. Computers and Mathematics with Applications, 50(1–2):179–185, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002415.

Ma:2008:SSF

[MG08a] Jingsen Ma and Wei Ge. Is standard symmetric formulation always better for smoothed particle hydrodynamics? Computers and Mathematics with Applications, 55(7):1503– 1513, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006311.

MirHassani:2008:MPP

[MG08b] S. A. MirHassani and M. Ghorbanalizadeh. The multi-product pipeline scheduling system. Computers and Mathematics with Applications, 56(4):891–897, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000904.

Mohamad:2008:UTS

[MG08c] Sannay Mohamad and K. Gopalsamy. A unified treatment for stability preservation in computer simulations of impulsive BAM networks. Computers and Mathematics with Applications, 55(9):2043– 2063, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006852.

Mahata:2005:JEL

[MGG05] G. C. Mahata, A. Goswami, and D. K. Gupta. A joint economic-lot-size model for purchaser and vendor in fuzzy sense. Computers and Mathematics with Applications, 50(10–12):1767–1790, November/December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105004232.

Mak:2002:NMG

[MH02] M. K. Mak and T. Harko. New method for generating general solution of Abel differential equation. Computers and Mathematics with Applications, 43(1–2):91–94, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002747.

Marinca:2007:PSS

[MH07] Vasile Marinca and Nicolae Herisanu. Periodic solutions for some strongly nonlinear oscillations by He’s variational iteration method. Computers and Mathematics with Applications, 54(7–8):1188–1196, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002891.

Marinca:2009:AAD

[MH09] Vasile Marinca and Nicolae Herisanu. An analytical approach to the dynamic analysis of a rotating electric machine. Computers and Mathematics with Applications, 58(11–12):2320–2324, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001588.

Matsunaga:2001:GAN

[MHS01] H. Matsunaga, T. Hara, and S. Sakata. Global attractivity for a nonlinear difference equation with variable delay. Computers and Mathematics with Applications, 41(5–6):543– 551, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002972.

Mattila:2008:CIL

[MHTR08] Keijo Mattila, Jari Hyväluoma, Jussi Timonen, and Tuomo Rossi. Comparison of implementations of the lattice-Boltzmann method. Computers and Mathematics with Applications, 55(7): 1514–1524, April 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107006232.

Mickens:2003:NFD

[Mic03] R. E. Mickens. A nonstandard finite difference scheme for a Fisher PDE having nonlinear diffusion. Computers and Mathematics with Applications, 45 (1–3):429–436, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103800287.

Michel:2008:RCC

[Mic08] Christian J. Michel. A 2006 review of circular codes in genes. Computers and Mathematics with Applications, 55(5):984–988, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005561.

Mickens:2009:RCM

[Mic09] Ronald E. Mickens. Reanalysis of a castor model limit cycle calculation. Computers and Mathematics with Applications, 57(5):740–743, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006342.

Mieussens:2001:CDV

[Mie01]  L. Mieussens.  Conver-

gence of a discrete-velocity model for the Boltzmann– BGK equation. Computers and Mathematics with Applications, 41(1–2):83–96, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101850082.

Migorski:2006:EHI

[Mig06]  S. Migórski.  Evolution

hemivariational inequality for a class of dynamic viscoelastic nonmonotone frictional contact problems. Computers and Mathematics with Applications, 52(5):677–698, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002483.

Mincsovics:2008:QAO

[Min08] Miklós Mincsovics. Qualitative analysis of onestep iterative methods and consistent matrix splittings. Computers and Mathematics with Applications, 55(10):2338–2345, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007316.

Manob:2004:STS

[MIPK04] A. H. Manob, T. M. Islam, M. T. Parvez, and M. Kaykobad. Search trees and Stirling numbers. Computers and Mathematics with Applications, 48(5–6):747–754, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003104.

Montalvo:2008:PSO

[MIPT08] Idel Montalvo, Joaquín Izquierdo, Rafael Pérez, and Michael M. Tung. Particle swarm optimization applied to the design of water supply systems. Computers and Mathematics with Applications, 56(3):769–776, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000771.

Mishra:2005:CMP

[Mis05]  S. K. Mishra.  Com-

plex minimax programming under generalized Type-I functions. Computers and Mathematics with Applications, 50(1–2): 1–11, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002270.

Mousa:2002:SIB

[MJ02] M. A. M. Ali Mousa and Z. F. Jaheen. Statistical inference for the Burr model based on progressively censored data. Computers and Mathematics with Applications, 43(10–11):1441–1449, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001104.

Masjed-Jamei:2008:GFT

[MJD08] Mohammad Masjed-Jamei and Mehdi Dehghan. A generalization of Fourier trigonometric series. Computers and Mathematics with Applications, 56(11): 2941–2947, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004549.

Mannan:2003:BHC

[MK03] M. Abdul Mannan and M. Kaykobad. Block Huffman coding. Computers and Mathematics with Applications, 46(10– 11):1581–1587, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901933.

Mukhopadhyay:2008:SGT

[MK08] Santwana Mukhopadhyay and Roushan Kumar. A study of generalized thermoelastic interactions in an unbounded medium with a spherical cavity. Computers and Mathematics with Applications, 56(9):2329–2339, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003714.

Mallet:2009:MWF

[MKF09] V. Mallet, D. E. Keyes, and F. E. Fendell. Modeling wildland fire propagation with level set methods. Computers and Mathematics with Applications, 57(7):1089–1101, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006329.

Moon:2008:ALM

[MKK08] Kyoung-Sook Moon, Won-Jung Kim, and Hongjoong Kim. Adaptive lattice methods for multi-asset models. Computers and Mathematics with Applications, 56(2):352–366, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108000217.

Michalis:2008:MMF

[MKSB08] V. K. Michalis, A. N.

Kalarakis, E. D. Skouras, and V. N. Burganos. Mesoscopic modeling of flow and dispersion phenomena in fractured solids. Computers and Mathematics with Applications, 55(7):1525– 1540, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006451.

Li:2008:NSC

[mLpW08] Yu mei Li and Xue ping Wang. Necessary and sufficient conditions for existence of maximal solutions for inf-α composite fuzzy relational equations. Computers and Mathematics with Applications, 55(9):1961– 1973, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006645.

Mo:2009:CDP

[MLW09] Yuanbin Mo, Hetong Liu, and Qin Wang. Conjugate direction particle swarm optimization solving systems of nonlinear equations. Computers and Mathematics with Applications, 57(11–12):1877– 1882, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005385.

Mariamma:2000:FNF

[MM00] N. K. Mariamma and S. N. Majhi. Flow of a Newtonian fluid in a blood vessel with permeable wall — a theoretical model. Computers and Mathematics with Applications, 40(12):1419–1432, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002509.

Mahapatra:2005:DPM

[MM05a] N. K. Mahapatra and M. Maiti. Decision process for multiobjective, multiitem production-inventory system via interactive fuzzy satisficing technique. Computers and Mathematics with Applications, 49(5–6): 805–821, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0898122105000933.

Miletics:2005:IET

[MM05b] E. Miletics and G. Molnárka.

Implicit extension of Taylor series method with numerical derivatives for initial value problems. Computers and Mathematics with Applications, 50(7):1167–1177, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003445.

Misra:2006:ISS

[MM06] J. C. Misra and A. Mitra.

Instabilities in single-species and host-parasite systems: Period-doubling bifurcations and chaos. Computers and Mathematics with Applications, 52(3–4):525– 538, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002203.

Makinde:2007:TSS

[MM07a] O. D. Makinde and P. Y.

Mhone. Temporal stability of small disturbances in MHD Jeffery– Hamel flows. Computers and Mathematics with Applications, 53(1):128–136, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001745.

Momoniat:2007:STF

[MM07b] E. Momoniat and D. P. Mason. Spreading of a thin film with suction or blowing including surface tension effects. Computers and Mathematics with Applications, 53(2):198–208, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700079X.

Maity:2008:NAM

[MM08a] K. Maity and M. Maiti. A numerical approach to a multi-objective optimal inventory control problem for deteriorating multi-items under fuzzy inflation and discounting. Computers and Mathematics with Applications, 55(8):1794–1807, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006529.

Mishra:2008:PDE

[MM08b] S. S. Mishra and P. P. Mishra. Price determination for an EOQ model for deteriorating items under perfect competition. Computers and Mathematics with Applications, 56(4):1082– 1101, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001107.

Misra:2008:SAT

[MM08c] J. C. Misra and A. Mitra. Synchronization among tumour-like cell aggregations coupled by quorum sensing: a theoretical study. Computers and Mathematics with Applications, 55(8):1842– 1853, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006566.

Mursaleen:2008:RCC

[MM08d] M. Mursaleen and S. A.

Mohiuddine. Regularly σconservative and σ-coercive four dimensional matrices. Computers and Mathematics with Applications, 56(6):1580–1586, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001843.

Maaranen:2004:QRI

[MMM04] H. Maaranen, K. Miettinen, and M. M. Mäkelä.

Quasi-random initial population for genetic algorithms. Computers and Mathematics with Applications, 47(12):1885–1895, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840240.

Moore:2001:ISN

[MN01] C. Moore and B. V. C. Nnoli. Iterative solution of nonlinear equations involving set-valued uniformly accretive operators. Computers and Mathematics with Applications, 42(1–2):131– 140, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001389.

Mizutani:2006:OPT

[MNIS06] Satoshi Mizutani, Toshio Nakagawa, Kodo Ito, and Hiroaki Sandoh. Optimal periodic testing policy for circuit with selftesting. Computers and Mathematics with Applications, 51(2):363–370, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004748.

Mydlarczyk:2001:PSN

[MO01] W. Mydlarczyk and W. Okrasiński.

Positive solutions to a nonlinear Abel-type integral equation on the whole line. Computers and Mathematics with Applications, 41(7–8):835–842, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003230.

Momani:2007:CBH

[MO07] Shaher Momani and Zaid Odibat. Comparison between the homotopy perturbation method and the variational iteration method for linear fractional partial differential equations. Computers and Mathematics with Applications, 54(7–8):910– 919, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002520.

Mobedi:2008:VHT

[MO08] Moghtada Mobedi and Hakan F. Oztop. Visualization of heat transport using dimensionless heatfunction for natural convection and conduction in an enclosure with thick solid ceiling. Computers and Mathematics with Applications, 56(10):2596–2608, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003970.

Moore:2002:DSI

[Moo02]  C. Moore.  A double-

sequence iteration process for fixed points of continuous pseudocontractions. Computers and Mathematics with Applications, 43(12):1585–1589, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001219.

Menegatto:2006:SPD

[MOP06] V. A. Menegatto, C. P. Oliveira, and A. P. Peron. Strictly positive definite kernels on subsets of the complex plane. Computers and Mathematics with Applications, 51(8):1233– 1250, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000769.

Moscariello:2008:QAS

[Mos08] G. Moscariello. Qualitative analysis and simulation of a 10 degrees of freedom model of an electrically operated vehicle. Computers and Mathematics with Applications, 55(8):1815– 1831, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006542.

Moudafi:2002:MEP

[Mou02] A. Moudafi. Mixed equilibrium problems: Sensitivity analysis and algorithmic aspect. Computers and Mathematics with Applications, 44(8– 9):1099–1108, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002183.

Mao:2002:EID

[MP02a] Guiyou Mao and L. R.

Petzold. Efficient integration over discontinuities for differential-algebraic systems. Computers and Mathematics with Applications, 43(1–2):65–79, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002723.

Misra:2002:PTB

[MP02b] J. C. Misra and S. K.

Pandey. Peristaltic transport of blood in small vessels: study of a mathematical model. Computers and Mathematics with Applications, 43(8–9):1183–1193, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800220.

Mishra:2005:GLO

[MP05] D. Mishra and P. Panigrahi. Graceful lobsters obtained by partitioning and component moving of branches of diameter four trees. Computers and Mathematics with Applications, 50(3–4):367–380, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002816.

Mainar:2007:GCB

[MP07a] E. Mainar and J. M. Peña. A general class of Bernsteinlike bases. Computers and Mathematics with Applications, 53(11):1686–1703, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001411.

Misra:2007:SSW

[MP07b] J. C. Misra and M. K. Patra. A study of solitary waves in a tapered aorta by using the theory of solitons. Computers and Mathematics with Applications, 54(2):242– 254, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002106.

Massoudi:2008:PFB

[MP08a] Mehrdad Massoudi and Tran X. Phuoc. Pulsatile flow of blood using a modified second-grade fluid model. Computers and Mathematics with Applications, 56(1):199– 211, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000059.

Mishra:2008:SGL

[MP08b] Debdas Mishra and Pratima Panigrahi. Some graceful lobsters with all three types of branches incident on the vertices of the central path. Computers and Mathematics with Applications, 56(5):1382–1394, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001624.

Michel:2008:VCF

[MPP08] Christian J. Michel, Giuseppe Pirillo, and Mario A. Pirillo. Varieties of commafree codes. Computers and Mathematics with Applications, 55(5):989–996, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005597.

Mathioudakis:2004:ISE

[MPS04] E. N. Mathioudakis, E. P. Papadopoulou, and Y. G. Saridakis. Iterative solution of elliptic collocation systems on a cognitive parallel computer. Computers and Mathematics with Applications, 48(5–6):951–970, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003268.

Mourrain:2006:DNR

[MPTV06] B. Mourrain, N. G. Pavlidis, D. K. Tasoulis, and M. N. Vrahatis. Determining the number of real roots of polynomials through neural networks. Computers and Mathematics with Applications, 51(3–4):527–536, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005195.

Matic:2000:IFG

[MPU00] M. Matić, J. Pecarić, and N. Ujević. Improvement and further generalization of inequalities of Ostrowski-Grüss type. Computers and Mathematics with Applications, 39(3–4):161–175, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003429.

Mishev:2002:RDE

[MPV02] D. P. Mishev, W. T. Patula, and H. D. Voulov. A reciprocal difference equation with maximum. Computers and Mathematics with Applications, 43(8–9):1021–1026, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800104.

Melamed:2007:SIG

[MPW07] Benjamin Melamed, Shuo Pan, and Yorai Wardi. Simulation of IPA gradients in hybrid network systems. Computers and Mathematics with Applications, 54(2): 161–182, July 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107002039.

Mrozek:2008:HAB

[MPZ08] Marian Mrozek, Pawel Pilarczyk, and Natalia Zelazna. Homology algorithm based on acyclic subspace. Computers and Mathematics with Applications, 55(11):2395–2412, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006943.

Ma:2005:SOS

[MQ05] Hang Ma and Qing-Hua Qin. A second-order scheme for integration of onedimensional dynamic analysis. Computers and Mathematics with Applications, 49(2–3):239–252, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000295.

Momani:2006:EMS

[MQ06] S. Momani and R. Qaralleh. An efficient method for solving systems of fractional integro-differential equations. Computers and Mathematics with Applications, 52(3–4):459–470, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002124.

Moges:2006:DLS

[MR06] M. A. Moges and T. G.

Robertazzi. Divisible load scheduling and Markov chain models. Computers and Mathematics with Applications, 52(10– 11):1529–1542, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003373.

Mora:2008:NPS

[MRÁR08] Cibeles Mora, María José Rodríguez-Álvarez, and José Vicente Romero. New pixellation scheme for CT algebraic reconstruction to exploit matrix symmetries. Computers and Mathematics with Applications, 56(3):715–726, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000746.

Maji:2002:ASS

[MRB02] P. K. Maji, A. R. Roy, and R. Biswas. An application of soft sets in a decision making problem.

Computers and Mathematics with Applications, 44 (8–9):1077–1083, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210200216X.

Muniz:2000:ETB

[MRdCV00] W. B. Muniz, F. M. Ramos, and H. F. de Campos Velho. Entropy- and Tikhonov-based regularization techniques applied to the backwards heat equation. Computers and Mathematics with Applications, 40 (8–9):1071–1084, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850178.

Moretti:2002:EVI

[MRMRM02] A. C. Moretti, M. A. Rojas-Medar, and M. Drina Rojas-Medar. The equations of a viscous incompressible chemically active fluid: existence and uniqueness of strong solutions in an unbounded domain. Computers and Mathematics with Applications, 44(3–4):287– 299, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001487.

Montuori:2008:ITA

[MRP08] Alfonso Montuori, Giovanni Raimondo, and Eros Pasero. An information theoretic approach for improving data driven prediction of protein model quality. Computers and Mathematics with Applications, 55(5):997–1006, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005585.

Marinkovic:2008:IST

[MRS08] Sladjana Marinković, Predrag Rajković, and Miomir Stanković. The inequalities for some types of qintegrals. Computers and Mathematics with Applications, 56(10):2490–2498, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003854.

Malaguti:2004:BSW

[MRT04] L. Malaguti, P. Rehák, and V. Taddei. Bounded solutions and wavefronts for discrete dynamics. Computers and Mathematics with Applications, 47(6–7):1079– 1094, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104900880.

Martinsson:2005:FAI

[MRT05] P. G. Martinsson, V. Rokhlin, and M. Tygert. A fast algorithm for the inversion of general Toeplitz matrices. Computers and Mathematics with Applications, 50(5–6):741–752, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003147.

Murthy:2000:TDE

[MS00] P. V. S. N. Murthy and P. Singh. Thermal dispersion effects on non-Darcy convection over a cone. Computers and Mathematics with Applications, 40(12):1433–1444, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002510.

Meng:2002:OLH

[MS02] Fanwei Meng and Yuangong Sun. Oscillation of linear Hamiltonian systems. Computers and Mathematics with Applications, 44(10– 11):1467–1477, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002717.

Mendivil:2003:CGW

[MS03] F. Mendivil and J. Silver.

Chaos games for wavelet analysis and wavelet synthesis. Computers and Mathematics with Applications, 46(5–6):863–873, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901490.

Maischak:2004:LSC

[MS04a] M. Maischak and E. P.

Stephan. A least squares coupling method with finite elements and boundary elements for transmission problems. Computers and Mathematics with Applications, 48(7–8):995–1016, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003426.

Mueller:2004:NSG

[MS04b] J. L. Mueller and T. S.

Shores. A new sinc-Galerkin method for convection– diffusion equations with mixed boundary conditions.

Computers and Mathematics with Applications, 47 (4–5):803–822, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900661.

Mallet:2008:AQM

[MS08] Vivien Mallet and Bruno Sportisse. Air quality modeling: From deterministic to stochastic approaches. Computers and Mathematics with Applications, 55(10):2329–2337, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700733X.

Mosheiov:2009:SMA

[MS09] Gur Mosheiov and Assaf Sarig. Scheduling a maintenance activity to minimize total weighted completion-time. Computers and Mathematics with Applications, 57(4):619–623, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006846.

Milovanovic:2000:MVM

[MSN^{+}00] E. I. Milovanović, M. K. Stojčev, N. M. Novaković, I. Ž. Milovanović, and T. I. Tokić. Matrix–vector multiplication on a fixed-size linear systolic array. Computers and Mathematics with Applications, 40(10– 11):1189–1203, November/ December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002315.

Montano:2008:NMP

[MSS08] Emedin Montaño, Mario Salas, and Ricardo L. Soto. Nonnegative matrices with prescribed extremal singular values. Computers and Mathematics with Applications, 56(1):30– 42, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700822X.

Moon:2000:MNM

[MT00] I. H. Moon and R. P.

Tewarson. Multi-nephron and multi-vasa recta models of the inner medullary renal concentrating mechanism. Computers and Mathematics with Applications, 40(2–3):345–357, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001656.

Mosbach:2009:QPI

[MT09] Sebastian Mosbach and Amanda G. Turner. A quantitative probabilistic investigation into the accumulation of rounding errors in numerical ODE solution. Computers and Mathematics with Applications, 57(7):1157–1167, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000431.

Melnik:2003:DGA

[MUHdH03] R. V. N. Melnik, A. Uhlherr, J. Hodgkin, and F. de Hoog. Distance geometry algorithms in molecular modelling of polymer and composite systems. Computers and Mathematics with Applications, 45(1–3): 515–534, January/February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103800342.

Murea:2005:BAN

[Mur05a] C. M. Murea. The BFGS algorithm for a nonlinear least squares problem arising from blood flow in arteries. Computers and Mathematics with Applications, 49(2–3):171–186, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500026X.

Muroya:2005:GSS

[Mur05b] Y. Muroya. Global stability for separable nonlinear delay differential equations. Computers and Mathematics with Applications, 49(11–12):1913– 1927, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002087.

Murio:2006:SNE

[Mur06] D. A. Murio. On the stable numerical evaluation of Caputo fractional derivatives. Computers and Mathematics with Applications, 51(9–10):1539– 1550, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001003.

Murio:2007:SNS

[Mur07a]  Diego A. Murio.  Sta-

ble numerical solution of a fractional-diffusion inverse heat conduction problem. Computers and Mathematics with Applications, 53(10):1492–1501, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001174.

Muroya:2007:GAD

[Mur07b] Yoshiaki Muroya. Global attractivity for discrete models of nonautonomous logistic equations. Computers and Mathematics with Applications, 53(7):1059– 1073, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000636.

Murio:2008:IFD

[Mur08a] Diego A. Murio. Implicit finite difference approximation for time fractional diffusion equations. Computers and Mathematics with Applications, 56(4):1138– 1145, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001156.

See [fDxZ11].

Murio:2008:TFI

[Mur08b] Diego A. Murio. Time fractional IHCP with Caputo fractional derivatives. Computers and Mathematics with Applications, 56(9):2371–2381, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003763.

Muslim:2006:ASH

[Mus06]  M. Muslim.  Approx-

imation of solutions to history-valued neutral functional differential equations. Computers and Mathematics with Applications, 51(3–4):537–550, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005201.

Marinov:2008:IPC

[MV08a] Tchavdar T. Marinov and Aghalaya S. Vatsala. Inverse problem for coefficient identification in the Euler–Bernoulli equation. Computers and Mathematics with Applications, 56(2): 400–410, July 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108000254.

Metsidik:2008:EVP

[MV08b] Metrose Metsidik and Elkin Vumar. Edge vulnerability parameters of bisplit graphs. Computers and Mathematics with Applications, 56(7):1741–1747, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002526.

Mollica:2007:ICE

[MVS^{+}07] Francesco Mollica, Maurizio Ventre, Fernando Sarracino, Luigi Ambrosio, and Luigi Nicolais. Implicit constitutive equations in the modeling of bimodular materials: an application to biomaterials. Computers and Mathematics with Applications, 53(2):209–218, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000855.

Mistiri:2000:EPI

[MW00] F. S. Mistiri and A. P. Wang. Estimation problems in an input-and-output system. Computers and Mathematics with Applications, 39(12): 29–42, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001279.

Mai:2008:HHG

[MWH08] Ta-Cheng Mai, Jeng-Jung Wang, and Lih-Hsing Hsu. Hyper-Hamiltonian generalized Petersen graphs. Computers and Mathematics with Applications, 55(9): 2076–2085, May 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107006876.

Wang:2009:NES

[mWmZ09] Tong ming Wang and Jia min Zhu. New explicit solutions of the Klein–Gordon equation using the variational iteration method combined with the Exp-function method. Computers and Mathematics with Applications, 58(11–12):2444–2448, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001874.

Ma:2009:UTS

[MWZ09] Lin Ma, Kuanquan Wang, and David Zhang. A universal texture segmentation and representation scheme based on ant colony optimization for iris image processing. Computers and Mathematics with Applications, 57(11–12):1862– 1868, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005415.

Ma:2009:CDD

[MXZL09] Runnian Ma, Yu Xie, Shengrui Zhang, and Wenbin Liu. Convergence of discrete delayed Hopfield neural networks. Computers and Mathematics with Applications, 57(11–12):1869– 1876, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005397.

Ma:2005:GA

[MY05] Manjun Ma and Jianshe Yu. Global attractivity of x_{n+1} = (1 − α)x_{n} + βe^{−γx_{n−k}}. Computers and Mathematics with Applications, 49(9–10):1397– 1402, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001586.

Ma:2009:AEM

[MYG09] Hong-Cai Ma, Yao-Dong Yu, and Dong-Jie Ge. The auxiliary equation method for solving the Zakharov– Kuznetsov (ZK) equation. Computers and Mathematics with Applications, 58(11–12):2523–2527, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001977.

Moges:2009:GSD

[MYR09] M. A. Moges, D. Yu, and T. G. Robertazzi. Grid scheduling divisible loads from two sources. Computers and Mathematics with Applications, 58(6):1081– 1092, September 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109004921.

Marlewski:2004:IMP

[MZ04] A. Marlewski and P. Zarzycki. Infinitely many positive solutions of the Diophantine equation x^{2}−kxy+ y^{2} + x = 0. Computers and Mathematics with Applications, 47(1):115–121, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900107.

Mei:2007:CTV

[MZ07] Shu-Li Mei and Sen-Wen Zhang. Coupling technique of variational iteration and homotopy perturbation methods for nonlinear matrix differential equations. Computers and Mathematics with Applications, 54(7–8):1092–1100, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700288X.

Ma:2009:EPS

[MZ09] Ruyun Ma and Bao Zhu. Existence of positive solutions for a semipositone boundary value problem on the half-line. Computers and Mathematics with Applications, 58(8):1672–1686, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004210.

Ma:2009:SKF

[MZD09] Xueling Ma, Jianming Zhan, and Wiesław A. Dudek. Some kinds of (\overline{∈}, \overline{∈} ∨ \overline{q})-fuzzy filters of BL-algebras. Computers and Mathematics with Applications, 58(2):248– 256, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002909.

Nabli:2004:PAD

[Nab04]  H. Nabli.  Performabil-

ity: asymptotic distribution and moment computation. Computers and Mathematics with Applications, 48(1– 2):1–8, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839944.

Nadarajah:2005:EBD

[Nad05]  S. Nadarajah.  Expo-

nentiated beta distributions. Computers and Mathematics with Applications, 49(7–8):1029–1035, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001197.

Nakai:2006:PJS

[Nak06]  Toru Nakai.  Properties

of a job search problem on a partially observable Markov chain in a dynamic economy. Computers and Mathematics with Applications, 51(2):189–198, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004566.

Nambiar:2000:IST

[Nam00]  K. K. Nambiar.  Intu-

itive set theory. Computers and Mathematics with Applications, 39(1–2):183– 185, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003223.

Nambiar:2001:TSE

[Nam01a] K. K. Nambiar. Theory of search engines. Computers and Mathematics with Applications, 42(12):1523– 1526, December 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122101002590.

Nambiar:2001:VIS

[Nam01b]  K. K. Nambiar.  Visu-

alization of intuitive set theory. Computers and Mathematics with Applications, 41(5–6):619–626, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003059.

Nambiar:2002:WHB

[Nam02] K. K. Nambiar. White hole, black whole, and the book. Computers and Mathematics with Applications, 44 (10–11):1325–1327, November/December 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102002596.

Nicolau:2008:SSC

[NB08] Dan V. Nicolau, Jr. and Kevin Burrage. Stochastic simulation of chemical reactions in spatially complex media. Computers and Mathematics with Applications, 55(5):1007– 1018, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700569X.

Nie:2009:SAG

[NC09] Xiaobing Nie and Jinde Cao. Stability analysis for the generalized Cohen– Grossberg neural networks with inverse Lipschitz neuron activations. Computers and Mathematics with Applications, 57(9):1522– 1536, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000327.

Nunokawa:2008:DIC

[NCK^{+}08] Mamoru Nunokawa, Nak Eun Cho, Oh Sang Kwon, Shigeyoshi Owa, and Hitoshi Saitoh. Differential inequalities for certain analytic functions. Computers and Mathematics with Applications, 56(11):2908–2914, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004859.

Namekata:2001:BPN

[ND01] T. Namekata and T. S. H. Driessen. Bargaining property of nucleolus and τvalue in a class of TUgames. Computers and Mathematics with Applications, 41(5–6):703–721, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003151.

Nadarajah:2005:CPT

[ND05a] S. Nadarajah and D. K. Dey. Convolutions of the Pearson Type VII distribution. Computers and Mathematics with Applications, 50(3–4):339–346, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002786.

Nadarajah:2005:CD

[ND05b] S. Nadarajah and D. K. Dey. Convolutions of the T distribution. Computers and Mathematics with Applications, 49(5–6):715–721, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000854.

Nesemann:2001:ILF

[Nes01]  T. Nesemann.  Invari-

ants and Liapunov functions for nonautonomous systems. Computers and Mathematics with Applications, 42(3–5):385–392, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001638.

Neta:2005:SSS

[Net05] B. Neta. P -stable symmetric super-implicit methods for periodic initial value problems. Computers and Mathematics with Applications, 50(5–6):701–705, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500310X.

Neta:2006:P

[Net06] B. Neta. Preface. Computers and Mathematics with Applications, 52(8–9): xiii–xv, October/November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0898122106003130.

Neta:2007:PSH

[Net07]  Beny Neta.  P-stable

high-order super-implicit and Obrechkoff methods for periodic initial value problems. Computers and Mathematics with Applications, 54(1):117–126, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001915.

Neta:2003:OVS

[NF03] B. Neta and T. Fukushima. Obrechkoff versus superimplicit methods for the solution of first- and secondorder initial value problems. Computers and Mathematics with Applications, 45(1–3):383–390, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210380024X.

Nakagawa:2003:OCI

[NFI03] S. Nakagawa, S. Fukumoto, and N. Ishii. Optimal checkpointing intervals for a double modular redundancy with signatures. Computers and Mathematics with Applications, 46(7):1089–1094, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901234.

Nadarajah:2004:C

[NG04] S. Nadarajah and A. K.

Gupta. Convolutions of K/(1 + χ^{n}). Computers and Mathematics with Applications, 48(5–6):813–822, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003153.

Nicolaides:2004:MCG

[Nic04] R. Nicolaides. A method for complex geometries in finite-difference solutions of Maxwell’s equations. Computers and Mathematics with Applications, 48(7–8):1111–1119, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003499.

Niederhausen:2001:GSS

[Nie01a]  H. Niederhausen.  Gen-

eralized Sheffer sequences satisfying piecewise functional conditions. Computers and Mathematics with Applications, 41(9):1155– 1171, May 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101000888.

Niell:2001:SAP

[Nie01b] A. M. Niell. The simultaneous approximation of polynomial roots. Computers and Mathematics with Applications, 41(1–2):1–14, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101850021.

Nie:2006:NPM

[Nie06]  Pu-Yan Nie.  A new

penalty method for nonlinear programming. Computers and Mathematics with Applications, 52(6– 7):883–896, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210600263X.

Niezgoda:2009:NIO

[Nie09] Marek Niezgoda. A new inequality of Ostrowski-Grüss type and applications to some numerical quadrature rules. Computers and Mathematics with Applications, 58(3):589–596, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002521.

Nikolic:2006:SMD

[Nik06] Z. S. Nikolic. A simple model for 3-D visualization of model topology due to rearrangement in liquid phase sintering. Computers and Mathematics with Applications, 51(3–4):551–558, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005213.

Neta:2008:HON

[NJ08] B. Neta and Anthony N.

Johnson. High-order nonlinear solver for multiple roots. Computers and Mathematics with Applications, 55(9):2012–2017, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006694.

Nambiar:2001:GDD

[NK01] K. K. Nambiar and V. Kannoth. Generic dependencies and database design. Computers and Mathematics with Applications, 41(3–4):281–288, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002728.

Narupiyakul:2008:FET

[NKCS08] Lalita Narupiyakul, Vlado Keselj, Nick Cercone, and Booncharoen Sirinaovakul. Focus to emphasize tone analysis for prosodic generation. Computers and Mathematics with Applications, 55(8):1735–1753, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006487.

Nkonga:2000:ILP

[Nko00] B. Nkonga. Implicit linearity preservation type schemes for moving meshes. Computers and Mathematics with Applications, 39 (9–10):201–220, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100000961.

Naadimuthu:2007:AAN

[NLL07] G. Naadimuthu, D. M. Liu, and E. S. Lee. Application of an adaptive neural fuzzy inference system to thermal comfort and group technology problems. Computers and Mathematics with Applications, 54(11–12):1395–1402, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003756.

See comments [ES08b, NLL08].

Naadimuthu:2008:CCA

[NLL08] G. Naadimuthu, D. M. Liu, and E. S. Lee. Comment on ‘Comment on: “Application of an adaptive neural fuzzy inference system to thermal comfort and group technology problems” [Comput. Math. Appl. 54 (2007) 1395–1402] by Tarkan Erdik and Zekal Sen’. Computers and Mathematics with Applications, 56(7):1903–1904, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003143.

See [NLL07, ES08b].

Niu:2009:MLP

[NLLL09] Dongxiao Niu, Jinchao Li, Jinying Li, and Da Liu. Middle-long power load forecasting based on particle swarm optimization. Computers and Mathematics with Applications, 57(11– 12):1883–1889, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S089812210800535X.

Nakao:2000:NVS

[NLR00] M. T. Nakao, S. H. Lee, and Cheon Seoung Ryoo. Numerical verification of solutions for elasto-plastic torsion problems. Computers and Mathematics with Applications, 39(3–4):195–204, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003454.

Noor:2007:EMF

[NMD07] Muhammad Aslam Noor and Syed Tauseef Mohyud-Din. An efficient method for fourth-order boundary value problems. Computers and Mathematics with Applications, 54(7–8):1101–1111, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002945.

Noor:2008:HPM

[NMD08] Muhammad Aslam Noor and Syed Tauseef Mohyud-Din. Homotopy perturbation method for solving sixth-order boundary value problems. Computers and Mathematics with Applications, 55(12):2953– 2972, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008176.

Noor:2009:VIM

[NMD09] Muhammad Aslam Noor and Syed Tauseef Mohyud-Din. Variational iteration method for unsteady flow of gas through a porous medium using He’s polynomials and Pade approximants. Computers and Mathematics with Applications, 58(11–12):2182–2189, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001485.

Nobakhtian:2006:IFN

[Nob06]  S. Nobakhtian.  Infine

functions and nonsmooth multiobjective optimization problems. Computers and Mathematics with Applications, 51(9–10):1385– 1394, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001076.

Nunokawa:2008:DSA

[NON^{+}08] Mamoru Nunokawa, Shigeyoshi Owa, Junichi Nishiwaki, Kazuo Kuroki, and Toshio Hayami. Differential subordination and argumental property. Computers and Mathematics with Applications, 56(10):2733–2736, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004136.

Noor:2000:IMG

[Noo00a]  M. A. Noor.  An it-

erative method for general mixed variational inequalities. Computers and Mathematics with Applications, 40(2–3):171–176, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001516.

Noor:2000:PSA

[Noo00b] M. A. Noor. Projectionsplitting algorithms for monotone variational inequalities. Computers and Mathematics with Applications, 39(9–10):73–79, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000870.

Noor:2002:SAF

[Noo02a] M. A. Noor. Sensitivity analysis framework for general quasi-variational inclusions. Computers and Mathematics with Applications, 44(8–9):1175–1181, October/November 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102002249.

Noor:2002:SMP

[Noo02b] M. A. Noor. Stability of the modified projected dynamical systems. Computers and Mathematics with Applications, 44(1–2): 1–5, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001256.

Navarra:2001:RGC

[NOY01] A. Navarra, P. L. Odell, and D. M. Young. A representation of the general common solution to the matrix equations A_{1}XB_{1} = C_{1} and A_{2}XB_{2} = C_{2} with applications. Computers and Mathematics with Applications, 41(7–8):929– 935, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003308.

Novotny:2007:SEC

[NP07] A. Novotný and M. Pokorný.

Stabilization to equilibria of compressible Navier– Stokes equations with infinite mass. Computers and Mathematics with Applications, 53(3–4):437–451, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001873.

Nakamura:2003:OMT

[NQHN03] S. Nakamura, C. Qian, I. Hayashi, and T. Nakagawa. An optimal maintenance time of automatic monitoring system of ATM with two kinds of breakdowns. Computers and Mathematics with Applications, 46(7):1095–1101, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901246.

Natalini:2001:NSR

[NR01] P. Natalini and P. E. Ricci. Newton sum rules of polynomials defined by a threeterm recurrence relation. Computers and Mathematics with Applications, 42 (6–7):767–771, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001961.

Natalini:2004:EBP

[NR04] P. Natalini and P. E.

Ricci. An extension of the bell polynomials. Computers and Mathematics with Applications, 47 (4–5):719–725, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900594.

Nieto:2000:EAS

[NRL00] J. J. Nieto and R. Rodríguez-López. Existence and approximation of solutions for nonlinear functional differential equations with periodic boundary value conditions. Computers and Mathematics with Applications, 40(4–5):433–442, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001711.

Nieto:2006:MMF

[NRL06] J. J. Nieto and R. Rodríguez-López. Monotone method for first-order functional differential equations. Computers and Mathematics with Applications, 52(3–4):471– 484, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002136.

Nieto:2008:BVP

[NRL08] Juan J. Nieto and Rosana Rodríguez-López. Boundary value problems for a class of impulsive functional equations. Computers and Mathematics with Applications, 55(12):2715– 2731, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007717.

Nabli:2000:PAD

[NS00] H. Nabli and B. Sericola. Performability analysis for degradable computer systems. Computers and Mathematics with Applications, 39(3–4):217–234, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003478.

Nadau:2007:NSS

[NS07] Lionel Nadau and Adélia Sequeira. Numerical simulations of shear dependent viscoelastic flows with a combined finite element-finite volume method. Computers and Mathematics with Applications, 53(3–4):547–568, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001733.

Nechaev:2008:MIS

[NSB08] O. V. Nechaev, E. P. Shurina, and M. A. Botchev. Multilevel iterative solvers for the edge finite element solution of the 3D Maxwell equation. Computers and Mathematics with Applications, 55(10):2346– 2362, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007328.

Necco:2002:UEN

[NSQ02] C. Necco, M. Silvestri, and L. Quintas. Uniqueness and efficiency of Nash equilibrium in a family of randomly generated repeated games. Computers and Mathematics with Applications, 43(10–11):1425–1440, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001098.

Ng:2001:NSA

[NST01] E. Y. K. Ng, W. L. Siauw, and H. M. Tsai. Nonlinear solution adaptive structured grids via heat-conduction analogies. Computers and Mathematics with Applications, 41(1–2):193–213, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101850161.

Novak:2002:RGT

[NT02] Á. B. Novák and Zs. Tuza. Reconstruction graphs and testing their properties in a relational spatial database. Computers and Mathematics with Applications, 43 (10–11):1391–1406, May/ June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001074.

Napoli:2008:DNE

[NT08] Gaetano Napoli and Stefano Turzi. On the determination of nontrivial equilibrium configurations close to a bifurcation point. Computers and Mathematics with Applications, 55(2):299–306, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003550.

Nie:2009:PSN

[NTHP09] Linfei Nie, Zhidong Teng, Lin Hu, and Jigen Peng. Permanence and stability in non-autonomous predator– prey Lotka–Volterra systems with feedback controls. Computers and Mathematics with Applications, 58(3):436–448, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003277.

Naresh:2009:SAT

[NTTS09] Ram Naresh, Agraj Tripathi, J. M. Tchuenche, and Dileep Sharma. Stability analysis of a time delayed SIR epidemic model with nonlinear incidence rate. Computers and Mathematics with Applications, 58(2): 348–359, July 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109002880.

Naruse:2006:OCI

[NUN06] Kenichiro Naruse, Shizuka Umemura, and Sayori Nakagawa. Optimal checkpointing interval for two-level recovery schemes. Computers and Mathematics with Applications, 51(2):371–376, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/089812210500475X.

Noutsos:2008:SCP

[NV08] D. Noutsos and P. Vassalos. Superlinear convergence for PCG using band plus algebra preconditioners for Toeplitz systems. Computers and Mathematics with Applications, 56(5):1255–1270, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001491.

Natesan:2003:NAS

[NVAR03] S. Natesan, J. Vigo-Aguiar, and N. Ramanujam. A numerical algorithm for singular perturbation problems exhibiting weak boundary layers. Computers and Mathematics with Applications, 45(1–3):469–479, January/February 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103800317.

Noor:2009:SIM

[NW09] Muhammad Aslam Noor and Muhammad Waseem. Some iterative methods for solving a system of nonlinear equations. Computers and Mathematics with Applications, 57(1):101–106, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006196.

Ngatchou-Wandji:2008:LPC

[NWL08] Joseph Ngatchou-Wandji and Nâamane Laïb. Local power of a Cramér-von Mises type test for parametric autoregressive models of order one. Computers and Mathematics with Applications, 56(4):918–929, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000941.

Oger:2008:SCO

[OAVB08] L. Oger, M. Ammi, A. Valance, and D. Beladjine. Study of the collision of one rapid sphere on 3D packings: Experimental and numerical results. Computers and Mathematics with Applications, 55(2):132–148, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003422.

Ocalan:2007:OFN

[Öca07]  Özkan Öcalan.  Oscilla-

tion of forced neutral differential equations with positive and negative coefficients. Computers and Mathematics with Applications, 54(11–12):1411–1421, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003318.

See erratum [KÖ08a].

Okamura:2000:OPC

[ODO00] H. Okamura, T. Dohi, and S. Osaki. Optimal policies for a controlled queueing system with removable server under a random vacation circumstance. Computers and Mathematics with Applications, 39(1–2):215– 227, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003259.

Okamura:2003:SAM

[ODO03] H. Okamura, T. Dohi, and S. Osaki. A structural approximation method to generate the optimal autosleep schedule for computer systems. Computers and Mathematics with Applications, 46(7):1103–1110, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901258.

Otero-Espinar:2009:UER

[OEV09] Victoria Otero-Espinar and Dolores R. Vivero. Uniqueness and existence results for initial value problems on time scales through a reciprocal problem and applications. Computers and Mathematics with Applications, 58(4):700–710, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002387.

Ouyang:2008:SIM

[OF08] Yao Ouyang and Jinxuan Fang. Sugeno integral of monotone functions based on Lebesgue measure. Computers and Mathematics with Applications, 56(2):367–374, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108000229.

Osilike:2000:WSC

[OI00] M. O. Osilike and D. I. Igbokwe. Weak and strong convergence theorems for fixed points of pseudocontractions and solutions of monotone type operator equations. Computers and Mathematics with Applications, 40(4–5):559–567, August/September 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100001796.

Onyango:2008:RHT

[OIL08] T. T. M. Onyango, D. B. Ingham, and D. Lesnic. Reconstruction of heat transfer coefficients using the boundary element method. Computers and Mathematics with Applications, 56(1):114– 126, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107008310.

Onyango:2009:RBC

[OIL09] T. T. M. Onyango, D. B. Ingham, and D. Lesnic. Reconstruction of boundary condition laws in heat conduction using the boundary element method. Computers and Mathematics with Applications, 57(1):153–168, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004744.

Osabe:2006:OCW

[OK06] Minoru Osabe and Mutsuto Kawahara. An optimal control of water level considering time-delay system. Computers and Mathematics with Applications, 52 (8–9):1317–1324, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003208.

Onishi:2008:LBS

[OKCO08] J. Onishi, A. Kawasaki, Y. Chen, and H. Ohashi. Lattice Boltzmann simulation of capillary interactions among colloidal particles. Computers and Mathematics with Applications, 55(7):1541–1553, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006463.

Oh:2001:SCD

[OLAS01] M. Oh, Y. Lee, S. V. Azarov, and V. I. Shin. Suboptimal continuous-discrete filtering based on decomposition of observations. Computers and Mathematics with Applications, 42(1–2):189– 197, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001432.

Ohm:2006:EEF

[OLS06] M. R. Ohm, H. Y. Lee, and J. Y. Shin. Error estimates of finite-element approximations for a fourth-order differential equation. Computers and Mathematics with Applications, 52(3–4):283– 288, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001969.

Odibat:2009:VIM

[OM09] Zaid Odibat and Shaher Momani. The variational iteration method: an efficient scheme for handling fractional partial differential equations in fluid mechanics. Computers and Mathematics with Applications, 58(11–12):2199–2208, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001436.

Oden:2001:GOE

[OP01] J. T. Oden and S. Prudhomme. Goal-oriented error estimation and adaptivity for the finite element method. Computers and Mathematics with Applications, 41(5–6):735– 756, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003175.

Osman:2001:MSC

[OR01] E.-S. Osman and S. Z.

Rida. Multisoliton solution of a class of coupled soliton equations. Computers and Mathematics with Applications, 41(10–11):1425–1430, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001079.

Oros:2002:IOS

[Oro02]  G. Oros.  Integral op-

erator and starlike functions. Computers and Mathematics with Applications, 43(10–11):1257–1260, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102000974.

Oliveri:2008:WHC

[OS08a] F. Oliveri and M. P.

Speciale. Wave hierarchies in continua with scalar microstructure in the plane and spherical symmetry. Computers and Mathematics with Applications, 55(2):285–298, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003549.

Ozden:2008:IFE

[OS08b] Hacer Ozden and Yilmaz Simsek. Interpolation function of the (h, q)- extension of twisted Euler numbers. Computers and Mathematics with Applications, 56(4):898–908, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000928.

Osilike:2009:EAC

[OS09] M. O. Osilike and Y. Shehu. Explicit averaging cyclic algorithm for common fixed points of a finite family of asymptotically strictly pseudocontractive maps in Banach spaces. Computers and Mathematics with Applications, 57(9):1502– 1510, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000455.

Osaki:2003:P

[Osa03]  Shunji Osaki.  Preface.

Computers and Mathematics with Applications, 46(7): xiii, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901118.

Ogden:2007:CAW

[OSS07] Ray W. Ogden, Giuseppe Saccomandi, and Ivonne Sgura. Computational aspects of worm-like-chain interpolation formulas. Computers and Mathematics with Applications, 53(2): 276–286, January 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107001058.

Owa:2004:GPS

[OSSY04] S. Owa, H. Saitoh, H. M.

Srivastava, and R. Yamakawa. Geometric properties of solutions of a class of differential equations. Computers and Mathematics with Applications, 47(10–11):1689–1696, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839828.

Oduntan:2008:MTS

[OTB^{+}08] Idowu O. Oduntan, Michel Toulouse, Richard Baumgartner, Christopher Bowman, Ray Somorjai, and Teodor G. Crainic. A multilevel tabu search algorithm for the feature selection problem in biomedical data. Computers and Mathematics with Applications, 55(5):1019–1033, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005664.

Ou:2008:APS

[Ou08] Chunxia Ou. Anti-periodic solutions for high-order Hopfield neural networks. Computers and Mathematics with Applications, 56(7): 1838–1844, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108002630.

Ong:2006:BPA

[OY06] Ghim Hwee Ong and Kai Yang. A binary partitioning approach to image compression using weighted finite automata for large images. Computers and Mathematics with Applications, 51(11):1705–1714, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001374.

Ozis:2007:AHS

[ÖY07a] Turgut Özis and Ahmet Yildirim. Application of He’s semi-inverse method to the nonlinear Schrödinger equation. Computers and Mathematics with Applications, 54(7–8):1039–1042, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002702.

Ozis:2007:DFA

[ÖY07b] Turgut Özis and Ahmet Yildirim. Determination of the frequencyamplitude relation for a Duffing-harmonic oscillator by the energy balance method. Computers and Mathematics with Applications, 54(7–8):1184–1187, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002878.

Ozis:2008:CBA

[ÖY08] Turgut Özis and Ahmet Yildirim. Comparison between Adomian’s method and He’s homotopy perturbation method. Computers and Mathematics with Applications, 56(5):1216–1224, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001235.

Ozkan:2009:GTI

[ÖY09] Umut Mutlu Özkan and Hüseyin Yildirim. Grüss type inequalities for double integrals on time scales. Computers and Mathematics with Applications, 57 (3):436–444, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108006810.

Ozbekler:2007:FOS

[ÖZ07] A. Özbekler and A. Zafer. Forced oscillation of superhalf-linear impulsive differential equations. Computers and Mathematics with Applications, 54(6):785–792, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002696.

Ozer:2008:ASG

[Öze08] Teoman Özer. An application of symmetry groups to nonlocal continuum mechanics. Computers and Mathematics with Applications, 55(9):1923–1942, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700661X.

Ouyang:2005:OCN

[OZY05] Zigen Ouyang, Shengfan Zhou, and Fuqi Yin. Oscillation for a class of neutral parabolic differential equations. Computers and Mathematics with Applications, 50(1–2):145– 155, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002385.

Padula:2006:RFW

[Pad06] M. Padula. The role of free work identity in viscous compressible flows. Computers and Mathematics with Applications, 51(5):849–856, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000216.

Palamides:2001:PSH

[Pal01] P. K. Palamides. Positive solutions for higher-order Lidstone boundary value problems. a new approach via Sperner’s lemma. Computers and Mathematics with Applications, 42(1–2):75– 89, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001328.

Pan:2000:PSA

[Pan00] Ping-Qi Pan. A projective simplex algorithm using LU decomposition. Computers and Mathematics with Applications, 39(1–2):187– 208, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003235.

Pan:2001:NPT

[Pan01]  V. Y. Pan.  A new

proximity test for polynomial zeros. Computers and Mathematics with Applications, 41(12):1559– 1560, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001201.

Pan:2003:MSL

[Pan03] V. Y. Pan. Matrix structure and loss-resilient encoding/decoding. Computers and Mathematics with Applications, 46(2–3):493–499, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900411.

Pan:2005:ADP

[Pan05a] V. Y. Pan. The amended DSeSC power method for polynomial root-finding. Computers and Mathematics with Applications, 49(9– 10):1515–1524, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105001653.

Pan:2005:CFA

[Pan05b] V. Y. Pan. Coefficientfree adaptations of polynomial root-finders. Computers and Mathematics with Applications, 50(1–2):263– 269, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002488.

Pang:2005:SCF

[Pan05c]  Chin-Tzong Pang.  Si-

multaneously controllable fuzzy matrices. Computers and Mathematics with Applications, 50(10– 12):1647–1658, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004153.

Pan:2007:NHF

[Pan07a] Victor Y. Pan. New homotopic/factorization and symmetrization techniques for Newton’s and Newton/ structured iteration. Computers and Mathematics with Applications, 54(5): 721–729, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107002507.

Pang:2007:APP

[Pan07b] Chin-Tzong Pang. On the asymptotic period of powers of a fuzzy matrix. Computers and Mathematics with Applications, 54(2):310– 318, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002179.

Panyanak:2007:MII

[Pan07c] Bancha Panyanak. Mann and Ishikawa iterative processes for multivalued mappings in Banach spaces. Computers and Mathematics with Applications, 54(6): 872–877, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107003331. See

erratum [SW08].

Pao:2001:FDS

[Pao01a] C. V. Pao. Finite difference solutions of reaction diffusion equations with continuous time delays. Computers and Mathematics with Applications, 42(3–5):399–412, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001651.

Pao:2001:NMN

[Pao01b]  C. V. Pao.  Numeri-

cal methods for nonlinear integro-parabolic equations of Fredholm type. Computers and Mathematics with Applications, 41(7–8):857– 877, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003254.

Pao:2003:AMI

[Pao03]  C. V. Pao.  Acceler-

ated monotone iterations for numerical solutions of nonlinear elliptic boundary value problems. Computers and Mathematics with Applications, 46(10– 11):1535–1544, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901891.

Paraskevopoulos:2004:RAS

[Par04a] A. G. Paraskevopoulos. A recursive approach to the solution of abstract linear equations and the tau method. Computers and Mathematics with Applications, 47(10–11):1753–1774, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104839890.

Park:2004:DDG

[Par04b]  Ju H. Park.  Delay-

dependent guaranteed cost stabilization criterion for neutral delay-differential systems: matrix inequality approach. Computers and Mathematics with Applications, 47(10–11):1507–1515, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104839683.

Pathak:2005:WGS

[Pat05] R. S. Pathak. Wavelets in a generalized Sobolev space. Computers and Mathematics with Applications, 49(5–6):823–839, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000945.

Pal:2001:LAF

[PB01] A. Pal and R. K. Bera.

Large amplitude free vibration of sandwich parabolic plates — revisited. Computers and Mathematics with Applications, 41(3–4):513– 522, February 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100002935.

Power:2002:CAB

[PB02] H. Power and V. Barraco. A comparison analysis between unsymmetric and symmetric radial basis function collocation methods for the numerical solution of partial differential equations. Computers and Mathematics with Applications, 43(3–5):551–583, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003054.

Pattaraintakorn:2008:FRS

[PC08] Puntip Pattaraintakorn and Nick Cercone. A foundation of rough sets theoretical and computational hybrid intelligent system for survival analysis. Computers and Mathematics with Applications, 56(7):1699–1708, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002472.

Pan:2009:P

[PC09] Linqiang Pan and Guangzhao Cui. Preface. Computers and Mathematics with Applications, 57(11–12): 1723, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005543.

Pluempitiwiriyawej:2009:LAC

[PCA09] Charnyote Pluempitiwiriyawej, Nick Cercone, and Xiangdong An. Lexical acquisition and clustering of word senses to conceptual lexicon construction. Computers and Mathematics with Applications, 57(9):1537– 1546, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000212.

Palhares:2005:DDR

[PCE^{+}05] R. M. Palhares, C. D. Campos, P. Ya. Ekel, M. C. R. Leles, and M. F. S. V. D’Angelo. Delay-dependent robust H_{∞} control of uncertain linear systems with time-varying delays. Computers and Mathematics with Applications, 50(1–2): 13–32, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002282.

Pavlovic:2009:GJI

[PCRR09] Mirjana Pavlović, Nenad Cakić, Miloje Rajović, and Stojan Radenović. A generalization of Jensen’s inequality for polynomials having concentration at low degrees. Computers and Mathematics with Applications, 57(2):332–337, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005932.

Platte:2004:CEE

[PD04] R. B. Platte and T. A.

Driscoll. Computing eigenmodes of elliptic operators using radial basis functions. Computers and Mathematics with Applications, 48(3–4):561–576, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210484077X.

Platte:2006:ESR

[PD06] R. B. Platte and T. A. Driscoll. Eigenvalue stability of radial basis function discretizations for timedependent problems. Computers and Mathematics with Applications, 51(8): 1251–1268, April 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122106000770.

Prost-Domasky:2003:AVF

[PDBH03] S. A. Prost-Domasky, C. L. Brooks, and K. T. Honeycutt. The application of pversion finite-element methods to fracture-dominated problems encountered in engineering practice. Computers and Mathematics with Applications, 46(1):125– 139, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210390085X.

Pardo:2007:OPD

[PdlF07] María José Pardo and David de la Fuente. Optimizing a priority-discipline queueing model using fuzzy set theory. Computers and Mathematics with Applications, 54(2):267–281, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700212X.

Pardo:2009:NTO

[PdlF09] María José Pardo and David de la Fuente. A new technique to optimize the functions of fuzzy profit of queuing models: Application to a queuing model with publicity and renouncement. Computers and Mathematics with Applications, 57(5):850–864, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006652.

Paszynski:2005:VGO

[PDP05] M. Paszyński, L. Demkowicz, and D. Pardo. Verification of goal-oriented HP-adaptivity. Computers and Mathematics with Applications, 50(8– 9):1395–1404, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003895.

Ponsoda:2008:SNM

[PDRR08] Enrique Ponsoda, Emilio Defez, María Dolores Roselló, and José Vicente Romero. A stable numerical method for solving variable coefficient advection-diffusion models. Computers and Mathematics with Applications, 56(3):754–768, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800076X.

Pei:2002:NTG

[Pei02]  Daowu Pei.  Notes on

the theory of granular lattices. Computers and Mathematics with Applications, 43(8–9):975–980, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800074.

Peng:2002:OTS

[Pen02] Mingshu Peng. Oscillation theorems of secondorder nonlinear neutral delay difference equations with impulses. Computers and Mathematics with Applications, 44(5–6):741–748, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001876.

Peng:2003:NCP

[Pen03] Mingshu Peng. A note correcting the proof of a lemma in a recent paper. Computers and Mathematics with Applications, 45 (4–5):783–784, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000397.

Peng:2009:RAB

[Pen09] Lequn Peng. RETRACTED:

Asymptotic behavior of solutions to a differential equation with statedependent delay. Computers and Mathematics with Applications, 57(9):1511– 1514, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000364.

Petcu:2001:EOS

[Pet01]  D. Petcu.  Experiments

with an ODE solver on a multiprocessor system. Computers and Mathematics with Applications, 42 (8–9):1189–1199, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002322.

Peterson:2004:UCT

[Pet04] J. Peterson. Using the Cayley transform and finite elements to solve linear time dependent problems. Computers and Mathematics with Applications, 48(7–8):1101–1109, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003487.

Petcu:2005:PPI

[Pet05]  D. Petcu.  The perfor-

mance of parallel iterative solvers. Computers and Mathematics with Applications, 50(7):1179–1189, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003457.

Papageorgiou:2001:CRK

[PF01] G. Papageorgiou and I. Th. Famelis. Continuous Runge– Kutta–Nyström methods for initial value problems with periodic solutions. Computers and Mathematics with Applications, 42 (8–9):1165–1176, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002309.

Peng:2000:OCS

[PG00] Mingshu Peng and Weigao Ge. Oscillation criteria for second-order nonlinear differential equations with impulses. Computers and Mathematics with Applications, 39(5–6):217– 225, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000559.

Palamides:2004:PVP

[PG04] P. K. Palamides and G. N.

Galanis. Periodic value problems of differential systems on infinite-dimensional spaces and applications to differential geometry. Computers and Mathematics with Applications, 47(12): 1809–1815, June 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104840185.

Pineda:2006:GOG

[PG06] M. Delgado Pineda and E. A. Galperin. Global optimization over general compact sets by the beta algorithm: a MAPLE code. Computers and Mathematics with Applications, 52(1– 2):33–54, July 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0898122106001805.

Pineda:2008:MCG

[PG08] M. Delgado Pineda and E. A. Galperin. Maple code of gamma algorithm for global optimization of uncertain functions over compact robust sets. Computers and Mathematics with Applications, 56(1):90– 103, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107008292.

Parrilla:2008:HOO

[PGH^{+}08] Eduardo Parrilla, Damián Ginestar, José Luis Hueso, Jaime Riera, and Juan Ramón Torregrosa. Handling occlusion in optical flow algorithms for object tracking. Computers and Mathematics with Applications, 56(3):733–742, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000758.

Pang:2008:SNB

[PGT08] Huihui Pang, Weigao Ge, and Min Tian. Solvability of nonlocal boundary value problems for ordinary differential equation of higher order with a p-Laplacian. Computers and Mathematics with Applications, 56(1):127– 142, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008322.

Pirbodaghi:2009:DEC

[PHAF09] T. Pirbodaghi, S. H. Hoseini, M. T. Ahmadian, and G. H. Farrahi. Duffing equations with cubic and quintic nonlinearities. Computers and Mathematics with Applications, 57(3):500–506, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006433.

Phung:2002:NCH

[Phu02] K.-D. Phung. Null controllability of the heat equation as singular limit of the exact controllability of dissipative wave equation under the Bardos– Lebeau–Rauch geometric control condition. Computers and Mathematics with Applications, 44(10– 11):1289–1296, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002560.

Peng:2003:LOF

[PHW03] D. H. Peng, M. A. Han, and H. Y. Wang. Linearized oscillations of first-order nonlinear neutral delay difference equations. Computers and Mathematics with Applications, 45(12):1785– 1796, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900009.

Piejko:2009:GCC

[Pie09]  Krzysztof Piejko.  Gen-

eralized convolutions and classes of functions defined by subordination. Computers and Mathematics with Applications, 57(2):302–307, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800641X.

Pan:2008:APA

[PIM^{+}08] Victor Y. Pan, Dmitriy Ivolgin, Brian Murphy, Rhys Eric Rosholt, Islam Taj-Eddin, Yuqing Tang, and Xiaodong Yan. Additive preconditioning and aggregation in matrix computations. Computers and Mathematics with Applications, 55(8):1870–1886, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006608.

Piotrowski:2000:BMU

[Pio00] J. A. Piotrowski. Building a model of a useful Turing machine. Computers and Mathematics with Applications, 39(1–2):127– 143, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003193.

Petkovic:2004:CFP

[PIR04] M. S. Petković, S. Ilić, and L. Rancić. The convergence of a family of parallel zerofinding methods. Computers and Mathematics with Applications, 48(3–4):455– 467, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840689.

Pituk:2001:LSN

[Pit01]  M. Pituk.  The lim-

its of the solutions of a nonautonomous linear delay difference equation. Computers and Mathematics with Applications, 42(3–5):543–550, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001766.

Pituk:2003:GAS

[Pit03] M. Pituk. Global asymptotic stability in a perturbed higher-order linear difference equation. Computers and Mathematics with Applications, 45 (6–9):1195–1202, March/ May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000841.

Ponsoda:2004:NSP

[PJJP04] E. Ponsoda, L. Jódar, S. Jerez, and A. E. Posso. Numerical solution with a priori error bounds of coupled time dependent hyperbolic systems. Computers and Mathematics with Applications, 47 (2–3):233–246, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210490020X.

Pan:2006:LTD

[PKM^{+}06] V. Y. Pan, M. Kunin, B. Murphy, R. E. Rosholt, Y. Tang, X. Yan, and W. Cao. Linking the TPR1, DPR1 and arrow-head matrix structures. Computers and Mathematics with Applications, 52(10– 11):1603–1608, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003439.

Panda:2008:MIE

[PKM08] Debdulal Panda, Samarjit Kar, and Manoranjan Maiti. Multi-item EOQ model with hybrid cost parameters under fuzzy/fuzzy-stochastic resource constraints: a geometric programming approach. Computers and Mathematics with Applications, 56(11):2970–2985, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004884.

Park:2009:ASK

[PKP09] Jong Yeoul Park, Yong Han Kang, and Sun Hye Park. Adaptive stabilization of a von Kármán plate equation with a boundary output feedback control. Computers and Mathematics with Applications, 58(9):1742– 1754, November 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0898122109005847.

Petrova:2008:SFV

[PKSP08] Silvia Petrova, Hristina Kirova, Dimiter Syrakov, and Maria Prodanova. Some fast variants of TRAP scheme for solving advection equation

— comparison with other schemes. Computers and Mathematics with Applications, 55(10):2363–2380, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007304.

Pathak:2007:CFP

[PKT07] H. K. Pathak, M. S. Khan, and Rakesh Tiwari. A common fixed point theorem and its application to nonlinear integral equations. Computers and Mathematics with Applications, 53(6):961–971, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000508.

Pai:2001:AFS

[PL01a] P.-F. F. Pai and E. S. Lee. Adaptive fuzzy systems in group technology. Computers and Mathematics with Applications, 42(10– 11):1393–1400, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002486.

Pai:2001:PCS

[PL01b] Ping-Feng Pai and E. S.

Lee. Parts clustering by self-organizing map neural network in a fuzzy environment. Computers and Mathematics with Applications, 42(1–2):179– 188, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001420.

Pedersen:2001:CDS

[PL01c] M. Pedersen and Zhigui Lin. Coupled diffusion systems with localized nonlinear reactions. Computers and Mathematics with Applications, 42(6– 7):807–816, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002000.

Pelaez:2003:NMC

[PL03] J. I. Peláez and M. T.

Lamata. A new measure of consistency for positive reciprocal matrices. Computers and Mathematics with Applications, 46(12):1839–1845, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902409.

Pang:2005:PRL

[PL05a] Tian-Xiao Pang and Zheng-Yan Lin. Precise rates in the law of logarithm for i.i.d. random variables. Computers and Mathematics with Applications, 49(7–8):997–1010, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001161.

Peng:2005:SAP

[PL05b] Jian Wen Peng and Xian Jun Long. Sensitivity analysis for parametric completely generalized strongly nonlinear implicit quasivariational inclusions. Computers and Mathematics with Applications, 50(5–6): 869–880, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105003226.

Platis:2006:GFP

[Pla06] A. Platis. A generalized formulation for the performability indicator. Computers and Mathematics with Applications, 51(2):239–246, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/089812210500461X.

Plaza:2007:ETL

[Pla07] Angel Plaza. The eighttetrahedra longest-edge partition and Kuhn triangulations. Computers and Mathematics with Applications, 54(3):427–433, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002258.

Papalexandris:2002:UAM

[PLD02] M. V. Papalexandris, A. Leonard, and P. E. Dimotakis. Unsplit algorithms for multidimensional systems of hyperbolic conservation laws with source terms. Computers and Mathematics with Applications, 44(1–2):25– 49, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001281.

Pang:2007:MPS

[PLG07] Huihui Pang, Hairong Lian, and Weigao Ge. Multiple positive solutions for second-order four-point boundary value problem. Computers and Mathematics with Applications, 54(9– 10):1267–1275, November 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107003355.

Pei:2009:DSE

[PLG^{+}09] Yongzhen Pei, Shaoying Liu, Shujing Gao, Shuping Li, and Changguo Li. A delayed SEIQR epidemic model with pulse vaccination and the quarantine measure. Computers and Mathematics with Applications, 58(1):135– 145, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002296.

Patra:2003:DSL

[PLL03] A. K. Patra, A. Laszloffy, and J. Long. Data structures and load balancing for parallel adaptive hp finiteelement methods. Computers and Mathematics with Applications, 46(1):105– 123, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900848.

Park:2009:CIS

[PLP09] Jong Hyun Park, Guee Sang Lee, and Soon Young Park. Color image segmentation using adaptive mean shift and statistical modelbased methods. Computers and Mathematics with Applications, 57(6):970–980, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005798.

Piziak:2001:HFS

[PM01] R. Piziak and J. J. Mitchell.

Hamiltonian formalism and the state of a physical system. Computers and Mathematics with Applications, 42 (6–7):793–805, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001997.

Petrila:2005:NNA

[PM05a] T. Petrila and S. Maksay. A new numerical approach of a sliding boundary layer. Computers and Mathematics with Applications, 50(1–2):113– 121, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500235X.

Poranen:2005:RTO

[PM05b] T. Poranen and E. Mäkinen. Remarks on the thickness and outerthickness of a graph. Computers and Mathematics with Applications, 50(1–2):249– 254, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002464.

Pratico:2008:FWR

[PM08] Filippo G. Praticò and Antonino Moro. Flow of water in rigid solids: Development and experimental validation of models for tests on asphalts. Computers and Mathematics with Applications, 55(2):235–244, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003501.

Pan:2009:NEF

[PMQR09] V. Y. Pan, B. Murphy, G. Qian, and R. E. Rosholt. A new errorfree floating-point summation algorithm. Computers and Mathematics with Applications, 57(4):560–564, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006718.

Pan:2008:ESR

[PMR^{+}08] V. Y. Pan, B. Murphy, R. E. Rosholt, Y. Tang, X. Wang, and A. Zheng. Eigen-solving via reduction to DPR1 matrices. Computers and Mathematics with Applications, 56(1):166– 171, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000023.

Patel:2007:CMA

[PMS07] J. Patel, A. K. Mishra, and H. M. Srivastava. Classes of multivalent analytic functions involving the Dziok– Srivastava operator. Computers and Mathematics with Applications, 54(5): 599–616, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107001629.

Piziak:2004:AP

[PO04] R. Piziak and P. L. Odell.

Affine projections. Computers and Mathematics with Applications, 48(1–2):177– 190, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840082.

Polat:2009:GAM

[Pol09] Mustafa Polat. Global attractor for a modified Swift– Hohenberg equation. Computers and Mathematics with Applications, 57(1):62– 66, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005890.

Posner:2000:CN

[Pos00] E. C. Posner. A code in the nose. Computers and Mathematics with Applications, 39(11):127, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001188.

Pasko:2003:MSP

[POS03] A. Pasko, O. Okunev, and V. Savchenko. Minkowski sums of point sets defined by inequalities. Computers and Mathematics with Applications, 45(10–11):1479–1487, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001317.

Postolica:2008:CBE

[Pos08] Vasile Postolica. Choquet boundaries and efficiency. Computers and Mathematics with Applications, 55 (3):381–391, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107003690.

Pan:2001:PAG

[PP01a] Ping-Qi Pan and Yunpeng Pan. A phase-1 approach for the generalized simplex algorithm. Computers and Mathematics with Applications, 42(10– 11):1455–1464, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002541.

Perez-Palomares:2001:PPB

[PP01b] A. Pérez-Palomares. Preservation properties by Bernsteintype operators. a probabilistic approach. Computers and Mathematics with Applications, 42(1–2):141– 155, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001390.

Philos:2004:BSL

[PP04] Ch. G. Philos and I. K. Purnaras. The behavior of solutions of linear Volterra difference equations with infinite delay. Computers and Mathematics with Applications, 47(10–11):1555–1563, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839725.

Perri:2006:IIC

[PP06] F. Perri and L. Pandolfi. Input identification to a class of nonlinear input-output causal systems. Computers and Mathematics with Applications, 51(12):1773– 1788, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001295.

Paparella:2008:AIC

[PP08] Francesco Paparella and Giuseppe Passoni. Absence of inelastic collapse for a 1D gas of grains with an internal degree of freedom. Computers and Mathematics with Applications, 55(2):218–229, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003483.

Panchuk:2009:SNA

[PP09] A. Panchuk and T. Puu.

Stability in a non-autonomous iterative system: an application to oligopoly. Computers and Mathematics with Applications, 58(10):2022– 2034, November 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109004891.

Pathak:2008:MGS

[PPA08] R. S. Pathak, Gireesh Pandey, and Ryuichi Ashino. Multiwavelets in the generalized Sobolev space H_{W}^{ω} (R^{n}). Computers and Mathematics with Applications, 55 (3):423–440, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107003744.

Petkovic:2003:GCL

[PPI03] M. Petković, L. Petković, and S. Ilić. The guaranteed convergence of Laguerrelike method. Computers and Mathematics with Applications, 46(2–3):239–251, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900289.

Power:2006:APE

[PPP^{+}06] P. W. Power, C. C. Pain, M. D. Piggott, F. Fang, G. J. Gorman, A. P. Umpleby, A. J. H. Goddard, and I. M. Navon. Adjoint a posteriori error measures for anisotropic mesh optimisation. Computers and Mathematics with Applications, 52 (8–9):1213–1242, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003154.

Palomares:2002:SBB

[PPRG02] A. Palomares, M. Pasadas, V. Ramírez, and M. Ruiz Galán. Schauder bases in Banach spaces: application to numerical solutions of differential equations. Computers and Mathematics with Applications, 44(5–6):619–622, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001761.

Przeworska-Rolewicz:2001:PLT

[PR01] D. Przeworska-Rolewicz. Postmodern logarithmotechnia. Computers and Mathematics with Applications, 41(9):1143–1154, May 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101000876.

Petkovic:2006:FRF

[PR06] M. S. Petković and L. Z.

Rancić. A family of root-finding methods with accelerated convergence. Computers and Mathematics with Applications, 51 (6–7):999–1010, March/ April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000320.

Posta:2007:OCN

[PR07] Miroslav Posta and Tomás Roubícek. Optimal control of Navier–Stokes equations by Oseen approximation. Computers and Mathematics with Applications, 53(3–4):569–581, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001654.

Peterson:2001:LBM

[PRA01] G. Peterson, J. Reif, and S. Azhar. Lower bounds for multiplayer noncooperative games of incomplete information. Computers and Mathematics with Applications, 41(7–8):957– 992, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003333.

Peterson:2002:DAM

[PRA02] G. Peterson, J. Reif, and S. Azhar. Decision algorithms for multiplayer noncooperative games of incomplete information. Computers and Mathematics with Applications, 43(1–2):179– 206, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002826.

Prazak:2007:RCE

[Pra07] Dalibor Prazák. A remark on characterization of entropy solutions. Computers and Mathematics with Applications, 53(3–4):453–460, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700185X.

Prusa:2009:IBC

[Pru09] Vít Prusa. On the influence of boundary condition on stability of Hagen– Poiseuille flow. Computers and Mathematics with Applications, 57(5):763–771, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800624X.

Petropoulou:2001:BSA

[PS01] E. N. Petropoulou and P. D.

Siafarikas. Bounded solutions and asymptotic stability of nonlinear difference equations in the complex plane II. Computers and Mathematics with Applications, 42(3–5):427–452, August/September 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101001687.

Patel:2003:PCM

[PS03a] J. Patel and P. Sahoo. Properties of a class of multivalent analytic functions. Computers and Mathematics with Applications, 46 (10–11):1633–1644, November/December 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103901982.

Petropoulou:2003:SND

[PS03b] E. N. Petropoulou and P. D. Siafarikas. Solutions of nonlinear delay and advanced partial difference equations in the space l_{N×N}^{1}. Computers and Mathematics with Applications, 45(6–9):905–934, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000658.

Paternoster:2004:AGM

[PS04] B. Paternoster and L. Shaikhet. Application of the general method of Lyapunov functionals construction for difference Volterra equations. Computers and Mathematics with Applications, 47(8–9):1165–1176, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901113.

Peer:2007:FSP

[PS07a] S. K. Peer and Dinesh K. Sharma. Finding the shortest path in stochastic networks. Computers and Mathematics with Applications, 53(5):729–740, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000260.

Peretz:2007:BTM

[PS07b] Ronen Peretz and Reuven Segev. Bounds on the trace mapping of LDfields. Computers and Mathematics with Applications, 53(3–4):665–684, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001162.

Parthasarathy:2008:TSM

[PS08a] P. R. Parthasarathy and R. Sudhesh. Transient solution of a multiserver Poisson queue with N-policy. Computers and Mathematics with Applications, 55 (3):550–562, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107004907.

Peer:2008:HCI

[PS08b] S. K. Peer and Dinesh K.

Sharma. Human-computer interaction design with multi-goal facilities layout model. Computers and Mathematics with Applications, 56(9):2164–2174, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003519.

Pogany:2009:SMT

[PS09] T. K. Pogány and H. M. Srivastava. Some Mathieutype series associated with the Fox–Wright function. Computers and Mathematics with Applications, 57 (1):127–140, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108004288.

Pickmann:2007:IEP

[PSES07] Hubert Pickmann, Ricardo L. Soto, J. Egaña, and Mario Salas. An inverse eigenvalue problem for symmetrical tridiagonal matrices. Computers and Mathematics with Applications, 54(5):699–708, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002489.

Parthasarathy:2002:ETA

[PSM02] P. R. Parthasarathy, N. Selvaraju, and G. Manimaran. Exact transient analysis of a circulant queuing network. Computers and Mathematics with Applications, 44(10– 11):1297–1306, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002572.

Pang:2007:MPC

[PSS07] Li-Ping Pang, Jie Shen, and He-Shan Song. A modified predictor–corrector algorithm for solving nonconvex generalized variational inequality. Computers and Mathematics with Applications, 54(3):319–325, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001332.

Pandey:2009:EAS

[PSS09] Rajesh K. Pandey, Om P.

Singh, and Vineet K. Singh. Efficient algorithms to solve singular integral equations of Abel type. Computers and Mathematics with Applications, 57(4):664–676, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006172.

Petkovic:2008:EPM

[PST08] Marko D. Petković, Predrag S. Stanimirović, and Milan B. Tasić. Effective partitioning method for computing weighted Moore– Penrose inverse. Computers and Mathematics with Applications, 55(8):1720– 1734, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006475.

Peile:2000:SPP

[PT00] R. E. Peile and H. Taylor. Sets of points with pairwise distinct slopes. Computers and Mathematics with Applications, 39(11):109– 115, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001152.

Pisarn:2007:HBM

[PT07] C. Pisarn and T. Theeramunkong. An HMM-based method for Thai spelling speech recognition. Computers and Mathematics with Applications, 54(1):76– 95, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001812.

Petkovic:2000:CEL

[PV00] M. S. Petković and D. V. Vranić. The convergence of Euler-like method for the simultaneous inclusion of polynomial zeros. Computers and Mathematics with Applications, 39(7–8):95– 105, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000687.

Pereira:2001:DBE

[PV01] E. Pereira and J. Vitória.

Deflation for block eigenvalues of block partitioned matrices with an application to matrix polynomials of commuting matrices. Computers and Mathematics with Applications, 42 (8–9):1177–1188, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002310.

Platkowski:2000:APD

[PW00] T. Platkowski and W. Waluś.

An acceleration procedure for discrete velocity approximation of the Boltzmann collision operator. Computers and Mathematics with Applications, 39(5–6):151– 163, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000511.

Peng:2005:LSS

[PW05] Yahong Peng and Ya-Guang Wang. Large solution of a semilinear elliptic problem. Computers and Mathematics with Applications, 49(9–10):1387– 1395, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001574.

Plubtieng:2006:SCT

[PW06] S. Plubtieng and R. Wangkeeree. Strong convergence theorems for threestep iterations with errors for non-lipschitzian nonself-mappings in Banach spaces. Computers and Mathematics with Applications, 51(6–7):1093–1102, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000393.

Pang:2008:FOO

[PWX08] Li-Ping Pang, Ming-Zheng Wang, and Zun-Quan Xia. First-order optimality conditions for two classes of generalized nonsmooth semi-infinite optimization. Computers and Mathematics with Applications, 56(5): 1457–1464, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108001715.

Peng:2009:NCO

[PXGL09] Y. Peng, X. Xiang, Y. Gong, and G. Liu. Necessary conditions of optimality for a class of optimal control problems on time scales. Computers and Mathematics with Applications, 58(10):2035–2045, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005719.

Peng:2009:SON

[PXW09] Y. Peng, X. Xiang, and W. Wei. Second-order nonlinear impulsive integrodifferential equations of mixed type with timevarying generating operators and optimal controls on Banach spaces. Computers and Mathematics with Applications, 57(1):42–53, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005919.

Peng:2006:OFC

[PXXW06] Y. Peng, X-Xiang, and W. Wei. Optimal feedback control for a class of strongly nonlinear impulsive evolution equations. Computers and Mathematics with Applications, 52(5):759–768, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002549.

Peng:2009:TEM

[PY09] Jian-Wen Peng and Jen-Chih Yao. Two extragradient methods for generalized mixed equilibrium problems, nonexpansive mappings and monotone mappings. Computers and Mathematics with Applications, 58(7):1287–1301, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004805.

Peng:2008:SML

[PYLX08] Dezhong Peng, Zhang Yi, Jian Cheng Lv, and Yong Xiang. A stable MCA learning algorithm. Computers and Mathematics with Applications, 56(4):847–860, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000862.

Peng:2007:ESC

[PZ07] Jian Wen Peng and Dao Li Zhu. Existence of solutions and convergence of iterative algorithms for a system of generalized nonlinear mixed quasi-variational inclusions. Computers and Mathematics with Applications, 53(5):693–705, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700017X.

Qian:2006:CAP

[QC06] Tao Qian and Qiuhui Chen.

Characterization of analytic phase signals. Computers and Mathematics with Applications, 51(9–10):1471– 1482, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000940.

Qiu:2003:WRS

[QFZ03] Chun-Yu Qiu, Chu-Li Fu, and You-Bin Zhu. Wavelets and regularization of the sideways heat equation. Computers and Mathematics with Applications, 46 (5–6):821–829, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103901453.

Qian:2008:MSC

[Qia08] Tiezheng Qian. Monolayer spreading in confined immiscible fluids: a molecular dynamics simulation. Computers and Mathematics with Applications, 55(7):1554– 1559, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006438.

Qiu:2008:DBV

[Qiu08] Zhipeng Qiu. Dynamical behavior of a vector-host epidemic model with demographic structure. Computers and Mathematics with Applications, 56(12):3118– 3129, December 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004732.

Qu:2009:CTR

[QJZ09] Shao-Jian Qu, Su-Da Jiang, and Ying Zhu. A conic trust-region method and its convergence properties. Computers and Mathematics with Applications, 57 (4):513–528, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108006731.

Qian:2008:CAR

[QLD08a] Yuhua Qian, Jiye Liang, and Chuangyin Dang. Converse approximation and rule extraction from decision tables in rough set theory. Computers and Mathematics with Applications, 55(8):1754–1765, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006499.

Qian:2008:IOI

[QLD08b] Yuhua Qian, Jiye Liang, and Chuangyin Dang. Interval ordered information systems. Computers and Mathematics with Applications, 56(8):1994–2009, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002770.

Qian:2003:RMR

[QNN03] C. Qian, S. Nakamura, and T. Nakagawa. Replacement and minimal repair policies for a cumulative damage model with maintenance. Computers and Mathematics with Applications, 46(7):1111–1118, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390126X.

Qian:2003:PFV

[QS03] J. Qian and W. W. Symes. A paraxial formulation for the viscosity solution of quasi-P eikonal equations. Computers and Mathematics with Applications, 46 (10–11):1691–1701, November/December 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103902033.

Qin:2008:NOC

[QSL08] Huizeng Qin, Nina Shang, and Youmin Lu. A note on oscillation criteria of second order nonlinear neutral delay differential equations. Computers and Mathematics with Applications, 56(12):2987–2992, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800477X.

Quapp:2001:SMD

[Qua01] W. Quapp. Searching minima of an N-dimensional surface: a robust valley following method. Computers and Mathematics with Applications, 41(3–4):407–414, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002832.

Qu:2009:PRA

[QW09] Bo Qu and Kerui Weng.

Path relinking approach for multiple allocation hub maximal covering problem. Computers and Mathematics with Applications, 57(11–12):1890– 1894, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005348.

Wu:2007:PGP

[qWjH07] Ke qing Wu and Nan jing Huang. Properties of the generalized f-projection operator and its applications in Banach spaces. Computers and Mathematics with Applications, 54(3):399–406, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002222.

Qi:2005:LPC

[QZ05] Jiangang Qi and Songping Zhou. Limit-point criterion for singular linear Dirac differential systems. Computers and Mathematics with Applications, 49(5–6):765–775, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000908.

Radenovic:2009:CFP

[Rad09] Stojan Radenović. Common fixed points under contractive conditions in cone metric spaces. Computers and Mathematics with Applications, 58(6):1273–1278, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004726.

Ramachandran:2005:BIM

[Ram05] P. A. Ramachandran. Boundary integral method for multispecies 1-D convection– diffusion-reaction problems. Computers and Mathematics with Applications, 49(9– 10):1575–1583, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122105001707.

Rappoport:2001:CVP

[Rap01] J. M. Rappoport. Canonical vector polynomials for the computation of complex order Bessel functions with the tau method. Computers and Mathematics with Applications, 41(3–4):399–406, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002820.

Rashid:2006:CAT

[Ras06]  A. Rashid.  Conver-

gence analysis of threelevel Fourier pseudospectral method for Korteweg– de Vries Burgers equation. Computers and Mathematics with Applications, 52(5): 769–778, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106002550.

Razgulin:2000:FDD

[Raz00] A. V. Razgulin. Finitedimensional dynamics of distributed optical systems with delayed feedback. Computers and Mathematics with Applications, 40(12):1405–1418, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002492.

Roychoudhuri:2006:TEI

[RB06] S. K. Roychoudhuri and N. Bandyopadhyay. Thermoelastic interactions in an infinite elastic solid due to distributed time-dependent heat sources in generalized thermo-elasticity III. Computers and Mathematics with Applications, 51(9– 10):1611–1624, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106001052.

Roychoudhuri:2007:IDB

[RB07] S. K. Roychoudhuri and Nupur Bandyopadhyay. Interactions due to body forces in generalized thermoelasticity III. Computers and Mathematics with Applications, 54(11–12):1341– 1352, December 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107002143.

Ren:2007:PSC

[RC07] Fengli Ren and Jinde Cao. Periodic solutions for a class of higher-order Cohen– Grossberg type neural networks with delays. Computers and Mathematics with Applications, 54(6):826–839, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002830.

Ren:2009:HOD

[RC09] Jingli Ren and Zhibo Cheng.

On high-order delay differential equation. Computers and Mathematics with Applications, 57(2):324–331, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006251.

Rafei:2007:AVI

[RD07] M. Rafei and H. Daniali.

Application of the variational iteration method to the Whitham–Broer–Kaup equations. Computers and Mathematics with Applications, 54(7–8):1079–1085, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002817.

Richardson:2000:EVS

[RDS00] A. D. Richardson, J. R. Dormand, and M. H. B. M. Shariff. Efficient variable stiffness methods for cooling of hot-rolled steel sections. Computers and Mathematics with Applications, 40(4–5):453–470, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001735.

Reed:2000:BHD

[Ree00] I. S. Reed. A brief history of the development of error correcting codes. Computers and Mathematics with Applications, 39(11): 89–93, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001127.

Ren:2003:SNE

[Ren03] Yan-Xia Ren. The structure of nonlinear elliptic equations on unbounded domains in dimensions 1 and 2- a probabilistic approach. Computers and Mathematics with Applications, 45 (4–5):771–782, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000385.

Rahman:2009:SSN

[RF09] Mohammad M. Rahman and David R. Fearn. A spectral solution of nonlinear mean field dynamo equations: With inertia. Computers and Mathematics with Applications, 58(3):422–435, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003204.

Roman-Flores:2003:CFS

[RFCCRM03] H. Román-Flores, Y. Chalco-Cano, and M. Rojas-Medar. Convolution of fuzzy sets and applications. Computers and Mathematics with Applications, 46 (8–9):1245–1251, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390215X.

Ramani:2001:SFS

[RGTT01] A. Ramani, B. Grammaticos, T. Tamizhmani, and K. M. Tamizhmani. Special function solutions of the discrete Painlevé equations. Computers and Mathematics with Applications, 42(3–5):603–614, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001808.

Ramani:2003:RDA

[RGTT03] A. Ramani, B. Grammaticos, T. Tamizhmani, and K. M. Tamizhmani. The road to the discrete analogue of the Painlevé property: Nevanlinna meets singularity confinement. Computers and Mathematics with Applications, 45 (6–9):1001–1012, March/ May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000762.

Ran:2009:IAB

[RH09] Rui-Sheng Ran and Ting-Zhu Huang. An inversion algorithm for a banded matrix. Computers and Mathematics with Applications, 58(9):1699–1710, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109005902.

Reif:2008:FDP

[RKS08] John H. Reif, Simon Kasif, and Deepak Sherlekar. Formula dissection: a parallel algorithm for constraint satisfaction. Computers and Mathematics with Applications, 55(6):1170–1177, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005846.

Reichel:2008:CSN

[RL08] Bartosz Reichel and Sergey Leble. On convergence and stability of a numerical scheme of coupled nonlinear Schrödinger equations. Computers and Mathematics with Applications, 55(4):745–759, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005068.

Rajagopal:2007:MWSa

[RM07a] K. R. Rajagopal and J. Malek.

Mechanics: the well-spring of mathematics (volume 1). Computers and Mathematics with Applications, 53(2): 155, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000648.

Rajagopal:2007:MWSb

[RM07b] K. R. Rajagopal and J. Malek.

Mechanics: the well-spring of mathematics (volume 2). Computers and Mathematics with Applications, 53 (3–4):345, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107001964.

Ricca:2008:MFP

[RM08] Renzo L. Ricca and Francesca Maggioni. Multiple folding and packing in DNA modeling. Computers and Mathematics with Applications, 55(5):1044–1053, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005688.

Re:2008:GTA

[RMC08] Angela Re, Ivan Molineris, and Michele Caselle. Graph theory analysis of genomics problems: Community analysis of fragile sites correlations and of pseudogenes alignments. Computers and Mathematics with Applications, 55(5):1034– 1043, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005676.

Ryoo:2002:NVSa

[RN02] C. S. Ryoo and M. T. Nakao. Numerical verification of solutions for variational inequalities of the second kind. Computers and Mathematics with Applications, 43(10–11):1371–1380, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001050.

Razali:2000:NCM

[RNM00] M. R. M. Razali, M. Z.

Nashed, and A. H. M. Murid. Numerical conformal mapping via the Bergman kernel using the generalized minimum residual method. Computers and Mathematics with Applications, 40(1):157– 164, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001498.

Rodin:2000:BR

[Rod00] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 40(8–9):1109–1116, October/November 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210085021X.

Rodin:2001:BRa

[Rod01a] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 42(1–2):265–272, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210100150X.

Rodin:2001:BRb

[Rod01b] ErvinY. Rodin. Book reports. Computers and Mathematics with Applications, 42(10–11):1497–1504, November/December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122101002577.

Rodin:2002:BRa

[Rod02a] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 43(1–2):267–274, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210100284X.

Rodin:2002:BRb

[Rod02b] ErvinY. Rodin. Book reports. Computers and Mathematics with Applications, 43(10–11):1469–1476, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001128.

Rodin:2004:BR

[Rod04] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 48(5–6):971–994, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210400327X.

Rodin:2005:BR

[Rod05] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 49(4):633–640, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105000568.

Rodin:2006:BR

[Rod06] Ervin Y. Rodin. BOOK REPORTS. Computers and Mathematics with Applications, 52(10–11):1615–1622, November/December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106003452.

Roeger:2009:GNF

[Roe09] Lih-Ing W. Roeger. General nonstandard finitedifference schemes for differential equations with three fixed-points. Computers and Mathematics with Applications, 57(3):379–383, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006809.

Rojo:2008:NAT

[Roj08]  Oscar Rojo.  A new

algebra of Toeplitz-plus-Hankel matrices and applications. Computers and Mathematics with Applications, 55(12):2856–2869, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007833.

Romanenko:2001:DDE

[Rom01] E. Yu. Romanenko. Differentialdifference equations reducible to difference and qdifference equations. Computers and Mathematics with Applications, 42(3–5): 615–626, August/September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210100181X.

Roop:2008:NAO

[Roo08]  John Paul Roop.  Nu-

merical approximation of a one-dimensional space fractional advection-dispersion equation with boundary layer. Computers and Mathematics with Applications, 56(7):1808–1819, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002617.

Rossini:2003:DOH

[Ros03] M. Rossini. Detecting objects hidden in the subsoil by a mathematical method. Computers and Mathematics with Applications, 45 (1–3):299–307, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800202.

Rodriguez:2001:TVO

[ROY01] L. Rodríguez, H. Ogata, and Y. Yano. A temporal versioned object-oriented data schema model. Computers and Mathematics with Applications, 41(1–2):177– 192, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210185015X.

Romkes:2003:PEA

[RPO03] A. Romkes, S. Prudhomme, and J. T. Oden. A priori error analyses of a stabilized discontinuous Galerkin method. Computers and Mathematics with Applications, 46 (8–9):1289–1311, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902203.

Rong:2008:SCF

[RQH^{+}08] Xiaoying Rong, Dewei Qi, Guowei He, JunYong Zhu, and Tim Scott. Single curved fiber sedimentation under gravity. Computers and Mathematics with Applications, 55(7):1560– 1567, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006414.

Revelli:2003:SCI

[RR03a] R. Revelli and L. Ridolfi. Sinc collocationinterpolation method for the simulation of nonlinear waves. Computers and Mathematics with Applications, 46(8–9):1443–1453, October/November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S089812210390232X.

Rojo:2003:CSN

[RR03b] O. Rojo and H. Rojo. Constructing symmetric nonnegative matrices via the Fast Fourier Transform. Computers and Mathematics with Applications, 45 (10–11):1655–1672, May/ June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001457.

Revelli:2008:GCM

[RR08] R. Revelli and L. Ridolfi. Generalized collocation method for twodimensional reaction-diffusion problems with homogeneous Neumann boundary conditions. Computers and Mathematics with Applications, 56(9):2360–2370, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003751.

Radenovic:2009:FPT

[RR09a] Stojan Radenović and B. E.

Rhoades. Fixed point theorem for two non-self mappings in cone metric spaces. Computers and Mathematics with Applications, 57(10):1701–1707, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002119.

Rafiq:2009:SMS

[RR09b] Arif Rafiq and Muhammad Rafiullah. Some multi-step iterative methods for solving nonlinear equations. Computers and Mathematics with Applications, 58(8):1589–1597, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004672.

Reyes:2008:ANS

[RRM08] Elia Reyes, Francisco Rodríguez, and José Antonio Martín. Analytic-numerical solutions of diffusion mathematical models with delays. Computers and Mathematics with Applications, 56(3):743–753, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000710.

Ramachandran:2007:ERW

[RRR07] Prabhu Ramachandran, M. Ramakrishna, and S. C. Rajan. Efficient random walks in the presence of complex two-dimensional geometries. Computers and Mathematics with Applications, 53(2):329–344, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001976.

Rinsaka:2006:SMA

[RS06] Koichiro Rinsaka and Hiroaki Sandoh. A stochastic model on an additional warranty service contract. Computers and Mathematics with Applications, 51(2):179–188, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004554.

Radha:2009:LRM

[RS09] R. Radha and S. Sivananthan. Local reconstruction method and voice system. Computers and Mathematics with Applications, 58(1): 74–79, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002144.

Rajovic:2009:TLN

[RSD09] Miloje Rajović, Rade Stojiljković, and Dragan Dimitrovski. Transformation of linear non-homogeneous differential equations of the second order to homogeneous. Computers and Mathematics with Applications, 57(4):604–611, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800686X.

Rojo:2001:NCP

[RSE01] O. Rojo, R. Soto, and J. Egaña. A note on the construction of a positive oscillatory matrix with a prescribed spectrum. Computers and Mathematics with Applications, 41(3–4):353–361, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002789.

Ryoo:2002:NVSb

[RSK02] Cheon Seoung Ryoo, Howan Song, and Sang Dong Kim. Numerical verification of solutions for some unilateral boundary value problems. Computers and Mathematics with Applications, 44(5–6):787–797, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001918.

Rojo:2000:BSE

[RSR00] O. Rojo, R. Soto, and H. Rojo. Bounds for sums of eigenvalues and applications. Computers and Mathematics with Applications, 39(7–8):1–15, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000602.

Rojo:2001:FCS

[RSR01] O. Rojo, R. Soto, and H. Rojo. Fast construction of a symmetric nonnegative matrix with a prescribed spectrum. Computers and Mathematics with Applications, 42(10– 11):1379–1391, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002474.

Rahnamayan:2007:NPI

[RTS07] Shahryar Rahnamayan, Hamid R.

Tizhoosh, and Magdy M. A. Salama. A novel population initialization method for accelerating evolutionary algorithms. Computers and Mathematics with Applications, 53(10):1605– 1614, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001344.

Rayes:2005:FPC

[RTW05] M. O. Rayes, V. Trevisan, and P. S. Wang. Factorization properties of Chebyshev polynomials.

Computers and Mathematics with Applications, 50 (8–9):1231–1240, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003767.

Rubin:2007:SIN

[Rub07] M. B. Rubin. A simplified implicit Newmark integration scheme for finite rotations. Computers and Mathematics with Applications, 53(2):219–231, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000880.

Raciti:2001:EDA

[RV01] F. Raciti and E. Venturino. Eigenfrequency determination for arbitrary cross-section waveguides. Computers and Mathematics with Applications, 41(3–4):319–326, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002765.

Ratnarajah:2005:CSW

[RV05] T. Ratnarajah and R. Vaillancourt. Complex singular Wishart matrices and applications. Computers and Mathematics with Applications, 50(3–4):399–411, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002841.

Rodriguez-Velazquez:2009:RIC

[RVKDF09] Juan Alberto Rodríguez-Velázquez, Aida Kamisalić, and Josep Domingo-Ferrer. On reliability indices of communication networks. Computers and Mathematics with Applications, 58(7): 1433–1440, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210900491X.

Riviere:2003:PEE

[RW03] B. Rivière and M. F.

Wheeler. A posteriori error estimates for a discontinuous Galerkin method applied to elliptic problems. log number: R74. Computers and Mathematics with Applications, 46(1):141– 163, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900861.

Rundell:2006:P

[RXY06] William Rundell, Xiaoling Xiang, and Hong-Ming Yin. Preface. Computers and Mathematics with Applications, 52(5):xiii–xiv, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002586.

Ryoo:2000:NVSb

[Ryo00a] C. S. Ryoo. Numerical verification of solutions for a simplified Signorini problem. Computers and Mathematics with Applications, 40 (8–9):1003–1013, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100850117.

Ryoo:2000:NVSa

[Ryo00b] Cheon Seoung Ryoo. Numerical verification of solutions for obstacle problems using a Newton-like method. Computers and Mathematics with Applications, 39(3–4):185–194, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003442.

Ryoo:2003:SOP

[Ryo03] Cheon Seoung Ryoo. Solving obstacle problems with guaranteed accuracy. Computers and Mathematics with Applications, 45(4–5): 823–834, February/March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103000439.

Ryoo:2004:NVS

[Ryo04] Cheon Seoung Ryoo. A numerical verification of solutionsof free boundary problems. Computers and Mathematics with Applications, 48(3–4):429–435, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840665.

Ryoo:2007:ANV

[Ryo07] Cheon Seoung Ryoo. An approach to the numerical verification of solutions for obstacle problems. Computers and Mathematics with Applications, 53(5):842–850, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000387.

Rubinstein:2003:SDE

[RZ03] R. Rubinstein and Ye Zhou.

Schiestel’s derivation of the epsilon equation and twoequation modelling of rotating turbulence. Computers and Mathematics with Applications, 46(4):633–638, August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900216.

Rivera:2003:EPA

[RZH03] W. Rivera, Jianping Zhu, and D. Huddleston. An efficient parallel algorithm with application to computational fluid dynamics. Computers and Mathematics with Applications, 45(1–3):165–188, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800135.

Saker:2002:OGA

[SA02] S. H. Saker and S. Agarwal. Oscillation and global attractivity in a nonlinear delay periodic model of respiratory dynamics. Computers and Mathematics with Applications, 44(5–6):623–632, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001773.

Sirakaya:2003:GNN

[SA03] S. Sirakaya and N. M.

Alemdar. Genetic neural networks to approximate feedback Nash equilibria in dynamic games. Computers and Mathematics with Applications, 46(10– 11):1493–1509, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901866.

Sen:2005:MVT

[SA05] S. K. Sen and H. Agarwal. Mean value theorem for boundary value problems with given upper and lower solutions. Computers and Mathematics with Applications, 49(9–10):1499– 1514, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001641.

Sen:2008:GRS

[SA08] S. K. Sen and Ravi P.

Agarwal. Golden ratio in science, as random sequence source, its computation and beyond. Computers and Mathematics with Applications, 56(2):469– 498, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800031X.

Sweilam:2007:VIM

[SAB07] N. H. Sweilam and R. F. Al-Bar. Variational iteration method for coupled nonlinear Schrödinger equations. Computers and Mathematics with Applications, 54(7–8):993–999, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700260X.

Sahadevan:2001:NDD

[Sah01] R. Sahadevan. Nonlinear differential-difference and difference equations: integrability and exact solvability. Computers and Mathematics with Applications, 42(3–5):627–637, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001821.

Salkuyeh:2008:CVI

[Sal08] Davod Khojasteh Salkuyeh. Convergence of the variational iteration method for solving linear systems of ODEs with constant coefficients. Computers and Mathematics with Applications, 56(8):2027–2033, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003039.

Sakurai:2002:PES

[SALK02] K. Sakurai, T. Aoki, W.-H.

Lee, and K. Kato. Poisson equation solver with fourthorder accuracy by using interpolated differential operator scheme. Computers and Mathematics with Applications, 43(6–7):621–630, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210100308X.

Samohyl:2007:ATM

[Sam07] Ivan Samohýl. Application of Truesdell’s model of mixtures to an ionic liquid mixture. Computers and Mathematics with Applications, 53(2):182–197, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000776.

Sajid:2008:ISC

[SANH08] M. Sajid, M. Awais, S. Nadeem, and T. Hayat. The influence of slip condition on thin film flow of a fourth grade fluid by the homotopy analysis method. Computers and Mathematics with Applications, 56(8):2019–2026, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002794.

Sarkar:2003:ESS

[Sar03] S. Sarkar. The effect of stable stratification on turbulence anisotropy in uniformly sheared flow. Computers and Mathematics with Applications, 46(4): 639–646, August 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103900228.

Sarra:2006:DTV

[Sar06a] S. A. Sarra. Digital total variation filtering as postprocessing for radial basis function approximation methods. Computers and Mathematics with Applications, 52(6– 7):1119–1130, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002835.

Sarra:2006:IMR

[Sar06b] S. A. Sarra. Integrated multiquadric radial basis function approximation methods. Computers and Mathematics with Applications, 51(8):1283– 1296, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000848.

Sen:2006:INB

[SAS06] S. K. Sen, H. Agarwal, and T. Samanta. Interpolation for nonlinear BVP in circular membrane with known upper and lower solutions. Computers and Mathematics with Applications, 51(6–7):1021–1046, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000344.

Sequeira:2009:LDM

[SASHS09] A. Sequeira, A. M. Artoli, A. S. Silva-Herdade, and C. Saldanha. Leukocytes dynamics in microcirculation under shear-thinning blood flow. Computers and Mathematics with Applications, 58(5):1035–1044, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001084.

Sattayatham:2006:RCC

[Sat06] Pairote Sattayatham. Relaxed control for a class of strongly nonlinear impulsive evolution equations. Computers and Mathematics with Applications, 52(5): 779–790, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122106002562.

Sauer:2007:EHW

[Sau07]  Niko Sauer.  Extended

Helmholtz–Weyl decompositions. Computers and Mathematics with Applications, 53(3–4):521–529, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001770.

Savas:2007:FSI

[Sav07a]  Ekrem Savaš.  Factors

for |A|_{k} summability of infinite series. Computers and Mathematics with Applications, 53(7):1045– 1049, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000612.

Savas:2007:SFT

[Sav07b] Ekrem Savaš. A summability factor theorem involving an almost increasing sequence for generalized absolute summability. Computers and Mathematics with Applications, 54(11–12):1380–1386, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003707.

Savas:2008:DSS

[Sav08a] Ekrem Savaş. (A_{σ})-double sequence spaces defined by Orlicz function and double statistical convergence. Computers and Mathematics with Applications, 55(6):1293–1301, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005986.

Savas:2008:ASF

[Sav08b]  Ekrem Savas.  On ab-

solute summability factors of infinite series. Computers and Mathematics with Applications, 56(1):25– 29, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008218.

Simos:2003:P

[SAVA03] T. E. Simos, George Avdelas, and Jesús Vigo-Aguiar. Preface. Computers and Mathematics with Applications, 45(1–3):xv, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103800019.

Sawaki:2003:OPC

[Saw03] K. Sawaki. Optimal policies in continuous time inventory control models with limited supply. Computers and Mathematics with Applications, 46(7):1139–1145, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901295.

Sundar:2000:GEP

[SB00] S. Sundar and B. K. Bhagavan. Generalized eigenvalue problems: Lanczos algorithm with a recursive partitioning method. Computers and Mathematics with Applications, 39(7–8):211– 224, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000778.

Sinha:2003:EAS

[SB03] M. Sinha and R. K. Bera.

Eigenvalue approach to study the effect of rotation and relaxation time in generalised thermoelasticity. Computers and Mathematics with Applications, 46(5–6):783–792, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901416.

Stoyan:2006:CVD

[SB06] G. Stoyan and Á. Baran.

Crouzeix–Velte decompositions for higher-order finite elements. Computers and Mathematics with Applications, 51(6–7):967–986, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000307.

Sveshnikov:2000:CDW

[SBD^{+}00] A. G. Sveshnikov, A. N.

Bogolyubov, A. L. Delitsyn, A. V. Krasilnikova, and D. V. Minaev. Calculation of dielectric waveguide systems using finite-difference method. Computers and Mathematics with Applications, 40(12):1387–1395, December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002479.

Szabo:2001:P

[SBGS01] Barna Szabó, Katalin Balla, Aurél Galántai, and György Szeidl. Preface. Computers and Mathematics with Applications, 42(8–9):xiii, October/November 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122101002176.

Szabo:2005:P

[SBGS05] Barna Szabó, Katalin Balla, Aurél Galántai, and György Szeidl. Preface. Computers and Mathematics with Applications, 50(7):xiii, October 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003482.

Silvani:2008:BLF

[SBPD08] C. Silvani, S. Bonelli, P. Philippe, and T. Desoyer. Buoyancy and local friction effects on rockfill settlements: a discrete modelling. Computers and Mathematics with Applications, 55(2):208–217, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003471.

Sakthivel:2006:COT

[SBS06] K. Sakthivel, K. Balachandran, and S. S. Sritharan. Controllability and observability theory of certain parabolic integrodifferential equations. Computers and Mathematics with Applications, 52(8– 9):1299–1316, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003191.

Statnikov:2006:MAT

[SBSY06] R. Statnikov, A. Bordetsky, A. Statnikov, and I. Yanushkevich. Multicriteria analysis tools in real-life problems. Computers and Mathematics with Applications, 52(1–2): 1–32, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001799.

Sheng:2000:AGM

[SC00a] Q. Sheng and H. Cheng. An adaptive grid method for degenerate semilinear quenching problems. Computers and Mathematics with Applications, 39(9–10):57– 71, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000869.

Song:2000:HDG

[SC00b] Xinyu Song and Lansun Chen. Harmless delays and global attractivity for nonautonomous predator–prey system with dispersion. Computers and Mathematics with Applications, 39(5–6):33–42, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000444.

Sun:2002:WDC

[SC02] Hung-Min Sun and Bor-Liang Chen. Weighted decomposition construction for perfect secret sharing schemes. Computers and Mathematics with Applications, 43(6–7):877–887, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101003285.

Shi:2003:PSE

[SC03] Guoliang Shi and Shaozhu Chen. Positive solutions of even higher-order singular superlinear boundary value problems. Computers and Mathematics with Applications, 45(4–5):593– 603, February/March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103000191.

Sakthivel:2004:SCS

[SC04] R. Sakthivel and Q. H. Choi.

A study on controllability of semilinear integrodiferential systems in Banach spaces. Computers and Mathematics with Applications, 47 (4–5):519–527, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900429.

See [HPT05].

Sheng:2007:FRR

[SC07] Xingping Sheng and Guoliang Chen. Full-rank representation of generalized inverse A_{T,S}^{(2)} and its application. Computers and Mathematics with Applications, 54(11–12):1422–1430, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107004841.

Siu:2009:HOM

[SCF^{+}09] T. K. Siu, W. K. Ching, E. Fung, M. Ng, and X. Li. A high-order Markov-switching model for risk measurement. Computers and Mathematics with Applications, 58(1):1– 10, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002363.

Schloss:2000:AHC

[Sch00] H. S. Schloss. Accuracy of the Hebrew calendar. Computers and Mathematics with Applications, 39(11): 23–24, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001024.

Schoberl:2001:ECS

[Sch01] J. Schöberl. Efficient contact solvers based on domain decomposition techniques. Computers and Mathematics with Applications, 42 (8–9):1217–1228, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002346.

Stiriba:2003:NSP

[SD03] Y. Stiriba and R. Donat. A numerical study of postshock oscillations in slowly moving shock waves. Computers and Mathematics with Applications, 46(5–6):719–739, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901374.

Shakeri:2007:NSB

[SD07] Fatemeh Shakeri and Mehdi Dehghan. Numerical solution of a biological population model using He’s variational iteration method. Computers and Mathematics with Applications, 54(7– 8):1197–1209, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210700291X.

Shakeri:2008:MLS

[SD08] Fatemeh Shakeri and Mehdi Dehghan. The method of lines for solution of the one-dimensional wave equation subject to an integral conservation condition. Computers and Mathematics with Applications, 56(9):2175–2188, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003520.

Saadatmandi:2009:VIM

[SD09] Abbas Saadatmandi and Mehdi Dehghan. Variational iteration method for solving a generalized pantograph equation. Computers and Mathematics with Applications, 58(11–12):2190–2196, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001503.

Shin:2009:SEN

[SDB09] Byeong-Chun Shin, M. T. Darvishi, and A. Barati. Some exact and new solutions of the Nizhnik– Novikov–Vesselov equation using the Exp-function method. Computers and Mathematics with Applications, 58(11–12):2147–2151, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001400.

Su:2008:DCP

[SDCL08] Hua Su, Binxiang Dai, Yuming Chen, and Kaiwang Li. Dynamic complexities of a predator–prey model with generalized Holling type III functional response and impulsive effects. Computers and Mathematics with Applications, 56(7):1715–1725, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002496.

Shi:2008:NSS

[SDDC08] Baochang Shi, Bin Deng, Rui Du, and Xingwang Chen. A new scheme for source term in LBGK model for convection–diffusion equation. Computers and Mathematics with Applications, 55(7):1568–1575, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006220.

Savula:2001:HMM

[SDK01] Y. H. Savula, I. I. Dyyak, and V. V. Krevs. Heterogeneous mathematical models in numerical analysis of structures. Computers and Mathematics with Applications, 42(8– 9):1201–1216, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002334.

Saadatjoo:2009:SCL

[SDK09] F. Saadatjoo, Vali Derhami, and S. M. Karbassi. Simultaneous control of linear systems by state feedback. Computers and Mathematics with Applications, 58(1):154– 160, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002399.

Saadati:2009:CHV

[SDVS09] R. Saadati, M. Dehghan, S. M. Vaezpour, and M. Saravi. The convergence of He’s variational iteration method for solving integral equations. Computers and Mathematics with Applications, 58(11–12):2167–2171, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001448.

Shu:2006:NCL

[SDZ06] C. Shu, H. Ding, and N. Zhao. Numerical comparison of least square-based finite-difference (LSFD) and radial basis function-based finite-difference (RBFFD) methods. Computers and Mathematics with Applications, 51(8):1297–1310, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210600085X.

Serra:2000:ACP

[Ser00]  S. Serra.  The analy-

sis of a Chebyshev problem via spectral matrix theory. Computers and Mathematics with Applications, 39(1–2):55–68, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003132.

Saied:2009:GWE

[SERG09] E. A. Saied, Reda G. Abd El-Rahman, and Marwa I. Ghonamy. A generalized Weierstrass elliptic function expansion method for solving some nonlinear partial differential equations. Computers and Mathematics with Applications, 58(9):1725–1735, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005768.

Seshaiyer:2003:SCN

[Ses03]  P. Seshaiyer.  Stabil-

ity and convergence of nonconforming hp finiteelement methods. Computers and Mathematics with Applications, 46(1):165– 182, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900873.

Srivastava:2001:TPO

[SF01] V. K. Srivastava and A. Fahim. A two-phase optimization procedure for integer programming problems. Computers and Mathematics with Applications, 42(12):1585–1595, December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002644.

Srivastava:2007:OMS

[SF07] Vijaya K. Srivastava and Atef Fahim. An optimization method for solving mixed discrete-continuous programming problems. Computers and Mathematics with Applications, 53(10): 1481–1491, May 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107001010.

Song:2002:TMF

[SFL02] Shiji Song, Chunbo Feng, and E. S. Lee. Triple I method of fuzzy reasoning. Computers and Mathematics with Applications, 44(12):1567–1579, December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002791.

See erratum [Liu09a] and comments [hYL09].

Sadighi:2007:ESN

[SG07] A. Sadighi and D. D.

Ganji. Exact solutions of nonlinear diffusion equations by variational iteration method. Computers and Mathematics with Applications, 54(7–8):1112–1121, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002957.

Sikdar:2008:BAB

[SG08] K. Sikdar and U. C.

Gupta. On the batch arrival batch service queue with finite buffer under server’s vacation: M^{X}/G^{Y} /1/N queue. Computers and Mathematics with Applications, 56(11):2861–2873, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004641.

Safari:2009:AHV

[SGM09] M. Safari, D. D. Ganji, and M. Moslemi. Application of He’s variational iteration method and Adomian’s decomposition method to the fractional KdV–Burgers– Kuramoto equation. Computers and Mathematics with Applications, 58(11– 12):2091–2097, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S089812210900131X.

Sturmer:2009:FFS

[SGR^{+}09] Markus Stürmer, Jan Götz, Gregor Richter, Arnd Dörfler, and Ulrich Rüde. Fluid flow simulation on the Cell Broadband Engine using the lattice Boltzmann method. Computers and Mathematics with Applications, 58(5):1062–1070, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002442.

Snyman:2001:SQS

[SH01] J. A. Snyman and A. M. Hay. The spherical quadratic steepest descent (SQSD) method for unconstrained minimization with no explicit line searches. Computers and Mathematics with Applications, 42(1–2):169– 178, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001419.

Snyman:2002:DOM

[SH02] J. A. Snyman and A. M.

Hay. The Dynamic-Q optimization method: an alternative to SQP? Computers and Mathematics with Applications, 44(12):1589– 1598, December 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210200281X.

Shen:2006:CCT

[SH06] Shu-Qian Shen and Ting-Zhu Huang. Convergence and comparison theorems for double splittings of matrices. Computers and Mathematics with Applications, 51(12):1751– 1760, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001246.

Shampine:2002:VOA

[Sha02] L. F. Shampine. Variable order Adams codes. Computers and Mathematics with Applications, 44(5–6):749– 761, September 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122102001888.

Sharma:2004:FMS

[Sha04] J. R. Sharma. A family of methods for solving nonlinear equations using quadratic interpolation. Computers and Mathematics with Applications, 48(5–6):709–714, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003062.

Shang:2009:DMP

[Sha09] Yueqiang Shang. A distributed memory parallel Gauss–Seidel algorithm for linear algebraic systems. Computers and Mathematics with Applications, 57(8):1369–1376, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900042X.

Shen:2000:ADE

[She00] Zhizhang Shen. The average diameter and its estimation in nonlinear structures. Computers and Mathematics with Applications, 40(2–3):261–276, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001590.

Shen:2002:CAD

[She02] Zhizhang Shen. The calculation of average distance in mesh structures. Computers and Mathematics with Applications, 44(10– 11):1379–1402, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210200264X.

Shekhtman:2004:PI

[She04a]  B. Shekhtman.  Poly-

nomial interpolation in R_{3}. Computers and Mathematics with Applications, 48(9):1299–1304, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003761.

Shevchenko:2004:MDO

[She04b] I. Shevchenko. Minimizing the distance to one evader while chasing another. Computers and Mathematics with Applications, 47(12):1827–1855, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840203.

Sheen:2005:FEC

[She05] J. N. Sheen. Fuzzy evaluation of cogeneration alternatives in a petrochemical industry. Computers and Mathematics with Applications, 49(5–6):741–755, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210500088X.

Shi:2003:OSA

[Shi03]  Yuming Shi.  Oscilla-

tion of self-adjoint secondorder vector difference equations to the parameter. Computers and Mathematics with Applications, 45 (10–11):1591–1600, May/ June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210300138X.

Shih:2005:FAM

[Shi05] Hsu-Shih Shih. Fuzzy approach to multilevel knapsack problems. Computers and Mathematics with Applications, 49(7–8):1157– 1176, May/June 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210500129X.

Siddiqui:2009:TFT

[SHI09] A. M. Siddiqui, T. Haroon, and S. Irum. Torsional flow of third grade fluid using modified homotopy perturbation method. Computers and Mathematics with Applications, 58(11–12):2274–2285, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001631.

Saravanan:2008:BCC

[SHKL08] S. Saravanan, A. K. Abdul Hakeem, P. Kandaswamy, and J. Lee. Buoyancy convection in a cavity with mutually orthogonal heated plates. Computers and Mathematics with Applications, 55(12):2903– 2912, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008127.

Shou:2009:HPM

[Sho09a] Da-Hua Shou. The homotopy perturbation method for nonlinear oscillators.

Computers and Mathematics with Applications, 58(11–12):2456–2459, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001904.

Shou:2009:VAN

[Sho09b] Da-Hua Shou. Variational approach for nonlinear oscillators with discontinuities. Computers and Mathematics with Applications, 58(11–12):2416–2419, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001758.

Shih:2005:RSP

[SHS05] Hsu-Shih Shih, Liang-Chih Huang, and Huan-Jyh Shyur. Recruitment and selection processes through an effective GDSS. Computers and Mathematics with Applications, 50(10– 12):1543–1558, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105004268.

Shi:2005:CBA

[SHWF05] Yong Shi, Jing He, Lei Wang, and Wei Fan. Computerbased algorithms for multiple criteria and multiple constraint level integer linear programming. Computers and Mathematics with Applications, 49(5–6):903–921, April/ May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001008.

Sun:2009:ASQ

[SHWF09] Li Sun, Guoping He, Yongli Wang, and Liang Fang. An active set quasi-Newton method with projected search for bound constrained minimization. Computers and Mathematics with Applications, 58(1): 161–170, July 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109002478.

Shang:2008:NET

[SHY08] Yadong Shang, Yong Huang, and Wenjun Yuan. New exact traveling wave solutions for the Klein– Gordon–Zakharov equations. Computers and Mathematics with Applications, 56(5):1441–1450, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001697.

Sandoh:2003:OIP

[SI03] H. Sandoh and N. Igaki. Optimal inspection policies for a scale. Computers and Mathematics with Applications, 46(7):1119–1127, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901271.

Sidi:2008:VEM

[Sid08] Avram Sidi. Vector extrapolation methods with applications to solution of large systems of equations and to PageRank computations. Computers and Mathematics with Applications, 56(1): 1–24, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107008188.

Siegmund:2003:NFN

[Sie03] S. Siegmund. Normal forms for nonautonomous difference equations. Computers and Mathematics with Applications, 45(6–9):1059– 1073, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0898122103000853.

Simos:2001:BNF

[Sim01] T. E. Simos. Bessel and Neumann fitted methods for the numerical solution of the Schrödinger equation. Computers and Mathematics with Applications, 42 (6–7):833–847, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002024.

Sirotkin:2000:TLP

[Sir00]  V. Sirotkin.  Two-level

parallel Schwarz methods for singularly perturbed semilinear elliptic problems. Computers and Mathematics with Applications, 40(4–5):645–667, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001851.

Shi:2009:SCS

[SJJ09] Huan-Nan Shi, Yong-Ming Jiang, and Wei-Dong Jiang. Schur-convexity and Schurgeometrically concavity of Gini means. Computers and Mathematics with Applications, 57(2):266–274, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006585.

Syau:2008:CFM

[SJL08] Yu-Ru Syau, Lixing Jia, and E. Stanley Lee. Φ_{1}- concavity and fuzzy multiple objective decision making. Computers and Mathematics with Applications, 55(6):1181–1188, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005871.

Syau:2009:GCV

[SJL09] Yu-Ru Syau, Lixing Jia, and E. Stanley Lee. Generalizations of E-convex and B-vex functions. Computers and Mathematics with Applications, 58(4):711–716, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003241.

Shawagfeh:2004:SSP

[SK04] N. Shawagfeh and D. Kaya. Series solution to the Pochhammer–Chree equation and comparison with exact solutions. Computers and Mathematics with Applications, 47(12):1915– 1920, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840276.

Satow:2006:HTM

[SK06a] Takashi Satow and Hajime Kawai. Hit and target models for DNA damage with indirect action. Computers and Mathematics with Applications, 51(2):257–268, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/0898122105004633.

Stynes:2006:NAS

[SK06b] M. Stynes and N. Kopteva.

Numerical analysis of singularly perturbed nonlinear reaction-diffusion problems with multiple solutions. Computers and Mathematics with Applications, 51(5):857–864, March 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210600023X.

Suga:2006:ETC

[SK06c] Ryosuke Suga and Mutsuto Kawahara. Estimation of tidal current using Kalman filter finite-element method. Computers and Mathematics with Applications, 52 (8–9):1289–1298, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210600318X.

Sweilam:2009:ESS

[SK09] N. H. Sweilam and M. M. Khader. Exact solutions of some coupled nonlinear partial differential equations using the homotopy perturbation method. Computers and Mathematics with Applications, 58(11–12):2134–2141, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001382.

Smyrlis:2001:SAO

[SKG01] Y.-S. Smyrlis, A. Karageorghis, and G. Georgiou. Some aspects of the one-dimensional version of the method of fundamental solutions. Computers and Mathematics with Applications, 41(5–6):647– 657, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003084.

Sampo:2006:MTS

[SKHK06] J. Sampo, J.-K. Kamarainen, M. Heiliö, and H. Kälviäinen. Measuring translation shiftability of frames. Computers and Mathematics with Applications, 52(6– 7):1089–1098, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210600280X.

Sakawa:2001:OPD

[SKU01] M. Sakawa, K. Kato, and S. Ushiro. Operation planning of district heating and cooling plants through genetic algorithms for nonlinear 0–1 programming. Computers and Mathematics with Applications, 42(10– 11):1365–1378, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002462.

Schandl:2002:NBA

[SKW02] B. Schandl, K. Klamroth, and M. M. Wiecek. Norm-based approximation in multicriteria programming. Computers and Mathematics with Applications, 44(7):925–942, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002043.

Shen:2000:SOC

[SL00a] Jianhua Shen and Zhiguo Luo. Some oscillation criteria for difference equations. Computers and Mathematics with Applications, 40 (6–7):713–719, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001905.

Shyu:2000:APV

[SL00b] Shyong-Jian Shyu and B. M. T. Lin. An application of parallel virtual machine framework to film production problem. Computers and Mathematics with Applications, 39(12): 53–62, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001292.

Syau:2006:FCM

[SL06a] Yu-Ru Syau and E. Stanley Lee. Fuzzy convexity and multiobjective convex optimization problems. Computers and Mathematics with Applications, 52(3–4):351–362, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002021.

Syau:2006:FWT

[SL06b] Yu-Ru Syau and E. Stanley Lee. Fuzzy weirstrass theorem and convex fuzzy mappings. Computers and Mathematics with Applications, 51(12):1741–1750, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001222.

Syau:2006:PCF

[SL06c] Yu-Ru Syau and E. Stanley Lee. Preinvexity and Φ_{1}-convexity of fuzzy mappings through a linear ordering. Computers and Mathematics with Applications, 51(3–4):405–418, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005328.

Si:2007:LAS

[SL07a] Jianguo Si and Tongbo Liu. Local analytic solutions of a functional differential equation with a deviating argument depending on the state derivative near resonance. Computers and Mathematics with Applications, 54(6):750–762, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002635.

Sun:2007:EMP

[SL07b] Jian-Ping Sun and Wan-Tong Li. Existence and multiplicity of positive solutions to nonlinear firstorder PBVPs on time scales. Computers and Mathematics with Applications, 54(6): 861–871, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107003306.

Sigalotti:2008:AKE

[SL08] Leonardo Di G. Sigalotti and Hender López. Adaptive kernel estimation and SPH tensile instability. Computers and Mathematics with Applications, 55(1):23–50, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003288.

Sevli:2009:ASF

[SL09] H. Sevli̇ and L. Leindler. On the absolute summability factors of infinite series involving quasi-power-increasing sequences. Computers and Mathematics with Applications, 57(5): 702–709, March 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108006822.

Li:2009:TDF

[sLhYL09] Xiao shen Li, Xue hai Yuan, and E. Stanley Lee. The three-dimensional fuzzy sets and their cut sets. Computers and Mathematics with Applications, 58(7):1349–1359, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002417.

Syau:2004:CUS

[SLJ04] Yu-Ru Syau, E. S. Lee, and Lixing Jia. Convexity and upper semicontinuity of fuzzy sets. Computers and Mathematics with Applications, 48(1–2):117– 129, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840033.

Singer:2003:HAP

[SLL03] B. A. Singer, D. P. Lockard, and G. M. Lilley. Hybrid acoustic predictions.

Computers and Mathematics with Applications, 46 (4):647–669, August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S089812210390023X.

Slota:2007:DIO

[Slo07]  Damian Slota.  Direct

and inverse one-phase Stefan problem solved by the variational iteration method. Computers and Mathematics with Applications, 54(7–8):1139–1146, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003045.

Slota:2009:ESH

[Slo09] Damian Slota. Exact solution of the heat equation with boundary condition of the fourth kind by He’s variational iteration method. Computers and Mathematics with Applications, 58(11–12):2495–2503, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002004.

Shih:2009:BBP

[SLTH09] Yuan-Kang Shih, Cheng-Kuan Lin, Jimmy J. M.

Tan, and Lih-Hsing Hsu. The bipanpositionable bipancyclic property of the hypercube. Computers and Mathematics with Applications, 58(9):1722–1724, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006117.

Srivastava:2000:AFC

[SM00] H. M. Srivastava and A. K. Mishra. Applications of fractional calculus to parabolic starlike and uniformly convex functions. Computers and Mathematics with Applications, 39(3–4):57–69, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003338.

Sobol:2002:CEC

[SM02] I. M. Sobol’ and E. E.

Myshetskaya. Convergence estimates for crude approximations of a Pareto set. Computers and Mathematics with Applications, 44(7):877–886, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002006.

Spartalis:2006:HAB

[SM06] S. I. Spartalis and C. Mamaloukas. Hyperstructures associated with binary relations. Computers and Mathematics with Applications, 51(1):41–50, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105004499.

Sun:2008:LBS

[SM08] Chenghai Sun and Lance L. Munn. Lattice-Boltzmann simulation of blood flow in digitized vessel networks. Computers and Mathematics with Applications, 55(7):1594–1600, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006402.

Shen:2009:MMA

[SM09] Yue-Yun Shen and Lu-Feng Mo. The max-min approach to a relativistic equation. Computers and Mathematics with Applications, 58(11–12):2131–2133, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001370.

Smart:2001:CDF

[Sma01] N. P. Smart. A comparison of different finite fields for elliptic curve cryptosystems. Computers and Mathematics with Applications, 42(1–2):91– 100, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210100133X.

Sajid:2008:FES

[SMH08] M. Sajid, R. Mahmood, and T. Hayat. Finite element solution for flow of a third grade fluid past a horizontal porous plate with partial slip. Computers and Mathematics with Applications, 56(5):1236–1244, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001259.

Samarskii:2002:MDS

[SMMM02] A. A. Samarskii, P. P. Matus, V. I. Mazhukin, and I. E. Mozolevski. Monotone difference schemes for equations with mixed derivatives. Computers and Mathematics with Applications, 44(3–4):501–510, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001645.

Suresh:2003:ESD

[SMO03] S. Suresh, V. Mani, and S. N. Omkar. The effect of start-up delays in scheduling divisible loads on bus networks: an alternate approach. Computers and Mathematics with Applications, 46(10– 11):1545–1557, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901908.

Suresh:2005:END

[SMOK05] S. Suresh, V. Mani, S. N. Omkar, and H. J. Kim. An equivalent network for divisible load scheduling in nonblocking mode of communication. Computers and Mathematics with Applications, 49(9–10):1413– 1431, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001604.

Stojanovic:2004:MMM

[SMS^{+}04] N. M. Stojanović, E. I. Milovanović, I. Stojmenović, T. Milovanović, and T. I. Tokić. Mapping matrix multiplication algorithm onto fault-tolerant systolic array. Computers and Mathematics with Applications, 48(1–2):275– 289, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840148.

Sales-Mayor:2004:TSF

[SMW04] F. Sales-Mayor and R. E.

Wyatt. A two-stage filter for smoothing multivariate noisy data on unstructured grids. Computers and Mathematics with Applications, 47(6–7):877–891, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900727.

Safjan:2001:ICI

[SN01] A. Safjan and M. Newman. The ill-conditioning of infinite element stiffness matrices. Computers and Mathematics with Applications, 41(10–11):1263–1291, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101000979.

Safjan:2002:TDI

[SN02] A. Safjan and M. Newman. Three-dimensional infinite elements utilizing basis functions with compact support. Computers and Mathematics with Applications, 43(8–9):981–1002, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800086.

Shi:2003:OFC

[SN03] Peng Shi and Sing Kiong Nguang. H_{∞} output feedback control of fuzzy system models under sampled measurements. Computers and Mathematics with Applications, 46(5–6):705–717, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901362.

Sweilam:2009:EDS

[SNA09] N. H. Sweilam, A. M. Nagy, and M. H. Alnasr. An efficient dynamical systems method for solving singularly perturbed integral equations with noise. Computers and Mathematics with Applications, 58(7):1418–1424, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004945.

Sanchez:2006:PND

[SNG06] L. E. Sánchez, D. K. Nagar, and A. K. Gupta. Properties of noncentral Dirichlet distributions. Computers and Mathematics with Applications, 52(12):1671–1682, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003506.

Saberi-Nadjafi:2009:HHP

[SNG09] Jafar Saberi-Nadjafi and Asghar Ghorbani. He’s homotopy perturbation method: an effective tool for solving nonlinear integral and integro-differential equations. Computers and Mathematics with Applications, 58(11–12):2379–2390, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001709.

Saberi-Nadjafi:2008:VIM

[SNT08] Jafar Saberi-Nadjafi and Mohamadreza Tamamgar. The variational iteration method: a highly promising method for solving the system of integro-differential equations. Computers and Mathematics with Applications, 56(2):346–351, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000205.

Snyman:2000:LLF

[Sny00]  J. A. Snyman.  The

LFOPC leap-frog algorithm for constrained optimization. Computers and Mathematics with Applications, 40 (8–9):1085–1096, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210085018X.

Satow:2003:ORP

[SO03] T. Satow and S. Osaki. Optimal replacement policies for a two-unit system with shock damage interaction. Computers and Mathematics with Applications, 46(7):1129–1138, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901283.

Sohn:2005:NTM

[Soh05] Sung-Ik Sohn. A new TVD– MUSCL scheme for hyperbolic conservation laws. Computers and Mathematics with Applications, 50 (1–2):231–248, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105002452.

Solomon:2000:GED

[Sol00] G. Solomon. Golay encoding/decoding via BCH– Hamming. Computers and Mathematics with Applications, 39(11):103–108, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001140.

Song:2000:ECF

[Son00]  Hong-Yeop Song.  The

existence of circular Florentine arrays. Computers and Mathematics with Applications, 39(11):31– 35, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001048.

Song:2008:NSC

[Son08]  Yisheng Song.  New

strong convergence theorems for nonexpansive nonself-mappings without boundary conditions. Computers and Mathematics with Applications, 56(6): 1473–1478, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108001739.

Song:2009:PAP

[Son09] Yihong Song. Positive almost periodic solutions of nonlinear discrete systems with finite delay. Computers and Mathematics with Applications, 58(1):128– 134, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000522.

Soto:2001:SCE

[Sot01]  R. L. Soto.  A suffi-

cient condition for the existence of an M-matrix with prescribed spectrum. Computers and Mathematics with Applications, 42 (6–7):849–859, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002036.

Soto:2002:ISP

[Sot02] R. L. Soto. The inverse spectrum problem for positive generalized stochastic matrices. Computers and Mathematics with Applications, 43(6–7):641–656, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003108.

Sun:2004:IOT

[SOW04] Yuan Gong Sun, C. H. Ou, and J. S. W. Wong. Interval oscillation theorems for a second-order linear differential equation. Computers and Mathematics with Applications, 48(10– 11):1693–1699, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004225.

Shen:2008:BTA

[SP08a] Jie Shen and Li-Ping Pang.

A bundle-type auxiliary problem method for solving generalized variationallike inequalities. Computers and Mathematics with Applications, 55(12):2993– 2998, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008279.

Singh:2008:SCT

[SP08b] S. L. Singh and Bhagwati Prasad. Some coincidence theorems and stability of iterative procedures. Computers and Mathematics with Applications, 55(11):2512– 2520, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007523.

Stratford:2008:PSP

[SP08c] Kevin Stratford and Ignacio Pagonabarraga. Parallel simulation of particle suspensions with the lattice Boltzmann method. Computers and Mathematics with Applications, 55 (7):1585–1593, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210700644X.

Song:2009:GAT

[SP09] Wei Song and Soon Cheol Park. Genetic algorithm for text clustering based on latent semantic indexing. Computers and Mathematics with Applications, 57(11–12):1901– 1907, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005300.

Shidfar:2006:NMS

[SPE06] A. Shidfar, R. Pourgholi, and M. Ebrahimi. A numerical method for solving of a nonlinear inverse diffusion problem. Computers and Mathematics with Applications, 52(6– 7):1021–1030, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002756.

Serodio:2001:CZQ

[SPV01] R. Serôdio, E. Pereira, and J. Vitória. Computing the zeros of quaternion polynomials. Computers and Mathematics with Applications, 42(8– 9):1229–1237, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002358.

Stanek:2008:ANS

[SPW08] Svatoslav Stanek, Gernot Pulverer, and Ewa B. Weinmüller. Analysis and numerical simulation of positive and dead-core solutions of singular twopoint boundary value problems. Computers and Mathematics with Applications, 56(7):1820–1837, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002629.

Shanthi:2003:ANM

[SR03] V. Shanthi and N. Ramanujam. Asymptotic numerical methods for singularly perturbed fourthorder ordinary differential equations of reactiondiffusion type. Computers and Mathematics with Applications, 46(2–3):463–478, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900393.

Shanthi:2004:BVT

[SR04] V. Shanthi and N. Ramanujam. A boundary value technique for boundary value problems for singularly perturbed fourthorder ordinary differential equations. Computers and Mathematics with Applications, 47(10–11):1673–1688, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104839816.

Subramanian:2007:NFT

[SR07] Shankar C. Subramanian and K. R. Rajagopal. A note on the flow through porous solids at high pressures. Computers and Mathematics with Applications, 53(2):260–275, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001046.

Srivastava:2003:RSS

[Sri03]  H. M. Srivastava.  Re-

marks on some series expansions associated with certain products of the incomplete gamma functions. Computers and Mathematics with Applications, 46(12):1749–1759, December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902343.

Sun:2006:LPL

[SS06] Huaqing Sun and Yuming Shi. Limit-point and limitcircle criteria for singular second-order linear difference equations with complex coefficients. Computers and Mathematics with Applications, 52(3–4):539– 554, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002173.

Shi:2007:MGM

[SS07a] Zhen-Jun Shi and Jie Shen. Memory gradient method with Goldstein line search. Computers and Mathematics with Applications, 53(1):28–40, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000235.

Spradlin:2007:LET

[SS07b] Christina Spradlin and Greg Spradlin. Lanchester’s equations in three dimensions. Computers and Mathematics with Applications, 53(7):999–1011, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000557.

Sarkar:2009:HAD

[SS09a] Bikash Kanti Sarkar and Shib Sankar Sana. A hybrid approach to design efficient learning classifiers. Computers and Mathematics with Applications, 58(1): 65–73, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900234X.

Shankar:2009:AAE

[SS09b] Maddu Shankar and S. Sundar. Asymptotic analysis of extrapolation boundary conditions for LBM. Computers and Mathematics with Applications, 57 (8):1313–1323, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109000406.

Simic:2009:SFD

[SS09c] Slobodan K. Simić and Zoran Stanić. On some forests determined by their Laplacian or signless Laplacian spectrum. Computers and Mathematics with Applications, 58(1):171– 178, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002806.

Sinha:2009:DRO

[SS09d] Koushik Sinha and Bhabani P. Sinha. On the distribution of runs of ones in binary strings. Computers and Mathematics with Applications, 58(9):1816–1829, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005744.

Srivastava:2009:PSB

[SS09e] V. P. Srivastava and Rashmi Srivastava. Particulate suspension blood flow through a narrow catheterized artery.

Computers and Mathematics with Applications, 58(2): 227–238, July 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109002351.

Srinivasan:2007:ISO

[SSD07] Gopala Krishna Srinivasan, V. D. Sharma, and B. S. Desale. An integrable system of ODE reductions of the stratified Boussinesq equations. Computers and Mathematics with Applications, 53(2):296–304, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002003.

S:2009:NAS

[SSH09] Alvaro H. Salas S., Cesar A. Gomez S., and Jairo Ernesto Castillo Hernańdez. New abundant solutions for the Burgers equation. Computers and Mathematics with Applications, 58(3):514–520, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003319.

Shen:2007:SVF

[SSL07] Judong Shen, Yu-Ru Syau, and E. S. Lee. Support vector fuzzy adaptive network in regression analysis. Computers and Mathematics with Applications, 54(11–12):1353–1366, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002969.

Shi:2009:CEL

[SSLL09] Benlin Shi, Yang Song, Changchen Liu, and Yunfeng Luo. Competition and evolution of linear and two-part tariff. Computers and Mathematics with Applications, 57(11–12):1895– 1900, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005324.

Sahimi:2001:YFE

[SSSH01] M. S. Sahimi, E. Sundararajan, M. Subramaniam, and N. A. A. Hamid. The d’yakonov fully explicit variant of the iterative decomposition method. Computers and Mathematics with Applications, 42(10– 11):1485–1496, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002565.

Sang:2009:PSN

[SSX09] Yanbin Sang, Hua Su, and Fuyi Xu. Positive solutions of nonlinear mpoint BVP for an increasing homeomorphism and homomorphism with sign changing nonlinearity on time scales. Computers and Mathematics with Applications, 58(2):216– 226, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002867.

Shampine:2000:ELO

[ST00] L. F. Shampine and S. Thompson. Event location for ordinary differential equations. Computers and Mathematics with Applications, 39(5–6):43–54, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000456.

Shibata:2007:RCN

[ST07] Yoshihiro Shibata and Koumei Tanaka. Rate of convergence of non-stationary flow to the steady flow of compressible viscous fluid. Computers and Mathematics with Applications, 53(3–4):605–623, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700123X.

Stanek:2002:PSS

[Sta02] S. Stanek. Positive solutions of singular Dirichlet and periodic boundary value problems. Computers and Mathematics with Applications, 43(6–7):681–692, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003133.

Starovoitov:2007:SPS

[Sta07] Victor N. Starovoitov. Solvability of the problem of the self-propelled motion of several rigid bodies in a viscous incompressible fluid. Computers and Mathematics with Applications, 53(3–4):413–435, February 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001885.

Steppeler:2006:HDM

[Ste06] J. Steppeler. HLI, a direct method suitable for partial and fully implicit time integration of primitive equation meteorological models.

Computers and Mathematics with Applications, 52 (8–9):1357–1372, October/ November 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003233.

Steigmann:2007:AFS

[Ste07] David J. Steigmann. Asymptotic finite-strain thinplate theory for elastic solids. Computers and Mathematics with Applications, 53(2):287–295, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001101.

Stevic:2008:DE

[Ste08] Stevo Stević. On the difference equation x_{n+1} = α + x_{n−1}/x_{n}. Computers and Mathematics with Applications, 56(5):1159–1171, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800117X.

Sviercoski:2008:AEC

[STH08] Rosangela F. Sviercoski, Bryan J. Travis, and James M. Hyman. Analytical effective coefficient and a first-order approximation for linear flow through block permeability inclusions. Computers and Mathematics with Applications, 55(9):2118– 2133, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006918.

Stiebler:2008:ADL

[STK08] Maik Stiebler, Jonas Tölke, and Manfred Krafczyk. Advection-diffusion lattice Boltzmann scheme for hierarchical grids. Computers and Mathematics with Applications, 55(7):1576– 1584, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006359.

Santonja:2008:MMP

[STV08] Francisco J. Santonja, Ana C.

Tarazona, and Rafael J. Villanueva. A mathematical model of the pressure of an extreme ideology on a society. Computers and Mathematics with Applications, 56(3):836–846, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000497.

Sun:2000:SCH

[Sun00] Xingping Sun. Scattered cardinal Hermite interpolation by positive definite functions. Computers and Mathematics with Applications, 40(1):165– 169, July 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001504.

Sun:2005:EPS

[Sun05a]  Hong-Rui Sun.  Exis-

tence of positive solutions to second-order time scale systems. Computers and Mathematics with Applications, 49(1):131–145, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210500012X.

Sundarapandian:2005:NODa

[Sun05b] V. Sundarapandian. Nonlinear observer design for a general class of discretetime nonlinear systems with real parametric uncertainty. Computers and Mathematics with Applications, 49(7–8):1177–1194, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001306.

Sundarapandian:2005:NODb

[Sun05c] V. Sundarapandian. Nonlinear observer design for a general class of nonlinear systems with real parametric uncertainty. Computers and Mathematics with Applications, 49(7–8):1195– 1211, May/June 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105001318.

Sundarapandian:2006:NCLa

[Sun06a]  V. Sundarapandian.  A

necessary condition for local asymptotic stability of discrete-time nonlinear systems with exogenous disturbance. Computers and Mathematics with Applications, 51(3–4):661–666, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005298.

Sundarapandian:2006:NCLb

[Sun06b]  V. Sundarapandian.  A

necessary condition for local asymptotic stability of nonlinear systems with exogenous disturbance. Computers and Mathematics with Applications, 51(3–4):667–672, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105005304.

Sun:2009:TPS

[Sun09a]  Hong-Rui Sun.  Triple

positive solutions for p-Laplacian m-point boundary value problem on time scales. Computers and Mathematics with Applications, 58(9):1736–1741, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109006087.

Sun:2009:SRS

[Sun09b] Xie-Hua Sun. Some remarks on sequences of binomial type. Computers and Mathematics with Applications, 58(9):1859–1868, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109005963.

Surucu:2008:PCS

[Sür08] Baris Sürücü. A power comparison and simulation study of goodness-of-fit tests. Computers and Mathematics with Applications, 56(6):1617–1625, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001880.

Senato:2001:STI

[SV01] D. Senato and A. Venezia.

A set-theoretic interpretation of the umbral calculus. Computers and Mathematics with Applications, 41(9):1109–1124, May 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101000852.

Sarler:2006:MEL

[SV06] B. Sarler and R. Vertnik.

Meshfree explicit local radial basis function collocation method for diffusion problems. Computers and Mathematics with Applications, 51(8):1269– 1282, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000836.

Siboni:2008:SMA

[SV08] S. Siboni and C. Della Volpe. Some mathematical aspects of the Kelvin equation. Computers and Mathematics with Applications, 55(1):51–65, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700329X.

Sommariva:2009:CAF

[SV09] Alvise Sommariva and Marco Vianello. Computing approximate Fekete points by QR factorizations of Vandermonde matrices. Computers and Mathematics with Applications, 57(8):1324–1336, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000625.

Si:2002:ASS

[SW02a] Jian-Guo Si and Xin-Ping Wang. Analytic solutions of a second-order iterative functional differential equation. Computers and Mathematics with Applications, 43(1–2):81–90, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002735.

See erratum [SW02b].

Si:2002:EAS

[SW02b] Jian-Guo Si and Xin-Ping Wang. Erratum to “Analytic Solutions of a Second-Order Iterative Functional Differential Equation”. Computers and Mathematics with Applications, 44(5–6):833, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001955.

See [SW02a].

Shan:2004:OND

[SW04] Wenrui Shan and Ge Weigao.

Oscillation of neutral difference equations with positive and negative coefficients. Computers and Mathematics with Applications, 47(10–11):1647–1657, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839786.

Sun:2006:ADG

[SW06] Shuyu Sun and M. F.

Wheeler. Analysis of discontinuous Galerkin methods for multicomponent reactive transport problems. Computers and Mathematics with Applications, 52(5): 637–650, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106002458.

Shi:2007:EMB

[SW07] Junping Shi and Shin-Hwa Wang. Exact multiplicity of boundary blow-up solutions for a bistable problem. Computers and Mathematics with Applications, 54(9–10):1285–1292, November 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003665.

Song:2008:EMI

[SW08] Yisheng Song and Hongjun Wang. Erratum to “Mann and Ishikawa iterative processes for multivalued mappings in Banach spaces” [Comput. Math. Appl. 54 (2007) 872–877]. Computers and Mathematics with Applications, 55(12):2999– 3002, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008140.

See [Pan07c].

She:2009:AAF

[SW09] Yan-Hong She and Guo-Jun Wang. An axiomatic approach of fuzzy rough sets based on residuated lattices. Computers and Mathematics with Applications, 58(1):189– 201, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002648.

Sweilam:2007:FOI

[Swe07] N. H. Sweilam. Fourth order integro-differential equations using variational iteration method. Computers and Mathematics with Applications, 54(7–8):1086– 1091, October 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107002842.

Shang:2007:FNH

[SWK07] Pengjian Shang, Meng Wan, and Santi Kama. Fractal nature of highway traffic data. Computers and Mathematics with Applications, 54(1):107– 116, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001897.

Shih:2008:VBG

[SWKT08] Yuan-Kang Shih, Yi-Chien Wu, Shin-Shin Kao, and Jimmy J. M. Tan. Vertexbipancyclicity of the generalized honeycomb tori. Computers and Mathematics with Applications, 56 (11):2848–2860, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108004513.

Shih:2004:NNA

[SWL^{+}04] Hsu-Shih Shih, Ue-Pyng Wen, S. Lee, Kuen-Ming Lan, and Han-Chyi Hsiao. A neural network approach to multiobjective and multilevel programming problems. Computers and Mathematics with Applications, 48(1–2):95– 108, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210484001X.

Song:2005:AES

[SWL05] Shiji Song, Cheng Wu, and E. S. Lee. Asymptotic equilibrium and stability of fuzzy differential equations. Computers and Mathematics with Applications, 49(7–8):1267–1277, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001367.

Siddikov:2005:NSA

[SWS05] B. M. Siddikov, B. A. Wade, and D. H. Schultz. Numerical simulation of the active magnetic regenerator. Computers and Mathematics with Applications, 49(9–10):1525– 1538, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001665.

Song:2006:EUC

[SWX06] Shiji Song, Cheng Wu, and Xiaoping Xue. Existence and uniqueness of Cauchy problem for fuzzy differential equations under dissipative conditions. Computers and Mathematics with Applications, 51(9–10):1483– 1492, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000952.

Sun:2006:PCP

[SX06] Taixiang Sun and Hongjian Xi. The periodic character of positive solutions of the difference equation χ_{n+1} = (χ_{n}, χ_{n−k}). Computers and Mathematics with Applications, 51(9–10):1431– 1436, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000897.

Suri:2003:D

[SY03a] Manil Suri and Zohar Yosibash. Dedication. Computers and Mathematics with Applications, 46(1):xv– xvi, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900769.

Suri:2003:P

[SY03b] Manil Suri and Zohar Yosibash. Preface. Computers and Mathematics with Applications, 46(1): xiii, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900757.

Sheu:2005:WSA

[SY05] Shey-Huei Sheu and Shiueh-Ling Yu. Warranty strategy accounts for bathtub failure rate and random minimal repair cost. Computers and Mathematics with Applications, 49(7–8):1233–1242, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001331.

Su:2009:MSM

[SY09] Ke Su and Zhensheng Yu. A modified SQP method with nonmonotone technique and its global convergence. Computers and Mathematics with Applications, 57(2):240–247, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004148.

Syam:2001:RMA

[Sya01a] M. I. Syam. The resultants method for approximating real fixed points of polynomials. Computers and Mathematics with Applications, 41(7–8):879– 891, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003266.

Syau:2001:DCF

[Sya01b]  Yu-Ru Syau.  Differen-

tiability and convexity of fuzzy mappings. Computers and Mathematics with Applications, 41(1–2):73–81, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101850070.

Syam:2002:IPI

[Sya02]  M. I. Syam.  Inter-

polation predictors over implicitly defined curves. Computers and Mathematics with Applications, 44 (8–9):1067–1076, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002158.

Su:2006:DMD

[SYHC06] Chad-Ton Su, Chien-Hsin Yang, Kuang-Hung Hsu, and Wen-Ko Chiu. Data mining for the diagnosis of type II diabetes from three-dimensional body surface anthropometrical scanning data. Computers and Mathematics with Applications, 51(6–7):1075–1092, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000381.

Shang:2009:SOC

[SYJJ09] L. Shang, S. Y. Yu, X. Y. Jia, and Y. S. Ji. Selection and optimization of cut-points for numeric attribute values. Computers and Mathematics with Applications, 57(6):1018– 1023, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005658.

Song:2005:SMA

[SYL05] M. H. Song, Z. W. Yang, and M. Z. Liu. Stability of θ-methods for advanced differential equations with piecewise continuous arguments. Computers and Mathematics with Applications, 49(9–10):1295– 1301, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001501.

Synowiec:2008:JDM

[Syn08] Daniel Synowiec. Jumpdiffusion models with constant parameters for financial log-return processes. Computers and Mathematics with Applications, 56(8):2120–2127, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003477.

Sun:2005:MPS

[SZ05] Jian-Ping Sun and Ya-Hong Zhao. Multiplicity of positive solutions of a class of nonlinear fractional differential equations. Computers and Mathematics with Applications, 49(1):73–80, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000076.

Szidarovszky:2006:DOI

[SZ06] F. Szidarovszky and J. Zhao. Dynamic oligopolies with intertemporal demand interaction. Computers and Mathematics with Applications, 52(12):1623–1626, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003488.

Slota:2009:NAH

[SZ09a] Damian Slota and Adam Zielonka. A new application of He’s variational iteration method for the solution of the one-phase Stefan problem. Computers and Mathematics with Applications, 58(11–12):2489–2494, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001965.

Sun:2009:GCM

[SZ09b] Juhe Sun and Liwei Zhang. A globally convergent method based on Fischer–Burmeister operators for solving secondorder cone constrained variational inequality problems. Computers and Mathematics with Applications, 58(10):1936–1946, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006099.

Sun:2009:DBD

[SZJ09] Gui-Quan Sun, Guang Zhang, and Zhen Jin. Dynamic behavior of a discrete modified Ricker & Beverton–Holt model. Computers and Mathematics with Applications, 57(8): 1400–1412, April 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109000261.

Sun:2002:IBV

[SZL02] Ying-Fei Sun, Mei-Ke Zhou, and Yan-Da Li. The initial boundary value problem for thin-plate flow of incompressible non-Newtonian viscous fluids. Computers and Mathematics with Applications, 44(8– 9):1207–1217, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002274.

Thandapani:2001:FON

[TA01] E. Thandapani and I. M.

Arockiasamy. Fourt-order nonlinear oscillations of difference equations. Computers and Mathematics with Applications, 42(3–5):357– 368, August/September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122101001602.

Tiryaki:2004:OTC

[TA04] A. Tiryaki and B. Ayanlar. Oscillation theorems for certain nonlinear differential equations of second order. Computers and Mathematics with Applications, 47(1):149–159, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900132.

Thavaneswaran:2006:PNF

[TA06] A. Thavaneswaran and S. S. Appadoo. Properties of a new family of volatility sign models. Computers and Mathematics with Applications, 52 (6–7):809–818, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002598.

Tu:2009:DDS

[TaL09] Lilan Tu and Jun an Lu. Delay-dependent synchronization in general complex delayed dynamical networks. Computers and Mathematics with Applications, 57(1):28–36, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005907.

Tang:2000:ESD

[Tan00]  B. Tang.  Evaluation

of some DNA cloning strategies. Computers and Mathematics with Applications, 39(11):43–48, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001061.

Tang:2001:GKS

[Tan01]  Hua-Zhong Tang.  Gas-

kinetic schemes for compressible flow of real gases. Computers and Mathematics with Applications, 41 (5–6):723–734, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100003163.

Tang:2002:ABS

[Tan02a]  X. H. Tang.  Asymp-

totic behavior of solutions for neutral difference equations. Computers and Mathematics with Applications, 44(3–4):301–315, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001499.

Tang:2002:BOS

[Tan02b]  X. H. Tang.  Bounded

oscillation of second-order delay difference equations of unstable type. Computers and Mathematics with Applications, 44(8– 9):1147–1156, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002225.

Tang:2003:IPL

[Tan03a]  H.-C. Tang.  Inconsis-

tent property of Lee and Li fuzzy ranking method. Computers and Mathematics with Applications, 45 (4–5):709–713, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000300.

Tang:2003:SDA

[Tan03b] Hui-Chin Tang. Simulated division with approximate factoring for the multiple recursive generator with both unrestricted multiplier and non-Mersenne prime modulus. Computers and Mathematics with Applications, 46 (8–9):1173–1181, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902094.

Tang:2005:EER

[Tan04]  Hui-Chin Tang.  Effec-

tive and efficient restriction on producing the multipliers for the multiple recursive random number generator. Computers and Mathematics with Applications, 47(8–9):1309–1315, April/ May 16, 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901241.

Tang:2007:LES

[Tan07]  Zhongwei Tang.  On

the least energy solutions of nonlinear Schrödinger equations with electromagnetic fields. Computers and Mathematics with Applications, 54(5):627–637, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002416.

Tao:2006:NTM

[Tao06]  Yan Tao.  Newton-type

method for a class of mathematical programs with complementarity constraints. Computers and Mathematics with Applications, 52 (12):1627–1638, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122106003464.

Tao:2009:SBS

[Tao09] Zhao-Ling Tao. Solving the breaking soliton equation by He’s variational method. Computers and Mathematics with Applications, 58(11–12):2395–2397, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001771.

Thiagarajah:2007:OVM

[TAT07] K. Thiagarajah, S. S. Appadoo, and A. Thavaneswaran. Option valuation model with adaptive fuzzy numbers. Computers and Mathematics with Applications, 53(5):831–841, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000375.

Taylor:2000:P

[Tay00] Herbert Taylor. Preface.

Computers and Mathematics with Applications, 39 (11):xiii–xiv, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100000997.

Taylor:2001:UPP

[Tay01]  B. D. Taylor.  Umbral

presentations for polynomial sequences. Computers and Mathematics with Applications, 41(9):1085– 1098, May 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101000839.

Tadei:2008:SIM

[TB08] Roberto Tadei and Nicola Bellomo. Special issue on modeling and computational methods in genomic sciences. Computers and Mathematics with Applications, 55(5):863–866, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005536.

Talo:2009:DDC

[TB09] Özer Talo and Feyzi Basar.

Determination of the duals of classical sets of sequences of fuzzy numbers and related matrix transformations. Computers and Mathematics with Applications, 58(4):717–733, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900306X.

Tekitek:2009:TPM

[TBDL09] M. M. Tekitek, M. Bouzidi, F. Dubois, and P. Lallemand. Towards perfectly matching layers for lattice Boltzmann equation. Computers and Mathematics with Applications, 58(5): 903–913, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109000959.

Tiryaki:2005:ICO

[TBG05] A. Tiryaki, Y. Bašci, and I. Güleç. Interval criteria for oscillation of second-order functional differential equations. Computers and Mathematics with Applications, 50 (8–9):1487–1498, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003962.

Turek:2001:FIR

[TBR01] S. Turek, Ch. Becker, and A. Runge. The FEAST indices-realistic evaluation of modern software components and processor technologies. Computers and Mathematics with Applications, 41(10–11):1431–1464, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001080.

Tiryaki:2004:IAO

[TÇ04] A. Tiryaki and D. Çakmak.

Integral averages and oscillation criteria of secondorder nonlinear differential equations. Computers and Mathematics with Applications, 47(10–11):1495–1506, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839671.

Tarazaga:2009:PGM

[TC09] Pablo Tarazaga and Diego Cuellar. Preconditioners generated by minimizing norms. Computers and Mathematics with Applications, 57(8):1305– 1312, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900056X.

Tian:2007:APC

[TCH07] Y. X. Tian, S. S. Chang, and J. L. Huang. On the approximation problem of common fixed points for a finite family of non-self asymptotically quasi-nonexpansive-type mappings in Banach spaces. Computers and Mathematics with Applications, 53(12):1847– 1853, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001551.

Tian:2008:SCC

[TCX08] Chuanjun Tian, Guanrong Chen, and Shengli Xie. Stability and chaos in a class of finite-dimensional discrete spatiotemporal systems. Computers and Mathematics with Applications, 56(10):2515–2527, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800388X.

Tatari:2007:ICF

[TD07] Mehdi Tatari and Mehdi Dehghan. Identifying a control function in parabolic partial differential equations from overspecified boundary data. Computers and Mathematics with Applications, 53(12):1933– 1942, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001691.

Tatari:2009:IHV

[TD09] Mehdi Tatari and Mehdi Dehghan. Improvement of He’s variational iteration method for solving systems of differential equations. Computers and Mathematics with Applications, 58(11–12):2160–2166, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001424.

Tian:2009:MIC

[TDHL09] Xuan Tian, Xiaoyong Du, He Hu, and Haihua Li. Modeling individual cognitive structure in contextual information retrieval. Computers and Mathematics with Applications, 57(6):1048–1056, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005622.

Tung:2008:NSS

[TDS08] M. M. Tung, E. Defez, and J. Sastre. Numerical solutions of second-order matrix models using cubic-matrix splines. Computers and Mathematics with Applications, 56(10):2561–2571, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003933.

Tsang:2008:ARC

[TDY08] Eric C. C. Tsang, Chen Degang, and Daniel S. Yeung. Approximations and reducts with covering generalized rough sets. Computers and Mathematics with Applications, 56(1):279– 289, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000151.

Tarabia:2006:ETS

[TEB06] A. M. K. Tarabia and A. H. El-Baz. Exact transient solutions of nonempty Markovian queues. Computers and Mathematics with Applications, 52(6– 7):985–996, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002720.

Telyakovskiy:2008:RPA

[Tel08] Aleksey S. Telyakovskiy. Regarding polynomial approximation for ordinary differential equations. Computers and Mathematics with Applications, 55(6):1122– 1128, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005779.

Terceiro:2000:CKF

[Ter00]  J. Terceiro.  Comments

on “Kalman-filtering methods for computing information matrices for time-invariant, periodic, and generally time-varying VARMA models and samples”. Computers and Mathematics with Applications, 40(2–3):405–411, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001681.

See [ZM94].

Tian:2003:SCO

[TG03] Bo Tian and Yi-Tian Gao.

Symbolic computation and observable effect for the (2+ 1)-dimensional symmetric regularized-long-wave equation from strongly magnetized cold-electron plasmas. Computers and Mathematics with Applications, 45(4–5):731–735, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000336.

Tian:2002:SFN

[TGH02] Bo Tian, Yi-Tian Gao, and Woopyo Hong. The solitonic features of a nonintegrable (3 + 1)-dimensional Jimbo– Miwa equation. Computers and Mathematics with Applications, 44(3–4):525– 528, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001669.

Terdik:2002:BAT

[TGIM02] Gy. Terdik, Z. Gál, E. Iglói, and S. Molnár. Bispectral analysis of traffic in highspeed networks. Computers and Mathematics with Applications, 43(12):1575– 1583, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001207.

Tian:2007:PST

[TGS07] Yu Tian, Weigao Ge, and Wenrui Shan. Positive solutions for three-point boundary value problem on the half-line. Computers and Mathematics with Applications, 53(7):1029– 1039, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000582.

Tan:2002:OAP

[TH02] Zhiyi Tan and Yong He. Ordinal algorithms for parallel machine scheduling with nonsimultaneous machine available times. Computers and Mathematics with Applications, 43(12):1521– 1528, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001153.

Tseng:2008:GWO

[THD08] Kuei-Lin Tseng, Shiow Ru Hwang, and S. S. Dragomir. Generalizations of weighted Ostrowski type inequalities for mappings of bounded variation and their applications. Computers and Mathematics with Applications, 55(8):1785–1793, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006517.

Thompson:2001:TMS

[Tho01] H. B. Thompson. Topological methods for some boundary value problems. Computers and Mathematics with Applications, 42(3–5):487–495, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001729.

Tian:2001:GNI

[Tia01] You-Xian Tian. Generalized nonlinear implicit quasivariational inclusions with fuzzy mappings. Computers and Mathematics with Applications, 42(1–2):101– 108, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001341.

Tian:2005:CIT

[Tia05] You-Xian Tian. Convergence of an Ishikawa type iterative scheme for asymptotically quasi-nonexpansive mappings. Computers and Mathematics with Applications, 49(11–12):1905– 1912, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002075.

Tian:2008:ASN

[Tia08] Hongjiong Tian. Asymptotic stability of numerical methods for linear delay parabolic differential equations. Computers and Mathematics with Applications, 56(7):1758–1765, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800254X.

Tseng:2001:CLB

[TJ01] Yuh-Min Tseng and Jinn-Ke Jan. Cryptanalysis of Liaw’s broadcasting cryptosystem. Computers and Mathematics with Applications, 41(12):1575– 1578, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001225.

Tang:2002:LBS

[TK02] Hui-Chin Tang and Chiang Kao. Lower bounds in spectral tests for vectors of nonsuccessive values produced by multiple recursive generator with some zero multipliers. Computers and Mathematics with Applications, 43(8–9):1153–1159, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102800190.

Thandapani:2007:ODS

[TK07] E. Thandapani and P. Mohan Kumar. Oscillation of difference systems of the neutral type. Computers and Mathematics with Applications, 54(4):556–566, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002386.

Tolke:2009:SOI

[TK09] J. Tölke and M. Krafczyk. Second order interpolation of the flow field in the lattice Boltzmann method. Computers and Mathematics with Applications, 58(5): 898–902, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109000947.

Tadi:2002:IMP

[TKC02] M. Tadi, M. V. Klibanov, and Wei Cai. An inversion method for parabolic equations based on quasireversibility. Computers and Mathematics with Applications, 43(8–9):927–941, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800037.

Teng:2000:PAB

[TL00] Zhidong Teng and Zhiming Li. Permanence and asymptotic behavior of the N-species nonautonomous Lotka–Volterra competitive systems. Computers and Mathematics with Applications, 39(7–8):107– 116, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000699.

Tiago:2006:ARB

[TL06] C. M. Tiago and V. M. A.

Leitão. Application of radial basis functions to linear and nonlinear structural analysis problems. Computers and Mathematics with Applications, 51(8):1311– 1334, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000782.

Tang:2007:GDE

[TL07] Yilei Tang and Weigu Li.

Global dynamics of an epidemic model with an unspecified degree. Computers and Mathematics with Applications, 53(11):1704– 1717, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001423.

Tang:2008:NSC

[TL08] X. H. Tang and Xiaoyan Lin. Necessary and sufficient conditions for oscillation of first-order nonlinear neutral difference equations. Computers and Mathematics with Applications, 55(6):1279–1292, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005962.

Li:2008:SSA

[tLjW08] Yao tang Li and Wen jing Wu. Symmetric and skewantisymmetric solutions to systems of real quaternion matrix equations. Computers and Mathematics with Applications, 55(6):1142– 1147, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005809.

Tang:2008:NSD

[TLLM08] Hengsheng Tang, Wei Luo, Xu Li, and Manjun Ma. Nontrivial solutions of discrete elliptic boundary value problems. Computers and Mathematics with Applications, 55(8):1854–1860, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006578.

Tsai:2009:NAA

[TLP^{+}09] J. S. H. Tsai, F.-C. Lu, R. S.

Provence, L. S. Shieh, and Z. Han. A new approach for adaptive blind equalization of chaotic communication: the optimal linearization technique. Computers and Mathematics with Applications, 58(9):1687–1698, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005689.

Tang:2007:OTC

[TLS07] G.-Y. Tang, C. Li, and L. Sun. Optimal tracking control for large-scale interconnected systems with time-delays. Computers and Mathematics with Applications, 53(1):80–88, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000399.

Tsai:2008:DOB

[TLS^{+}08] J. S. H. Tsai, J. H. Lin, L. S. Shieh, C. R. Liu, and S. M. Guo. Decentralized observer-based tracker for analog systems with saturating actuator and state constraints. Computers and Mathematics with Applications, 55(1):1– 22, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002982.

Li:2008:AGE

[tLW08] Yao tang Li and Jiyu Wang. An analysis on the global exponential stability and the existence of periodic solutions for nonautonomous hybrid BAM neural networks with distributed delays and impulses. Computers and Mathematics with Applications, 56(9):2256–2267, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003593.

Thandapani:2003:ABN

[TMG03] E. Thandapani, S. L. Marian, and J. R. Graef. Asymptotic behavior of nonoscillatory solutions of neutral difference equations. Computers and Mathematics with Applications, 45(6–9):1461–1468, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210300107X.

Teraoka:2003:SST

[TOH03] Y. Teraoka, S. Osumi, and H. Hohjo. Some Stackelberg type location game. Computers and Mathematics with Applications, 46(7):1147–1154, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901301.

Topal:2004:SOP

[Top04] S. Gulsan Topal. Secondorder periodic boundary value problems on time scales. Computers and Mathematics with Applications, 48(3–4):637–648, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840835.

Toutounian:2005:DGV

[Tou05] F. Toutounian. Determination of a good value of the time step and preconditioned Krylov subspace methods for the Navier–Stokes equations. Computers and Mathematics with Applications, 49(7–8):1045–1057, May/ June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001215.

Tirani:2003:ASM

[TP03] R. Tirani and C. Paracelli. An algorithm for starting multistep methods. Computers and Mathematics with Applications, 45(1–3):123–129, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103800111.

Taberlet:2008:ESF

[TRD08] Nicolas Taberlet, Patrick Richard, and Renaud Delannay. The effect of sidewall friction on dense granular flows. Computers and Mathematics with Applications, 55(2):230–234, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107003495.

Thandapani:2001:OCT

[TRG01a] E. Thandapani, K. Ravi, and J. R. Graef. Oscillation and comparison theorems for half-linear secondorder difference equations. Computers and Mathematics with Applications, 42 (6–7):953–960, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002115.

Thandapani:2001:OTG

[TRG01b] E. Thandapani, K. Ravi, and J. R. Graef. Oscillation theorems for general quasilinear secondorder difference equations. Computers and Mathematics with Applications, 42(3–5):687–694, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001870.

Trzaska:2001:ESC

[Trz01] Z. W. Trzaska. Explicit solution of certain nonlinear difference equations by using power polynomials. Computers and Mathematics with Applications, 42 (6–7):981–991, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002140.

Takeda:2000:DEA

[TS00a] E. Takeda and J. Satoh. A data envelopment analysis approach to multicriteria decision problems with incomplete information. Computers and Mathematics with Applications, 39(9–10): 81–90, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000882.

Talahmeh:2000:ADR

[TS00b] S. Talahmeh and P. Siy. Arithmetic division in RNS using Galois Field GF(p). Computers and Mathematics with Applications, 39 (5–6):227–238, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100000560.

Thianwan:2006:CCN

[TS06] Sornsak Thianwan and Suthep Suantai. Convergence criteria of a new three-step iteration with errors for nonexpansive nonself-mappings. Computers and Mathematics with Applications, 52(6– 7):1107–1118, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002823.

Tsaur:2003:EIU

[Tsa03] Ruey-Chyn Tsaur. Extrapolating Internet users in Taiwan by risk assessment. Computers and Mathematics with Applications, 46 (10–11):1725–1734, November/December 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103902069.

Tsitouras:2002:ETS

[Tsi02] Ch. Tsitouras. Explicit twostep methods for secondorder linear IVPs. Computers and Mathematics with Applications, 43(8–9):943– 949, April/May 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102800049.

Tsitouras:2003:ENT

[Tsi03] Ch Tsitouras. Explicit Numerov type methods with reduced number of stages. Computers and Mathematics with Applications, 45(1–3):37–42, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103800056.

Turner:2009:CGT

[TSY09] Danny W. Turner, James D. Stamey, and Dean M. Young. Classic group testing with cost for grouping and testing. Computers and Mathematics with Applications, 58(10):1930–1935, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109006178.

Tan:2002:VVR

[TSZ02] Jieqing Tan, Baorui Song, and Gongqin Zhu. Vector valued rational interpolants over triangular grids. Computers and Mathematics with Applications, 44(10– 11):1357–1367, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002626.

Tadj:2003:DPQ

[TT03] L. Tadj and C. Tadj. The departure process of a quorum queueing system. Computers and Mathematics with Applications, 45 (4–5):655–664, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000257.

Tian:2009:ITT

[TT09] Yongge Tian and Yoshio Takane. The inverse of any two-by-two nonsingular partitioned matrix and three matrix inverse completion problems. Computers and Mathematics with Applications, 57(8):1294– 1304, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000546.

Tung:2005:FMF

[Tun05] Michael M. Tung. FORM matters: Fast symbolic computation under UNIX. Computers and Mathematics with Applications, 49(7–8):1127–1137, May/ June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001276.

Turner:2004:MCC

[Tur04] J. C. Turner, Jr. Modelling control of crystal growth processes. Computers and Mathematics with Applications, 48(7–8):1231–1243, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003578.

Turkyilmazoglu:2007:AIS

[Tur07] M. Turkyilmazoğlu. An asymptotic investigation into the stabilization effect of suction in the rotatingdisk boundary layer flow. Computers and Mathematics with Applications, 53 (5):750–759, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107000284.

Tornabene:2007:VAS

[TV07] Francesco Tornabene and Erasmo Viola. Vibration analysis of spherical structural elements using the GDQ method. Computers and Mathematics with Applications, 53(10):1538– 1560, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001228.

Tu:2000:CLR

[TWS00] Shih-Tong Tu, Tsu-Chen Wu, and H. M. Srivastava. Commutativity of the Leibniz rules in fractional calculus. Computers and Mathematics with Applications, 40(2–3):303–312, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001620.

Tian:2004:FOC

[TX04] Chuanjun Tian and Shengli Xie. Further oscillation criteria for delay partial difference equations. Computers and Mathematics with Applications, 47(12):1905– 1914, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840264.

Tang:2002:FEM

[TXC02] Yi-Fa Tang, Ai-Guo Xiao, and Jing-Bo Chen. Is the formal energy of the mid-point rule convergent? Computers and Mathematics with Applications, 43(8–9):1171–1181, April/ May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800219.

Tokuno:2003:RBS

[TY03] K. Tokuno and S. Yamada.

Relationship between software availability measurement and the number of restorations with imperfect debugging. Computers and Mathematics with Applications, 46(7):1155–1163, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103901313.

Tang:2000:ONN

[TYP00] X. H. Tang, J. S. Yu, and D. H. Peng. Oscillation and nonoscillation of neutral difference equations with positive and negative coefficients. Computers and Mathematics with Applications, 39(7–8):169– 181, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000730.

Thomas:2003:SET

[TYS03] T. G. Thomas, Y. F. Yao, and N. D. Sandham. Structure and energetics of a turbulent trailing edge flow. Computers and Mathematics with Applications, 46(4):671–680, August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103900241.

Tsaur:2005:FRA

[TYW05] Ruey-Chyn Tsaur, Jia-Chi O. Yang, and Hsiao-Fan Wang. Fuzzy relation analysis in fuzzy time series model. Computers and Mathematics with Applications, 49(4):539–548, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000489.

Tang:2001:NOC

[TZ01] X. H. Tang and R. Y.

Zhang. New oscillation criteria for delay difference equations. Computers and Mathematics with Applications, 42(10– 11):1319–1330, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002437.

Tang:2004:BTB

[TZ04a] Yilei Tang and Weinian Zhang. Bogdanov–Takens bifurcation of a polynomialdifferential system in biochemical reaction. Computers and Mathematics with Applications, 48(5–6): 869–883, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104003207.

Tian:2004:EAS

[TZ04b] Chuanjun Tian and Jihong Zhang. Exponential asymptotic stability of delay partial difference equations. Computers and Mathematics with Applications, 47(2–3):345–352, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900296.

Thomsen:2008:DDA

[TZ08] Per Grove Thomsen and Zahari Zlatev. Development of a data assimilation algorithm. Computers and Mathematics with Applications, 55(10):2381– 2393, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007298.

Tang:2003:VWF

[TZS03] Xiaowo Tang, Zongfang Zhou, and Yong Shi. The variable weighted functions of combined forecasting. Computers and Mathematics with Applications, 45 (4–5):723–730, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000324.

Ugur:2006:SLO

[UA06] Ö. Ugur and M. U. Akhmet.

The Sturm–Liouville operator on the space of functions with discontinuity conditions. Computers and Mathematics with Applications, 51(6–7):889–896, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000472.

Ubhaya:2004:ADA

[Ubh04] V. A. Ubhaya. An algorithm for discrete approximation by quasi-convex functions on R^{m}. Computers and Mathematics with Applications, 47(10–11):1707–1712, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839841.

Unal:2008:DAL

[ÜÇT08] Mehmet Ünal, Devrim Çakmak, and Aydin Tiryaki. A discrete analogue of Lyapunov-type inequalities for nonlinear systems. Computers and Mathematics with Applications, 55(11):2631–2642, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007626.

Fukhar-ud-din:2007:ACF

[udK07] Hafiz Fukhar ud din and Abdul Rahim Khan. Approximating common fixed points of asymptotically nonexpansive maps in uniformly convex Banach spaces. Computers and Mathematics with Applications, 53(9):1349– 1360, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700096X.

Uddin:2009:NSC

[UHuI09] Marjan Uddin, Sirajul Haq, and Siraj ul Islam. Numerical solution of complex modified Korteweg–de Vries equation by mesh-free collocation method. Computers and Mathematics with Applications, 58(3):566–578, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002405.

Ujevic:2003:NBF

[Uje03] N. Ujević. New bounds for the first inequality of Ostrowski-Grüss type and applications. Computers and Mathematics with Applications, 46(2–3):421–427, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900356.

Ujevic:2004:EIQ

[Uje04a] N. Ujević. Error inequalities for a quadrature formula and applications. Computers and Mathematics with Applications, 48(10– 11):1531–1540, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004079.

Ujevic:2004:SIS

[Uje04b] N. Ujević. Sharp inequalities of Simpson type and Ostrowski type. Computers and Mathematics with Applications, 48(1–2):145– 151, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840057.

Ujevic:2007:NEB

[Uje07] Nenad Ujević. New error bounds for the Simpson’s quadrature rule and applications. Computers and Mathematics with Applications, 53(1):64–72, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700034X.

Ugurlu:2008:SCH

[UK08] Yavuz Ugurlu and Dogan Kaya. Solutions of the Cahn–Hilliard equation. Computers and Mathematics with Applications, 56(12):3038–3045, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004446.

Uster:2000:CWA

[ÜL00] H. Üster and R. F. Love. The convergence of the Weiszfeld algorithm. Computers and Mathematics with Applications, 40(4–5): 443–451, August/September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122100001723.

Uster:2002:GRB

[ÜL02] H. Üster and R. F. Love. A generalization of the rectangular bounding method for continuous location models. Computers and Mathematics with Applications, 44(1–2):181–191, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001396.

Ungar:2000:HTE

[Ung00]  A. A. Ungar.  Hyper-

bolic trigonometry in the Einstein relativistic velocity model of hyperbolic geometry. Computers and Mathematics with Applications, 40(2–3):313–332, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001632.

Ungar:2001:HTA

[Ung01] A. A. Ungar. Hyperbolic trigonometry and its application in the Poincaré ball model of hyperbolic geometry. Computers and Mathematics with Applications, 41(1–2):135–147, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101850124.

Ungar:2003:SMD

[Ung03] A. A. Ungar. Seeing the Möbius disc-transformation group like never before. Computers and Mathematics with Applications, 45 (4–5):805–822, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000427.

Ungar:2005:ESR

[Ung05]  A. A. Ungar.  Ein-

stein’s special relativity: Unleashing the power of its hyperbolic geometry. Computers and Mathematics with Applications, 49 (2–3):187–221, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000271.

Ungar:2007:EVA

[Ung07] Abraham A. Ungar. Einstein’s velocity addition law and its hyperbolic geometry. Computers and Mathematics with Applications, 53(8):1228–1250, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700082X.

Valko:2004:CSA

[VA04] P. P. Valkó and J. Abate. Comparison of sequence accelerators for the Gaver method of numerical Laplace transform inversion. Computers and Mathematics with Applications, 48(3–4): 629–636, August 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104840823.

VandeVyver:2007:ENT

[Van07] Hans Van de Vyver. An explicit Numerov-type method for second-order differential equations with oscillating solutions. Computers and Mathematics with Applications, 53(9):1339– 1348, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000958.

Vigo-Aguiar:2003:MRP

[VAR03] J. Vigo-Aguiar and D. L.

Richardson. Modification of the Richardson–Panovsky methods for precise integration of satellite orbits. Computers and Mathematics with Applications, 45(1–3):25–36, January/ February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103800044.

Vespucci:2001:IDC

[VB01] M. T. Vespucci and C. G.

Broyden. Implementation of different computational variations of biconjugate residual methods.

Computers and Mathematics with Applications, 42 (8–9):1239–1253, October/ November 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210100236X.

VanDeun:2007:COR

[VBG07] J. Van Deun, A. Bultheel, and P. González Vera. Computing orthogonal rational functions with poles near the boundary. Computers and Mathematics with Applications, 53(9):1421– 1428, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001095.

Vurro:2003:NTI

[VC03] M. Vurro and L. Castellano. Numerical treatments of the interface discontinuity in solid-water mass transfer problems. Computers and Mathematics with Applications, 45 (4–5):785–788, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000403.

Villegas:2008:RCU

[VDMK08] Rossmary Villegas, Oliver Dorn, Miguel Moscoso, and Manuel Kindelan. Reservoir characterization using stochastic initializations and the level set method. Computers and Mathematics with Applications, 56(3):697–708, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000680.

Verma:2001:NCI

[Ver01a] R. U. Verma. A new class of iterative algorithms for approximation-solvability of nonlinear variational inequalities. Computers and Mathematics with Applications, 41(3–4):505–512, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002923.

Verma:2001:PMA

[Ver01b]  R. U. Verma.  Projec-

tion methods, algorithms, and a new system of nonlinear variational inequalities. Computers and Mathematics with Applications, 41(7–8):1025–1031, April 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003369.

Verma:2003:NIV

[Ver03]  R. U. Verma.  Nonlin-

ear implicit variational inequalities involving partially relaxed pseudomonotone mappings. Computers and Mathematics with Applications, 46(10– 11):1703–1709, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902045.

Verma:2008:RPP

[Ver08a] Ram U. Verma. The overrelaxed proximal point algorithm based on H-maximal monotonicity design and applications. Computers and Mathematics with Applications, 55(11):2673– 2679, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007663.

See corrigendum [Ver12].

Verma:2008:GEB

[Ver08b] Rom U. Verma. Generalized Eckstein–Bertsekas proximal point algorithm based on A-maximal monotonicity design. Computers and Mathematics with Applications, 56(7):1861–1875, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002666.

Verma:2009:NRP

[Ver09] Ram U. Verma. A new relaxed proximal point procedure and applications to nonlinear variational inclusions. Computers and Mathematics with Applications, 58(8):1631–1635, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004301.

Verma:2012:CRP

[Ver12] Ram U. Verma. Corrigendum to “The over-relaxed proximal point algorithm based on H-maximal monotonicity design and applications” [Comput. Math. Appl. 55 (2008) 2673– 2679]. Computers and Mathematics with Applications, 64(9):2961–2963, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122112005652.

Viswamurthy:2008:PSH

[VG08] S. R. Viswamurthy and R. Ganguli. Performance sensitivity of helicopter global and local optimal harmonic vibration controller. Computers and Mathematics with Applications, 56(10):2468–2480, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003830.

Vulanovic:2006:PEE

[VH06] R. Vulanović and G. Hovhannisyan. A posteriori error estimates for one-dimensional convectiondiffussion problems. Computers and Mathematics with Applications, 51(6–7): 915–926, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106000265.

Villanueva:2004:DVS

[VJ04] R. J. Villanueva and L. Jódar. Discrete vector Sturm– Liouville problems with non-self-adjoint boundary conditions: Eigenstructure, orthogonality, and eigenfunctions expansion. Computers and Mathematics with Applications, 47 (6–7):1123–1133, March/ April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900922.

Verma:2009:LBM

[VM09] Nishith Verma and Dieter Mewes. Lattice Boltzmann methods for simulation of micro and macrotransport in a packed bed of porous adsorbents under non-isothermal condition. Computers and Mathematics with Applications, 58(5):1003–1014, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001059.

Vieru:2008:NES

[VNFF08] D. Vieru, M. Nazar, Corina Fetecau, and C. Fetecau. New exact solutions corresponding to the first problem of Stokes for Oldroyd-b fluids. Computers and Mathematics with Applications, 55(8):1644– 1652, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006025.

VanJoolen:2004:SDW

[VNG04] V. Van Joolen, B. Neta, and D. Givoli. A stratified dispersive wave model with high-order nonreflecting boundary conditions. Computers and Mathematics with Applications, 48(7–8):1167–1180, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003530.

Valarmathi:2002:ANM

[VR02] S. Valarmathi and N. Ramanujam. An asymptotic numerical method for singularly perturbed thirdorder ordinary differential equations of convection– diffusion type. Computers and Mathematics with Applications, 44(5–6):693–710, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001839.

VanLeemput:2009:SIL

[VRJ09] Pieter Van Leemput, Martin Rheinländer, and Michael Junk. Smooth initialization of lattice Boltzmann schemes. Computers and Mathematics with Applications, 58(5):867–882, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000923.

Vieru:2008:EBF

[VSKF08] D. Vieru, I. Siddique, M. Kamran, and C. Fetecau. Energetic balance for the flow of a second-grade fluid due to a plate subject to a shear stress. Computers and Mathematics with Applications, 56(4):1128– 1137, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001144.

Vrankar:2004:CRB

[VTR04] L. Vrankar, G. Turk, and F. Runovc. Combining the radial basis function Eulerian and Lagrangian schemes with geostatistics for modeling of radionuclide migration through the geosphere. Computers and Mathematics with Applications, 48(10– 11):1517–1529, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004067.

Varga:2008:LLT

[VV08] K. Pásztor Varga and M. Várterész. Languages of logic and their applications. Computers and Mathematics with Applications, 55(8):1660–1669, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006050.

Wong:2001:FRF

[WA01] P. J. Y. Wong and R. P. Agarwal. Further results on fixed-sign solutions for a system of higherorder difference equations. Computers and Mathematics with Applications, 42(3–5):497–514, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001730.

Wei:2008:ESN

[WA08] Li Wei and Ravi P. Agarwal. Existence of solutions to nonlinear Neumann boundary value problems with generalized p-Laplacian operator. Computers and Mathematics with Applications, 56(2):530–541, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108000825.

Wachspress:2000:EEF

[Wac00] E. L. Wachspress. Evaluating elliptic functions and their inverses. Computers and Mathematics with Applications, 39(3–4):131–136, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003399.

Wachspress:2008:TLE

[Wac08] Eugene L. Wachspress. Trail to a Lyapunov equation solver. Computers and Mathematics with Applications, 55(8):1653–1659, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006037.

Wang:2000:SBF

[Wan00a] Guo-Jun Wang. Separable Boolean functions and generalized Fibonacci sequences. Computers and Mathematics with Applications, 39(3–4):205–216, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003466.

Wang:2000:TGL

[Wan00b] Guo-Jun Wang. Theory of granular lattices and its applications. Computers and Mathematics with Applications, 39(12): 1–9, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001243.

Wang:2000:SSW

[Wan00c] Qiao Wang. The shiftability of some wavelet bases. Computers and Mathematics with Applications, 40 (8–9):957–964, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100850063.

Wang:2000:OTF

[Wan00d] Qiru Wang. Oscillation theorems for first-order nonlinear neutral functional differential equations. Computers and Mathematics with Applications, 39(5–6):19– 28, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000420.

Wang:2000:AMI

[Wan00e]  Yuan-Ming Wang.  Ac-

celerated monotone iterative methods for a boundary value problem of second-order discrete equations. Computers and Mathematics with Applications, 39(3–4):85–94, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003351.

Wang:2001:GSD

[Wan01]  Wendi Wang.  Global

stability of discrete competition model. Computers and Mathematics with Applications, 42(6– 7):773–782, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001973.

Wan:2002:FIF

[Wan02a]  Zhong Wan.  Further

investigation on feasibility of mathematical programs with equilibrium constraints. Computers and Mathematics with Applications, 44(1–2):7– 11, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001268.

Wang:2002:BRS

[Wan02b] Mingxin Wang. Blow-up rate for a semilinear reaction diffusion system. Computers and Mathematics with Applications, 44(5–6):573–585, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001724.

Wang:2002:SCD

[Wan02c]  Song Wang.  Solving

convection-dominated anisotropic diffusion equations by an exponentially fitted finite volume method. Computers and Mathematics with Applications, 44(8– 9):1249–1265, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002304.

Wang:2002:CIP

[Wan02d] Weichung Wang. The convergence of an interiorpoint method using modified search directions in final iterations. Computers and Mathematics with Applications, 44(3–4):347– 356, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001530.

Wang:2002:NMI

[Wan02e] Xiaoqun Wang. A new measure of irregularity of distribution and quasi-Monte Carlo methods for global optimization. Computers and Mathematics with Applications, 43(6–7):657–669, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210100311X.

Wang:2003:NMS

[Wan03a]  Yuan-Ming Wang.  Nu-

merov’s method for strongly nonlinear two-point boundary value problems. Computers and Mathematics with Applications, 45(4–5): 759–763, February/March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103000361.

Wang:2003:PMI

[Wan03b] Yuan-Ming Wang. Parallel monotone iterative relaxation methods for a class of two-dimensional discrete boundary value problems. Computers and Mathematics with Applications, 45(6–9):887–903, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000646.

Wang:2004:MGQ

[Wan04a] Jinting Wang. An M/G/1 queue with second optional service and server breakdowns. Computers and Mathematics with Applications, 47(10–11):1713–1723, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104839853.

Wang:2004:OCS

[Wan04b]  Peiguang Wang.  Oscil-

lation criteria for secondorder neutral equations with distributed deviating arguments. Computers and Mathematics with Applications, 47(12):1935– 1946, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210484029X.

Wang:2005:BCS

[Wan05a] Qing-Wen Wang. Bisymmetric and centrosymmetric solutions to systems of real quaternion matrix equations. Computers and Mathematics with Applications, 49(5–6):641–650, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105000799.

Wang:2005:GSS

[Wan05b]  Qing-Wen Wang.  The

general solution to a system of real quaternion matrix equations. Computers and Mathematics with Applications, 49(5–6):665–675, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000817.

Wang:2006:SPC

[Wan06a] Mingxin Wang. Stationary patterns caused by crossdiffusion for a three-species prey–predator model. Computers and Mathematics with Applications, 52(5): 707–720, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106002501.

Wang:2006:ABSb

[Wan06b] Xiaoping Wang. Asymptotic behavior of solutions for neutral difference equations. Computers and Mathematics with Applications, 52(10– 11):1595–1602, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003427.

Wang:2006:ABSa

[Wan06c] Yuan-Ming Wang. Asymptotic behavior of solutions for a cooperation-diffusion model with a saturating interaction. Computers and Mathematics with Applications, 52(3–4):339–350, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210600201X.

Wang:2007:EIM

[Wan07a] Lin Wang. Explicit iteration method for common fixed points of a finite family of nonself asymptotically nonexpansive mappings. Computers and Mathematics with Applications, 53(7):1012– 1019, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000569.

Wang:2007:ESR

[Wan07b]  Xinmao Wang.  Effect

of small rank modification on the condition number of a matrix. Computers and Mathematics with Applications, 54(6):819–825, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002799.

Wang:2008:PSN

[Wan08a] Da-Bin Wang. Positive solutions for nonlinear firstorder periodic boundary value problems of impulsive dynamic equations on time scales. Computers and Mathematics with Applications, 56(6):1496–1504, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001764.

Wang:2008:SMS

[Wan08b] Ji-Bo Wang. Single-machine scheduling with general learning functions. Computers and Mathematics with Applications, 56(8):1941– 1947, October 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210800271X. See

note [YH11].

Wang:2009:SMC

[Wan09a] Deng-Shan Wang. A systematic method to construct Hirota’s transformations of continuous soliton equations and its applications. Computers and Mathematics with Applications, 58(1):146–153, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002120.

Wang:2009:SMSb

[Wan09b] Ji-Bo Wang. Single machine scheduling with decreasing linear deterioration under precedence constraints. Computers and Mathematics with Applications, 58(1): 95–103, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002892.

Wang:2009:GMB

[Wan09c] Mingjin Wang. Generalizations of Milne’s U(n + 1)qbinomial theorems. Computers and Mathematics with Applications, 58(1):80– 87, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900248X.

Wang:2009:VAN

[Wan09d] Shu-Qiang Wang. A variational approach to nonlinear two-point boundary value problems. Computers and Mathematics with Applications, 58(11–12):2452–2455, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001850.

Wang:2009:ABS

[Wan09e] Yuan-Ming Wang. Asymptotic behavior of solutions for a Lotka–Volterra mutualism reaction-diffusion system with time delays. Computers and Mathematics with Applications, 58(3):597–604, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002582.

Wazwaz:2000:ASB

[Waz00] A.-M. Wazwaz. Approximate solutions to boundary value problems of higher order by the modified decomposition method. Computers and Mathematics with Applications, 40 (6–7):679–691, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001875.

Wazwaz:2001:RAO

[Waz01] A. M. Wazwaz. A reliable algorithm for obtaining positive solutions for nonlinear boundary value problems. Computers and Mathematics with Applications, 41(10–11):1237–1244, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101000943.

Wazwaz:2004:VGK

[Waz04]  A.-M Wazwaz.  Vari-

ants of the generalized KdV equation with compact and noncompact structures. Computers and Mathematics with Applications, 47(4–5):583–591, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900478.

Wazwaz:2005:ESD

[Waz05a] A. M. Wazwaz. Exact solutions to the double sinh-Gordon equation by the tanh method and a variable separated ODE method. Computers and Mathematics with Applications, 50 (10–12):1685–1696, November/December 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105004256.

Wazwaz:2005:NVI

[Waz05b] A.-M. Wazwaz. Nonlinear variants of the improved Boussinesq equation with compact and noncompact structures. Computers and Mathematics with Applications, 49(4):565–574, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000507.

Wazwaz:2005:TSC

[Waz05c]  A.-M. Wazwaz.  The

tanh and the sine-cosine methods for the complex modified K dV and the generalized K dV equations. Computers and Mathematics with Applications, 49(7–8):1101–1112, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001252.

Wazwaz:2005:VTD

[Waz05d] A.-M. Wazwaz. Variants of the two-dimensional Boussinesq equation with compactons, solitons, and periodic solutions. Computers and Mathematics with Applications, 49(2–3): 295–301, January/February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105000349.

Wazwaz:2007:VIMc

[Waz07a] Abdul-Majid Wazwaz. The variational iteration method: a powerful scheme for handling linear and nonlinear diffusion equations. Computers and Mathematics with Applications, 54(7–8): 933–939, October 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107002556.

Wazwaz:2007:VIMb

[Waz07b] Abdul-Majid Wazwaz. The variational iteration method: a reliable analytic tool for solving linear and nonlinear wave equations. Computers and Mathematics with Applications, 54(7–8):926– 932, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002544.

Wazwaz:2007:VIMa

[Waz07c] Abdul-Majid Wazwaz. The variational iteration method for solving linear and nonlinear systems of PDEs. Computers and Mathematics with Applications, 54(7– 8):895–902, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107003033.

Weisz:2001:RAB

[WB01] J. Weisz and R. Bodnár. A refined “angle between normals” criterion for scattered data interpolation. Computers and Mathematics with Applications, 41(3–4):531– 534, February 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100002959.

Wang:2009:LBS

[WB09] Dinan Wang and Jörg Bernsdorf. Lattice Boltzmann simulation of steady non-Newtonian blood flow in a 3D generic stenosis case. Computers and Mathematics with Applications, 58(5):1030–1034, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001096.

Weiner:2008:ETS

[WBSP08] Rüdiger Weiner, Katja Biermann, Bernhard A. Schmitt, and Helmut Podhaisky. Explicit two-step peer methods. Computers and Mathematics with Applications, 55(4):609–619, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107004956.

Wang:2004:GPI

[WC04] Yong Wang and Wu Congxin. Global Poincaré inequalities for Green’s operator applied to the solutions of the nonhomogeneous A-harmonic equation. Computers and Mathematics with Applications, 47(10–11):1545–1554, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104839713.

Wang:2002:UPI

[WCL02] Jinting Wang, Jinhua Cao, and Bin Liu. Unreliable production-inventory model with a two-phase Erlang demand arrival process. Computers and Mathematics with Applications, 43(1–2):1–13, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002668.

Wong:2008:EPSa

[WCW08] F. H. Wong, T. G. Chen, and S. P. Wang. Existence of positive solutions for various boundary value problems. Computers and Mathematics with Applications, 56(4):953–958, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000977.

Wu:2005:OSB

[WD05] Shujin Wu and Yongrui Duan. Oscillation, stability, and boundedness of second-order differential systems with random impulses. Computers and Mathematics with Applications, 49(9–10):1375– 1386, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001562.

Wu:2007:GWC

[WD07a] Shanhe Wu and Lokenath Debnath. Generalization of the Wolstenholme cyclic inequality and its application. Computers and Mathematics with Applications, 53(1):104–114, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000600.

Wu:2007:ICS

[WD07b] Shanhe Wu and Lokenath Debnath. Inequalities for convex sequences and their applications. Computers and Mathematics with Applications, 54(4):525–534, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002350.

Wang:2008:ESG

[WD08] Dongqing Wang and Feng Ding. Extended stochastic gradient identification algorithms for Hammerstein– Wiener ARMAX systems. Computers and Mathematics with Applications, 56 (12):3157–3164, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210800432X.

Wang:2004:ACM

[WDLP04] Z. K. Wang, G. S. Djambazov, C.-H. Lai, and K. A. Pericleous. An acoustic correction method for extracting sound signals. Computers and Mathematics with Applications, 47(1):57–69, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900053.

Wang:2008:NIS

[WDLP08] Z.-K. Wang, G. S. Djambazov, C.-H. Lai, and K. A. Pericleous. Numerical investigation of a source extraction technique based on an acoustic correction method. Computers and Mathematics with Applications, 55(3):441–458, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107004129.

Wei:2000:DME

[Wei00] Yuchuan Wei. Dirichlet multiplication and easilygenerated function analysis. Computers and Mathematics with Applications, 39(9–10):173–199, May 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000095X.

Wei:2005:PNS

[Wei05]  Gengping Wei.  The

persistence of nonoscillatory solutions of difference equations under impulsive perturbations. Computers and Mathematics with Applications, 50(10– 12):1579–1586, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004098.

Weng:2000:EGS

[Wen00] Peixuan Weng. Existence and global stability of positive periodic solution of periodic integrodifferential systems with feedback controls. Computers and Mathematics with Applications, 40(6–7):747–759, September/October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001930.

Weng:2002:GAP

[Wen02] Pei-Xuan Weng. Global attractivity of periodic solution in a model of hematopoiesis. Computers and Mathematics with Applications, 44(8– 9):1019–1030, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002110.

Wen:2008:NKT

[Wen08] Kai Ting Wen. A new KKM theorem in L-convex spaces and some applications. Computers and Mathematics with Applications, 56(11):2781–2785, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004628.

Wu:2001:NHO

[WF01] Xinyuan Wu and Dongsheng Fu. New high-order convergence iteration methods without employing derivatives for solving nonlinear equations. Computers and Mathematics with Applications, 41(3–4):489–495, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210000290X.

Wang:2005:UFT

[WF05] G.-J. Wang and Li Fu.

Unified forms of triple I method. Computers and Mathematics with Applications, 49(5–6):923–932, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500101X.

Wang:2007:NCP

[WF07] Lin Wang and Xilin Fu.

A new comparison principle for impulsive differential systems with variable impulsive perturbations and stability theory. Computers and Mathematics with Applications, 54(5):730–736, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002659.

Wang:2004:PIC

[WFC04a] Quanfang Wang, Dexing Feng, and Daizhan Cheng. Parameter identification for a class of abstract nonlinear parabolic distributed parameter systems. Computers and Mathematics with Applications, 48(12):1847–1861, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104004444.

Waters:2004:MGP

[WFC04b] L. K. Waters, G. J. Fix, and C. L. Cox. The method of glowinski and pironneaufor the unsteady Stokes problem. Computers and Mathematics with Applications, 48(7–8):1191–1211, October 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104003554.

Wang:2005:GPA

[WFLC05] Yong Wang, Shu-Cherng Fang, J. E. Lavery, and Hao Cheng. A geometric programming approach for bivariate cubic L_{1} splines. Computers and Mathematics with Applications, 49(4):481–514, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000453.

Wihler:2003:ECD

[WFS03] T. P. Wihler, P. Frauenfelder, and C. Schwab. Exponential convergence of the hp -DGFEM for diffusion problems. Computers and Mathematics with Applications, 46(1):183– 205, July 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900885.

Wei:2009:RQP

[WFT^{+}09] Zhe Wei, Yixiong Feng, Jianrong Tan, Junhao Wu, Dandan Yang, and Jinlong Wang. Research on quality performance conceptual design based on SPEA2+. Computers and Mathematics with Applications, 57(11–12):1943– 1948, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005269.

Wei:2009:PPT

[WFTL09] Wei Wei, Yixiong Feng, Jianrong Tan, and Zhongkai Li. Product platform two-stage quality optimization design based on multiobjective genetic algorithm. Computers and Mathematics with Applications, 57(11–12):1929– 1937, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005294.

Wang:2009:PSP

[WFY09] Weibing Wang, Xiangling Fu, and Xuxin Yang. Positive solutions of periodic boundary value problems for impulsive differential equations. Computers and Mathematics with Applications, 58(8):1623–1630, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004465.

Wang:2006:PSD

[WG06] Youyu Wang and Weigao Ge. Periodic solutions for Duffing equations with a p-Laplacian-like operator. Computers and Mathematics with Applications, 52(6– 7):1079–1088, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002793.

Wang:2007:EST

[WG07a] Youyu Wang and Weigao Ge. Existence of solutions for a third order differential equation with integral boundary conditions. Computers and Mathematics with Applications, 53(1):144–154, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000211.

Wang:2007:ETP

[WG07b] Youyu Wang and Weigao Ge. Existence of triple positive solutions for multipoint boundary value problems with a one dimensional p-Laplacian. Computers and Mathematics with Applications, 54(6):793–807, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002751.

Wang:2008:TPS

[WG08a] Da-Bin Wang and Wen Guan. Three positive solutions of boundary value problems for p-Laplacian difference equations. Computers and Mathematics with Applications, 55(9): 1943–1949, May 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107006621.

Wu:2008:ANS

[WG08b] Shifeng Wu and Siqing Gan. Analytical and numerical stability of neutral delay integro-differential equations and neutral delay partial differential equations. Computers and Mathematics with Applications, 55(11):2426–2443, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006980.

Wang:2009:PCP

[WGF09] Bin Wang, Hai-Cheng Guo, and Dai-Ning Fang. A protocol for characterizing the performance of lightweight truss-core actuators. Computers and Mathematics with Applications, 58(11–12):2365–2372, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900176X.

Wu:2006:MSF

[WGZ06] Shujin Wu, Xiaolin Guo, and Yong Zhou. pmoment stability of functional differential equations with random impulses. Computers and Mathematics with Applications, 52(12):1683–1694, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003518.

Wu:2005:ESF

[WH05] S. J. Wu and D. Han. Exponential stability of functional differential systems with impulsive effect on random moments. Computers and Mathematics with Applications, 50(1–2):321– 328, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500252X.

Wu:2007:SSP

[WH07] Xu-Hong (Benn) Wu and Ji-Huan He. Solitary solutions, periodic solutions and compacton-like solutions using the Exp-function method. Computers and Mathematics with Applications, 54(7–8):966–986, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002581.

Wang:2008:MCG

[WH08a] Dongdong Wang and Hongbo Hua. Minimality considerations for graph energy over a class of graphs. Computers and Mathematics with Applications, 56(12):3181– 3187, December 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108004392.

Wu:2008:NCL

[WH08b] Yuhai Wu and Maoan Han. New configurations of 24 limit cycles in a quintic system. Computers and Mathematics with Applications, 55(9):2064– 2075, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006864.

Wu:2009:SBL

[WH09] Chi-Ye Wu and Ting-Zhu Huang. Stability of block LU factorization for block tridiagonal matrices. Computers and Mathematics with Applications, 57(3):339–347, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006792.

Wu:2009:SME

[WHD09] Ai-Guo Wu, Junqiang Hu, and Guang-Ren Duan. Solutions to the matrix equation AX − EXF = BY. Computers and Mathematics with Applications, 58(10):1891–1900, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006142.

Whiteman:2000:SBI

[Whi00] A. L. Whiteman. Some balanced incomplete block designs. Computers and Mathematics with Applications, 39(11):117–119, June 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001164.

Wong:2002:MMM

[WHL02] S. M. Wong, Y. C. Hon, and T. S. Li. A meshless multilayer model for a coastal system by radial basis functions. Computers and Mathematics with Applications, 43(3–5):585–605, February/ March 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101003066.

Wang:2009:CEC

[WHL^{+}09] Jianfeng Wang, Qiongxiang Huang, Yongzhi Liu, Ruying Liu, and Chengfu Ye. The cospectral equivalence classes of graphs having an isolated vertex. Computers and Mathematics with Applications, 57(10):1638– 1644, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002132.

Wang:2008:TFL

[WHS08] Guo-Jun Wang, Xiao-Jing Hui, and Jian-She Song. The R_{0}-type fuzzy logic metric space and an algorithm for solving fuzzy modus ponens. Computers and Mathematics with Applications, 55(9):1974– 1987, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006657.

Wen:2008:FLA

[WI08] Meilin Wen and Kakuzo Iwamura. Facility locationallocation problem in random fuzzy environment:

Using (α, β)-cost minimization model under the Hurewicz criterion. Computers and Mathematics with Applications, 55(4):704–713, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005032.

Winkler:2003:SAN

[Win03] J. R. Winkler. A statistical analysis of the numerical condition of multiple roots of polynomials. Computers and Mathematics with Applications, 45(1– 3):9–24, January/February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800032.

Wineman:2007:NVM

[Win07]  Alan Wineman.  Non-

linear viscoelastic membranes. Computers and Mathematics with Applications, 53(2):168–181, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000697.

Wang:2008:FOC

[WJA08] Yuan-Ming Wang, Hai-Yun Jiang, and Ravi P.

Agarwal. A fourth-order compact finite difference method for higher-order Lidstone boundary value problems. Computers and Mathematics with Applications, 56(2):499–521, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000321.

Wang:2009:SSM

[WJS09] Zhi-Gang Wang, Yue-Ping Jiang, and H. M. Srivastava. Some subclasses of meromorphically multivalent functions associated with the generalized hypergeometric function. Computers and Mathematics with Applications, 57(4): 571–586, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108006676.

Wang:2008:SRA

[WJW08] Weiping Wang, Cangzhi Jia, and Tianming Wang. Some results on the Apostol– Bernoulli and Apostol– Euler polynomials. Computers and Mathematics with Applications, 55(6):1322– 1332, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006013.

Wan:2004:NET

[WJX04] Aying Wan, Daqing Jiang, and Xiaojie Xu. A new existence theory for positive periodic solutions to functional differential equations. Computers and Mathematics with Applications, 47(8–9):1257–1262, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901204.

Weniger:2003:EMI

[WK03] E. J. Weniger and B. Kirtman. Extrapolation methods for improving the convergence of oligomer calculations to the infinite chain limit of quasi-one-dimensional stereoregular polymers. Computers and Mathematics with Applications, 45(1–3):189–215, January/February 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800147.

Wang:2004:FAD

[WK04] Hsiao-Fan Wang and Ching-Yi Kuo. Factor analysis in data mining. Computers and Mathematics with Applications, 48(10– 11):1765–1778, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104004262.

Wu:2005:HAL

[WK05] Xionghua Wu and Wenbin Kong. A highly accurate linearized method for free boundary problems. Computers and Mathematics with Applications, 50 (8–9):1241–1250, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003779.

Wang:2008:AFK

[WK08] Yu-Jie Wang and Chao-Shun Kao. An application of a fuzzy knowledge system for air cargo overbooking under uncertain capacity. Computers and Mathematics with Applications, 56(10):2666–2675, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004069.

Wertz:2006:RMS

[WKL06] J. Wertz, E. J. Kansa, and L. Ling. The role of the multiquadric shape parameters in solving elliptic partial differential equations. Computers and Mathematics with Applications, 51(8):1335– 1348, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000794.

Wang:2008:AAL

[WKZ08] Jinhuan Wang, Linghua Kong, and Sining Zheng. Asymptotic analysis for a localized nonlinear diffusion equation. Computers and Mathematics with Applications, 56(9):2294–2304, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003647.

Wang:2004:NSS

[WL04] Heyuan Wang and Kaitai Li. Numerical simulation of spherical Couette flow. Computers and Mathematics with Applications, 48(1–2):109–116, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840021.

Wang:2005:HFV

[WL05] Meiqing Wang and Choi-Hong Lai. A hybrid fractal video compression method. Computers and Mathematics with Applications, 50(3–4):611–621, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002683.

Wang:2006:RPE

[WL06] Yijie Wang and Sijun Li. Research and performance evaluation of data replication technology in distributed storage systems. Computers and Mathematics with Applications, 51(11):1625– 1632, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001313.

Wang:2007:MSS

[WL07a] Ren-Hong Wang and Feng-Gong Lang. Multivariate spline space over crosscut partition. Computers and Mathematics with Applications, 54(3):415–426, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002246.

Wang:2007:GTF

[WL07b] Yu-Jie Wang and Hsuan-Shih Lee. Generalizing TOPSIS for fuzzy multiplecriteria group decisionmaking. Computers and Mathematics with Applications, 53(11):1762–1772, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001514.

Wang:2008:GOS

[WL08a] Yanjun Wang and Ying Lan. Global optimization for special reverse convex programming. Computers and Mathematics with Applications, 55(6):1154– 1163, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005822.

Wang:2008:RMR

[WL08b] Yu-Jie Wang and Hsuan-Shih Lee. The revised method of ranking fuzzy numbers with an area between the centroid and original points. Computers and Mathematics with Applications, 55(9):2033– 2042, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006840.

Wu:2008:ESP

[WL08c] Huaiqin Wu and Yingwei Li. Existence and stability of periodic solution for BAM neural networks with discontinuous neuron activations. Computers and Mathematics with Applications, 56(8):1981–1993, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002769.

Wang:2009:ARF

[WL09a] Ying-Ming Wang and Ying Luo. Area ranking of fuzzy numbers based on positive and negative ideal points. Computers and Mathematics with Applications, 58(9):1769–1779, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109005896.

Wu:2009:PIC

[WL09b] Yusen Wu and Peiluan Li. Pseudo-isochronicity in a class of septic differential systems. Computers and Mathematics with Applications, 58(10):2069–2080, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005926.

Wu:2009:NSP

[WLD09] Kaining Wu, Chun Lu, and Xiaohua Ding. The numerical simulation of periodic solutions for a neural network. Computers and Mathematics with Applications, 58(11–12):2339–2346, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001655.

Wang:2006:APH

[WLL06] Meiqing Wang, Rong Liu, and Choi-Hong Lai. Adaptive partition and hybrid method in fractal video compression. Computers and Mathematics with Applications, 51(11):1715– 1726, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001386.

Walczak:2001:SEO

[WLS01] S. Walczak, U. Ledzewicz, and H. Schättler. Stability of elliptic optimal control problems. Computers and Mathematics with Applications, 41(10–11):1245–1256, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101000955.

Wang:2006:RSF

[WLS06] Pin-Yu Wang, Shy-Der Lin, and H. M. Srivastava. Remarks on a simple fractional-calculus approach to the solutions of the Bessel differential equation of general order and some of its applications. Computers and Mathematics with Applications, 51(1):105–114, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105004517.

Wu:2007:HPT

[WLT07] Chen-Feng Wu, Liang-Teh Lee, and Der-Fu Tao. An HMM prediction and throttling-based call admission control scheme for wireless multimedia networks. Computers and Mathematics with Applications, 54(3):364–378, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002209.

Wang:2007:USB

[WLW07] Zenggui Wang, Lishan Liu, and Yonghong Wu. The unique solution of boundary value problems for nonlinear second-order integraldifferential equations of mixed type in Banach spaces. Computers and Mathematics with Applications, 54(9–10):1293–1301, November 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003677.

Wen:2008:IIG

[WlW08a] Jiajin Wen and Wan lan Wang. Inequalities involving generalized interpolation polynomials. Computers and Mathematics with Applications, 56(4):1045– 1058, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001077.

Wu:2008:EPP

[WLW08b] Xinmin Wu, Jingwen Li, and Zhicheng Wang. Existence of positive periodic solutions for a generalized prey–predator model with harvesting term. Computers and Mathematics with Applications, 55(8):1895– 1905, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210700795X.

Wang:2009:FRR

[WLW09] Shuming Wang, Yan-Kui Liu, and Junzo Watada. Fuzzy random renewal process with queueing applications. Computers and Mathematics with Applications, 57(7):1232–1248, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000686.

Wang:2008:ISC

[WLWC08] Qing Wang, James Lam, Yimin Wei, and Tongwen Chen. Iterative solutions of coupled discrete Markovian jump Lyapunov equations. Computers and Mathematics with Applications, 55(4):843–850, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005159.

Wu:2007:NSC

[WLZ07] Xin-Min Wu, Jing-Wen Li, and Hou-Qing Zhou. A necessary and sufficient condition for the existence of positive periodic solutions of a model of hematopoiesis. Computers and Mathematics with Applications, 54(6):840–849, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002908.

Wang:2004:PPO

[WM04] Feng Wang and Zhien Ma.

Persistence and periodic orbits for an SIS model in a polluted environment. Computers and Mathematics with Applications, 47 (4–5):779–792, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900648.

Wang:2009:OLS

[WM09] Chih-Hsiung Wang and Fan-Chin Meng. Optimal lot size and offline inspection policy. Computers and Mathematics with Applications, 58(10):1921–1929, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006154.

Wang:2008:FSI

[WML08] De-Gang Wang, Yan-Ping Meng, and Hong-Xing Li. A fuzzy similarity inference method for fuzzy reasoning. Computers and Mathematics with Applications, 56(10):2445–2454, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003684.

Wang:2009:PBT

[WNS09] Jian Wang, Changyong Niu, and Ruimin Shen. Priority-based target coverage in directional sensor networks using a genetic algorithm. Computers and Mathematics with Applications, 57(11–12):1915– 1922, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005336.

Wong:2000:TPS

[Won00] P. J. Y. Wong. Triple positive solutions of conjugate boundary value problems II. Computers and Mathematics with Applications, 40(4–5):537–557, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001784.

Wong:2008:MFS

[Won08] Patricia J. Y. Wong. Multiple fixed-sign solutions for a system of higher order three-point boundaryvalue problems with deviating arguments. Computers and Mathematics with Applications, 55(3):516–534, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107004889.

Wang:2000:FAO

[WS00] Chih-Hsiung Wang and Shey-Heui Sheu. Fast approach to the optimal production/PM policy. Computers and Mathematics with Applications, 40(10– 11):1297–1314, November/ December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002406.

Wang:2004:FCF

[WS04a] Liya Wang and Yu-Ru Syau. Fuzzy Φ-convexity and fuzzy decision making. Computers and Mathematics with Applications, 47(10–11):1697–1705, May/ June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210483983X.

Wu:2004:DQD

[WS04b] Xionghua Wu and Ye Shen.

Differential quadrature domain decomposition method for a class of parabolic equations. Computers and Mathematics with Applications, 48(12):1819–1832, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004420.

Wong:2005:CSS

[WS05] P. J. Y. Wong and Y. C. Soh. Constant-sign solutions for a system of integral equations on time scales. Computers and Mathematics with Applications, 49 (2–3):271–280, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000325.

Wang:2007:GQS

[WS07] Xing-Yuan Wang and Yuan-Yuan Sun. The general quaternionic M–J sets on the mapping z ← z^{α} + c, (α ∈ N). Computers and Mathematics with Applications, 53(11):1718– 1732, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001447.

Wang:2008:MMC

[WS08a] Xia Wang and Xinyu Song. Mathematical models for the control of a pest population by infected pest. Computers and Mathematics with Applications, 56(1):266– 278, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800014X.

Wu:2008:NQN

[WS08b] Ting Wu and Linping Sun.

A new quasi-Newton pattern search method based on symmetric rank-one update for unconstrained optimization. Computers and Mathematics with Applications, 55(6):1201–1214, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005895.

Wu:2008:ELE

[WS08c] Yan Wu and P. N. Shivakumar. Eigenvalues of the Laplacian on an elliptic domain. Computers and Mathematics with Applications, 55(6):1129– 1136, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005780.

Wang:2008:MED

[WSG08] Dinan Wang, Jonathan L. Summers, and Philip H. Gaskell. Modelling of electrokinetically driven mixing flow in microchannels with patterned blocks. Computers and Mathematics with Applications, 55(7):1601– 1610, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006372.

Wang:2008:MPB

[WSL08] Weibing Wang, Jianhua Shen, and Zhiguo Luo. Multi-point boundary value problems for second-order functional differential equations. Computers and Mathematics with Applications, 56(8):2065–2072, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003416.

Wang:2008:VPA

[WSO08] Gaoxia Wang, Yi Shen, and Ming Ouyang. A vector partitioning approach to detecting community structure in complex networks. Computers and Mathematics with Applications, 55(12):2746– 2752, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007742.

Wu:2002:NII

[WSZ02] Xinyuan Wu, Rong Shao, and Yiran Zhu. New iterative improvement of a solution for an illconditioned system of linear equations based on a linear dynamic system. Computers and Mathematics with Applications, 44 (8–9):1109–1116, October/ November 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002195.

Wu:2003:JFC

[WSZ03] Xingyuan Wu, Rong Shao, and Yiran Zhu. Jacobi-free and complex-free method for finding simultaneously all zeros of polynomials having only real zeros. Computers and Mathematics with Applications, 46 (8–9):1387–1395, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902276.

Wang:2009:CCM

[WSZ09] Zhi-Gang Wang, Yong Sun, and Zhi-Hua Zhang. Certain classes of meromorphic multivalent functions. Computers and Mathematics with Applications, 58(7):1408–1417, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004817.

Wang:2000:BVS

[WT00] Hsiao-Fan Wang and Ruey-Chyn Tsaur. Bicriteria variable selection in a fuzzy regression equation. Computers and Mathematics with Applications, 40 (6–7):877–883, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002030.

Wu:2003:QPS

[WTY03] Xian Wu, B. Thompson, and George Xian-Zhi Yuan. Quasiequilibrium problem in H-spaces. Computers and Mathematics with Applications, 45(10–11):1629–1636, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001421.

Wu:2005:SIP

[WTZY05] C. Z. Wu, K. L. Teo, Yi Zhao, and W. Y. Yan. Solving an identification problem as an impulsive optimal parameter selection problem. Computers and Mathematics with Applications, 50(1–2):217– 229, July 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002440.

Wu:2000:NET

[Wu00] Xian Wu. New existence theorems for maximal elements in noncompact Hspaces with applications to equilibrium of games. Computers and Mathematics with Applications, 40 (8–9):1097–1106, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100850191.

Wu:2001:ELS

[Wu01]  Xian Wu.  Equilibria

of lower semicontinuous games. Computers and Mathematics with Applications, 42(1–2):13–22, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0898122101001262.

Wu:2003:LRA

[Wu03]  Hsien-Chung Wu.  Lin-

ear regression analysis for fuzzy input and output data using the extension principle. Computers and Mathematics with Applications, 45(12):1849–1859, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210390006X.

Wu:2004:EFO

[Wu04] Hsien-Chung Wu. Evaluate fuzzy optimization problems based on biobjective programming problems. Computers and Mathematics with Applications, 47(6–7):893–902, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900739.

Wu:2008:OCL

[Wu08a] Hsien-Chung Wu. Optimality conditions for linear programming problems with fuzzy coefficients. Computers and Mathematics with Applications, 55(12):2807– 2822, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007791.

Wu:2008:DEE

[Wu08b]  J. L. Wu.  Distribu-

tion and estimation for eigenvalues of real quaternion matrices. Computers and Mathematics with Applications, 55(9):1998– 2004, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006670.

Wu:2008:SIA

[Wu08c] Shanhe Wu. Some improvements of Aczél’s inequality and Popoviciu’s inequality. Computers and Mathematics with Applications, 56(5):1196–1205, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001211.

Wu:2009:AAS

[Wu09]  Zhao-Chun Wu.  Ap-

proximate analytical solutions of fifth-order boundary value problems by the variational iteration method. Computers and Mathematics with Applications, 58(11–12):2514–2517, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001916.

Wei:2000:EPW

[WW00] Yimin Wei and Hebing Wu. Expression for the perturbation of the weighted Moore– Penrose inverse. Computers and Mathematics with Applications, 39(5–6):13– 18, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000419.

Wang:2002:TCC

[WW02a] Hsiao-Fan Wang and Yu-Pin Wen. Time-constrained Chinese postman problems. Computers and Mathematics with Applications, 44(3–4):375–387, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001566.

Wu:2002:CSC

[WW02b] Jianhua Wu and Guangsheng Wei. Coexistence states for cooperative model with diffusion. Computers and Mathematics with Applications, 43(10–11):1277– 1290, May/June 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102001001.

Wang:2003:MIT

[WW03] Lianglong Wang and Zhicheng Wang. Monotone iterative techniques for parameterized BVPs of abstract semilinear evolution equations. Computers and Mathematics with Applications, 46 (8–9):1229–1243, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902148.

Wang:2004:OHP

[WW04a] Jing Wang and Ke Wang.

Optimal harvesting policy for single population with stage structure. Computers and Mathematics with Applications, 48(5–6):943–950, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003256.

Wang:2004:DCL

[WW04b] Yuanshi Wang and Hong Wu. Dynamics of competitive Lotka–Volterra systems that can be projected to a line. Computers and Mathematics with Applications, 47(8–9):1263–1271, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901216.

Wang:2008:ASS

[WW08a] Weiping Wang and Tianming Wang. Alternate Sylvester sums on the Frobenius set. Computers and Mathematics with Applications, 56(5):1328–1334, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001557.

Wang:2008:ESF

[WW08b] Zongyi Wang and Peixuan Weng. Existence of solutions for first order PBVPs with impulses on time scales. Computers and Mathematics with Applications, 56(8):2010–2018, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002782.

Wu:2008:AIP

[WW08c] Jinming Wu and Renhong Wang. Approximate implicitization of parametric surfaces by using compactly supported radial basis functions. Computers and Mathematics with Applications, 56(12):3064–3069, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004951.

Wu:2008:PPS

[WW08d] Jun Wu and Zhicheng Wang. Positive periodic solutions of second-order nonlinear differential systems with two parameters. Computers and Mathematics with Applications, 56(1): 43–54, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008231.

Wang:2009:GIB

[WW09a] Weiping Wang and Tianming Wang. General identities on Bell polynomials. Computers and Mathematics with Applications, 58(1): 104–118, July 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109002570.

Wang:2009:SLC

[WW09b] Zhengyu Wang and Xinyuan Wu. A semi-local convergence theorem for a robust revised Newton’s method. Computers and Mathematics with Applications, 58(7):1320–1327, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004969.

Wang:2009:NTP

[WW09c] Zhiwu Wang and Zhengguo Wang. A novel twophase heuristic method for vehicle routing problem with backhauls. Computers and Mathematics with Applications, 57(11–12):1923– 1928, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005312.

Wong:2008:EPSb

[WWC08] F. H. Wong, S. P. Wang, and T. G. Chen. Existence of positive solutions for second order functional differential equations. Computers and Mathematics with Applications, 56(10):2580–2587, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108003957.

Wei:2005:IDI

[WWH05] Yuchuan Wei, Ge Wang, and Jiang Hsieh. An intuitive discussion on the ideal ramp filter in computed tomography (I). Computers and Mathematics with Applications, 49(5–6):731–740, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000878.

Wu:2009:SFF

[WWHL09] Qing E. Wu, Tuo Wang, Yong Xuan Huang, and Ji Sheng Li. State fusion of fuzzy automata with application on target tracking. Computers and Mathematics with Applications, 57(6):949–960, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005804.

Wang:2006:GSV

[WWL06] Kaifa Wang, Wendi Wang, and Xianning Liu. Global stability in a viral infection model with lytic and nonlytic immune responses. Computers and Mathematics with Applications, 51(9–10):1593–1610, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001040.

Warne:2006:EPE

[WWS^{+}06] P. G. Warne, D. A. Polignone Warne, J. S. Sochacki, G. E. Parker, and D. C. Carothers. Explicit apriori error bounds and adaptive error control for approximation of nonlinear initial value differential systems. Computers and Mathematics with Applications, 52(12):1695–1710, December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210600352X.

Wang:2008:HEO

[WWW08] Peiguang Wang, Haixia Wu, and Yonghong Wu. Higher even-order convergence and coupled solutions for second-order boundary value problems on time scales. Computers and Mathematics with Applications, 55(8):1693– 1705, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006086.

Wang:2009:SMSa

[WWW^{+}09] Ji-Bo Wang, Dan Wang, Li-Yan Wang, Lin Lin, Na Yin, and Wei-Wei Wang. Single machine scheduling with exponential time-dependent learning effect and pastsequence-dependent setup times. Computers and Mathematics with Applications, 57(1):9–16, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005865.

Wang:2008:GAB

[WWWL08] Guangmin Wang, Zhongping Wan, Xianjia Wang, and Yibing Lv. Genetic algorithm based on simplex method for solving linearquadratic bilevel programming problem. Computers and Mathematics with Applications, 56(10):2550– 2555, November 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210800391X.

Wu:2004:OAN

[WWX04] Hong-Wu Wu, Qi-Ru Wang, and Yuan-Tong Xu. Oscillation and asymptotics for nonlinear second-order differential equations. Computers and Mathematics with Applications, 48(1–2):61– 72, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839981.

Wong:2009:EPS

[WWYL09] Fu-Hsiang Wong, Sheng-Ping Wang, Shiueh-Ling Yu, and Wei-Cheng Lian. Existence of positive solutions for an (n + 2)-order nonlinear BVP. Computers and Mathematics with Applications, 58(3):498–507, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003289.

Wang:2003:OCS

[WWZ03] Qi-Ru Wang, Xiao-Ming Wu, and Si-Ming Zhu. Oscillation criteria for secondorder nonlinear damped differential equations. Computers and Mathematics with Applications, 46(8– 9):1253–1262, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902161.

Wu:2000:ETS

[WX00a] Xin-Yuan Wu and Jian-Lin Xia. An explicit two-step method exact for the scalar test equation y′ = λy. Computers and Mathematics with Applications, 39(5–6):249–257, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000584.

Wu:2000:VFS

[WX00b] Xin-Yuan Wu and Jian-Lin Xia. The vector form of a sixth-order A-stable explicit one-step method for stiff problems. Computers and Mathematics with Applications, 39(3–4):247–257, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003491.

Wandzurat:2003:SQR

[WX03a] S. Wandzurat and H. Xiao.

Symmetric quadrature rules on a triangle. Computers and Mathematics with Applications, 45(12):1829– 1840, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900046.

Wong:2003:TSS

[WX03b] Patricia J. Y. Wong and Lihua Xie. Three symmetric solutions of Lidstone boundary value problems for difference and partial difference equations. Computers and Mathematics with Applications, 45(6–9):1445–1460, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001020.

Wu:2003:EAN

[WX03c] Xin-Yuan Wu and Jianlin Xia. Error analysis of a new transformation for multiple zeros finding free from derivative evaluations. Computers and Mathematics with Applications, 46 (8–9):1195–1200, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902112.

Wang:2009:WPM

[WX09a] Wen-Long Wang and Gen-Qi Xu. The well-posedness of an M/G/1 queue with second optional service and server breakdown. Computers and Mathematics with Applications, 57(5):729–739, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006263.

Wang:2009:GSW

[WX09b] Yue-Peng Wang and Da-Feng Xia. Generalized solitary wave solutions for the Klein–Gordon–Schrödinger equations. Computers and Mathematics with Applications, 58(11–12):2300–2306, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001564.

Wang:2009:RTO

[WXM09] Qi-Ru Wang, Yuan-Tong Xu, and Ronald M. Mathsen. Riccati techniques and oscillation for selfadjoint matrix Hamiltonian systems. Computers and Mathematics with Applications, 58(6):1211–1222, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109003721.

Wu:2001:QCM

[WXS01] Xin-Yuan Wu, Jian-Lin Xia, and Rong Shao. Quadratically convergent multiple roots finding method without derivatives. Computers and Mathematics with Applications, 42(1–2):115– 119, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001365.

Wang:2001:NVE

[WXW01] Yiju Wang, Naihua Xiu, and Changyu Wang. A new version of extragradient method for variational inequality problems. Computers and Mathematics with Applications, 42 (6–7):969–979, September/ October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002139.

Wei:2001:MTP

[WY01] Quanling Wei and Hong Yan. A method of transferring polyhedron between the intersection-form and the sum-form. Computers and Mathematics with Applications, 41(10–11):1327–1342, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001006.

Wei:2002:ABA

[WY02] Quanling Wei and Hong Yan. An algebra-based approach for linearly constrained concave minimization. Computers and Mathematics with Applications, 43(8–9):965–974, April/May 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102800062.

Wang:2006:ODE

[WY06] Youbin Wang and Jurang Yan. Oscillation of a differential equation with fractional delay and piecewise constant arguments. Computers and Mathematics with Applications, 52(6– 7):1099–1106, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002811.

Wang:2007:STP

[WY07] Lihe Wang and Fahuai Yi. Schauder theory for a parabolic equation in a wedge-shaped domain. Computers and Mathematics with Applications, 53(7):1107–1118, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000703.

Wang:2009:CIC

[WY09a] Xingyuan Wang and Canghai Yu. Cryptanalysis and improvement on a cryptosystem based on a chaotic map. Computers and Mathematics with Applications, 57(3):476–482, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006081.

Wei:2009:CDL

[WY09b] Zhouchao Wei and Qigui Yang. Controlling the diffusionless Lorenz equations with periodic parametric perturbation. Computers and Mathematics with Applications, 58(10):1979– 1987, November 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S089812210900577X.

Wang:2000:SCA

[WYH00] Shengli Wang, Hai Yang, and Bingsheng He. Solving a class of asymmetric variational inequalities by a new alternating direction method. Computers and Mathematics with Applications, 40 (8–9):927–937, October/ November 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210085004X.

Wang:2008:SST

[WYL08a] Deng-Shan Wang, Zhenya Yan, and Hongbo Li. Some special types of solutions of a class of the (N + 1)- dimensional nonlinear wave equations. Computers and Mathematics with Applications, 56(6):1569–1579, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001831.

Wang:2008:EAB

[WYL08b] Hong-Yong Wang, Shou-Zhi Yang, and Xiu-Juan Li. Error analysis for bivariate fractal interpolation functions generated by 3-D perturbed iterated function systems. Computers and Mathematics with Applications, 56(7):1684–1692, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002277.

Wang:2009:NPR

[WYL^{+}09] Anna Wang, Wenjing Yuan, Junfang Liu, Zhiguo Yu, and Hua Li. A novel pattern recognition algorithm: Combining ART network with SVM to reconstruct a multi-class classifier. Computers and Mathematics with Applications, 57(11–12):1908– 1914, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005282.

Wu:2009:HIS

[WYSS09] Ting Wu, Yingsha Yang, Linping Sun, and Hu Shao. A heuristic iterated-subspace minimization method with pattern search for unconstrained optimization. Computers and Mathematics with Applications, 58(10): 2051–2059, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109005860.

Wang:2005:SES

[WYW05] Chunpeng Wang, Jingxue Yin, and Zejia Wang. Similar entropy solutions of a singular diffusion equation. Computers and Mathematics with Applications, 49(7–8):1059–1068, May/June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001227.

Wei:2001:DIL

[WYZY01] Yuchuan Wei, Guangwen Yan, Qishan Zhang, and Kai Yao. Dirichlet inversion and lattice inversion problem. Computers and Mathematics with Applications, 41(5–6):641– 645, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003072.

Wang:2000:NEP

[WZ00] Z. C. Wang and R. Y.

Zhang. Nonexistence of eventually positive solutions of a difference inequality with multiple and variable delays and coefficients. Computers and Mathematics with Applications, 40 (6–7):705–712, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001899.

Wang:2006:ASM

[WZ06] Zhiyong Wang and Chengjian Zhang. An analysis of stability of Milstein method for stochastic differential equations with delay. Computers and Mathematics with Applications, 51(9–10):1445– 1452, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000915.

Wang:2007:ORP

[WZ07a] Guan Jun Wang and Yuan Lin Zhang. An optimal replacement policy for a two-component series system assuming geometric process repair. Computers and Mathematics with Applications, 54(2):192– 202, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002052.

Wang:2007:DTG

[WZ07b] Jinting Wang and Qing Zhao. A discrete-time Geo/G/1 retrial queue with starting failures and second optional service. Computers and Mathematics with Applications, 53(1):115–127, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001071.

Wu:2007:SFO

[WZ07c] Haihua Wu and Zhan Zhou. Stability for first order delay dynamic equations on time scales. Computers and Mathematics with Applications, 53(12):1820– 1831, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001502.

See remarks [Zho08].

Wang:2008:IIA

[WZ08a] Renhong Wang and Xiaolei Zhang. Interval iterative algorithm for computing the piecewise algebraic variety. Computers and Mathematics with Applications, 56(2):565– 571, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000850.

Wu:2008:LSD

[WZ08b] Qiang Wu and Ding-Xuan Zhou. Learning with sample dependent hypothesis spaces. Computers and Mathematics with Applications, 56(11):2896–2907, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004926.

Wu:2008:ABM

[WZ08c] Xiaonan Wu and Jiwei Zhang. Artificial boundary method for two-dimensional Burgers’ equation. Computers and Mathematics with Applications, 56(1):242– 256, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000126.

Wen:2009:ESG

[WZ09] Yi-Xin Wen and Xin-Wei Zhou. Exact solutions for the generalized nonlinear heat conduction equations using the Exp-function method. Computers and Mathematics with Applications, 58(11–12):2464–2467, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001989.

Wang:2006:ELS

[WZE^{+}06] Hong Wang, Weidong Zhao, R. E. Ewing, M. Al-Lawatia, M. S. Espedal, and A. S. Telyakovskiy. An Eulerian– Lagrangian solution technique for single-phase compositional flow in threedimensional porous media. Computers and Mathematics with Applications, 52(5): 607–624, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210600246X.

Wang:2000:OCN

[WZG00] Peiguang Wang, Jiuli Zhao, and Weigao Ge. Oscillation criteria of nonlinear hyperbolic equations with functional arguments. Computers and Mathematics with Applications, 40(4–5):513–521, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001760.

Wang:2004:GOG

[WZG04] Yanjun Wang, Kecun Zhang, and Yuelin Gao. Global optimization of generalized geometric programming. Computers and Mathematics with Applications, 48(10–11):1505–1516, November/December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004055.

Wei:2009:ASD

[WZZ09] Xiaopeng Wei, Dongsheng Zhou, and Qiang Zhang. On asymptotic stability of discrete-time non-autonomous delayed Hopfield neural networks. Computers and Mathematics with Applications, 57(11–12):1938– 1942, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005270.

Xu:2009:PCA

[XA09] Neng Xu and M. K. Aouf.

Properties of certain analytic multivalent functions defined by a linear operator. Computers and Mathematics with Applications, 58(6):1169–1179, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109004854.

Xu:2000:PST

[XC00] Rui Xu and Lansun Chen. Persistence and stability for a two-species ratiodependent predator–prey system with time delay in a two-patch environment. Computers and Mathematics with Applications, 40(4–5):577–588, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001814.

Xiong:2004:BFB

[XC04a] Chengji Xiong and Huaihui Chen. Bloch functions and Bloch constants. Computers and Mathematics with Applications, 47(6–7):931–946, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900776.

Xue:2004:ASO

[XC04b] Lian Xue and Xiao-Liang Cheng. An algorithm for solving the obstacle problems. Computers and Mathematics with Applications, 48(10– 11):1651–1657, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004183.

Xia:2006:GAP

[XC06] Yonghui Xia and Jinde Cao.

Global attractivity of a periodic ecological model with m-predators and n-preys by “Pure-delay type” system. Computers and Mathematics with Applications, 52 (6–7):829–852, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002628.

Xu:2007:SII

[XC07] Aimin Xu and Dezao Cui. Some integral inequalities for functions with ( n-1)st derivatives of bounded variation. Computers and Mathematics with Applications, 54(1):31– 37, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700171X.

Xu:2007:GCR

[XCD07] Rui Xu, M. A. J. Chaplain, and F. A. Davidson. Global convergence of a reaction-diffusion predator–prey model with stage structure and nonlocal delays. Computers and Mathematics with Applications, 53(5):770–788, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000302.

Xu:2007:RFI

[XCH07] Min Xu, Xujin Chen, and Xiaodong Hu. On the restricted forwarding index problem in communication networks. Computers and Mathematics with Applications, 53(11):1633– 1643, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001083.

Xiang:2008:PCC

[XCL^{+}08] Linying Xiang, Zengqiang Chen, Zhongxin Liu, Fei Chen, and Zhuzhi Yuan. Pinning control of complex dynamical networks with heterogeneous delays. Computers and Mathematics with Applications, 56(5):1423–1433, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001673.

Xu:2008:NDG

[XCS^{+}08] Xulin Xu, Zengqiang Chen, Guangya Si, Xiaofeng Hu, and Pi Luo. A novel definition of generalized synchronization on networks and a numerical simulation example. Computers and Mathematics with Applications, 56(11):2789–2794, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004586.

Xie:2002:HGA

[XD02] Jinxing Xie and Jiefang Dong. Heuristic genetic algorithms for general capacitated lot-sizing problems. Computers and Mathematics with Applications, 44(1–2):263–276, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001463.

Xing:2009:IGO

[XD09] Yuming Xing and Shusen Ding. Inequalities for Green’s operator with Lipschitz and BMO norms. Computers and Mathematics with Applications, 58(2): 273–280, July 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109002624.

Xie:2009:GBI

[XDD09] Li Xie, Jie Ding, and Feng Ding. Gradient based iterative solutions for general linear matrix equations. Computers and Mathematics with Applications, 58(7):1441–1448, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004684.

Xue:2007:NFP

[XG07] Chunyan Xue and Weigao Ge. New fixed point theorems and applications to boundary value problems with p-Laplacian. Computers and Mathematics with Applications, 53(5):706–716, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000223.

Xia:2002:CAC

[XH02] Bican Xia and Xiaorong Hou. A complete algorithm for counting real solutions of polynomial systems of equations and inequalities. Computers and Mathematics with Applications, 44(5–6):633–642, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001785.

Xia:2007:VIG

[XH07] Fu-Quan Xia and Nan-Jing Huang. Variational inclusions with a general H-monotone operator in Banach spaces. Computers and Mathematics with Applications, 54(1):24– 30, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700168X.

Xiang:2005:NLC

[XHZ05] Guanghui Xiang, Maoan Han, and Tonghua Zhang. The number of limit cycles for a family of polynomial systems. Computers and Mathematics with Applications, 49(11–12):1669– 1678, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001884.

Xiao:2007:EAV

[Xia07] Ai-Guo Xiao. Error analysis of variable stepsize Runge–Kutta methods for a class of multiply-stiff singular perturbation problems. Computers and Mathematics with Applications, 53(12):1854–1866, June 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001575.

Xie:2000:LSS

[Xie00]  Dong-Xiu Xie.  Least-

squares solution for inverse eigenpair problem of nonnegative definite matrices. Computers and Mathematics with Applications, 40 (10–11):1241–1251, November/December 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100002352.

Xie:2001:ACG

[Xie01]  Ping Ding Xie.  Ab-

stract convexity and generalizations of himmelberg type fixed-point theorems. Computers and Mathematics with Applications, 41(3–4):497–504, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002911.

Xie:2008:CBT

[Xie08]  Changzhen Xie.  Con-

struction of biorthogonal two-direction refinable function and two-direction wavelet with dilation factor m. Computers and Mathematics with Applications, 56(7):1845–1851, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002642.

Xiao:2001:MPR

[XL01] Ti-Jun Xiao and Jin Liang.

Multiplicative perturbations of C-regularized semigroups. Computers and Mathematics with Applications, 41(10–11):1215–1221, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210100092X.

Xu:2009:LFH

[XL09] Hang Xu and Shi-Jun Liao.

Laminar flow and heat transfer in the boundarylayer of non-Newtonian fluids over a stretching flat sheet. Computers and Mathematics with Applications, 57(9):1425–1431, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000698.

Xu:2003:GAS

[XLL03] Bingji Xu, Xinzhi Liu, and Xiaoxin Liao. Global asymptotic stability of highorder Hopfield type neural networks with time delays. Computers and Mathematics with Applications, 45(10–11):1729–1737, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001512.

Xu:2004:ASL

[XLL04] Bingji Xu, Xinzhi Liu, and Xiaoxin Liao. Absolute stability of Lurie systems with impulsive effects. Computers and Mathematics with Applications, 47(2–3):419–425, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210490034X.

Xiao:2008:LBR

[XLL08] Hongmin Xiao, Zehui Li, and Weiwei Liu. The limit behavior of a risk model based on entrance processes. Computers and Mathematics with Applications, 56(5):1434–1440, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001685.

Xu:2009:ASI

[XLT09a] Bingji Xu, Xiang Liu, and Kok Lay Teo. Asymptotic stability of impulsive highorder Hopfield type neural networks. Computers and Mathematics with Applications, 57(11–12):1968– 1977, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005257.

Xu:2009:GES

[XLT09b] Bingji Xu, Xiang Liu, and Kok Lay Teo. Global exponential stability of impulsive high-order Hopfield type neural networks with delays. Computers and Mathematics with Applications, 57(11–12):1959– 1967, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005208.

Xie:2002:ESS

[XLZ02] Fu-Ding Xie, Ming Li, and Yong Zhang. Exact solutions of some systems of nonlinear partial differential equations using symbolic computation. Computers and Mathematics with Applications, 44(5–6):711–716, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001840.

Xu:2008:FTC

[XM08] Luoshan Xu and Xuxin Mao. Formal topological characterizations of various continuous domains. Computers and Mathematics with Applications, 56(2): 444–452, July 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210800028X.

Xu:2009:FRA

[XMW09] F. F. Xu, D. Q. Miao, and L. Wei. Fuzzy-rough attribute reduction via mutual information with an application to cancer classification. Computers and Mathematics with Applications, 57(6):1010–1017, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800566X.

Xiao:2001:RPO

[XT01] Ai-Guo Xiao and Yi-Fa Tang. Regularity properties of one-leg methods for delay differential equations. Computers and Mathematics with Applications, 41(3–4):363–372, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002790.

Xiao:2002:SPM

[XT02] Ai-Guo Xiao and Yi-Fa Tang. Symplectic properties of multistep Runge– Kutta methods. Computers and Mathematics with Applications, 44(10– 11):1329–1338, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002602.

Xiao:2004:OPS

[XT04a] Ai-Guo Xiao and Yi-Fa Tang. Order properties of symplectic Runge– Kutta–Nyström methods. Computers and Mathematics with Applications, 47 (4–5):569–582, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900466.

Xie:2004:FOC

[XT04b] Sheng Li Xie and Chuan Jun Tian. Frequent oscillatory criteria for partial difference equations with several delays. Computers and Mathematics with Applications, 48(3–4):335–345, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840586.

Xu:2008:RSU

[XT08] Honglei Xu and Kok Lay Teo. Robust stabilization of uncertain impulsive switched systems with delayed control. Computers and Mathematics with Applications, 56(1):63– 70, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107008255.

Xie:2001:OCP

[XTX01] Shengli Xie, Chuanjun Tian, and Zhendong Xie. Oscillation of a class of partial difference equations with unbounded delay. Computers and Mathematics with Applications, 42(3–5):529–541, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001754.

Xue:2008:EVF

[XTZ08] Feng Xue, Wansheng Tang, and Ruiqing Zhao. The expected value of a function of a fuzzy variable with a continuous membership function. Computers and Mathematics with Applications, 55(6):1215– 1224, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005901.

Xu:2001:ABN

[Xu01]  Daoyi Xu.  Asymptotic

behavior of nonlinear difference equations with delays. Computers and Mathematics with Applications, 42(3–5):393–398, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210100164X.

Xu:2002:UEB

[Xu02a] Da Xu. Uniform L^{1} error bounds for the semidiscrete solution of a Volterra equation with completely monotonic convolution kernel. Computers and Mathematics with Applications, 43(10–11):1303–1318, May/ June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001025.

Xu:2002:NF

[Xu02b] Yan Xu. The α-normal functions. Computers and Mathematics with Applications, 44(3–4):357–363, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001542.

Xu:2003:IAS

[Xu03a] Daoyi Xu. Invariant and attracting sets of Volterra difference equations with delays. Computers and Mathematics with Applications, 45(6–9):1311–1317, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001044.

Xu:2003:MFS

[Xu03b]  Yan Xu.  Meromorphic

functions sharing one finite set. Computers and Mathematics with Applications, 45(10–11):1489–1495, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001329.

Xu:2004:CDL

[Xu04]  Guoliang Xu.  Conver-

gence of discrete Laplace– Beltrami operators over surfaces. Computers and Mathematics with Applications, 48(3–4):347–360, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840598.

Xu:2006:NIE

[Xu06] Aimin Xu. A note to inequalities of Euler–Grüss type. Computers and Mathematics with Applications, 52(6–7):941–952, September/October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002689.

Xu:2007:HHP

[Xu07a] Lan Xu. He’s homotopy perturbation method for a boundary layer equation in unbounded domain. Computers and Mathematics with Applications, 54(7–8):1067–1070, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002787.

Xu:2007:VIM

[Xu07b] Lan Xu. Variational iteration method for solving integral equations. Computers and Mathematics with Applications, 54(7–8):1071– 1078, October 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107002805.

Xu:2008:UMF

[Xu08a]  Yan Xu.  Uniqueness

of meromorphic functions sharing one value. Computers and Mathematics with Applications, 56(10):2692– 2699, November 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004094.

Xu:2008:OCS

[Xu08b] Zhiting Xu. Oscillation criteria for second order forced elliptic differential equations with mixed nonlinearities. Computers and Mathematics with Applications, 56(5):1225–1235, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001247.

Xie:2005:CAT

[XW05] Jinxing Xie and Xijun Wang. Complexity and algorithms for two-stage flexible flowshop scheduling with availability constraints. Computers and Mathematics with Applications, 50(10–12):1629–1638, November/December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210500413X.

Xing:2004:GEC

[XWD04] Yuming Xing, Baoling Wang, and Shusen Ding. Global estimates for compositions of operators applied to differential forms. Computers and Mathematics with Applications, 48 (12):1905–1913, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004481.

Xiang:2006:RTI

[XWL06] X. Xiang, Wei Wei, and Hongbo Liu. Relaxed trajectories of integrodifferential equations and optimal control on Banach space. Computers and Mathematics with Applications, 52(5):735–748, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002525.

Xiao:2008:LMB

[XWW08] Yunhai Xiao, Zengxin Wei, and Zhiguo Wang. A limited memory BFGStype method for largescale unconstrained optimization. Computers and Mathematics with Applications, 56(4):1001–1009, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001028.

Xiu:2004:MFP

[XWZ04] Naihua Xiu, Yiju Wang, and Xiangsun Zhang. Modified fixed-point equations and related iterative methods for variational inequalities. Computers and Mathematics with Applications, 47(6–7):913–920, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900752.

Xu:2002:GFA

[XX02] Zengkun Xu and Dapeng Xu. A general form of Alzer’s inequality. Computers and Mathematics with Applications, 44(3–4):365– 373, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001554.

Xie:2008:ALB

[XX08] Linsen Xie and Tingfan Xie. The approximation of localized Bernstein polynomials. Computers and Mathematics with Applications, 56(4):1040–1044, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001065.

Xu:2009:AIS

[XX09a] Liguang Xu and Daoyi Xu.

P -attracting and p-invariant sets for a class of impulsive stochastic functional differential equations. Computers and Mathematics with Applications, 57(1):54–61, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005889.

Xu:2009:DSG

[XX09b] Zhiqiang Xu and Guoliang Xu. Discrete schemes for Gaussian curvature and their convergence. Computers and Mathematics with Applications, 57(7):1187– 1195, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000480.

Xiao:2009:HQC

[XXC^{+}09] Jianhua Xiao, Jin Xu, Zhihua Chen, Kai Zhang, and Linqiang Pan. A hybrid quantum chaotic swarm evolutionary algorithm for DNA encoding. Computers and Mathematics with Applications, 57(11–12):1949– 1958, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005221.

Xiaoning:2006:PSS

[XXD06] Lin Xiaoning, Li Xiaoyue, and Jiang Daqing. Positive solutions to superlinear semipositone periodic boundary value problems with repulsive weak singular forces. Computers and Mathematics with Applications, 51(3–4):507–514, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005171.

Xie:2004:MDA

[XXLD04] Jinxing Xie, Wenxun Xing, Zhixin Liu, and Jiefang Dong. Minimum deviation algorithm for two-stage nowait flowshops with parallel machines. Computers and Mathematics with Applications, 47(12):1857– 1863, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840215.

Xia:2004:NRI

[XY04] Bican Xia and Lu Yang.

A new result on the pirreducibility of binding polynomials. Computers and Mathematics with Applications, 48(12):1811– 1817, December 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104004419.

Xu:2009:ETD

[XYC^{+}09] Fei Xu, Wei Yan, Yu-Lin Chen, Chuan-Qi Li, and Ya-Nan Zhang. Evaluation of two-dimensional ZK–MEW equation using the Exp-function method. Computers and Mathematics with Applications, 58(11–12):2307–2312, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900162X.

Xia:2006:RSI

[XZ06] Bican Xia and Ting Zhang. Real solution isolation using interval arithmetic. Computers and Mathematics with Applications, 52 (6–7):853–860, September/ October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002896.

Xu:2008:MMS

[XZ08a] Hongru Xu and Jinping Zeng. A multisplitting method for symmetrical affine second-order cone complementarity problem. Computers and Mathematics with Applications, 55 (3):459–469, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122107004828.

Xu:2008:CFR

[XZ08b] Jiuping Xu and Lihui Zhao. A class of fuzzy rough expected value multi-objective decision making model and its application to inventory problems. Computers and Mathematics with Applications, 56(8):2107–2119, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003465.

Xiong:2009:ESR

[XZ09a] Zhiping Xiong and Bing Zheng. On the eigenvalues of a specially rank-r updated complex matrix. Computers and Mathematics with Applications, 57(10):1645– 1650, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002247.

Xu:2009:VAA

[XZ09b] Lan Xu and Nan Zhang.

A variational approach to analyzing catalytic reactions in short monoliths. Computers and Mathematics with Applications, 58(11–12):2460–2463, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900193X.

Xiao:2009:RBI

[XZDD09] Yongsong Xiao, Yong Zhang, Jie Ding, and Jiyang Dai. The residual based interactive least squares algorithms and simulation studies. Computers and Mathematics with Applications, 58(6):1190–1197, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002910.

Xu:2001:GDH

[XZZ01] Daoyi Xu, Hongyong Zhao, and Hong Zhu. Global dynamics of Hopfield neural networks involving variable delays. Computers and Mathematics with Applications, 42(1–2):39– 45, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001286.

Yan:2001:OFO

[Yan01] Jurang Yan. Oscillation of first-order impulsive differential equations with advanced argument. Computers and Mathematics with Applications, 42(10– 11):1353–1363, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002450.

Yan:2002:IST

[Yan02] Guozheng Yan. Inverse scattering for transmission problem. Computers and Mathematics with Applications, 44(3–4):439–444, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001608.

Yan:2003:GTA

[Yan03a] Zhenya Yan. Generalized transformations and abundant new families of exact solutions for (2 + 1)- dimensional dispersive long wave equations. Computers and Mathematics with Applications, 46(8– 9):1363–1372, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902252.

Yang:2003:IMC

[Yan03b] Cheng-Hong Yang. An interactive Morse code emulation management system. Computers and Mathematics with Applications, 46(2–3):479–492, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210390040X.

Yang:2003:PSS

[Yan03c] Guang Chong Yang. Positive solutions of some second-order nonlinear singular differential equations. Computers and Mathematics with Applications, 45 (4–5):605–614, February/ March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000208.

Yang:2003:APC

[Yan03d] Laurence Tianruo Yang. Accuracy of preconditioned CG-type methods for least squares problems. Computers and Mathematics with Applications, 45(1–3):77– 96, January/February 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103800093.

Yang:2003:NSN

[Yan03e] Xiaojing Yang. Nonoscillatory solutions of nonlinear differential systems. Computers and Mathematics with Applications, 46 (8–9):1347–1362, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902240.

Yang:2003:GSC

[Yan03f]  Yu-Fei Yang.  Globally

and superlinearly convergent algorithms for the solution of box-constrained optimization. Computers and Mathematics with Applications, 45(12):1807– 1821, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900022.

Yan:2004:ISM

[Yan04a] Guozheng Yan. Inverse scattering by a multilayered obstacle. Computers and Mathematics with Applications, 48(12):1801–1810, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004407.

Yan:2004:UIS

[Yan04b] Guozheng Yan. Uniqueness of the inverse scattering problem of a locally perturbed half-plane. Computers and Mathematics with Applications, 48(3–4):411– 418, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840641.

Yan:2004:OPS

[Yan04c] Jurang Yan. Oscillation properties of a second-order impulsive delay differential equation. Computers and Mathematics with Applications, 47(2–3):253–258, January/February 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104900223.

Yang:2004:EPS

[Yan04d]  Xiaojing Yang.  Exis-

tence of positive solutions for quasi-linear differential equations. Computers and Mathematics with Applications, 47(8–9):1233–1239, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104901174.

Yan:2005:USP

[Yan05a] Guozheng Yan. Uniqueness of a scattering problem with L^{2} boundary values. Computers and Mathematics with Applications, 50 (8–9):1363–1370, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210500386X.

Yang:2005:EPS

[Yan05b] Xiaojing Yang. Existence of periodic solutions of planar systems. Computers and Mathematics with Applications, 50(3–4):645–652, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002695.

Yang:2006:PSS

[Yan06a] G. C. Yang. Positive solutions of singular Dirichlet boundary value problems with sign-changing nonlinearities. Computers and Mathematics with Applications, 51(9–10):1463– 1470, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000939.

Yang:2006:UAS

[Yan06b] Shijun Yang. A unified approach to some inequalities of Ostrowski-Grüss type. Computers and Mathematics with Applications, 51(6–7):1047–1056, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106000356.

Yang:2007:EUP

[Yan07]  Zhilin Yang.  Existence

and uniqueness of positive solutions for a higher order boundary value problem. Computers and Mathematics with Applications, 54(2):220–228, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107002088.

Yang:2008:NSS

[Yan08] Cheng-Fu Yang. A note on “Soft Set Theory” [Comput. Math. Appl. 45 (4–5) (2003) 555–562]. Computers and Mathematics with Applications, 56(7):1899–1900, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001946.

See [MBR03].

Yao:2004:EIS

[Yao04]  Qingliu Yao.  Existence

and iteration of n symmetric positive solutions for a singular two-point boundary value problem. Computers and Mathematics with Applications, 47(8–9):1195–1200, April/ May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901137.

Yao:2005:SIP

[Yao05] Qingliu Yao. Successive iteration and positive solution for nonlinear secondorder three-point boundary value problems. Computers and Mathematics with Applications, 50(3–4):433– 444, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002877.

Yao:2007:SIP

[Yao07] Qingliu Yao. Successive iteration of positive solution for a discontinuous third-order boundary value problem. Computers and Mathematics with Applications, 53(5):741–749, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000272.

Yaslan:2008:MPS

[Yas08] Ismail Yaslan. Multiple positive solutions for nonlinear three-point boundary value problems on time scales. Computers and Mathematics with Applications, 55(8):1861–1869, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700658X.

Yavin:2000:CTL

[Yav00] Y. Yavin. Control of a three-link manipulator with a constraint on the velocity of its end-effector. Computers and Mathematics with Applications, 40(10– 11):1263–1273, November/ December 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100002376.

Yavin:2005:CMD

[Yav05a] Y. Yavin. Control of the motion of a disk rollin on a curve in R. Computers and Mathematics with Applications, 50(5–6):855–868, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105003214.

Yavin:2005:DCM

[Yav05b] Y. Yavin. Directional control of the motion of a rolling disk by using a rotor fixed in the disk’s plane. Computers and Mathematics with Applications, 50(10– 12):1659–1668, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105004165.

Yavin:2005:PPC

[Yav05c] Y. Yavin. Point-to-point and collision avoidance control of the motion of an autonomous bicycle. Computers and Mathematics with Applications, 50(10– 12):1525–1542, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500427X.

Yavin:2005:SMD

[Yav05d] Y. Yavin. Stabilization of the motion of a disk rolling on a horizontal plane by using a translational moving rod along its axis. Computers and Mathematics with Applications, 50(3–4):347– 357, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002798.

Yavin:2006:DKM

[Yav06]  Y. Yavin.  The deriva-

tion of a kinematic model from the dynamic model of the motion of a riderless bicycle. Computers and Mathematics with Applications, 51(6–7):865–878, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000253.

Yavin:2007:CMD

[Yav07a] Y. Yavin. Control of the motion of a disk rolling on a plane curve. Computers and Mathematics with Applications, 54(11–12):1329–1340, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001289.

Yavin:2007:MCM

[Yav07b]  Y. Yavin.  Modelling

and control of the motion of a riderless bicycle rolling on a moving plane. Computers and Mathematics with Applications, 54(11–12):1319–1328, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001113.

Yanev:2008:LAC

[YAVB08] Nicola Yanev, Rumen Andonov, Philippe Veber, and Stefan Balev. Lagrangian approaches for a class of matching problems in computational biology. Computers and Mathematics with Applications, 55(5):1054– 1067, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005640.

Yusufoglu:2007:AVI

[YB07a] Elcin Yusufoğlu and Ahmet Bekir. Application of the variational iteration method to the regularized long wave equation. Computers and Mathematics with Applications, 54(7–8):1154–1161, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002829.

Yusufoglu:2007:NSE

[YB07b] Elcin Yusufoğlu and Ahmet Bekir. Numerical simulation of equal-width wave equation. Computers and Mathematics with Applications, 54(7–8):1147–1153, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107003057.

Yusufoglu:2008:SCN

[YB08] E. Yusufoğlu and A. Bekir.

Symbolic computation and new families of exact travelling solutions for the Kawahara and modified Kawahara equations. Computers and Mathematics with Applications, 55(6):1113– 1121, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005767.

Yan:2006:MFM

[YC06] Chang-Dau Yan and Wei-Hua Chieng. Method for finding multiple roots of polynomials. Computers and Mathematics with Applications, 51(3–4):605–620, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105005250.

Yao:2007:NAA

[YC07a] Ming-Jong Yao and Weng-Ming Chu. A new approximation algorithm for obtaining the probability distribution function for project completion time. Computers and Mathematics with Applications, 54(2):282– 295, July 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002131.

You:2007:DPS

[YC07b] Peng-Sheng You and Ta Cheng Chen. Dynamic pricing of seasonal goods with spot and forward purchase demands. Computers and Mathematics with Applications, 54(4):490–498, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002325.

Yang:2002:CAS

[YCCY02] Cheng-Huei Yang, Pau-Choo Chung, Li-Yeh Chuang, and Cheng-Hong Yang. Communication aid system for users with physical impairments. Computers and Mathematics with Applications, 43(6–7):901–910, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003303.

Yuan:2008:GRS

[YD08] Yongxin Yuan and Hua Dai.

Generalized reflexive solutions of the matrix equation AXB = D and an associated optimal approximation problem. Computers and Mathematics with Applications, 56(6):1643–1649, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001910.

Yao:2001:ELS

[YE01] Ming-Jong Yao and S. E.

Elmaghraby. The economic lot scheduling problem under power-of-two policy. Computers and Mathematics with Applications, 41(10–11):1379–1393, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001031.

Ye:2002:FCM

[Ye02] Xingde Ye. The Fourier collocation method for the Cahn–Hilliard equation. Computers and Mathematics with Applications, 44(1–2):213–229, July 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001426.

Yenicerioglu:2008:SPS

[Yen08] Ali Fuat Yeniçerioğlu. Stability properties of second order delay integrodifferential equations. Computers and Mathematics with Applications, 56(12): 3109–3117, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004501.

Yavin:2004:CTL

[YF04] Y. Yavin and C. Frangos. Control of a threelink manipulator: collision avoidance of a moving obstacle. Computers and Mathematics with Applications, 47(10–11):1627–1639, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839762.

Yi:2004:NNB

[YFT04] Zhang Yi, Yan Fu, and Hua Jin Tang. Neural networks based approach for computing eigenvectors and eigenvalues of symmetric matrix. Computers and Mathematics with Applications, 47(8–9):1155–1164, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901101.

Yu:2008:SAS

[YG08] Huidan Yu and Sharath S.

Girimaji. Study of axisswitching and stability of laminar rectangular jets using lattice Boltzmann method. Computers and Mathematics with Applications, 55(7):1611–1619, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006335.

Yin:2009:CAR

[YGH09] Yunfei Yin, Guanghong Gong, and Liang Han. Control approach to rough set reduction. Computers and Mathematics with Applications, 57(1):117–126, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005841.

Yuan:2003:CPD

[YH03] Zhaohui Yuan and Lihong Huang. Convergence and periodicity in a discretetime network of two neurons with self-connections. Computers and Mathematics with Applications, 46 (8–9):1337–1345, October/ November 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103902239.

Yang:2004:DMS

[YH04] Qingzhi Yang and Yu Hong. Discretization method for semi-definite programming. Computers and Mathematics with Applications, 48 (12):1937–1945, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104004511.

Yang:2009:NDD

[YH09] Degang Yang and Chunyan Hu. Novel delaydependent global asymptotic stability condition of Hopfield neural networks with delays. Computers and Mathematics with Applications, 57(11–12):1978– 1984, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005245.

Yin:2011:NSM

[YH11] Na Yin and Xue Huang.

A note on “Single-machine scheduling with general learning functions”. Computers and Mathematics with Applications, 61(9): 2929–2930, May 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122111002318. See

[Wan08b].

Yan:2008:SAP

[YHS08] Xihong Yan, Deren Han, and Wenyu Sun. A selfadaptive projection method with improved step-size for solving variational inequalities. Computers and Mathematics with Applications, 55(4):819–832, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107005135.

Yang:2004:SID

[YHT04] Gou-Sheng Yang, Dah-Yan Hwang, and Kuei-Lin Tseng. Some inequalities for differentiable convex and concave mappings. Computers and Mathematics with Applications, 47(2–3):207–216, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210490017X.

Yuksel:2006:WAC

[YI06] I. Yuksel and N. Ispir. Weighted approximation by a certain family of summation integral-type operators. Computers and Mathematics with Applications, 52 (10–11):1463–1470, November/December 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122106003336.

Yildirim:2008:SBF

[Yil08] Ahmet Yildirim. Solution of BVPs for fourth-order integro-differential equations by using homotopy perturbation method. Computers and Mathematics with Applications, 56(12): 3175–3180, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122108004422.

Yildirim:2009:AHH

[Yil09] Ahmet Yildirim. Application of He’s homotopy perturbation method for solving the Cauchy reactiondiffusion problem. Computers and Mathematics with Applications, 57(4):612–618, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006755.

Yao:2008:QFI

[YJCZ08] Yuqin Yao, Jie Ji, Dengyuan Chen, and Yunbo Zeng. The quadratic-form identity for constructing the Hamiltonian structures of the discrete integrable systems. Computers and Mathematics with Applications, 56(11):2874–2882, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800463X.

Yavin:2000:AEI

[YK00] Y. Yavin and P. D. Kemp.

The application of extended inverse dynamics control. Computers and Mathematics with Applications, 40(4–5):669–677, August/ September 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100001863.

Yuan:2004:DCF

[YL04a] Xue-Hai Yuan and E. S. Lee. The definition of convex fuzzy subset. Computers and Mathematics with Applications, 47(1):101–113, January 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900090.

Yuan:2004:FGB

[YL04b] Xuehai Yuan and E. S. Lee. Fuzzy group based on fuzzy binary operation. Computers and Mathematics with Applications, 47 (4–5):631–641, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210490051X.

See [Ano05f].

Yue:2004:SRM

[YL04c] Dong Yue and James Lam. Suboptimal robust mixed H_{2}H_{∞} controller design for uncertain descriptor systems with distributed delays. Computers and Mathematics with Applications, 47(6–7):1041–1055, March/ April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900855.

Yuan:2005:CFF

[YL05] Xuehai Yuan and E. S. Lee.

The category FGrp of fuzzy groups is not equivalent to the category Grp of classical groups. Computers and Mathematics with Applications, 49(11–12):1953– 1956, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105002129.

Lan:2006:CFP

[yL06]  Heng you Lan.  Com-

mon fixed-point iterative processes with errors for generalized asymptotically quasi-nonexpansive mappings. Computers and Mathematics with Applications, 52(10–11):1403–1412, November/December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106003282.

Yu:2007:SVO

[YL07] Guolin Yu and Sanyang Liu. Some vector optimization problems in Banach spaces with generalized convexity. Computers and Mathematics with Applications, 54(11–12):1403–1410, December 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107004816.

Lan:2008:ASF

[yL08a] Heng you Lan. An approach for solving fuzzy implicit variational inequalities with linear membership functions. Computers and Mathematics with Applications, 55(3):563–572, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107004919.

Lan:2008:SIP

[yL08b] Heng you Lan. Stability of iterative processes with errors for a system of nonlinear (A, η)-accretive variational inclusions in Banach spaces. Computers and Mathematics with Applications, 56(1):290– 303, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000163.

Yuan:2008:NBI

[YL08c] Gonglin Yuan and Xiwen Lu. A new backtracking inexact BFGS method for symmetric nonlinear equations. Computers and Mathematics with Applications, 55(1):116–129, January 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700363X.

Yang:2009:MMM

[YL09a] Haijian Yang and Qingguo Li. A multiplicative multisplitting method for solving the linear complementarity problem. Computers and Mathematics with Applications, 58(10):1970–1978, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006051.

Yang:2009:SIM

[YL09b] Miin-Shen Yang and Der-Chen Lin. On similarity and inclusion measures between type-2 fuzzy sets with an application to clustering. Computers and Mathematics with Applications, 57(6):896–907, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005713.

Lan:2009:ASN

[yL09c]  Heng you Lan.  Ap-

proximation solvability of nonlinear random (A, η)- resolvent operator equations with random relaxed cocoercive operators. Computers and Mathematics with Applications, 57(4):624–632, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108006093.

Yoon:2006:EMC

[YLB06] Young-Cheol Yoon, Sang-Ho Lee, and T. Belytschko. Enriched meshfree collocation method with diffuse derivatives for elastic fracture. Computers and Mathematics with Applications, 51(8):1349– 1366, April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000800.

You:2008:PCC

[YLD08] Surong You, Jiajin Le, and Xiaodong Ding. Pricing a contingent claim with random interval or fuzzy random payoff in one-period setting. Computers and Mathematics with Applications, 56(8):1905–1917, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002599.

Yousefi:2009:HVI

[YLD09] S. A. Yousefi, A. Lotfi, and Mehdi Dehghan. He’s variational iteration method for solving nonlinear mixed Volterra–Fredholm integral equations. Computers and Mathematics with Applications, 58(11–12):2172–2176, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001473.

Yuan:2008:IEP

[YLL08] Shifang Yuan, Anping Liao, and Yuan Lei. Inverse eigenvalue problems of tridiagonal symmetric matrices and tridiagonal bisymmetric matrices. Computers and Mathematics with Applications, 55(11):2521– 2532, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007535.

Yang:2009:CIV

[YLY^{+}09] Xibei Yang, Tsau Young Lin, Jingyu Yang, Yan Li, and Dongjun Yu. Combination of interval-valued fuzzy set and soft set. Computers and Mathematics with Applications, 58(3):521–527, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003228.

See note [JY11].

Yang:2008:PAL

[YLZ08a] Xiao-Rong Yang, Wei-Dong Liu, and Li-Xin Zhang. Precise asymptotics in the law of logarithm under dependence assumptions. Computers and Mathematics with Applications, 56(6):1634–1642, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001909.

Yuan:2008:SVM

[YLZ08b] Xue-Hai Yuan, Hong-Xing Li, and Cheng Zhang. The set-valued mapping based on ample fields. Computers and Mathematics with Applications, 56(8):1954–1965, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002733.

Yao:2001:ART

[YLZG01] Hengshen Yao, Zhiping Li, Jianjun Zhang, and Hongjiang Ge. Application of random theory in determining the optimum producing well numbers for oil and gas reservoirs. Computers and Mathematics with Applications, 41(10–11):1465–1471, May/June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001092.

Yi:2004:IST

[YM04a] Z. Yi and D. A. Murio. Identification of source terms in 2-D IHCP. Computers and Mathematics with Applications, 47(10–11):1517–1533, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839695.

Yi:2004:STI

[YM04b] Zh. Yi and D. A. Murio.

Source term identification in 1-D IHCP. Computers and Mathematics with Applications, 47(12):1921– 1933, June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840288.

Ye:2009:HVM

[YM09] Yan-Hai Ye and Lu-Feng Mo. He’s variational method for the Benjamin– Bona–Mahony equation and the Kawahara equation. Computers and Mathematics with Applications, 58(11–12):2420–2422, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001783.

Yoneyama:2006:PEP

[YMI06] Kanji Yoneyama, Tatsuya Matsukawa, and Hiroaki Ishii. Performance evaluation of product form cellular mobile communications networks. Computers and Mathematics with Applications, 51(2):219–228, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0898122105004591.

Yu:2003:ORN

[YMMC03] Jimin Yu, Chunlai Mu, Xiaowu Mu, and Shuwei Chen. Output regulation of nonlinear singularly perturbed systems. Computers and Mathematics with Applications, 46(10– 11):1571–1579, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901921.

Yang:2005:DPH

[YMTE05] Xiaofan Yang, G. M. Megson, Yuanyan Tang, and D. J. Evans. Diameter of parallelogramic honeycomb torus. Computers and Mathematics with Applications, 50(8– 9):1477–1486, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003949.

Nie:2008:LSF

[yNyLjZaZ08] Pu yan Nie, Ming yong Lai, Shu jin Zhu, and Pei ai Zhang. A line search filter approach for the system of nonlinear equations. Computers and Mathematics with Applications, 55(9):2134– 2141, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700692X.

Yang:2006:SCC

[YO06] Dingong Yang and Shigeyoshi Owa. Sufficient conditions for Carathéodory functions. Computers and Mathematics with Applications, 51(3–4):467–474, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005134.

Yong:2009:OLR

[Yon09] Wen-An Yong. An Onsagerlike relation for the lattice Boltzmann method. Computers and Mathematics with Applications, 58(5): 862–866, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109000911.

You:2004:SBI

[You04]  Yuncheng You.  Spec-

tral barriers and inertial manifolds for time- discretized dissipative equations. Computers and Mathematics with Applications, 48(9):1351–1368, November 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003803.

Yuan:2004:SCE

[YS04] Guangwei Yuan and Longjun Shen. Stability and convergence of the explicitimplicit conservative domain decomposition procedure for parabolic problems. Computers and Mathematics with Applications, 47(4–5):793–801, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210490065X.

Yeh:2003:IAM

[YSH03] Her-Tyan Yeh, Hung-Min Sun, and Tzonelih Hwang. Improved authenticated multiplekey agreement protocol. Computers and Mathematics with Applications, 46 (2–3):207–211, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103900253.

Yang:2008:MUL

[YSP^{+}08] Guowu Yang, Xiaoyu Song, Marek A. Perkowski, William N. N. Hung, and Chang-Jun Seo. Minimal universal library for n × n reversible circuits. Computers and Mathematics with Applications, 56(1):160– 165, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008358.

Yovanof:2000:SDD

[YT00] G. S. Yovanof and H. Taylor. B_{2}-sequences and the distinct distance constant. Computers and Mathematics with Applications, 39 (11):37–42, June 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S089812210000105X.

Yang:2008:GAS

[YTC08] Xiaofan Yang, Yuan Yan Tang, and Jianqiu Cao. Global asymptotic stability of a family of difference equations. Computers and Mathematics with Applications, 56(10):2643–2649, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004033.

Young:2008:MFS

[YTCF08] D. L. Young, C. C. Tsai, C. W. Chen, and C. M. Fan. The method of fundamental solutions and condition number analysis for inverse problems of Laplace equation. Computers and Mathematics with Applications, 55(6):1189–1200, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005883.

Young:2002:SSF

[YTEC02] D. L. Young, C. C. Tsai, T. I. Eldho, and A. H.- D. Cheng. Solution of Stokes flow using an iterative DRBEM based on compactly-supported, positive-definite radial basis function. Computers and Mathematics with Applications, 43(3–5):607–619, February/March 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101003078.

Yu:2004:GEP

[Yu04] Dong Won Yu. The generalized Euler process for exponentially dominant systems. Computers and Mathematics with Applications, 47(2–3):161–181, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900144.

Yusufoglu:2009:IHP

[Yus09] Elçin Yusufoğlu. An improvement to homotopy perturbation method for solving system of linear equations. Computers and Mathematics with Applications, 58(11–12):2231–2235, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001497.

See notes [BG11].

Yang:2005:NIG

[YV05] Jie Yang and A. S. Vatsala. Numerical investigation of generalized quasilinearization method for reaction diffusion systems. Computers and Mathematics with Applications, 50(3–4):587– 598, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002713.

Yang:2003:NQF

[YW03a] Shijun Yang and Xinghua Wang. New quadrature formulas based on the zeros of the Chebyshev polynomials of the second kind. Computers and Mathematics with Applications, 45(10–11):1611–1618, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001408.

Yu:2003:FQM

[YW03b] Wenjian Yu and Zeyi Wang. A fast quasi-multiple medium method for 3-D BEM calculation of parasitic capacitance. Computers and Mathematics with Applications, 45(12):1883– 1894, June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103900095.

Yang:2007:GFP

[YW07] Chunming Yang and Peixuan Weng. Green functions and positive solutions for boundary value problems of third-order difference equations. Computers and Mathematics with Applications, 54(4):567–578, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002398.

Yang:2008:EMP

[YW08a] Guoying Yang and Mingxin Wang. Existence of multiple positive solutions for a p-Laplacian system with sign-changing weight functions. Computers and Mathematics with Applications, 55(4):636–653, February 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700497X.

Ye:2008:TMF

[YW08b] Chengfu Ye and Jianfeng Wang. Trees with mmatchings and the fourth and fifth minimal Hosoya index. Computers and Mathematics with Applications, 56(2):387–399, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122108000242.

You:2008:EFV

[YW08c] Cuilian You and Meilin Wen. The entropy of fuzzy vectors. Computers and Mathematics with Applications, 56(6):1626–1633, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001892.

Yang:2009:RRS

[YW09] Rui-Hua Yang and Shan-Yuan Wang. Resonance in a rotor-spun composite yarn spinning process obtained using the variational iteration method. Computers and Mathematics with Applications, 58(11–12):2486–2488, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001928.

Yu:2008:OCE

[YWC08] Zhi-Jian Yu, Jun-De Wu, and Min-Hyung Cho. Operation continuity of effect algebras. Computers and Mathematics with Applications, 56(8):2054–2057, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003398.

Yu:2007:EPS

[YWYL07] Shiueh-Ling Yu, Fu-Hsiang Wong, Cheh-Chih Yeh, and Shang-Wen Lin. Existence of positive solutions for n + 2 order p-Laplacian BVP. Computers and Mathematics with Applications, 53(9):1367– 1379, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000995.

Yu:2003:IDM

[YWZ03] Yumei Yu, Wendi Wang, and Yong Zhang. An innovation diffusion model for three competitive products. Computers and Mathematics with Applications, 46 (10–11):1473–1481, November/December 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103901842.

Yang:2007:RSU

[YX07] Zhichun Yang and Daoyi Xu. Robust stability of uncertain impulsive control systems with timevarying delay. Computers and Mathematics with Applications, 53(5):760–769, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000296.

Yang:2009:TDP

[YX09a] Shouzhi Yang and Yanmei Xue. Two-direction polyscale refinability. Computers and Mathematics with Applications, 58(1):119– 127, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109002612.

Yu:2009:TBA

[YX09b] Long Yu and Jian Xiao. Trade-off between accuracy and interpretability: Experience-oriented fuzzy modeling via reducedset vectors. Computers and Mathematics with Applications, 57(6):885–895, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005725.

Yuan:2009:GRG

[YXZ09] Wenjun Yuan, Bing Xiao, and Jianjun Zhang. The general result of Gol’dberg’s theorem concerning the growth of meromorphic solutions of algebraic differential equations. Computers and Mathematics with Applications, 58(9):1788–1791, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006026.

Yu:2009:DVP

[YXZL09] Yong Yu, Chunxiang Xu, Xiaosong Zhang, and Yongjian Liao. Designated verifier proxy signature scheme without random oracles. Computers and Mathematics with Applications, 57 (8):1352–1364, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S089812210900039X.

Yang:2007:FDA

[YY07] Hui Yang and Xiaofan Yang.

A fast diagnosis algorithm for locally twisted cube multiprocessor systems under the MM ^{∗} model. Computers and Mathematics with Applications, 53(6):918–926, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000442.

Yoshida:2003:MFS

[YYNK03] Y. Yoshida, M. Yasuda, J. Nakagami, and M. Kurano. A multiobjective fuzzy stopping in a stochastic and fuzzy environment. Computers and Mathematics with Applications, 46(7):1165–1172, October 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103901325.

Yang:2003:ICO

[YYZ03] Qigui Yang, Lijun Yang, and Siming Zhu. Interval criteria for oscillation of second-order nonlinear neutral differential equations. Computers and Mathematics with Applications, 46(5–6):903–918, September 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/S0898122103901520.

See comment [Liu08b].

Yao:2007:AMC

[YYZ07] Yonghong Yao, Jen-Chih Yao, and Haiyun Zhou. Approximation methods for common fixed points of infinite countable family of nonexpansive mappings. Computers and Mathematics with Applications, 53(9):1380–1389, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001009.

Yan:2002:FFO

[YZ02a] Guozheng Yan and Huijiang Zhao. The far field operator for a multilayered scatterer. Computers and Mathematics with Applications, 43(6–7):631–639, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003091.

Yan:2002:CFS

[YZ02b] Zhen-Ya Yan and Hong-Qing Zhang. Constructing families of soliton-like solutions to a (2 + 1)- dimensional breaking soliton equation using symbolic computation. Computers and Mathematics with Applications, 44(10– 11):1439–1444, November/ December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102002687.

Yang:2009:BLP

[YZHY09] Jun Yang, Min Zhang, Bo He, and Chao Yang. Bi-level programming model and hybrid genetic algorithm for flow interception problem with customer choice. Computers and Mathematics with Applications, 57(11–12):1985– 1994, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005233.

Yang:2009:NSP

[YZLL09] Junjie Yang, Jianzhong Zhou, Li Liu, and Yinghai Li. A novel strategy of pareto-optimal solution searching in multi-objective particle swarm optimization (MOPSO). Computers and Mathematics with Applications, 57(11–12):1995– 2000, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800521X.

Yu:2009:AFS

[YZWL09] Jian Yu, Shuigeng Zhou, Lipo Wang, and Jingsheng Lei. Advances in fuzzy sets and knowledge discovery. Computers and Mathematics with Applications, 57(6):865–866, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005762.

Yuan:2008:EDE

[YZWY08] Xiaohui Yuan, Yongchuan Zhang, Liang Wang, and Yanbin Yuan. An enhanced differential evolution algorithm for daily optimal hydro generation scheduling. Computers and Mathematics with Applications, 55(11):2458–2468, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007018.

Yu:2008:ICL

[YZZ^{+}08] Yongbin Yu, Fengli Zhang, Qishui Zhong, Xiaofeng Liao, and Juebang Yu. Impulsive control of Lurie systems. Computers and Mathematics with Applications, 56(11):2806–2813, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800494X.

Yong:2009:SCJ

[YZZC09] Xuelin Yong, Xin Zeng, Zhiyong Zhang, and Yufu Chen. Symbolic computation of Jacobi elliptic function solutions to nonlinear differential-difference equations. Computers and Mathematics with Applications, 57(7):1107–1114, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000248.

Zhang:2003:OSD

[ZA03a] B. G. Zhang and R. P. Agarwal. The oscillation and stability of delay partial difference equations. Computers and Mathematics with Applications, 45(6–9):1253–1295, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103000993.

Zhang:2003:CMA

[ZA03b] Weinian Zhang and R. P. Agarwal. Construction of mappings with attracting cycles. Computers and Mathematics with Applications, 45(6–9):1213–1219, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000889.

Zahran:2009:CDF

[ZAAESG09] A. M. Zahran, M. Azab Abd-Allah, Kamal El-Saady, and A. Ghareeb. The category of double fuzzy preproximity spaces. Computers and Mathematics with Applications, 58(8):1558–1572, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109005069.

Zalmai:2002:ECD

[Zal02] G. J. Zalmai. Efficiency conditions and duality models for multiobjective fractional subset programming problems with generalized ([digamma], α, ρ, θ)-V - convex functions. Computers and Mathematics with Applications, 43(12):1489– 1520, June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001141.

Zalmai:2003:PFS

[Zal03] G. J. Zalmai. Parameterfree sufficient optimality conditions and duality models for minmax fractional subset programming problems with generalized (F ρ, θ)-convex functions. Computers and Mathematics with Applications, 45(10–11):1507–1535, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001342.

Zeng:2008:GIA

[ZAY08] Lu-Chuan Zeng, Qamrul Hasan Ansari, and Jen-Chih Yao. General iterative algorithms for solving mixed quasi-variational-like inclusions. Computers and Mathematics with Applications, 56(10):2455–2467, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003787.

Zreikat:2003:PMN

[ZBS03] A. I. Zreikat, G. Bolch, and J. Sztrik. Performance modelling of nonhomogeneous unreliable multiserver systems using MOSEL. Computers and Mathematics with Applications, 46(2–3): 293–312, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122103900320.

Zhu:2001:EGO

[ZC01] Dongmei Zhu and Shaozhu Chen. Exponential growth and oscillation of nonlinear homogeneous difference equations with periodic coefficients. Computers and Mathematics with Applications, 41(3–4):415–422, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002844.

Zeng:2006:ABF

[ZC06] Xiao-Ming Zeng and X. J.

Chen. Approximation for bounded functions and absolutely continuous functions by beta operators. Computers and Mathematics with Applications, 51(9– 10):1585–1592, May 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122106001039.

Zhu:2008:SAM

[ZC08] Zhi-Qiang Zhu and Sui Sun Cheng. Stability analysis for multistep computational schemes. Computers and Mathematics with Applications, 55(12):2753– 2761, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107007754.

Zhao:2009:TPS

[ZC09a] Yulin Zhao and Haibo Chen. Triple positive solutions for nonlinear boundary value problems in Banach space. Computers and Mathematics with Applications, 58(9):1780–1787, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109005811.

Zhao:2009:FPD

[ZC09b] Zengqin Zhao and Xiangping Chen. Fixed points of decreasing operators in ordered Banach spaces and applications to nonlinear second order elliptic equations. Computers and Mathematics with Applications, 58(6):1223–1229, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109005057.

Zhuang:2004:RPB

[ZCC04] Wan Zhuang, Yubo Chen, and Jie Chen. Remarks on the periodic boundary value problems for integrodifferential equations of Volterra type. Computers and Mathematics with Applications, 47(10–11):1641–1646, May/June 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104839774.

Zanivan:2008:VDV

[ZCCB08] Sara Zanivan, Davide Corà, Michele Caselle, and Federico Bussolino. VRG: a database of vascular dysfunctions related genes. Computers and Mathematics with Applications, 55 (5):1068–1073, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122107005573.

Zhou:2000:CCC

[ZCK00] Haiyun Zhou, Yeol Je Cho, and Shin Min Kang. Characteristic conditions for convergence of generalized steepest descent approximation to multivalued accretive operator equations. Computers and Mathematics with Applications, 39(3–4):1–11, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003284.

Zhu:2005:ECC

[ZCK05] Jiang Zhu, Yeol Je Cho, and Shin Min Kang. Equivalent contractive conditions in symmetric spaces. Computers and Mathematics with Applications, 50(10– 12):1621–1628, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105004128.

Zhang:2008:EMF

[ZCL08] Xiao-Yu Zhang, Jun-Fan Chen, and Wei-Chuan Lin. Entire or meromorphic functions sharing one value. Computers and Mathematics with Applications, 56(7):1876–1883, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002678.

Zhan:2001:ANS

[ZCM01a] Shenghe Zhan, C. Coles, and D. A. Murio. Automatic numerical solution of generalized 2-D IHCP by discrete mollification. Computers and Mathematics with Applications, 41(1–2):15–38, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101850033.

Ziadi:2001:GON

[ZCM01b] A. Ziadi, Y. Cherruault, and G. Mora. Global optimization: a new variant of the alienor method. Computers and Mathematics with Applications, 41(1–2):63–71, January 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101850069.

Zeng:2004:NOT

[ZCS04] Guoping Zeng, I. Chlamtac, and Yi Su. Nonlinear observers for two classes of perturbed systems. Computers and Mathematics with Applications, 48(3–4):387– 398, August 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104840628.

Zhu:2006:MPS

[ZCW06] Yuhong Zhu, Hexin Chen, and Shigang Wang. Multiple positive solutions to superlinear repulsive singular impulsive differential equations. Computers and Mathematics with Applications, 52(6–7):1031–1044, September/October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106002768.

Zhang:2001:OSS

[ZCZ01] Zhenguo Zhang, Jianfeng Chen, and Caishun Zhang. Oscillation of solutions for second-order nonlinear difference equations with nonlinear neutral term. Computers and Mathematics with Applications, 41(12):1487–1494, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001134.

Zhang:2001:SCP

[ZD01] Bing Gen Zhang and Xing Hua Deng. The stability of certain partial difference equations. Computers and Mathematics with Applications, 42(3–5):419–425, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001675.

Zhang:2003:SPD

[ZD03] Binggen Zhang and Xinghua Deng. The stability of partial difference systems with retarded arguments.

Computers and Mathematics with Applications, 45 (6–9):1227–1233, March/ May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000920.

Zhu:2008:MSD

[ZDLH08] Ming-Fang Zhu, Ting Dai, Sung-Yoon Lee, and Chun-Pyo Hong. Modeling of solutal dendritic growth with melt convection. Computers and Mathematics with Applications, 55(7):1620– 1628, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107006219.

Zhan:2008:GFH

[ZDS08] Jianming Zhan, Bijan Davvaz, and K. P. Shum. Generalized fuzzy hyperideals of hyperrings. Computers and Mathematics with Applications, 56(7):1732–1740, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002514.

Zeng:2000:RAB

[Zen00] Xiao-Ming Zeng. Rates of approximation of bounded variation functions by two generalized Meyer-König and zeller type operators. Computers and Mathematics with Applications, 39(9– 10):1–13, May 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100000821.

Zeng:2009:EMP

[Zen09] Zhijun Zeng. Existence and multiplicity of positive periodic solutions for a class of higher-dimension functional differential equations with impulses. Computers and Mathematics with Applications, 58(10):1911–1920, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006075.

Zhang:2005:ENS

[ZFYS05] Weiping Zhang, Wei Feng, Juran Yan, and Jinsheng Song. Existence of nonoscillatory solutions of firstorder linear neutral delay differential equations. Computers and Mathematics with Applications, 49(7–8):1021–1027, May/ June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001185.

Zeng:2002:RCB

[ZG02] Xiao-Ming Zeng and V. Gupta.

Rate of convergence of Baskakov–Bézier type operators for locally bounded functions. Computers and Mathematics with Applications, 44(10–11):1445–1453, November/December 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102002699.

Zhang:2009:PSCb

[ZG09a] Xuemei Zhang and Weigao Ge. Positive solutions for a class of boundaryvalue problems with integral boundary conditions. Computers and Mathematics with Applications, 58(2): 203–215, July 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109002788.

Zhao:2009:PSH

[ZG09b] Junfang Zhao and Weigao Ge. Positive solutions for a higher-order fourpoint boundary value problem with a p-Laplacian. Computers and Mathematics with Applications, 58(6): 1103–1112, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122109004908.

Gu:2008:INS

[zGtM08] Neng zhu Gu and Jiang tao Mo. Incorporating nonmonotone strategies into the trust region method for unconstrained optimization. Computers and Mathematics with Applications, 55(9):2158– 2172, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006967.

Zeng:2005:CMO

[ZGY05a] Lu-Chuan Zeng, Sy-Ming Guu, and Jen-Chih Yao. Characterization of Hmonotone operators with applications to variational inclusions. Computers and Mathematics with Applications, 50(3–4):329–337, August 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105002774.

Zeng:2005:IAC

[ZGY05b] Lu-Chuan Zeng, Sy-Ming Guu, and Jen-Chih Yao. Iterative algorithm for completely generalized setvalued strongly nonlinear mixed variational-like inequalities. Computers and Mathematics with Applications, 50(5–6):935–945, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003263.

Zeng:2007:IMG

[ZGY07a] Lu-Chuan Zeng, Sy-Ming Guu, and Jen-Chih Yao. An iterative method for generalized nonlinear set-valued mixed quasi-variational inequalities with H-monotone mappings. Computers and Mathematics with Applications, 54(4):476–483, August 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002301.

Zeng:2007:TSI

[ZGY07b] Lu-Chuan Zeng, Sy-Ming Guu, and Jen-Chih Yao. Three-step iterative algorithms for solving the system of generalized mixed quasi-variational-like inclusions. Computers and Mathematics with Applications, 53(10):1572–1581, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001290.

Zhou:2002:PSC

[ZH02] Yuying Zhou and Yisheng Huang. Positive solutions for a class of elliptic equations. Computers and Mathematics with Applications, 43(10–11):1271–1276, May/June 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102000998.

Zhu:2004:DCN

[ZH04] Huiyan Zhu and Lihong Huang. Dynamics of a class of nonlinear discrete-time neural networks. Computers and Mathematics with Applications, 48(1–2):85– 94, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104840008.

Zhang:2007:RSY

[ZH07] Li-Na Zhang and Ji-Huan He. Resonance in sirospun yarn spinning using a variational iteration method. Computers and Mathematics with Applications, 54(7– 8):1064–1066, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0898122107002763.

Zhang:2009:NAD

[ZH09] Wenxing Zhang and Deren Han. A new alternating direction method for cocoercive variational inequality problems. Computers and Mathematics with Applications, 57(7):1168– 1178, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006664.

Zhang:2001:GEA

[Zha01a] Hui Zhang. Global existence and asymptotic behaviour of the solution of a generalized burger’s equation with viscosity. Computers and Mathematics with Applications, 41(5–6):589– 596, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100003023.

Zhang:2001:BND

[Zha01b] Shunian Zhang. Boundedness of neutral delay difference systems. Computers and Mathematics with Applications, 42(10– 11):1255–1265, November/ December 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101002371.

Zhang:2001:SND

[Zha01c] Shunian Zhang. Stability of neutral delay difference systems. Computers and Mathematics with Applications, 42(3–5):291–299, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001547.

Zhang:2001:MPA

[Zha01d] Yongmin Zhang. Multilevel projection algorithm for solving obstacle problems. Computers and Mathematics with Applications, 41(12):1505–1513, June 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122101001158.

Zhang:2002:IIN

[Zha02a] Jianna J. Zhang. Introducing IVSA: a new concept learning algorithm. Computers and Mathematics with Applications, 43(6–7):821–832, March/ April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003236.

Zhang:2002:NTS

[Zha02b] Shunian Zhang. A new technique in stability of infinite delay differential equations. Computers and Mathematics with Applications, 44 (10–11):1275–1287, November/December 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122102002559.

Zhao:2002:ETJ

[Zha02c] Cheng-Dong Zhao. Existence of trajectories joining critical points for a symbiotic system via the Conley index. Computers and Mathematics with Applications, 43(6–7):755–766, March/April 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101003194.

Zhang:2003:GSA

[Zha03a] Jiye Zhang. Global stability analysis in delayed cellular neural networks. Computers and Mathematics with Applications, 45 (10–11):1707–1720, May/ June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122103001494.

Zhang:2003:NSS

[Zha03b] Jun Zhang. Numerical simulation of 2D square driven cavity using fourth-order compact finite difference schemes. Computers and Mathematics with Applications, 45(1–3):43–52, January/February 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122103800068.

Zhang:2004:NSG

[Zha04a] Chengjian Zhang. NGP(α)- stability of general linear methods for NDDEs. Computers and Mathematics with Applications, 47 (6–7):1105–1113, March/ April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104900909.

Zhang:2004:ASA

[Zha04b] Jiye Zhang. Absolute stability analysis in cellular neural networks with variable delays and unbounded delay. Computers and Mathematics with Applications, 47(2–3):183–194, January/ February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900156.

Zhang:2006:CMO

[Zha06]  Zhuhong Zhang.  Con-

strained multiobjective optimization immune algorithm: Convergence and application. Computers and Mathematics with Applications, 52(5):791–808, September 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002574.

Zhang:2007:VAS

[Zha07a]  Juan Zhang.  Varia-

tional approach to solitary wave solution of the generalized Zakharov equation. Computers and Mathematics with Applications, 54(7–8):1043–1046, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107002726.

Zhang:2007:GAE

[Zha07b] Sheng Zhang. A generalized auxiliary equation method and its application to (2 + 1)-dimensional Korteweg–de Vries equations. Computers and Mathematics with Applications, 54(7–8):1028–1038, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002684.

Zhang:2008:SPR

[Zha08a] Jiancheng Zhang. Some properties of the roots of theories in propositional logic systems. Computers and Mathematics with Applications, 55(9):2086– 2093, May 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107006888.

Zhang:2008:UTE

[Zha08b]  Jilong Zhang.  Unique-

ness theorems for entire functions concerning fixed points. Computers and Mathematics with Applications, 56(12):3079–3087, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004355.

Zhang:2008:GPR

[Zha08c] Yuan Lin Zhang. A geometrical process repair model for a repairable system with delayed repair. Computers and Mathematics with Applications, 55(8):1629– 1643, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107005974.

Zhao:2008:SGF

[Zha08d]  Zengqin Zhao.  Solu-

tions and Green’s functions for some linear secondorder three-point boundary value problems. Computers and Mathematics with Applications, 56(1):104– 113, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008309.

Zhang:2009:AHA

[Zha09a] Hui-Li Zhang. Application of He’s amplitude-frequency formulation to a nonlinear oscillator with discontinuity. Computers and Mathematics with Applications, 58(11–12):2197–2198, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001539.

Zhang:2009:PSS

[Zha09b]  Hui-Li Zhang.  Peri-

odic solutions for some strongly nonlinear oscillations by He’s energy balance method. Computers and Mathematics with Applications, 58(11–12):2480–2485, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001898.

Zhang:2009:NSF

[Zha09c] Juan Zhang. The numerical solution of fifth-order boundary value problems by the variational iteration method. Computers and Mathematics with Applications, 58(11–12):2347–2350, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001692.

Zhang:2009:SIP

[Zha09d] Xingqiu Zhang. Successive iteration and positive solutions for a secondorder multi-point boundary value problem on a half-line. Computers and Mathematics with Applications, 58(3):528–535, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109003290.

Zhang:2009:DFD

[Zha09e] Yang Zhang. A discontinuous finite difference streamline diffusion method for time-dependent hyperbolic problems. Computers and Mathematics with Applications, 58(10):1901–1910, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109006105.

Zhao:2009:DTC

[Zha09f]  Bin Zhao.  Dynamics

of two classes of recursive sequences. Computers and Mathematics with Applications, 58(2):329– 333, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900279X.

Zhao:2009:HFA

[Zha09g] Ling Zhao. He’s frequencyamplitude formulation for nonlinear oscillators with an irrational force. Computers and Mathematics with Applications, 58(11–12):2477– 2479, December 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109001886.

Zhou:2005:ENN

[ZHKW05] Qian Zhou, Rui Huang, Yuanyuan Ke, and Yifu Wang. Existence of the nontrivial nonnegative periodic solutions for the quasilinear parabolic equation with nonlocal terms. Computers and Mathematics with Applications, 50 (8–9):1293–1302, October/ November 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003810.

Zhu:2004:CPS

[ZHL04] Huiyan Zhu, Lihong Huang, and Xinyuan Liao. Convergence and periodicity of solutions for a class of delay difference equations. Computers and Mathematics with Applications, 48(10– 11):1477–1484, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210400402X.

Zhang:2005:MIM

[ZHL05] Yong Zhang, Ting-Zhu Huang, and Xing-Ping Liu. Modified iterative methods for nonnegative matrices and M-matrices linear systems. Computers and Mathematics with Applications, 50(10– 12):1587–1602, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004104.

Zhou:2001:OHO

[Zho01]  Yong Zhou.  Oscilla-

tions of higher-order linear difference equations. Computers and Mathematics with Applications, 42(3–5):323–331, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001572.

Zhou:2003:POD

[Zho03] Zhan Zhou. Periodic orbits on discrete dynamical systems. Computers and Mathematics with Applications, 45(6–9):1155–1161, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000750.

Zhou:2006:SRA

[Zho06]  Y. Y. Zhou.  Some

results about perturbed maximal monotone mappings. Computers and Mathematics with Applications, 51(3–4):487–496, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105005158.

Zhou:2007:VTP

[Zho07]  Xin-Wei Zhou.  Varia-

tional theory for physiological flow. Computers and Mathematics with Applications, 54(7–8):1000–1002, October 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002611.

Zhou:2008:RSF

[Zho08] Zhan Zhou. Remarks on “Stability for first order delay dynamic equations on time scales” [Comput. Math. Appl. 53 (2007) 1820–1831]. Computers and Mathematics with Applications, 56(4):1155–1156, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108001478.

See [WZ07c] and errata

[KÖ08b].

Zhu:2005:MNM

[Zhu05] Ling Zhu. A modified Newton method in parallel circular iteration of singlestep and double-step. Computers and Mathematics with Applications, 50(10– 12):1513–1524, November/ December 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004062.

Zhu:2008:GRJ

[Zhu08a] Ling Zhu. A general refinement of Jordan-type inequality. Computers and Mathematics with Applications, 55(11):2498–2505, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700750X.

Zhu:2008:IAS

[Zhu08b] Wei Zhu. Invariant and attracting sets of impulsive delay difference equations with continuous variables. Computers and Mathematics with Applications, 55(12):2732–2739, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007729.

Zhu:2009:SNI

[Zhu09a]  Ling Zhu.  Some new

inequalities of the Huygens type. Computers and Mathematics with Applications, 58(6):1180–1182, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210900488X.

Zhu:2009:SIC

[Zhu09b] Ling Zhu. A source of inequalities for circular functions. Computers and Mathematics with Applications, 58(10):1998–2004, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109005999.

Zhang:2005:DTF

[ZhYL05] Cheng Zhang, Xue hai Yuan, and E. Stanley Lee. Duality theory in fuzzy mathematical programming problems with fuzzy coefficients. Computers and Mathematics with Applications, 49(11–12):1709– 1730, June 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122105001914.

Zhang:2000:WCB

[ZHZ00] Weinian Zhang, Xiaorong Hou, and Zhenbing Zeng. Weak centers and bifurcation of critical periods in reversible cubic systems. Computers and Mathematics with Applications, 40 (6–7):771–782, September/ October 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100001954.

Zang:2006:NDL

[ZHZT06] Hong Zang, Maoan Han, Tonghua Zhang, and M. O. Tadé. The number and distributions of limit cycles for a class of quintic near-Hamiltonian systems. Computers and Mathematics with Applications, 52(10– 11):1577–1594, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106003415.

Zhang:2008:ULS

[ZJ08] Qiumei Zhang and Daqing Jiang. Upper and lower solutions method and a second order three-point singular boundary value problem. Computers and Mathematics with Applications, 56(4):1059–1070, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001089.

Zhang:2006:NET

[ZJLW06] Xiaoying Zhang, Daqing Jiang, Xiaoyue Li, and Ke Wang. A new existence theory for single and multiple positive periodic solutions to Volterra integrodifferential equations with impulse effects. Computers and Mathematics with Applications, 51(1):17–32, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210500444X.

Zhang:2002:SBM

[ZL02] C. J. Zhang and X. X. Liao. Stability of BDF methods for nonlinear Volterra integral equations with delay. Computers and Mathematics with Applications, 43(1–2):95–102, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002759.

Zhang:2003:PSN

[ZL03] Shunian Zhang and Wan-Tong Li. Periodic solutions of neutral difference equations. Computers and Mathematics with Applications, 45(6–9):1245–1251, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000968.

Zhang:2004:ROS

[ZL04] Zhengce Zhang and Kaitai Li. Radial oscillatory solutions of some quasilinear elliptic equations. Computers and Mathematics with Applications, 47(8–9):1327–1334, April/ May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901265.

Zhu:2005:SIR

[ZL05] Yuanguo Zhu and Baoding Liu. Some inequalities of random fuzzy variables with application to moment convergence. Computers and Mathematics with Applications, 50(5–6):719–727, September 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105003123.

Zhang:2006:VWS

[ZL06] Yu-Zhuo Zhang and Hong-Xing Li. Variable weighted synthesis inference method for fuzzy reasoning and fuzzy systems. Computers and Mathematics with Applications, 52(3–4):305– 322, August 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001982.

Zhang:2007:QPD

[ZL07a] Qi Zhang and Yirong Liu. A quintic polynomial differential system with eleven limit cycles at the infinity. Computers and Mathematics with Applications, 53(10):1518– 1526, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107001204.

Zhao:2007:RTM

[ZL07b] Zhihong Zhao and Yongjin Li. Reverse triple I method of fuzzy reasoning for the implication operator R_{L}. Computers and Mathematics with Applications, 53(7):1020– 1028, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107000570.

Zou:2007:PBL

[ZL07c] Bin Zou and Luoqing Li. The performance bounds of learning machines based on exponentially strongly mixing sequences. Computers and Mathematics with Applications, 53(7):1050– 1058, April 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000624.

Zeng:2008:ECM

[ZL08a] Xiao-Ming Zeng and Bo-Yong Lian. An estimate on the convergence of MKZ-Bézier operators. Computers and Mathematics with Applications, 56(12):3023–3028, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004471.

Zhang:2008:UTM

[ZL08b] Tongdui Zhang and Weiran Lü. Uniqueness theorems on meromorphic functions sharing one value. Computers and Mathematics with Applications, 55(12):2981– 2992, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008206.

Zhang:2008:EFO

[ZL08c] Xinguang Zhang and Lishan Liu. Eigenvalue of fourth-order m-point boundary value problem with derivatives. Computers and Mathematics with Applications, 56(1):172– 185, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000035.

Zhang:2008:NSH

[ZL08d] Xinguang Zhang and Lishan Liu. Nontrivial solutions for higher order multi-point boundary value problems. Computers and Mathematics with Applications, 56(4):861–873, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108000874.

Zhang:2009:ICC

[ZL09a] Qun-Jiao Zhang and Jun-An Lu. Impulsively control complex networks with different dynamical nodes to its trivial equilibrium. Computers and Mathematics with Applications, 57(7):1073–1079, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109000352.

Zhang:2009:RSU

[ZL09b] Zhigang Zhang and Xinzhi Liu. Robust stability of uncertain discrete impulsive switching systems. Computers and Mathematics with Applications, 58(2):380– 389, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109002569.

Zheng:2009:HVM

[ZL09c] Xiangwei Zheng and Hong Liu. A hybrid vertical mutation and self-adaptation based MOPSO. Computers and Mathematics with Applications, 57(11–12):2030– 2038, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005154.

Zlatev:2001:POS

[Zla01] Z. Zlatev. Partitioning ODE systems with an application to air pollution models. Computers and Mathematics with Applications, 42(6–7):817–832, September/October 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002012.

Zhang:2008:RUU

[ZLCM08] Liguo Zhang, Ying Lu, Yangzhou Chen, and Nikos E. Mastorakis. Robust uniformly ultimate boundedness control for uncertain switched linear systems. Computers and Mathematics with Applications, 56(7):1709–1714, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002484.

Zhou:2008:CGB

[ZLD08] Bin Zhou, James Lam, and Guang-Ren Duan. Convergence of gradient-based iterative solution of coupled Markovian jump Lyapunov equations. Computers and Mathematics with Applications, 56(12):3070–3078, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004367.

Zhang:2007:EES

[ZLH07] Fuxing Zhang, Bingwen Liu, and Lihong Huang. Existence and exponential stability of periodic solutions for a class of Cohen– Grossberg neural networks with bounded and unbounded delays. Computers and Mathematics with Applications, 53(9):1325– 1338, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000946.

Zhang:2006:MPS

[ZLJW06] Xiaoying Zhang, Xiaoyue Li, Daqing Jiang, and Ke Wang. Multiplicity positive solutions to periodic problems for first-order impulsive differential equations. Computers and Mathematics with Applications, 52(6–7):953–966, September/October 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106002690.

Zhou:2008:GEP

[ZLLL08] Tiejun Zhou, Yuehua Liu, Xiaoping Li, and Yirong Liu. Global exponential periodicity for BAM neural network with periodic coefficients and continuously distributed delays. Computers and Mathematics with Applications, 55(11):2689– 2698, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007687.

Zheng:2006:IFI

[ZLN06] Yang Zheng, Guanrong Liu, and Xiaoxiao Niu. An improved fractal image compression approach by using iterated function system and genetic algorithm. Computers and Mathematics with Applications, 51(11):1727– 1740, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106001398.

Zhang:2006:ESA

[ZLS06] Cong-Jun Zhang, Jinlu Li, and Min Sun. Existence of solutions for abstract economic equilibrium problems and algorithm. Computers and Mathematics with Applications, 52(10– 11):1471–1482, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003348.

Zhang:2006:NBV

[ZLY06] Fengqin Zhang, Meili Li, and Jurang Yan. Nonhomogeneous boundary value problem for first-order impulsive differential equations with delay. Computers and Mathematics with Applications, 51(6–7):927–936, March/April 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106000277.

Zhang:2009:WWG

[ZLY09] Yuanping Zhang, Xiaogang Liu, and Xuerong Yong.

Which wheel graphs are determined by their Laplacian spectra? Computers and Mathematics with Applications, 58(10):1887–1890, November 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004635.

Zhang:2009:PSCa

[ZLZ09] Liang Zhang, Hong-Xu Li, and Xiao-Bing Zhang. Periodic solutions of competition Lotka–Volterra dynamic system on time scales. Computers and Mathematics with Applications, 57(7):1204–1211, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000418.

Zadrozny:1994:KFM

[ZM94] P. A. Zadrozny and S. Mittnik. Kalman-filtering methods for computing information matrices for time-invariant, periodic, and generally time-varying VARMA models and samples. Computers and Mathematics with Applications, 28(4):107–119, August 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/0898122194001324.

See comments [Ter00].

Zhen:2002:PSD

[ZM02] Jin Zhen and Zhien Ma. Periodic solutions for delay differential equations model of plankton allelopathy. Computers and Mathematics with Applications, 44(3–4):491–500, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001633.

Zhang:2004:TPB

[ZM04a] Guang Zhang and R. Medina. Three-point boundary value problems for difference equations. Computers and Mathematics with Applications, 48(12):1791–1799, December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122104004390.

Zhang:2004:IIS

[ZM04b] Wen-Xiu Zhang and Ju-Sheng Mi. Incomplete information system and its optimal selections. Computers and Mathematics with Applications, 48(5–6):691–698, September 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104003049.

Zhang:2009:EDA

[ZMLG09] Zhen Zhang, Siliang Ma, Hui Liu, and Yuexin Gong. An edge detection approach based on directional wavelet transform. Computers and Mathematics with Applications, 57(8):1265– 1271, April 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109000492.

Zhu:2007:FFD

[ZMM07] Huaning Zhu, Morteza M.

Mehrabadi, and Mehrdad Massoudi. The frictional flow of a dense granular material based on the dilatant double shearing model. Computers and Mathematics with Applications, 53(2):244–259, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001034.

Zhang:2001:LOS

[ZP01] Zhenguo Zhang and Bi Ping.

Linear oscillation of secondorder nonlinear difference equations with damped term. Computers and Mathematics with Applications, 41(5–6):659–667, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100003096.

Zegeye:2002:SCA

[ZP02] H. Zegeye and E. Prempeh. Strong convergence of approximants to fixed points of Lipschitzian pseudocontractive maps. Computers and Mathematics with Applications, 44(3–4):339– 346, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001529.

Zhao:2000:AKE

[ZQ00] Ping Fu Zhao and Meng Zhao Qin. Approximations for KdV equation as a Hamiltonian system. Computers and Mathematics with Applications, 39(5–6):1–11, March 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122100000407.

Zheng:2008:EBL

[ZQC08] Zhaowen Zheng, Jiangang Qi, and Shaozhu Chen. Eigenvalues below the lower bound of minimal operators of singular Hamiltonian expressions. Computers and Mathematics with Applications, 56(11):2825–2833, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004562.

Zdunek:2005:TAS

[ZRS05] A. Zdunek, W. Rachowicz, and N. Sehlstedt. Toward hp-adaptive solution of 3D electromagnetic scattering from cavities. Computers and Mathematics with Applications, 49(1):23–38, January 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000040.

Zhang:2004:PSI

[ZS04] Yu Zhang and Jitao Sun.

Practical stability of impulsive functional differential equations in terms of two measurements. Computers and Mathematics with Applications, 48(10– 11):1549–1556, November/ December 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104004092.

Zhang:2005:OSO

[ZS05] B. G. Zhang and Zhu Shanliang. Oscillation of second-order nonlinear delay dynamic equations on time scales. Computers and Mathematics with Applications, 49(4):599–609, February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000532.

Zhang:2007:TNS

[ZS07] Le Zhang and Andras Z. Szeri. Transportation of neutral solute in deformable, anisotropic, soft tissue. Computers and Mathematics with Applications, 53(2):232–243, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000983.

Zhang:2008:EGA

[ZS08a] Lijuan Zhang and Ligeng Si. Existence and global attractivity of almost periodic solution for DCNNs with time-varying coefficients. Computers and Mathematics with Applications, 55(8):1887–1894, April 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107007857.

Zhu:2008:SNR

[ZS08b] Ling Zhu and Jinju Sun. Six new Redheffer-type inequalities for circular and hyperbolic functions. Computers and Mathematics with Applications, 56(2):522– 529, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000813.

Zhang:2001:RRS

[ZSG01] D. C. Zhang, B. Shi, and M. J. Gai. A rational recursive sequence. Computers and Mathematics with Applications, 41(3–4):301–306, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100002741.

Zeng:2004:DPL

[ZSG04] X. Y. Zeng, B. Shi, and M. J. Gai. A discrete periodic Lotka– Volterra system with delays. Computers and Mathematics with Applications, 47(4–5):491–500, February/ March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104900405.

Zhang:2003:CAS

[ZSKZ03] D. C. Zhang, B. Shi, C. Kent, and X. Y. Zeng. Counterexamples to the article “Stability properties of nonlinear difference equations and conditions for boundedness”. Computers and Mathematics with Applications, 46(10– 11):1735–1739, November/ December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103902070.

Zhang:2008:APB

[ZSPZ08] Zhonghua Zhang, Yaohong Suo, Jigeng Peng, and Juan Zhang. Analysis of a periodic bacteria-immunity model with delayed quorum sensing. Computers and Mathematics with Applications, 56(11):2834–2847, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004847.

Zhao:2006:EPS

[ZSZ06] Z. J. Zhao, Q. K. Song, and J. Y. Zhang. Exponential periodicity and stability of neural networks with reaction-diffusion terms and both variable and unbounded delays. Computers and Mathematics with Applications, 51(3–4):475–486, February 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105005146.

Zhang:2004:OCC

[ZT04] B. G. Zhang and C. J. Tian.

Oscillation criteria of a class of partial difference equations with delays. Computers and Mathematics with Applications, 48(1–2):291– 303, July 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210484015X.

Zhou:2007:PSK

[ZT07] Yinggao Zhou and Xianhua Tang. Periodic solutions for a kind of Rayleigh equation with a deviating argument. Computers and Mathematics with Applications, 53(5):825–830, March 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107000351.

Zhao:2008:PAC

[ZT08] Yuexu Zhao and Jingxuan Tao. Precise asymptotics in complete moment convergence for self-normalized sums. Computers and Mathematics with Applications, 56(7):1779–1786, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108002575.

Zhang:2009:NFC

[ZTH09] Yong Zhang, Gong-You Tang, and Nai-Ping Hu. Non-fragile control for nonlinear networked control systems with long timedelay. Computers and Mathematics with Applications, 57(10):1630–1637, May 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210900203X.

Zhang:2009:EFM

[ZTW09] Sheng Zhang, Jing-Lin Tong, and Wei Wang. Expfunction method for a nonlinear ordinary differential equation and new exact solutions of the dispersive long wave equations. Computers and Mathematics with Applications, 58(11–12):2294–2299, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109001576.

Zhang:2003:NTD

[ZW03a] Bao-Lin Zhang and Zheng-Su Wan. New techniques in designing finitedifference domain decomposition algorithm for the heat equation. Computers and Mathematics with Applications, 45(10–11):1695–1705, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103801268.

Zhou:2003:SPO

[ZW03b] Zhan Zhou and Jianhong Wu. Stable periodic orbits in nonlinear discrete-time neural networks with delayed feedback. Computers and Mathematics with Applications, 45(6–9):935–942, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210300066X.

Zhang:2004:LTB

[ZW04] Hui Zhang and Xiaohong Wang. Large-time behavior of smooth solutions to a nonuniformly parabolic equation. Computers and Mathematics with Applications, 47(2–3):353–363, January/February 2004. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122104900302.

Zheng:2006:BRP

[ZW06] Sining Zheng and Lidong Wang. Blow-up rate and profile for a degenerate parabolic system coupled via nonlocal sources. Computers and Mathematics with Applications, 52(10– 11):1387–1402, November/ December 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122106003270.

Zhu:2008:SRL

[ZW08] Wen-Tao Zhu and Chuan-Kun Wu. Security of the redefined Liaw’s broadcasting cryptosystem. Computers and Mathematics with Applications, 56(7):1665–1667, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108002253.

Zhang:2009:FOA

[ZW09a] Qingjie Zhang and Wenqia Wang. A four-order alternating segment Crank– Nicolson scheme for the dispersive equation. Computers and Mathematics with Applications, 57(2):283–289, January 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006305.

Zhang:2009:IRR

[ZW09b] Xiaolei Zhang and Renhong Wang. Isolating the real roots of the piecewise algebraic variety. Computers and Mathematics with Applications, 57(4):565–570, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210800669X.

Zhang:2009:ANI

[ZWC^{+}09] Xuncai Zhang, Yanfeng Wang, Guangzhao Cui, Ying Niu, and Jin Xu. Application of a novel IWO to the design of encoding sequences for DNA computing. Computers and Mathematics with Applications, 57(11–12):2001– 2008, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005191.

Zhang:2000:PSS

[ZWCW00] R. Y. Zhang, Z. C. Wang, Y. Chen, and J. Wu. Periodic solutions of a single species discrete population model with periodic harvest/stock. Computers and Mathematics with Applications, 39(1–2):77–90, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122199003156.

Zhang:2004:PSN

[ZWW04] Zhengqiu Zhang, Jun Wu, and Zhicheng Wang. Periodic solutions of nonautonomous stage-structured cooperative system. Computers and Mathematics with Applications, 47(4–5): 699–706, February/March 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122104900570.

Zhao:2008:OCS

[ZWY08] Aimin Zhao, Yanmei Wang, and Jurang Yan. Oscillation criteria for secondorder nonlinear differential equations with nonlinear damping. Computers and Mathematics with Applications, 56(2):542– 555, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108000837.

Zhan:2008:STG

[ZX08] Jianming Zhan and Yang Xu. Some types of generalized fuzzy filters of BLalgebras. Computers and Mathematics with Applications, 56(6):1604–1616, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001879.

Zhang:2009:OPF

[ZX09] Ying Zhang and Ying-Tao Xu. A one-parameter filled function method applied to nonsmooth constrained global optimization. Computers and Mathematics with Applications, 58(6):1230–1238, September 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122109004763.

Zheng:2007:CQG

[ZXL07] Xiaojin Zheng, Zengkun Xu, and Cheng Li. Constraint qualification in a general class of nondifferentiable mathematical programming problems. Computers and Mathematics with Applications, 53(1):21– 27, January 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S089812210700020X.

Zhang:2009:ESN

[ZXlD09] Ya-Nan Zhang, Fei Xu, and Ling ling Deng. Exact solution for nonlinear Schrödinger equation by He’s frequency formulation. Computers and Mathematics with Applications, 58(11–12):2449–2451, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122109001795.

Zhao:2009:RCC

[ZXTL09] Shuanfeng Zhao, Guanghua Xu, Tangfei Tao, and Lin Liang. Real-coded chaotic quantum-inspired genetic algorithm for training of fuzzy neural networks. Computers and Mathematics with Applications, 57(11–12):2009– 2015, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800518X.

Zhou:2005:UFK

[ZXX05] S. Zhou, H. Xu, and W. Xiao. A unified formulation of Kautz network and generalized hypercube. Computers and Mathematics with Applications, 49(9–10):1403–1411, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105001598.

Zhang:2002:NLS

[ZY02] Guochuan Zhang and Deshi Ye. A note on on-line scheduling with partial information. Computers and Mathematics with Applications, 44(3–4):539–543, August 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122102001682.

Zhang:2005:GAB

[ZY05] Xiaosheng Zhang and Jurang Yan. Global asymptotic behavior of nonlinear difference equations. Computers and Mathematics with Applications, 49(9– 10):1335–1345, May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122105001537.

Zhou:2007:EAR

[ZY07] Ling Zhou and Stephen Yau. Efficient association rule mining among both frequent and infrequent items. Computers and Mathematics with Applications, 54(6):737–749, September 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122107002519.

Zhang:2008:SIG

[ZY08a] Zhizheng Zhang and Hanqing Yang. Several identities for the generalized Apostol–Bernoulli polynomials. Computers and Mathematics with Applications, 56(12):2993–2999, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800480X.

Zhou:2008:OBH

[ZY08b] Xiaoliang Zhou and Rong Yu. Oscillatory behavior of higher order nonlinear neutral forced differential equations with oscillating coefficients. Computers and Mathematics with Applications, 56(6):1562–1568, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210800182X.

Zheng:2009:CCS

[ZYCI09] Bing Zheng, Lijuan Ye, and Dragana S. Cvetkovic-Ilic. The *congruence class of the solutions of some matrix equations. Computers and Mathematics with Applications, 57(4):540–549, February 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108006627.

Zhang:2007:NDE

[ZYEZ07] Yu Zhang, Xiaofan Yang, David J. Evans, and Ce Zhu. On the nonlinear difference equation system x_{n+1} = A + y_{n−m}/x_{n}, y_{n+1} = A + x_{n−m}/y_{n}. Computers and Mathematics with Applications, 53(10):1561– 1566, May 2007. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107001253.

Zhai:2008:PSO

[ZYG08] Cheng-Bo Zhai, Chen Yang, and Chun-Mei Guo. Positive solutions of operator equations on ordered Banach spaces and applications. Computers and Mathematics with Applications, 56(12):3150–3156, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108004811.

Zhang:2009:ACA

[ZYG^{+}09] Tiancheng Zhang, Dejun Yue, Yu Gu, Yi Wang, and Ge Yu. Adaptive correlation analysis in stream time series with sliding windows. Computers and Mathematics with Applications, 57(6):937–948, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005671.

Zhou:2001:ABD

[ZYH01] Zhan Zhou, Jianshe Yu, and Lihong Huang. Asymptotic behavior of delay difference systems. Computers and Mathematics with Applications, 42(3–5):283–290, August/September 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122101001535.

Zhang:2006:CFM

[ZYL06] Cheng Zhang, Xue-Hai Yuan, and E. Stanley Lee. Convex fuzzy mapping and operations of convex fuzzy mappings. Computers and Mathematics with Applications, 51(1):143–152, January 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105004505.

Zhang:2008:EPS

[ZYW08] Mingchuan Zhang, Yanmin Yin, and Zhongli Wei. Existence of positive solution for singular semipositone (k, n − k) conjugate m-point boundary value problem. Computers and Mathematics with Applications, 56(4):1146– 1154, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108001168.

Zhang:2008:IHS

[ZYY08] Xiaopan Zhang, Xiaohui Yuan, and Yanbin Yuan. Improved hybrid simulated annealing algorithm for navigation scheduling for the two dams of the three gorges project. Computers and Mathematics with Applications, 56(1):151– 159, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122107008346.

Zhao:2009:LAM

[ZYY09] Xiang Zhao, Yuan Yao, and Liping Yan. Learning algorithm for multimodal optimization. Computers and Mathematics with Applications, 57(11–12):2016– 2021, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108005178.

Zhang:2008:CAC

[ZYZ08] Haixian Zhang, Zhang Yi, and Lei Zhang. Continuous attractors of a class of recurrent neural networks. Computers and Mathematics with Applications, 56(12):3130–3137, December 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122108004380.

Zhang:2000:ONS

[ZZ00a] B. G. Zhang and Yong Zhou. Oscillation and nonoscillation for second-order linear difference equations. Computers and Mathematics with Applications, 39(1–2): 1–7, January 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122199003089.

Zhou:2000:ODD

[ZZ00b] Yong Zhou and B. G. Zhang. Oscillations of delay difference equations in a critical state. Computers and Mathematics with Applications, 39(7–8):71– 80, April 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122100000663.

Zhang:2001:OKT

[ZZ01a] B. G. Zhang and Yong Zhou.

Oscillation of a kind of twovariables functional equation. Computers and Mathematics with Applications, 42(3–5):369–378, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001614.

Zhang:2001:SSN

[ZZ01b] B. G. Zhang and Yong Zhou. Semicycles of solutions of nonlinear difference equations with several delays. Computers and Mathematics with Applications, 42(3–5):413–418, August/ September 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101001663.

Zhou:2001:ENT

[ZZ01c] Yong Zhou and B. G.

Zhang. An estimate of numbers of terms of semicycles of delay difference equations. Computers and Mathematics with Applications, 41(5–6):571–578, March 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S089812210000300X.

Zhang:2002:ODE

[ZZ02a] B. G. Zhang and Yong Zhou. Oscillations of difference equations with several delays. Computers and Mathematics with Applications, 44(5–6):817–821, September 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102001931.

Zilinskas:2002:GOB

[ZZ02b] A. Zilinskas and J. Zilinskas. Global optimization based on a statistical model and simplicial partitioning. Computers and Mathematics with Applications, 44(7):957–967, October 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0898122102002067.

Zhang:2003:NAT

[ZZ03a] Ju-Liang Zhang and Xiang-Sun Zhang. A nonmonotone adaptive trust region method and its convergence. Computers and Mathematics with Applications, 45(10–11):1469–1477, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001305.

Zhou:2003:ENS

[ZZ03b] Yong Zhou and B. G. Zhang.

Existence of nonoscillatory solutions of higherorder neutral delay difference equations with variable coefficients. Computers and Mathematics with Applications, 45(6–9):991–1000, March/May 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103000749.

Zheng:2005:GES

[ZZ05] W. Zheng and J. Zhang.

Global exponential stability of a class of neural networks with variable delays. Computers and Mathematics with Applications, 49(5–6):895–902, April/May 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122105000994.

Zilinskas:2006:PHA

[ZZ06] A. Zilinskas and J. Zilinskas. Parallel hybrid algorithm for global optimization of problems occurring in MDS-based visualization. Computers and Mathematics with Applications, 52(1–2):211–224, July 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122106001933.

Zhang:2009:PPB

[ZZ09a] Chungou Zhang and Zhihui Zhu. Preservation properties of the Baskakov– Kantorovich operators. Computers and Mathematics with Applications, 57(9): 1450–1455, May 2009. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109000649.

Zhao:2009:SLL

[ZZ09b] Lingxiao Zhao and Zhenyue Zhang. Supervised locally linear embedding with probability-based distance for classification. Computers and Mathematics with Applications, 57(6):919–926, March 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108005695.

Zhongxue:2009:STT

[ZZ09c] Lü Zhongxue and Liu Zuhan. Sharp thresholds of two-components Bose– Einstein condensates. Computers and Mathematics with Applications, 58(8): 1608–1614, October 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122109004490.

Zou:2008:GHA

[ZZD08] L. Zou, Z. Zong, and G. H.

Dong. Generalizing homotopy analysis method to solve Lotka–Volterra equation. Computers and Mathematics with Applications, 56(9):2289–2293, November 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122108003635.

Zhou:2003:LSS

[ZZH03] Fu-Zhao Zhou, Lei Zhang, and Xi-Yan Hu. Leastsquare solutions for inverse problems of centrosymmetric matrices. Computers and Mathematics with Applications, 45(10–11):1581–1589, May/June 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122103001378.

Zhang:2004:MNM

[ZZJ^{+}04] Qiang Zhang, Mengping Zhang, Guodong Jin, Dayou Liu, and Chi-Wang Shu. Modeling, numerical methods, and simulation for particle-fluid two-phase flow problems. Computers and Mathematics with Applications, 47(8–9):1437–1462, April/May 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122104901368.

Zhou:2002:MIA

[ZZZ02] Shuzi Zhou, Jinping Zeng, and Wuping Zhan. Monotonic iterative algorithms for an implicit two-sided obstacle problem. Computers and Mathematics with Applications, 43(1–2):31–40, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0898122101002681.

Zhao:2009:MPS

[ZZZ09] Yuxin Zhao, Wei Zu, and Haitao Zeng. A modified particle swarm optimization via particle visual modeling analysis. Computers and Mathematics with Applications, 57(11–12):2022– 2029, June 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:// REFERENCES 764

www.sciencedirect.com/
science/article/pii/S0898122108005166.